A Bibliography of Papers in Lecture Notes in Computer Science (2001) (Part 3 of 4)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

16 March 2017 Version 1.09

Title word cross-reference

#1 [1041].

(1 + ϵ) [1396]. (2^{m}) [1898]. (H, C, K) [901].

(p) [1898]. (t, n) [194]. (Z_{3}; +) [1010]. + [1614]. 0 − 1 [1281]. 0.863 [503]. 0/1 [269]. 1 [1484, 499, 729, 993]. 1.5 [1003]. 2 [499, 1853, 918, 1141, 993, 426, 1493]. 3 [456, 2504, 458, 906, 1456, 2557, 334, 703, 2618, 920, 811, 504, 2494, 808, 1289, 845, 355, 2547, 348, 835, 740, 1182]. 3/2 [499]. 3^{rd} [196]. 4 [1731, 920]. 6 [1731]. ^{3} [742]. ^{K} [2293]. ^{∘} [677]. b [508]. E [2275]. : = [1614]. d [1602]. ϵ [575]. ϵ_{0} [2592]. NP _{W} [899]. K [234, 1447, 583, 897, 262, 883, 1842, 581, 2105]. k > 2 [2528]. L [511]. λ [1130, 2589, 1817].

M [361, 16]. Z^{2} [1008]. A [2285]. N [1010, 1335]. ν [2084]. O(Δ log n) [1864]. O(k^{2}) [2524]. O[n^{3}/z^{3}] [431]. ω [1109]. π [1628]. P ≠ NP [983]. Q [2085]. R [1567, 1538]. ρμ [1321]. θ [1702]. × [1614]. ε [362]. X [2280, 1769]. y [1892]. Z [1567, 1553]. Z_{4} [261].

-Approximation [1396, 906, 499, 1853, 503]. -Band [361]. -Bit [1335]. -Borsuk [269]. -Calculus [1130, 2589, 1628]. -Coded [1321]. -Coloring [901]. -Commerce [16]. -Connected [262]. -Coordinate [1892]. -D [456, 334, 703, 811, 2494, 835]. -Dimensional [1602, 729, 1289]. -Edge-Coloring [1864]. -Game [1817]. -Hard [899]. -Homogeneous [918]. -in- [1731]. -Interval [2105]. -Learning [2085].

-Machines [1769]. -Means [1447, 583]. -Median [897]. -Minimal [2280]. -Nearest [234]. -Norm [1484]. -Partite [508]. -Regular [920]. -Renaming [2524]. -Representable [1109]. -RES [2275]. -SAT [883]. -Segments [581]. -Server [1842]. -Subsumption [1702]. -Support [575, 2084]. -system [2285]. -Trees [1538]. -Uniform [362]. -Way [1003].

/MOM [1069].

1 [872]. 1.4 [2629]. 1.82Gbits [1886]. 1.82Gbits/sec [1886]. 100 [2686]. 178B [2689].

2 [313, 701]. 2-Based [313]. 200 [1357]. 2000 [2490, 164]. 2001 [2655, 1564]. 2002 [2395]. 21 [1733]. 2i [2546].

360 [677].

4 [2215, 2195].

61508 [2692].

7 [305, 300, 302, 304, 301, 303].

802.11 [2179].

A-Prolog [2272]. Abduction [1694, 2346, 2285]. Abductive [2302]. Abelian [191, 1055]. Ability [620, 1798, 556]. able [2324]. Absence [1765]. Abstract [1244, 90, 1561, 96, 1339, 412, 408, 519, 2516, 1206, 1316, 2226, 411, 395, 1112, 1115]. Abstracting [415]. Abstraction [1577, 1637, 2564]. Abstraction-Based [1637]. Abstractions [273]. Accelerated [1338]. Accelerating [1303, 541]. Accelerators [1333]. Accepting [887]. Access [2238, 784, 1094, 2173, 2179, 1607, 1556, 783, 2018, 2180, 1545, 1951, 447, 778, 2219, 1715, 428, 2178, 2251]. Accesses [2606]. Accessible [437]. Accompanist [2340]. According [1366]. Accordion [1475]. Account [1779]. Accumulation [310]. Accuracy [1869, 1491, 2060, 1488, 301, 622]. Accurate [2009, 2254]. Achieving [696]. Acids [1399]. ACISP [164]. ACL2 [1584]. Acquired [2046]. Acquisition [68, 73, 2032, 689, 381, 348, 690]. Across [1394]. Act [1828]. Action [2621, 1994, 1656]. Actions [2356, 2275, 2322, 63]. Activating [2202]. Activation [2144, 2318]. Active [600, 2239, 332, 2202, 579, 1554, 1345, 333, 835, 2112, 2158]. Activity [2371, 1782, 2626, 2624, 91, 207, 607, 2625]. Acts [1655]. Actual [1108]. Acyclic [504]. Ad [2535, 1096, 510, 1730]. Ad-Hoc [1730, 2535, 1096, 510]. AdaBoost.M2 [2067]. Adaptability [1768, 1477]. Adaptable [2198, 665, 1090, 1970]. Adaptation [1744, 2338, 2662, 1727]. Adapted [2364, 2361, 2034]. Adapting [1374]. Adaption [1794]. Adaptive [657, 646, 2397, 1745, 1793, 1781, 1964, 13, 2196, 2213, 1596, 1723, 1293, 542, 627, 297, 759, 159, 2500, 2524, 2033, 1666, 1697, 2516, 443, 761, 579, 724, 315, 1344, 329, 2197, 366, 643, 520, 1730, 1330, 2045, 664]. Adaptive-Rate [2196]. Adding [1743, 2538, 712, 2459]. Addition [1885]. Addition-Subtraction [1885]. Address [769, 865, 2377]. Addressing [2317]. ADEPT [1947]. Adequately [1195]. ADI [2513, 730]. Adjacency [129]. adjoint [2282]. Adjusting [1508]. Adjustment [652, 2040]. Admissible [2103]. Admission [1669, 2185, 2199, 2182, 1671, 1670, 1668]. ADOME [157]. Adults [68]. Advanced [1308, 2575, 371, 1349, 1309]. Advances [1984]. Advection [2469]. ADVISOR [2388, 2295]. AES [1906, 1886]. Aesthetic [1813]. Affair [94]. Affective [1803]. Affine [322]. Afford [93]. Affordable [1687]. AFPs [987]. Afterward [802]. Against [1272, 2653, 1906, 177, 1911, 1885, 188, 1079, 2148, 960]. Age [118]. Agent [1760, 1961, 2373, 2325, 1524, 20, 2257, 1971, 2569, 2229, 2584, 1980, 48, 2563, 626, 1967, 766, 754, 2613, 1968, 198, 763, 758, 1970, 761, 762, 707, 2575, 1977, 2650, 2359, 2230, 2298, 2581, 2231, 2560, 1966, 755, 59, 2045, 1979, 12, 85, 1962, 1982, 2323, 2357, 2227, 2228, 2567, 624, 1748, 1969, 2578, 2574, 639, 625]. Agent-Based [1961, 1524, 1971, 1968, 2575, 1977, 1966, 2045, 1967]. Agent-Mediated [762]. Agent-Oriented [2325, 2650]. Agents [1974, 2583, 1978, 1743, 2579, 1769, 2565, 2573, 1972, 2577, 2354, 2585, 2568, 1963, 1969, 2580, 29, 2528, 759, 1976, 2557, 1981, 1958, 2556, 1798, 1959, 1786, 756, 2566, 2090, 2558, 2091, 651, 1973, 2582, 118, 2562, 1009, 1960, 1975, 2353, 1957, 1965, 2232, 1730, 2044, 1790, 767, 753]. Agglomerative [608]. Aggregate [1992, 1553]. Aggregated [2189, 291]. Aggregates [2124, 2136]. Aggregation [1679, 1004, 2200, 1700, 2468, 253]. Aggregation-Based [1679]. Aggregation/Disaggregation [2468]. Aggregations [140]. Agreement [2525, 2003, 505]. Agricultural [2428]. Aguri [1679]. AGVs [455]. AI [2297, 2329]. Aid [982]. Aided [2444, 359, 2417]. Air [2487, 2497, 2488, 2492, 2493, 2494, 2469, 2177]. Airport [2540]. ajka [2010]. Alarms [1975]. Alexandria [1947]. Alfred [1101]. Algal [1752]. Algebra [1566, 1619, 2305, 1996, 1993, 1034, 477, 1614, 971, 1612, 450]. Algebraic [1625, 484, 876, 1856, 321, 2420, 2478, 754, 2024, 987, 1257, 170, 1110, 1622, 266]. Algebras [1127, 1995]. Algorithm [1277, 432, 1166, 2473, 312, 495, 976, 2182, 1691, 732, 341, 496, 1484, 1908, 1905, 1864, 2065, 357, 2475, 1862, 1863, 2624, 906, 2334, 499, 2531, 320, 385, 1004, 1854, 1446, 2127, 1085, 1541, 1548, 1853, 1633, 336, 1886, 2082, 502, 424, 280, 503, 1887, 1308, 1855, 941, 1807, 315, 445, 539, 1910, 1213, 1892, 467, 2142, 813, 1079, 371, 808, 1788, 1289, 356, 1867, 1365, 622, 581, 688, 957, 379, 804, 1150, 1631]. Algorithmic [1837, 896, 877, 2077, 2109, 2155, 99, 1600]. Algorithms [1095, 2280, 1094, 1097, 833, 1449, 1431, 827, 1017, 893, 1245, 1592, 2119, 384, 877, 1739, 2532, 2120, 905, 2122, 1598, 1468, 1023, 1872, 500, 1839, 2124, 1092, 2476, 1398, 881, 968, 2517, 1544, 2137, 463, 1850, 1871, 1847, 1673, 1412, 465, 1855, 466, 2480, 924, 319, 1586, 1009, 266, 353, 1787, 883, 1846, 1868, 998, 1677, 1012, 268, 556, 2687, 972, 1386, 376, 835, 2112, 2158, 1924, 1879, 2469, 1644, 1163, 1443, 1483, 793, 792]. Alife [1816, 1782]. Alignment [1379, 934]. Alignments [1380]. All-Optical [1967]. Allocation [1091, 228, 1971, 2191, 2425, 1009, 1776, 1512]. Almost [1353]. along [1383]. Alphabets [923]. Alter [1100]. Alternating [756]. Alternating-Offers [756]. Alternation [926]. Alternative [1960]. Alternatives [41]. Amalgamability [914]. Ambainis [976]. Ambainis-Freivalds [976]. Ambients [1616, 1123, 412, 411]. Ambiguity [364, 47, 935]. AMBR [50]. Amino [1399]. Amoeboid [1801]. among [2573, 2091, 2232]. Amorphous [1811]. Analog [1673]. Analogy [2610]. Analyse [343]. Analyser [2010]. Analysers [592]. Analyses [402, 404, 396]. Analysing [411]. Analysis [2419, 800, 950, 954, 398, 1210, 1634, 889, 1097, 1220, 420, 965, 1424, 1066, 2440, 2691, 225, 1986, 1359, 403, 2633, 2429, 400, 1865, 2308, 2067, 412, 216, 731, 1901, 213, 2345, 393, 1053, 2596, 414, 2635, 214, 364, 102, 1911, 1604, 1450, 358, 2030, 2479, 335, 370, 635, 547, 397, 465, 425, 568, 394, 2507, 395, 727, 2089, 2683, 347, 329, 601, 1553, 2436, 1078, 1227, 389, 2443, 373, 1441, 2053, 1493, 1423, 2320, 700, 281, 604, 544, 2004, 572, 413, 376, 656, 367, 521, 2625, 674, 799, 2406, 107]. Analytic [1394]. Analytical [2196, 23, 1490]. Analyzed [502]. Analyzing [407, 628, 2387, 2682, 30, 1948]. Anaphora [65, 58]. Anaphoric [52]. Anatomy [1113, 1403]. Anchoring [2366]. Animal [1776]. Animals [104]. Animated [1651]. Animation [1823, 1824]. Animations [359]. Annealing [1167, 824, 488]. Annotated [2361, 683]. Annotation [2615, 2567, 2015]. Announcement [2395, 2394]. Anonymity [954, 953]. Anonymous [2518, 2527, 197, 27, 964]. Answer [2288, 2295, 2287, 2307, 2278, 2273]. Answer-Set [2287, 2307]. Answering [2360]. Ant [1739, 1773, 1807]. Anthropological [2603]. Anti [267]. Anti-symmetric [267]. Antibodies [1754]. Anticipation [642]. Anticipation-Based [642]. Antigens [1754]. Antithetic [1158]. Antitonic [2284]. Any [191]. Anytime [1069, 1191, 1150]. Anywhere [1069]. AOP [2241]. AP [742]. Aperiodic [1510]. API [1075]. Appendix [2168, 1274, 1273, 1833]. Appliances [871, 614]. Applicable [2682]. Application [2343, 2590, 975, 228, 2428, 2116, 550, 460, 1575, 1180, 73, 819, 1723, 2327, 871, 2671, 294, 359, 1487, 1073, 2316, 2030, 1701, 207, 757, 1219, 296, 1216, 1705, 961, 739, 1706, 19, 1516, 366, 2432, 452, 355, 281, 807, 594, 835, 587, 182]. Applications [1667, 1927, 2056, 2363, 1217, 305, 3, 1424, 1964, 1428, 1661, 2213, 302, 1067, 737, 2628, 304, 1734, 715, 438, 1480, 1534, 2243, 1508, 968, 2491, 716, 697, 2680, 1198, 2377, 399, 1728, 2195, 2176, 717, 2022, 707, 1453, 331, 236, 301, 1713, 2, 699, 743, 273, 449, 544, 2045, 1101, 1790, 540, 186, 11, 2251, 431, 1069]. Applicative [2458]. Applied [2065, 2512, 715, 2328, 16, 295, 2688, 2447, 601, 1241]. Applying [1974, 611, 2401, 2084, 394, 614]. Appointed [1050]. Approach [973, 1760, 2058, 569, 2437, 418, 1400, 151, 2365, 863, 833, 1652, 65, 1482, 2246, 2634, 2431, 13, 2693, 2337, 461, 2430, 2119, 2632, 1519, 2505, 2257, 1069, 45, 532, 73, 1971, 2070, 2464, 114, 2127, 805, 1354, 48, 2499, 2416, 224, 754, 210, 254, 2083, 842, 185, 2682, 1603, 2607, 1355, 2506, 834, 1817, 2493, 54, 529, 1585, 2359, 2472, 22, 478, 728, 2608, 2, 329, 1819, 673, 1657, 564, 2181, 2616, 2333, 1332, 1362, 363, 2273, 2388, 281, 604, 2187, 454, 829, 344, 948, 664, 150, 1802, 2339, 156, 2660]. Approaches [611, 2019, 1594, 1922, 2109, 2155, 266, 1682, 286, 1182]. Appropriate [44, 387]. Approximability [1848, 933, 513]. Approximate [1590, 1875, 1282, 1085, 1849, 1118, 1838]. Approximating [1851, 1862, 979, 487, 970, 2300, 679]. Approximation [973, 38, 1431, 497, 2120, 1863, 1396, 905, 906, 499, 1023, 1872, 500, 537, 968, 1853, 558, 1847, 552, 2506, 503, 993, 505, 150]. Approximations [321, 561, 250, 1174]. AQUILA [1680, 2212]. Arc [1556]. Arc-Tree [1556]. Architectural [2369, 914, 1886]. Architecture [2623, 1313, 2373, 1458, 1094, 1676, 1925, 237, 2248, 2371, 1359, 523, 249, 1500, 714, 1720, 1969, 1898, 642, 1488, 1751, 1940, 2212, 849, 1265, 424, 236, 1909, 718, 19, 1334, 1715, 1454, 1496, 640, 1406, 2045, 1790, 1549, 2659]. Architectures [2192, 2363, 1352, 2483, 2225, 623, 1461, 1295, 2405, 2635, 1447, 1462, 1298, 2687, 1349, 2668]. Archive [1949, 1954]. Archives [588, 1948]. Archiving [282]. ArcIMS [249]. Area [1522, 384, 711, 844, 1715, 2251]. Area-Based [384]. Argument [2573, 755]. Argument-Based [2573, 755]. Argumentation [1201, 1202]. Arguments [2358]. Arising [1097, 1600]. Arithmetic [1576, 1882, 1108, 1025, 1896, 1363, 1358]. Arm [652, 653]. ARMiner [134]. Arms [2576]. Army [769]. Arrakeen [2620].

Arrangement [89]. Array [1352, 1605, 469, 1399]. Arrays [422, 1318, 1307, 1298, 1441]. Arrhythmia [1706]. Arrhythmias [2147]. ARSOM [661, 2370]. Art [2203, 2542, 2582, 1368, 542]. ART-Type [542]. ARTCP [432]. Artefact [112]. Articles [1934]. Articulated [813]. Artifact [111, 82]. Artificial [2437, 1794, 1813, 2077, 1783, 1751, 1762, 1816, 1786, 1757, 1812, 1759, 1742, 2438]. ARTMAP [550]. Ascent [496]. Asian [1869]. ASL [914]. ASM [2192]. ASMs [2413]. ASP [2277]. Aspect [2664, 2662, 2401, 2663, 2399]. Aspect-Oriented [2662]. AspectJ [2661]. Aspecto [205]. Aspecto-Temporal [205]. Aspects [2573, 2493, 112]. aspps [2287]. ASR [2028, 2023]. Assemblies [1401]. Assembly [1381]. Assertion [2419]. Assertion-Based [2419]. Assessing [1385, 1650]. Assessment [2676, 707, 298]. Assessments [2674]. Assets [2660]. Assignment [508, 2186, 539]. Assignments [1093]. Assist [1981]. Assistant [2415]. Assisted [1143, 2451]. Association [2118, 963, 2126, 2133, 2148, 2320, 2152, 2004, 134, 2125]. Associations [2644, 1701, 2646]. Associative [585, 635, 445, 1137]. Assumption [1203, 1043, 2005]. Assumption-Based [1203]. Assumptions [18]. Assurance [413]. Assured [2185]. Asymmetric [508, 1592, 1165, 18, 1041, 578]. Asymmetries [90]. Asymptotic [2417]. Asynchronous [1058, 1360, 1243, 1452, 1626]. Atmospheric [2491, 2496]. Atomix [2313]. ATP [2458]. ATRACT [2013]. ATT [35]. ATT-Meta [35]. Attachment [2362]. Attack [187, 1041, 960, 1904]. Attacks [1906, 954, 1899, 163, 550, 1913, 1885, 188, 166, 1079, 1900]. Attempts [330]. Attention [375, 680]. Attraction [380].

Attractors [636]. Attribute [154].

Attributes [155, 197]. Auction [797, 180]. Auctions [2270, 2571, 2569, 500, 1870, 759, 757, 2570]. Audio [1667, 962, 941]. Audit [30]. Auditory [671, 680, 2014]. Audits [2695]. Augmentation [1020]. Augmented [346, 90, 349, 2014]. Augmenting [499, 1653]. AURORA [2493]. Authenticated [199]. Authenticating [799]. Authentication [772, 800, 791, 790, 17, 181, 1046, 944, 2218, 183, 182]. Authenticity [2432]. Author [1983, 2462, 33, 79, 123, 161, 202, 229, 257, 270, 299, 388, 417, 471, 486, 518, 693, 751, 768, 803, 846, 874, 936, 966, 1027, 1061, 1083, 1099, 1142, 1214, 1250, 1292, 1367, 1378, 1402, 1533, 1562, 1589, 1608, 1647, 1663, 1688, 1710, 1737, 1822, 1880, 1914, 1956, 1997, 2057, 2113, 2159, 2216, 2223, 2256, 2296, 2332, 2374, 2396, 2411, 2515, 2539, 2555, 2586, 2587, 2619, 2656, 2672, 2696, 1939]. Authorization [1068]. Authorizations [185]. Autism [85]. AutoCAD [2546]. Autocalibration [385]. Automata [2414, 887, 1618, 481, 457, 976, 1220, 407, 895, 1001, 896, 1125, 909, 1579, 911, 912, 913, 244, 990, 1774, 1137, 1765, 1007, 1587, 997, 1013, 932, 1249, 423]. Automata-Theoretic [997]. Automate [2131, 287, 2]. Automated [1566, 2233, 1067, 1315, 2342, 1696, 2053, 348, 1651, 349]. Automatic [2115, 2056, 1925, 892, 1236, 284, 2021, 2013, 2002, 2340, 2145, 2003, 1846, 955, 21, 2050]. Automatically [2046, 1948]. Automation [1824, 2675]. Automaton [621, 1002, 2447]. Autonomous [2368, 2565, 1001, 2354, 2569, 650, 651, 2353]. Autoplex [2233]. Autopoiesis [1818]. Autopoietic [1786]. Availability [190, 1509, 1962]. Available [1069]. Average [978, 1022, 1854, 501, 1672, 2533, 2150, 1380]. Averaged [838]. Aversion [1764]. AVI [1315]. Avoidance [1957, 645]. Avoiding [925]. Aware [1974, 1727, 236, 1716, 2562, 2651]. Axiomatic [2058, 967, 184]. Axioms [888].

B [1274]. B2B [22]. Back [782, 781]. Background [2489, 2161]. Backjumping [2345]. Backpropagation [2087]. Backscattering [2475]. Bad [521]. Bagged [534]. Bagging [526]. Baire [894]. Balanced [1758]. Balancing [1425, 1429, 1430, 1479, 2492, 443, 718]. Baldwinian [1779]. Band [361]. Bandera [1613]. Bandwidth [2208, 2186, 512, 2191, 702, 1465, 1900]. Bandwidth-Optimal [1900]. Bank [2037, 1338]. Barbed [1628]. Bargaining [756]. Barriers [1481]. Base [2428, 2390, 2666, 1832, 2393]. Based [1296, 2419, 841, 1978, 1961, 1166, 2654, 800, 787, 1203, 2059, 125, 584, 398, 1523, 457, 151, 100, 35, 2453, 2056, 2303, 1992, 2182, 2428, 1207, 65, 2017, 1040, 199, 1311, 2634, 2573, 1918, 1424, 2431, 1524, 1346, 2337, 2371, 1430, 157, 1679, 437, 2610, 1353, 1343, 384, 2210, 1325, 9, 2391, 222, 20, 1068, 2392, 2513, 2611, 313, 1734, 2640, 1472, 412, 819, 2327, 1971, 1558, 765, 2012, 2433, 1720, 2457, 1560, 297, 1637, 351, 1147, 2229, 383, 337, 642, 2128, 1377, 1719, 159, 2133, 2423, 2077, 293, 276, 2369, 1942, 2314, 2416, 275, 2517]. Based [619, 1941, 1541, 1544, 1300, 386, 1940, 684, 1797, 224, 336, 2543, 1243, 1191, 2613, 597, 1645, 1736, 2386, 316, 1700, 825, 2638, 638, 1968, 1871, 51, 2643, 2387, 1959, 138, 185, 2184, 287, 401, 2180, 464, 2028, 2607, 2085, 2086, 52, 1944, 1233, 1583, 170, 1287, 2575, 2055, 2240, 2007, 614, 1977, 1939, 2359, 1356, 236, 539, 1818, 1910, 1357, 2340, 2291, 2142, 479, 2608, 536, 166, 347, 374, 274, 2380, 739, 329, 1417, 298, 2041, 307, 2026, 1389, 1546, 2016, 2436, 30, 2092, 4, 78, 47, 1542, 2181, 373, 528, 2432, 677, 273, 409, 2318]. Based [883, 1504, 2053, 1920, 1520, 2254, 938, 2232, 1218, 314, 1337, 1677, 2408, 2175, 1363, 640, 1364, 1966, 1674, 644, 281, 196, 570, 2187, 755, 1755, 285, 2250, 470, 565, 2151, 622, 1223, 2045, 1790, 1477, 960, 664, 286, 653, 144, 12, 239, 1547, 1327, 1082, 179, 674, 455, 156, 767, 799, 2106, 1982, 134, 129, 625, 2668, 2406, 1399, 1946, 1967, 763, 1671]. Bases [2412, 1184, 2310, 1283, 2385, 298]. Basic [1096, 2682, 1230]. Basis [649, 205, 2579, 967, 580, 158, 288, 184, 546, 536, 2144, 733, 543, 606, 545, 851, 2661, 1650]. Basket [1869]. Baukasten^{≪} [475, 473, 474]. Bayes [593, 1198, 2088, 2105, 2107]. Bayesian [2036, 2676, 537, 1697, 1698, 2084, 1157, 291, 601, 2025, 2301, 1158, 576, 556]. BDDs [1220, 893, 1624]. BDI [2355]. Be [85]. Bed [2209]. Beds [457]. Behavior [1745, 1995, 1781, 1461, 75, 48, 1762, 1916, 2574, 479, 860, 2318, 2648, 2687]. Behavior-Based [2318]. Behavioral [534, 2640, 1306, 410]. Behaviour [1793, 1806, 672, 2129, 650, 1776]. Behavioural [2061, 1794, 1800]. Belief [1162, 1171, 67, 1187, 1173, 2350, 2676, 2312, 1169, 1170, 1172, 1174, 1149, 1163]. Beliefs [1192, 627]. Bellman [445]. Bellman-Ford [445]. Benchmark [2426]. Benchmarking [961]. Benchmarks [702]. Beowulf [722, 540]. Best [978, 1668]. Bethe [804]. Better [1675, 1451, 545]. Between [1605, 2623, 2579, 493, 1803, 2265, 1192, 876, 2036, 34, 1865, 1882, 1840, 575, 1701, 1922, 89, 77, 1658, 934, 1392, 1182, 128]. Beyond [1450, 1987, 1371, 1137, 99]. Bézier [1314]. Bi [1482]. Bi-knapsack [1482]. Bias [1805, 538]. Bias-Variance [538]. Biased [1607, 2102]. Bibliographic [1939]. Bicartesian [2593]. BiCMOS [1299]. Bid [757]. Bidding [2569]. Bidirectionality [926]. Big [1904]. Bigraphical [1610]. Biharmonic [733]. Bilateral [1097, 19]. Bilateral-Contract [1097]. Bilinear [1007]. Bilingualism [101]. Billing [787]. Bin [1841]. Binary [534, 685, 2153, 269]. Binary-Encoded [2153]. Binaural [677]. Binding [1400, 1411, 1399]. Binocular [380]. Bio [1756]. Bio-Language [1756]. Bioinformatics [1696]. Biological [1743, 2437, 573, 2132, 465]. Biologically [2436]. Biomass [1752]. Biomedical [331]. Biometric [1072]. Biometrics [1080]. Biomorphic [2499]. Biosequence [1454]. Bipedal [1795]. Bisection [488, 920]. Bisimulation [809]. Bit [1335, 1898, 1005, 2038, 314]. Bit-Serial [1898]. Bitprobe [1859]. Bits [1039, 1005]. Bivariate [891]. Black [777, 2527, 1031]. Black-Box [777, 1031]. Blight [367]. Blind [590, 1125, 597, 591, 589]. Block [1484, 168, 1230, 1681]. blocking [2536, 2522]. Blocks [2457]. Blood [332]. Bloomy [2149]. Blueberry [2547]. Blur [960]. Blur/Deblur [960]. Board [1373]. Board-Laying [1373]. Body [733, 324]. Book [2601]. Bookmark [2561, 766]. Boolean [513, 2345, 1882, 1797, 210, 263, 516, 211]. Boom [352]. Boosting [2341, 2069, 2334, 525, 1693, 1382, 2096, 2101]. Booting [173]. Border [666]. Border-Ownership [666]. Born [81]. Borsuk [269]. Bottleneck [1465]. Bound [1856, 1092, 2525, 2516]. Boundaries [2036]. Boundary [562, 456, 2504, 663, 2471]. Bounded [506, 1036, 1472, 1390, 2271, 1623, 970, 913, 1395, 1218, 1867, 935]. Bounding [264]. Bounds [888, 893, 900, 904, 2078, 262, 558, 920, 517, 1868, 556]. Box [777, 1031]. BP [619]. Braid [1056]. Brain [1760, 519, 2144, 631]. Branch [1460, 1092, 1462, 1459]. Branching [974, 2305, 2069]. Break [964]. Breaking [2400]. Breakpoints [1341]. Breast [596]. Brick [1021]. Bridging [1179, 2064, 52]. Bright [777]. Broadcast [1058, 1872, 1548, 427, 1493]. Broadcasted [255]. Broadcasting [1873]. Broadening [789, 788]. Broker [2186]. Bronchial [330]. Brownian [2466]. Browser [1917]. Browsing [146, 2609, 225, 2141, 2614]. BSP [1418, 1481]. Bubble [1327]. Büchi [1124]. Büchi-Elgot-Trakhtenbrot [1124]. Buckets [1451]. Budget [500]. Budget-Constrained [500]. Buffer [240]. Buffering [1538, 1728]. Bugs [1222]. Build [1658]. Building [1660, 1472, 1521, 2457, 695, 105, 2542, 1949, 1360, 2018, 2092, 215, 223]. Buildings [2497]. Built [2548, 219]. Bundles [757]. Burst [670]. Bus [1241, 1962]. Business [2628, 1, 11, 2253]. Business-to-Business [2253]. Buyers [2572]. Buying [1843]. Bypass [1780]. Byte [1305]. Bytecode [1074].

C [400, 439, 709, 914, 1944, 2408]. C-Merge [1944]. Cache [1086, 1544, 1439, 687, 1231]. Caching [855, 1839, 24, 1312, 235]. Cactus [1517, 1516]. CAL [2451]. Calculational [1569]. Calculations [733, 470]. Calculi [1127, 2306, 2221]. Calculus [1130, 2589, 1573, 2367, 2595, 1578, 1104, 1580, 2449, 2314, 410, 1628, 1221, 1202, 1615]. Calendars [244]. Calibration [811]. Call [1668]. CAM [631]. CAM-Brain [631]. Camera [2618, 1830, 811]. Can [1992, 1793, 526, 93, 1755, 85]. Cancellation [600]. Candidates [965]. Cannot [801]. Canonical [568, 572]. Can’t [1780]. Capabilities [2368]. Capability [1734]. Capacity [1177, 959]. Capture [1106]. Capturing [1543]. Card [1065, 1066, 1067, 1068, 1070, 1074, 1077, 764, 1079]. Cardiac [335, 1706]. Cardinal [232, 245]. Cards [1064, 1069, 1076, 1078, 1082]. CardS4 [1071]. Carlo [2473, 2475, 2334, 2466, 2467, 2477, 2479, 538, 466, 539, 996, 533, 2333]. Carried [1436]. Carriers [2479]. Carrying [177]. Cart [2445]. Cartesian [882, 88]. CAs [1761]. Case [2059, 1497, 2295, 2349, 1086, 1947, 2407, 2000, 978, 1576, 2124, 2128, 227, 1937, 76, 1732, 1361, 2574, 2653, 924, 2678, 274, 996, 743, 273, 2382, 1950, 1423, 738, 11]. Case-Based [2059, 2128, 274, 273]. Cases [2654, 901, 986, 78]. Cases-Experiences [2654]. CAST [2437, 2436]. CAST-mechanism [2436]. CAST.FSM [2414]. Castes [752]. Catadioptric [387]. Catalog [23, 21]. Categorical [1120, 1136]. Categorically [429]. Categories [1573, 2593, 2612, 2434, 318]. Categorization [2342, 542, 2135, 2141, 2096, 57, 63]. Categorizing [32]. Category [894, 49, 994, 1818, 875]. Causal [1162, 1166, 1522, 1160, 1161, 527, 2347]. Causality [1204]. Cautious [2052]. Caveats [1160]. CBC [1045, 163]. CBMs [631]. ccNUMA [2483]. CCS [1629]. CDMA [2177]. Cell [126, 1801]. Cells [667, 1759]. Cellular [457, 896, 423, 913, 424, 1765, 1007, 426, 2175]. Cellular-Pipelined [424]. Centralized [2239]. Centred [2389, 54]. Centroid [841]. Centroid-Based [841]. Cephalometric [349]. Ceramics [2499]. Certain [201]. Certificates [784, 1065, 783, 194]. Certification [798]. Certified [1583]. Certifying [1574]. Cervantes [1921]. CG [227]. CGs [222, 206]. Chain [1862, 1388, 731, 1004, 539, 533, 679]. Chain-Positioning [1388]. Chains [2532, 1614, 1987, 2468, 1885, 88]. Challenge [2111, 2157]. Challenged [1326]. Challenges [2192, 696, 969, 2022, 1600]. Change [2116, 1187, 982, 136, 2398]. Changeable [193]. Changepoint [2036]. Changes [1161, 524]. Changing [1429]. Channel [1609, 1913, 962, 2680, 1527, 601, 1177, 609]. Channels [940, 1775]. Chaos [675, 406]. Chaotic [636, 617]. Chapter [2168, 2161, 2160, 2163, 2166, 2164, 2162, 2167, 2165, 1272, 1257, 1270, 1271, 1264, 1253, 1254, 1265, 1269, 1252, 1258, 1256, 1259, 1260, 1266, 1251, 1255, 1261, 1262, 1263, 1267, 1268]. Character [267, 1950]. Characterisations [1111]. Characterising [2352]. Characteristics [8, 2468, 2235, 740]. Characterization [2615, 1210, 2657, 2518, 1625, 1706, 935]. Characterizations [2281, 2598, 381]. Characterizing [2431, 136, 429]. Characters [289, 106, 688]. Charge [2479]. CHARME [1564]. Check [1613, 267, 9]. Checker [685]. Checking [1217, 1220, 1990, 1123, 1638, 400, 2247, 1239, 408, 1637, 641, 2271, 1623, 1614, 1985, 2259, 914, 399, 1987, 1231, 1219, 2677, 1227, 1554, 1218, 1222]. Checkpointing [1414, 864]. Chemical [2349, 951, 1389, 2496]. Chemical-Distance-Based [1389]. Chemistry [1757, 1759]. Chernoff [517]. Children [68]. China [1952]. Chinese [1825, 958, 2039]. Chip [1311, 1243, 1887, 1308, 1366, 1358]. Chip-Based [1311]. Choice [1147, 1183, 1777]. Chosen [1041]. Chronicle [1706]. Cinematic [350]. Cipher [168, 1907]. Ciphers [1045]. Ciphertext [1041]. Circuit [890, 615]. Circuits [1343, 1297, 1360, 514, 882, 1327]. Circular [988]. Circumscriptions [1189]. Citation [145]. Cladistics [2119]. Claim [1532]. Class [1212, 2666, 169, 1895, 992, 2145, 574, 2102, 1249, 1150, 2183]. Classes [326, 475, 2450, 995, 478]. Classical [1127, 1190, 2435, 673]. Classification [8, 2365, 596, 623, 2222, 2070, 1931, 2123, 152, 2075, 386, 766, 288, 602, 1444, 563, 2139, 296, 614, 2140, 274, 319, 2093, 826, 1920, 2418, 2149, 570, 830, 583, 327, 2100, 2102, 135, 587, 2339, 2689]. Classifications [326, 1944]. Classifier [2060]. Classifiers [655, 2073, 291, 565, 2105, 2106]. Classify [1782, 131]. Classifying [1692, 2450, 167]. clausal [1212]. Clause [2097]. Cleanness [400]. Clearing [1097]. Client [791, 790, 2195, 707]. Client-Agent-Server [707]. Client-Server [2195]. Climbing [2519, 488]. Clinical [1933, 2673, 272, 2151]. Cliques [831]. Clock [1225]. Clocked [1243]. Clocks [1475, 1471]. Clones [2061]. Close [574]. Close-Class-Set [574]. Closest [1538]. Closings [360]. Closure [995, 728]. Cluster [723, 1422, 1502, 1523, 2453, 458, 1424, 437, 1432, 725, 701, 440, 1510, 1513, 736, 703, 1361, 546, 744, 1511, 1713, 733, 749, 1504, 1507, 1410, 540]. Cluster-Based [2453]. Cluster-to-Cluster [1713]. Clustering [1448, 279, 549, 588, 534, 2068, 2121, 2122, 1598, 819, 1447, 2127, 1942, 145, 1541, 530, 2079, 1501, 318, 280, 1799, 1807, 2013, 2007, 1951, 2142, 468, 2092, 2146, 533, 608, 1730, 285, 583, 144, 133, 587]. Clusters [1506, 435, 151, 317, 1503, 1398, 1508, 1514, 698, 718, 1417, 699, 1712, 470, 722]. Cluttered [660]. CNOP [1591]. Co [74, 2139, 1341, 242]. Co-debug [1341]. Co-evolution [2139]. Co-location [242]. Co-supposed [74]. Coalesce [1437]. Coalition [760]. Coarsening [1174]. Cocolourings [979]. CODD [2605]. Code [1457, 1436, 1435, 391, 703, 958, 742, 867]. Codebooks [312]. Codecs [307]. Coded [673, 1950, 1321]. Coder [311]. Codes [498, 261, 2135, 2049, 490, 1305, 269, 177]. Coding [313, 310, 666, 1498, 2038, 314, 308, 1363, 668]. Coefficient [314, 1349]. Coefficients [2025]. Coevolutionary [1771]. Cognition [101, 102]. Cognitive [205, 84, 110, 86, 624, 103, 80, 114, 111, 122, 112, 676, 54, 57, 625]. Coherence [419, 1168, 1191, 1231]. Coherence-Based [1191]. Coherent [1171, 1153, 1155, 2593]. Cohesion [2017]. Coin [1988, 1037, 274, 949, 451]. Coin-Tossing [1037]. COiN-Video [949]. Coinductive [1141]. Collaborate [2091]. Collaboration [2354, 2383]. Collaborations [2639]. Collaborative [1769, 765, 2580, 29, 1655, 2561, 764, 1725, 2380, 2241, 2331, 2153, 639]. Collapse [1109]. Collecting [735]. Collection [1848, 2018, 1241]. Collections [1932, 1917]. Collector [857, 858]. Collision [1497]. Collisions [1738]. Collusion [958]. Collusion-Secure [958]. Colonies [1773, 1774]. Color [1858, 818, 584, 657, 302, 310, 824, 365, 1831, 331, 375, 372, 363, 306, 587]. Color-Spanning [1858]. Coloring [1864, 901, 1019, 1841, 517, 269]. Colour [382]. Colouring [1024]. Combination [1170, 347, 2101]. Combinatorial [2270, 1238, 877, 1004, 266]. Combinatory [1127]. Combine [1519]. Combined [114, 365, 1917]. Combining [1698, 51, 2109, 2155, 1025, 1703, 78, 583]. Combustion [2497]. Comfortableness [89]. Commands [748]. Commerce [2343, 17, 16, 12, 1549, 14, 162, 11, 24, 2575, 22, 4]. Commercial [2575]. Commit [2522]. Commitments [1029]. Committee [576]. Committees [2092]. Common [1282]. CommonKADS [2424]. Communicating [889, 2044]. Communication [973, 1654, 1422, 436, 1489, 390, 1648, 1675, 421, 2358, 2535, 1096, 74, 2584, 2662, 1655, 1513, 2557, 962, 2618, 2027, 2695, 1657, 1960, 1775, 702, 1493, 451, 1427, 1665, 1778, 1809, 1653, 108, 755]. Communication-Computation [436]. Communications [706, 2465, 701, 1046, 1481, 1366]. Communities [1521, 2378, 221]. Community [726, 765, 119]. Commutation [969, 986]. Commutative [1137, 1105]. commutativity [2489, 1104]. Compact [1605, 1084, 1241, 355]. Companies [2377]. Company [2384]. Comparative [1179, 384, 700, 2266]. Comparing [2591, 1401, 1384, 141, 2088, 1193, 2321]. Comparison [458, 40, 827, 2511, 605, 42, 655, 861, 295, 1505, 2146, 469, 2687, 1182, 1924]. Compatibility [1193]. Compatible [1390]. Compensation [2025, 354]. Competition [75, 634]. Competitive [1843, 2519, 1839, 1870, 629]. Compilation [1304, 1346, 1439, 1233, 453]. Compiler [2482, 1438, 2289, 1364]. Compilers [1433]. Complement [2472]. Complementarity [1210]. Complementary [519, 834]. Complete [1130, 1021, 1211, 1032, 1136]. Completion [1022, 501]. Complex [1761, 580, 1499, 648, 487, 667, 1816, 635, 238, 2426, 1766, 12, 1810]. Complex-Valued [635]. Complexity [973, 2588, 2591, 890, 1449, 876, 2061, 2431, 1106, 2430, 2429, 1876, 903, 905, 2526, 910, 986, 2516, 2259, 1024, 916, 1672, 1850, 1895, 1026, 925, 553, 997, 622]. Complexity-Theoretic [2430, 997]. Compliant [643]. Complicated [140]. Component [593, 2657, 2225, 590, 603, 1480, 1404, 2405, 2662, 2607, 601, 592, 1520, 1337, 2402, 2406]. Component-Based [2607, 1520]. Component-Oriented [2405, 2402]. Components [2397, 2404, 732, 2669, 1345, 1520, 589, 2670, 2651, 2659, 720]. Composable [1029]. Compose [1639]. Composing [245, 454]. Composite [992, 1896]. Composites [31]. Composition [1964, 7, 2401, 1535, 1791, 2631, 2244]. Compositional [1642, 1804, 1540, 1231, 396, 1000, 1632]. Compositions [2406]. Compounds [2349]. Comprehensive [2391]. Compressed [981]. Compressible [2498]. Compressing [247, 150]. Compression [312, 148, 310, 1605, 1337]. Compression-Decompression [1337]. Compressor [1357]. Computability [2518, 1111]. Computable [894, 929]. Computation [2414, 975, 436, 833, 876, 2278, 1032, 1455, 2191, 1033, 880, 754, 1620, 404, 1037, 2521, 416, 728, 2442, 2451, 1520, 2486, 324]. Computational [974, 886, 101, 2257, 1407, 903, 1521, 2297, 122, 2493, 529, 1757, 447, 700, 622, 688, 1742, 942, 1810]. Computationally [2496]. Computations [863, 1636, 2485]. Computer [1566, 1458, 100, 2390, 550, 42, 2497, 2444, 2490, 359, 1143, 1823, 1824, 121, 2417, 467, 971, 2217, 2451, 450, 2418, 108, 97, 83, 664, 2469]. Computer-Aided [2444, 359, 2417]. Computer-Based [100]. Computer-Supported [2390]. Computers [1300, 1453, 1496]. Computing [2603, 1738, 2344, 1517, 2280, 434, 1502, 1756, 1391, 1866, 891, 1167, 160, 1346, 460, 461, 1430, 437, 510, 2118, 1575, 737, 1187, 1519, 1407, 1387, 1518, 1295, 1293, 1404, 2484, 2465, 1299, 1354, 442, 1513, 1450, 1673, 149, 464, 2468, 424, 1344, 2665, 925, 447, 1515, 1504, 1597, 832, 715]. Concentration [616]. Concept [950, 225, 208, 216, 213, 2137, 1735, 2320, 2352, 83, 223, 2269]. Conception [1720]. Concepts [279, 2366, 2300, 301]. Conceptual [205, 1171, 209, 2367, 1551, 2605, 218, 213, 220, 217, 280, 2146, 226, 2004]. ConcernBASE [2632]. Concerns [2660]. Conclusion [2167]. Concrete [1901]. Concurrency [1445, 2246, 990]. Concurrent [482, 390, 1613, 2606, 444, 1476, 446, 2648, 753]. Condition [176]. Condition-Driven [176]. Conditional [1153, 2306, 1155, 815, 1462, 2005, 1200, 1172]. Conditionally [593]. Conditionals [1195]. Conditioning [536, 673]. Conditions [1804, 1719, 914, 394]. Cones [987, 821]. Conference [2655, 167]. Confidence [963, 1693, 132, 2148]. Confidence-Rated [1693]. Configurable [1295, 114]. Configuration [1972, 2311, 2662, 1307, 1736, 2351, 1328, 1312]. Configurations [1640, 2328]. Configurator [2310].

Confirmer [195]. Conflict [35, 1245, 2416]. Conflicts [1205, 1192]. Conformal [330]. Congestion [2198, 2208, 2196, 2190, 1726, 1728, 1957, 1677, 1665]. Congruences [1992]. Conjugate [2463]. Connected [732, 262, 1512]. Connecting [1407]. Connection [2306, 1962]. Connectionist [48, 599]. Connectionist-Symbolic [48]. Connectionist/Markovian [599]. Connections [2204, 678]. Connectivity [499]. Conquer [719]. Consciousness [1783]. Consensus [421, 2542, 2079]. Consequence [1694]. Consequence-Finding [1694]. Consequences [2692]. Conservation [1270]. Consider [41]. Considered [441, 860]. Considering [1427]. Consistency [2640, 1636]. Constant [1034, 1843, 514, 1037, 1314, 1494, 554, 1349]. Constant-Round [1037]. Constants [1209]. Constrained [1902, 243, 1874, 500, 1377, 1693, 1591, 1665, 1116]. Constraint [398, 2303, 2371, 159, 1736, 1702, 1547]. Constraint-Based [2303, 2371, 159, 1736, 1547]. Constraints [2247, 2328, 393, 1773, 1847, 401, 377, 2240, 2648, 1559, 153]. Construct [1159, 195]. Constructing [1661, 159]. Construction [1567, 1045, 2540, 2529, 2543, 1699, 2137, 1161, 479, 651, 2145, 1328, 21]. Constructions [1859]. Constructive [1120, 100, 2600]. Constructors [1141, 2097]. Constructs [2405, 1328]. Consultancy [2384]. Consumer [2574, 1062]. Consumption [2179]. Contact [1082]. Contact-Less [1082]. Contacts [2534]. Content [2615, 305, 2233, 2610, 765, 297, 1946, 1541, 2613, 1871, 24, 1727, 2240, 298, 1536, 1729]. Content-Aware [1727]. Content-Based [765, 297, 1541, 2613, 1871, 2240, 298]. Contention [1427]. Contents [285]. Contests [1776]. Context [64, 1974, 35, 36, 38, 39, 66, 40, 2364, 1066, 68, 1353, 69, 41, 42, 34, 43, 70, 127, 1069, 45, 72, 73, 74, 214, 1384, 1981, 49, 50, 51, 53, 922, 54, 527, 2562, 77, 47, 1134, 58, 61, 935]. Context-Aware [1974]. Context-Based [35, 47]. Context-Dependent [69]. Context-Driven [70, 73]. Context-Free [1353, 1384, 935]. Context-Sensitive [922]. Contexts [44, 71, 75, 48, 60]. Contextual [67, 484, 1546, 57, 59, 62, 63]. Contextualised [78]. Contextualising [76]. Contingent [1532]. Continuous [2198, 237, 828, 580, 485, 2071, 2466, 635, 2282, 839, 630, 599, 2029]. Continuous-Time [630]. Continuously [246]. Contour [840, 332, 336, 825, 2552, 356]. Contract [1097, 2401, 2651, 147]. Contraction [2304]. Contracts [2403]. Contrast [41, 1336, 656]. Contrast-Enhanced [656]. Contribution [320, 2207]. Contributions [2445]. Control [2198, 646, 2481, 1669, 2185, 2199, 2182, 2173, 1445, 2179, 2208, 2693, 2196, 1714, 2679, 1181, 636, 642, 650, 2190, 634, 652, 2617, 1726, 1797, 2329, 2554, 2680, 2184, 1670, 1527, 1808, 1975, 2445, 2443, 2442, 2446, 1712, 643, 2219, 1611, 1681, 1715, 2319, 1677, 1665, 1012, 2273, 1668, 1801, 645, 2152, 1790, 2551, 1646, 1671]. Controllability [1342]. Controllable [1767]. Controlled [1728, 607, 345, 2183]. Controllers [2441, 644]. Controlling [672, 2606]. Conundrum [1197]. Convection [1487]. Convection-Dominated [1487]. Conventions [1776]. Convergence [600, 2078, 2477, 1809, 2508]. Convergent [2122]. Conversation [89]. Conversor [1314]. Converting [909]. CONVEX [1863, 1586]. Conveyor [906]. Convolution [1286, 322]. Cooperation [1654, 2579, 2228, 634, 2234, 720]. Cooperative [1648, 2568, 634, 1659, 1798, 25, 1009, 4, 2232]. Cooperatively [219]. Cooperatively-Built [219]. Coordinate [634, 1892]. Coordinated [1651]. Coordinating [2250]. Coordination [437, 71, 1713, 2581]. Coprocessor [1361, 1889]. Copy [1924]. Copy-Detection [1924]. Copyright [18]. Copyrighted [949]. Coq [2456, 1587, 1574]. CORBA [1519, 464, 1520, 449]. CORBA-Based [464]. Core [1242, 1005, 1332]. Cores [1352, 1340, 1337]. Corner [2551]. CORONET [2380]. CORONET-Train [2380]. Corpora [66, 2000, 2002, 2011]. Corporate [2391, 2388]. Corpus [683, 2002, 2053]. Correct [1571, 1233]. Corrected [462]. Correcting [2135, 490]. Correction [2618, 2002]. Correctness [1247, 785, 786, 1105]. Correlation [818, 2071, 1053, 568, 1546, 572, 540, 674, 128]. Cortex [678]. Cortical [670]. Coscheduling [712, 1507]. Cost [1561, 1057, 2082, 749, 239, 2251]. Cost-Sensitive [2082]. Costs [1851]. COTS [2685]. Count [2533, 132]. Counter [1125, 1078]. Counter-Measures [1078]. Counterexamples [408]. Counterfactual [46]. Countermeasure [1885, 188]. Counting [1473, 1014]. Coupled [563, 681]. Coupling [1846]. Course [273]. Courtship [1777]. Covariance [818]. Cover [326, 1863]. Coverability [1227]. Coverage [2362]. Covering [968]. Covers [1866]. Covert [962]. CoWing [2561]. CPPvm [709]. Crashes [2526]. Crawling [2258, 2336]. Creating [2378, 246, 6]. Creation [102, 1076, 2053]. Credit [760, 536]. Credulous [2304, 1201]. Credulously [1194]. crisp [2122, 1323]. Criteria [2677, 360, 2574]. Criterion [1114, 1105, 1371, 2508]. Critical [2682, 2348, 2674]. Critique [120]. Cross [1923, 1980, 1941, 1707, 1924]. Cross-Entropy [1980]. Cross-Language [1941]. Cross-Referencing [1924]. Cross-Validation [1707]. Cryptanalysis [1052, 164, 165]. Crypto [785, 786, 793, 792, 1889]. Crypto-algorithms [793, 792]. Crypto-systems [793, 792]. Cryptographic [178, 415, 2571, 2430, 170, 188]. Cryptography [777, 975, 1911, 1894, 200]. Cryptology [1288]. Cryptoprocessor [1899]. Cryptosystem [169, 1055, 164]. Cryptosystems [1287, 1892]. Crystals [1485]. CSP [2316]. CTMC [1985]. Cube [1602]. Cubes [737]. Cuckoo [1845]. Culling [1497]. Cultural [1804, 1918, 1809]. Culture [2227]. Cumulants [591]. Current [1517, 1999, 1955]. Curse [1601]. Curvature [836, 522]. Curvatures [577]. Curve [836, 1039, 199, 1911, 1892, 1909, 1079, 201, 1891, 200]. Curved [980]. Curves [1038, 1913, 557, 362, 1893, 581]. Custom [1244, 1346]. Customizable [1918, 1921]. Cut [1162, 2591]. Cut-Elimination [2591]. cut/Max [827]. Cuts [1878]. Cutwidth [1867]. Cyber [187, 1815]. Cyberspace [774, 773]. Cyborgs [81]. Cycle [1866, 1450, 1824]. Cycles [1866, 1878, 310, 262, 1290]. Cyclosporine [616]. Czech [2000, 2006, 2033, 2019, 2009, 2010, 2015].

D [456, 2504, 458, 2557, 334, 703, 2618, 811, 2494, 426, 808, 1493, 845, 355, 2547, 348, 835, 740]. D-Objects [355]. Danish [2490, 2050]. Darboux [378]. Darwinian [1389]. Darwinist [1745]. Data [562, 2627, 205, 1095, 15, 1836, 735, 2215, 148, 2116, 2338, 1625, 1883, 1088, 2037, 96, 160, 1383, 2337, 1478, 461, 1438, 1908, 2062, 1184, 1475, 2117, 1714, 737, 534, 1173, 2121, 438, 878, 819, 2464, 1419, 939, 152, 337, 2127, 2553, 628, 2132, 2563, 1556, 1604, 530, 2079, 1940, 1442, 254, 2386, 2083, 2138, 2564, 2226, 2260, 551, 139, 142, 1440, 154, 141, 710, 2109, 2155, 263, 1807, 1939, 746, 1357, 1100, 2142, 252, 253, 2145, 282, 132, 1555, 298, 311, 1089, 1459, 699, 533, 272, 273, 428, 852, 1012, 520, 1364, 281, 140, 2110]. Data [2156, 583, 6, 2383, 2102, 135, 2103, 2549, 2153, 348, 609, 156, 1709, 134, 867, 1399, 158, 2128]. Data-Dependent [1419]. Data-Intensive [699]. Data-Parallel [1438, 438]. Data-Sharing [852]. Databank [726]. Database [125, 2604, 2224, 2245, 1561, 5, 2628, 127, 741, 2548, 985, 2606, 1552, 24, 2028, 2636, 1444, 1550, 2695, 1554, 1101, 143, 1559]. Database-Driven [24]. Databases [2233, 1445, 240, 124, 2605, 1526, 2024, 1442, 149, 233, 1454, 1109, 2150, 2151, 153, 248, 1443, 2125]. Dataflow [2607, 450]. Datastreams [1328]. Dates [1606]. DAVID [2173]. DBMS [1405]. DBSCAN [1448]. DCT [315]. Deactivating [2202]. Dead [1537]. Deadline [494]. Deadlines [915]. Deadlock [1492]. Dealing [1527]. Debate [2685]. Deblur [960]. Debug [438, 1341]. Debugging [2481, 1237, 207, 2321]. Decentralized [2191, 1802]. Deceptive [2358]. Decidability [1123]. Decidable [986, 1122]. Decision [886, 1135, 876, 321, 1165, 2066, 1180, 284, 1147, 2423, 1143, 2079, 1126, 2082, 2564, 1583, 2087, 1565, 2149, 2100]. Decisions [2673, 2077, 1755]. Declarative [2270, 2258, 2285]. Decodability [1650]. Decomposition [456, 1159, 1487, 1156, 882, 1842]. Decompositions [879, 2415]. Decompression [1337]. Decoupled [2303]. Dedekind [1014]. Dedicated [501, 1775, 453, 1507]. Deduction [1325, 2597, 1121, 1136, 1708]. Default [1168, 1197, 2263, 2268, 2262, 2269]. Defeasible [2265]. Defect [351]. Defense [2315]. Deferring [545]. Defined [1769, 2124, 2161, 2162, 2165, 191]. Definite [2007]. Definition [2355, 2212, 2236, 930]. Definitions [1551]. Deformable [336]. Degraded [1719, 350]. Degree [1279, 1390, 262, 970, 504, 1867, 431].

Degree- [504]. Degrees [506]. Deictic [1658]. Deictics [2050]. Delaunay [1861]. Delay [2206]. Delayed [86]. Delegation [2557]. Delivery [155]. Demand [1297, 1699, 1717, 2253]. Demand-Driven [1699]. Democratic [192]. Demographic [2116]. Demonstrating [1337]. Demonstration [2447]. Dempster [1170]. Denial [794, 188]. Denial-of-Service [188]. DenK [1659]. Denoising [605, 805]. Denotational [1248]. Dense [151, 985, 994]. Densities [535]. Dependability [2694]. Dependence [38, 1441]. Dependencies [610]. Dependency [1999, 49, 575, 728, 2663]. Dependent [2302, 2048, 69, 1023, 1419, 2273, 235]. Deploying [698, 2231]. Deployment [2205]. Deposit [1932]. Deposition [2495]. Depressing [671]. Depth [514, 554]. Deques [1087]. Derandomization [514]. Deregulated [1097]. Derivation [1135]. Deriving [132, 2237, 1221]. Descent [531, 730]. Describing [2385]. Description [435, 305, 661, 2370, 1989, 2292, 2293, 2420, 2300, 425, 2007, 1818, 301, 2269]. Descriptions [2059, 2361, 214]. Descriptor [839]. Descriptors [302, 304, 303]. Design [2441, 2641, 950, 2349, 2245, 2428, 84, 1561, 110, 2413, 1947, 3, 1532, 2037, 2693, 1242, 2430, 2605, 1723, 2205, 2444, 1558, 2405, 1320, 422, 856, 1087, 1937, 2669, 2401, 1664, 1264, 2636, 2415, 1216, 1716, 2197, 2443, 1229, 1332, 2421, 452, 1966, 644, 1897, 83, 1223, 1338, 2661, 1342, 521, 2631, 379, 248]. Design-Level [1342]. designated [198]. Designed [2541]. Designing [1016, 809, 1088, 2673, 93, 816, 2398, 2631]. Desires [1192]. Desktop [2546]. Despite [1472]. Destructible [1767]. Detailed [670, 1318]. Detecting [2116, 659, 275, 1554]. Detection [956, 1497, 1245, 550, 663, 1475, 655, 1180, 294, 2005, 2083, 546, 683, 2002, 291, 366, 2616, 356, 671, 955, 183, 1924].

Detection/Correction [2002]. Detector [427, 2094]. Detectors [2036]. Determinant [266]. Determinants [258]. Determination [2032, 347, 1399]. Determining [2074, 2044, 431]. Deterministic [2531, 1490, 1859, 1685, 172]. Detour [1862]. Develop [2575]. Developing [76]. Development [1517, 2654, 433, 2428, 1540, 2431, 2483, 1953, 2405, 1320, 2243, 2671, 2491, 2047, 2387, 1952, 2650, 2, 4, 1345, 2546, 2388, 119, 2379]. Developments [1929]. Deviating [174]. Device [1326, 1734, 200]. Devices [2583, 1902, 2585, 1076, 1062]. DFA [1084, 999]. DFT [308]. Diagnosing [2272]. Diagnosis [2310, 2327, 2314, 1143, 2347]. Diagram [2137]. Diagrammatic [1649]. Diagrams [2626, 2642, 992, 1156, 1649, 1597, 2625]. Dialectical [1201]. Dialog [2558]. Dialogos [2048]. Dialogue [2056, 1655, 2047, 2046, 2052, 2055, 2043, 2049, 2616, 2053, 2054, 2045]. Diameter [1395]. Diaphony [2474]. Diatoms [284]. Dictators [516]. Dictionaries [1860]. Dictionary [2015]. DICUT [503]. Difference [2507, 1658, 376]. Differences [72]. Different [1994, 141, 2273, 2106]. Differential [1911, 993]. Differentiated [1675, 2209, 1720, 1678, 1682, 2178]. Differentiating [591]. Differentiation [1721, 1677, 2194]. Difficult [1739, 1658]. Difficulty [1779]. Diffie [1039]. DiffServ [2182, 2186, 2200]. Diffusion [418, 2466, 2520, 839, 844]. Digit [2067, 654, 1350]. Digit-Recognition-Example [2067]. Digit-Serial [1350]. Digital [1933, 2441, 1927, 1923, 1945, 1947, 2540, 1936, 1929, 951, 1931, 1870, 1930, 102, 1946, 2617, 16, 1940, 369, 2018, 1952, 1916, 1919, 1935, 1955, 1932, 1079, 938, 350, 1926, 959, 960, 2549, 367]. Digitization [1950]. Digraphs [912]. Dijkstra [1855, 1365]. DiKe [2327].

Dilemma [538]. Dimension

[69, 1286, 554, 1821]. Dimensional

[1602, 126, 2513, 1770, 2120, 1500, 2548, 1487, 358, 729, 925, 1765, 1289, 2496, 150, 2283]. Dimensionality [1601, 528]. Dimensioning [1684]. Dimensions [38, 110, 1815, 1672, 111, 269]. Diophantine [321]. DIPORSI [747]. Direct [1296, 1491, 535, 1658, 1273]. Directed [897, 445, 2670, 822]. Direction [245, 241]. Directional [577]. Directions [232, 1894]. Director [1829]. Disaggregation [2468]. Disambiguating [1926]. Disambiguation [1998, 2019, 2003]. Discourse [2001, 2050]. Discovering [1400, 610, 318, 1701, 2139, 242, 2103, 2104]. Discovery [2233, 1524, 2117, 70, 2130, 2129, 1442, 2098, 2150, 1386, 2111, 2157, 2112, 2158, 2244]. Discrete [973, 828, 610, 2109, 2155, 1356, 618, 1324, 201]. Discrete-Time [618]. Discrete/Continuous [828]. Discretization [2105]. Discriminant [537, 2030, 2089]. Discriminating [292]. Discrimination [633, 574]. Discriminative [528]. Discussion [797, 772, 800, 789, 786, 801, 775, 769, 777, 782, 787, 771, 780, 776, 793, 791, 796, 784, 798, 774, 794, 2394, 778]. Disjoint [1872, 928]. Disjunctive [2263, 2281, 2266]. Disk [492, 1874]. Disks [1840]. Displacement [739]. Display [2614]. Disruption [86]. Dissection [2545]. Dissemination [160, 949]. Dissimilarity [274]. Dissimilarity-Based [274]. Distance [1961, 1383, 2121, 1874, 2071, 1875, 662, 1849, 1389, 832, 2670]. Distances [2135, 1148, 993, 1392, 1838]. Distinguishing [2144]. Distortion [939, 314, 309]. Distortion-Free [939]. Distributed [1928, 784, 1927, 1769, 2182, 855, 2455, 459, 1445, 732, 2693, 2239, 2225, 1034, 1864, 1470, 1519, 865, 2519, 1865, 2532, 249, 1530, 2311, 731, 1723, 1468, 1971, 747, 476, 2229, 2530, 1508, 442, 1949, 745, 1557, 1967, 783, 1442, 2211, 1630, 764, 1413, 416, 2230, 1716, 1009, 2460, 1553, 2231, 450, 1520, 449, 194, 1979, 133, 143]. Distributing [1814]. Distribution [1497, 178, 2215, 2481, 833, 32, 2497, 1004, 377, 264, 167, 2432, 1717, 2102, 286, 183]. Distributions [2466, 364, 1987]. Diversiform [688]. Diversity [2106]. Divide [719, 1639]. Divide-and-Conquer [719]. Divider [1232]. Divisibility [989]. Division [1773, 266]. Division-Free [266]. Divisor [169]. Divisors [1282]. DJ [1434]. DJ-Graphs [1434]. DLs [1953]. DLV [2292, 2293]. Do [2526, 2377]. DO-178B [2689]. Do-All [2526]. Document [361, 588, 26, 1942, 687, 155, 136, 1951, 2031, 2143, 1917, 1920, 955, 131, 285, 960, 1559]. Documents [1923, 937, 1938, 1595, 2250, 6, 1708]. Dog [119]. Domain [456, 735, 2048, 1918, 1576, 2505, 73, 2356, 2568, 1487, 2046, 2360, 207, 2143, 2235, 1934, 2273, 2098, 344, 1291, 206, 1978]. Domain-Dependent [2273]. Domain-Independent [2568, 1934]. Domain-Oriented [207]. Domains [1225, 17, 43, 2376, 2316, 1243, 1151, 2107]. Dominated [1487]. DOMINATING [885]. Don’t [2076]. DOS-Resistant [791, 790]. Double [2068, 759, 763, 804]. Double-Loop [804]. Doubt [1200]. Downbeats [691]. DPA [1912, 1884]. DQL [2085]. Drawing [1876, 1377]. Dreyfus [94]. Drift [82]. Driven [1561, 893, 2037, 1304, 70, 73, 1699, 24, 760, 176, 1600]. Driving [86]. DSA [1035, 1285]. DSA-Like [1285]. DSMC [462]. DSP [1347, 1332, 1337]. DSS [1685]. DTD [137]. DTDs [5]. Dtrees [1159]. Dual [493, 496, 2680, 502, 2506]. Duality [1840]. Duckling [796, 795]. Duplex [1357]. Duplication [391]. Duration [1221]. Durations [1994]. during [1492, 330]. Dutch [2031]. DV [2199]. DWE [1935]. Dyadic [2474]. Dynamic [1974, 2509, 723, 2373, 1618, 2242, 2372, 1482, 1995, 341, 1964, 661, 2370, 2367, 719, 2210, 332, 1492, 2274, 712, 213, 2376, 1514, 1938, 2283, 1552, 24, 2255, 1409, 1467, 2347, 718, 2353, 1512, 1957, 2503, 1012, 452, 2040, 817, 2398, 678, 668, 656]. Dynamical [889, 675, 2442]. Dynamically [1429, 1352, 1795, 1340, 1334]. Dynamically-Rearranging [1795]. Dynamics [2504, 1744, 896, 573, 627, 626, 831, 2098].

E-check [9]. E-Commerce

[12, 162, 24, 2575, 22, 4]. E-Government [115]. E-marketplaces [2244]. E-Markets [760]. E-Speak [26]. e-Texts [2014]. E2 [487]. Early [666]. Earth [1947]. Earthmoving [2543]. Earthwork [2540]. Easily [698]. Easy [901, 1590, 2262]. ECC [1912]. ECG [605, 335, 2147]. ECN [1677]. Eco [1743]. Eco-Grammars [1743]. Ecosystem [1750]. Ecosystems [2576, 2553]. Eden [1750]. Edge [1864, 499, 980, 337, 825]. Edge-Based [337]. Edge-Connectivity [499]. Edit [832]. Editing [1329]. Edition [1921]. Editions [1924]. Editor [2324]. EEG [602, 607, 608]. EEG-Segments [608]. Effect [43, 2489, 414, 2263, 1768]. Effective [2518, 1125, 393, 702]. Effectiveness [1372, 2688]. Effects [2196, 86, 1784, 1941, 50, 1444, 119]. Efficiency [945, 1092]. Efficient [432, 1091, 975, 2534, 1352, 2214, 1220, 1058, 126, 2535, 1638, 1460, 1237, 1003, 1461, 1526, 1049, 1038, 981, 2584, 1991, 617, 633, 1087, 1509, 1633, 250, 1871, 404, 1970, 1452, 193, 1892, 252, 743, 1896, 1441, 1907, 1890, 27, 1718, 1707, 251, 324, 2496, 150, 1349, 1249]. Efficiently [2074, 149, 1527, 832, 1559]. Effort [2654, 1668]. Eigenvalue [1485]. Elastic [1669, 2475]. Elasticity [2501]. Election [2523, 2531]. Elections [1183]. Electric [614]. Electro [1530].

Electro-Magnetic [1530]. Electromagnetic [1901, 2608, 470, 1078]. Electron [2475, 2476, 2480]. Electronic [2343, 8, 17, 9, 2249, 1921, 138, 2574, 19, 1080, 1330, 1549, 1915, 1082, 11]. Element [2509, 460, 461, 340, 1487, 443, 2506, 2472, 750, 1324]. Elements [137]. Elevator [2329]. Elgot [1124]. Elicitation [2386]. Eligibility [2064]. Elimination [2591, 390, 1159, 1435]. Eliminations [2597]. Ellipse [1180]. Elliptic [1039, 199, 1038, 2470, 1913, 1911, 2471, 1892, 1909, 1079, 1893, 201, 1891, 200]. Elliptical [546]. EMA [1078]. Embedded [1881, 800, 1326, 2585, 907, 647, 1501, 2642, 1298, 1334, 1495, 799]. Embedding [939, 1582, 406, 1477]. Embeddings [511, 512]. Embodied [92]. Embodiment [87]. Emergence [1781, 1772, 1798, 1764, 1757, 1742, 637]. Emergent [1774, 83, 1810]. Empirical [1652, 1164, 1372, 1374, 2081, 1603, 215]. Empirically [1650]. Employing [1972, 2243]. Emptiness [1639]. Emulator [1295]. Enable [1974]. Enabled [25, 848]. Enabling [2583, 1403, 1722]. Encoded [2153]. Encoding [2061, 2274, 319]. Encodings [1140]. Encryption [1040, 1908, 189, 1046, 1041, 1308, 1309, 1903, 1896]. End [94, 2199, 1721, 2206, 2286, 2052, 1593, 2551, 182]. End-System [1721]. End-to-End [2199, 2206, 1593, 182]. Endomorphisms [1038]. Endpoint [2185, 2183]. Energy [827, 2179, 840, 829, 804]. Enforcing [260]. Engine [1325, 861, 26]. Engineering [2645, 1232, 24, 2607, 2393, 2544, 115, 1674, 355]. English [180]. Enhance [741]. Enhanced [159, 189, 2618, 656, 521]. Enhancement [1336, 601]. Enhancements [1670]. Enhancing [1923, 139, 1939, 2015, 1539]. Enrichment [2020]. Enrolment [1072]. Ensemble [616, 549, 276, 1546]. Ensembles [2106]. Entailments [1189]. Enterprise [2242, 2248]. Enterprises [2581, 1535, 2383].

Entertainment [1790]. Entropy [1980, 533]. Enumerable [1010]. Environment [2373, 433, 100, 1744, 1660, 2482, 1532, 1746, 1964, 1524, 2483, 1237, 719, 9, 1770, 1374, 1518, 1720, 871, 2548, 442, 2447, 717, 1935, 1409, 2055, 746, 1725, 651, 345, 2178, 1755, 1962]. Environmental [761]. Environments [2323, 2372, 1522, 2239, 86, 1408, 1748, 2671, 2584, 2557, 91, 16, 697, 464, 1344, 1764, 2241, 30, 1516, 733, 1529]. Epilogue [873]. Epipolar [381]. Episodic [676]. EProMS [4]. Equation [2513, 2098, 653, 2103]. Equational [1137, 1136, 930, 2319]. Equations [2511, 2498, 1455, 838, 2449, 2478, 1486, 921, 539]. Equilibria [2572]. Equilibrium [1160]. Equitable [517]. Equivalence [493, 992, 1539, 2262]. Equivalences [2306, 1628]. Equivalent [2687]. Erlanger [1272]. Erroneous [337]. Error [310, 2071, 2078, 2135, 683, 958, 490, 565, 183, 2106]. Error-Correcting [490]. Errors [2002, 2508]. Escrow [797, 1050]. Essence [2622]. Essential [56]. Essentially [977]. EST [468]. Establishing [1778]. Esterel [1224]. Estimates [2208, 2495]. Estimating [2654, 2060, 403, 1595, 286, 1392]. Estimation [806, 2473, 833, 1484, 2036, 340, 535, 603, 805, 1297, 238, 342, 362, 606, 1485, 344, 622, 594, 239, 367]. Estimators [2665]. Estuarine [729]. Eternity [952]. Ethernet [1422]. Ethnography [1937]. Euclidean [512, 2120, 2141]. Euler [653]. Eulerian [2490, 912]. Europe [2495]. European [1869, 1366]. EuroWordNet [2039]. Evaluate [2196]. Evaluating [747, 2658, 857, 1915]. Evaluation [890, 2694, 1244, 2182, 1947, 1602, 240, 1428, 1372, 1343, 1993, 2488, 1173, 1374, 2209, 1931, 385, 1720, 2684, 174, 1614, 662, 1557, 1548, 1937, 2047, 2081, 1416, 1476, 1348, 2336, 961, 733, 272, 1324, 1471, 1223, 2100, 2668, 1443]. Evaluator [2459]. Event [870, 2116, 1470, 1503, 1940, 994, 448]. Event-Based [1940]. Events [1205, 2275, 1206, 63]. Every [1005]. Evidence [1184, 2084]. Evoked [603, 606]. Evolution [872, 300, 1780, 1804, 1339, 1782, 157, 2225, 1794, 2249, 1813, 158, 1783, 1784, 1762, 2507, 1764, 1788, 1809, 632, 1821, 147, 631, 2139]. Evolutionarily [1749]. Evolutionary [833, 1781, 1806, 2433, 627, 1354, 2423, 295, 1772, 2082, 1785, 1750, 756, 2301, 1787, 1392]. Evolvability [1766, 1789, 1791]. Evolvable [1790]. Evolved [1768]. Evolving [1795, 1748, 650, 1786, 1775, 1821]. Exact [887, 509, 1575, 1868, 1838, 1392]. Exact-IEBP [1392]. Examine [438, 102]. Example [2067]. Examples [1692, 325, 295, 807, 660]. Exchange [199, 2578, 1645]. Exclusion [2517, 2516]. Exclusive [446]. Executable [1435, 867]. Execute [2326]. Executing [1653]. Execution [2372, 2455, 1964, 2624, 724, 1464, 430]. Exhibitions [1946]. Existential [1107, 2300]. Existing [1076]. Expansion [67, 2018]. Experience [1686, 1569, 2390, 2391, 2692, 92, 2385, 2393, 852]. Experiences [1457, 2654, 1087, 143]. Experiment [1456, 92]. Experimental [1095, 1097, 117, 827, 1602, 1096, 385, 2684, 439, 1198, 1594, 52, 1393, 586, 1443]. Experiments [1448, 2325, 1391, 1661, 1596, 1598, 1752, 122, 1587, 2687]. Expert [1938, 2616]. Explanation [62]. Explanatory [1207]. Explicit [260, 463, 467, 1859, 1753, 1778, 1481, 59]. Explicitly [2269]. Explicitness [782, 781]. Exploiting [2338, 2306, 1236, 1093, 1427]. Exploration [422]. Exploring [2064, 1295, 292]. Exponent [1278]. Exponential [630]. Exponentials [1157]. Exponentiation [1905]. Express [1145]. Expressed [468]. Expressibility [2264]. Expressing [1994]. Expression [530, 2089, 1084, 2410, 2111, 2157]. Expressions [1856, 1659]. Expressive [1103, 2104]. Expressiveness [2263]. Extended [1313, 90, 407, 2286, 2516, 1643, 399, 2226, 2236, 2269]. Extending [1212, 2645, 2228, 2567, 1636, 2636]. Extends [517]. Extensible [1090, 1539]. Extension [433, 2540, 1466, 177, 1194, 222, 1125, 628, 1126]. Extensionality [1114]. Extensions [711, 227, 541]. External [1836, 1450]. Externally [1767]. Extract [2337, 684]. Extracting [562, 5, 1541, 667, 2318]. Extraction [593, 36, 2258, 2592, 692, 2072, 2123, 2131, 825, 598, 154, 568, 2140, 2041, 2552, 327, 2101]. Extradimensional [1780]. Extrusion [2551]. Eye [655].

F [670]. F5 [957]. Face [341, 291]. Faces [113, 1288, 660]. Faceted [28]. Facilities [433]. Facility [496, 1414, 502]. Fact [82]. Factor [1400, 544]. Factoring [1281]. Factors [1942, 166, 2681]. Factory [2391]. Facts [2318]. Fail [2680, 195]. Fail-Safe [2680]. Fail-Stop [195]. Failure [953, 427, 1718]. Fair [407, 2191, 1645]. Fairness [2190, 756, 1622]. Falling [2507]. False [1381]. FAM [659, 275]. FAM-Based [275]. Families [510, 985]. Family [886, 2660]. Fast [562, 1869, 2510, 1883, 1343, 901, 1862, 1280, 2122, 1297, 1084, 342, 742, 733, 808, 522, 368, 722, 309, 594, 1574]. Faster [600, 1852, 1038, 1985, 1626]. Faster-than [1626]. Fault [789, 1746, 708, 2691, 1343, 1873, 2684, 2686, 788, 1504, 2687]. Fault-Tolerant [1746, 1873]. Feasibility [1108]. Feasible [1098]. Featherweight [2600]. Feature [2058, 2365, 339, 580, 1958, 283, 598, 568, 374, 319, 2041, 1546, 2544, 844, 27, 327, 2106]. Featureless [296]. Features [2595, 365, 687, 1699, 2093, 322, 83, 292]. FEC [1667, 2213]. Federated [2245, 1521, 1930, 1939, 143]. Federation [870]. Feedback [279, 86, 2387, 130]. Feedforward [830]. FEM [2470]. fetching [1544]. Few [660, 162]. fGREP [1297]. Fiber [2187]. Fibre [1527]. Fictitious [2505]. Field [561, 823, 1696, 365, 1307, 1896, 201, 470, 812]. Fields [810, 1898, 294, 1895, 170, 678]. Figure [812]. Figure-Ground [812]. File [2519, 1865, 1971, 702]. File-I [702]. File-I/O [702]. Films [2507]. Filter [2037, 2026, 368, 1338]. Filtering [836, 765, 1666, 1871, 2018, 155, 764, 2140, 2143, 1362, 2153, 156]. Filtration [1387]. Financial [627, 536]. Finding [151, 840, 1279, 1390, 1398, 1980, 2313, 1694, 1849, 445, 2148, 998, 1393, 2011, 660]. Finds [691]. Fine [2490, 919]. Fine-Grid [2490]. Fingered [641]. Fingerprint [2339]. Fingerprinting [958, 964]. Finite [2509, 398, 621, 460, 461, 340, 2532, 907, 1636, 1898, 1487, 911, 912, 191, 443, 1895, 2506, 2507, 1584, 921, 2472, 750, 1055, 1587, 201, 376, 884]. Finite-Difference [2507]. Finite-State [621, 2532]. Finite-Tree [398]. FIPA [2583, 1982]. FIR [619]. Fire [745]. Firewall [1070]. First [1210, 2299, 2066, 2466, 1190, 1118, 278, 170, 1635, 1213, 426, 1820]. First- [2466]. First-Order [1210, 2299, 1190, 1118, 278, 426]. Fit [67]. Fitness [1747, 1813, 2071, 627]. Fitting [1180, 293, 502]. Fixed [1052, 1119, 1681]. Fixed-Pattern [1052]. Fixing [1846]. Fixpoint [2281]. FLASH [1231]. Flatness [358]. Flexibility [1654]. Flexible [2363, 1065, 2371, 17, 2405, 1930, 856, 423, 1726, 352, 1077, 1555, 1801]. Flight [1808]. Floating [1240, 1575, 402, 1232, 1284]. Floating-Point [1575, 402, 1232]. Flow [1619, 494, 2497, 2505, 1020, 906, 2222, 294, 842, 2506, 2502, 342, 640, 827]. Flows [492, 2198, 458, 2191, 2500, 1678, 1868, 845, 1549]. Fluctuations [557]. Fluents [2302]. Fluid [2506, 2503]. Flux [845, 2317]. fMRI [2144]. fMRI-Measured [2144]. Foam [2504]. Focus [41]. Focused [2336]. Focussed [396]. Foe [1933]. Foil [1884]. Fold [296]. Folding [577]. Forager [1753]. Force [643]. Ford [445]. Forecast [1457]. Forecasting [1322]. Forensic [942]. Forest [745, 505]. Forests [886, 284]. Forgery [163]. Forgetting [1942]. Form [1912, 1892, 1891, 2379, 1287]. Formal [2641, 1618, 419, 1240, 2690, 1245, 2629, 2691, 225, 1237, 213, 1246, 984, 1232, 1243, 2638, 2642, 1175, 1075, 1585, 1818, 1553, 1229, 2320, 14, 752]. Formalisation [2433, 1578]. Formalism [1763]. Formalisms [222]. Formalization [780, 779, 2001, 2649]. Formalizations [40]. Formalizing [1197, 1586]. Formally [1481, 1223]. Formally-Based [1223]. Format [104]. Formation [760, 1759, 1800, 653, 674]. Formations [1802]. Forms [1280, 2273]. Formulae [516]. Formulas [1212, 2345]. Formulation [1267, 1268]. Forum [2394]. Forward [1047]. Forward-Secure [1047]. Forwarding [2185]. Foundation [695, 1110, 2402]. Foundations [2600]. Founded [2265, 1133, 2266]. Four [2548, 467]. Fourier [1574, 2143]. FPGA [1346, 1353, 1343, 1325, 1306, 1297, 1300, 1336, 1355, 1887, 1308, 1356, 1345, 1331, 1324, 1363, 1342, 1351, 1349]. FPGA-Based [1353, 1343, 1325, 1300, 1356, 1363]. FPGAs [1296, 1304, 1299, 1347, 1317, 1329, 1348, 1364]. FPTAS [1851]. Fractal [1086]. Fractional [968, 1473]. Fragment [1117]. Fragments [1401]. Frame [1315, 2274, 2446]. Frame-Grabber [1315]. Frames [1258]. Framework [1247, 2487, 1153, 2182, 2355, 2196, 2062, 1661, 1996, 9, 70, 289, 1480, 747, 1578, 2312, 1320, 2314, 2578, 2491, 2084, 1175, 411, 1412, 1077, 2665, 301, 1733, 820, 1294, 480, 1674, 2100, 848, 11]. Frameworks [1179, 1201, 2671, 2088]. Free [893, 1353, 939, 1033, 1384, 479, 831, 265, 266, 1013, 935, 804]. Freeing [88]. Freivalds [976]. French [205]. Frequency [1492, 2041]. Friend [1933]. Friendly [3, 2196, 1726, 1665]. Fringe [1215, 1564, 347]. Front [460, 2286, 2052]. Front-End [2286]. Frozen [2595]. Frustum [1497]. FSM [2416]. Full [1971, 1824, 1075, 1357, 2008, 1332, 129]. Full-Duplex [1357]. Full-Rate [1332]. Full-Text [129]. Fully [1130]. Fun [1]. Function [2290, 537, 530, 288, 546, 1348, 536, 733, 606, 594, 1172, 2163]. Functional [2453, 2455, 2429, 2454, 2460, 1504, 1342]. Functionalities [221]. Functionality [2239]. Functionals [821]. Functions [649, 818, 1091, 1171, 1044, 1918, 1054, 580, 2679, 364, 983, 2162, 985, 1005, 1169, 552, 263, 1170, 529, 947, 2146, 543, 1174, 1709]. Fusarium [367]. Fusion [623, 1188, 571, 1173, 2128, 1922, 2456, 446]. Future [1339, 694, 2167, 1955, 1954, 83, 1687]. Fuzziness [1543]. Fuzzy [1209, 279, 550, 1430, 1180, 2126, 2133, 275, 2139, 1009, 30, 382, 2125]. FX [625]. FX-Market [625].

Gabor [844, 1338]. Galois [223]. Gambit [1318]. Game [877, 542, 567, 2297, 1645, 1817, 831]. Game-Based [1645]. Game-Theoretic [542]. Games [2344, 877, 2315, 639, 1644]. Gap [1865]. Gaphyl [2119]. Gaps [2064]. Garbage [857, 858]. Gas [663]. Gate [1318, 1307]. Gate-Arrays [1318]. Gathering [1981, 1970]. Gaussian [569, 571, 603, 557, 2417, 560, 543, 606]. Gaussians [320]. Gaz [762]. Gaz-Guide [762]. Gaze [2618]. Gazettes [762]. GCA [423]. Geb [1782]. Gene [1398, 530, 2089]. General [1572, 1088, 461, 1154, 1877, 1022, 2664, 2134, 256, 2425, 1189, 2677, 1842, 1015].

General-Purpose [2664]. Generalisable [637]. Generalisation [1132, 536]. Generalization [2058, 1124, 555, 620, 1440, 556, 545, 2106]. Generalized [507, 481, 209, 535, 1021, 359, 2597, 568, 528, 612]. Generalizing [223]. Generally [1630]. Generate [206]. Generated [2497, 2255]. Generating [2358, 2671, 1893, 1649]. Generation [457, 2056, 1067, 1051, 1889, 2669, 463, 1316, 1035, 2279, 1954, 2043, 2049, 961, 2552, 2201, 688, 2549, 382, 1651, 2668]. Generative [1345, 2023]. Generator [2657, 1053, 165]. Generators [1592, 898, 1890]. Generic [2404, 2224, 1575, 2289, 1931, 1297, 1763, 1076, 2240, 2422, 1344, 2665, 1109, 2201]. Genetic [1691, 2119, 325, 2139, 1213, 319, 813, 1382, 1788, 688, 835, 379]. Genome [1383]. Genomes [1392]. Genomic [1381, 2567]. Genomics [2140]. Genotype [1766]. Genotype-Phenotype [1766]. Gentzen [2592]. Geodesic [837]. Geographic [1947, 1926]. Geographical [1544]. Geometric [1400, 1857, 1025, 2551, 2107]. Geometry [1811, 1090, 2447, 381, 2547]. Geomorphometrical [2545]. Georeferenced [1953]. German [2002]. Gestures [1657, 1658]. GF [1898, 1909]. Gibbs [556]. Gigabit [1422]. Gigahertz [1299]. Giganet [701]. Gilgamesh [1406]. GIS [2229, 2545, 2608, 851]. GIS-Based [2608]. Given [499]. Global [2246, 1641, 2484, 358, 842, 829, 830, 423]. Globally [143]. Go [567]. Goal [130]. Goal-Oriented [130]. Goals [2220]. Going [1711]. Good [2430, 521]. Goods [1870]. Govern [1980]. Government [115, 778]. GPFS [1457]. GPRS [2178]. GPS [2543]. GPS-Based [2543]. Grabber [1315]. Gradient [2463, 531]. Grain [367]. Grammar [1384, 1269, 401, 2055, 1789]. Grammars [1743, 1769, 1353, 277, 2403]. Grammatical [2131, 2003]. Granularities [244]. Graph [435, 1634, 833, 893, 317, 1876, 499, 2638, 834, 728, 2016, 1868, 1759, 1015, 948]. Graph-Driven [893]. Graph-Matching [317]. Graphic [2443]. Graphical [1179, 1165, 1143, 1541]. Graphics [58, 1651]. Graphs [885, 506, 205, 146, 209, 1878, 1877, 208, 218, 1021, 2624, 1874, 213, 220, 1370, 567, 501, 204, 262, 210, 2137, 1451, 920, 922, 445, 504, 925, 479, 265, 1867, 1434, 430, 269, 1879, 1838]. Grasping [641]. Gravity [2142]. Greedoid [1026]. Greedy [502, 1879]. Greenstone [1927]. Grey [341]. Grid [2464, 2490, 925, 2238, 1517, 1523, 1519, 1518, 1403, 1404, 442, 697, 699, 1515, 1516]. Grids [579, 700]. GRM [1523]. Gröbner [2412]. Grooming [2174]. Ground [812]. Groundness [397]. Group [100, 855, 1424, 2121, 2517, 191, 547, 1735, 467, 166, 1109]. Group-Based [1424]. Grouped [48]. Grouping [840, 1852, 1255, 820, 821, 822]. Groups [2381, 1938, 169, 1056, 1261, 1262, 1263, 2425, 1055]. Growing [2601, 588, 1807, 329]. Growth [2308, 82]. GRT [1375]. GSM [1073, 1332]. GTM [577]. Guarantee [1863]. Guaranteed [1188]. Guarantees [2193, 1963]. Guards [1629]. Guesser [2006]. Guest [534]. Guide [1980, 762]. Guide/Govern [1980]. Guided [375]. Guiding [11]. Gurmukhi [323].

H.323 [1719]. H.323-Based [1719]. Halfedge [1088]. Hammerstein [649]. Hamming [1849, 269]. Hand [641, 1981]. Hand-off [1981]. Handling [1680, 35, 2684, 1560, 1503, 1853, 1195, 2501, 1434]. Handwriting [688]. Handwritten [289, 654, 290]. Hard [899, 901, 1590, 970, 1005, 2333]. Harddisk [1903]. Harder [2226]. Hardware [1303, 1883, 1421, 1888, 1320, 1354, 1228, 1233, 315, 1216, 1909, 1234, 392, 1365, 1338, 1305]. Hardware-Software [1354]. Harmful [441]. Harness [714]. Hash [1045, 1005]. Hash-CBC [1045]. Hashes [2432]. Hashing [1844, 941, 1845]. Haskell [2461, 1478]. Hausdorff [2426]. HAZOPs [2674]. HBT [1299]. HDMR [2496]. Head [367]. Health [595]. Heaps [1451]. Hearing [677]. Heated [2507]. Heavy [1456, 1752]. Hebbian [673]. Hedges [1708]. Heideggerian [94]. Height [1602]. Hellman [1039]. HELM [1568]. Help [98, 2055]. Helps [1990]. Herbrand [1132]. Heritage [1918]. Hermeneutical [1945]. Hermite [1287]. Hessian [1913, 1891]. Heterogeneous [2239, 2257, 626, 1922, 744, 2241, 2237, 1520, 734, 430]. Heterogeneous-Agent [626]. Heuristic [1855, 1193]. Heuristics [2278, 1590, 2316, 1219, 1546, 1368]. Hidden [2114, 809, 2338, 2454, 1384, 1301, 2340]. Hiding [963, 940, 941]. Hierarchical [2414, 735, 2453, 1217, 588, 2310, 1598, 1579, 647, 2079, 1967, 1235, 369, 2415, 1219, 2142, 608, 1410, 133, 587, 812]. Hierarchies [997]. Hierarchy [661, 2370, 1125, 908, 913, 2145, 929]. High [126, 332, 1407, 2120, 1404, 856, 1488, 189, 1509, 1433, 2329, 697, 763, 849, 24, 464, 1887, 1453, 748, 698, 132, 447, 1328, 1363, 604, 1897, 2496, 150, 1082, 1962]. High-Accuracy [1488]. High-Availability [1509]. High-Dimensional [126, 2120, 150]. High-Level [1328]. High-Performance [849, 1453, 748, 698, 447]. High-Radix [1897]. High-Resolution [332, 604]. High-Speed [189, 1082]. High-tech [763]. Higher [2588, 1114, 1573, 1576, 1577, 2449, 1140, 1565, 1234]. Higher-Level [1234]. Higher-Order [1114, 1573, 2449, 1140, 1565]. Highlight [824]. Highly [1607, 1719]. Higraphs [1110]. Hilbert [296]. Hill [2519, 488]. Hill-Climbing [2519, 488]. Hip [334]. Hippocampus [336, 676]. Histogram [584]. Historical [1812, 1926, 1821]. Histories [2318]. History [2314, 82, 1787]. History-Based [2314]. HLA [459]. HMM [1511]. HNM [2038]. Hoc [1730, 2535, 1096, 510]. Hodgkin [669, 679]. Hoeffding [517]. HOL [1578, 1580, 1235, 1585, 1588]. HOL/NuPRL [1585]. Holarchical [2415]. Hold’em [2065]. Holding [1953]. Hole [2527]. Holism [1794, 55]. Home [1366]. Homogeneous [918]. Homogenization [2499]. Homomorphisms [1000]. Hop [2533]. Hopcount [1672]. Hopfield [630]. Horizontal [677, 2469]. Horn [1212, 1103, 393]. Horn-Like [1212]. Horses [177]. Hosting [1514]. Hosts [1973]. Hot [1490]. Hot-Spot [1490]. House [2032]. Household [614]. HPC [1480]. Hub [489]. Hull [1586]. Human [1197, 42, 101, 1051, 75, 2021, 652, 111, 2682, 2425, 1199, 118, 2681, 1777, 664, 348]. Human-Machine [2682]. Humans [1803, 104]. Hunting [119]. Huxley [669, 679]. Hybrid [2419, 418, 2365, 1571, 459, 617, 1751, 1235, 2001, 2506, 1157, 282, 311, 1382, 1454, 2418, 599, 143, 1249]. HYDRA MPI [742]. Hyper [1400, 2447, 867, 1116]. Hyper-Automaton [2447]. Hyper-Code [867]. Hyper-Geometric [1400]. Hypercube [925]. Hypercycles [1741]. Hyperedge [1617]. Hypergraph [366]. Hypergraphs [916]. Hypermedia [13, 1558, 2433]. HyperMirror [89]. Hyperspace [151]. Hypertext [31]. Hypertree [879]. Hypertrees [1024]. Hypotheses [295, 876]. Hypothesis [104].

I/O [1525, 1526, 1528, 702]. IBM [1889, 703]. iButton [1072]. IC [764]. IC2D [2481]. ICA [596, 598, 594]. ICE [14]. iCoIN [2391]. Iconic [124, 386]. IDEA [1908]. Ideal [119]. Ideas [1101]. Identification [649, 775, 2513, 284, 941, 2050, 374, 1934, 2147, 353, 137]. Identify [1806, 391]. Identity [2590, 1040, 1050, 181, 118]. Identity-Based [1040]. IEBP [1392]. IEC [2692]. IEEE [2179]. If [983]. IFS [1457]. II [706, 1254, 1262, 2480, 1646]. III [1263]. Illuminate [1793]. Illumination [818]. ILP [1696, 1699, 1701, 1703, 1706, 1707]. Im [1028]. Image [125, 2361, 124, 339, 2610, 357, 828, 2611, 1499, 310, 824, 297, 337, 364, 805, 102, 662, 1336, 365, 736, 369, 1498, 658, 842, 746, 375, 329, 298, 814, 743, 333, 826, 373, 821, 1496, 290, 1337, 350, 308, 355, 570, 2614, 306, 381, 367]. Imagery [1813, 747]. Images [2603, 361, 836, 332, 939, 2130, 823, 358, 370, 811, 331, 330, 311, 356, 363, 387, 817, 660, 382]. Imaginary [296]. Imaging [328, 640]. IMAGO [1959]. Impact [1538, 403, 189]. Impaired [2055, 2054]. Impediment [2030]. Imperative [390]. Imperatives [1656]. Imperfect [1709]. Imperfectly [370]. Implement [1145]. Implementation [1906, 2441, 1302, 950, 872, 706, 2510, 1484, 384, 2287, 475, 2209, 1888, 1720, 2490, 2405, 2536, 856, 1606, 462, 2166, 1087, 2212, 1886, 444, 1314, 1308, 1786, 170, 1910, 1716, 748, 2494, 467, 1896, 371, 1907, 1324, 1597, 938, 1507, 1330, 1338, 1820, 143, 248]. Implementations [2185, 1908, 1887, 1309, 1079]. Implementing [1506, 863, 1088, 1996, 861, 114, 868, 1301]. Implication [2133]. Implication-Based [2133]. Implications [876]. Implicit [2227, 260, 1552, 2598, 194, 2322]. Importance [1203, 624, 1158, 2094, 1809]. Impossible [1188]. Improve [2075, 1373, 2695, 1342]. Improved [888, 1114, 600, 326, 2475, 2477, 1379, 1048, 165]. Improvement [900, 325, 546, 536, 2382, 622]. Improvements [945, 227]. Improving [1491, 190, 1445, 146, 2061, 1092, 1528, 2557, 1462, 2277, 1287, 1527, 2101, 474, 1380].

Improvise [1651]. Impulsive [366]. Incidents [2683]. Incomplete [977, 1211, 1144, 337, 2315, 1673, 2268]. Incompleteness [408]. Incompressible [458, 2500]. Inconsistency [141]. Incorporating [765]. Incorporation [692]. Increase [2246, 946]. Increasing [1346, 449, 880]. Incremental [2362, 498, 2066, 1479, 1376, 564, 223]. Incrementality [2052]. Increments [978]. Independence [2005, 1172]. Independent [593, 1431, 332, 2568, 1320, 547, 601, 1934, 592]. Independently [1243]. Index [1983, 2462, 33, 79, 123, 161, 202, 229, 257, 270, 299, 388, 417, 471, 486, 518, 693, 751, 768, 803, 846, 874, 936, 966, 1027, 1061, 1083, 1099, 1142, 1214, 1250, 1292, 1367, 1378, 1402, 1533, 1562, 1589, 1608, 1647, 1663, 1688, 1710, 1737, 1822, 1880, 1914, 1956, 1997, 2057, 2113, 2159, 2216, 2223, 2256, 2296, 2332, 2374, 2396, 2411, 2515, 2539, 2555, 2586, 2587, 2619, 2656, 2672, 2696, 1276, 1662, 1835, 126, 255, 150]. Indexed [2594]. Indexical [56]. Indexing [2610, 1806, 1949, 659, 2613, 256, 155, 150, 129]. Indices [139]. Indirect [1460]. Indiscernibility [1210]. Individual [1784, 111, 1821]. Induction [2059, 2592, 2063, 2594, 1246, 1694, 278, 2335, 1131, 2268, 1704, 2095, 2150, 1708]. Induction-Oriented [1246]. Induction-Recursion [2594]. Inductive [967, 34, 1698, 1141, 2261, 2097]. Industrial [355]. Industry [2654, 1097, 763]. Inference [593, 967, 2304, 1066, 2131, 275, 2259, 1920, 556]. Inferences [52]. Inferring [2316, 394, 2322]. Infinitary [1127]. Infinite [2290, 1813, 903, 1636, 991, 923, 1644]. Infinite-State [1644]. Infinitely [888]. Infinity [511]. Inflationary [1119]. Inflectional [2003]. Influence [2005, 1156, 2026]. Information [1928, 2238, 1619, 2579, 1210, 2604, 2258, 2565, 67, 1540, 2248, 2576, 2364, 2577, 896, 1943, 2063, 177, 10, 2391, 2611, 127, 1211, 2585, 2222, 1144, 1560, 2568, 871, 114, 217, 2131, 2423, 537, 2315, 634, 689, 1941, 1541, 1544, 530, 16, 2556, 76, 940, 2001, 1198, 2018, 1673, 155, 404, 1970, 762, 98, 2520, 53, 941, 2255, 1951, 1939, 2140, 2422, 995, 924, 25, 1765, 22, 2234, 2235, 2562, 1819, 2016, 1079, 2237, 1542, 2231, 2421, 452, 1950, 281, 119, 130, 61, 306, 2659, 14, 1182, 179, 186, 128, 2438]. Information-Carrying [177]. Information-Statistical [281]. Information-Theoretic [1819]. Infrastructure [2224, 2205, 122, 2195, 1413, 1477]. Infrastructures [1975, 2581]. Inheritance [5, 1257, 2004]. Inhibitory [633]. Initial [1359, 1531, 631, 107]. Initialization [338]. Initiative [1954]. Injection [1343, 2666, 2686, 2687]. Innateness [1779]. Inner [543]. Inner-Product [543]. Innocence [1113]. Innovation [92, 763]. Input [1618, 1326, 1472, 575, 536]. Input/Output [1618]. Inputs [672]. Insect [1808]. Insecurity [1285]. Inspection [1315]. Inspired [1808]. Installation [2578]. Instance [2062, 1592, 684]. Instance-Based [684]. Instances [985, 970]. Instantiation [2407]. Instantiators [2277]. Instruction [1458, 1499, 1322, 1463, 1323, 1437, 1464, 469]. Instruction-Level [1458]. Instrument [92, 80, 121]. Integer [892, 400, 1388, 1282]. Integral [2504, 1291]. Integrate [713, 1960]. Integrated [2441, 1669, 708, 2391, 603, 2193, 1317, 1937, 721, 2443, 2388, 135, 11]. Integrating [1566, 2192, 2632, 141, 2236]. Integration [2565, 2130, 2423, 2477, 1509, 2556, 2554, 2138, 1361, 2226, 2260, 138, 154, 176, 22, 2234, 2253, 2469]. Integrity [2247, 1735, 2432, 1559, 153]. Intel [1423]. Intellectual [943, 944, 942]. Intelligence [1794, 104, 2077, 529]. Intelligent [2323, 2579, 1969, 2416, 1959, 53, 2320, 59]. IntelliMedia [1660]. Intel(R) [1241]. Intended [39]. Intensive [699]. Intentions [1192, 1151]. Inter [2248, 1981]. Inter-enterprise [2248]. Inter-technology [1981]. Interacting [106, 753]. Interaction [2214, 2227, 1803, 72, 758, 112, 1413, 2503, 2045, 664, 85]. Interactions [1974, 2243, 1958, 2682, 553]. Interactive [775, 2481, 2358, 1579, 2556, 111, 1728, 2195, 82, 112, 2683, 2016, 1934, 19, 1790, 2549, 1547, 1060]. Interchange [26]. Interclass [382]. Interconnected [2173]. Interconnection [1491, 2242, 1489, 1246]. Interdependence [1681]. Interdisciplinary [115, 1536]. Interesting [2125]. Interestingness [2126, 2134]. Interface [23, 2652, 451, 2044, 705, 2177]. Interfaces [2056, 1417, 108, 58, 2659]. Interference [1753, 420]. Interferometric [828]. Interior [250]. Interleaving [2653]. Interlinked [1921]. Intermediary [1624]. Internal [1883, 1797, 1089, 998]. International [1908]. Internet [2143, 2215, 2179, 2248, 2213, 1068, 249, 1722, 1814, 185, 761, 2027, 2574, 1973, 2197, 2181, 2194, 1683, 1081]. Interoperability [1971, 1930, 1585, 61, 1549]. Interpolants [733]. Interpolating [837]. Interpolation [359, 2597]. Interpretation [412, 2001, 411, 1657, 1920]. Interpretation-Based [412]. Interpreters [1461]. Interpreting [2367]. Interprocedural [404]. interrupted [371]. Intersection [1026, 1128]. Intersections [1596]. Interstate [115]. Interval [501, 254, 256, 1121, 2105]. Intervals [1167]. Intonation [2035]. Intra [2163]. Intra-function [2163]. Intrinsically [1007]. Introducing [146, 546, 2574]. Introduction [390, 2412, 1479, 880, 881, 2160, 427, 1406, 2170]. intrusive [614]. Invariant [818, 824, 2080, 524, 322]. Inverse [340]. Inversion [1393]. Invertible [983]. Investigating [2249]. Investigation [1164, 2686, 426]. Investigations [272]. Invocation [1077]. Invocations [446].

Involving [2683]. IO [1457, 715]. Ion [1456]. IP [1686, 2192, 1680, 1352, 2214, 2199, 2171, 2209, 1720, 2188, 2212, 2195, 2180, 2176, 1727, 1337, 2175, 1668]. IP-Based [2180]. IP-Cores [1352]. IP-over-WDM [2171]. IPv4 [1731]. IPv4/v6 [1731]. IR [2099]. Iris [551]. Irrational [1567]. Irrationalities [321]. Irreducible [1434]. Irregular [735, 308, 355]. ISA [440]. Isabelle [1569, 1578, 1580]. Isabelle/HOL [1578, 1580]. Isabelle/Isar [1569]. Isar [1569]. Isolated [1514]. Isomorphic [2450]. Isomorphism [1015]. Issue [20, 1463]. Issues [1686, 1676, 178, 93, 1952, 2043, 2031, 2197, 1712]. iState [2647]. Itanium^{T M} [1241]. Iterate [2599]. Iterated [1186]. Iteration [1150, 1149]. Iterative [1947, 2068, 2526, 2124, 2492, 1439, 2468, 467].

Jacobi [1912]. Jakarta [1063]. JaMake [2482]. Java [2465, 2485, 2486, 1506, 433, 2482, 1065, 1066, 1069, 1070, 1071, 2671, 1613, 1074, 868, 1077, 1441, 2600, 2410, 848, 1305, 455]. Java-Programs [433]. JavaCard [1063, 1563, 1075]. Java^{T M} [849]. JBits [1360, 1328]. JBits^{T M} [1309]. JCCM [1065]. Jini [1430, 455]. Jobs [1510]. Join [1557, 2277, 1443]. Join-Ordering [2277]. Joins [241]. Joint [334, 1564, 653]. Jointrees [1159]. JPEG [1337]. Judgment [1199, 211]. Jump [533]. Junction [356]. Junctures [2551]. Juntas [516]. Just [2410, 1]. JVM [2455].

K2 [1164]. Kappa [1202]. KASUMI [168]. KBs [219]. Kernel [708, 177, 714, 475, 2327, 1090, 614, 1505, 566, 570, 559, 565, 572]. Kernels [1068, 1499]. Key [793, 2690, 199, 792, 1051, 985, 169, 1059, 1055, 947, 167, 938, 955, 131, 1082, 183, 171, 164]. Key-Based [938]. Keyboard [607]. Keyless [944]. Keynote [769, 2377]. Keys [778]. Keywords [128]. Kinematic [645].

Kinematics [2439]. Kinetic [2496]. Kinetically [1326]. Kinetics [2476, 2480]. Kleptographic [1900]. Knapsack [1025, 1482]. Knife [769]. Know [1755]. Knowledge [956, 1998, 2428, 2048, 2361, 2337, 68, 1184, 977, 2117, 1060, 2384, 2310, 73, 1211, 692, 2381, 982, 2389, 2369, 1093, 2378, 1442, 2386, 2046, 2330, 2283, 185, 1176, 1832, 207, 2575, 2393, 2268, 77, 78, 1542, 2319, 2273, 119, 2322, 2098, 2379, 2150, 2151, 59, 60, 2269, 2111, 2157, 63, 2128]. Knowledge-Based [2337, 185]. Knowledge-Bases [1184]. Known [1285]. Kodama [29]. Kolmogorov [2431, 2432]. KQML [755]. Kripke [1120]. Kronecker [1991]. KSIX [723]. Kullback [537, 2135].

L*MASS [2357]. Label [1175, 2117]. Labeled [682, 551]. Labeling [1874, 1875, 546, 821, 832]. Labelled [1121]. Labelling [2316]. Labels [1627]. Laboratory [272]. Labour [1773]. Lack [782, 781]. Lagrange [2505, 1370]. Lambda [1127]. Lambda-Calculi [1127]. LAMGAC [717]. LAN [717]. Land [326, 2546]. Landforms [2541]. Landmarking [349]. Landscape [1747]. Landscaping [476]. Language [950, 435, 2357, 1756, 228, 1804, 2000, 1989, 2628, 2626, 74, 2457, 689, 1941, 51, 2642, 1825, 2086, 1951, 2144, 2051, 927, 2236, 47, 203, 409, 1234, 453, 2040, 2646, 686, 755, 231, 2410, 1639, 14, 2029, 1810]. Language-Based [2086, 409]. Languages [1474, 398, 1625, 42, 2664, 984, 1143, 986, 989, 922, 416, 923, 515, 2003, 2460, 935]. Large [1837, 2487, 2310, 1004, 2465, 152, 1232, 250, 2485, 2184, 2409, 219, 2494, 1917, 1512, 1907, 852, 2011, 2029]. Large-Scale [2465, 2485, 219, 2494, 852]. Laser [353]. Last [2466]. Last-Passage [2466]. Latency [2522, 1314, 1417, 1464]. Latency-Throughput [1417]. Latent [1806]. Lateral [678]. Lattice [258, 1277, 1302, 1638, 1283, 1287, 882, 1289].

Lattices [2137, 1286, 1288, 1290, 223]. Laws [1270]. Layer [313, 629, 2414, 548]. Layered [2114, 1876, 1726, 22, 2197]. Layering [704]. Laying [1373]. Layouts [2533]. Lazy [2059, 2449, 2459]. LC [280]. Leader [2523, 2531]. Leads [667]. League [1369]. Leakage [1079]. Learn [531, 91, 2377]. Learned [2385, 1536, 1809]. Learner [2341, 2066]. Learners [2081]. Learning [1162, 2115, 682, 2059, 100, 1690, 2326, 279, 2062, 1936, 2362, 2064, 2065, 675, 2108, 2154, 289, 2066, 44, 2068, 2375, 2342, 277, 1771, 573, 648, 567, 525, 383, 585, 2376, 1693, 531, 2075, 2076, 102, 2129, 91, 1784, 49, 2080, 613, 283, 325, 295, 1700, 288, 1798, 2082, 638, 763, 551, 287, 557, 2085, 579, 1703, 1764, 560, 2558, 2091, 1705, 2380, 673, 426, 2301, 564, 814, 524, 2451, 553, 2319, 2382, 2333, 290, 2388, 830, 807, 565, 545, 2269, 2104, 2111, 2157, 639, 2339, 595, 1709, 206, 1163]. Least [600, 930, 572]. Least-Squares [600]. Leave [565]. Leave-One-Out [565]. Legacy [1960]. Legal [2355, 2016]. Leibler [537, 2135]. Lemmas [1988]. Length [527, 1371, 362]. Less [1844, 1082]. Lessons [2021, 2385, 1536]. Letopi [1950]. Leukaemia [73]. Level [2463, 1458, 2453, 2354, 1456, 1347, 1509, 2472, 1713, 2051, 1505, 1234, 1328, 864, 705, 1342]. Level- [1456]. Levels [2623, 2564, 2435]. Leverage [122]. Lexical [205]. LHC [1456]. Libraries [1933, 1936, 1931, 1930, 2617, 1940, 1952, 1832]. Library [1927, 1923, 1945, 1415, 1947, 1575, 1813, 1916, 1955, 1344, 750, 1505, 1926]. Lie [801]. Life [1353, 1751, 1824, 1812, 1129]. Lifelike [348]. Lifetimes [1472]. Lifting [307]. Lifting-Based [307]. Light [1831, 1588]. Lightweight [1670, 1520, 2153]. Like [1212, 1285, 1755]. Likelihood [806, 815, 820]. Limited [2309, 1825]. Limiting [1127]. LIN3 [487]. Line [2473, 904, 1921, 1942, 915, 2658, 683, 560, 1366, 583]. Linear [2588, 346, 895, 1861, 1034, 1854, 293, 1103, 2478, 191, 575, 1486, 1008, 996, 821, 1013, 589, 376, 629, 884, 1249, 597, 2501]. Lines [2407, 2544, 2552]. Linguistic [2337, 1779]. Link [144]. Linked [2536]. Linked-Lists [2536]. Linking [1928, 1923, 1925, 1387, 63]. LINUX [470, 725, 1361, 698, 1507]. Lip [338]. Liquid [2502, 343]. List [1470, 1427]. Lists [1607, 2536]. Literary [1924]. Literature [106]. Little [2410]. Lived [2524]. Liveness [1622]. Lives [1821]. Living [1742]. Lixto [2258]. LLL [1283, 1284]. LLL-Reduction [1283, 1284]. LMS [1330]. Load [1425, 1497, 1429, 1430, 1435, 1479, 2492, 443, 1670, 718]. Loads [2509, 1437]. Local [493, 2017, 1372, 289, 1374, 1370, 2500, 380, 1373, 325, 842, 552, 342, 1635, 1100, 2230, 1376, 55, 564, 2348, 883, 356, 999, 1163]. Local-Ratio [493]. Locality [1236, 1107, 1440]. Localization [677]. Location [496, 1092, 181, 502, 235, 242]. Location-Dependent [235]. Locked [671]. Locomotion [1795]. Locus [1102]. Log [1550, 344, 1380]. Log-Only [1550]. Log-Polar [344]. Logarithmic [1863]. Logarithms [201]. Logging [1522]. Logic [1127, 1120, 1209, 398, 2270, 1690, 1114, 1168, 2267, 2265, 1339, 1641, 2284, 1208, 208, 1119, 2289, 2257, 2286, 1577, 2342, 2274, 1102, 1580, 1190, 1103, 1697, 1698, 1797, 210, 1118, 2001, 2688, 1643, 1361, 2346, 397, 2281, 1117, 2264, 1583, 2282, 1301, 394, 1105, 1213, 30, 448, 1128, 2261, 2451, 1145, 2599, 1920, 1565, 1337, 1115, 480, 2276, 2218, 2097, 2262, 2266, 884, 211]. Logic-Based [398, 1920]. Logical [1179, 2573, 1475, 34, 1211, 216, 2421, 230, 1729, 1014, 2595]. Logics [2300, 1121, 2269, 1122]. Logit [628]. LogP [1426]. Logs [1695, 1545]. Logspace [918]. Long [2534, 2524, 262, 2495, 1683]. Long-Lived [2524]. Long-Term [2495, 1683]. Look [1112]. Looking [777, 72, 2203]. Loop [1333, 1436, 763, 804].

Loop-Carried [1436]. Loops [2387, 1767, 1434]. Lossless [310, 1357, 311, 307, 1363]. Lossy [1860, 307]. Lossy-to-Lossless [307]. Lotos [1228]. Low [1278, 2522, 1895, 2038, 749, 314, 269]. Low-Cost [749]. Low-Latency [2522]. Lower [1171, 893, 904, 2525, 1868]. Lowercase [289]. LSA [2099]. LSO [2395]. LSTM [611, 613]. LTL [1990, 1239, 2271]. Lumen [332]. LUT [1296, 1327]. LUT-Based [1296, 1327]. LVQ [528, 612].

m [1672, 2605, 742]. M-CODD [2605]. M-SPH [742]. MAC [163, 1904]. Machine [94, 1745, 1561, 509, 1003, 904, 1461, 711, 2682, 2255, 564, 1011, 576, 2111, 2157, 2339, 849]. Machine-Generated [2255]. Machines [562, 1851, 1769, 889, 732, 1346, 573, 988, 2638, 991, 563, 574, 1115, 572, 631, 88, 695]. Macro [1466, 1405]. Macro- [1405]. Macrocell [1298]. Made [1599, 2262]. Magnetic [332, 1530, 2502]. Maintaining [698]. Maintenance [142, 2393, 2695]. Major [1952]. Make [1738, 1711]. Makespan [264]. Making [780, 779, 1780, 1143, 345, 221]. Malicious [1973]. Malleable [1431, 1847]. Mammograms [596]. Mammography [656]. Man [94]. Manage [254]. Manageable [698]. Management [1978, 2424, 723, 2224, 2390, 1311, 1524, 2210, 1470, 2384, 1724, 2381, 2389, 1946, 2416, 2561, 1967, 1968, 2330, 1550, 2663, 2581, 47, 1536, 1966, 2331, 1800, 1729, 2045, 135, 1962, 787]. Manager [450, 2054]. Manager-Worker [450]. Managing [15, 2211, 1975, 2379]. Maneuvering [2473]. Manifolds [293, 577]. Manipulation [2224, 1551, 1655, 1658]. Manipulations [400]. Manipulators [1256, 645]. Manner [621]. Manufacturing [1264]. Many [1212, 888, 900, 214, 2281, 1855, 1494, 733, 1290]. Many-to-One [1494]. Many-Valued [1212, 214, 2281]. Map [1381, 582, 588, 249, 585, 1377, 662, 2369, 578, 676, 583, 674, 808]. Maple [1566, 1531]. Mapping [1296, 2623, 1352, 90, 159, 578, 2545, 743, 1382, 1766, 430, 540, 1559, 153]. Mappings [1922, 267]. Maps [1599, 580, 1498, 278, 2141, 681, 608, 678]. MARIAN [1930]. Market [534, 759, 638, 2359, 625]. Market-Based [638]. Marketplace [2248]. Marketplaces [2575, 2244]. Markets [10, 2249, 627, 760]. Marking [1678, 1677]. Markov [2114, 836, 809, 2532, 731, 810, 1004, 1614, 1384, 823, 365, 1139, 1987, 2468, 1301, 539, 2340, 533, 679, 812]. Markovian [599]. Marks [800, 799]. MAS [752]. Masking [1882]. Massively [422, 602]. Master [1508]. Master-Worker [1508]. Match [124]. Matching [1313, 833, 317, 981, 838, 1085, 380, 736, 28, 834, 1855, 811, 1764, 831, 137, 387, 2150, 1386, 221]. Matchings [1590]. Matchmaking [2230]. MatchPRO [1357]. Mate [1401, 1777]. Mate-Pairs [1401]. Material [353]. Math [1568]. Math-Web [1568]. Mathematica [2443]. Mathematical [1207, 2489, 369, 1253, 1254, 1265, 2217, 1012, 215, 674]. Mathematics [268]. Matrices [1170, 27]. Matrix [1303, 1086, 1484, 814, 1010, 522, 734, 540, 129]. Matrix-Vector [522]. Matroid [1026]. Matroid-Greedoid [1026]. Matter [1714]. MAWC [2601]. Max [1993, 900, 2191, 1085, 602, 487, 503, 2348]. Max-flow [827]. Max-Min [2191]. Max-Plus [1993]. Max-Shift [1085]. Maximal [1620, 831, 629]. Maximising [815, 2093]. Maximization [689]. Maximizing [845]. Maximum [806, 1017, 1590, 1390, 805, 1026, 504, 820, 505, 431]. Mbytes [1357]. Mbytes/s [1357]. MC [1394]. MCMC [817]. MDG [1235]. MDS [305]. Meaning [94, 109, 1778, 97]. Means [2667, 1447, 102, 2557, 53, 350, 583, 2496, 2438]. Measure [976, 627, 2134, 2086, 2533, 1789, 637]. Measure-Once [976]. Measured [2144]. Measurement [2182, 2392, 351, 2203, 2387, 2181, 1600]. Measurement-Based [2182]. Measurement-Driven [1600]. Measurements [2206, 1593]. Measures [1203, 2430, 2126, 662, 1816, 141, 1078, 1330]. Measuring [2012]. Mechanical [1570, 1264]. Mechanising [1064]. Mechanism [228, 2199, 1731, 542, 2580, 1513, 652, 721, 1716, 63, 2436]. Mechanisms [1667, 434, 1097, 2684, 1725, 2219]. Mechanistic [2379]. Media [1738, 2198, 2506, 66]. Median [1397, 897, 1393, 368]. Mediated [113, 762, 1951]. Mediation [1557]. Mediator [1540, 85]. Medical [2604, 2673, 2032, 138, 464, 329, 298, 311, 328, 333, 273, 2150]. Meditations [94]. Mediterranean [2491]. Medium [84]. Meeting [2232]. Meets [1566]. MEG [609]. MELL [1104]. Melodies [1541]. Membership [984, 1511, 1859, 179]. Membranes [1756]. Memories [1244, 635]. Memory [1836, 1094, 419, 732, 86, 2309, 725, 1636, 713, 1447, 585, 610, 1300, 1797, 1298, 676, 467, 1089, 860, 1465, 540, 1346]. Memoryless [1987]. MemTo [725]. Mental [65, 106, 640]. Mercury [2491]. Merge [1944]. Mergers [771, 770]. Merging [1184, 549, 1944, 1183]. Merit [1546]. Mesh [443, 1494, 1512, 2503]. Mesh-Connected [1512]. Meshes [340, 247, 843, 374]. Meso [2488]. Meso-scale [2488]. Message [945, 1522, 2633, 713, 2529, 721, 1505, 743]. Message-Passing [2529, 743]. Messages [1527]. Messaging [1505]. Meta [35, 2595, 2640, 158, 715, 2128, 159, 1817, 1368, 2651]. Meta-computing [715]. Meta-data [158, 2128]. Meta-game [1817]. Meta-heuristics [1368]. Meta-knowledge [2128]. Meta-logical [2595]. Meta-model [2640, 2651]. Meta-search [159]. Metabolic [1758]. Metacomputing [436, 1524, 1515]. Metadata [1929]. Metal [1752]. Metamodel [2623]. Metamodeling [2622, 220, 25]. Metaphor [108]. Metaphorical [35]. MetaPL [425]. Metaprogramming [2667, 2409]. Metasystem [1414]. Metering [943]. Method [1296, 2509, 2463, 36, 2504, 460, 461, 828, 549, 1882, 1980, 647, 462, 294, 1942, 2164, 1556, 2477, 2617, 1204, 684, 443, 335, 369, 51, 538, 591, 1355, 2425, 1807, 2507, 2020, 2558, 2009, 333, 844, 1324, 574, 343, 1668, 1014, 1015, 2153]. Methodological [2181]. Methodologies [2658]. Methodology [1972, 187, 1558, 2574, 2435, 2380, 1241, 1915]. Methods [841, 2591, 456, 1474, 2511, 1238, 605, 340, 2498, 1374, 1888, 2012, 213, 2467, 2277, 2468, 1286, 1395, 342, 560, 446, 2336, 1553, 882, 730, 608, 817, 223, 586, 595, 1632]. Metric [1164, 807]. Metrics [2100]. Metro [2173]. Metropolitan [1715]. MIC [2470]. Micro [787, 1405]. Micro-management [787]. Micro-parallelism [1405]. Microcellular [2499]. Micromobility [2175]. Microprocessor [1247]. Microprocessors [541]. Middleware [723, 1724, 1723, 1725, 852]. MIDI [1595]. Migration [1095, 719]. MIMD [1496]. Mimicking [669]. Min [1166, 827, 2191, 602, 1016]. Min-Based [1166]. Min-cut [827]. Min-cut/Max-flow [827]. Min-Max [602]. Mind [1814, 108, 99, 1814, 1820]. Mind-Prosthesis [108]. Minimal [1130, 2280, 809, 1844, 1032, 910, 911, 2259, 1728, 844]. Minimalist [1799]. Minimisation [1879]. Minimising [1426]. Minimization [827, 828, 512, 906, 2121, 842, 2038, 829, 653]. Minimize [804, 2106]. Minimizing [492, 2179, 840, 1022, 501, 821]. Minimum [979, 1872, 2644, 1848, 2529, 533, 1863]. Minimum-Entropy [533]. Minimum-Time [1872]. Mining [878, 2464, 2124, 1446, 2132, 1604, 1442, 2386, 2109, 2155, 2089, 282, 272, 2602, 27, 281, 251, 2151, 2099, 2110, 2156, 271, 2383, 135, 133, 1547, 134]. Mirror [290, 690]. Mission [2550]. MISTY [168]. MISTY-Type [168]. MIX [946]. MIX-Net [946]. Mixed [160, 326, 313, 2471, 1373, 1024, 916, 2506]. Mixed-Hybrid [2506]. Mixin [2408]. Mixin-Based [2408]. Mixture [596, 603, 525, 2083, 544]. Mixtures [320, 597, 1157, 592, 589]. Mizar [1588]. ML [1394]. MLS [175]. MML [2621]. MMX [1423]. Mobile [1974, 1978, 1616, 160, 1972, 1123, 2535, 1096, 2527, 412, 2585, 1731, 1963, 1969, 2635, 2229, 642, 1980, 1981, 1797, 2613, 1968, 198, 411, 1959, 1977, 236, 2582, 233, 1960, 1975, 1965, 2231, 428, 1730, 1966, 1979, 1962, 767]. Mobile-Agent [2229]. Mobility [1615, 437, 2221, 1617, 2188]. Möbius [1996]. Modal [1120, 1210, 317, 1119, 1637, 1071, 1117, 58]. Mode [1872, 1801]. Model [649, 669, 1743, 2437, 1618, 2242, 1676, 419, 1217, 1153, 2224, 1168, 809, 1990, 32, 1781, 460, 1123, 1638, 1002, 1877, 670, 694, 675, 332, 1036, 1747, 571, 1173, 1239, 2327, 2514, 2490, 408, 1637, 713, 75, 871, 2664, 1613, 2271, 1614, 2128, 610, 423, 628, 1976, 1655, 293, 2369, 1384, 2315, 626, 2491, 823, 1941, 652, 365, 1985, 666, 2259, 1785, 2387, 1987, 1490, 1161, 729, 2335, 1231, 1301, 1059, 1379, 676, 1219, 54, 756, 2422, 236, 2494, 2427, 329, 1555, 2663, 1859, 4, 1227, 1554, 366, 743, 2652, 679, 677, 428, 1520]. Model [27, 1218, 1984, 314, 1777, 140, 544, 949, 599, 2151, 231, 2023, 851, 2661, 286, 688, 2496, 680, 674, 767, 1222, 812, 2251, 753, 2640, 2651]. Model-Based [2327, 236, 329]. Model-Check [1613]. Model-Checking [1638, 1239, 408]. Model-Theoretic [1168]. Modeling [2487, 1642, 2428, 187, 1996, 34, 43, 1865, 2595, 2553, 29, 48, 334, 1175, 2574, 2575, 54, 527, 1977, 2235, 2681, 2181, 2098, 2496, 63, 625, 1810]. Modelling [13, 2691, 1181, 2635, 476, 152, 982, 2491, 1763, 2493, 2502, 311, 1776, 226, 1363, 2646, 2004, 14, 1708, 1989]. Models [1130, 2114, 2627, 1491, 2280, 1474, 1744, 65, 2290, 95, 596, 561, 1160, 2392, 624, 2342, 603, 907, 2204, 525, 476, 2271, 1991, 1143, 2686, 2492, 1938, 2259, 1439, 2024, 51, 1316, 2138, 2643, 552, 377, 557, 112, 811, 1951, 1954, 2340, 1704, 2291, 2051, 2092, 2217, 2219, 826, 1920, 1753, 1012, 2321, 2630, 2402, 1477, 1152, 12, 2103, 2549, 2659, 348, 11, 129, 2469, 1650]. Modern [1461, 541]. Modes [189]. MoDeST [1989]. Modified [467, 368]. Modular [602, 683, 1286, 1897]. Module [2026, 410]. Modules [1319]. Moduli [975]. Modulus [798]. Moiré [816]. MOLEN [1321]. MOM [1069]. Moments [535, 324]. Monadic [1103, 2459]. Money [797, 1082]. Monitor [1421]. Monitoring [1523, 616, 2205, 725, 2643, 614, 2608, 722, 720, 595]. Monocular [677]. Monoid [479]. Monoids [989, 918, 921]. Monotone [1141, 831]. Monotonic [2288, 1208, 2279]. Monotonically [929]. Monte [2473, 2475, 2334, 2466, 2467, 2477, 2479, 538, 466, 539, 996, 533, 2333]. Monte-Carlo [996]. Montgomery [1910, 1892]. Montgomery-Form [1892]. Morphisms [482]. Morphogenesis [1763]. Morphogenetic [2439]. Morphological [2006, 359, 1336, 278, 2003, 2010]. Morphology [1207, 1780, 369]. Mosaic [359]. Most [2300]. Motifs [370]. Motion [340, 304, 1259, 330, 2436, 2602, 343, 344]. Motivation [1781, 1332]. Motivational [64]. Movement [906]. Movements [686]. Movies [2616]. Moving [806, 237, 2239, 234, 230, 246, 2150, 231]. MPEG [2215, 305, 300, 2196, 302, 304, 313, 2195, 301, 1733, 2616, 303]. MPEG-21 [1733]. MPEG-4 [2215, 2195]. MPEG-7 [305, 300, 302, 304, 301, 303]. MPI [1457, 2510, 1415, 708, 715, 701, 696, 704, 1528, 710, 727, 750, 739, 700, 705]. MPI-2 [701]. MPI-I [1528]. MPI-I/O [1528]. MPI Connect [715]. MPLS [2172, 1684, 1674]. MPP [458]. MRI [336, 656]. Multi [2414, 1978, 2323, 2198, 2357, 1761, 2227, 66, 2228, 1346, 2117, 2567, 1069, 624, 714, 1032, 1748, 1969, 1306, 1033, 1946, 2578, 2164, 1853, 548, 2283, 758, 2574, 757, 2282, 1799, 2566, 2145, 298, 1957, 2616, 2183, 1718, 1950, 2149, 1801, 58, 2661, 653, 639, 248, 241, 625]. Multi-adjoint [2282]. Multi-agent [2323, 2357, 2227, 2228, 2567, 624, 1748, 1969, 2578, 2574, 639, 625]. Multi-Agents [1957]. Multi-applications [1069]. Multi-cell [1801]. Multi-class [2183]. Multi-criteria [2574]. Multi-dimensional [2283]. Multi-domain [1978]. Multi-expert [2616]. Multi-FPGA-Prototype [1306]. Multi-label [2117]. Multi-layer [2414, 548]. Multi-media [66]. Multi-memory-FPGA-Based [1346]. Multi-modal [58]. Multi-Object [757, 1799]. Multi-objective [2149]. Multi-operator [2164]. Multi-paradigm [2661]. Multi-party [1032, 1033]. Multi-Personal-Agent [758]. Multi-presentation [1946]. Multi-resolution [1718]. Multi-scale [248]. Multi-script [1950]. Multi-segmented [298]. Multi-site [2566]. Multi-stage [639]. Multi-state [2198, 1761]. Multi-threaded [714]. Multi-valued [2145]. Multi-vehicle [1853]. Multi-way [241]. Multiagent [2355, 2356, 2416, 1772]. Multicast [2192, 1734, 2193, 2189, 2191, 2190, 2202, 1726, 1871]. Multicasting [1018, 1735]. Multichannel [2032]. Multiclass [563]. Multiclock [1224]. Multicommodity [492, 1868]. Multicomputers [1512]. Multicriteria [1147]. Multidatabase [2246, 2247]. Multidimensional [13, 1560, 1556, 369, 140]. Multigrid [2498, 2500]. Multilevel [2622, 496, 2467, 739]. Multilingual [1542]. Multimedia [2245, 2609, 124, 2605, 127, 1500, 1723, 1722, 1949, 2606, 1732, 1728, 2607, 1711, 1735, 1716, 1733, 1465, 1665, 541, 2194, 1495, 1651]. Multimodal [1654, 1648, 1652, 1660, 1661, 1659, 1657, 1653, 1658]. Multinomial [628]. Multipartitioning [1438]. Multiple [1225, 1313, 17, 1638, 2062, 840, 20, 310, 810, 1870, 2612, 2138, 2564, 1727, 2570, 2232, 1539, 2097, 2152, 2549, 381, 348, 1924]. Multiple-Resolution [2612]. Multiple-Valued [1638]. Multiplication [1086, 1038, 1910, 1892, 734, 1351, 1349]. Multiplicity [2644]. Multiplier [1898, 1895, 1897]. Multipliers [2505, 1370, 1350]. Multipliers/Fictitious [2505]. Multipoint [1963]. Multipoint-to-Point [1963]. Multiprocessor [419, 1022, 1501]. Multirelational [1700]. Multiresolution [247]. Multiscale [361, 2150]. Multisensor [1173]. Multisequence [1291]. Multiset [2458]. Multispectral [826]. Multithreaded [1424, 389]. Multithreading [1462]. Multiuser [2243]. Murecine [2603]. Museum [1946]. Music [32, 1541, 1271, 1948, 1331, 2104]. Musical [2111, 2157]. Mutation [1805, 2686]. Mutual [2528, 2517, 530, 2516]. MXML [1560]. Mystification [2110, 2156]. Myth [1687]. Mythical [1813]. Myths [441].

Naive [593, 2105, 2107]. Naive-Bayes [2105]. Nakagami [286]. Name [1617]. Names [1926]. Naming [772, 1469]. Narrative [104]. Narratives [104]. Narrowing [1865, 2449]. Natural [950, 81, 74, 667, 1825, 2597, 2144, 1121, 47, 1788, 1109, 686]. Natural-Born [81]. Natural-Language [74]. Nature [2577]. Navier [2498]. Navigation [2326, 146, 1377, 1122]. nBrains [1792]. NC [898]. Near [506, 310]. Near-Lossless [310]. Near-optimum [506]. Nearest [259, 243, 1603, 234]. Necessary [757]. Need [2110, 2156]. Needed [409]. Needs [2260]. Negation [2263]. Negative [2151]. Negotiation [2579, 2573, 760, 1718, 2560]. Negotiation-Credit [760]. Negotiations [20, 19, 2232]. Neighbor [259, 243, 1603, 234]. Neighborhood [1469, 586]. Neocognitron [654]. Nepal [1478]. Nervous [1740]. Nested [1572, 1478, 2247]. Net [481, 472, 946, 475, 1239, 473, 628, 287, 478, 2331, 644, 429, 474, 147]. Nets [1642, 484, 483, 2440, 477, 485, 439, 633, 1633, 2435, 479, 630, 480, 686, 474]. Network [492, 435, 1680, 2173, 1319, 550, 1684, 1430, 187, 670, 2391, 2206, 691, 711, 1795, 2071, 2204, 2200, 2676, 2203, 741, 1503, 672, 1719, 1740, 617, 1967, 597, 558, 658, 1316, 1678, 546, 1591, 676, 1527, 1939, 2675, 291, 1417, 524, 1965, 2207, 606, 452, 541, 1668, 645, 830, 705, 1593, 664, 656, 1962]. Networked [1320]. Networking [1600]. Networks [1162, 2414, 649, 1837, 1491, 432, 1166, 2437, 584, 1429, 2534, 1690, 2174, 1489, 1675, 2338, 420, 2518, 600, 2171, 616, 663, 621, 2535, 1096, 1873, 2210, 2186, 1165, 535, 590, 523, 665, 2538, 692, 542, 1018, 1246, 1748, 2191, 620, 1980, 48, 2529, 647, 537, 189, 619, 1762, 1698, 1967, 684, 1763, 613, 1736, 288, 2330, 2184, 2195, 2180, 602, 683, 563, 352, 1807, 1157, 536, 618, 947, 615, 673, 426, 1158, 2183, 1611, 1681, 1715, 553, 554, 1717, 1677, 1766, 2175, 520, 1730, 632, 949, 2187, 734, 622, 545, 2170, 231, 430, 629, 1687, 2438]. Networks [1163]. Neural [649, 2437, 584, 646, 1690, 600, 616, 663, 621, 590, 665, 1795, 692, 2071, 672, 48, 628, 647, 650, 617, 537, 619, 652, 1762, 684, 613, 597, 558, 287, 602, 546, 683, 563, 352, 2435, 681, 618, 947, 615, 673, 426, 524, 643, 2094, 554, 520, 640, 644, 645, 686, 632, 830, 622, 545, 664, 286, 668, 656, 680, 2438]. Neural-Net [644]. Neural-Network [558]. Neural-Network-Based [664]. Neuro [101]. Neuro-Psycho-Computational [101]. NeuroCluster [434]. Neurocomputing [2337]. Neurocontrollers [1768]. Neuromodulations [1768]. Neuron [669, 2338, 672, 2439, 679, 690]. Neurons [569, 670, 573, 553]. Neutral [1766]. Neutron [740]. News [917, 1934]. Newton [591]. Next [1881, 796, 695, 1316, 1954, 2043, 795]. Nilpotent [1006]. NLDA [523]. NLP [2363]. NNs [527]. No [899, 310]. Node [1434]. Noise [603, 575, 2038, 2025, 366, 363, 609]. Noisy [1091, 940]. Nomadic [1964]. Nominal [2221, 2107]. NoMoRe [2288]. Non [2368, 1212, 2288, 1570, 420, 1060, 513, 2122, 2489, 1104, 2536, 880, 2522, 597, 1645, 463, 2501, 198, 710, 172, 1105, 614, 2279, 2141, 1055, 371, 1507, 1621]. Non-approximability [513]. Non-blocking [2536, 2522]. Non-Boolean [513]. Non-clausal [1212]. Non-commutative [1105]. Non-commutativity [2489, 1104]. Non-crisp [2122]. Non-dedicated [1507]. Non-designated [198]. Non-deterministic [172]. Non-Euclidean [2141]. Non-interactive [1060]. Non-interference [420]. Non-interrupted [371]. Non-intrusive [614]. Non-linear [597, 2501]. Non-Monotonic [2288, 2279]. Non-persistent [710]. Non-repudiation [1570, 1645]. Non-sequential [1621]. Non-size-increasing [880]. Non-stationary [463]. Non-traditional [2368]. Nonbinary [2078]. Nonces [1285].

Nondeterministic [909, 721]. Noninvertible [983]. Nonlinear [1738, 646, 1054, 2512, 573, 627, 2471, 288, 568, 2443, 2442, 553, 586, 609]. Nonlinearized [2030]. Nonmonotonic [2261]. Norm [1484]. Normal [1287]. Normalised [1302]. Normalized [1138]. Normative [1204]. Notation [425]. Notations [110, 111]. Note [2590, 911, 916]. Notebooks [2443]. Notes [2221]. Novel [1731, 685, 1487, 1556, 1670, 1735, 319, 379]. Novelty [2083]. NOWs [1492]. NP [876, 970, 928]. NP-Hard [970]. NP-Pairs [928]. NPACI [698]. NTRU [1902]. NUMA [1488]. Number [1034, 1390, 1889, 170, 925, 1890]. Numbers [1575, 1387, 2478, 1111, 929, 1014]. Numeral [290]. Numeric [2443]. Numerical [731, 2514, 2507, 2502, 1483]. NuPRL [1585]. NURBS [1314]. Nyberg [1285]. Nyberg-Rueppel [1285].

O [1525, 1526, 1528, 702]. OAEP [1041, 1044, 1043, 1042]. Obfuscating [1028]. Object [806, 2487, 872, 1561, 855, 1066, 2440, 5, 2633, 127, 665, 2664, 856, 254, 256, 316, 1552, 2636, 757, 843, 1413, 1550, 2, 2460, 372, 2236, 813, 677, 848, 1122, 1443, 1799]. Object-Oriented [2487, 1561, 2440, 5, 2633, 1552, 2, 2460]. Object-Relational [127, 254, 256, 2636]. Objective [529, 2146, 2149]. Objects [1928, 1858, 2473, 1652, 237, 461, 1519, 158, 216, 2612, 1701, 710, 1543, 233, 230, 355, 231]. Obligations [1192]. Oblivious [1036, 1418]. Observability [1342]. Observable [638]. Observations [1761, 2275, 246]. Observed [2103]. Obstacle [645]. Ocean [587]. OCL [2627, 2639, 2629, 2628, 2648]. Ocsid [1582]. ODMG [870]. Off [1605, 1981]. Offer [1780]. Offers [756]. Official [762]. Offline [1874, 1048]. Offs [954, 34, 2309]. Often [1098]. OilEd [2324]. OLAP [252]. Old [95].

Oligonucleotides [1385]. Olympic [1676]. Omega [2276]. Omega-Restricted [2276]. Omnidirectional [2368]. Omniscience [34]. On-Card [1074]. On-Line [2473, 904, 1921, 1942, 915, 683, 560, 583]. Once [976]. One [706, 965, 892, 421, 900, 1590, 701, 913, 23, 1765, 180, 1494, 601, 565]. One-Channel [601]. One-Dimensional [1765]. One-Page [23]. One-Sided [706, 701]. One-Time [180]. One-Way [913]. Online [178, 494, 1045, 495, 32, 561, 1020, 1874, 145, 917, 1841, 618, 1890, 1048, 21, 2438]. Online/Offline [1048]. Only [1051, 1550]. Onset [671]. onto [1352, 430, 540]. Ontological [1945]. Ontologies [2568, 2423, 1922, 2255, 2352]. Ontology [2645, 2579, 2324, 203, 2039]. Open [776, 1642, 2267, 2239, 1073, 2360, 236, 1954, 1733, 949, 405]. Openings [360]. OpenMP [2483]. Operating [1477]. Operation [189, 1358]. Operational [1155, 1070, 1248]. Operations [2601, 1551, 402, 252]. Operator [1689, 1805, 2489, 2164]. Operators [125, 1184, 2165, 1186]. Opportunistic [2579]. Opportunities [2582]. Optical [2171, 2172, 1018, 1967, 842, 342, 1950, 640, 2170]. Optimal [2179, 1873, 1057, 1020, 1526, 2531, 1510, 2313, 1047, 288, 1664, 2520, 1041, 2493, 2298, 1494, 1417, 1393, 1012, 1900, 269]. Optimality [1771, 2570]. Optimally [2148]. Optimised [1089]. Optimising [1415]. Optimization [2303, 1238, 1325, 403, 1739, 2071, 260, 2499, 1886, 2434, 1591, 529, 1537, 739, 2301, 1364, 817, 2614, 344, 430, 405, 848]. Optimize [2365, 687]. Optimized [1322, 2135, 315, 1789, 1434]. Optimizer [1539]. Optimizing [2463, 436, 2278, 1488, 716, 1545, 2235, 2426]. optimum [506]. Options [1869]. Oracle [1067]. Oracles [427]. Order [2302, 1210, 1114, 1573, 2299, 1576, 2066, 1577, 1190, 2449, 1118, 1111, 1046, 278, 1635, 426, 2093, 1893, 1140, 1109, 1565, 196, 491]. Order-Generic [1109]. Ordered [497, 2289, 2072, 1169, 370, 1109]. Ordering [312, 2277]. Orderings [1162, 895, 1163]. Orders [1159, 265]. Ordinal [1576, 2070, 1146]. Organic [2379]. Organisation [68, 1774, 637]. Organisational [2382]. Organisations [2380, 4, 1758]. Organising [1802]. Organisms [1816]. Organization [434, 1762, 1801, 678]. Organizational [2376, 2426]. Organizations [2375, 1403]. Organize [1935]. Organized [584, 636, 2415]. Organizing [2238, 582, 588, 662, 578, 1807, 2141, 1948, 608, 632, 583, 1410]. Orientation [839, 678, 674]. Oriented [2487, 2325, 1561, 1675, 2440, 461, 5, 2633, 127, 1246, 2405, 2127, 2662, 1552, 1916, 207, 2650, 2559, 2, 961, 2460, 1536, 2402, 130, 1358]. Orienteering [1023]. Origin [2490]. Origins [681, 1810]. Orthogonal [1206, 864]. Orthogonalization [1284]. Orthogonally [849, 868]. Oscillator [680]. Oscillators [691]. Oscillatory [672, 658, 343]. Ossa [226]. Other [104, 1285]. Outliers [2083, 387]. Output [1618, 591]. Overhearing [2228]. Overlap [436, 2551]. Overrepresented [1385]. Overview [1277, 869, 850, 866, 865, 304, 473, 859, 854, 847, 1823, 862]. Ownership [956, 666]. Oz [2558].

p [1909, 2238, 876]. P-Grid [2238]. P-NP-Hypotheses [876]. P2P [2238]. Package [1591, 750]. Packet [432, 2171, 2172, 2193, 2220, 1664, 1362, 1677, 2187, 2170]. Packing [497, 1878, 968]. PACX [715]. Padding [1052, 1041]. Page [23, 1594]. Pagers [864]. Pages [2018]. Paging [1843]. Pains [2637]. Paintbrush [817]. Pair [1193]. Pair-Wise [1193]. Pairing [1040]. Pairs [1538, 1401, 928].

Pairwise [1383, 819, 295, 563, 2092]. Palm [200]. Panel [794, 2378, 778]. Panoramas [381]. Paper [937]. PAPI [1505]. Parabolic [2513]. Paradigm [2610, 2633, 1635, 2661]. Paradigms [2404, 1960]. Paragraphs [2017]. Parallel [2588, 492, 1851, 1448, 1497, 434, 435, 1474, 457, 2510, 458, 1525, 732, 460, 461, 1428, 1438, 1484, 1908, 737, 1519, 715, 438, 711, 1455, 1404, 1419, 2465, 439, 1479, 441, 1446, 1447, 422, 423, 442, 1840, 462, 2162, 1487, 745, 1509, 716, 443, 1442, 736, 703, 463, 1895, 1037, 602, 717, 1444, 465, 425, 1451, 466, 727, 746, 748, 2494, 467, 468, 1531, 742, 749, 730, 1454, 1496, 449, 454, 734, 2687, 1338, 1310, 455, 431, 1443]. Parallelisable [1498]. Parallelism [1458, 390, 1478, 2163, 450, 431, 1405]. Parallelization [456, 2464, 750, 2348, 740]. Parallizing [729]. Parameter [892, 2513, 2301]. Parameterization [2026]. Parameterized [1640, 1876, 475, 404, 1231, 1348, 478]. Parameterizing [397]. Parameters [2338, 1734, 335, 286, 2551]. Parametric [407, 2120, 405]. Parametrisation [334]. Parenthesis [515]. Pareto [1771, 1098]. Parity [893, 907]. Parse [1703]. Parser [1353]. Parsing [2363, 2362, 2009, 2008, 2097]. Part [383, 2222, 2476, 2480, 1646]. Part-Based [383]. Partial [1127, 2302, 2118, 737, 477, 662, 1093, 1476, 995, 924, 1635, 265, 931, 1328, 309]. Partiality [38, 1572]. Partially [1169, 638, 1285, 345]. Participants [89]. Particle [339, 703, 343]. Particulate [2505]. Partite [508]. Partition [1106, 857]. Partitioners [1428]. Partitioning [719, 1526, 636, 1354, 1219, 1868, 1327]. Partly [337]. Parts [223]. Party [1037, 1035, 1032, 1033]. Pass [2033, 1233]. Passage [2466]. Passing [1782, 713, 2529, 721, 1505, 743]. Passive [2205]. Password [199]. Passwords [1051].

Past [1339, 694]. Patch [1799]. patent [763]. Path [2303, 905, 1021, 819, 648, 2466, 1854, 1980, 1853, 2533, 1365]. Paths [258, 840, 1872, 262, 1451, 1098, 445]. Pattern [1052, 72, 212, 981, 2123, 633, 49, 2080, 1545, 296, 353, 1386, 1547]. Pattern-Category [49]. Patterns [2690, 953, 2400, 1607, 71, 2072, 2129, 816, 278, 2089, 347, 242, 998, 1777, 577, 251, 2631]. Pay [1711]. Payment [8, 9, 1076, 1082]. PBAC [1671]. PC [736, 749, 470]. PC-Cluster [749]. PC-Clusters [470]. PCA [293]. PCA-Based [293]. pCG [228]. PCP [2199]. PCP-DV [2199]. PCS [2448]. PCTRAN [703]. PDE [1488]. PDES [718, 738, 1487]. Peano [132]. Pedro [1650]. Peircean [203]. Pelvis [334]. Penalization [2346]. Penalty [495]. Pentium(R)4 [1232]. PENTRAN^{T M} [703]. People [96, 1714, 2076, 98, 345]. PEPA [1991]. Peptide [1399]. Perceived [63]. Perception [87, 675, 2021, 63]. Perceptron [69]. Perceptrons [555, 548]. Perceptual [840, 43, 941]. Percolation [457]. PerDiS [852]. Perfect [1844]. Performance [1422, 458, 2185, 2182, 1086, 600, 1424, 384, 1325, 1421, 2512, 1499, 1407, 2067, 1863, 2209, 726, 1404, 2204, 1419, 1418, 2635, 152, 741, 856, 1614, 1488, 759, 1528, 1433, 2329, 716, 703, 697, 849, 24, 2176, 464, 1416, 425, 1887, 1527, 727, 1409, 1453, 1395, 748, 698, 447, 2093, 1505, 1516, 1681, 1465, 449, 1423, 1471, 2104, 521, 2177, 1443]. Performances [555, 1467]. Performing [2566]. Periodic [2213]. Periodical [2090]. Periods [931]. Permutations [1450]. Persistence [853, 2399, 848]. Persistence-Enabled [848]. Persistent [863, 861, 710, 868, 2461, 860, 851, 849]. Person [54]. Person-Centred [54]. Personal [871, 173, 766, 758, 2562]. Persons [1326]. Perspective [38, 1702, 2685, 2398]. Pertinence [2291]. Petascale [695]. Petri [2440, 472, 477, 485, 475, 473, 439, 1633, 479, 478, 480, 474]. PEVM [849]. Pfaffian [266]. pGOLOG [2312]. Phase [1057, 347, 2147, 671]. PHBs [2185]. Phenomena [57]. Phenomenological [1788]. Phenotype [2117, 1766]. Philosophical [60]. Phone [1974]. Phonectic [2042]. Phoneme [2028, 2023]. Phoneme-Threephone [2023]. Photogrammetry [347]. Photonic [1485]. Photos [2369]. Phrase [1917, 131]. Phylogenetic [1395]. Physical [1686, 2272, 2420]. Physics [1270]. Pi [1578, 410]. Pi-Calculus [410]. Picard [2508]. Picard-Uzawa [2508]. Picture [282]. PID [644]. Piece [98]. Piecewise [999]. Pigeonhole [888]. Pilot [2002]. PIM [1406]. Pinpoint [29]. Pipeline [1464, 1229]. Pipelined [1839, 424, 1330]. Pipelines [724, 1420]. Pirates [1031]. PIV [1355]. Pivoting [217, 834]. pixel [825]. Pixels [338]. PJama [872]. PKCS [1041]. PKI [784, 783]. PKIs [184]. Place [2359]. Placed [1297]. Placement [1296, 2519, 1865, 1436, 1317, 1318]. Placing [1383, 1329]. Plan [2669, 1371, 2560]. Plan-Length [1371]. Planar [885, 1759]. Planner [1375]. Planning [2373, 2302, 2295, 2372, 2303, 1599, 2293, 1374, 648, 1144, 1370, 1373, 2329, 1830, 1829, 1831, 1828, 2425, 1376, 2427, 2235, 2273]. Plans [2326, 2392, 1955, 2322]. Plant [1092]. Planted [488]. Plates [2509]. Platform [872, 1927, 1723, 1320, 1073, 2211, 1725, 1337, 470, 2659, 1982]. Platform-Independent [1320]. Platforms [458, 745]. Platypus [856]. Plausibility [107]. Plausible [1471]. Playback [2197]. Player [2065]. Playing [877, 2090]. Playout [1664]. PLCG [2104]. Plot [1828, 2152]. plp [2289, 2026]. Plural [58]. Plus [1993, 2269]. PMI [2099]. PMOS [857]. PNN [309]. Pockets [1597]. Point [1240, 2305, 1575, 1038, 1963, 402, 1232, 1284, 377, 930, 234]. Pointer [403, 404].

Points [1602, 1119, 467]. Poisson [653]. Poker [2065]. Polar [344]. Pole [2445]. Pole-Cart [2445]. Policies [1374, 1594, 857, 2219]. Policy [1978, 1968, 1944, 2090, 763]. Policy-Based [1944]. Politics [117]. Polling [1417]. Pollutants [2497]. Pollution [2487, 2488, 2492, 2494, 2608, 2469]. Polygonal [1862, 1597]. Polyhedrization [358]. Polylingual [853]. Polymorphism [1199]. Polynomial [321, 535, 1390, 910, 1486, 424, 1583, 995, 996, 998, 1867]. Polynomial-Based [1583]. Polynomiality [1212]. Polynomially [935]. Polynomials [1279, 1281]. POMDPs [1150]. Population [1777, 2098]. Porous [2506]. Portability [1086]. Portable [1519, 742]. Portal [116, 2330, 330, 447]. POS [2051]. Position [347]. Positioning [582, 1388, 352]. Positions [641]. Positive [1389, 2151, 1709]. Positives [1381]. Possibilistic [1209, 1166, 1179, 1177]. Possibilities [2002]. Possibility [1178, 1028]. Possible [2304, 1188]. Post [94, 988, 2596]. Post-Turing [2596]. Posteriori [1923, 805]. Postman [495]. Posture [348]. Potential [696, 2506, 606, 1742]. Potentials [603]. Power [974, 886, 1097, 891, 34, 1455, 1103, 987, 1885, 479, 1627, 932, 1366]. Powerful [991]. PowerPC [1244]. Practical [800, 1397, 1585, 180, 1657, 959, 61, 171, 799]. Practice [2397, 2309, 2069, 2329, 1827, 1098]. Pragmatic [37, 1197, 45, 2536]. Pragmatics [37, 52]. Prague [1999]. PRAM [1094]. Pre [1352, 1544, 253]. Pre-aggregation [253]. Pre-fetching [1544]. Pre-synthesized [1352]. Preacher [2620]. Precedence [1847]. Precision [402]. Preconditioner [1487]. Preconditioning [2470, 2471, 1486]. Preconditionings [2463]. Predator [1796]. Predecessor [1089]. Predeclared [1445]. Predicate [208, 1581]. Predicates [513, 241]. Predict [2204, 565]. Predictable [611]. Prediction [2308, 573, 1419, 1418, 984, 2078, 617, 1462, 619, 1798, 1416, 536, 618, 1459]. Prediction-Preserving [984]. Predictions [1339, 52, 1516, 640]. Predictive [2060, 652, 2081, 1507]. Predictor [1460, 1330]. Predicts [1777]. Prefer [1196]. Preference [1146, 1148, 28, 1145]. Preferences [1205, 2362]. Preferential [1189]. Preferred [1201]. Prefetching [1839, 1840]. Prefixed [930]. Prefixes [1993, 1636]. Preliminary [1193]. Preparing [2106]. Presence [91, 1490]. Present [1339, 694]. Presentation [2364, 1558, 1949, 989, 1650, 1946]. Presentations [2609, 1651]. Presenting [2678]. Preservation [1943, 1929, 586]. Preserving [984]. Pressure [1072]. Presupposed [74]. Preventing [1912]. Previsions [1155]. Price [2571]. Pricing [1869, 2184]. Pricing-Based [2184]. Primal [493, 2506]. Primal-Dual [493]. Primary [678]. Prime [1893]. Primes [888]. Primitive [1005]. Primitives [1883, 1032, 1260]. Principal [590, 581]. Principals [771, 770]. Principle [888, 2401, 1139, 1131, 883, 453, 653]. Principled [2354]. Principles [582, 2104]. Priori [1185, 59, 1765]. Prioritizing [1503]. Priority [2303, 1629]. Privacy [181]. Private [178, 945, 2210, 955, 179]. Privileges [784, 783]. ProASIC [1357]. ProASIC-Based [1357]. Probabilistic [1619, 1203, 1168, 1746, 1601, 1154, 1986, 1181, 2073, 610, 825, 1620, 2281, 2109, 2155, 395, 1213, 604, 1202, 544, 1632, 1631]. Probabilities [887, 1743, 1171, 1994, 1167]. Probability [1167, 2077, 1544]. Probability-Based [1544]. Probe [1671]. Probe-Based [1671]. Probes [1383]. Problem [973, 1277, 507, 1851, 508, 228, 495, 1482, 1017, 496, 1532, 1397, 897, 1592, 1388, 1590, 2274, 2071, 1092, 1848, 1369, 986, 1853, 50, 989, 992, 2658, 2317, 2506, 207, 993, 1420, 2140, 504, 539, 1213, 505, 1842, 1846, 1868, 1218, 56, 933, 1290, 1529, 179, 269, 107]. Problem-Solving [1532]. Problems [1016, 1686, 1837, 456, 735, 889, 833, 497, 509, 1165, 1739, 905, 1021, 1590, 1852, 488, 968, 2471, 970, 463, 1850, 2434, 918, 1025, 1376, 1531, 2333, 2486, 60, 1879]. Procedural [2547]. Procedure [1209, 1126, 1583, 431]. Procedures [1135, 892, 1565]. Process [1619, 569, 723, 836, 2242, 2365, 228, 2428, 484, 1995, 2431, 1996, 1993, 580, 2392, 476, 1623, 1614, 2389, 1269, 1622, 4, 1612, 2382, 2388]. Process-Algebraic [1622]. Process-Centred [2389]. Process-Integrated [2388]. Processes [418, 420, 1994, 1640, 624, 571, 2673, 476, 1452, 1630, 1626, 557, 2417, 560, 1812, 2144, 1531, 429, 1621]. Processing [1303, 2245, 1466, 96, 828, 1499, 906, 747, 741, 255, 2162, 250, 2052, 1316, 139, 2422, 746, 743, 47, 1965, 428, 1496, 1337, 350, 453, 520, 355, 686, 609, 235]. Processor [1426, 440, 422, 1216, 1909, 1512, 1465, 1241, 1321]. Processors [1411, 172, 748, 1323, 1345, 1294]. Proclaiming [516]. Produce [1432]. Product [1303, 2349, 2407, 582, 23, 1298, 2658, 2351, 22, 543, 554, 2660]. Product-Line [2658]. Products [1273, 522, 83]. Profile [2641, 1304, 1734, 1459, 1674, 1380]. Profile-Based [1674]. Profile-Driven [1304]. Profile-Profile [1380]. Profiled [2559]. Profiler [1679]. Profiles [155, 1076]. Profiling [2343]. Prognostic [273]. Progol [1705]. Program [2592, 2483, 2069, 1563, 1272, 425, 2049, 468, 867]. Programmable [1339, 1307, 1736, 1301, 1909, 1337]. Programmers [2055, 2054]. Programming [1209, 1474, 2270, 1482, 461, 42, 1519, 2287, 2307, 1388, 2278, 487, 713, 1479, 441, 174, 1580, 1698, 325, 697, 2346, 717, 2264, 2282, 2139, 2055, 448, 2261, 1145, 2408, 2273, 2097, 1081, 2285, 1477, 1310, 1249]. Programs [974, 433, 1028, 1690, 390, 1114, 2455, 2267, 2483, 2284, 1208, 2289, 400, 1770, 2286, 1419, 439, 1418, 1697, 721, 444, 397, 2281, 1476, 394, 395, 727, 1955, 1100, 389, 1221, 2276, 97, 454, 2262, 2266]. Progressive [2074, 2170, 1150]. Project [2424, 2390, 2692, 1921, 224, 1536, 2111, 2157, 2540]. Projection [1446, 577, 586]. Projections [212, 490]. Projects [2387]. Prolog [2272, 415, 227, 1959]. Prolog-Based [1959]. Pronoun [2008]. Proof [956, 1209, 950, 2590, 2592, 1201, 2596, 1232, 1585]. Proof-of-Concept [950]. Proof-Theoretic [2590]. Proofs [1570]. Propagation [745, 360]. Properties [84, 216, 2222, 476, 1613, 370, 2417, 1241, 2418, 2352, 1226, 2107]. Property [1857, 652, 943, 944, 942]. Proportional [2105]. Proposal [2424, 1972, 2617, 2650]. Proposed [164]. Propositional [2287, 2307, 1139, 1148, 1583, 1705]. Propositionalisation [2136]. Propositionalization [1691]. Prosody [2034, 527]. Prospects [2486]. Prosthesis [108]. Protecting [1881, 1031, 1046, 1062, 1973]. Protection [2215, 18, 1076, 2187, 960, 942]. Protection-Based [2187]. Protections [1911]. Protein [296, 469, 1386]. Proteus [1359]. Proto [1759]. Proto-cells [1759]. Protocol [1609, 432, 2192, 797, 1570, 1064, 1319, 1726, 1670, 1231, 1713, 1241, 12, 147, 182]. Protocols [775, 769, 2173, 1571, 415, 199, 2571, 1058, 1096, 414, 2578, 1645, 170, 167, 2201, 2218]. Prototype [1947, 1306, 1979]. Prototypes [2290]. Prototyping [2457, 1294]. Provable [1869]. Provably [1233, 194]. Provably-Correct [1233]. Prover [2448, 2452]. Provers [1579]. Provide [100, 1722, 78, 2432]. Providers [2063].

Providing [954, 452]. Proving

[1566, 1571, 1049, 1071, 476, 1005, 2319]. Provision [2193, 2180]. Provisioning [2189, 1685]. Proximities [578]. Proximity [149]. Proxy [198]. Proxylets [1736]. Pruning [1447, 1705, 1218]. Pseudo [1889]. Pseudo-random [1889]. Pseudorandom [898, 490]. Pseudorandomness [168, 1056, 491]. psmodels [2291]. Pspace [2598, 1134]. Psycho [101]. Psychological [1197]. Psychology [1198]. PTAS [1022, 1606]. PTASs [1016]. Public [793, 792, 169, 962, 1059, 1055, 947, 944, 1082, 164]. Public-Key [793, 792, 169, 1059]. Public-Key-Based [1082]. Publish [2240]. Publish/Subscribe [2240]. Publishers [1954]. Publishing [1954]. PuMA [1306]. Purchase [628, 2570]. Pure [120, 853, 268]. Purpose [1088, 2664]. Purposeful [1781]. Putative [1400]. Puzzles [791, 790]. PVCs [1972]. PVD [1597]. PVFS [1528]. PVM [719, 712, 695, 709, 745, 710, 707, 728, 718, 700]. PVM/MPI [700]. PVS [2419, 1566, 1582]. Pythagorean [1006].

QoS [1680, 1686, 1724, 1734, 2193, 1720, 2189, 1963, 189, 2617, 2212, 1672, 1673, 2180, 1711, 1716, 1718, 2187, 1685, 1687]. QoS-Aware [1716]. QoS-Based [1720]. Quadratic [1280, 1105, 539]. Qualification [2317]. Qualitative [2129, 635, 536, 2446, 2095, 1202]. Quality [2392, 1181, 2563, 2493, 1727, 2197, 1682, 2181]. Quality-Adaptive [2197]. Quantification [1238, 332, 1819]. Quantified [2345]. Quantifier [907]. Quantifiers [209]. Quantify [351]. Quantifying [1995]. Quantiser [1335]. Quantitative [1760, 1178, 2125, 1631]. Quantization [657, 548]. Quantum [974, 887, 1002, 1003, 2476, 881, 2480, 1011, 972, 183]. Quasi [2693, 1432, 2548, 1190, 2477].

Quasi-Classical [1190]. Quasi-Four [2548]. Quasi-Monte [2477]. Quasi-Synchronous [2693]. Quasi-workload [1432]. Quasirandom [2478]. Queries [237, 2440, 240, 1538, 1857, 243, 984, 250, 238, 1545, 2566, 230, 239]. Query [2245, 2610, 297, 741, 255, 159, 1537, 31, 1539, 234, 128, 235]. Querying [2627, 237, 2609, 2252, 1555, 233]. Question [2360]. Question/Answering [2360]. Questions [776]. Queue [1417]. Queue-Based [1417]. Queued [1840]. Queueing [2204]. Queues [1494]. Queuing [2530]. Quiescent [427]. Quorum [2517]. Quorum-Based [2517].

R&D [2381]. Rabin [1044]. Race [2576, 1475]. Radar [2550]. Radial [649, 580, 288, 536, 733, 543, 606, 545]. Radiation [1382]. Radicals [2039]. Radio [2180, 2574]. Radiographs [349]. Radionuclides [2514]. Radiotherapy [330]. Radix [1897]. Ramification [1204]. Ramified [2588]. Ramsey [2350]. Random [508, 1602, 2512, 1988, 810, 823, 365, 1884, 958, 265, 1890, 363, 812, 1879, 1399, 1889]. Random-Error [958]. Randomised [2528]. Randomization [1990]. Randomized [1019, 2532, 260, 1885, 1621]. Randomness [902, 191]. Range [2534]. Rank [907, 1141]. Ranking [124, 1147]. Rankings [1196]. Rapid [1328, 1809, 1295]. Rat [676]. Rate [2198, 2196, 2190, 676, 2038, 673, 2041, 1332, 314]. Rate-Coded [673]. Rated [1693]. Rates [1394, 558]. Ratio [493, 900, 1843]. Rational [1153, 989, 922, 1109]. Rats [1762]. RBF [2338, 535, 541]. RDB [2130]. RDD [2686]. RDD-100 [2686]. RDFS [222]. Re [1364]. Re-use [1364]. Reachability [1220, 1986, 1623, 1620]. Reaction [418, 2514]. Reaction-Diffusion [418]. Reactive [1642, 2174, 482, 2444, 1610]. Reactivity [2052]. Reader [2042]. Real [2641, 36, 1856, 1353, 953, 655, 1315, 438, 2200, 2553, 1740, 1336, 283, 1111, 2643, 2195, 1355, 929, 338, 1011, 1109, 1221, 1337, 356, 722, 405, 2547]. Real-Life [1353]. Real-Time [2641, 36, 655, 1315, 2553, 2643, 1355, 1011, 1221, 1337, 356, 722, 405, 2547]. Real-World [438, 283]. Realities [226]. Reality [346, 90, 87, 2022, 349, 1687]. Realizability [1112]. Realization [482, 479]. Rearranging [1795]. Reason [2324, 1194]. Reason-able [2324]. Reasoning [2288, 1203, 35, 1063, 1569, 37, 2325, 1168, 2290, 2267, 1197, 1616, 1154, 1160, 46, 2356, 2580, 2612, 2275, 1195, 1191, 1118, 1206, 2260, 184, 1176, 2279, 1100, 274, 2353, 273, 1151, 1202]. Reber [621]. Receiver [1330]. Receivers [1664, 1030]. Receptive [678]. Reciprocals [891]. Recognition [818, 2037, 326, 289, 2067, 654, 2021, 2033, 2080, 316, 278, 1301, 2013, 296, 2031, 2025, 2041, 2026, 1706, 353, 574, 290, 1950, 1755, 599, 382]. Recognition-Based [1755]. Recognizable [1013]. Recognizer [1008, 2029]. Recommend [1938]. Recommendation [2254]. Recommender [28]. Recommending [2560]. Reconfigurable [1303, 1352, 1319, 1311, 1335, 1315, 1333, 1326, 1888, 1299, 1354, 1300, 1361, 1340, 1323, 1341, 1334, 1294, 1362, 1365, 1350, 1310, 1358]. Reconfiguration [1492, 1320, 1322]. Reconsidered [1042]. Reconstructing [1057]. Reconstruction [1057, 585, 2369, 370, 843, 811, 2544, 2147, 333, 360, 835]. Record [138]. Record&Replay [721]. Recording [1846]. Recoverability [1252]. Recoverable [872]. Recovery [1892, 1718, 171]. Recurrence [629]. Recurrent [621, 2072, 620, 672, 613, 618, 645]. Recursion [2588, 1572, 2594, 1140]. Recursive [2258, 892, 1159, 1133]. Recursively [1010]. Recv [441]. RED [2185]. Redesign [315]. Reduced [958].

Reducibility [902, 984, 928]. Reducing [526, 141, 201, 2320]. Reduction [584, 1280, 2079, 1283, 1284, 2138, 1286, 528, 1464, 1788, 1289, 363, 1351, 609, 431]. Reductions [1239, 995, 1635]. Redundancy [1435]. Redundant [645]. Rees [1010]. Refactoring [2630]. Reference [1925, 1652, 1473, 1258, 2402]. References [2169, 1275, 1834]. Referencing [1924]. Referential [1659, 153]. Refined [885, 2337]. Refinement [1689, 1220, 218, 1577, 1580, 2500, 1243, 443, 2145, 1639]. Refinement-Based [1243]. Refinements [1986, 408]. Refining [1481]. Reflect [72]. Reflection [2667]. Reflective [2589]. Refuse [1780]. Regard [2689]. Region [2491, 329, 2495]. Regions [2493, 373, 308, 246]. Register [1225, 1884]. Registers [196]. Registration [346, 1540, 337, 842, 180]. Regression [532, 276, 575, 2084, 552, 2087, 830]. Regressors [566]. Regular [2308, 401, 920, 1084, 923, 1879]. Regularity [1137]. Regulatory [1763]. Reification [82]. Reinforcement [2065, 2108, 2154, 2066, 638, 2085, 1764]. Relabelling [2079]. Related [178, 1823, 2417, 1934, 2011]. Relating [1656, 997]. Relation [985, 1626, 2418]. Relational [2059, 2341, 127, 2066, 1693, 254, 256, 139, 377, 2636, 2333, 6, 1000, 1559, 153]. Relations [1207, 2355, 1133, 1206, 2458, 245]. Relationship [2265, 89]. Relationships [256, 149, 77, 230, 822]. Relative [1259]. Relaxation [508, 691, 825, 832]. Relaxed [1193]. Relaxing [810]. Release [1606, 1853, 2546]. Relevance [279, 45, 1941, 52, 1819, 109, 272, 612]. Relevance-Based [1941, 52]. Reliability [1491, 1203, 67, 946, 2488, 1173, 2676]. Reliable [1472, 2584, 754, 427]. Relief [2545]. Relief-Dissection [2545].

Remaindering [958]. Remote

[735, 855, 1530, 447]. Remotely [747]. Removal [387]. Removing [1537]. Renaming [2524, 1884]. Rendering [2553, 817]. Reordering [155]. Repair [2202]. Repairman [917]. Repeated [379]. Repetitions [1849]. Replaceability [2670]. Replacement [1617, 1544, 1594]. Replanning [1372]. Replica [2519, 1865]. Replicated [2537]. Replicating [1767]. Replication [1746, 1716]. Replicator [539]. Replicators [1741]. Report [1213]. Representable [1109]. Representation [735, 1179, 1001, 2367, 977, 1146, 1991, 1148, 2283, 1084, 2268, 1542, 2496, 2014, 2244]. Representations [1609, 205, 481, 1638, 383, 1268, 814, 1984, 246]. Represented [237, 2612]. Representing [2048, 624, 1579, 2642, 58]. Reproducing [1770, 1757, 1222]. repudiation [1570, 1645]. Reputation [946]. Requests [2235]. Required [191]. Requirements [2690, 1426, 1407, 1072, 1728]. RES [2275]. Resampling [2341]. Research [2603, 1686, 1952, 1823, 1916, 2043, 2582, 1600]. Research-Oriented [1916]. Resemblance [1595]. Reservation [1981]. Reset [569]. Residential [1717]. Residuated [1208]. Residue [2450]. Resilience [958]. Resilient [1054]. Resistant [791, 790, 1883, 174]. Resolution [2299, 357, 332, 2490, 2612, 1659, 2493, 2007, 2008, 604, 1718]. Resolving [1192, 1958]. Resonance [669, 332]. Resource [1091, 1744, 1524, 70, 1020, 1377, 1981, 1944, 2425, 2235, 1776, 1712, 2183, 1351]. Resource-Constrained [1377]. Resources [1407, 1521, 98, 1935, 1950]. Respect [1171]. Response [2479]. Restoration [350]. Restricted [2276]. Restrictions [2300]. Result [1788]. Results [887, 905, 1021, 878, 1901, 2081, 1594, 1193, 993, 1375, 505, 371, 1681, 242, 1393, 144, 631]. Resurrecting [796, 795]. Retrieval [2604, 1918, 124, 2610, 1596, 2611, 297, 662, 1941, 1541, 1940, 2613, 762, 1939, 2031, 2016, 749, 130, 807, 306, 179, 128, 2438]. Retrieving [2670]. Retro [527]. Retro-causal [527]. Reusable [1724]. Reuse [2657, 2413, 2390, 2392, 1464]. Reversal [1397]. Reversals [1391, 1396]. Reverse [217, 2544, 355]. Reversibility [1002]. Reversible [307, 533]. Review [1517, 784, 1929, 783, 727, 186]. Revise [2349]. Revised [908]. Revision [1188, 1185, 2350, 1186, 1182]. Revisit [142]. Revisited [1569, 1278, 2305, 1643, 2043, 867]. Revocation [784, 783, 1030]. Revolution [300]. Reward [1992]. Rewards [2637]. Reweighting [2341]. Rewriting [480, 1789]. RHW [1358]. Rhythm [335]. Rhythms [691]. Riddles [2603]. Rigid [806]. Rigidity [842]. Rijndael [1309, 1888, 1886, 1887, 1896]. Ring [2527, 1334]. Rings [2173, 1640, 905, 2531, 1006, 1226]. Rippling [1576]. Risk [787, 1764, 536, 2570]. Risk-Aversion [1764]. RLS [1302]. RMM [1558]. RMM-Based [1558]. RNA [934]. RNS [1910]. Road [231, 382]. Robin [1873]. Robot [648, 641, 642, 1740, 1796, 1797, 2680, 1256, 1801]. Robotic [2416, 1800, 85]. Robotics [2366, 2439]. Robots [2368, 1793, 1803, 650, 1799, 643, 2318]. Robust [1017, 1746, 1060, 2611, 2122, 532, 337, 962, 941, 2009, 2025, 599, 379]. Robustness [2061, 1033, 2052]. Rocks [698]. Rogers [54]. Role [2603, 1781, 2577, 68, 627, 91, 2669, 1831, 54, 85, 752]. Room [228]. Rotated [2470]. Rough [1176, 2151, 156]. Round [1873, 1037]. Rounding [1019, 1025]. Rounds [1034]. Router [1319]. Routines [2161]. Routing [1961, 2534, 1489, 980, 1963, 2188, 1297, 1317, 1318, 1672, 1329, 1673, 1490, 1494, 1674, 1365]. Row [151]. RSA [798, 1278, 1044, 1052, 1043, 1041, 1910]. RT [1315, 1361]. RT-AVI [1315]. RT-Linux [1361]. RTCA [2689]. RTCA/DO [2689]. RTCA/DO-178B [2689]. RTL [1230, 1305]. Rueppel [1285]. Ruin [507]. Rule [1692, 2640, 692, 1147, 2130, 276, 2134, 2082, 278, 2150, 135, 2104, 382]. Rule-Based [2640, 1147, 276]. Rules [415, 1197, 2337, 321, 2062, 2118, 963, 2628, 1102, 105, 174, 2126, 2129, 2133, 2263, 684, 2139, 132, 1554, 78, 2148, 2421, 2320, 2152, 2269, 134, 2125]. Run [1325, 1322]. Run-Time [1325, 1322].

s [1357]. S4 [1120]. SA [308]. SADL [2632]. Safe [483, 2680, 2410]. Safety [2690, 2679, 2673, 2682, 2678, 2677, 2674]. Safety-Critical [2682]. SAFL [1216]. Salesman [508, 1592, 1590, 993]. Salt [188]. Same [274]. Sammon [578]. Sample [2074]. Sampling [509, 2074, 362, 2041, 882, 1158, 2094, 2112, 2158]. SAMR [1428]. SANs [1527]. SAT [1325, 2348, 883, 1218]. SAT-Based [1218]. Satellite [2130, 2176]. Satisfaction [1702]. Satisfiability [2307, 2345, 921, 1213]. Satisfying [2572, 1093]. Scalability [313, 2112, 2158]. Scalable [872, 1522, 403, 1403, 2071, 2189, 748, 1909, 2254, 1897, 559, 1685, 722, 767]. Scalar [1335, 1892]. Scale [2487, 2465, 1146, 2485, 219, 2494, 852, 2488, 248]. Scan [1342]. Scanner [354]. Scanning [761, 1454]. SCC [1639]. Scenario [2643, 2668, 2406]. Scenario-Based [2643, 2668, 2406]. Scenarios [32, 1187]. Scene [382, 521]. Scenes [661, 2370, 2367, 2369, 2353, 2616]. Scheduled [2695]. Scheduler [1469, 1467]. Schedulers [2200, 1664]. Schedules [1426, 1230]. Scheduling [1016, 1425, 1851, 494, 1431, 1346, 2371, 1238, 509, 904, 1852, 1022, 1369, 1510, 1606, 1508, 1548, 1853, 224, 1373, 915, 1847, 1412, 264, 392, 2232, 1427]. Schema [2590, 493, 2224, 2245, 1551, 137]. Schemas [1561, 5]. Schemata [2287, 2236]. Scheme [2198, 1039, 1050, 2208, 2213, 1721, 719, 2186, 1068, 1931, 1049, 181, 2190, 386, 193, 18, 964, 1718, 194, 1685, 2689]. Schemes [205, 2303, 497, 1875, 2492, 191, 463, 1285, 1030, 195, 1048, 2503, 196, 2508, 171]. Schlesinger [320]. Schnorr [194]. Scholarship [1921]. Schur [2472]. SCI [1361, 470]. SCI-Based [470]. SCI-Cluster [1361]. Science [2203, 114, 971, 2418]. Scientific [2485, 1954, 2486]. Scope [789, 1346, 788]. Score [594]. Scoring [2143, 1380]. Scrambling [1883]. Screening [1399]. script [1950]. Scripts [1695, 1081]. SDMI [965]. Sea [98]. Seal [1615]. Sealed [757]. Seals [2502]. Seamless [2554]. Search [1162, 885, 2174, 126, 1372, 104, 2527, 2308, 2309, 1374, 878, 861, 1370, 1369, 2537, 1377, 2528, 489, 1373, 2046, 1916, 1605, 1084, 1635, 1376, 2559, 31, 2348, 883, 2319, 234, 59, 1979, 309, 144, 2112, 2158, 159]. Searching [216, 29, 1603, 744, 1765, 1163]. sec [1886]. Seceder [1747]. Second [2571, 1111, 961, 2093]. Second-Order [1111]. Second-Price [2571]. Secrecy [420]. Secret [1278, 1057, 900]. Secrets [900]. Sector [2377, 1076]. Secure [1906, 789, 798, 178, 2571, 1058, 1034, 1032, 1043, 1889, 1047, 173, 1307, 1073, 1046, 198, 1037, 193, 1076, 958, 788, 194, 1982, 177]. Security [769, 1616, 187, 1068, 2222, 2221, 414, 2220, 16, 1005, 1076, 1062, 416, 2675, 176, 2217, 409, 1612, 2674, 186]. Security-Typed [416]. SEE [2389]. Seed [375]. Seeking [2255]. Segment [2024, 1283, 1284]. Segmentation [361, 2017, 332, 534, 665, 824, 819, 815, 365, 336, 369, 658, 323, 375, 372, 820, 338, 373]. Segmentations [823]. segmented [298]. Segments [608, 581]. Segregation [812]. Selected [461, 2377]. Selecting [1734, 2572]. Selection [2365, 1813, 523, 2127, 2128, 293, 283, 23, 857, 375, 2298, 319, 1389, 1546]. Selective [510]. Selectivity [238]. Self [2238, 2114, 584, 1599, 1746, 582, 588, 1770, 2531, 636, 1469, 2530, 1508, 2529, 1488, 662, 1031, 1970, 578, 1807, 925, 1757, 2141, 2207, 608, 1767, 644, 1801, 632, 1683, 583, 1410, 678, 1802, 1226, 637, 292, 165]. Self-Adaptable [1970]. Self-Adjusting [1508]. Self-Avoiding [925]. Self-Exploring [292]. Self-Made [1599]. Self-Optimizing [1488]. Self-Organisation [637]. Self-Organising [1802]. Self-Organization [1801, 678]. Self-Organized [584]. Self-Organizing [2238, 582, 588, 662, 1807, 2141, 608, 632, 583, 1410]. Self-Replicating [1767]. Self-Replication [1746]. Self-Reproducing [1770, 1757]. Self-Shrinking [165]. Self-Similar [2114]. Self-Similarity [2207, 1683]. Self-Stabilizing [2531, 2529, 1469]. Self-Timed [1226]. Self-Tuning [644]. Sellers [2572]. Semantic [1568, 2615, 2645, 2324, 37, 148, 90, 2246, 1806, 2457, 1190, 1111, 2330, 2141, 2559, 2235, 2051, 2097, 2004, 235]. Semantically [2011]. Semantico [205]. Semantico-Cognitive [205]. Semantics [1120, 2621, 2288, 37, 2304, 2265, 484, 1201, 2629, 1551, 1178, 1070, 2644, 1623, 1248, 2638, 1175, 2281, 2282, 2007, 875, 396, 2044, 2266]. Semi [2115, 225, 2068, 525, 1273, 1987, 2050, 349, 2237, 964, 1708, 987]. Semi-AFPs [987]. Semi-anonymous [964]. Semi-automated [349]. Semi-automatic [2050]. Semi-direct [1273]. Semi-Markov [1987]. Semi-structured [225, 2237, 1708]. Semi-supervised [2115, 2068, 525]. Semiconcept [210]. Semiconductors [2476, 2480]. Semidefinite [487]. Semigroups [1010, 884]. Semiring [927]. Semirings [1013]. Send [441]. Send-Recv [441]. Sense [1998, 780, 779, 2019, 98]. Sensed [747]. Sensing [2322]. Sensitive [2001, 2082, 922]. Sensor [1173, 386, 622, 583]. Sensors [1784]. Sentence [2140]. Sentences [682, 2053, 206]. Separable [2510, 1724]. Separating [320, 2660]. Separation [1856, 590, 597, 591, 1233, 589]. Sequence [1381, 663, 1446, 1072, 2132, 2642, 1440, 468, 618, 469, 354]. Sequences [2474, 1391, 2116, 357, 665, 2078, 254, 465, 372, 350, 640]. Sequential [2473, 889, 1165, 585, 2077, 1547, 1327, 639, 1621]. Serial [1908, 1898, 1350]. Serializability [2246]. Series [891, 623, 611, 2123, 2130, 2129, 617, 987, 612, 2150]. Server [2453, 249, 1696, 2195, 707, 1842]. Servers [1472, 2202, 1718]. Service [2242, 1676, 952, 945, 1524, 1430, 1721, 794, 1719, 2522, 2211, 1511, 188, 1535, 1682, 2181, 1677, 2194, 2244]. Services [789, 190, 1675, 1069, 2209, 1720, 2200, 1722, 2386, 1968, 1678, 2211, 176, 2562, 788, 2178]. Sesquicartesian [2593]. Session [869, 850, 866, 1051, 859, 854, 847, 862]. Session-Key [1051]. Sessions [1727, 1668]. Set [2288, 2295, 1017, 1602, 840, 1596, 1499, 2287, 2307, 499, 2278, 2127, 1613, 2525, 204, 1584, 1323, 479, 574, 2273, 2151, 2452, 156, 885, 780, 779]. Sets [785, 786, 1899, 899, 951, 152, 1581, 1176, 1009, 1950]. Setting [1933]. Several [458]. SGI [2490]. Shading [824, 378, 376]. Shallow [2052]. Shape [841, 318, 315, 839, 347, 2026, 378, 807, 376, 835]. Shape-Adaptive [315]. Shape-from-Shading [378]. Shaped [843]. Shared [1094, 419, 713, 1447, 467, 1712]. Shared-Memory [1094]. SharedPlans [1653]. Sharing [1066, 900, 191, 548, 852]. Sheets [351]. Shell [1695]. Shellsort [978]. Shift [1085, 196]. Shifting [297, 1740]. SHMEM [704]. Shock [318]. Shop [906]. Shops [800, 799]. Short [1162, 798, 1866, 2412, 1527, 686]. Shortest [1277, 1854, 1451, 1098, 445, 1365, 1290]. Shot [659, 275]. Shouldn’t [1]. Shrinking [1053, 165]. Shuffle [1049]. Shuffling [151]. Shuttle [2550]. Side [777, 1388, 1913]. Side-Channel [1913]. Sided [706, 701, 1481]. Sides [274]. Sieve [1277].

SiGe [1299]. Sigmoidal [647]. Signal [1303, 2479, 938, 606, 453, 609]. Signalling [1740]. Signals [605, 602]. Signature [800, 198, 1285, 1076, 1079, 195, 1048, 1080, 196, 799]. Signature-Embedded [800, 799]. Signatures [1052, 126, 1047, 1035, 194]. Signed [2432]. Significance [1385, 1498]. Signing [1047]. Silico [1752]. SIMD [1466, 1465]. Similar [2114, 137]. Similarities [2237, 807]. Similarity [2344, 125, 126, 2120, 2012, 232, 1816, 141, 2086, 298, 2092, 2207, 1683]. Similarity-Based [125]. Simple [205, 1400, 312, 496, 1844, 2070, 1854, 1092, 210, 1785, 1774, 150]. Simpler [1852, 909]. Simplex [843]. Simplified [1044, 2065, 2618]. Simply [1130]. SIMs [1073]. SIMspeak [1073]. Simulated [1167, 2067, 824, 488, 1773, 651]. Simulating [50, 63]. Simulation [418, 2504, 458, 459, 187, 10, 1470, 2497, 1865, 2514, 75, 2476, 745, 1751, 2034, 763, 2480, 2359, 2347, 541, 455]. Simulations [2488, 1530, 2417]. Simulators [615]. Simultaneous [1205, 2569, 1318, 1152, 2103]. Simultaneously [322]. Single [1869, 492, 2590, 1194, 865, 904, 603, 2190, 1509, 2618, 1887, 1308, 589, 609]. Single-Chip [1887, 1308]. Single-Machine [904]. Single-Rate [2190]. Single-Stock [1869]. Singular [975]. SITA [1973]. site [2566]. Sites [1400, 1394, 24]. Situated [2357, 72]. Situation [2367, 2314, 112, 2318]. Situations [2032]. Size [332, 2074, 524, 932, 2035, 880]. Skeletonization [371]. Skeptically [1194]. SKETHIC [177]. Skew [1444]. Skewed [1607]. Skills [2384, 85]. Skip [1607]. SLA [2210]. SLA-Based [2210]. Sleep [604]. Slicing [391, 2521]. Sliding [1904]. Slots [1427]. Slovak [2028]. Slovenian [2029]. Slow [667]. SlugBot [1796]. Small [2583, 495, 1279, 2479, 988, 166, 4, 1867].

Small-Signal [2479]. Smaller [1624]. Smallest [1858, 1804]. Smart [1064, 1067, 1068, 1069, 2032, 1076, 764, 1078, 1079, 2112, 2158, 1082]. Smart-Card [1067]. Smart-Card-Based [1068]. Smartcard [1072, 1081]. Smartcards [1065, 1071, 1080]. SMIL [1949]. Smodels [2294]. Smooth [1805, 1509]. Smoothing [378]. SMP [1424, 701]. SMS [2042]. SNPs [1850]. SOAP [1069]. Sobol [2474]. Sobolev [2471]. Soccer [1798, 1800]. Social [2577, 104, 114, 1183, 2090, 2426, 2421, 85]. Societies [1954]. Society [761]. Soft [548, 1670, 392, 581]. Soft-State [1670]. Software [2654, 2407, 2431, 96, 2400, 20, 2384, 2375, 1181, 1404, 2405, 174, 2635, 2381, 1320, 1354, 1613, 2387, 82, 2607, 2650, 2380, 118, 4, 2432, 2382, 1496, 2560, 2321, 2402, 2379, 948, 413, 12, 2670, 720, 2631, 1962, 2689, 2668]. Solid [1266, 343]. Solum [1102]. Solution [2270, 1397, 1724, 2350, 1487]. Solutions [1394, 1279, 1187, 2274, 2313, 276, 2022, 1076, 115]. Solvable [1390]. Solve [1739]. Solver [2510, 2513, 2449, 2500]. Solvers [1325, 2512, 1488, 740]. Solving [1482, 1532, 1590, 648, 1369, 2478, 50, 401, 207, 2472, 1531, 1529]. Solving-Oriented [207]. SOM [661, 2370, 585]. SOM/ARSOM [661, 2370]. Some [1906, 1837, 178, 384, 878, 983, 2021, 1784, 1186, 2081, 1198, 2468, 1193, 466, 77, 505, 1681, 1]. Sonic [1750]. Sorption [1752]. Sorter [1242]. Sorting [1396, 1799]. Sound [600, 1882, 2032]. Soundness [1059]. Source [590, 2553, 391, 597, 2049, 589, 867, 2237]. Sources [1998, 67, 1540, 2361, 2196, 141, 591, 1939, 2237, 6, 2549]. Southeastern [2495]. SP2 [703]. SPA [1912]. SPA/DPA [1912]. Space [90, 84, 1844, 865, 1003, 2309, 1806, 2531, 106, 981, 908, 422, 2471, 1087, 296, 814, 2147, 844, 1150, 1163]. Space-Efficient [1003, 1087].

Space-Progressive [1150]. Spaces [2367, 1936, 2220, 2141, 543]. Spanish [2008]. Spanning [1858, 2529]. Sparse [899, 2464, 566]. SPART [335]. Spatial [2344, 240, 326, 124, 1526, 255, 2612, 1762, 250, 335, 254, 1773, 149, 1701, 238, 252, 253, 132, 373, 242, 281, 306, 239, 1182, 241]. Spatial-Match [124]. Spatially [236, 1753]. Spatio [2121, 823, 233]. Spatio-Temporal [823, 233, 2121]. Spatiotemporal [1543, 251, 231, 851]. Speak [26]. Speaking [2034]. Special [924]. Specialization [1170]. Specially [1486]. Specific [2389, 2300, 1332, 206]. Specification [2270, 2690, 2628, 2454, 2626, 2640, 1306, 111, 1075, 1553, 2652, 1241, 453, 752]. Specifications [780, 779, 1244, 914, 1221, 1366]. Specifying [71, 2124, 1228, 2648]. Speckle [286]. Spectral [2308, 1604, 2025]. Spectrally [326]. Spectroscopy [353]. Spectrum [962, 757, 1485]. Speculative [1462, 626]. Speech [2037, 2036, 2021, 2033, 2030, 2024, 2027, 2022, 1301, 2038, 2031, 601, 2025, 2041, 1657, 599, 2045, 2014, 2029]. SpeechDat [2028]. Speed [189, 1363, 1082]. Speeding [2066]. Spell [1601, 685]. SPH [742]. Spherical [843]. Spider [1976, 438]. Spike [670]. Spiking [569, 669, 553]. Splicing [2252]. Splines [837, 1267]. Splitting [1861, 2489, 1434]. SPMD [433, 435, 694]. SPMD-Tasks [435]. Spoken [2047, 2043, 2031]. Spontaneous [607, 1759]. Sport [345]. Sports [2615, 1369]. Spot [1490]. Spread [962]. Squares [600, 572]. SSE [1423]. SSL [1046]. SSM [2192]. Stability [1804, 321, 1758]. Stabilization [2445]. Stabilize [526]. Stabilizing [2531, 2530, 2537, 2529, 1469]. Stable [2523, 2290, 1017, 2271, 1487, 1726, 2264, 1749, 1704, 2291, 1597]. Stage [1180, 639]. Stall [492]. Stamping [190]. Standard [2658, 1308, 1309, 1733, 1337, 14]. Standardization [154]. Standardized [1041]. Star [843]. Star-Shaped [843]. Stars [1021]. Starting [2074]. State [2473, 889, 621, 2532, 2638, 1552, 1670, 1725, 2582, 2231, 1115, 2322, 1368, 1644, 2198, 1761]. Statechart [2647]. Statecharts [2677, 2649]. Statei [1950]. Stateless [1030]. Statements [1247]. States [1007]. Static [1634, 1744, 420, 1304, 1066, 412, 1479, 402, 399, 394, 396, 413]. Stationary [1004, 463]. Statistical [1385, 2476, 2080, 2688, 1198, 51, 1594, 1379, 197, 2480, 2665, 311, 281, 2110, 2156, 286, 379]. Statistics [1782, 294]. Statistification [2110, 2156]. Status [1171, 2375, 1999, 1955]. Stealing [2183]. Steel [351]. Steepest [730]. Steering [1413]. Steganographic [957]. Steganography [938]. Steiner [933]. Step [421]. Steps [1166, 1921, 2524, 2052]. Stepwise [2335]. Stereo [2368, 1313, 384, 810, 380, 835]. Stereoscopic [665]. Still [660]. Stimuli [668]. Stochastic [569, 669, 457, 1996, 1993, 1989, 1374, 1384, 1630, 2051, 1163]. Stock [1869]. Stokes [2498]. Stop [195]. Stop Consonant [574]. Stop Consonant-Vowel [574]. Stopping [2508]. Storage [870, 1525, 1036, 861]. Store [856, 1509, 2351, 860]. Stores [1437, 848]. Stories [103]. Story [1827, 1826]. Straightness [351]. Strand [2220]. Stratagems [1939]. Strategic [758]. Strategies [20, 2519, 2200, 2572, 1557, 2686, 1937, 1678, 2234, 2175, 521]. Strategy [2179, 160, 2535, 2085, 1749, 2104, 175, 639]. Stratification [1193]. Stratified [2072, 1134]. Stream [1769, 2211, 1907, 343]. Streamed [1499]. Streaming [1669, 1732]. Streams [1940, 1727, 949]. Strength [2590, 2430, 880]. Stress [2032]. Stretching [1692]. Strict [2269]. Strictness [393]. Strikes [782, 781]. String [400, 1085, 141, 1789]. Strong [952, 198, 2262, 2104]. Strongly [1839, 983]. Structural [2499, 1582, 1699, 739].

Structurally [444]. Structure

[2238, 809, 1088, 126, 951, 1450, 2001, 1266, 1161, 352, 2426, 2421, 2503, 1012, 1801, 1755, 1350, 1802]. Structured [661, 1019, 1748, 174, 2638, 1486, 31, 520, 225, 2237, 1708]. Structures [1836, 2428, 1746, 2121, 212, 1104, 207, 263, 994, 1100, 166, 1089, 1000, 934, 12, 1122]. Structuring [2413, 2385, 2320]. Students [1947, 91]. Studies [68, 227, 76, 111, 1757, 11]. Study [1738, 1095, 1497, 2295, 1086, 1947, 2407, 1096, 10, 42, 384, 1576, 2124, 1762, 1937, 1741, 1361, 2180, 1603, 2678, 2002, 2142, 743, 2382, 1423, 738, 2687, 586, 2266]. Studying [2512]. Styles [2034]. Sub [825, 2237]. Sub-pixel [825]. Sub-source [2237]. Subcategorization [2005]. Subclasses [2075]. Subgraph [504]. Subgroup [179]. Subject [1540, 1536]. Subject-Oriented [1536]. Subqueries [1537]. Subscribe [2240]. Subset [499]. Subsets [1008, 1010, 1013, 2106]. Subspace [2511, 614, 1149]. Subspaces [667]. Substation [2675]. Subsumption [1702]. Subtraction [1885]. Subtyping [1581]. Successes [696]. Successive [1986]. Succumbs [1904]. Suffix [1605]. Suggestion [1197]. Suicide [1749]. Suitability [707]. Suites [2652]. Sulfur [2495]. SUM [1016]. Summarization [2115, 682, 1934]. Summary [880, 2167, 242]. Super [357, 1742]. Super-Resolution [357]. Super-Turing [1742]. Superimposition [1941]. Superlinear [554]. Superposition [973, 1135]. Superquadric [316]. Superquadric-Based [316]. Superscalar [1465, 2687]. Superscalars [1463]. Supervised [2338, 621, 815, 2115, 2068, 525]. Supervision [2032]. Supply [2331]. Support [562, 2667, 2645, 2227, 1675, 459, 2540, 2371, 1438, 2691, 963, 2108, 2154, 715, 2209, 573, 765, 712, 1408, 1147, 2423, 575, 2084, 53, 563, 564, 1682, 2148, 576, 574, 1539, 215, 1330, 2194, 705, 83, 58, 572, 2383, 720, 2631]. Supported [2390, 10, 2613]. Supporting [433, 2565, 13, 1921, 1720, 2381, 2195, 1161, 2351, 2559, 2241, 19, 1668, 140, 1549]. supposed [74]. Surface [1260, 811]. Surfaces [2475, 2071]. Surgery [2680]. Survey [534, 879, 463, 197, 1838]. Survival [1804, 1747]. SWAN [1824]. Swap [1437]. Swapping [1312]. Sweep [1268]. Swiss [769]. Switch [738]. Switched [432]. Switches [2193]. Switching [2171, 2172, 1882, 2170]. Sylvester [2511]. Symbiosis [2404]. Symbiotic [1791]. Symbolic [1244, 2367, 48, 2612, 1985, 1620, 2089, 527, 618, 1227, 1554, 2443, 2442, 1809, 1644]. Symbolic-Graphic-Numeric [2443]. Symbolic-Sequence [618]. Symbols [2290, 1274]. Symmetric [793, 792, 810, 1246, 1556, 918, 630, 267]. Symmetric-Key [793, 792]. Symmetry [2400, 928]. Synapses [671]. Synaptic [553]. Synchronization [670, 1452, 1222]. Synchronized [1617]. Synchronizing [912]. Synchronous [2693, 2526, 2243, 2525, 759, 1646]. Synchrony [1641]. Synonyms [2099]. Syntactic [1353, 2012, 927]. Syntactic-Based [2012]. Syntactically [137]. Syntax [690, 1810]. Synthesis [489, 1300, 2038, 1216, 1234, 1291]. Synthesized [444, 1352]. Synthesizer [1331]. Synthesizing [1305]. System [2368, 2288, 787, 2641, 35, 2325, 2245, 2634, 2358, 437, 1721, 2567, 2292, 946, 865, 127, 2454, 2519, 1865, 2293, 158, 711, 765, 1455, 2433, 1560, 351, 1377, 1347, 2578, 2416, 145, 2561, 1509, 1967, 1659, 2052, 1968, 2642, 1552, 1959, 138, 139, 1970, 2447, 2027, 2028, 1817, 425, 1301, 2055, 764, 1951, 1356, 1716, 282, 2016, 410, 2445, 2446, 1542, 1345, 2451, 450, 428, 2648, 2254, 226, 2418, 1966, 355, 2294, 119, 1801, 2054, 755, 1366, 722, 1477, 2101, 61, 135, 133, 690, 835, 521, 143, 1082, 767, 2285]. System-Mediated [1951]. Systems [2238, 2419, 8, 492, 2323, 1743, 1881, 125, 1618, 954, 1210, 1642, 1634, 2357, 889, 646, 482, 1945, 855, 459, 2246, 1660, 2412, 2248, 13, 160, 2693, 977, 953, 1315, 1986, 9, 1596, 2512, 2429, 127, 2257, 2247, 485, 2278, 1211, 1468, 1971, 2444, 2433, 1578, 2272, 2470, 2484, 192, 1637, 114, 217, 2423, 1740, 2420, 2606, 619, 1544, 2478, 1501, 76, 2047, 1243, 2688, 1772, 111, 2501, 1486, 1912, 28, 82, 2176, 2682, 1444, 1610, 921, 2435, 1977, 2675, 2422, 2472, 2279, 25, 1550, 2043, 681, 2347, 2683, 2695, 1553, 1960, 78, 2442, 1957]. Systems [1907, 1334, 2231, 267, 2426, 1778, 1627, 1658, 2250, 559, 1366, 1495, 960, 405, 1310, 1742, 2674, 2014, 2631, 625, 2438, 1632, 1646, 1631, 793, 792, 1989]. Systolic [1334, 469].

T3E [1415]. Table [2165, 996]. Tableau [1190]. Tableaux [1116]. Tables [1445]. Tabu [1369]. Tactical [75]. Tactics [1798]. Tag [468]. Tagged [2002]. Tamper [1883, 174]. Tamper-Resistant [174]. TAO XML [2607]. Targets [1855, 1076]. Tarski [1101]. Task [1311, 1419, 2389, 2557, 283, 746, 1009, 430, 1310]. Task-Parallel [1310]. Task-Specific [2389]. Tasks [435, 1431, 1918, 1022, 648, 501, 386, 1847, 1935, 1504]. Taught [85]. Taxon [1394]. Taxonomy [1816]. TBL [1703]. TCP [2208, 2196, 2204, 1726, 1732, 1678, 2176, 1681, 2207, 1665, 2177]. TCP-Friendly [2196, 1726, 1665]. tech [763]. Technical [2429]. Technique [885, 657, 1634, 493, 146, 461, 2605, 1165, 323, 2007, 1392]. Techniques [2303, 2413, 1343, 1852, 1624, 1373, 1025, 1233, 394, 2521, 1705, 2094, 449, 1218, 2170, 720, 942]. Technological [763]. Technologies [2657, 93, 2554]. Technology [84, 110, 1525, 460, 120, 101, 2384, 72, 1293, 1404, 80, 2563, 2613, 2022, 271, 1981].

Tectogrammatical [2015]. Telephone [2026]. Telephone-Based [2026]. Telephony [2199, 1668]. Television [16]. Tell [801]. Template [806, 2667, 2248, 838, 1323]. Templates [2314, 2410]. Temporal [205, 148, 2116, 2299, 2580, 2130, 2129, 823, 1643, 111, 1440, 1550, 233, 673, 273, 281, 2150, 884, 1226, 2121]. Temporal-Spatial [281]. Ten [1881]. Tensor [890]. Term [2058, 224, 1298, 2013, 2495, 1683, 2101]. Term-Based [224]. Terminal [1129]. Termination [394, 395, 1554]. Terms [537, 2013]. Ternary [1280, 2644]. Terraffectors [2542]. Terrain [2553, 2549, 2547]. Tertiary [2453]. Test [1782, 1067, 2512, 2209, 2350, 1848, 2669, 1389]. Test-Bed [2209]. Testability [999]. Testbed [1686, 707]. Testing [1769, 2634, 1857, 2688, 2643, 1630, 52, 516, 515]. Tests [1592, 1890]. Texas [2065]. Text [2115, 682, 2365, 2017, 1596, 2342, 2135, 323, 744, 2141, 2093, 749, 1787, 2096, 285, 271, 2014, 1924, 129]. Text-to-Speech [2014]. Textbooks [1915]. Texts [225, 1649, 2014]. Textual [2018, 1649]. Texture [818, 2611, 819, 815, 365, 327, 303]. Textures [326]. TH [708]. TH-MPI [708]. Their [481, 482, 302, 304, 1173, 212, 189, 2137, 2022, 207, 221]. Theme [1541]. Theorem [1566, 1132, 894, 1124, 908, 1071, 1107, 1579, 1198, 919, 2319, 1842]. Theorema [2452, 2448]. Theorems [321]. Theoretic [2590, 1168, 2430, 542, 1819, 997, 948]. Theoretical [2081, 673, 376]. Theories [1689, 1201, 2268, 1137, 2262]. Theorists [2412]. Theory [1203, 2397, 1449, 1572, 896, 1943, 1184, 877, 1178, 2309, 2354, 2433, 2596, 2077, 91, 2297, 2329, 1139, 2501, 1257, 1253, 1254, 1265, 1255, 1006, 1630, 1826, 2521, 1584, 1818, 2142, 215, 1101, 2452, 156]. Therapy [2030, 85]. There [977, 899]. Thermography [333]. Thesaurus [1917].

Thinking [83]. Thinning [331]. Third [204]. Thought [92, 2226, 2110, 2156, 2410]. Thoughts [162]. Thread [1477]. Thread-Based [1477]. threaded [714]. Threads [1411, 1467]. Three [1394, 313, 1770, 641, 145, 1487, 358, 2019, 2385, 2146]. Three-Dimensional [1770, 1487, 358]. Three-Fingered [641]. Three-Layer [313]. Three-Taxon [1394]. Three-Tier [145]. Threephone [2023]. Threshold [191, 193, 194, 999, 629]. Thresholding [605, 373]. Throughput [1417]. Tier [145]. Tight [900, 915, 558]. Tightly [1317]. Tile [2612, 1544]. Tiling [1482, 1333]. Time [492, 190, 2641, 36, 494, 2305, 1861, 655, 1315, 1325, 623, 2309, 1020, 1022, 1023, 1872, 1390, 2200, 981, 611, 2123, 1854, 2553, 982, 610, 2130, 2129, 1322, 617, 1336, 2516, 913, 991, 244, 2643, 552, 2195, 1355, 994, 995, 180, 1008, 618, 2025, 996, 338, 1011, 998, 1221, 630, 1337, 1427, 356, 612, 738, 1867, 2150, 722, 430, 405, 2547, 2469]. Time-Dependent [1023]. Time-Integration [2469]. Time-Series [2130, 2150]. Time-Space [2309]. Time-Stamping [190]. Time-Varying [552]. Time-Window [611]. Time/Space [981]. Timed [1220, 1625, 1989, 990, 429, 1226]. Times [1853, 1605]. Timestamps [1410]. Timing [1228, 2648, 2625]. Tissue [596]. TM [1244]. TMOS [858]. TMR [1472]. TMR-Based [1472]. TOAST [1723]. Today [328]. TOEFL [2099]. Together [98, 83]. Tolerance [789, 1343, 1719, 788, 1504]. Tolerant [1746, 708, 1873]. Tomography [2206, 1593]. Tomorrow [328]. Tool [1411, 1154, 2457, 103, 725, 80, 1613, 1318, 2047, 716, 1235, 464, 1944, 941, 282, 452, 396, 134]. Toolkit [1421]. Tools [305, 2428, 110, 3, 861, 1408, 1347, 1762, 98, 1233, 727, 25, 698, 2442, 2689]. Toolset [1063]. Top [1457, 704, 710, 2039]. Topic [1425, 1458, 1474, 1502, 1489, 1449, 1525, 1468, 1408, 1433, 1442, 1416, 1453, 1515, 1495, 1529, 1483]. Topically [1934]. Topics [1823]. Topographic [278, 374]. Topography [2550]. Topological [587]. Topologies [21, 1386]. Topology [1429, 632]. TOPPER [716]. Tori [1493]. Tossing [1037]. Total [2121]. Totals [2118]. Trace [922]. Traceability [18]. Tracefiles [726]. Traces [2064, 2459]. Tracing [1031, 1030, 119]. Tracker [338]. Tracking [841, 341, 339, 330, 345, 813]. Trade [954, 34, 2309, 1605, 1973, 2148]. Trade-Offs [954, 34, 2309]. Tradeoffs [1908, 1417]. Trader [10]. Trader-Supported [10]. Trading [759]. traditional [2368]. Traffic [954, 1680, 2174, 1669, 1679, 2193, 386, 1490, 2207, 1674, 2194, 1683]. Trails [30]. Train [2380]. Trained [621]. Training [562, 600, 647, 2083, 2094, 2053]. Traitor [1031]. Trajectories [1806, 2466, 2220, 2025]. Trajectory [1244, 653]. Trakhtenbrot [1124]. Transaction [785, 786, 1675, 1899, 428]. Transactional [855, 104, 858]. Transactions [483, 2247, 12]. Transcript [797, 772, 800, 789, 801, 775, 769, 777, 782, 787, 771, 780, 776, 793, 791, 796, 784, 798, 774]. Transcription [1400, 1787]. Transducer [2063]. Transductive [551, 2096]. Transfer [1036, 1253, 1254, 1251]. Transferable [1173]. Transfinite [2592]. Transform [1574]. Transformation [1634, 1700, 2638, 2607, 322, 829, 61]. Transformation-Based [1700]. Transformations [1225, 7, 168, 1230]. Transforms [1356, 307]. Transients [630]. Transition [485, 1236, 1637, 2038, 1627, 1631]. Transitioning [1731]. Transitions [659, 275, 1785]. Transitive [728]. Translational [344]. Translator [439, 1585, 2647]. Translatory [1808]. Transmission [2437, 1726, 1725, 1778].

Transparency [1971]. Transparent [93, 1903, 1496, 864, 1477]. Transport [432, 2171, 2514, 703, 2195, 1713, 2201, 740]. Transport-Level [1713]. Transpositions [1396]. transsys [1763]. Travel [1599, 2560]. Traveling [508, 1592, 1590, 917, 993]. Treatment [330, 2469]. Tree [885, 398, 2066, 2308, 1390, 1446, 2529, 1556, 1498, 401, 1137, 1587, 998, 2149, 933, 832]. Treebank [1703, 1999]. Trees [1167, 1394, 897, 1538, 905, 1598, 1390, 2072, 255, 2537, 2202, 2079, 2564, 318, 2335, 2087, 1395, 831, 132, 2095, 2100]. Treewidth [1867]. Trends [1407, 1293, 1952]. Trial [603]. Triangle [247, 265]. Triangle-Free [265]. Triangular [340, 374]. Triangulation [1861]. Trigger [1456]. Trilinear [2470]. Trip [2560]. Triples [1006]. Triticale [367]. Trivial [2345]. Trojan [177]. Tropos [2650]. Troubleshooting [1152]. True [1890, 1465]. Truly [753]. Truncated [1157, 1351]. Trust [787, 774, 2577, 184, 18, 773]. Trust-Based [787]. Trusted [186]. Trustworthy [937]. Truth [1945, 2345, 996]. Truth-Table [996]. Truths [441]. TTIG [743]. Tumor [330]. Tunable [658]. Tuning [1173, 2354, 1409, 1420, 2695, 644]. Turing [1003, 2596, 1011, 632, 1742]. Turning [1069]. Tutorials [2655]. Two [1166, 2463, 40, 484, 2065, 2513, 1500, 1888, 861, 909, 672, 2033, 2052, 1678, 551, 1037, 1117, 1035, 2472, 1288, 925, 1007, 22, 274, 2051, 826, 273, 1481, 2321, 934]. Two-Differentiated [1678]. Two-Dimensional [2513, 925]. Two-Layered [22]. Two-Level [2463, 2472, 2051]. Two-Party [1037, 1035]. Two-Pass [2033]. Two-Player [2065]. Two-Sided [1481]. Two-States [1007]. Two-Steps [1166]. Two-Variable [1117]. Two-Way [909]. TwoKey [2152]. Type [2588, 1572, 542, 1578, 414, 168, 1139, 1129, 2035]. Typed [1130, 416]. Types [1138, 1539]. Typicalness [2088]. Typing [1615, 440]. Typogenetics [1741].

Ubiquitous [1534]. UDTO [2166]. Ulisse [865]. Ultrasound [808]. UML [2627, 2641, 2645, 2634, 2637, 2626, 2640, 2624, 2644, 2635, 2655, 2638, 2643, 2636, 2650, 2677, 2648, 2630, 2651, 2625, 2631, 2649]. UML-Based [2634]. UML-Statecharts [2649]. Unary [909]. Unaveraged [609]. Unavoidable [1640]. Unbalanced [1451]. Unbounded [1782]. Uncalibrated [2369]. Uncertain [1197, 1195, 1764, 1151]. Uncertainty [35, 1180, 1144, 1198, 264, 1199, 2427, 1543]. Unconditional [1033, 2520]. Unconditionally [193]. Uncovered [1692]. Understanding [1760, 1805, 2637, 2073, 91, 1827, 1826, 1825, 2051, 131, 58, 1683, 1810]. Undirected [1878, 1877]. Unfair [1469]. Unfittest [1747]. Unfolding [1633, 1261, 1262, 1263]. Unicast [2198, 2206, 2193]. UNICORE [1518, 442]. Unidirectional [2531]. Unification [1006, 1134]. Unified [1661, 1898, 2646, 2100]. Uniform [1135, 1094, 362, 478, 427]. Uniform-Memory-Access [1094]. Unifying [1154, 480, 867]. Unique [1469]. Unit [1240, 554, 1363, 2035]. Units [2036, 1757, 2051]. Universal [506, 1905, 2078, 2077, 1005, 988, 1007, 971, 1759]. Universality [557]. Universally [1029]. Universe [1815]. University [1947, 2381]. Unix [748]. Unknown [529]. Unlabeled [682]. Unmanned [2446, 1611]. Unordered [998]. Unpredictability [1039]. Unrelated [1851]. Unremitting [1810]. Unrestricted [1353]. Unsupervised [2115, 621, 665, 2068, 383, 613]. Unused [1427]. Update [2286, 2312]. Updating [2085]. Upper [920]. Urban [2493, 1975]. USA [2295]. USA-Advisor [2295]. Usability [474]. Usage [27]. Use [2654, 1524, 2066, 1806, 2204, 1958, 295, 1855, 2653, 2255, 345, 1553, 1080, 119, 1364]. Used [100]. User [2343, 146, 3, 1918, 2364, 2361, 1734, 2124, 29, 181, 2161, 2162, 2165, 112, 2559, 1505, 1496, 864, 1080, 2044]. User-Adapted [2361]. User-Artefact [112]. User-Defined [2124, 2161, 2162, 2165]. User-Level [864]. User-Profiled [2559]. Users [1974, 774, 111, 773, 2241, 130]. Uses [227]. Using [841, 361, 1978, 2424, 649, 818, 492, 1457, 2115, 682, 797, 800, 793, 1313, 836, 870, 1094, 1642, 2270, 1923, 228, 1220, 1244, 2048, 2455, 2639, 1482, 1319, 341, 663, 317, 1383, 460, 461, 326, 126, 582, 1238, 1638, 792, 596, 5, 1993, 2036, 2118, 225, 1538, 1159, 963, 535, 1068, 590, 580, 2513, 127, 499, 438, 824, 284, 603, 810, 1299, 1637, 487, 1051, 1963, 75, 214, 741, 672, 1613, 1980, 1693, 531, 2075, 2131, 662, 1949, 1401, 1322, 2033, 1541, 2478, 652, 365, 2079, 169, 962, 548, 391, 1235, 687, 250, 335, 1700]. Using [288, 369, 598, 2083, 2138, 2618, 198, 1912, 1176, 602, 683, 502, 2607, 591, 2458, 843, 1799, 568, 2493, 1935, 2545, 1287, 352, 2507, 2020, 857, 1939, 1216, 342, 2566, 1055, 2143, 2235, 718, 319, 132, 2241, 2051, 2026, 1437, 31, 1975, 1230, 378, 1459, 1227, 1554, 2353, 2236, 2147, 533, 743, 1957, 749, 1158, 353, 1787, 574, 2648, 1868, 1565, 1337, 1328, 1677, 606, 671, 608, 955, 541, 1730, 738, 1789, 2096, 1801, 1729, 1202, 645, 2097, 360, 577, 130, 2098, 2044, 832, 2150, 59, 2269, 2100, 309, 2101, 1338, 688, 653, 306, 2670, 2549, 629, 1342, 2674]. Using [1349, 720, 367, 1249, 1082, 455, 799, 1924, 2106]. Utility [1667, 1153]. Utilizing [2046, 128]. Utterance [39, 2035]. Utterances [574]. Uzawa [2508].

v2.0 [1041]. v6 [1731]. Valency [2015]. Validated [1650]. Validating [2251]. Validation [1381, 2641, 1707, 285]. Valuation [1532, 1165]. Value [456, 2471, 154, 1531, 2318, 1150, 1149, 1810]. Valued [1212, 1638, 214, 635, 2281, 2145].

Values [1169, 2092]. VAMP [1240]. Variability [2080]. Variable [1394, 547, 1117, 2506, 527, 2041, 1464, 1846]. Variables [2595, 1988, 1158, 998]. Variance [526, 538]. Variational [532, 805, 342, 826]. Variations [919]. Variorum [1921]. Various [2435, 166, 1677]. Varying [552]. VC [69, 554]. VC-Dimension [69]. Vector [562, 1277, 1961, 312, 497, 1054, 1500, 573, 575, 2084, 563, 564, 522, 576, 574, 368, 1290, 572]. Vectors [2108, 2154, 294, 1281]. vehicle [1853]. Vehicles [2446, 1611]. Velocimetry [339]. Verb [2005]. Verbs [2015]. Verifiable [193]. Verification [1609, 419, 1217, 1240, 892, 1245, 2440, 1242, 1237, 1236, 1246, 1072, 1563, 1232, 1243, 1235, 1645, 1074, 1231, 1622, 1230, 1787, 1080, 1342, 1249, 2689, 1631]. Verifiers [1050]. Verify [1244]. Verifying [1047, 2458]. Verilog [1248]. Version [654, 439, 1190, 445, 2663]. Versions [894]. Versus [996, 876, 2679, 450, 2099]. Vertex [1872]. Vertex-Disjoint [1872]. Very [2038]. Vessel [332]. via [2413, 2179, 400, 2466, 489, 813, 844, 1464, 1962, 1380, 1422]. Vibration [2512]. Victorian [96]. Video [178, 2215, 663, 2196, 2213, 313, 665, 659, 386, 1666, 2617, 1664, 2618, 372, 2197, 1717, 314, 1718, 1729, 949, 949]. Video-Based [386]. Video-on-Demand [1717]. Videomicroscopy [1180]. Videos [2615, 667]. View [1497, 1155, 1215, 1564, 142, 2236]. Viewpoint [2429]. Views [15]. Virtex [1302, 1336, 1330]. Virtual [2323, 800, 2233, 2214, 90, 84, 117, 87, 2210, 86, 1461, 711, 1403, 106, 695, 2557, 91, 1329, 849, 2351, 2359, 607, 1725, 2533, 2241, 2581, 1535, 226, 2383, 799, 221]. vis [111]. vis-à-vis [111]. Viscoelastic [652]. Viscous [343]. Vision [2368, 827, 351, 666, 677, 835]. Visions [95]. Visitor [2665]. Visual [2603, 2615, 1315, 86, 641, 325, 295, 658, 23, 1808, 301, 375, 2608, 2436, 2110, 2156, 678, 521].

Visualisation [2547]. Visualising [1805]. Visualization [2639, 580, 1530, 1412, 1355, 1807]. Visualize [577]. Visually [2055, 2054, 1790]. Visually-Based [1790]. VizzScheduler [1412]. VLSI [1343, 1886]. Vocabulary [2033, 2029]. Vocoders [1332]. VoD [2453]. Voice [2209]. VoIP [1719]. Volume [2437, 1740, 373]. Voronoi [1597]. Voting [326, 549, 294]. Voting-Merging [549]. Vowel [681, 574]. Vs [2460, 2341, 1184, 900, 1188, 488, 2506, 1434]. VSS [1057].

Wagner [1125]. Wake [1463]. Wake-Up [1463]. Walk [363]. Walks [925]. Wall [1021]. Want [2076]. Warehouse [2563]. Warehouses [142, 252, 253]. Warehousing [737]. Warp [738]. Warping [2618]. Was [100, 96, 2410]. Watchpoint [396]. Watermark [956, 1919]. Watermarked [955]. Watermarking [950, 951, 947, 944, 948, 960]. Watermarks [959]. Wave [460, 2507]. Wavelet [605, 1666, 342, 1356, 307, 239]. Wavelet-Based [239]. Wavelets [361]. Waves [1738]. Way [1432, 1003, 909, 913, 241]. Ways [117]. WDM [2174, 2173, 2171, 2170]. Weak [888, 2341]. Weaker [888]. Weather [1457]. Web [489, 800, 15, 1568, 2645, 2056, 2258, 2324, 3, 2063, 157, 437, 225, 1877, 9, 2391, 70, 1534, 1560, 2229, 29, 2131, 159, 1514, 1544, 1938, 2556, 687, 2386, 2046, 2018, 24, 136, 2360, 1545, 2559, 2, 2380, 30, 1682, 2432, 27, 2254, 1754, 131, 2250, 59, 2099, 144, 133, 455, 799]. Web-Accessible [437]. Web-Based [800, 2056, 157, 2391, 1560, 2386, 2380, 30, 2432, 2254, 2250, 455, 799]. Web-Oriented [2559]. Web-Pages [2018]. WebCluster [437]. Weight [1017, 548, 2040]. Weighted [494, 317, 1022, 2191, 1473, 149, 1855, 844, 1677, 832]. Weil [1040]. Well [2265, 1133, 2266]. Well-Founded [2265, 1133, 2266]. WES [1937]. WFMS [157]. WFQ [2200]. Wheeled [1793]. While [2064, 86]. Whole [1392]. Whom [56]. Wide [1522, 2362, 1814, 1463, 1754, 1820, 2251]. Wide-Area [1522]. Wideband [2177]. Width [920]. Wigner [364]. Wilf [919]. Wind [561]. Wind-Field [561]. Window [611, 1677]. Window-Based [1677]. Windows [810, 387, 1904]. Wireless [2179, 1873, 127, 2584, 1715, 2178, 1730, 182, 1687]. Wireless-Oriented [127]. Wise [1193]. Within [2121, 2452, 205, 789, 237, 2576, 1996, 1194, 1173, 788, 2659]. Within-Group [2121]. without [1866, 526, 18, 1775, 1129, 1778]. Wizard [2558]. WLAN [717]. WMPI [706]. Word [1998, 2012, 204, 2019, 989, 918]. WordNet [2020]. Words [903, 2006, 1175, 931, 2011]. WordSieve [36]. Work [116, 2167, 2212, 1935, 1536, 1549, 221]. Workbench [2013, 1660]. Worker [1508, 450]. Workflow [157, 2626, 53, 30, 2581, 1966, 175, 767]. Workflows [2252, 1535]. Working [100, 86]. workload [1432]. Works [106, 1924]. Workshops [2655]. Workspace [116]. Workstations [741, 1503, 744, 734]. World [341, 953, 438, 283, 1814, 1754, 1820]. World-Wide-Mind [1814, 1820]. Worldmaking [117]. Worlds [2214, 2304, 638]. Woven [103]. Wrappers [1319, 1243]. Wrapping [2258, 2063]. Wreath [1267, 1268]. Writing [1840]. WS1S [1126]. WSD [2020]. WTA [633]. WWW [146, 860].

X [1357]. X-MatchPRO [1357]. XML [2048, 237, 2252, 5, 127, 7, 277, 26, 138, 155, 139, 25, 1081, 6, 1559, 153]. XML-Represented [237]. XML-Transformations [7]. XMLSpaces [2250]. XSL [1946]. XSL-based [1946].

XSLT [1923]. XVCL [2660].

Year [2375]. Years [1881]. Yoda [2254].

Zero [956, 483, 1060, 1498]. Zero-Knowledge [956]. Zero-Safe [483]. Zero-Tree [1498]. Zhurnal’nykh [1950]. Zürich [2540].

References

Werthner:2001:JBS

[1] Hannes Werthner. Just business — shouldn’t we have some fun? Lecture Notes in Computer Science, 2115: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                              tronic). URL http://link.springer-
                                              ny.com/link/service/series/0558/
                                              bibs/2115/21150001.htm;   http:/
                                              /link.springer-ny.com/link/service/
                                              series/0558/papers/2115/21150001.
                                              pdf.

Pastor:2001:OOA

[2] Oscar Pastor, Silvia Abrahão, and Joan Fons. An object-oriented approach to automate Web applications development. Lecture Notes in Computer Science, 2115:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                              tronic). URL http://link.springer-
                                              ny.com/link/service/series/0558/
                                              bibs/2115/21150016.htm;   http:/
                                              /link.springer-ny.com/link/service/
                                              series/0558/papers/2115/21150016.
                                              pdf.

Brisaboa:2001:TDU

[3] Nieves R. Brisaboa, Miguel R. Penabad, Ángeles S. Places, and Francisco J. Rodríguez. Tools for the design of user friendly Web applications. Lecture Notes in Computer Science, 2115: 29–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150029.
pdf.

Rashid:2001:ECB

[4] Awais Rashid, Ruzanna Chitchyan, Andreas Speck, and Elke Pulvermueller. EProMS: An E-commerce based process model for cooperative software development in small organisations. Lecture Notes in Computer Science, 2115:39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2115/21150039.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2115/21150039.pdf.

Chung:2001:EOO

[5] Tae-Sun Chung, Sangwon Park, Sang-Young Han, and Hyoung-Joo Kim. Extracting object-oriented database schemas from XML DTDs using inheritance. Lecture Notes in Computer Science, 2115:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150049.
pdf.

Vittori:2001:CXD

[6] César M. Vittori, Carina F. Dorneles, and Carlos A. Heuser. Creating XML documents from relational data sources. Lecture Notes in Computer Science, 2115:60–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150060.
                                        pdf.

Eder:2001:CXT

[7] Johann Eder and Walter Strametz. Composition of XML-transformations. Lecture Notes in Computer Science, 2115:71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2115/21150071.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150071.
                                        pdf.

Abrazhevich:2001:CCE

[8] Dennis Abrazhevich. Classification and characteristics of electronic payment systems. Lecture Notes in Computer Science, 2115:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150081.
                                        pdf.

Dani:2001:CFE

[9] A. R. Dani and P. Radha Krishna. An E-check framework for electronic payment systems in the Web based environment. Lecture Notes in Computer Science, 2115:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150091.
pdf.

Christoffel:2001:TSI

[10] Michael Christoffel, Thorwald Franke, and Stefan Kotkamp. Tradersupported information markets — A simulation study. Lecture Notes in Computer Science, 2115:101–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150101.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150101.
pdf.

Yu:2001:IFB

[11] Chien-Chih Yu. An integrated framework of business models for guiding electronic commerce applications and case studies. Lecture Notes in Computer Science, 2115:111–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150111.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2115/21150111.
                                        pdf.

Wang:2001:MPS

[12] Guandong Wang and Amitabha Das. Models and protocol structures for software agent based complex Ecommerce transactions. Lecture Notes in Computer Science, 2115:121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150121.
                                        pdf.

Cannataro:2001:MAM

[13] Mario Cannataro, Alfredo Cuzzocrea, and Andrea Pugliese. A multidimensional approach for modelling and supporting adaptive hypermedia systems. Lecture Notes in Computer Science, 2115:132–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150132.
                                        pdf.

Wombacher:2001:MIS

[14] Andreas Wombacher and Karl Aberer. Modelling the ICE standard with a formal language for information commerce. Lecture Notes in Computer Science, 2115:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2115/21150142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150142.
pdf.

Arantes:2001:MWD

[15] Álisson R. Arantes, Alberto H. F. Laender, Paulo B. Golgher, and Altigran S. da Silva. Managing Web data through views. Lecture Notes in Computer Science, 2115:154–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150154.
pdf.

Katzenbeisser:2001:AIS

[16] Stefan Katzenbeisser and Philipp Tomsich. Applied information security for m-commerce and digital television environments. Lecture Notes in Computer Science, 2115:166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150166.
pdf.

Chang:2001:FAM

[17] Kyung-Ah Chang, Byung-Rae Lee, and Tai-Yun Kim. Flexible authentication with multiple domains of electronic commerce. Lecture Notes in Computer Science, 2115:176–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150176.
                                        pdf.

Magkos:2001:ATS

[18] Emmanouil Magkos, Panayiotis Kotzanikolaou, and Vassilios Chrissikopoulos. An asymmetric traceability scheme for copyright protection without trust assumptions. Lecture Notes in Computer Science, 2115:186–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150186.
                                        pdf.

Rebstock:2001:AAS

[19] Michael Rebstock. An application architecture for supporting interactive bilateral electronic negotiations. Lecture Notes in Computer Science, 2115:196– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150196.
                                        pdf.

Deschner:2001:SSA

[20] Dominik Deschner, Florian Lang, and Freimut Bodendorf. Strategies for software agent based multiple issue negotiations. Lecture Notes in Computer Science, 2115:206–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150206.
pdf.

Sung:2001:ACO

[21] Wing-Kin Sung, David Yang, Siu-Ming Yiu, Wai-Shing Ho, David Cheung, and Tak-Wah Lam. Automatic construction of online catalog topologies. Lecture Notes in Computer Science, 2115: 216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150216.
pdf.

Omelayenko:2001:TLI

[22] Borys Omelayenko and Dieter Fensel. A two-layered integration approach for product information in B2B Ecommerce. Lecture Notes in Computer Science, 2115:226–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2115/21150226.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2115/21150226.pdf.

Lee:2001:VOP

[23] Juhnyoung Lee, Priscilla Wang, and Ho Soo Lee. A visual onepage catalog interface for analytical product selection. Lecture Notes in Computer Science, 2115:240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150240.
pdf.

Li:2001:EHP

[24] Wen-Syan Li, K. Selçuk Candan, Wang-Pin Hsiung, Oliver Po, and Divyakant Agrawal. Engineering high performance database-driven Ecommerce Web sites through dynamic content caching. Lecture Notes in Computer Science, 2115:250–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150250.
pdf.

Nicolle:2001:XEM

[25] Christophe Nicolle and Kokou Yétongnon. XML enabled metamodeling and tools for cooperative information systems. Lecture Notes in Computer Science, 2115:260–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2115/21150260.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2115/21150260.
pdf.

Graupner:2001:SXD

[26] Sven Graupner, Wooyoung Kim, Akhil Sahai, and Dmitry Lenkov. E-Speak — an XML document interchange engine. Lecture Notes in Computer Science, 2115:270–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150270.
pdf.

Shahabi:2001:FMM

[27] Cyrus Shahabi, Farnoush Banaei-Kashani, Jabed Faruque, and Adil Faisal. Feature matrices: A model for efficient and anonymous Web usage mining. Lecture Notes in Computer Science, 2115:280–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2115/21150280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2115/21150280.
pdf.

Loney:2001:FPM

[28] Fred N. Loney. Faceted preference matching in recommender systems. Lecture Notes in Computer Science, 2115:295–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2115/21150295.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2115/21150295.
pdf.

Helmy:2001:PWS

[29] Tarek Helmy, Satoshi Amamiya, and Makoto Amamiya. Pinpoint Web searching and user modeling on the collaborative kodama agents. Lecture Notes in Computer Science, 2115:305– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150305.
                                        pdf.

Quirchmayr:2001:AWA

[30] Gerald Quirchmayr, Beate List, and A. Min Tjoa. Analyzing workflow audit trails in Web-based environments with fuzzy logic. Lecture Notes in Computer Science, 2115:315–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2115/21150315.
                                        pdf.

Qiu:2001:UHC

[31] Zhanzi Qiu, Matthias Hemmje, and Erich J. Neuhold. Using hypertext composites in structured query and search. Lecture Notes in Computer Science, 2115:326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2115/21150326.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2115/21150326.
pdf.

Buhse:2001:CDM

[32] Willms Buhse. Categorizing distribution model scenarios for online music. Lecture Notes in Computer Science, 2115:337–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2115/21150337.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2115/21150337.
pdf.

Anonymous:2001:AIfe

[33] Anonymous. Author index. Lecture Notes in Computer Science, 2115:347– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2115/2115auth.pdf.

Costa:2001:TOB

[34] Horacio Arló Costa. Trade-offs between inductive power and logical omniscience in modeling context. Lecture Notes in Computer Science, 2116:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160001.
pdf.

Barnden:2001:UCH

[35] John A. Barnden. Uncertainty and conflict handling in the ATT-meta context-based system for metaphorical reasoning. Lecture Notes in Computer Science, 2116:15–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160015.
pdf.

Bauer:2001:WMR

[36] Travis Bauer and David B. Leake. WordSieve: A method for real-time context extraction. Lecture Notes in Computer Science, 2116:30–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160030.
pdf.

Bell:2001:PRP

[37] John Bell. Pragmatic reasoning. pragmatic semantics and semantic pragmatics. Lecture Notes in Computer Science, 2116:45–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160045.
pdf.

Benerecetti:2001:DCD

[38] Massimo Benerecetti, Paolo Bouquet, and Chiara Ghidini. On the dimensions of context dependence: Partiality, approximation, and perspective. Lecture Notes in Computer Science, 2116:59– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160059.
pdf.

Bianchi:2001:CUI

[39] Claudia Bianchi. Context of utterance and intended context. Lecture Notes in Computer Science, 2116:73– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160073.
pdf.

Bouquet:2001:TFC

[40] P. Bouquet and L. Serafini. Two formalizations of context: A comparison. Lecture Notes in Computer Science, 2116:87–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2116/21160087.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2116/21160087.
pdf.

Cohan:2001:CAF

[41] Jocelyn Cohan. Consider the alternatives: Focus in contrast and context. Lecture Notes in Computer Science, 2116:102–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2116/21160102.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160102.
                                        pdf.

Connolly:2001:CSH

[42] John H. Connolly. Context in the study of human languages and computer programming languages: A comparison. Lecture Notes in Computer Science, 2116:116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160116.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160116.
                                        pdf.

Dastani:2001:MCE

[43] Mehdi Dastani and Bipin Indurkhya. Modeling context effect in perceptual domains. Lecture Notes in Computer Science, 2116:129–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160129.
                                        pdf.

Edmonds:2001:LAC

[44] Bruce Edmonds. Learning appropriate contexts. Lecture Notes in Computer Science, 2116:143–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160143.
pdf.

Ekbia:2001:CRP

[45] Hamid R. Ekbia and Ana G. Maguitman. Context and relevance: A pragmatic approach. Lecture Notes in Computer Science, 2116:156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160156.
pdf.

Ferrario:2001:CR

[46] Roberta Ferrario. Counterfactual reasoning. Lecture Notes in Computer Science, 2116:170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160170.
pdf.

Romacker:2001:CBA

[47] Martin Romacker and Udo Hahn. Context-based ambiguity management for natural language processing. Lecture Notes in Computer Science, 2116: 184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2116/21160184.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160184.
                                        pdf.

Henninger:2001:CSA

[48] Amy E. Henninger, Avelino J. Gonzalez, Michael Georgiopoulos, and Ronald F. DeMara. A connectionistsymbolic approach to modeling agent behavior: Neural networks grouped by contexts. Lecture Notes in Computer Science, 2116:198–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2116/21160198.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2116/21160198.pdf.

Juttner:2001:CDP

[49] Martin Jüttner and Ingo Rentschler. Context dependency of patterncategory learning. Lecture Notes in Computer Science, 2116:210–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160210.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160210.
                                        pdf.

Kokinov:2001:SCE

[50] Boicho Kokinov and Maurice Grinberg. Simulating context effects in problem solving with AMBR. Lecture Notes in Computer Science, 2116:221– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2116/21160221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160221.
pdf.

Langlois:2001:NMB

[51] David Langlois, Kamel Smaïli, and Jean-Paul Haton. A new method based on context for combining statistical language models. Lecture Notes in Computer Science, 2116:235– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160235.
pdf.

Matsui:2001:EPT

[52] Tomoko Matsui. Experimental pragmatics: Towards testing relevancebased predictions about anaphoric bridging inferences. Lecture Notes in Computer Science, 2116:248–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160248.
pdf.

Maus:2001:WCM

[53] Heiko Maus. Workflow context as a means for intelligent information support. Lecture Notes in Computer Science, 2116:261–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160261.
                                        pdf.

Motschnig-Pitrik:2001:RMC

[54] Renate Motschnig-Pitrik and Ladislav Nykl. The role and modeling of context in a cognitive model of Rogers’ person-centred approach. Lecture Notes in Computer Science, 2116:275– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160275.
                                        pdf.

Penco:2001:LH

[55] Carlo Penco. Local holism. Lecture Notes in Computer Science, 2116:290– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160290.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160290.
                                        pdf.

Stojanovic:2001:WPE

[56] Isidora Stojanovic. Whom is the problem of the essential indexical a problem for? Lecture Notes in Computer Science, 2116:304–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2116/21160304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160304.
pdf.

Tijus:2001:CCC

[57] Charles Tijus. Contextual categorization and cognitive phenomena. Lecture Notes in Computer Science, 2116:316– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160316.
pdf.

Turner:2001:RGC

[58] Elise H. Turner and Roy M. Turner. Representing the graphics context to support understanding plural anaphora in multi-modal interfaces. Lecture Notes in Computer Science, 2116:330– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160330.
pdf.

Turner:2001:UEP

[59] Roy M. Turner, Elise H. Turner, Thomas A. Wagner, Thomas J. Wheeler, and Nancy E. Ogle. Using explicit, A priori contextual knowledge in an intelligent Web search agent. Lecture Notes in Computer Science, 2116: 343–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160343.
                                        pdf.

Vassallo:2001:CPP

[60] Nicla Vassallo. Contexts and philosophical problems of knowledge. Lecture Notes in Computer Science, 2116: 353–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160353.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160353.
                                        pdf.

Wache:2001:PCT

[61] Holger Wache and Heiner Stuckenschmidt. Practical context transformation for information system interoperability. Lecture Notes in Computer Science, 2116:367–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160367.
                                        pdf.

Young:2001:EC

[62] R. A. Young. Explanation as contextual. Lecture Notes in Computer Science, 2116:381–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2116/21160381.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2116/21160381.pdf.

Zibetti:2001:CCM

[63] Elisabetta Zibetti, Vicenç Quera, Francesc Salvador Beltran, and Charles Tijus. Contextual categorization: A mechanism linking perception and knowledge in modeling and simulating perceived events as actions. Lecture Notes in Computer Science, 2116: 395–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160395.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160395.
pdf.

Adams:2001:MC

[64] William A. Adams. The motivational context. Lecture Notes in Computer Science, 2116:409–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160409.
pdf.

Boella:2001:AAB

[65] Guido Boella and Leonardo Lesmo. An approach to anaphora based on mental models. Lecture Notes in Computer Science, 2116:413–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160413.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160413.
                                        pdf.

Bosco:2001:CMM

[66] Cristina Bosco and Carla Bazzanella. Context and multi-media corpora. Lecture Notes in Computer Science, 2116: 417–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160417.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160417.
                                        pdf.

Bovens:2001:BEC

[67] Luc Bovens and Stephan Hartmann. Belief expansion, contextual fit, and the reliability of information sources. Lecture Notes in Computer Science, 2116:421–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160421.
                                        pdf.

Chevalier:2001:RCA

[68] Aline Chevalier and Laure Martinez. The role of the context in the acquisition and in the organisation of knowledge: Studies from adults and from children. Lecture Notes in Computer Science, 2116:425–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2116/21160425.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160425.
pdf.

Ciskowski:2001:VDC

[69] Piotr Ciskowski. VC-dimension of a context-dependent perceptron. Lecture Notes in Computer Science, 2116:429– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160429.
pdf.

Dichev:2001:FCD

[70] Christo Dichev. A framework for context-driven Web resource discovery. Lecture Notes in Computer Science, 2116:433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2116/21160433.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2116/21160433.
pdf.

Etcheverry:2001:SCC

[71] Patrick Etcheverry, Philippe Lopistéguy, and Pantxika Dagorret. Specifying contexts for coordination patterns. Lecture Notes in Computer Science, 2116:437–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2116/21160437.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2116/21160437.pdf.

Fayard:2001:LST

[72] Anne-Laure Fayard and Austin Henderson. Looking at “situated” technology: Differences in pattern of interaction reflect differences in context. Lecture Notes in Computer Science, 2116: 441–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160441.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160441.
pdf.

Fernandez-Breis:2001:CDA

[73] J. T. Fernández-Breis, Rafael Valencia-García, Rodrigo Martínez-Béjar, and Pascual Cantos-Gómez. A contextdriven approach for knowledge acquisition: Application to a leukaemia domain. Lecture Notes in Computer Science, 2116:445–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2116/21160445.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2116/21160445.pdf.

Fetzer:2001:CNL

[74] Anita Fetzer. Context in naturallanguage communication: Presupposed or co-supposed? Lecture Notes in Computer Science, 2116:449–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2116/21160449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160449.
pdf.

Gonzalez:2001:UCC

[75] Avelino J. Gonzalez and Shinya Saeki. Using contexts competition to model tactical human behavior in a simulation. Lecture Notes in Computer Science, 2116:453–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160453.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160453.
pdf.

Klemke:2001:CSD

[76] Roland Klemke and Achim Nick. Case studies in developing contextualising information systems. Lecture Notes in Computer Science, 2116:457–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2116/21160457.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2116/21160457.
pdf.

Pomerol:2001:ASR

[77] J.-C. Pomerol and P. Brézillon. About some relationships between knowledge and context. Lecture Notes in Computer Science, 2116:461–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2116/21160461.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160461.
                                        pdf.

Richards:2001:CCR

[78] Debbie Richards. Combining cases and rules to provide contextualised knowledge based systems. Lecture Notes in Computer Science, 2116:465– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2116/21160465.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2116/21160465.
                                        pdf.

Anonymous:2001:AIff

[79] Anonymous. Author index. Lecture Notes in Computer Science, 2116:471– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2116/2116auth.pdf.

Gorayska:2001:CTT

[80] Barbara Gorayska, Jonathon P. Marsh, and Jacob L. Mey. Cognitive technology: Tool or instrument? Lecture Notes in Computer Science, 2117:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170001.pdf.

Clark:2001:NBC

[81] Andy Clark. Natural-born cyborgs? Lecture Notes in Computer Science, 2117:17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170017.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170017.pdf.

Loomes:2001:FAR

[82] Martin Loomes and Chrystopher L. Nehaniv. Fact and artifact: Reification and drift in the history and growth of interactive software systems. Lecture Notes in Computer Science, 2117:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170025.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170025.pdf.

Tuikka:2001:TTC

[83] Tuomo Tuikka and Kari Kuutti. Thinking together in concept design for future products — emergent features for computer support. Lecture Notes in Computer Science, 2117:40–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170040.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170040.pdf.

Biocca:2001:SCT

[84] Frank Biocca. The space of cognitive technology: The design medium and cognitive properties of virtual space. Lecture Notes in Computer Science, 2117:55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170055.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170055.pdf.

Werry:2001:CSI

[85] Iain Werry, Kerstin Dautenhahn, Bernard Ogden, and William Harwin. Can social interaction skills be taught by a social agent? the role of a robotic mediator in autism therapy. Lecture Notes in Computer Science, 2117:57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170057.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170057.pdf.

Day:2001:CED

[86] Philip N. Day, Patrik O’Brian Holt, and George T. Russell. The cognitive effects of delayed visual feedback: Working memory disruption while driving in virtual environments. Lecture Notes in Computer Science, 2117:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170075.
htm;   http://link.springer.de/

link/service/series/0558/papers/
2117/21170075.pdf.

Chan:2001:EPV

[87] Melanie Chan. Embodiment, perception, and virtual reality. Lecture Notes in Computer Science, 2117: 83–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2117/21170083.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2117/21170083.
pdf.

teBoekhorst:2001:FMC

[88] I. René J. A. te Boekhorst. Freeing machines from Cartesian chains. Lecture Notes in Computer Science, 2117:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170095.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170095.pdf.

Morikawa:2001:RBA

[89] Osamu Morikawa and Takanori Maesako. The relationship between the arrangement of participants and the comfortableness of conversation in HyperMirror. Lecture Notes in Computer Science, 2117:109–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170109.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170109.pdf.

Biocca:2001:MSA

[90] Frank Biocca, David Lamas, Ping Gai, and Robert Brady. Mapping the semantic asymmetries of virtual and augmented reality space (extended abstract). Lecture Notes in Computer Science, 2117:117–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170117.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170117.pdf.

Jelfs:2001:PRA

[91] Anne Jelfs and Denise Whitelock. Presence and the role of activity theory in understanding: How students learn in virtual learning environments. Lecture Notes in Computer Science, 2117:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170123.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170123.pdf.

Gooding:2001:EII

[92] David C. Gooding. Experiment as an instrument of innovation: Experience and embodied thought. Lecture Notes in Computer Science, 2117:130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170130.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170130.pdf.

Halloran:2001:CWA

[93] John Halloran. Can we afford it? issues in designing transparent technologies. Lecture Notes in Computer Science, 2117:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170141.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170141.pdf.

Ali:2001:SED

[94] Syed Mustafa Ali. “The End of The (Dreyfus) Affair”: (post)Heideggerian meditations on man, machine, and meaning. Lecture Notes in Computer Science, 2117:149–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170149.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170149.pdf.

Chimir:2001:NVO

[95] Igor Chimir and Mark Horney. New visions of old models. Lecture Notes in Computer Science, 2117: 157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2117/21170157.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2117/21170157.
pdf.

Campbell-Kelly:2001:AVD

[96] Martin Campbell-Kelly. Victorian data processing — when software was people (abstract). Lecture Notes in Computer Science, 2117:164, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170164.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170164.pdf.

Tenenberg:2001:MCP

[97] Josh Tenenberg. On the meaning of computer programs. Lecture Notes in Computer Science, 2117: 165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2117/21170165.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2117/21170165.
pdf.

Lunzer:2001:SSR

[98] Aran Lunzer and Yuzuru Tanaka. Sense from a sea of resources: Tools to help people piece information together. Lecture Notes in Computer Science, 2117:175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170175.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170175.pdf.

Talbott:2001:BAM

[99] Steve Talbott. Beyond the algorithmic mind. Lecture Notes in Computer Science, 2117:190–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170190.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170190.pdf.

Barker:2001:HGW

[100] Trevor Barker and Janet Barker. How group working was used to provide a constructive computerbased learning environment. Lecture Notes in Computer Science, 2117: 203–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2117/21170203.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2117/21170203.
pdf.

Derkach:2001:NPC

[101] Lydia Derkach. Neuro-psycho-computational technology in human cognition under bilingualism. Lecture Notes in Computer Science, 2117:214–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170214.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170214.pdf.

Hokanson:2001:DIC

[102] Brad Hokanson. Digital image creation and analysis as a means to examine learning and cognition. Lecture Notes in Computer Science, 2117:226–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170226.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170226.pdf.

Gerdt:2001:WSC

[103] Petri Gerdt, Piet Kommers, Chee-Kit Looi, and Erkki Sutinen. Woven stories as a cognitive tool. Lecture Notes in Computer Science, 2117:233–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170233.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170233.pdf.

Dautenhahn:2001:NIH

[104] Kerstin Dautenhahn. The narrative intelligence hypothesis: In search of the transactional format of narratives in humans and other social animals. Lecture Notes in Computer Science, 2117:248–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170248.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170248.pdf.

Goldstein:2001:BR

[105] Ronnie Goldstein, Ivan Kalas, Richard Noss, and Dave Pratt. Building rules. Lecture Notes in Computer Science, 2117:267–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/

series/0558/bibs/2117/21170267.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170267.pdf.

Galitsky:2001:VMS

[106] Boris Galitsky. Virtual mental space: Interacting with the characters of works of literature. Lecture Notes in Computer Science, 2117:282–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170282.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170282.pdf.

duBoulay:2001:PPI

[107] Benedict du Boulay and Rosemary Luckin. The plausibility problem: An initial analysis. Lecture Notes in Computer Science, 2117: 289–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2117/21170289.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2117/21170289.
pdf.

Stojanov:2001:CIC

[108] Georgi Stojanov and Kire Stojanoski. Computer interfaces: From communication to mind-prosthesis metaphor. Lecture Notes in Computer Science, 2117:301–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170301.

                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170301.pdf.

Riedl:2001:MR

[109] Reinhard Riedl. Meaning and relevance. Lecture Notes in Computer Science, 2117:311–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170311.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170311.pdf.

Blackwell:2001:CDN

[110] A. F. Blackwell, C. Britton, A. Cox, T. R. G. Green, C. Gurr, G. Kadoda, M. S. Kutar, M. Loomes, C. L. Nehaniv, M. Petre, C. Roast, C. Roe, A. Wong, and R. M. Young. Cognitive dimensions of notations: Design tools for cognitive technology. Lecture Notes in Computer Science, 2117:325–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170325.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170325.pdf.

Kutar:2001:CDA

[111] Maria S. Kutar, Chrystopher L. Nehaniv, Carol Britton, and Sara Jones. The cognitive dimensions of an artifact vis-à-vis individual human users: Studies with notations for the temporal specification of interactive systems. Lecture Notes in Computer Science, 2117:342–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170342.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170342.pdf.

MeurigBeynon:2001:ISM

[112] Meurig Beynon, Chris Roe, Ashley Ward, and Allan Wong. Interactive situation models for cognitive aspects of user-artefact interaction. Lecture Notes in Computer Science, 2117:356–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170356.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170356.pdf.

Donath:2001:MF

[113] Judith Donath. Mediated faces. Lecture Notes in Computer Science, 2117:373–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170373.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170373.pdf.

Gurr:2001:ICI

[114] Corin Gurr and Gillian Hardstone. Implementing configurable information systems: A combined social science and cognitive science approach. Lecture Notes in Computer Science, 2117:391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170391.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170391.pdf.

Riedl:2001:IEI

[115] Reinhard Riedl. Interdisciplinary engineering of interstate E-government solutions. Lecture Notes in Computer Science, 2117:405–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170405.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170405.pdf.

Brophy:2001:WWW

[116] Richard Brophy and Will Venters. Work, workspace, and the workspace portal. Lecture Notes in Computer Science, 2117:421–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170421.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170421.pdf.

Borders:2001:EPW

[117] Max Borders and Doug Bryan. Experimental politics: Ways of virtual worldmaking. Lecture Notes in Computer Science, 2117:432–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170432.
                                        htm;   http://link.springer.de/

link/service/series/0558/papers/
2117/21170432.pdf.

Pickering:2001:HIA

[118] John Pickering. Human identity in the age of software agents. Lecture Notes in Computer Science, 2117:442–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170442.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170442.pdf.

Syrjanen:2001:TIH

[119] Anna-Liisa Syrjänen. Tracing for the ideal hunting dog: Effects of development and use of information system on community knowledge. Lecture Notes in Computer Science, 2117:452–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2117/21170452.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2117/21170452.pdf.

Chan:2001:CPT

[120] Ho Mun Chan and Barbara Gorayska. Critique of pure technology. Lecture Notes in Computer Science, 2117: 463–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2117/21170463.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2117/21170463.
pdf.

MeurigBeynon:2001:CI

[121] Meurig Beynon, Yih-Chang Ch’en, Hsing-Wen Hseu, Soha Maad, Suwanna Rasmequan, Chris Roe, Jaratsri Rungrattanaubol, Steve Russ, Ashley Ward, and Allan Wong. The computer as instrument. Lecture Notes in Computer Science, 2117: 476–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/bibs/2117/21170476.htm;   http:
                                        //link.springer.de/link/service/
                                        series/0558/papers/2117/21170476.
                                        pdf.

Landauer:2001:CIE

[122] Christopher Landauer and Kirstie L. Bellman. Computational infrastructure for experiments in cognitive leverage. Lecture Notes in Computer Science, 2117:490–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2117/21170490.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2117/21170490.pdf.

Anonymous:2001:AIfg

[123] Anonymous. Author index. Lecture Notes in Computer Science, 2117: 521–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer.de/link/service/series/
                                        0558/papers/2117/2117auth.pdf.

Chang:2001:SMI

[124] Jae-Woo Chang and Yeon-Jung Kim. Spatial-match iconic image retrieval with ranking in multimedia databases. Lecture Notes in Computer Science, 2118:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180003.
pdf.

Atnafu:2001:SBO

[125] Solomon Atnafu, Lionel Brunie, and Harald Kosch. Similarity-based operators in image database systems. Lecture Notes in Computer Science, 2118: 14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180014.
pdf.

Chang:2001:EHD

[126] Jae-Woo Chang and Kwang-Taek Song. An efficient high-dimensional index structure using cell signatures for similarity search. Lecture Notes in Computer Science, 2118:26–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180026.
pdf.

Dittberner:2001:UOR

[127] Christoph Dittberner, Thomas Courvoisier, and Guntram Flach. Using object-relational database systems and XML in the context of a wireless-oriented multimedia information system. Lecture Notes in Computer Science, 2118:34–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180034.
pdf.

Yoshida:2001:UCB

[128] Tetsuya Yoshida, Daiki Shinkai, and Shogo Nishida. Utilizing the correlation between query keywords for information retrieval. Lecture Notes in Computer Science, 2118:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180049.
pdf.

Zhou:2001:AMB

[129] Shuigeng Zhou, Jihong Guan, Yunfa Hu, Jiangtao Hu, and Aoying Zhou. Adjacency matrix based full-text indexing models. Lecture Notes in Computer Science, 2118:60–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180060.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2118/21180060.
pdf.

Toda:2001:GOI

[130] Hiroyuki Toda, Toshifumi Enomoto, and Tetsuji Satoh. Goal-oriented information retrieval using feedback from users. Lecture Notes in Computer Science, 2118:72–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180072.
pdf.

Tang:2001:CWD

[131] Changjie Tang, Tong Li, Changyu Liu, and Yin Ge. Classify Web document by key phrase understanding. Lecture Notes in Computer Science, 2118:80– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180080.
pdf.

Perrizo:2001:DHC

[132] William Perrizo, Qin Ding, Qiang Ding, and Amalendu Roy. Deriving high confidence rules from spatial data using Peano count trees. Lecture Notes in Computer Science, 2118:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180091.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180091.
                                        pdf.

Wen:2001:DHC

[133] Catherine W. Wen, Huan Liu, Wilson X. Wen, and Jeffery Zheng. A distributed hierarchical clustering system for Web mining. Lecture Notes in Computer Science, 2118:103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180103.
                                        pdf.

Zhou:2001:ADM

[134] Haofeng Zhou, Beijun Ruan, Jianqiu Zhu, Yangyong Zhu, and Baile Shi. ARMiner: A data mining tool based on association rules. Lecture Notes in Computer Science, 2118:114– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180114.
                                        pdf.

Wang:2001:ICR

[135] Daling Wang, Yubin Bao, Xiao Ji, Guoren Wang, and Baoyan Song. An integrated classification rule management system for data mining. Lecture Notes in Computer Science, 2118:122– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2118/21180122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180122.
pdf.

Lim:2001:CWD

[136] Lipyeow Lim, Min Wang, Sriram Padmanabhan, Jeffrey Scott Vitter, and Ramesh Agarwal. Characterizing Web document change. Lecture Notes in Computer Science, 2118:133– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180133.
pdf.

Su:2001:ISS

[137] Hong Su, Sriram Padmanabhan, and Ming-Ling Lo. Identification of syntactically similar DTD elements for schema matching. Lecture Notes in Computer Science, 2118:145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180145.
pdf.

Li:2001:XBE

[138] Hongyan Li, Shiwei Tang, Dongqing Yang, and Yi Yu. An XML based electronic medical record integration system. Lecture Notes in Computer Science, 2118:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180160.
                                        pdf.

Liang:2001:EXD

[139] Yuqi Liang, Aoying Zhou, Shihui Zheng, Long Zhang, and Wenyun Ji. Enhancing XML data processing in relational system with indices. Lecture Notes in Computer Science, 2118:168– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180168.
                                        pdf.

Tan:2001:MDM

[140] Hongxing Tan, Wenwen Qi, and Longxiang Zhou. A multidimensional data model supporting complicated aggregations. Lecture Notes in Computer Science, 2118:179–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180179.
                                        pdf.

Lujan-Mora:2001:CSS

[141] Sergio Luján-Mora and Manuel Palomar. Comparing string similarity measures for reducing inconsistency in integrating data from different sources. Lecture Notes in Computer Science, 2118:191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180191.
pdf.

Liang:2001:RVM

[142] Weifa Liang and Jeffrey X. Yu. Revisit on view maintenance in data warehouses. Lecture Notes in Computer Science, 2118:203–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180203.
pdf.

Xu:2001:EHI

[143] Xuebiao Xu, Andrew C. Jones, Nick Pittas, W. Alex Gray, Nick J. Fiddian, Richard J. White, John Robinson, Frank A. Bisby, and Sue M. Brandt. Experiences with a hybrid implementation of a globally distributed federated database system. Lecture Notes in Computer Science, 2118:212– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180212.
pdf.

Wang:2001:LBC

[144] Yitong Wang and Masaru Kitsuregawa. Link based clustering of Web search results. Lecture Notes in Computer Science, 2118:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180225.
                                        pdf.

Jiang:2001:TTC

[145] Haifeng Jiang, Wenwu Lou, and Wei Wang. Three-tier clustering: An online citation clustering system. Lecture Notes in Computer Science, 2118:237– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180237.
                                        pdf.

Bouras:2001:ING

[146] Christos Bouras, Agisilaos Konidaris, Eleni Konidari, and Afrodite Sevasti. Introducing navigation graphs as a technique for improving WWW user browsing. Lecture Notes in Computer Science, 2118:249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180249.
                                        pdf.

Xu:2001:ECN

[147] Lai Xu and Hans Weigand. The evolution of the contract net protocol. Lecture Notes in Computer Science, 2118: 257–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180257.
pdf.

Bettini:2001:SCT

[148] Claudio Bettini. Semantic compression of temporal data. Lecture Notes in Computer Science, 2118:267–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180267.
pdf.

Lin:2001:ECW

[149] Xuemin Lin, Xiaomei Zhou, Chengfei Liu, and Xiaofang Zhou. Efficiently computing weighted proximity relationships in spatial databases. Lecture Notes in Computer Science, 2118:279– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180279.
pdf.

Wang:2001:CIS

[150] Shuguang Wang, Cui Yu, and Beng Chin Ooi. Compressing the index — A simple and yet efficient approximation approach to highdimensional indexing. Lecture Notes in Computer Science, 2118:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180291.
pdf.

Barbara:2001:FDC

[151] Daniel Barbará and Xintao Wu. Finding dense clusters in hyperspace: An approach based on row shuffling. Lecture Notes in Computer Science, 2118: 305–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180305.
pdf.

Gu:2001:MCP

[152] Baohua Gu, Feifang Hu, and Huan Liu. Modelling classification performance for large data sets. Lecture Notes in Computer Science, 2118:317– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180317.
pdf.

Yang:2001:MRI

[153] Xiaochun Yang and Guoren Wang. Mapping referential integrity constraints from relational databases to XML. Lecture Notes in Computer Science, 2118:329–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2118/21180329.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2118/21180329.pdf.

Lu:2001:AVE

[154] Hongjun Lu and Zengping Tian. Attribute value extraction and standardization in data integration. Lecture Notes in Computer Science, 2118:343– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180343.
pdf.

Lian:2001:IAR

[155] Wang Lian, David W. Cheung, and S. M. Yiu. Indexing attributes and reordering profiles for XML document filtering and information delivery. Lecture Notes in Computer Science, 2118: 355–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180355.
pdf.

Yin:2001:ADF

[156] Xuri Yin, Zhihua Zhou, Ning Li, and Shifu Chen. An approach for data filtering based on rough set theory. Lecture Notes in Computer Science, 2118: 367–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180367.
                                        pdf.

Chiu:2001:WBW

[157] Dickson K. W. Chiu, Qing Li, and Kamalakar Karlapalem. Web-based workflow evolution in ADOME–WFMS. Lecture Notes in Computer Science, 2118:377–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2118/21180377.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180377.
                                        pdf.

Estrella:2001:MDO

[158] Florida Estrella, Zsolt Kovacs, Jean-Marie Le Goff, Richard McClatchey, and Norbert Toth. Meta-data objects as the basis for system evolution. Lecture Notes in Computer Science, 2118: 390–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2118/21180390.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2118/21180390.
                                        pdf.

Huang:2001:AWM

[159] Lieming Huang, Ulrich Thiel, Matthias Hemmje, and Erich J. Neuhold. Adaptive Web meta-search enhanced by constraint-based query constructing and mapping. Lecture Notes in Computer Science, 2118:400–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180400.
pdf.

Cao:2001:MDD

[160] Guohong Cao, Yiqiong Wu, and Bo Li. A mixed data dissemination strategy for mobile computing systems. Lecture Notes in Computer Science, 2118:408– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2118/21180408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2118/21180408.
pdf.

Anonymous:2001:AIfh

[161] Anonymous. Author index. Lecture Notes in Computer Science, 2118:417– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2118/2118auth.pdf.

Yacobi:2001:FTC

[162] Yacov Yacobi. A few thoughts on E-commerce. Lecture Notes in Computer Science, 2119:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190001.
pdf.

Brincat:2001:NCM

[163] Karl Brincat and Chris J. Mitchell. New CBC–MAC forgery attacks. Lecture Notes in Computer Science, 2119: 3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190003.
pdf.

Youssef:2001:CPK

[164] Amr Youssef and Guang Gong. Cryptanalysis of a public key cryptosystem proposed at ACISP 2000. Lecture Notes in Computer Science, 2119:15– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190015.
pdf.

Zenner:2001:ICS

[165] Erik Zenner, Matthias Krause, and Stefan Lucks. Improved cryptanalysis of the self-shrinking generator. Lecture Notes in Computer Science, 2119:21– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190021.
pdf.

Pavlovski:2001:ABS

[166] Chris Pavlovski and Colin Boyd. Attacks based on small factors in various group structures. Lecture Notes in Computer Science, 2119:36–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190036.
pdf.

Saeednia:2001:CCK

[167] Shahrokh Saeednia, Rei Safavi-Naini, and Willy Susilo. On classifying conference key distribution protocols. Lecture Notes in Computer Science, 2119: 51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190051.
pdf.

Kang:2001:PMT

[168] Ju-Sung Kang, Okyeon Yi, Dowon Hong, and Hyunsook Cho. Pseudorandomness of MISTY-type transformations and the block cipher KASUMI. Lecture Notes in Computer Science, 2119:60–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2119/21190060.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190060.
                                        pdf.

Kim:2001:NPK

[169] Hwankoo Kim and SangJae Moon. New public-key cryptosystem using divisor class groups. Lecture Notes in Computer Science, 2119:74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190074.
                                        pdf.

Meyer:2001:FIC

[170] Andreas Meyer, Stefan Neis, and Thomas Pfahler. First implementation of cryptographic protocols based on algebraic number fields. Lecture Notes in Computer Science, 2119:84– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190084.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190084.
                                        pdf.

Yen:2001:PKR

[171] Sung-Ming Yen. Practical key recovery schemes. Lecture Notes in Computer Science, 2119:104–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190104.
pdf.

May:2001:NDP

[172] David May, Henk L. Muller, and Nigel P. Smart. Non-deterministic processors. Lecture Notes in Computer Science, 2119:115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2119/21190115.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2119/21190115.pdf.

Itoi:2001:PSB

[173] Naomaru Itoi, William A. Arbaugh, Samuela J. Pollack, and Daniel M. Reeves. Personal secure booting. Lecture Notes in Computer Science, 2119: 130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190130.
pdf.

Goto:2001:ETR

[174] Hideaki Goto, Masahiro Mambo, Hiroki Shizuya, and Yasuyoshi Watanabe. Evaluation of tamper-resistant software deviating from structured programming rules. Lecture Notes in Computer Science, 2119:145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190145.
                                        pdf.

Wietrzyk:2001:SMW

[175] Vlad Ingar Wietrzyk, Makoto Takizawa, and Vijay Varadharajan. A strategy for MLS workflow. Lecture Notes in Computer Science, 2119:159– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190159.
                                        pdf.

Neuman:2001:CDI

[176] Clifford Neuman. Condition-driven integration of security services. Lecture Notes in Computer Science, 2119:176– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190176.
                                        pdf.

Cho:2001:SSK

[177] Eun-Sun Cho, Sunho Hong, Sechang Oh, Hong-Jin Yeh, Manpyo Hong, Cheol-Won Lee, Hyundong Park, and Chun-Sik Park. SKETHIC: Secure Kernel Extension against Trojan Horses with Information-Carrying Codes. Lecture Notes in Computer Science, 2119:177–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190177.
pdf.

Bao:2001:SPD

[178] Feng Bao, Robert Deng, Peirong Feng, Yan Guo, and Hongjun Wu. Secure and private distribution of online video and some related cryptographic issues. Lecture Notes in Computer Science, 2119: 190–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190190.
pdf.

Yamamura:2001:PIR

[179] Akihiro Yamamura and Taiichi Saito. Private information retrieval based on the subgroup membership problem. Lecture Notes in Computer Science, 2119:206–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2119/21190206.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2119/21190206.
pdf.

Omote:2001:PEA

[180] Kazumasa Omote and Atsuko Miyaji. A practical English auction with onetime registration. Lecture Notes in Computer Science, 2119:221–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190221.
                                        pdf.

Hirose:2001:UAS

[181] Shouichi Hirose and Susumu Yoshida. A user authentication scheme with identity and location privacy. Lecture Notes in Computer Science, 2119:235– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190235.
                                        pdf.

Yang:2001:EEA

[182] Jong-Phil Yang, Weon Shin, and Kyung-Hyune Rhee. An end-to-end authentication protocol in wireless application protocol. Lecture Notes in Computer Science, 2119:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190247.
                                        pdf.

Yamamura:2001:EDA

[183] Akihiro Yamamura and Hirokazu Ishizuka. Error detection and authentication in quantum key distribution. Lecture Notes in Computer Science, 2119:260–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2119/21190260.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2119/21190260.
pdf.

Liu:2001:ABR

[184] Chuchang Liu, Maris Ozols, and Tony Cant. An axiomatic basis for reasoning about trust in PKIs. Lecture Notes in Computer Science, 2119:274– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190274.
pdf.

Lin:2001:KBA

[185] Along Lin. A knowledge-based approach to Internet authorizations. Lecture Notes in Computer Science, 2119: 292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190292.
pdf.

Yesberg:2001:ATR

[186] John Yesberg and Marie Henderson. Applications of trusted review to information security. Lecture Notes in Computer Science, 2119:305–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190305.
                                        pdf.

Chi:2001:NSM

[187] Sung-Do Chi, Jong Sou Park, Ki-Chan Jung, and Jang-Se Lee. Network security modeling and cyber attack simulation methodology. Lecture Notes in Computer Science, 2119:320– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190320.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190320.
                                        pdf.

Park:2001:CSC

[188] DongGook Park, JungJoon Kim, Colin Boyd, and Ed Dawson. Cryptographic salt: A countermeasure against denial-of-service attacks. Lecture Notes in Computer Science, 2119:334–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190334.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190334.
                                        pdf.

Jung:2001:EMO

[189] Oliver Jung, Sven Kuhn, Christoph Ruland, and Kai Wollenweber. Enhanced modes of operation for the encryption in high-speed networks and their impact on QoS. Lecture Notes in Computer Science, 2119:344–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190344.
pdf.

Ansper:2001:IAT

[190] Arne Ansper, Ahto Buldas, Märt Saarepera, and Jan Willemson. Improving the availability of timestamping services. Lecture Notes in Computer Science, 2119:360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190360.
pdf.

King:2001:RRL

[191] Brian King. Randomness required for linear threshold sharing schemes defined over any finite Abelian group. Lecture Notes in Computer Science, 2119:376–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190376.
pdf.

Ghodosi:2001:DS

[192] Hossein Ghodosi and Josef Pieprzyk. Democratic systems. Lecture Notes in Computer Science, 2119:392–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190392.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190392.
                                        pdf.

Maeda:2001:EUS

[193] Ayako Maeda, Atsuko Miyaji, and Mitsuru Tada. Efficient and unconditionally secure verifiable threshold changeable scheme. Lecture Notes in Computer Science, 2119:403–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190403.
                                        pdf.

Stinson:2001:PSD

[194] Douglas R. Stinson and Reto Strobl. Provably secure distributed Schnorr signatures and a (t, n) threshold scheme for implicit certificates. Lecture Notes in Computer Science, 2119:417– 434, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190417.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190417.
                                        pdf.

Safavi-Naini:2001:HCF

[195] Rei Safavi-Naini, Willy Susilo, and Huaxiong Wang. How to construct failstop confirmer signature schemes. Lecture Notes in Computer Science, 2119: 435–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190435.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190435.
pdf.

Tan:2001:SSB

[196] Chik How Tan, Xun Yi, and Chee Kheong Siew. Signature schemes based on 3^{rd} order shift registers. Lecture Notes in Computer Science, 2119: 445–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190445.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190445.
pdf.

Nakanishi:2001:ASS

[197] Toru Nakanishi and Yuji Sugiyama. Anonymous statistical survey of attributes. Lecture Notes in Computer Science, 2119:460–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2119/21190460.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2119/21190460.pdf.

Lee:2001:SMA

[198] Byoungcheon Lee, Heesun Kim, and Kwangjo Kim. Secure mobile agent using strong non-designated proxy signature. Lecture Notes in Computer Science, 2119:474–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190474.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190474.
                                        pdf.

Boyd:2001:ECB

[199] Colin Boyd, Paul Montague, and Khanh Nguyen. Elliptic curve based password authenticated key exchange protocols. Lecture Notes in Computer Science, 2119:487–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190487.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190487.
                                        pdf.

Weimerskirch:2001:ECC

[200] André Weimerskirch, Christof Paar, and Sheueling Chang Shantz. Elliptic curve cryptography on a palm OS device. Lecture Notes in Computer Science, 2119:502–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2119/21190502.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2119/21190502.
                                        pdf.

Shim:2001:RCE

[201] Kyungah Shim. Reducing certain elliptic curve discrete logarithms to logarithms in a finite field. Lecture Notes in Computer Science, 2119:514– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2119/21190514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2119/21190514.
pdf.

Anonymous:2001:AIfi

[202] Anonymous. Author index. Lecture Notes in Computer Science, 2119:521– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2119/2119indx.pdf.

Sarbo:2001:POL

[203] Janos Sarbo and József Farkas. A peircean ontology of language. Lecture Notes in Computer Science, 2120: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200001.
pdf.

Hoede:2001:WGT

[204] C. Hoede and L. Zhang. Word graphs: The third set. Lecture Notes in Computer Science, 2120:15– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200015.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2120/21200015.
                                        pdf.

Amghar:2001:ATD

[205] Tassadit Amghar, Thierry Charnois, and Delphine Battistelli. Aspectotemporal data and lexical representations in French within simple conceptual graphs on the basis of semanticocognitive schemes. Lecture Notes in Computer Science, 2120:29–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2120/21200029.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2120/21200029.
                                        pdf.

Zhang:2001:LGC

[206] Lei Zhang and Yong Yu. Learning to generate CGs from domain specific sentences. Lecture Notes in Computer Science, 2120:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2120/21200044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2120/21200044.
                                        pdf.

Masmoudi:2001:SOD

[207] Slim Masmoudi. Solving-oriented and domain-oriented knowledge structures: Their application to debugging problem solving activity. Lecture Notes in Computer Science, 2120:58– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2120/21200058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200058.
pdf.

Dau:2001:CGP

[208] Frithjof Dau. Concept graphs and predicate logic. Lecture Notes in Computer Science, 2120:72–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200072.
pdf.

Cao:2001:GQC

[209] Tru H. Cao. Generalized quantifiers and conceptual graphs. Lecture Notes in Computer Science, 2120:87– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200087.
pdf.

Klinger:2001:SSG

[210] Julia Klinger. Simple semiconcept graphs: A Boolean logic approach. Lecture Notes in Computer Science, 2120: 101–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200101.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2120/21200101.
pdf.

Wille:2001:BJL

[211] Rudolf Wille. Boolean judgment logic. Lecture Notes in Computer Science, 2120:115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2120/21200115.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2120/21200115.
pdf.

Ganter:2001:PST

[212] Bernhard Ganter and Sergei O. Kuznetsov. Pattern structures and their projections. Lecture Notes in Computer Science, 2120:129–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200129.
pdf.

Ganter:2001:FCA

[213] Bernhard Ganter and Sebastian Rudolph. Formal concept analysis methods for dynamic conceptual graphs. Lecture Notes in Computer Science, 2120:143–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200143.
pdf.

Gugisch:2001:MVC

[214] Ralf Gugisch. Many-valued context analysis using descriptions. Lecture Notes in Computer Science, 2120:157– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200157.
pdf.

Strahringer:2001:MSE

[215] Selma Strahringer, Rudolf Wille, and Uta Wille. Mathematical support for empirical theory building. Lecture Notes in Computer Science, 2120:169– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200169.
pdf.

Ferre:2001:SOP

[216] Sébastien Ferré and Olivier Ridoux. Searching for objects and properties with logical concept analysis. Lecture Notes in Computer Science, 2120:187– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200187.
pdf.

Hereth:2001:RPC

[217] Jo Hereth and Gerd Stumme. Reverse pivoting in conceptual information systems. Lecture Notes in Computer Science, 2120:202–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2120/21200202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2120/21200202.
                                        pdf.

Dibie-Barthelemy:2001:RCG

[218] Juliette Dibie-Barthélemy, Ollivier Haemmerlé, and Stéphane Loiseau. Refinement of conceptual graphs. Lecture Notes in Computer Science, 2120:216– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2120/21200216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2120/21200216.
                                        pdf.

Martin:2001:LSC

[219] Philippe Martin and Peter Eklund. Large-scale cooperatively-built KBs. Lecture Notes in Computer Science, 2120:231–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2120/21200231.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2120/21200231.
                                        pdf.

Gerbe:2001:CGM

[220] Olivier Gerbé, Guy W. Mineau, and Rudolf K. Keller. Conceptual graphs and metamodeling. Lecture Notes in Computer Science, 2120:245–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200245.
pdf.

deMoor:2001:MVC

[221] Aldo de Moor and Willem-Jan van den Heuvel. Making virtual communities work: Matching their functionalities. Lecture Notes in Computer Science, 2120:260–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200260.
pdf.

Delteil:2001:ERB

[222] Alexandre Delteil, Catherine Faron-Zucker, and Rose Dieng. Extension of RDFS based on the CGs formalisms. Lecture Notes in Computer Science, 2120:275–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200275.
pdf.

Valtchev:2001:BCG

[223] Petko Valtchev and Rokia Missaoui. Building concept (Galois) lattices from parts: Generalizing the incremental methods. Lecture Notes in Computer Science, 2120:290–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2120/21200290.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2120/21200290.
                                        pdf.

Kim:2001:TBA

[224] Pok-Son Kim and Manfred Schmidt-Schauß. A term-based approach to project scheduling. Lecture Notes in Computer Science, 2120:304–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2120/21200304.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2120/21200304.
                                        pdf.

Cole:2001:BSS

[225] Richard Cole and Peter Eklund. Browsing semi-structured Web texts using formal concept analysis. Lecture Notes in Computer Science, 2120:319– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2120/21200319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2120/21200319.
                                        pdf.

Southey:2001:OCM

[226] Finnegan Southey and James G. Linders. Ossa — A conceptual modelling system for virtual realities. Lecture Notes in Computer Science, 2120:333– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200333.
pdf.

Kabbaj:2001:UIE

[227] Adil Kabbaj, Bernard Moulin, Jeremi Gancet, David Nadeau, and Olivier Rouleau. Uses, improvements, and extensions of Prolog+CG: Case studies. Lecture Notes in Computer Science, 2120:346–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2120/21200346.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2120/21200346.
pdf.

Benn:2001:APM

[228] David Benn and Dan Corbett. An application of the process mechanism to a room allocation problem using the pCG language. Lecture Notes in Computer Science, 2120:360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2120/21200360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2120/21200360.
pdf.

Anonymous:2001:AIfj

[229] Anonymous. Author index. Lecture Notes in Computer Science, 2120:377– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2120/2120auth.pdf.

Su:2001:MOL

[230] Jianwen Su, Haiyan Xu, and Oscar H. Ibarra. Moving objects: Logical relationships and queries. Lecture Notes in Computer Science, 2121:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210003.
                                        pdf.

Vazirgiannis:2001:SML

[231] Michalis Vazirgiannis and Ouri Wolfson. A spatiotemporal model and language for moving objects on road networks. Lecture Notes in Computer Science, 2121:20–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210020.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210020.
                                        pdf.

Goyal:2001:SCD

[232] Roop K. Goyal and Max J. Egenhofer. Similarity of cardinal directions. Lecture Notes in Computer Science, 2121:

36–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210036.
pdf.

Porkaew:2001:QMO

[233] Kriengkrai Porkaew, Iosif Lazaridis, and Sharad Mehrotra. Querying mobile objects in spatio-temporal databases. Lecture Notes in Computer Science, 2121:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2121/21210059.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2121/21210059.
pdf.

Song:2001:NNS

[234] Zhexuan Song and Nick Roussopoulos. K-nearest neighbor search for moving query point. Lecture Notes in Computer Science, 2121:79–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210079.
pdf.

Zheng:2001:SCL

[235] Baihua Zheng and Dik Lun Lee. Semantic caching in location-dependent query processing. Lecture Notes in Computer Science, 2121:97–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210097.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210097.
                                        pdf.

Nicklas:2001:MBO

[236] Daniela Nicklas, Matthias Großmann, Thomas Schwarz, Steffen Volz, and Bernhard Mitschang. A model-based, open architecture for mobile, spatially aware applications. Lecture Notes in Computer Science, 2121:117– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210117.
                                        pdf.

Brinkhoff:2001:CQW

[237] Thomas Brinkhoff and Jürgen Weitkämper. Continuous queries within an architecture for querying XML-represented moving objects. Lecture Notes in Computer Science, 2121:136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210136.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210136.
                                        pdf.

Mamoulis:2001:SEC

[238] Nikos Mamoulis and Dimitris Papadias. Selectivity estimation of complex spatial queries. Lecture Notes in Computer Science, 2121:155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210155.
pdf.

Wang:2001:WBC

[239] Min Wang, Jeffrey Scott Vitter, Lipyeow Lim, and Sriram Padmanabhan. Wavelet-based cost estimation for spatial queries. Lecture Notes in Computer Science, 2121:175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210175.
pdf.

Chan:2001:EBQ

[240] Edward P. F. Chan. Evaluation of buffer queries in spatial databases. Lecture Notes in Computer Science, 2121: 197–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210197.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210197.
pdf.

Zhu:2001:MWS

[241] Hongjun Zhu, Jianwen Su, and Oscar H. Ibarra. On multi-way spatial joins with direction predicates. Lecture Notes in Computer Science, 2121:

217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210217.
                                        pdf.

Shekhar:2001:DSC

[242] Shashi Shekhar and Yan Huang. Discovering spatial co-location patterns: A summary of results. Lecture Notes in Computer Science, 2121:236–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210236.
                                        pdf.

Ferhatosmanoglu:2001:CNN

[243] Hakan Ferhatosmanoglu, Ioanna Stanoi, Divyakant Agrawal, and Amr El Abbadi. Constrained nearest neighbor queries. Lecture Notes in Computer Science, 2121:257–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2121/21210257.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2121/21210257.pdf.

Lago:2001:CTG

[244] Ugo Dal Lago and Angelo Montanari. Calendars, time granularities, and automata. Lecture Notes in Computer Science, 2121:279–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210279.
pdf.

Skiadopoulos:2001:CCD

[245] Spiros Skiadopoulos and Manolis Koubarakis. Composing cardinal direction relations. Lecture Notes in Computer Science, 2121:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210299.
pdf.

Tossebro:2001:CRC

[246] Erlend Tøssebro and Ralf Hartmut Güting. Creating representations for continuously moving regions from observations. Lecture Notes in Computer Science, 2121:321–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210321.
pdf.

Danovaro:2001:CMT

[247] Emanuele Danovaro, Leila De Floriani, Paola Magillo, and Enrico Puppo. Compressing multiresolution triangle meshes. Lecture Notes in Computer Science, 2121:345–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210345.
                                        pdf.

Zhou:2001:DIM

[248] Sheng Zhou and Christopher B. Jones. Design and implementation of multi-scale databases. Lecture Notes in Computer Science, 2121:365– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210365.
                                        pdf.

East:2001:AAD

[249] Russell East, Roop Goyal, Art Haddad, Alexander Konovalov, Andrea Rosso, Mike Tait, and Jay Theodore. The architecture of ArcIMS, a distributed Internet map server. Lecture Notes in Computer Science, 2121:387– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210387.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210387.
                                        pdf.

Kothuri:2001:EPL

[250] Ravi Kanth Kothuri and Siva Ravada. Efficient processing of large spatial queries using interior approximations.

Lecture Notes in Computer Science, 2121:404–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2121/21210404.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2121/21210404.
pdf.

Tsoukatos:2001:EMS

[251] Ilias Tsoukatos and Dimitrios Gunopulos. Efficient mining of spatiotemporal patterns. Lecture Notes in Computer Science, 2121:425–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210425.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210425.
pdf.

Papadias:2001:EOO

[252] Dimitris Papadias, Panos Kalnis, Jun Zhang, and Yufei Tao. Efficient OLAP operations in spatial data warehouses. Lecture Notes in Computer Science, 2121:443–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2121/21210443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2121/21210443.
pdf.

Pedersen:2001:PAS

[253] Torben Bach Pedersen and Nektaria Tryfona. Pre-aggregation in spatial data warehouses. Lecture Notes in Computer Science, 2121:460–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210460.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210460.
                                        pdf.

Kriegel:2001:ISO

[254] Hans-Peter Kriegel, Marco Pötke, and Thomas Seidl. Interval sequences: An object-relational approach to manage spatial data. Lecture Notes in Computer Science, 2121:481–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210481.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210481.
                                        pdf.

Hambrusch:2001:QPB

[255] Susanne Hambrusch, Chuan-Ming Liu, Walid G. Aref, and Sunil Prabhakar. Query processing in broadcasted spatial index trees. Lecture Notes in Computer Science, 2121:502– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2121/21210502.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2121/21210502.
                                        pdf.

Kriegel:2001:ORI

[256] Hans-Peter Kriegel, Marco Pötke, and Thomas Seidl. Object-relational indexing for general interval relationships. Lecture Notes in Computer Science, 2121:522–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2121/21210522.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2121/21210522.
pdf.

Anonymous:2001:AIfk

[257] Anonymous. Author index. Lecture Notes in Computer Science, 2121:543– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2121/2121auth.pdf.

Aigner:2001:LPD

[258] Martin Aigner. Lattice paths and determinants. Lecture Notes in Computer Science, 2122:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2122/21220001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2122/21220001.
pdf.

Alt:2001:NN

[259] Helmut Alt. The nearest neighbor. Lecture Notes in Computer Science, 2122: 13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2122/21220013.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2122/21220013.
                                        pdf.

Gartner:2001:EIE

[260] Bernd Gärtner and Emo Welzl. Explicit and implicit enforcing — randomized optimization. Lecture Notes in Computer Science, 2122:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2122/21220025.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2122/21220025.
                                        pdf.

Helleseth:2001:CZ

[261] Tor Helleseth. Codes over Z_{4}. Lecture Notes in Computer Science, 2122: 47–55, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2122/21220047.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2122/21220047.
                                        pdf.

Jung:2001:DBL

[262] Heinz Adolf Jung. Degree bounds for long paths and cycles in kconnected graphs. Lecture Notes in Computer Science, 2122:56–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2122/21220056.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2122/21220056.
                                        pdf.

Meinel:2001:DSB

[263] Christoph Meinel and Christian Stangier. Data structures for Boolean functions. Lecture Notes in Computer Science, 2122:61–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2122/21220061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2122/21220061.
pdf.

Mohring:2001:SUU

[264] Rolf H. Möhring. Scheduling under uncertainty: Bounding the makespan distribution. Lecture Notes in Computer Science, 2122:79–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2122/21220079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2122/21220079.
pdf.

Promel:2001:RGR

[265] Hans Jürgen Prömel and Anusch Taraz. Random graphs, random triangle-free graphs, and random partial orders. Lecture Notes in Computer Science, 2122:98–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2122/21220098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2122/21220098.
pdf.

Rote:2001:DFA

[266] Günter Rote. Division-free algorithms for the determinant and the Pfaffian: Algebraic and combinatorial approaches. Lecture Notes in Computer Science, 2122:119–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2122/21220119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2122/21220119.
                                        pdf.

Schulz:2001:CCS

[267] Ralph-Hardo Schulz. Check character systems and anti-symmetric mappings. Lecture Notes in Computer Science, 2122:136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2122/21220136.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2122/21220136.
                                        pdf.

Stroth:2001:APM

[268] Gernot Stroth. Algorithms in pure mathematics. Lecture Notes in Computer Science, 2122:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2122/21220148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2122/21220148.
                                        pdf.

Ziegler:2001:CHG

[269] Günter M. Ziegler. Coloring Hamming graphs, optimal binary codes, and the 0/1-Borsuk problem in low dimensions. Lecture Notes in Computer Science, 2122:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2122/21220159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2122/21220159.
pdf.

Anonymous:2001:AIfl

[270] Anonymous. Author index. Lecture Notes in Computer Science, 2122:173– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2122/2122auth.pdf.

Visa:2001:TTM

[271] A. Visa. Technology of text mining. Lecture Notes in Computer Science, 2123:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2123/21230001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2123/21230001.
pdf.

Sack:2001:ECR

[272] U. Sack and M. Kamprad. Evaluation of clinical relevance of clinical laboratory investigations by data mining. Lecture Notes in Computer Science, 2123:12–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2123/21230012.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2123/21230012.
pdf.

Schmidt:2001:TAC

[273] R. Schmidt and L. Gierl. Temporal abstractions and case-based reasoning for medical course data: Two prognostic applications. Lecture Notes in Computer Science, 2123:23–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2123/21230023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2123/21230023.
pdf.

Perner:2001:CBR

[274] P. Perner. Are case-based reasoning and dissimilarity-based classification two sides of the same coin? Lecture Notes in Computer Science, 2123:35– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2123/21230035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2123/21230035.
pdf.

Jang:2001:FBF

[275] S.-W. Jang, G.-Y. Kim, and H.-I. Choi. FAM-based fuzzy inference for detecting shot transitions. Lecture Notes in Computer Science, 2123:52– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2123/21230052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2123/21230052.
pdf.

Indurkhya:2001:RBE

[276] N. Indurkhya and S. M. Weiss. Rulebased ensemble solutions for regression. Lecture Notes in Computer Science, 2123:62–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2123/21230062.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2123/21230062.
pdf.

Fernau:2001:LXG

[277] H. Fernau. Learning XML grammars. Lecture Notes in Computer Science, 2123:73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2123/21230073.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2123/21230073.
pdf.

Malerba:2001:FOR

[278] D. Malerba, F. Esposito, A. Lanza, and F. A. Lisi. First-order rule induction for the recognition of morphological patterns in topographic maps. Lecture Notes in Computer Science, 2123:88– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2123/21230088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230088.
                                        pdf.

Bhanu:2001:CLF

[279] B. Bhanu and A. Dong. Concepts learning with fuzzy clustering and relevance feedback. Lecture Notes in Computer Science, 2123:102–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230102.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230102.
                                        pdf.

Martinez-Trinidad:2001:LCC

[280] J. F. Martínez-Trinidad and G. Sánchez-Díaz. LC: A conceptual clustering algorithm. Lecture Notes in Computer Science, 2123:117–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230117.
                                        pdf.

Sy:2001:DMA

[281] B. K. Sy and A. K. Gupta. Data mining approach based on informationstatistical analysis: Application to temporal-spatial data. Lecture Notes in Computer Science, 2123:128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2123/21230128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2123/21230128.
pdf.

Perner:2001:HTD

[282] P. Perner and T. Belikova. A hybrid tool for data mining in picture archiving system. Lecture Notes in Computer Science, 2123:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2123/21230141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2123/21230141.
pdf.

Kollmar:2001:FSR

[283] D. Kollmar and D. H. Hellmann. Feature selection for a real-world learning task. Lecture Notes in Computer Science, 2123:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2123/21230157.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2123/21230157.pdf.

Fischer:2001:AID

[284] S. Fischer and H. Bunke. Automatic identification of diatoms using decision forests. Lecture Notes in Computer Science, 2123:173–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2123/21230173.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2123/21230173.
                                        pdf.

Toivonen:2001:VTC

[285] J. Toivonen, A. Visa, T. Vesanen, B. Back, and H. Vanharanta. Validation of text clustering based on document contents. Lecture Notes in Computer Science, 2123:184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230184.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230184.
                                        pdf.

Wachowiak:2001:SNA

[286] M. P. Wachowiak and R. Smolíková. Statistical and neural approaches for estimating parameters of a speckle model based on the nakagami distribution. Lecture Notes in Computer Science, 2123:196–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230196.
                                        pdf.

Linder:2001:HAN

[287] R. Linder and S. J. Pöppl. How to automate neural net based learning. Lecture Notes in Computer Science, 2123: 206–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230206.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2123/21230206.
pdf.

Krzyzak:2001:NFL

[288] A. Krzyzak. Nonlinear function learning and classification using optimal radial basis function networks. Lecture Notes in Computer Science, 2123:217– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2123/21230217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2123/21230217.
pdf.

Dong:2001:LLF

[289] J. x. Dong, A. Krzyzak, and C. Y. Suen. Local learning framework for recognition of lowercase handwritten characters. Lecture Notes in Computer Science, 2123:226–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2123/21230226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2123/21230226.
pdf.

Shi:2001:MIL

[290] M. Shi, T. Wakabayashi, W. Ohyama, and F. Kimura. Mirror image learning for handwritten numeral recognition. Lecture Notes in Computer Science, 2123:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2123/21230239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230239.
                                        pdf.

Pham:2001:FDA

[291] T. V. Pham, M. Worring, and A. W. M. Smeulders. Face detection by aggregated Bayesian network classifiers. Lecture Notes in Computer Science, 2123: 249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230249.
                                        pdf.

Wu:2001:TSE

[292] Y. Wu and T. S. Huang. Towards selfexploring discriminating features. Lecture Notes in Computer Science, 2123: 263–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230263.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230263.
                                        pdf.

Imiya:2001:PBM

[293] A. Imiya and H. Ootani. PCAbased model selection and fitting for linear manifolds. Lecture Notes in Computer Science, 2123:278–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230278.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2123/21230278.
pdf.

Imiya:2001:SFV

[294] A. Imiya and K. Iwawaki. Statistics of flow vectors and its application to the voting method for the detection of flow fields. Lecture Notes in Computer Science, 2123:293–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2123/21230293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2123/21230293.
pdf.

Krawiec:2001:UPC

[295] K. Krawiec. On the use of pairwise comparison of hypotheses in evolutionary learning applied to learning from visual examples. Lecture Notes in Computer Science, 2123:307– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2123/21230307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2123/21230307.
pdf.

Mottl:2001:FPR

[296] V. Mottl, S. Dvoenko, O. Seredin, C. Kulikowski, and I. Muchnik. Featureless pattern recognition in an imaginary Hilbert space and its application to protein fold classification. Lecture Notes in Computer Science, 2123:322– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230322.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230322.
                                        pdf.

Giacinto:2001:AQS

[297] G. Giacinto, F. Roli, and G. Fumera. Adaptive query shifting for contentbased image retrieval. Lecture Notes in Computer Science, 2123:337–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230337.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230337.
                                        pdf.

Potamias:2001:CBS

[298] G. Potamias. Content-based similarity assessment in multi-segmented medical image data bases. Lecture Notes in Computer Science, 2123:347– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2123/21230347.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2123/21230347.
                                        pdf.

Anonymous:2001:AIfm

[299] Anonymous. Author index. Lecture Notes in Computer Science, 2123:363– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2123/2123auth.pdf.

Bober:2001:MER

[300] Mirosław Bober. MPEG-7: Evolution or revolution? Lecture Notes in Computer Science, 2124:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240001.
pdf.

Ohm:2001:MVD

[301] Jens-Rainer Ohm. The MPEG-7 visual description framework — concepts, accuracy, and applications. Lecture Notes in Computer Science, 2124: 2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240002.
pdf.

Cieplinski:2001:MCD

[302] Leszek Cieplinski. MPEG-7 color descriptors and their applications. Lecture Notes in Computer Science, 2124: 11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240011.
pdf.

Wu:2001:TDM

[303] Peng Wu, Yong Man Ro, Chee Sun Won, and Yanglim Choi. Texture descriptors in MPEG-7. Lecture Notes in Computer Science, 2124:21– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240021.
pdf.

Divakaran:2001:OMM

[304] Ajay Divakaran. An overview of MPEG-7 motion descriptors and their applications. Lecture Notes in Computer Science, 2124:29–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240029.
pdf.

Benitez:2001:MMC

[305] Ana B. Benitez, Di Zhong, Shih-Fu Chang, and John R. Smith. MPEG-7 MDS content description tools and applications. Lecture Notes in Computer Science, 2124:41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240041.
pdf.

Walczak:2001:IRU

[306] Krzysztof Walczak. Image retrieval using spatial color information. Lecture Notes in Computer Science, 2124:53– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240053.
pdf.

Przelaskowski:2001:LBR

[307] Artur Przelaskowski. Lifting-based reversible transforms for lossy-to-lossless wavelet codecs. Lecture Notes in Computer Science, 2124:61–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240061.
pdf.

Stasinski:2001:CII

[308] Ryszard Stasinski. Coding of irregular image regions by SA DFT. Lecture Notes in Computer Science, 2124:71– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240071.
pdf.

Virmajoki:2001:FPU

[309] Olli Virmajoki, Pasi Fränti, and Timo Kaukoranta. Fast PNN using partial distortion search. Lecture Notes in Computer Science, 2124:77– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240077.
                                        pdf.

Domanski:2001:NLC

[310] Marek Domanski and Krzysztof Rakowski. Near-lossless color image compression with no error accumulation in multiple coding cycles. Lecture Notes in Computer Science, 2124:85– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240085.
                                        pdf.

Przelaskowski:2001:HLC

[311] Artur Przelaskowski. Hybrid lossless coder of medical images with statistical data modelling. Lecture Notes in Computer Science, 2124:92– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240092.
                                        pdf.

Bartkowiak:2001:SAO

[312] Maciej Bartkowiak and Adam Łuczak. A simple algorithm for ordering and compression of vector codebooks. Lecture Notes in Computer Science, 2124: 102–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240102.
pdf.

Domanski:2001:MBV

[313] Marek Domański and Sławomir Maćkowiak. MPEG 2-based video coding with three-layer mixed scalability. Lecture Notes in Computer Science, 2124:110– 117, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240110.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240110.
pdf.

Siemek:2001:CBR

[314] Grzegorz Siemek. The coefficient based rate distortion model for the low bit rate video coding. Lecture Notes in Computer Science, 2124:118– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240118.
pdf.

Mroczek:2001:SAD

[315] Krzysztof Mroczek. Shape-adaptive DCT algorithm — hardware optimized redesign. Lecture Notes in Computer Science, 2124:125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240125.
                                        pdf.

Krivic:2001:SBO

[316] Jaka Krivic and Franc Solina. Superquadricbased object recognition. Lecture Notes in Computer Science, 2124:134– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240134.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240134.
                                        pdf.

Carcassoni:2001:WGM

[317] Marco Carcassoni and Edwin R. Hancock. Weighted graph-matching using modal clusters. Lecture Notes in Computer Science, 2124:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240142.
                                        pdf.

Luo:2001:DSC

[318] B. Luo, A. Robles-Kelly, A. Torsello, R. C. Wilson, and E. R. Hancock. Discovering shape categories by clustering shock trees. Lecture Notes in Computer Science, 2124:152–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2124/21240152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240152.
pdf.

Pernkopf:2001:FSC

[319] Franz Pernkopf and Paul O’Leary. Feature selection for classification using genetic algorithms with a novel encoding. Lecture Notes in Computer Science, 2124:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240161.
pdf.

Franc:2001:CSA

[320] Vojtěch Franc and Václav Hlaváč. A contribution to the Schlesinger’s algorithm separating mixtures of Gaussians. Lecture Notes in Computer Science, 2124:169–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2124/21240169.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2124/21240169.pdf.

Chernov:2001:DAA

[321] Vladimir M. Chernov. Diophantine approximations of algebraic irrationalities and stability theorems for polynomial decision rules. Lecture Notes in Computer Science, 2124:177–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240177.
                                        pdf.

Suk:2001:FIS

[322] Tomáš Suk and Jan Flusser. Features invariant simultaneously to convolution and affine transformation. Lecture Notes in Computer Science, 2124: 183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240183.
                                        pdf.

Lehal:2001:TSG

[323] G. S. Lehal and Chandan Singh. A technique for segmentation of gurmukhi text. Lecture Notes in Computer Science, 2124:191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240191.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240191.
                                        pdf.

Tuzikov:2001:ECB

[324] Alexander V. Tuzikov, Stanislav A. Sheynin, and Pavel V. Vasiliev. Efficient computation of body moments. Lecture Notes in Computer Science, 2124:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2124/21240201.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2124/21240201.
pdf.

Krawiec:2001:GPL

[325] Krzysztof Krawiec. Genetic programming with local improvement for visual learning from examples. Lecture Notes in Computer Science, 2124:209– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240209.
pdf.

Chan:2001:IRS

[326] Jonathan C.-W. Chan, Ruth S. De-Fries, and John R. G. Townshend. Improved recognition of spectrally mixed land cover classes using spatial textures and voting classifications. Lecture Notes in Computer Science, 2124: 217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240217.
pdf.

Verma:2001:TFE

[327] B. Verma and S. Kulkarni. Texture feature extraction and classification. Lecture Notes in Computer Science, 2124: 228–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240228.
                                        pdf.

Przelaskowski:2001:TTM

[328] Artur Przelaskowski. Today’s and tomorrow’s medical imaging. Lecture Notes in Computer Science, 2124:236– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240236.
                                        pdf.

Pohle:2001:NAM

[329] Regina Pohle and Klaus D. Toennies. A new approach for modelbased adaptive region growing in medical image analysis. Lecture Notes in Computer Science, 2124:238–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240238.
                                        pdf.

Orkisz:2001:ABT

[330] Maciej Orkisz, Anne Frery, Olivier Chapet, Françoise Mornex, and Isabelle E. Magnin. Attempts to bronchial tumor motion tracking in portal images during conformal radiotherapy treatment. Lecture Notes in Computer Science, 2124:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240247.
pdf.

Nedzved:2001:CTA

[331] A. Nedzved, Y. Ilyich, S. Ablameyko, and S. Kamata. Color thinning with applications to biomedical images. Lecture Notes in Computer Science, 2124: 256–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240256.
pdf.

Desbleds-Mansard:2001:DAC

[332] Catherine Desbleds-Mansard, Alfred Anwander, Linda Chaabane, Maciej Orkisz, Bruno Neyran, Philippe C. Douek, and Isabelle E. Magnin. Dynamic active contour model for size independent blood vessel lumen segmentation and quantification in high-resolution magnetic resonance images. Lecture Notes in Computer Science, 2124:264–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240264.
pdf.

Ruminski:2001:MAT

[333] Jacek Ruminski, Mariusz Kaczmarek, and Antoni Nowakowski. Medical active thermography — A new image reconstruction method. Lecture Notes in Computer Science, 2124:274– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240274.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240274.
                                        pdf.

Jedrzejek:2001:DMP

[334] Czesław Jedrzejek, Andrzej Łempicki, Rafał Renk, and Jakub Radziulis. 3-D modeling and parametrisation of pelvis and hip joint. Lecture Notes in Computer Science, 2124:282–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240282.
                                        pdf.

Kowalski:2001:CRA

[335] Henryk A. Kowalski, Andrzej Skorupski, Zbigniew Szymanski, Wojciech Ziembla, Dariusz Wojciechowski, Piotr Sionek, and Piotr Jedrasik. Cardiac rhythm analysis using spatial ECG parameters and SPART method. Lecture Notes in Computer Science, 2124:290– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240290.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2124/21240290.
pdf.

Klemencic:2001:DCB

[336] Jan Klemenčič, Vojko Valenčič, and Nuška Pečarič. Deformable contour based algorithm for segmentation of the hippocampus from MRI. Lecture Notes in Computer Science, 2124:298– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240298.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240298.
pdf.

Gut:2001:EBR

[337] Piotr Gut, Leszek Chmielewski, Paweł Kukołowicz, and Andrzej Dabrowski. Edge-based robust image registration for incomplete and partly erroneous data. Lecture Notes in Computer Science, 2124:309–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240309.
pdf.

Sadeghi:2001:RTS

[338] Mohammad Sadeghi, Josef Kittler, and Kieron Messer. Real time segmentation of lip pixels for lip tracker initialization. Lecture Notes in Computer Science, 2124:317–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240317.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240317.
                                        pdf.

Chetverikov:2001:PIV

[339] Dmitry Chetverikov. Particle image velocimetry by feature tracking. Lecture Notes in Computer Science, 2124:325– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240325.
                                        pdf.

Condell:2001:EMT

[340] J. V. Condell, B. W. Scotney, and P. J. Morrow. Estimation of motion through inverse finite element methods with triangular meshes. Lecture Notes in Computer Science, 2124:333– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240333.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240333.
                                        pdf.

Buenaposada:2001:FTU

[341] José M. Buenaposada, David Sopeña, and Luis Baumela. Face tracking using the dynamic grey world algorithm. Lecture Notes in Computer Science, 2124: 341–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2124/21240341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240341.
pdf.

Neckels:2001:FLE

[342] Kai Neckels. Fast local estimation of optical flow using variational and wavelet methods. Lecture Notes in Computer Science, 2124:349–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240349.
pdf.

Suchecki:2001:MAM

[343] Witold Suchecki and Krzysztof Urbaniec. A method to analyse the motion of solid particle in oscillatory stream of a viscous liquid. Lecture Notes in Computer Science, 2124:357– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240357.
pdf.

Traver:2001:OAT

[344] V. Javier Traver and Filiberto Pla. An optimization approach for translational motion estimation in log-polar domain. Lecture Notes in Computer Science, 2124:365–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240365.
                                        pdf.

Pers:2001:TPS

[345] Janez Perš and Stanislav Kovačič. Tracking people in sport: Making use of partially controlled environment. Lecture Notes in Computer Science, 2124:374–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2124/21240374.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240374.
                                        pdf.

Ansar:2001:LAR

[346] Adnan Ansar and Kostas Daniilidis. Linear augmented reality registration. Lecture Notes in Computer Science, 2124:383–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240383.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240383.
                                        pdf.

Pawlowski:2001:SPD

[347] Michał Pawłowski and Małgorzata Kujawinska. Shape and position determination based on combination of photogrammetry with phase analysis of fringe patterns. Lecture Notes in Computer Science, 2124:391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240391.
pdf.

Wingbermuhle:2001:AAL

[348] Jochen Wingbermühle, Claus-Eberhard Liedtke, and Juri Solodenko. Automated acquisition of lifelike 3D human models from multiple posture data. Lecture Notes in Computer Science, 2124:400–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2124/21240400.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2124/21240400.
pdf.

Romaniuk:2001:ARS

[349] B. Romaniuk, M. Desvignes, J. Robiaille, M. Revenu, and M. J. Deshayes. Augmented reality and semiautomated landmarking of cephalometric radiographs. Lecture Notes in Computer Science, 2124:410–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240410.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240410.
pdf.

Skoneczny:2001:RDC

[350] Slawomir Skoneczny and Marcin Iwanowski. On restoration of degraded cinematic sequences by means of digital image processing. Lecture Notes in Computer Science, 2124:419– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240419.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240419.
pdf.

Gonzalez:2001:VBM

[351] Rafael C. González, Raul Valdés, and Jose A. Cancelas. Vision based measurement system to quantify straightness defect in steel sheets. Lecture Notes in Computer Science, 2124:427– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240427.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240427.
pdf.

Mielikainen:2001:PFB

[352] Jarno Mielikäinen, Ilkka Koskinen, Heikki Handroos, Pekka Toivanen, and Heikki Kälviäinen. Positioning of flexible boom structure using neural networks. Lecture Notes in Computer Science, 2124:435–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240435.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240435.
pdf.

Samek:2001:MIU

[353] Ota Samek, Vladislav Krzyžánek, David C. S. Beddows, Helmut H. Telle, Josef Kaiser, and Miroslav Liška. Material identification using laser spectroscopy and pattern recognition algorithms. Lecture Notes in Computer Science, 2124:443–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240443.
pdf.

Toczyski:2001:SSC

[354] Tomasz Toczyski and Slawomir Skoneczny. Scanner sequence compensation. Lecture Notes in Computer Science, 2124: 451–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240451.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240451.
pdf.

Svirsky:2001:IAI

[355] Dmitry Svirsky and Yuri Polozkov. The industrial application of the irregular 3D-objects image processing in the compact reverse engineering system. Lecture Notes in Computer Science, 2124:457–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240457.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2124/21240457.
                                        pdf.

Sluzek:2001:LAR

[356] Andrzej Sluzek. A local algorithm for real-time junction detection in contour images. Lecture Notes in Computer Science, 2124:465–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240465.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240465.
                                        pdf.

Dekeyser:2001:NAS

[357] Fabien Dekeyser, Patrick Bouthemy, and Patrick Pérez. A new algorithm for super-resolution from image sequences. Lecture Notes in Computer Science, 2124:473–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240473.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240473.
                                        pdf.

Kenmochi:2001:FAT

[358] Yukiko Kenmochi, Li Chunyan, and Kazunori Kotani. Flatness analysis of three-dimensional images for global polyhedrization. Lecture Notes in Computer Science, 2124:482–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240482.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2124/21240482.
pdf.

Iwanowski:2001:GMM

[359] Marcin Iwanowski. Generalized morphological mosaic interpolation and its application to computeraided animations. Lecture Notes in Computer Science, 2124:493–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240493.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240493.
pdf.

Terol-Villalobos:2001:OCR

[360] Iván R. Terol-Villalobos and Damián Vargas-Vázquez. Openings and closings by reconstruction using propagation criteria. Lecture Notes in Computer Science, 2124:502–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240502.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240502.
pdf.

Acharyya:2001:MSD

[361] Mausumi Acharyya and Malay K. Kundu. Multiscale segmentation of document images using M-band wavelets. Lecture Notes in Computer Science, 2124:510–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/2124/21240510.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2124/21240510.pdf.

Noakes:2001:LECa

[362] Lyle Noakes, Ryszard Kozera, and Reinhard Klette. Length estimation for curves with ε-uniform sampling. Lecture Notes in Computer Science, 2124: 518–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240518.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240518.
                                        pdf.

Smolka:2001:RWA

[363] B. Smolka, M. Szczepanski, K. N. Plataniotis, and A. N. Venetsanopoulos. Random walk approach to noise reduction in color images. Lecture Notes in Computer Science, 2124:527– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240527.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240527.
                                        pdf.

Hahn:2001:WDA

[364] Stefan L. Hahn and Kajetana M. Snopek. Wigner distributions and ambiguity functions in image analysis. Lecture Notes in Computer Science, 2124:537–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2124/21240537.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2124/21240537.
pdf.

Kato:2001:MRF

[365] Zoltan Kato and Ting-Chuen Pong. A Markov random field image segmentation model using combined color and texture features. Lecture Notes in Computer Science, 2124:547–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240547.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240547.
pdf.

Rital:2001:AAH

[366] Soufiane Rital, Alain Bretto, Driss Aboutajdine, and Hocine Cherifi. Application of adaptive hypergraph model to impulsive noise detection. Lecture Notes in Computer Science, 2124:555– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240555.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240555.
pdf.

Wiwart:2001:EFH

[367] Marian Wiwart, Irena Koczowska, and Andrzej Borusiewicz. Estimation of fusarium head blight of triticale using digital image analysis of grain. Lecture Notes in Computer Science, 2124: 563–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240563.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240563.
                                        pdf.

Smolka:2001:FMV

[368] B. Smolka, M. Szczepanski, K. N. Plataniotis, and A. N. Venetsanopoulos. Fast modified vector median filter. Lecture Notes in Computer Science, 2124: 570–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240570.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240570.
                                        pdf.

Kukielka:2001:HMD

[369] Grzegorz Kukiełka and Jerzy Woznicki. Hierarchical method of digital image segmentation using multidimensional mathematical morphology. Lecture Notes in Computer Science, 2124:581– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240581.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240581.
                                        pdf.

Krzyzanek:2001:IIO

[370] Vladislav Krzyžánek, Ota Samek, and Rudolf Reichelt. Images of imperfectly ordered motifs: Properties, analysis, and reconstruction. Lecture Notes in Computer Science, 2124:589– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240589.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240589.
pdf.

Saeed:2001:IAR

[371] Khalid Saeed, Mariusz Rybnik, and Marek Tabedzki. Implementation and advanced results on the non-interrupted skeletonization algorithm. Lecture Notes in Computer Science, 2124:601–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240601.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240601.
pdf.

Porikli:2001:OSC

[372] Fatih Murat Porikli. Object segmentation of color video sequences. Lecture Notes in Computer Science, 2124: 610–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240610.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240610.
pdf.

Sankowski:2001:TIS

[373] Dominik Sankowski and Volodymyr Mosorov. Thresholding image segmentation based on the volume analysis of spatial regions. Lecture Notes in Computer Science, 2124:620–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240620.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240620.
                                        pdf.

Pedrini:2001:TFI

[374] Hélio Pedrini and William Robson Schwartz. Topographic feature identification based on triangular meshes. Lecture Notes in Computer Science, 2124: 621–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240621.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240621.
                                        pdf.

Ouerhani:2001:VAG

[375] Nabil Ouerhani, Neculai Archip, Heinz Hügli, and Pierre-Jean Erard. Visual attention guided seed selection for color image segmentation. Lecture Notes in Computer Science, 2124:630– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240630.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240630.
                                        pdf.

Wei:2001:TAF

[376] Tiangong Wei and Reinhard Klette. Theoretical analysis of finite difference algorithms for linear shape from shading. Lecture Notes in Computer Science, 2124:638–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2124/21240638.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2124/21240638.pdf.

Luo:2001:RCP

[377] Bin Luo and Edwin R. Hancock. Relational constraints for point distribution models. Lecture Notes in Computer Science, 2124:646–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240646.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240646.
pdf.

Ragheb:2001:SSU

[378] Hossein Ragheb and Edwin R. Hancock. Shape-from-shading using Darboux smoothing. Lecture Notes in Computer Science, 2124:657–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240657.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240657.
pdf.

Yuen:2001:NRS

[379] Shiu Yin Yuen, Hoi Shan Lam, and Chun Ki Fong. A novel robust statistical design of the repeated genetic algorithm. Lecture Notes in Computer Science, 2124:668–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240668.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240668.
                                        pdf.

Jahn:2001:BSM

[380] Herbert Jahn. Binocular stereo matching by local attraction. Lecture Notes in Computer Science, 2124:676–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240676.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240676.
                                        pdf.

Wei:2001:CIA

[381] Shou-Kang Wei, Fay Huang, and Reinhard Klette. Characterizations of image acquisition and epipolar geometry of multiple panoramas. Lecture Notes in Computer Science, 2124:684– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240684.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240684.
                                        pdf.

Wilson:2001:IFR

[382] Malcolm Wilson. Interclass fuzzy rule generation for road scene recognition from colour images. Lecture Notes in Computer Science, 2124:692– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240692.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240692.
pdf.

Guillamet:2001:ULP

[383] David Guillamet and Jordi Vitrià. Unsupervised learning of part-based representations. Lecture Notes in Computer Science, 2124:700–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240700.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240700.
pdf.

Cyganek:2001:CSP

[384] Bogusław Cyganek and Jan Borgosz. A comparative study of performance and implementation of some areabased stereo algorithms. Lecture Notes in Computer Science, 2124:709– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2124/21240709.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2124/21240709.
pdf.

Fusiello:2001:NAA

[385] Andrea Fusiello. A new autocalibration algorithm: Experimental evaluation. Lecture Notes in Computer Science, 2124:717–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240717.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240717.
                                        pdf.

Kasprzak:2001:ICS

[386] Włodzimierz Kasprzak. An iconic classification scheme for video-based traffic sensor tasks. Lecture Notes in Computer Science, 2124:725–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240725.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240725.
                                        pdf.

Svoboda:2001:MCI

[387] Tomáš Svoboda and Tomáš Pajdla. Matching in catadioptric images with appropriate windows, and outliers removal. Lecture Notes in Computer Science, 2124:733–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2124/21240733.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2124/21240733.
                                        pdf.

Anonymous:2001:AIfn

[388] Anonymous. Author index. Lecture Notes in Computer Science, 2124:741– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/2124/2124auth.pdf.

Rinard:2001:AMP

[389] Martin Rinard. Analysis of multithreaded programs. Lecture Notes in Computer Science, 2126:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260001.
pdf.

Bertran:2001:CPI

[390] Miquel Bertran, Francesc Babot, August Climent, and Miquel Nicolau. Communication and parallelism introduction and elimination in imperative concurrent programs. Lecture Notes in Computer Science, 2126:20– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260020.
pdf.

Komondoor:2001:USI

[391] Raghavan Komondoor and Susan Horwitz. Using slicing to identify duplication in source code. Lecture Notes in Computer Science, 2126:40– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260040.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2126/21260040.
                                        pdf.

Sharp:2001:SSH

[392] Richard Sharp and Alan Mycroft. Soft scheduling for hardware. Lecture Notes in Computer Science, 2126:57– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260057.
                                        pdf.

Glynn:2001:ESA

[393] Kevin Glynn, Peter J. Stuckey, and Martin Sulzmann. Effective strictness analysis with HORN constraints. Lecture Notes in Computer Science, 2126: 73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260073.
                                        pdf.

Mesnard:2001:ASA

[394] Fred Mesnard and Ulrich Neumerkel. Applying static analysis techniques for inferring termination conditions of logic programs. Lecture Notes in Computer Science, 2126:93–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260093.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2126/21260093.
pdf.

Monniaux:2001:AAP

[395] David Monniaux. An abstract analysis of the probabilistic termination of programs. Lecture Notes in Computer Science, 2126:111–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260111.
pdf.

Spoto:2001:WST

[396] Fausto Spoto. Watchpoint semantics: A tool for compositional and focussed static analyses. Lecture Notes in Computer Science, 2126:127–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260127.
pdf.

Lu:2001:PGA

[397] Lunjin Lu. Parameterizing a groundness analysis of logic programs. Lecture Notes in Computer Science, 2126: 146–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260146.
pdf.

Bagnara:2001:FTA

[398] Roberto Bagnara, Roberta Gori, Patricia M. Hill, and Enea Zaffanella. Finite-tree analysis for constraint logicbased languages. Lecture Notes in Computer Science, 2126:165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260165.
                                        pdf.

Leino:2001:AES

[399] K. Rustan M. Leino. Applications of extended static checking. Lecture Notes in Computer Science, 2126:185– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260185.
                                        pdf.

Dor:2001:CCS

[400] Nurit Dor, Michael Rodeh, and Mooly Sagiv. Cleanness checking of string manipulations in C programs via integer analysis. Lecture Notes in Computer Science, 2126:194– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260194.
                                        pdf.

Liu:2001:SRT

[401] Yanhong A. Liu, Ning Li, and Scott D. Stoller. Solving regular tree grammar based constraints. Lecture Notes in Computer Science, 2126:213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260213.
pdf.

Goubault:2001:SAP

[402] Eric Goubault. Static analyses of the precision of floatingpoint operations. Lecture Notes in Computer Science, 2126:234–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260234.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260234.
pdf.

Das:2001:EIS

[403] Manuvir Das, Ben Liblit, Manuel Fähndrich, and Jakob Rehof. Estimating the impact of scalable pointer analysis on optimization. Lecture Notes in Computer Science, 2126:260– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260260.
pdf.

Liang:2001:ECP

[404] Donglin Liang and Mary Jean Harrold. Efficient computation of parameterized pointer information for interprocedural analyses. Lecture Notes in Computer Science, 2126:279–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260279.
                                        pdf.

Wang:2001:POO

[405] Farn Wang and Hsu-Chun Yen. Parametric optimization of open real-time systems. Lecture Notes in Computer Science, 2126:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260299.
                                        pdf.

Sidorova:2001:EC

[406] Natalia Sidorova and Martin Steffen. Embedding chaos. Lecture Notes in Computer Science, 2126:319– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260319.
                                        pdf.

Bouajjani:2001:AFP

[407] Ahmed Bouajjani, Aurore Collomb-Annichini, Yassine Lakhnech, and Mihaela Sighireanu. Analyzing fair parametric extended automata. Lecture Notes in Computer Science, 2126:335– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260335.
pdf.

Giacobazzi:2001:ICR

[408] Roberto Giacobazzi and Elisa Quintarelli. Incompleteness, counterexamples, and refinements in abstract model-checking. Lecture Notes in Computer Science, 2126:356–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2126/21260356.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2126/21260356.
pdf.

Schneider:2001:LBS

[409] Fred Schneider. Language-based security: What’s needed and why. Lecture Notes in Computer Science, 2126:374– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260374.
pdf.

Rajamani:2001:BMS

[410] Sriram K. Rajamani and Jakob Rehof. A behavioral module system for the pi-calculus. Lecture Notes in Computer Science, 2126:375–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260375.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260375.
                                        pdf.

Levi:2001:AIF

[411] Francesca Levi and Sergio Maffeis. An abstract interpretation framework for analysing mobile ambients. Lecture Notes in Computer Science, 2126:395– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2126/21260395.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2126/21260395.
                                        pdf.

Feret:2001:AIB

[412] Jérôme Feret. Abstract interpretationbased static analysis of mobile ambients. Lecture Notes in Computer Science, 2126:412–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2126/21260412.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2126/21260412.pdf.

Wagner:2001:SAS

[413] David Wagner. Static analysis and software assurance. Lecture Notes in Computer Science, 2126:431–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260431.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260431.
pdf.

Gordon:2001:TEA

[414] Andrew D. Gordon and Alan Jeffrey. A type and effect analysis of security protocols. Lecture Notes in Computer Science, 2126:432–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260432.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260432.
pdf.

Blanchet:2001:ACP

[415] Bruno Blanchet. Abstracting cryptographic protocols by Prolog rules. Lecture Notes in Computer Science, 2126: 433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2126/21260433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2126/21260433.
pdf.

Myers:2001:STL

[416] Andrew Myers. Security-typed languages and distributed computation.

Lecture Notes in Computer Science, 2126:437–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2126/21260437.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2126/21260437.
pdf.

Anonymous:2001:AIfo

[417] Anonymous. Author index. Lecture Notes in Computer Science, 2126:439– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2126/21260439.pdf.

Bandman:2001:HAR

[418] O. Bandman. A hybrid approach to reaction-diffusion processes simulation. Lecture Notes in Computer Science, 2127:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270001.
pdf.

Barrio-Solorzano:2001:FVC

[419] M. Barrio-Solórzano, M. E. Beato, C. E. Cuesta, and P. de la Fuente. Formal verification of coherence for a shared memory multiprocessor model. Lecture Notes in Computer Science, 2127:17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

0558/bibs/2127/21270017.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2127/21270017.
pdf.

Bodei:2001:SAS

[420] C. Bodei, P. Degano, F. Nielson, and H. Riis Nielson. Static analysis for secrecy and non-interference in networks of processes. Lecture Notes in Computer Science, 2127:27– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270027.
pdf.

Brasileiro:2001:COC

[421] F. Brasileiro, F. Greve, A. Mostefaoui, and M. Raynal. Consensus in one communication step. Lecture Notes in Computer Science, 2127:42– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270042.
pdf.

Hannig:2001:DSE

[422] F. Hannig and J. Teich. Design space exploration for massively parallel processor arrays. Lecture Notes in Computer Science, 2127:51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2127/21270051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270051.
                                        pdf.

Hoffmann:2001:GGC

[423] R. Hoffmann, K.-P. Völkmann, S. Waldschmidt, and W. Heenes. GCA: Global Cellular Automata. A flexible parallel model. Lecture Notes in Computer Science, 2127:66–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270066.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270066.
                                        pdf.

Markova:2001:CPA

[424] V. Markova. Cellular-pipelined algorithm architecture for polynomial computing. Lecture Notes in Computer Science, 2127:74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270074.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270074.
                                        pdf.

Mazzocca:2001:MNS

[425] N. Mazzocca, M. Rak, and U. Villano. MetaPL: A notation system for parallel program description and performance analysis. Lecture Notes in Computer Science, 2127:80– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2127/21270080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270080.
pdf.

Pudov:2001:FOC

[426] S. Pudov. First-order 2D cellular neural networks investigation and learning. Lecture Notes in Computer Science, 2127:94–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2127/21270094.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2127/21270094.
pdf.

Raynal:2001:QUR

[427] M. Raynal. Quiescent uniform reliable broadcast as an introduction to failure detector oracles. Lecture Notes in Computer Science, 2127:98– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270098.
pdf.

Segun:2001:TPM

[428] K. Segun, A. R. Hurson, and A. Spink. A transaction processing model for the mobile data access system. Lecture Notes in Computer Science, 2127:112– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270112.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270112.
                                        pdf.

Virbitskaite:2001:CTN

[429] I. B. Virbitskaite. Characterizing timed net processes categorically. Lecture Notes in Computer Science, 2127:128– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270128.
                                        pdf.

Vodovoz:2001:MHT

[430] N. Vodovoz. Mapping heterogeneous task graphs onto networks: Execution time optimization. Lecture Notes in Computer Science, 2127:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270142.
                                        pdf.

Zotov:2001:ZRP

[431] I. V. Zotov and V. S. Titov. An O[n^{3}/z^{3}] reduction procedure for determining the maximum degree of parallelism in parallel applications. Lecture Notes in Computer Science, 2127:150– 158, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270150.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2127/21270150.
pdf.

Alekseev:2001:AEA

[432] I. V. Alekseev and V. A. Sokolov. ARTCP: Efficient algorithm for transport protocol for packet switched networks. Lecture Notes in Computer Science, 2127:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2127/21270159.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2127/21270159.pdf.

Avetisyan:2001:EJE

[433] A. Avetisyan, S. Gaissaryan, and O. Samovarov. Extension of Java environment by facilities supporting development of SPMD Java-programs. Lecture Notes in Computer Science, 2127: 175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270175.
pdf.

Babenko:2001:MPC

[434] L. K. Babenko, A. G. Chefranov, P. A. Fedorov, A. Yu. Korobko, and O. B. Makarevich. Mechanisms of parallel computing organization for NeuroCluster. Lecture Notes in Computer Science, 2127:181– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2127/21270181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270181.
                                        pdf.

Babenko:2001:PST

[435] L. K. Babenko, A. G. Chefranov, and R. V. Trotsenko. Parallel SPMDtasks graph description language for network clusters. Lecture Notes in Computer Science, 2127:186–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270186.
                                        pdf.

Baude:2001:OMC

[436] F. Baude, D. Caromel, N. Furmento, and D. Sagnol. Optimizing metacomputing with communicationcomputation overlap. Lecture Notes in Computer Science, 2127:190–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270190.
                                        pdf.

Ciancarini:2001:WWA

[437] P. Ciancarini and D. Rossi. WebCluster: A Web-accessible cluster computing system based on coordination and mobility. Lecture Notes in Computer Science, 2127:205–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270205.
pdf.

Fahringer:2001:USE

[438] T. Fahringer, K. Sowa-Piekło, J. Luitz, and H. Moritsch. On using SPi-DER to examine and debug real-world data-parallel applications. Lecture Notes in Computer Science, 2127:211– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270211.
pdf.

Golenkov:2001:EVP

[439] E. A. Golenkov, A. S. Sokolov, G. V. Tarasov, and D. I. Kharitonov. Experimental version of parallel programs translator from Petri nets to C++. Lecture Notes in Computer Science, 2127: 226–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270226.
pdf.

Goossens:2001:TIC

[440] B. Goossens. Typing the ISA to cluster the processor. Lecture Notes in Computer Science, 2127:232–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270232.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270232.
                                        pdf.

Gorlatch:2001:SRC

[441] S. Gorlatch. Send-recv considered harmful? myths and truths about parallel programming. Lecture Notes in Computer Science, 2127:243–257, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270243.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270243.
                                        pdf.

Huber:2001:UGC

[442] V. Huber. UNICORE: A Grid computing environment for distributed and parallel computing. Lecture Notes in Computer Science, 2127:258–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270258.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270258.
                                        pdf.

Kopyssov:2001:PAM

[443] S. Kopyssov and A. Novikov. Parallel adaptive mesh refinement with load balancing for finite element method. Lecture Notes in Computer Science, 2127:266–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2127/21270266.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2127/21270266.
pdf.

Lammermann:2001:CIS

[444] S. Lämmermann, E. Tyugu, and V. Vlassov. Concurrent implementation of structurally synthesized programs. Lecture Notes in Computer Science, 2127:277–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270277.
pdf.

Nepomniaschaya:2001:AVB

[445] A. S. Nepomniaschaya. An associative version of the Bellman-Ford algorithm for finding the shortest paths in directed graphs. Lecture Notes in Computer Science, 2127:285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270285.
pdf.

Oyama:2001:FCI

[446] Y. Oyama, K. Taura, and A. Yonezawa. Fusion of concurrent invocations of exclusive methods. Lecture Notes in Computer Science, 2127:293–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270293.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270293.
                                        pdf.

Prokhorov:2001:CPR

[447] V. V. Prokhorov. Computational portal: Remote access to highperformance computing. Lecture Notes in Computer Science, 2127:308–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270308.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270308.
                                        pdf.

Ramirez:2001:ELP

[448] R. Ramirez and A. E. Santosa. Event logic programming. Lecture Notes in Computer Science, 2127:314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270314.
                                        pdf.

Shevchenko:2001:TIP

[449] R. Shevchenko and A. Doroshenko. Techniques for increasing performance of CORBA parallel distributed applications. Lecture Notes in Computer Science, 2127:319–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2127/21270319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270319.
pdf.

Schreiner:2001:MWP

[450] W. Schreiner. Manager-worker parallelism versus dataflow in a distributed computer algebra system. Lecture Notes in Computer Science, 2127:329– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270329.
pdf.

Sidorov:2001:CIC

[451] E. Sidorov, S. Bobkov, and S. Aryashev. Communication interface CoIn. Lecture Notes in Computer Science, 2127:344–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2127/21270344.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2127/21270344.
pdf.

Sosonkina:2001:DTP

[452] M. Sosonkina and G. Chen. Design of a tool for providing dynamic network information to an application. Lecture Notes in Computer Science, 2127:350– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270350.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270350.
                                        pdf.

Soula:2001:CPS

[453] J. Soula, Ph. Marquet, A. Demeure, and J.-L. Dekeyser. Compilation principle of a specification language dedicated to signal processing. Lecture Notes in Computer Science, 2127:358– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270358.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270358.
                                        pdf.

Thorelli:2001:ACP

[454] L.-E. Thorelli and V. Vlassov. An approach to composing parallel programs. Lecture Notes in Computer Science, 2127:371–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270371.
                                        pdf.

Ye:2001:WBP

[455] R. Ye, W.-J. Hsu, and Z.-H. Liu. Webbased parallel simulation of AGVs using Java and JINI. Lecture Notes in Computer Science, 2127:379–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270379.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2127/21270379.
pdf.

Babin:2001:PDD

[456] V. N. Babin, V. P. Il’in, and A. S. Pylkin. On the parallelization of domain decomposition methods for 3-D boundary value problems. Lecture Notes in Computer Science, 2127:385– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270385.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270385.
pdf.

Bandini:2001:PGP

[457] S. Bandini, G. Mauri, and G. Pavesi. Parallel generation of percolation beds based on stochastic cellular automata. Lecture Notes in Computer Science, 2127:391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2127/21270391.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2127/21270391.
pdf.

Bessonov:2001:PSI

[458] O. Bessonov, D. Fougère, and B. Roux. Parallel simulation of 3D incompressible flows and performance comparison for several MPP and cluster platforms. Lecture Notes in Computer Science, 2127:401–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270401.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270401.
                                        pdf.

Borshchev:2001:DSH

[459] A. Borshchev, Y. Karpov, and P. Lebedev. Distributed simulation of hybrid systems with HLA support. Lecture Notes in Computer Science, 2127:410– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270410.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270410.
                                        pdf.

Chambarel:2001:APC

[460] A. Chambarel and H. Bolvin. Application of the parallel computing technology to a wave front model using the finite element method. Lecture Notes in Computer Science, 2127:421– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270421.
                                        pdf.

Chambarel:2001:GPC

[461] A. Chambarel and D. Fougère. A general parallel computing approach using the finite element method and the objects oriented programming by selected data technique. Lecture Notes in Computer Science, 2127:428–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270428.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270428.
pdf.

Ignattieva:2001:PIC

[462] S. Ignattieva and V. Memnonov. Parallel implementation of a corrected DSMC method. Lecture Notes in Computer Science, 2127:436–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270436.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270436.
pdf.

Laevsky:2001:PAN

[463] Yu.M. Laevsky, P. V. Banushkina, S. A. Litvinenko, and A. A. Zotkevich. Parallel algorithms for nonstationary problems: Survey of new generation of explicit schemes. Lecture Notes in Computer Science, 2127: 442–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270442.
pdf.

Ludwig:2001:TEC

[464] T. Ludwig, M. Lindermeier, A. Stamatakis, and G. Rackl. Tool environments in CORBA-based medical high performance computing. Lecture Notes in Computer Science, 2127:447– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270447.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270447.
                                        pdf.

Mauri:2001:PAA

[465] G. Mauri and G. Pavesi. Parallel algorithms for the analysis of biological sequences. Lecture Notes in Computer Science, 2127:456–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270456.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270456.
                                        pdf.

Mikhailov:2001:SPM

[466] G. A. Mikhailov. Some parallel Monte Carlo algorithms. Lecture Notes in Computer Science, 2127:469– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270469.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270469.
                                        pdf.

Othman:2001:IPF

[467] M. Othman and A. R. Abdullah. Implementation of the parallel four points modified explicit group iterative algorithm on shared memory parallel computer. Lecture Notes in Computer Science, 2127:480–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270480.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270480.
pdf.

Pedretti:2001:PES

[468] K. Pedretti, T. Scheetz, T. Braun, C. Roberts, N. Robinson, and T. Casavant. A parallel expressed sequence tag (EST) clustering program. Lecture Notes in Computer Science, 2127: 490–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2127/21270490.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2127/21270490.
pdf.

Schmidt:2001:PSC

[469] B. Schmidt, H. Schröder, and M. Schimmler. Protein sequence comparison on the instruction systolic array. Lecture Notes in Computer Science, 2127:498–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2127/21270498.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2127/21270498.pdf.

Trinitis:2001:SBL

[470] C. Trinitis, M. Schulz, M. Eberl, and W. Karl. SCI-based LINUX PCclusters as a platform for electromagnetic field calculations. Lecture Notes in Computer Science, 2127:510– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2127/21270510.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2127/21270510.
                                        pdf.

Anonymous:2001:AIfp

[471] Anonymous. Author index. Lecture Notes in Computer Science, 2127:515– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2127/2127auth.pdf.

Desel:2001:WPN

[472] Jörg Desel and Gabriel Juhás. “What Is a Petri Net?”. Lecture Notes in Computer Science, 2128:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2128/21280001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2128/21280001.
                                        pdf.

Gajewsky:2001:TNB

[473] Meike Gajewsky and Hartmut Ehrig. The ^{≫}Petri Net Baukasten^{≪}: An overview. Lecture Notes in Computer Science, 2128:26–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2128/21280026.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2128/21280026.pdf.

Weber:2001:IUP

[474] Herbert Weber, Sabine Lembke, and Alexander Borusan. Improving the usability of Petri nets with the ^{≫}Petri Net Baukasten^{≪}. Lecture Notes in Computer Science, 2128:54–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2128/21280054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2128/21280054.
pdf.

Ermel:2001:IPN

[475] Claudia Ermel and Michael Weber. Implementation of parameterized net classes with the Petri net kernel of the ^{≫}Petri Net Baukasten^{≪}. Lecture Notes in Computer Science, 2128:79– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2128/21280079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2128/21280079.
pdf.

Gruhn:2001:PLM

[476] Volker Gruhn and Ursula Wellen. Process landscaping: Modelling distributed processes and proving properties of distributed process models. Lecture Notes in Computer Science, 2128: 103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2128/21280103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2128/21280103.
                                        pdf.

Desel:2001:PNP

[477] Jörg Desel, Gabriel Juhás, and Robert Lorenz. Petri nets over partial algebra. Lecture Notes in Computer Science, 2128:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2128/21280126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2128/21280126.
                                        pdf.

Padberg:2001:PNC

[478] Julia Padberg and Hartmut Ehrig. Parameterized net classes: A uniform approach to Petri net classes. Lecture Notes in Computer Science, 2128:173– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2128/21280173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2128/21280173.
                                        pdf.

Padberg:2001:BRC

[479] Julia Padberg, Hartmut Ehrig, and Grzegorz Rozenberg. Behavior and realization construction for Petri nets based on free monoid and power set graphs. Lecture Notes in Computer Science, 2128:230–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2128/21280230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2128/21280230.
pdf.

Stehr:2001:RLU

[480] Mark-Oliver Stehr, José Meseguer, and Peter Csaba Ölveczky. Rewriting logic as a unifying framework for Petri nets. Lecture Notes in Computer Science, 2128:250–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2128/21280250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2128/21280250.
pdf.

Badouel:2001:GAT

[481] Eric Badouel, Marek A. Bednarczyk, and Philippe Darondeau. Generalized automata and their net representations. Lecture Notes in Computer Science, 2128:304–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2128/21280304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2128/21280304.
pdf.

Bednarczyk:2001:CRR

[482] Marek A. Bednarczyk and Andrzej M. Borzyszkowski. On concurrent realization of reactive systems and their morphisms. Lecture Notes in Computer Science, 2128:346–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2128/21280346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2128/21280346.
                                        pdf.

Bruni:2001:TZS

[483] Roberto Bruni and Ugo Montanari. Transactions and zero-safe nets. Lecture Notes in Computer Science, 2128: 380–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2128/21280380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2128/21280380.
                                        pdf.

Bruni:2001:TAP

[484] Roberto Bruni and Vladimiro Sassone. Two algebraic process semantics for contextual nets. Lecture Notes in Computer Science, 2128:427– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2128/21280427.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2128/21280427.
                                        pdf.

Droste:2001:CPN

[485] Manfred Droste and R. M. Shortt. Continuous Petri nets and transition systems. Lecture Notes in Computer Science, 2128:457–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2128/21280457.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2128/21280457.
pdf.

Anonymous:2001:AIfq

[486] Anonymous. Author index. Lecture Notes in Computer Science, 2128:485– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2128/2128auth.pdf.

Goemans:2001:UCS

[487] Michel X. Goemans. Using complex semidefinite programming for approximating MAX E2–LIN3. Lecture Notes in Computer Science, 2129:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290001.
pdf.

Impagliazzo:2001:HCV

[488] Russell Impagliazzo. Hill-climbing vs. simulated annealing for planted bisection problems. Lecture Notes in Computer Science, 2129:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290002.
pdf.

Karlin:2001:WSH

[489] Anna R. Karlin. Web search via hub synthesis. Lecture Notes in Computer Science, 2129:6–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2129/21290006.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290006.
                                        pdf.

Trevisan:2001:ECC

[490] Luca Trevisan. Error-correcting codes and pseudorandom projections. Lecture Notes in Computer Science, 2129: 7–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2129/21290007.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290007.
                                        pdf.

Vadhan:2001:OP

[491] Salil P. Vadhan. Order in pseudorandomness. Lecture Notes in Computer Science, 2129:10–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2129/21290010.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2129/21290010.pdf.

Albers:2001:MST

[492] Susanne Albers and Carsten Witt. Minimizing stall time in single and parallel disk systems using multicommodity network flows. Lecture Notes in Computer Science, 2129:12– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290012.
pdf.

Bar-Yehuda:2001:EBP

[493] Reuven Bar-Yehuda and Dror Rawitz. On the equivalence between the primal-dual schema and the localratio technique. Lecture Notes in Computer Science, 2129:24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290024.
pdf.

Becchetti:2001:OWF

[494] Luca Becchetti, Stefano Leonardi, Alberto Marchetti-Spaccamela, and Kirk R. Pruhs. Online weighted flow time and deadline scheduling. Lecture Notes in Computer Science, 2129:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290036.
pdf.

Berman:2001:OAP

[495] Piotr Berman and Junichiro Fukuyama. An online algorithm for the postman problem with a small penalty. Lecture Notes in Computer Science, 2129:48– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2129/21290048.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290048.
                                        pdf.

Bumb:2001:SDA

[496] Adriana Bumb and Walter Kern. A simple dual ascent algorithm for the multilevel facility location problem. Lecture Notes in Computer Science, 2129:55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2129/21290055.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290055.
                                        pdf.

Caprara:2001:ASO

[497] Alberto Caprara, Hans Kellerer, and Ulrich Pferschy. Approximation schemes for ordered vector packing problems. Lecture Notes in Computer Science, 2129:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2129/21290063.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290063.
                                        pdf.

Dodis:2001:IC

[498] Yevgeniy Dodis and Shai Halevi. Incremental codes. Lecture Notes in Computer Science, 2129:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290075.
pdf.

Even:2001:AAA

[499] Guy Even, Jon Feldman, Guy Kortsarz, and Zeev Nutov. A 3/2approximation algorithm for augmenting the edge-connectivity of a graph from 1 to 2 using a subset of a given edge set. Lecture Notes in Computer Science, 2129:90– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290090.
pdf.

Garg:2001:AAB

[500] Rahul Garg, Vijay Kumar, and Vinayaka Pandit. Approximation algorithms for budget-constrained auctions. Lecture Notes in Computer Science, 2129:102–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2129/21290102.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2129/21290102.
pdf.

Halldorsson:2001:MAC

[501] Magnús M. Halldórsson, Guy Kortsarz, and Hadas Shachnai. Minimizing average completion of dedicated tasks and interval graphs. Lecture Notes in Computer Science, 2129:114– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290114.
pdf.

Mahdian:2001:GFL

[502] Mohammad Mahdian, Evangelos Markakis, Amin Saberi, and Vijay Vazirani. A greedy facility location algorithm analyzed using dual fitting. Lecture Notes in Computer Science, 2129:127–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290127.
pdf.

Matuura:2001:AAM

[503] Shiro Matuura and Tomomi Matsui. 0.863-approximation algorithm for MAX DICUT. Lecture Notes in Computer Science, 2129:138–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290138.
pdf.

Newman:2001:MAS

[504] Alantha Newman. The maximum acyclic subgraph problem and degree-3 graphs. Lecture Notes in Computer Science, 2129:147–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290147.
pdf.

Rodrigues:2001:SAR

[505] Estela Maris Rodrigues, Marie-France Sagot, and Yoshiko Wakabayashi. Some approximation results for the maximum agreement forest problem. Lecture Notes in Computer Science, 2129:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290159.
pdf.

Alon:2001:NOU

[506] Noga Alon, Michael Capalbo, Yoshiharu Kohayakawa, Vojtech Rödl, Andrzej Rucinski, and Endre Szemerédi. Near-optimum universal graphs for graphs with bounded degrees. Lecture Notes in Computer Science, 2129:170– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290170.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2129/21290170.
                                        pdf.

Amano:2001:GRP

[507] Kazuyuki Amano, John Tromp, Paul M. B. Vitányi, and Osamu Watanabe. On a generalized ruin problem. Lecture Notes in Computer Science, 2129: 181–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2129/21290181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290181.
                                        pdf.

Baltz:2001:PRA

[508] Andreas Baltz, Tomasz Schoen, and Anand Srivastav. On the bpartite random asymmetric traveling salesman problem and its assignment relaxation. Lecture Notes in Computer Science, 2129:192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2129/21290192.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290192.
                                        pdf.

Cho:2001:ESM

[509] Sung woo Cho and Ashish Goel. Exact sampling in machine scheduling problems. Lecture Notes in Computer Science, 2129:202–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2129/21290202.
                                        htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2129/21290202.pdf.

Clementi:2001:CAH

[510] Andrea E. F. Clementi, Pilu Crescenzi, Angelo Monti, Paolo Penna, and Riccardo Silvestri. On computing adhoc selective families. Lecture Notes in Computer Science, 2129:211–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290211.
pdf.

Coppersmith:2001:IE

[511] Don Coppersmith. L infinity embeddings. Lecture Notes in Computer Science, 2129:223–228, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290223.
pdf.

Dunagan:2001:EEB

[512] John Dunagan and Santosh Vempala. On Euclidean embeddings and bandwidth minimization. Lecture Notes in Computer Science, 2129:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290229.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2129/21290229.
                                        pdf.

Engebretsen:2001:NAN

[513] Lars Engebretsen. The non-approximability of non-Boolean predicates. Lecture Notes in Computer Science, 2129:241– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2129/21290241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290241.
                                        pdf.

Klivans:2001:DCD

[514] Adam R. Klivans. On the derandomization of constant depth circuits. Lecture Notes in Computer Science, 2129: 249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2129/21290249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290249.
                                        pdf.

Parnas:2001:TPL

[515] Michal Parnas, Dana Ron, and Ronitt Rubinfeld. Testing parenthesis languages. Lecture Notes in Computer Science, 2129:261–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2129/21290261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2129/21290261.
                                        pdf.

Parnas:2001:PDJ

[516] Michal Parnas, Dana Ron, and Alex Samorodnitsky. Proclaiming dictators and juntas or testing Boolean formulae. Lecture Notes in Computer Science, 2129:273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290273.
pdf.

Pemmaraju:2001:ECE

[517] Sriram V. Pemmaraju. Equitable coloring extends Chernoff– Hoeffding bounds. Lecture Notes in Computer Science, 2129:285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2129/21290285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2129/21290285.
pdf.

Anonymous:2001:AIfr

[518] Anonymous. Author index. Lecture Notes in Computer Science, 2129:297– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2129/21290297.pdf.

Grossberg:2001:CBA

[519] Stephen Grossberg. The complementary brain (abstract). Lecture Notes in Computer Science, 2130:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300003.
pdf.

Sperduti:2001:NNA

[520] Alessandro Sperduti. Neural networks for adaptive processing of structured data. Lecture Notes in Computer Science, 2130:5–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300005.
pdf.

Worgotter:2001:BDG

[521] Florentin Wörgötter. Bad design and good performance: Strategies of the visual system for enhanced scene analysis. Lecture Notes in Computer Science, 2130:13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300013.
pdf.

Schraudolph:2001:FCM

[522] Nicol N. Schraudolph. Fast curvature matrix-vector products. Lecture Notes in Computer Science, 2130:19– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300019.
pdf.

Dorronsoro:2001:ASN

[523] José R. Dorronsoro, Ana M. González, and Carlos Santa Cruz. Architecture selection in NLDA networks. Lecture Notes in Computer Science, 2130:27– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300027.
pdf.

RuizdeAngulo:2001:NLI

[524] Vicente Ruiz de Angulo and Carme Torras. Neural learning invariant to network size changes. Lecture Notes in Computer Science, 2130:33– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300033.
pdf.

Grandvalet:2001:BMM

[525] Yves Grandvalet, Florence d’Alché Buc, and Christophe Ambroise. Boosting mixture models for semisupervised learning. Lecture Notes in Computer Science, 2130:41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300041.
                                        pdf.

Grandvalet:2001:BCS

[526] Yves Grandvalet. Bagging can stabilize without reducing variance. Lecture Notes in Computer Science, 2130:49– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300049.
                                        pdf.

Muller:2001:SPM

[527] Achim F. Müller and Hans Georg Zimmermann. Symbolic prosody modeling by causal retro-causal NNs with variable context length. Lecture Notes in Computer Science, 2130:57– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300057.
                                        pdf.

Sato:2001:DDR

[528] Atsushi Sato. Discriminative dimensionality reduction based on generalized LVQ. Lecture Notes in Computer Science, 2130:65–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2130/21300065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300065.
pdf.

Nakayama:2001:CIA

[529] Hirotaka Nakayama, Masao Arakawa, and Rie Sasaki. A computational intelligence approach to optimization with unknown objective functions. Lecture Notes in Computer Science, 2130:73– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300073.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300073.
pdf.

Kaski:2001:CGE

[530] Samuel Kaski, Janne Sinkkonen, and Janne Nikkilä. Clustering gene expression data by mutual information with gene function. Lecture Notes in Computer Science, 2130:81– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300081.
pdf.

Hochreiter:2001:LLU

[531] Sepp Hochreiter, A. Steven Younger, and Peter R. Conwell. Learning to learn using gradient descent. Lecture Notes in Computer Science, 2130:87– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300087.
                                        pdf.

Faul:2001:VAR

[532] Anita C. Faul and Michael E. Tipping. A variational approach to robust regression. Lecture Notes in Computer Science, 2130:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300095.
                                        pdf.

Roberts:2001:MED

[533] Stephen J. Roberts, Christopher Holmes, and Dave Denison. Minimumentropy data clustering using reversible jump Markov chain Monte Carlo. Lecture Notes in Computer Science, 2130: 103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300103.
                                        pdf.

Dolnicar:2001:BMS

[534] Sara Dolničar and Friedrich Leisch. Behavioral market segmentation of binary guest survey data with bagged clustering. Lecture Notes in Computer Science, 2130:111–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2130/21300111.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2130/21300111.pdf.

Dermatas:2001:DEP

[535] Evangelos Dermatas. Direct estimation of polynomial densities in generalized RBF networks using moments. Lecture Notes in Computer Science, 2130: 119–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300119.
pdf.

Parra:2001:GIR

[536] Xavier Parra, Núria Agell, and Xari Rovira. Generalisation improvement of radial basis function networks based on qualitative input conditioning for financial credit risk prediction. Lecture Notes in Computer Science, 2130:127– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300127.
pdf.

Ito:2001:ABD

[537] Yoshifusa Ito and Cidambi Srinivasan. Approximation of Bayesian discriminant function by neural networks in terms of Kullback–Leibler information. Lecture Notes in Computer Science, 2130:135–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2130/21300135.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300135.
                                        pdf.

Mark:2001:BVD

[538] Zlochin Mark and Yoram Baram. The bias-variance dilemma of the Monte Carlo method. Lecture Notes in Computer Science, 2130:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300141.
                                        pdf.

Nishiyama:2001:MCM

[539] Takehiro Nishiyama, Kazuo Tsuchiya, and Katsuyoshi Tsujita. A Markov chain Monte Carlo algorithm for the quadratic assignment problem based on replicator equations. Lecture Notes in Computer Science, 2130:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300148.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300148.
                                        pdf.

Weeks:2001:MCM

[540] Michael Weeks, Jim Austin, Anthony Moulds, Aaron Turner, Zygmunt Ulanowski, and Julian Young. Mapping correlation matrix memory applications onto a Beowulf cluster. Lecture Notes in Computer Science, 2130: 156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300156.
pdf.

Strey:2001:ARN

[541] Alfred Strey and Martin Bange. Accelerating RBF network simulation by using multimedia extensions of modern microprocessors. Lecture Notes in Computer Science, 2130:164–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300164.
pdf.

Fung:2001:GTA

[542] Wai keung Fung and Yun hui Liu. A game-theoretic adaptive categorization mechanism for ARTtype networks. Lecture Notes in Computer Science, 2130:170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300170.
pdf.

Sandberg:2001:GRB

[543] Irwin W. Sandberg. Gaussian radial basis functions and innerproduct spaces. Lecture Notes in Computer Science, 2130:177–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300177.
                                        pdf.

Tanaka:2001:MPF

[544] Masahiro Tanaka. Mixture of probabilistic factor analysis model and its applications. Lecture Notes in Computer Science, 2130:183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300183.
                                        pdf.

Valls:2001:DLB

[545] José María Valls, Pedro Isasi, and Inés María Galván. Deferring the learning for better generalization in radial basis neural networks. Lecture Notes in Computer Science, 2130:189– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300189.
                                        pdf.

Lurette:2001:ICD

[546] Christophe Lurette and Stéphane Lecoeuche. Improvement of cluster detection and labeling neural network by introducing elliptical basis function. Lecture Notes in Computer Science, 2130:196–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300196.
pdf.

Lagus:2001:IVG

[547] Krista Lagus, Esa Alhoniemi, and Harri Valpola. Independent variable group analysis. Lecture Notes in Computer Science, 2130:203–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300203.
pdf.

Koksal:2001:WQM

[548] Fatih Köksal, Ethem Alpaydin, and Günhan Dündar. Weight quantization for multi-layer perceptrons using soft weight sharing. Lecture Notes in Computer Science, 2130:211– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300211.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2130/21300211.
                                        pdf.

Dimitriadou:2001:VME

[549] Evgenia Dimitriadou, Andreas Weingessel, and Kurt Hornik. Voting-merging: An ensemble method for clustering. Lecture Notes in Computer Science, 2130:217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2130/21300217.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300217.
                                        pdf.

Cannady:2001:AFA

[550] James Cannady and Raymond C. Garcia. The application of fuzzy ARTMAP in the detection of computer network attacks. Lecture Notes in Computer Science, 2130:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300225.
                                        pdf.

Li:2001:TLL

[551] Chun Hung Li and Pong Chi Yuen. Transductive learning: Learning iris data with two labeled data. Lecture Notes in Computer Science, 2130:231– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300231.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2130/21300231.
pdf.

Lewandowski:2001:ATV

[552] Achim Lewandowski and Peter Protzel. Approximation of time-varying functions with local regression models. Lecture Notes in Computer Science, 2130: 237–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300237.
pdf.

Schmitt:2001:CLN

[553] Michael Schmitt. Complexity of learning for networks of spiking neurons with nonlinear synaptic interactions. Lecture Notes in Computer Science, 2130:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2130/21300247.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2130/21300247.
pdf.

Schmitt:2001:PUN

[554] Michael Schmitt. Product unit neural networks with constant depth and superlinear VC dimension. Lecture Notes in Computer Science, 2130:253– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300253.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2130/21300253.
pdf.

Gavin:2001:GPP

[555] Gérald Gavin. Generalization performances of perceptrons. Lecture Notes in Computer Science, 2130:259– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300259.
pdf.

Teytaud:2001:BGA

[556] Olivier Teytaud and Hélène Paugam-Moisy. Bounds on the generalization ability of Bayesian inference and Gibbs algorithms. Lecture Notes in Computer Science, 2130:265–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300265.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300265.
pdf.

Malzahn:2001:LCG

[557] Dörthe Malzahn and Manfred Opper. Learning curves for Gaussian processes models: Fluctuations and universality. Lecture Notes in Computer Science, 2130:271–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300271.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2130/21300271.
pdf.

Kurkova:2001:TBR

[558] Věra Kǔrková and Marcello Sanguineti. Tight bounds on rates of neural-network approximation. Lecture Notes in Computer Science, 2130:277– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300277.
pdf.

Tresp:2001:SKS

[559] Volker Tresp and Anton Schwaighofer. Scalable kernel systems. Lecture Notes in Computer Science, 2130:285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300285.
pdf.

Oba:2001:LLM

[560] Shigeyuki Oba, Masa aki Sato, and Shin Ishii. On-line learning methods for Gaussian processes. Lecture Notes in Computer Science, 2130:292– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300292.
pdf.

Csato:2001:OAW

[561] Lehel Csató, Dan Cornford, and Manfred Opper. Online approximations for wind-field models. Lecture Notes in Computer Science, 2130:300–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300300.
                                        pdf.

Abe:2001:FTS

[562] Shigeo Abe and Takuya Inoue. Fast training of support vector machines by extracting boundary data. Lecture Notes in Computer Science, 2130:308– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300308.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300308.
                                        pdf.

Mayoraz:2001:MCP

[563] Eddy Nicolas Mayoraz. Multiclass classification with pairwise coupled neural networks or support vector machines. Lecture Notes in Computer Science, 2130:314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300314.
                                        pdf.

Ralaivola:2001:ISV

[564] Liva Ralaivola and Florence d’Alché Buc. Incremental support vector machine learning: A local approach. Lecture Notes in Computer Science, 2130: 322–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300322.
pdf.

Tsuda:2001:LPL

[565] Koji Tsuda, Gunnar Rätsch, Sebastian Mika, and Klaus-Robert Müller. Learning to predict the leave-one-out error of kernel based classifiers. Lecture Notes in Computer Science, 2130: 331–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300331.
pdf.

Roth:2001:SKR

[566] Volker Roth. Sparse kernel regressors. Lecture Notes in Computer Science, 2130:339–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2130/21300339.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2130/21300339.
pdf.

Graepel:2001:LGG

[567] Thore Graepel, Mike Goutrié, Marco Krüger, and Ralf Herbrich. Learning on graphs in the game of go. Lecture Notes in Computer Science, 2130:347– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300347.
pdf.

Melzer:2001:NFE

[568] Thomas Melzer, Michael Reiter, and Horst Bischof. Nonlinear feature extraction using generalized canonical correlation analysis. Lecture Notes in Computer Science, 2130:353–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300353.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300353.
pdf.

Amemori:2001:GPA

[569] Ken ichi Amemori and Shin Ishii. Gaussian process approach to stochastic spiking neurons with reset. Lecture Notes in Computer Science, 2130:361– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300361.
pdf.

Teytaud:2001:KBI

[570] Olivier Teytaud and David Sarrut. Kernel based image classification. Lecture Notes in Computer Science, 2130: 369–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300369.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300369.
pdf.

El-Beltagy:2001:GPM

[571] Mohammed A. El-Beltagy and W. Andy Wright. Gaussian processes for model fusion. Lecture Notes in Computer Science, 2130:376–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2130/21300376.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2130/21300376.pdf.

VanGestel:2001:KCC

[572] Tony Van Gestel, Johan A. K. Suykens, Jos De Brabanter, Bart De Moor, and Joos Vandewalle. Kernel canonical correlation analysis and least squares support vector machines. Lecture Notes in Computer Science, 2130:384– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300384.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300384.
pdf.

Frontzek:2001:LPN

[573] Thomas Frontzek, Thomas Navin Lal, and Rolf Eckmiller. Learning and prediction of the nonlinear dynamics of biological neurons with support vector machines. Lecture Notes in Computer Science, 2130:390–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300390.
pdf.

Sekhar:2001:CCS

[574] Chellu Chandra Sekhar, Kazuya Takeda, and Fumitada Itakura. Closeclass-set discrimination method for recognition of stop consonant-vowel utterances using support vector machines. Lecture Notes in Computer Science, 2130:399–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2130/21300399.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2130/21300399.pdf.

Kwok:2001:LDB

[575] James T. Kwok. Linear dependency between ϵ and the input noise in ϵsupport vector regression. Lecture Notes in Computer Science, 2130:405– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300405.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2130/21300405.
pdf.

Schwaighofer:2001:BCS

[576] Anton Schwaighofer and Volker Tresp. The Bayesian committee support vector machine. Lecture Notes in Computer Science, 2130:411–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300411.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300411.
pdf.

Tino:2001:UDC

[577] Peter Tiňo, Ian Nabney, and Yi Sun. Using directional curvatures to visualize folding patterns of the GTM projection manifolds. Lecture Notes in Computer Science, 2130:421–428, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300421.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300421.
pdf.

Martin-Merino:2001:SOM

[578] Manuel Martin-Merino and Alberto Muñoz. Self organizing map and sammon mapping for asymmetric proximities. Lecture Notes in Computer Science, 2130:429–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300429.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300429.
                                        pdf.

Milano:2001:ALA

[579] Michele Milano, Jürgen Schmidhuber, and Petros Koumoutsakos. Active learning with adaptive grids. Lecture Notes in Computer Science, 2130:436– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300436.
                                        pdf.

Diaz:2001:CPV

[580] Ignacio Díaz, Alberto B. Diez, and Abel A. Cuadrado Vega. Complex process visualization through continuous feature maps using radial basis functions. Lecture Notes in Computer Science, 2130:443–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300443.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300443.
                                        pdf.

Verbeek:2001:SSA

[581] Jakob J. Verbeek, Nikos Vlassis, and Ben Kröse. A soft k-segments algorithm for principal curves. Lecture Notes in Computer Science, 2130:450– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2130/21300450.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300450.
pdf.

Charalambous:2001:PPU

[582] Chris Charalambous, George C. Hadjinicola, and Eitan Muller. Product positioning using principles from the self-organizing map. Lecture Notes in Computer Science, 2130:457–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300457.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300457.
pdf.

VanLaerhoven:2001:CSO

[583] Kristof Van Laerhoven. Combining the self-organizing map and Kmeans clustering for on-line classification of sensor data. Lecture Notes in Computer Science, 2130:464–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300464.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300464.
pdf.

Atsalakis:2001:HBC

[584] Antonios Atsalakis, Ioannis Andreadis, and Nikos Papamarkos. Histogram based color reduction through selforganized neural networks. Lecture Notes in Computer Science, 2130:470– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300470.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300470.
                                        pdf.

Hattori:2001:SLS

[585] Motonobu Hattori, Hiroya Arisumi, and Hiroshi Ito. Sequential learning for SOM associative memory with map reconstruction. Lecture Notes in Computer Science, 2130:477–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300477.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300477.
                                        pdf.

Venna:2001:NPN

[586] Jarkko Venna and Samuel Kaski. Neighborhood preservation in nonlinear projection methods: An experimental study. Lecture Notes in Computer Science, 2130:485–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300485.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300485.
                                        pdf.

Yacoub:2001:THC

[587] Méziane Yacoub, Fouad Badran, and Sylvie Thiria. A topological hierarchical clustering: Application to ocean color classification. Lecture Notes in Computer Science, 2130:492– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300492.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300492.
pdf.

Dittenbach:2001:HCD

[588] Michael Dittenbach, Dieter Merkl, and Andreas Rauber. Hierarchical clustering of document archives with the growing hierarchical selforganizing map. Lecture Notes in Computer Science, 2130:500–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300500.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300500.
pdf.

Vollgraf:2001:BSS

[589] Roland Vollgraf, Ingo Schießl, and Klaus Obermayer. Blind source separation of single components from linear mixtures. Lecture Notes in Computer Science, 2130:509–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300509.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300509.
pdf.

Diamantaras:2001:BSS

[590] Konstantinos I. Diamantaras. Blind source separation using principal component neural networks. Lecture Notes in Computer Science, 2130:515–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300515.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300515.
                                        pdf.

Martin-Clemente:2001:BSS

[591] Rubén Martín-Clemente, José I. Acha, and Carlos G. Puntonet. Blind separation of sources by differentiating the output cumulants and using Newton’s method. Lecture Notes in Computer Science, 2130:521–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300521.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300521.
                                        pdf.

Roberts:2001:MIC

[592] Stephen J. Roberts and William D. Penny. Mixtures of independent component analysers. Lecture Notes in Computer Science, 2130:527–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300527.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300527.
                                        pdf.

Akaho:2001:CIC

[593] Shotaro Akaho. Conditionally independent component extraction for naive Bayes inference. Lecture Notes in Computer Science, 2130:535–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300535.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300535.
pdf.

Vlassis:2001:FSF

[594] Nikos Vlassis. Fast score function estimation with application in ICA. Lecture Notes in Computer Science, 2130: 541–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300541.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300541.
pdf.

Ypma:2001:HML

[595] Alexander Ypma, Co Melissant, Ole Baunbæk-Jensen, and Robert P. W. Duin. Health monitoring with learning methods. Lecture Notes in Computer Science, 2130:547–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300547.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300547.
pdf.

Christoyianni:2001:BTC

[596] Ioanna Christoyianni, Athanasios Koutras, Evangelos Dermatas, and George Kokkinakis. Breast tissue classification in mammograms using ICA mixture models. Lecture Notes in Computer Science, 2130:554–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300554.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300554.
pdf.

Koutras:2001:NNB

[597] Athanasios Koutras, Evangelos Dermatas, and George Kokkinakis. Neural network based blind source separation of non-linear mixtures. Lecture Notes in Computer Science, 2130:561– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300561.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300561.
pdf.

Kwak:2001:FEU

[598] Nojun Kwak, Chong-Ho Choi, and Jin Young Choi. Feature extraction using ICA. Lecture Notes in Computer Science, 2130:568–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300568.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300568.
pdf.

Trentin:2001:CSR

[599] Edmondo Trentin and Marco Gori. Continuous speech recognition with a robust connectionist/Markovian hybrid model. Lecture Notes in Computer Science, 2130:577–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2130/21300577.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2130/21300577.pdf.

Bouchard:2001:FCI

[600] Martin Bouchard. Faster convergence and improved performance in leastsquares training of neural networks for active sound cancellation. Lecture Notes in Computer Science, 2130:583– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300583.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300583.
pdf.

Potamitis:2001:BIC

[601] Ilyas Potamitis, Nikos Fakotakis, and George Kokkinakis. Bayesian independent component analysis as applied to one-channel speech enhancement. Lecture Notes in Computer Science, 2130: 593–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300593.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300593.
pdf.

Lu:2001:MPC

[602] Bao-Liang Lu, Jonghan Shin, and Michinori Ichikawa. Massively parallel classification of EEG signals using minmax modular neural networks. Lecture Notes in Computer Science, 2130:601– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300601.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300601.
                                        pdf.

Flexer:2001:STE

[603] Arthur Flexer, Herbert Bauer, Claus Lamm, and Georg Dorffner. Single trial estimation of evoked potentials using Gaussian mixture models with integrated noise component. Lecture Notes in Computer Science, 2130:609– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300609.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300609.
                                        pdf.

Sykacek:2001:PAH

[604] Peter Sykacek, Stephen Roberts, Iead Rezek, Arthur Flexer, and Georg Dorffner. A probabilistic approach to high-resolution sleep analysis. Lecture Notes in Computer Science, 2130:617– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300617.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2130/21300617.
pdf.

Cherkassky:2001:CWT

[605] Vladimir Cherkassky and Steven Kilts. Comparison of wavelet thresholding methods for denoising ECG signals. Lecture Notes in Computer Science, 2130:625–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300625.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300625.
pdf.

Sita:2001:EPS

[606] G. Sita and A. G. Ramakrishnan. Evoked potential signal estimation using Gaussian radial basis function network. Lecture Notes in Computer Science, 2130:630–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300630.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300630.
pdf.

Obermaier:2001:VKC

[607] Bernhard Obermaier, Gernot Müller, and Gert Pfurtscheller. ‘virtual keyboard’ controlled by spontaneous EEG activity. Lecture Notes in Computer Science, 2130:636–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2130/21300636.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300636.
                                        pdf.

Sommer:2001:CES

[608] David Sommer and Martin Golz. Clustering of EEG-segments using hierarchical agglomerative methods and self-organizing maps. Lecture Notes in Computer Science, 2130:642–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300642.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300642.
                                        pdf.

Woon:2001:NSP

[609] Wei Lee Woon and David Lowe. Nonlinear signal processing for noise reduction of unaveraged single channel MEG data. Lecture Notes in Computer Science, 2130:650–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300650.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300650.
                                        pdf.

Hochreiter:2001:DPM

[610] Sepp Hochreiter and Michael C. Mozer. A discrete probabilistic memory model for discovering dependencies in time. Lecture Notes in Computer Science, 2130:661–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300661.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300661.
pdf.

Gers:2001:ALT

[611] Felix A. Gers, Douglas Eck, and Jürgen Schmidhuber. Applying LSTM to time series predictable through timewindow approaches. Lecture Notes in Computer Science, 2130:669–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300669.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300669.
pdf.

Strickert:2001:GRL

[612] Marc Strickert, Thorsten Bojer, and Barbara Hammer. Generalized relevance LVQ for time series. Lecture Notes in Computer Science, 2130:677– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300677.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300677.
pdf.

Klapper-Rybicka:2001:ULL

[613] Magdalena Klapper-Rybicka, Nicol N. Schraudolph, and Jürgen Schmidhuber. Unsupervised learning in LSTM recurrent neural networks. Lecture Notes in Computer Science, 2130:684– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300684.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300684.
                                        pdf.

Murata:2001:AKB

[614] Hiroshi Murata and Takashi Onoda. Applying kernel based subspace classification to a non-intrusive monitoring for household electric appliances. Lecture Notes in Computer Science, 2130: 692–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300692.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300692.
                                        pdf.

Plebe:2001:NNC

[615] Alessio Plebe, A. Marcello Anile, and Salvatore Rinaudo. Neural networks in circuit simulators. Lecture Notes in Computer Science, 2130:699–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300699.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300699.
                                        pdf.

Camps:2001:NNE

[616] Gustavo Camps, Emilio Soria, José D. Martín, Antonio J. Serrano, Juan J. Ruixo, and N. Víctor Jiménez. Neural networks ensemble for cyclosporine concentration monitoring. Lecture Notes in Computer Science, 2130:706– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300706.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300706.
pdf.

Inoue:2001:EHN

[617] Hirotaka Inoue, Yoshinobu Fukunaga, and Hiroyuki Narihisa. Efficient hybrid neural network for chaotic time series prediction. Lecture Notes in Computer Science, 2130:712–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300712.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300712.
pdf.

Perez-Ortiz:2001:OSS

[618] Juan Antonio Pérez-Ortiz, Jorge Calera-Rubio, and Mikel L. Forcada. Online symbolic-sequence prediction with discrete-time recurrent neural networks. Lecture Notes in Computer Science, 2130:719–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300719.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300719.
pdf.

Kaleta:2001:PSB

[619] Stanislav Kaleta, Daniel Novotný, and Peter Sinčák. Prediction systems based on FIR BP neural networks. Lecture Notes in Computer Science, 2130:725– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300725.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300725.
                                        pdf.

Hammer:2001:GAR

[620] Barbara Hammer. On the generalization ability of recurrent networks. Lecture Notes in Computer Science, 2130: 731–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300731.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300731.
                                        pdf.

Cernansky:2001:FSR

[621] Michal Čerňanský and Lubica Beňušková. Finite-state Reber automaton and the recurrent neural networks trained in supervised and unsupervised manner. Lecture Notes in Computer Science, 2130:737–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2130/21300737.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300737.
                                        pdf.

Turchenko:2001:ECC

[622] Volodymyr Turchenko, Volodymyr Kochan, and Anatoly Sachenko. Estimation of computational complexity of sensor accuracy improvement algorithm based on neural networks. Lecture Notes in Computer Science, 2130: 743–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300743.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300743.
pdf.

Dietrich:2001:FAC

[623] Christian Dietrich, Friedhelm Schwenker, and Günther Palm. Fusion architectures for the classification of time series. Lecture Notes in Computer Science, 2130:749–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2130/21300749.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2130/21300749.pdf.

Edmonds:2001:IRC

[624] Bruce Edmonds and Scott Moss. The importance of representing cognitive processes in multi-agent models. Lecture Notes in Computer Science, 2130: 759–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300759.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300759.
pdf.

Zimmermann:2001:MAF

[625] Georg Zimmermann, Ralph Neuneier, and Ralph Grothmann. Multiagent FX-market modeling based on cognitive systems. Lecture Notes in Computer Science, 2130:767–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300767.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300767.
pdf.

Kaizoji:2001:SDH

[626] Taisei Kaizoji. Speculative dynamics in a heterogeneous-agent model. Lecture Notes in Computer Science, 2130: 775–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300775.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300775.
pdf.

Gaunersdorfer:2001:NAB

[627] Andrea Gaunersdorfer and Cars H. Hommes. Nonlinear adaptive beliefs and the dynamics of financial markets: The role of the evolutionary fitness measure. Lecture Notes in Computer Science, 2130:782–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300782.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300782.
pdf.

Hruschka:2001:APD

[628] Harald Hruschka, Werner Fettes, and Markus Probst. Analyzing purchase data by a neural net extension of the multinomial logit model. Lecture Notes in Computer Science, 2130:790– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300790.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300790.
pdf.

Wersing:2001:UMR

[629] Heiko Wersing and Helge Ritter. Using maximal recurrence in linear threshold competitive layer networks. Lecture Notes in Computer Science, 2130: 799–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300799.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300799.
pdf.

Sima:2001:ETC

[630] Jiří Šíma and Pekka Orponen. Exponential transients in continuous-time symmetric Hopfield nets. Lecture Notes in Computer Science, 2130:806– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300806.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300806.
pdf.

deGaris:2001:IER

[631] Hugo de Garis, Andrzej Buller, Leo de Penning, Tomasz Chodakowski, and Derek Decesare. Initial evolution results on CAM-brain machines (CBMs). Lecture Notes in Computer Science, 2130:814–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300814.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300814.
pdf.

Teuscher:2001:SOT

[632] Christof Teuscher and Eduardo Sanchez. Self-organizing topology evolution of Turing neural networks. Lecture Notes in Computer Science, 2130:820–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300820.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300820.
pdf.

Jain:2001:EPD

[633] Brijnesh J. Jain and Fritz Wysotzki. Efficient pattern discrimination with inhibitory WTA nets. Lecture Notes in Computer Science, 2130:827–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300827.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300827.
pdf.

Kamimura:2001:CIC

[634] Ryotaro Kamimura and Taeko Kamimura. Cooperative information control to coordinate competition and cooperation. Lecture Notes in Computer Science, 2130:835–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2130/21300835.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2130/21300835.
pdf.

Kuroe:2001:QAC

[635] Yasuaki Kuroe, Naoki Hashimoto, and Takehiro Mori. Qualitative analysis of continuous complex-valued associative memories. Lecture Notes in Computer Science, 2130:843–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300843.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300843.
pdf.

Goerke:2001:SOP

[636] Nils Goerke, Florian Kintzler, and Rolf Eckmiller. Self organized partitioning of chaotic attractors for control. Lecture Notes in Computer Science, 2130: 851–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300851.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300851.
pdf.

Wright:2001:GMS

[637] W. Andy Wright, Robert E. Smith, Martin Danek, and Pillip Greenway. A generalisable measure of selforganisation and emergence. Lecture Notes in Computer Science, 2130:857– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300857.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300857.
pdf.

Kwee:2001:MBR

[638] Ivo Kwee, Marcus Hutter, and Jürgen Schmidhuber. Market-based reinforcement learning in partially observable worlds. Lecture Notes in Computer Science, 2130:865–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300865.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300865.
pdf.

Wright:2001:SSL

[639] W. Andy Wright. Sequential strategy for learning multi-stage multiagent collaborative games. Lecture Notes in Computer Science, 2130:874– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300874.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300874.
pdf.

Stephan:2001:NAM

[640] Volker Stephan and Horst-Michael Gross. Neural architecture for mental imaging of sequences based on optical flow predictions. Lecture Notes in Computer Science, 2130:885–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300885.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300885.
pdf.

Heidemann:2001:VCG

[641] Gunther Heidemann and Helge Ritter. Visual checking of grasping positions of a three-fingered robot hand. Lecture Notes in Computer Science, 2130: 891–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300891.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300891.
pdf.

Heinze:2001:ABC

[642] Andrea Heinze and Horst-Michael Gross. Anticipation-based control architecture for a mobile robot. Lecture Notes in Computer Science, 2130:899– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300899.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300899.
pdf.

Saadia:2001:NAFb

[643] N. Saadia, Y. Amirat, J. Pontnaut, and A. Ramdane-Cherif. Neural adaptive force control for compliant robots. Lecture Notes in Computer Science, 2130: 906–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300906.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300906.
                                        pdf.

Suzuki:2001:DNN

[644] Michiyo Suzuki, Toru Yamamoto, Kazuo Kawada, and Hiroyuki Sogo. A design of neural-net based selftuning PID controllers. Lecture Notes in Computer Science, 2130:914–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300914.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300914.
                                        pdf.

Tang:2001:KCO

[645] Wai Sum Tang, Cherry Miu Ling Lam, and Jun Wang. Kinematic control and obstacle avoidance for redundant manipulators using a recurrent neural network. Lecture Notes in Computer Science, 2130:922–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300922.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2130/21300922.
pdf.

Baruch:2001:ANC

[646] Ieroham Baruch, Jose Martin Flores, Federico Thomas, and Ruben Garrido. Adaptive neural control of nonlinear systems. Lecture Notes in Computer Science, 2130:930–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300930.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300930.
pdf.

Hu:2001:HMT

[647] Jinglu Hu and Kotaro Hirasawa. A hierarchical method for training embedded sigmoidal neural networks. Lecture Notes in Computer Science, 2130: 937–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300937.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300937.
pdf.

Frontzek:2001:TLP

[648] Thomas Frontzek, Thomas Navin Lal, and Rolf Eckmiller. Towards learning path planning for solving complex robot tasks. Lecture Notes in Computer Science, 2130:943–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300943.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300943.
                                        pdf.

Al-Duwaish:2001:HMI

[649] Hussain N. Al-Duwaish and S. Saad Azhar Ali. Hammerstein model identification using radial basis functions neural networks. Lecture Notes in Computer Science, 2130:951–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300951.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300951.
                                        pdf.

Hulse:2001:ENB

[650] Martin Hülse, Bruno Lara, Frank Pasemann, and Ulrich Steinmetz. Evolving neural behaviour control for autonomous robots. Lecture Notes in Computer Science, 2130:957–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300957.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300957.
                                        pdf.

Panangadan:2001:CAA

[651] Anand Panangadan and Michael G. Dyer. Construction by autonomous agents in a simulated environment. Lecture Notes in Computer Science, 2130: 963–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2130/21300963.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300963.
pdf.

Katayama:2001:NCM

[652] Masazumi Katayama. A neural control model using predictive adjustment mechanism of viscoelastic property of the human arm. Lecture Notes in Computer Science, 2130:971– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300971.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300971.
pdf.

Wada:2001:MJA

[653] Yasuhiro Wada, Yuichi Kaneko, and Eri Nakano. Multi joint arm trajectory formation based on the minimization principle using the Euler– Poisson equation. Lecture Notes in Computer Science, 2130:977–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21300977.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21300977.
pdf.

Fukushima:2001:NNV

[654] Kunihiko Fukushima. Neocognitron of a new version: Handwritten digit recognition. Lecture Notes in Computer Science, 2130:987–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300987.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300987.
                                        pdf.

Cozzi:2001:CCR

[655] Alex Cozzi, Myron Flickner, Jainchang Mao, and Shivakumar Vaithyanathan. A comparison of classifiers for realtime eye detection. Lecture Notes in Computer Science, 2130:993–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21300993.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21300993.
                                        pdf.

Wismuller:2001:NNA

[656] Axel Wismüller, Oliver Lange, Dominik R. Dersch, Klaus Hahn, and Gerda L. Leinsinger. Neural network analysis of dynamic contrastenhanced MRI mammography. Lecture Notes in Computer Science, 2130:1000– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301000.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301000.
                                        pdf.

Atsalakis:2001:NAC

[657] Antonios Atsalakis, Nikos Papamarkos, and Charalambos Strouthopoulos. A new adaptive color quantization technique. Lecture Notes in Computer Science, 2130:1006–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301006.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301006.
pdf.

Kuzmina:2001:TON

[658] Margarita G. Kuzmina, Eduard A. Manykin, and Irina I. Surina. Tunable oscillatory network for visual image segmentation. Lecture Notes in Computer Science, 2130:1013–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301013.
pdf.

Jang:2001:DST

[659] Seok-Woo Jang, Gye-Young Kim, and Hyung-Il Choi. Detecting shot transitions for video indexing with FAM. Lecture Notes in Computer Science, 2130:1020–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301020.
pdf.

Wieghardt:2001:FFC

[660] Jan Wieghardt and Hartmut S. Loos. Finding faces in cluttered still images with few examples. Lecture Notes in Computer Science, 2130:1026– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301026.
pdf.

Chella:2001:DDS

[661] Antonio Chella, Maria Donatella Guarino, and Roberto Pirrone. Description of dynamic structured scenes by a SOM/ARSOM hierarchy. Lecture Notes in Computer Science, 2130:1034– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301034.
pdf.

Huang:2001:EDM

[662] Yin Huang, Ponnuthurai N. Suganthan, Shankar M. Krishnan, and Xiang Cao. Evaluation of distance measures for partial image retrieval using self-organizing map. Lecture Notes in Computer Science, 2130:1042– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301042.
pdf.

Cao:2001:VSB

[663] Xiang Cao and Ponnuthurai N. Suganthan. Video sequence boundary detection using neural gas networks. Lecture Notes in Computer Science, 2130:1048– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301048.
pdf.

Votsis:2001:NNB

[664] George Votsis, Nikolaos D. Doulamis, Anastasios D. Doulamis, Nicolas Tsapatsoulis, and Stefanos D. Kollias. A neural-network-based approach to adaptive human computer interaction. Lecture Notes in Computer Science, 2130:1054–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301054.
pdf.

Doulamis:2001:ANN

[665] Anastasios D. Doulamis, Klimis S. Ntalianis, Nikolaos D. Doulamis, and Stefanos D. Kollias. Adaptable neural networks for unsupervised video object segmentation of stereoscopic sequences. Lecture Notes in Computer Science, 2130:1060–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2130/21301060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301060.
                                        pdf.

Kikuchi:2001:MBO

[666] Masayuki Kikuchi and Youhei Akashi. A model of border-ownership coding in early vision. Lecture Notes in Computer Science, 2130:1069–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301069.
                                        pdf.

Kayser:2001:ESS

[667] Christoph Kayser, Wolfgang Einhäuser, Olaf Dümmer, Peter König, and Konrad Körding. Extracting slow subspaces from natural videos leads to complex cells. Lecture Notes in Computer Science, 2130:1075–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301075.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301075.
                                        pdf.

Wilke:2001:NCD

[668] Stefan D. Wilke. Neural coding of dynamic stimuli. Lecture Notes in Computer Science, 2130:1081–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2130/21301081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301081.
pdf.

Amemori:2001:RSS

[669] Ken ichi Amemori and Shin Ishii. Resonance of a stochastic spiking neuron mimicking the Hodgkin– Huxley model. Lecture Notes in Computer Science, 2130:1087–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301087.
pdf.

Curuklu:2001:SBS

[670] Baran Çürüklü and Anders Lansner. Spike and burst synchronization in a detailed cortical network model with I–F neurons. Lecture Notes in Computer Science, 2130:1095–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301095.
pdf.

Smith:2001:UDS

[671] Leslie S. Smith. Using depressing synapses for phase locked auditory onset detection. Lecture Notes in Computer Science, 2130:1103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301103.
                                        pdf.

Haschke:2001:COB

[672] Robert Haschke, Jochen J. Steil, and Helge Ritter. Controlling oscillatory behaviour of a two neuron recurrent neural network using inputs. Lecture Notes in Computer Science, 2130:1109– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301109.
                                        pdf.

Porr:2001:THL

[673] Bernd Porr and Florentin Wörgötter. Temporal Hebbian learning in ratecoded neural networks: A theoretical approach towards classical conditioning. Lecture Notes in Computer Science, 2130:1115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301115.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301115.
                                        pdf.

Yamazaki:2001:MAC

[674] Tadashi Yamazaki. A mathematical analysis of a correlation based model for the orientation map formation. Lecture Notes in Computer Science, 2130:1121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301121.
pdf.

Dauce:2001:LCM

[675] Emmanuel Daucé. Learning from chaos: A model of dynamical perception. Lecture Notes in Computer Science, 2130:1129–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2130/21301129.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2130/21301129.pdf.

Misjak:2001:EMC

[676] Fanni Misják, Máté Lengyel, and Péter Érdi. Episodic memory and cognitive map in a rate model network of the rat hippocampus. Lecture Notes in Computer Science, 2130:1135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301135.
pdf.

Schauer:2001:MHO

[677] Carsten Schauer and Horst-Michael Gross. A model of horizontal 360^{∘} object localization based on binaural hearing and monocular vision. Lecture Notes in Computer Science, 2130:1141– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301141.
                                        pdf.

Weber:2001:SOO

[678] Cornelius Weber. Self-organization of orientation maps, lateral connections, and dynamic receptive fields in the primary visual cortex. Lecture Notes in Computer Science, 2130:1147– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301147.
                                        pdf.

Sakumura:2001:MCM

[679] Yuichi Sakumura, Norio Konno, and Kazuyuki Aihara. Markov chain model approximating the Hodgkin– Huxley neuron. Lecture Notes in Computer Science, 2130:1153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301153.
                                        pdf.

Wrigley:2001:NOM

[680] Stuart N. Wrigley and Guy J. Brown. A neural oscillator model of auditory attention. Lecture Notes in Computer Science, 2130:1163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301163.
pdf.

Oudeyer:2001:CNM

[681] Pierre yves Oudeyer. Coupled neural maps for the origins of vowel systems. Lecture Notes in Computer Science, 2130:1171–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301171.
pdf.

Amini:2001:LTS

[682] Massih-Reza Amini and Patrick Gallinari. Learning for text summarization using labeled and unlabeled sentences. Lecture Notes in Computer Science, 2130:1177–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301177.
pdf.

Ma:2001:LED

[683] Qing Ma, Bao-Liang Lu, Masaki Murata, Michnori Ichikawa, and Hitoshi Isahara. On-line error detection of annotated corpus using modular neural networks. Lecture Notes in Computer Science, 2130:1185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301185.
                                        pdf.

Kim:2001:IBM

[684] DaeEun Kim and Jaeho Lee. Instancebased method to extract rules from neural networks. Lecture Notes in Computer Science, 2130:1193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301193.
                                        pdf.

Hodge:2001:NBS

[685] Victoria J. Hodge and Jim Austin. A novel binary spell checker. Lecture Notes in Computer Science, 2130:1199– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301199.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301199.
                                        pdf.

Taylor:2001:NNS

[686] Neill Taylor and John Taylor. Neural nets for short movements in natural language processing. Lecture Notes in Computer Science, 2130:1205– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301205.
pdf.

Koskela:2001:UDF

[687] Timo Koskela, Jukka Heikkonen, and Kimmo Kaski. Using document features to optimize Web cache. Lecture Notes in Computer Science, 2130:1211– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301211.
pdf.

Wada:2001:GDC

[688] Yasuhiro Wada, Kei Ohkawa, and Keiichi Sumita. Generation of diversiform characters using a computational handwriting model and a genetic algorithm. Lecture Notes in Computer Science, 2130:1217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301217.
pdf.

Kamimura:2001:IML

[689] Ryotaro Kamimura and Taeko Kamimura. Information maximization and language acquisition. Lecture Notes in Computer Science, 2130:1225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301225.
                                        pdf.

Womble:2001:MNS

[690] Steve Womble and Stefan Wermter. A mirror neuron system for syntax acquisition. Lecture Notes in Computer Science, 2130:1233–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301233.
                                        pdf.

Eck:2001:NRO

[691] Douglas Eck. A network of relaxation oscillators that finds downbeats in rhythms. Lecture Notes in Computer Science, 2130:1239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2130/21301239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2130/21301239.
                                        pdf.

Fukumi:2001:KIR

[692] Minoru Fukumi, Yasue Mitsukura, and Norio Akamatsu. Knowledge incorporation and rule extraction in neural networks. Lecture Notes in Computer Science, 2130:1248–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2130/21301248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2130/21301248.
pdf.

Anonymous:2001:AIfs

[693] Anonymous. Author index. Lecture Notes in Computer Science, 2130:1255– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2130/2130auth.pdf.

Darema:2001:SMP

[694] Frederica Darema. The SPMD model: Past, present and future. Lecture Notes in Computer Science, 2131:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310001.
pdf.

Geist:2001:BFN

[695] G. Al Geist. Building a foundation for the next PVM: Petascale Virtual Machines. Lecture Notes in Computer Science, 2131:2– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310002.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310002.
                                        pdf.

Gropp:2001:CSA

[696] William D. Gropp. Challenges and successes in achieving the potential of MPI. Lecture Notes in Computer Science, 2131:7–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310007.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310007.
                                        pdf.

Laforenza:2001:PHP

[697] Domenico Laforenza. Programming high performance applications in Grid environments. Lecture Notes in Computer Science, 2131:8– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310008.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310008.
                                        pdf.

Papadopoulos:2001:NRC

[698] Philip M. Papadopoulos, Mason J. Katz, and Greg Bruno. NPACI rocks clusters: Tools for easily deploying and maintaining manageable highperformance Linux clusters. Lecture Notes in Computer Science, 2131:10– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2131/21310010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310010.
pdf.

Reinefeld:2001:CDI

[699] Alexander Reinefeld. Clusters for dataintensive applications in the Grid. Lecture Notes in Computer Science, 2131: 12–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310012.
pdf.

Sunderam:2001:CAP

[700] Vaidy Sunderam and Zsolt Németh. A comparative analysis of PVM/MPI and computational Grids. Lecture Notes in Computer Science, 2131:14– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310014.
pdf.

Golbiewski:2001:MOS

[701] Maciej Gołbiewski and Jesper Larsson Träff. MPI-2 one-sided communications on a Giganet SMP cluster. Lecture Notes in Computer Science, 2131: 16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310016.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2131/21310016.
                                        pdf.

Rabenseifner:2001:ECF

[702] Rolf Rabenseifner and Alice E. Koniges. Effective communication and file-I/O bandwidth benchmarks. Lecture Notes in Computer Science, 2131: 24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310024.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310024.
                                        pdf.

Kucukboyaci:2001:PPT

[703] Vefa Kucukboyaci, Alireza Haghighat, and Glenn E. Sjoden. Performance of PENTRAN^{T M} 3-D parallel particle transport code on the IBM SP2 and PCTRAN cluster. Lecture Notes in Computer Science, 2131:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310036.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310036.
                                        pdf.

Huse:2001:LST

[704] Lars Paul Huse. Layering SHMEM on top of MPI. Lecture Notes in Computer Science, 2131:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310044.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2131/21310044.
pdf.

Tourancheau:2001:SMN

[705] Bernard Tourancheau and Roland Westrelin. Support for MPI at the network interface level. Lecture Notes in Computer Science, 2131:52– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310052.
pdf.

Baptista:2001:IOS

[706] Tiago Baptista, Hernani Pedroso, and João Gabriel Silva. The implementation of one-sided communications for WMPI II. Lecture Notes in Computer Science, 2131:61– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310061.
pdf.

Matuszek:2001:APS

[707] Mariusz R. Matuszek. Assessment of PVM suitability to testbed clientagent-server applications. Lecture Notes in Computer Science, 2131:69– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310069.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2131/21310069.
                                        pdf.

Chen:2001:TMK

[708] Yu Chen, Qian Fang, Zhihui Du, and Sanli Li. TH-MPI: OS kernel integrated fault tolerant MPI. Lecture Notes in Computer Science, 2131:75– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310075.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310075.
                                        pdf.

Gorzig:2001:CCP

[709] Steffen Görzig. CPPvm — C++ and PVM. Lecture Notes in Computer Science, 2131:83–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2131/21310083.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2131/21310083.pdf.

Manis:2001:PNP

[710] G. Manis. Persistent and nonpersistent data objects on top of PVM and MPI. Lecture Notes in Computer Science, 2131:91–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310091.
                                        pdf.

Fischer:2001:SAN

[711] Markus Fischer. System area network extensions to the parallel virtual machine. Lecture Notes in Computer Science, 2131:98–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310098.
pdf.

Gaito:2001:ADC

[712] A. Gaito, M. Rak, and U. Villano. Adding dynamic coscheduling support to PVM. Lecture Notes in Computer Science, 2131:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310106.
pdf.

Gonzalez:2001:MIM

[713] J. A. González, C. León, C. Rodríguez, and F. Sande. A model to integrate message passing and shared memory programming. Lecture Notes in Computer Science, 2131:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310114.
pdf.

Elwasif:2001:AMT

[714] Wael R. Elwasif, David E. Bernholdt, James A. Kohl, and G. A. Geist. An architecture for a multithreaded harness kernel. Lecture Notes in Computer Science, 2131:126– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310126.
pdf.

Fagg:2001:PIS

[715] Graham E. Fagg, Edgar Gabriel, Michael Resch, and Jack J. Dongarra. Parallel IO support for metacomputing applications: MPI Connect IO applied to PACX–MPI. Lecture Notes in Computer Science, 2131:135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310135.
pdf.

Konstantinou:2001:TTO

[716] Dimitris Konstantinou, Nectarios Koziris, and George Papakonstantinou. TOPPER: A tool for optimizing the performance of parallel applications. Lecture Notes in Computer Science, 2131:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310148.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2131/21310148.
pdf.

Macias:2001:PPA

[717] Elsa M. Macías, Alvaro Suárez, C. N. Ojeda-Guerra, and E. Robayna. Programming parallel applications with LAMGAC in a LAN–WLAN environment. Lecture Notes in Computer Science, 2131:158–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2131/21310158.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2131/21310158.pdf.

Pears:2001:DLB

[718] Arnold N. Pears and Nicola Thong. A dynamic load balancing architecture for PDES using PVM on clusters. Lecture Notes in Computer Science, 2131: 166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310166.
pdf.

Czarnul:2001:DPD

[719] Pawel Czarnul, Karen Tomko, and Henryk Krawczyk. Dynamic partitioning of the divide-and-conquer scheme with migration in PVM environment. Lecture Notes in Computer Science, 2131:174–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2131/21310174.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310174.
                                        pdf.

Wismuller:2001:UMT

[720] Roland Wismüller. Using monitoring techniques to support the cooperation of software components. Lecture Notes in Computer Science, 2131:183– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310183.
                                        pdf.

Kranzlmuller:2001:IRM

[721] Dieter Kranzlmüller, Christian Schaubschläger, and Jens Volkert. An integrated record&replay mechanism for nondeterministic message passing programs. Lecture Notes in Computer Science, 2131:192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2131/21310192.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2131/21310192.pdf.

Uthayopas:2001:FSR

[722] Putchong Uthayopas and Sugree Phatanapherom. Fast and scalable real-time monitoring system for Beowulf clusters. Lecture Notes in Computer Science, 2131:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2131/21310201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310201.
pdf.

Angskun:2001:DPM

[723] Thara Angskun, Putchong Uthayopas, and Arnon Rungsawang. Dynamic process management in KSIX cluster middleware. Lecture Notes in Computer Science, 2131:209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310209.
pdf.

Moreno:2001:AEP

[724] Luz Marina Moreno, Francisco Almeida, Daniel González, and Casiano Rodríguez. Adaptive execution of pipelines. Lecture Notes in Computer Science, 2131: 217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310217.
pdf.

Gine:2001:MMM

[725] Francesc Giné, Francesc Solsona, Xavi Navarro, Porfidio Hernández, and Emilio Luque. MemTo: A memory monitoring tool for a Linux cluster. Lecture Notes in Computer Science, 2131:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2131/21310225.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310225.
                                        pdf.

Ferschweiler:2001:CDP

[726] Ken Ferschweiler, Mariacarla Calzarossa, Cherri Pancake, Daniele Tessera, and Dylan Keon. A community databank for performance tracefiles. Lecture Notes in Computer Science, 2131:233– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310233.
                                        pdf.

Moore:2001:RPA

[727] Shirley Moore, David Cronk, Kevin London, and Jack Dongarra. Review of performance analysis tools for MPI parallel programs. Lecture Notes in Computer Science, 2131:241– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310241.
                                        pdf.

Pagourtzis:2001:PCT

[728] Aris Pagourtzis, Igor Potapov, and Wojciech Rytter. PVM computation of the transitive closure: The dependency graph approach. Lecture Notes in Computer Science, 2131:249– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310249.
pdf.

Luo:2001:PDE

[729] Jun Luo, Sanguthevar Rajasekaran, and Chenxia Qiu. Parallizing 1dimensional estuarine model. Lecture Notes in Computer Science, 2131:257– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310257.
pdf.

Schevtschenko:2001:PAS

[730] I. V. Schevtschenko. A parallel ADI and steepest descent methods. Lecture Notes in Computer Science, 2131:265– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310265.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310265.
pdf.

Fischer:2001:DNM

[731] Markus Fischer and Peter Kemper. Distributed numerical Markov chain analysis. Lecture Notes in Computer Science, 2131:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310272.
                                        pdf.

Bu:2001:PAC

[732] Libor Bu and Pavel Tvrdík. A parallel algorithm for connected components on distributed memory machines. Lecture Notes in Computer Science, 2131: 280–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310280.
                                        pdf.

Roussos:2001:BMB

[733] George Roussos and B. J. C. Baxter. Biharmonic many body calculations for fast evaluation of radial basis function interpolants in cluster environments. Lecture Notes in Computer Science, 2131:288–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310288.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310288.
                                        pdf.

Tinetti:2001:HNW

[734] Fernando Tinetti, Antonio Quijano, Armando De Giusti, and Emilio Luque. Heterogeneous networks of workstations and the parallel matrix multiplication. Lecture Notes in Computer Science, 2131:296–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310296.
pdf.

Baiardi:2001:CRD

[735] Fabrizio Baiardi, Paolo Mori, and Laura Ricci. Collecting remote data in irregular problems with hierarchical representation of the domain. Lecture Notes in Computer Science, 2131:304– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310304.
pdf.

Krawczyk:2001:PIM

[736] Henryk Krawczyk and Jamil Saif. Parallel image matching on PC cluster. Lecture Notes in Computer Science, 2131:312–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2131/21310312.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2131/21310312.
pdf.

Dehne:2001:CPD

[737] Frank Dehne, Todd Eavis, and Andrew Rau-Chaplin. Computing partial data cubes for parallel data warehousing applications. Lecture Notes in Computer Science, 2131:319–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310319.
                                        pdf.

Suppi:2001:PCS

[738] Remo Suppi, Fernando Cores, and Emilio Luque. PDES: A case study using the switch time warp. Lecture Notes in Computer Science, 2131:327– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310327.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310327.
                                        pdf.

Plunkett:2001:AMD

[739] Craig L. Plunkett, Alfred G. Striz, and J. Sobieszczanski-Sobieski. Application of MPI in displacement based multilevel structural optimization. Lecture Notes in Computer Science, 2131:335– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310335.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310335.
                                        pdf.

Wu:2001:PCS

[740] Guang Jun Wu and Robert Roy. Parallelization of characteristics solvers for 3D neutron transport. Lecture Notes in Computer Science, 2131:344– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310344.
pdf.

Haddad:2001:UNW

[741] Mohammed Al Haddad and Jerome Robinson. Using a network of workstations to enhance database query processing performance. Lecture Notes in Computer Science, 2131:352–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310352.
pdf.

Pringle:2001:TPF

[742] Gavin J. Pringle, Steven P. Booth, Hugh M. P. Couchman, Frazer R. Pearce, and Alan D. Simpson. Towards a portable, fast parallel AP^{3}M-SPH code: HYDRA MPI. Lecture Notes in Computer Science, 2131:360– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310360.
pdf.

Roig:2001:EMM

[743] Concepció Roig, Ana Ripoll, Javier Borrás, and Emilio Luque. Efficient mapping for message-passing applications using the TTIG model: A case study in image processing. Lecture Notes in Computer Science, 2131:370– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310370.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310370.
pdf.

Michailidis:2001:TSH

[744] Panagiotis D. Michailidis and Konstantinos G. Margaritis. Text searching on a heterogeneous cluster of workstations. Lecture Notes in Computer Science, 2131:378–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310378.
pdf.

Jorba:2001:SFF

[745] Josep Jorba, Tomàs Margalef, and Emilio Luque. Simulation of forest fire propagation on parallel & distributed PVM platforms. Lecture Notes in Computer Science, 2131:386– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310386.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310386.
pdf.

Nicolescu:2001:DTP

[746] Cristina Nicolescu and Pieter Jonker. A data and task parallel image processing environment. Lecture Notes in Computer Science, 2131:393–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310393.
pdf.

Gallud:2001:EDF

[747] J. A. Gallud, J. García-Consuegra, J. M. García, and L. Orozco. Evaluating the DIPORSI framework: Distributed processing of remotely sensed imagery. Lecture Notes in Computer Science, 2131:401–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310401.
pdf.

Ong:2001:SUC

[748] Emil Ong, Ewing Lusk, and William Gropp. Scalable Unix commands for parallel processors: A highperformance implementation. Lecture Notes in Computer Science, 2131:410– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310410.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2131/21310410.
                                        pdf.

Rungsawang:2001:LCP

[749] A. Rungsawang, A. Laohakanniyom, and M. Lertprasertkune. Low-cost parallel text retrieval using PC-cluster. Lecture Notes in Computer Science, 2131:419–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2131/21310419.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310419.
                                        pdf.

Okulicka-Dluzewska:2001:PFE

[750] Felicja Okulicka-Dłuzewska. Parallelization of finite element package by MPI library. Lecture Notes in Computer Science, 2131:427–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2131/21310427.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2131/21310427.
                                        pdf.

Anonymous:2001:AIft

[751] Anonymous. Author index. Lecture Notes in Computer Science, 2131:437– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2131/2131auth.pdf.

Zhu:2001:RCF

[752] Hong Zhu. The role of castes in formal specification of MAS. Lecture Notes in Computer Science, 2132:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2132/21320001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2132/21320001.
pdf.

deVries:2001:TCM

[753] Wieke de Vries, Frank S. de Boer, Wiebe van der Hoek, and John-Jules Ch. Meyer. A truly concurrent model for interacting agents. Lecture Notes in Computer Science, 2132:16– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2132/21320016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2132/21320016.
pdf.

Kinny:2001:RAC

[754] David Kinny. Reliable agent computation: An algebraic approach. Lecture Notes in Computer Science, 2132:31– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2132/21320031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2132/21320031.
pdf.

Toda:2001:ABA

[755] Yoshirou Toda, Masashi Yamashita, and Hajime Sawamura. An argumentbased agent system with KQML as an agent communication language. Lecture Notes in Computer Science, 2132: 48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2132/21320048.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2132/21320048.
                                        pdf.

Nawa:2001:FAO

[756] Norberto Eiji Nawa, Katsunori Shimohara, and Osamu Katai. On fairness in an alternating-offers bargaining model with evolutionary agents. Lecture Notes in Computer Science, 2132: 63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2132/21320063.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2132/21320063.
                                        pdf.

Matsui:2001:SBM

[757] Tomomi Matsui and Takahiro Watanabe. Sealed bid multi-object auctions with necessary bundles and its application to spectrum auctions. Lecture Notes in Computer Science, 2132:78– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2132/21320078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2132/21320078.
                                        pdf.

Lin:2001:SMP

[758] Aizhong (Alan) Lin. Strategic multipersonal-agent interaction. Lecture Notes in Computer Science, 2132:93– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2132/21320093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2132/21320093.
pdf.

Hsu:2001:MPA

[759] Wei-Tek Hsu and Von-Wun Soo. Market performance of adaptive trading agents in synchronous double auctions. Lecture Notes in Computer Science, 2132:108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2132/21320108.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2132/21320108.
pdf.

Lin:2001:NCD

[760] Yu-Hsin Lin and Soe-Tsyr Yuan. Negotiation-credit driven coalition formation in E-markets. Lecture Notes in Computer Science, 2132:122–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2132/21320122.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2132/21320122.
pdf.

Liu:2001:AAS

[761] Rey-Long Liu. An adaptive agent society for environmental scanning through the Internet. Lecture Notes in Computer Science, 2132:139–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2132/21320139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2132/21320139.
                                        pdf.

Liu:2001:GGA

[762] Jyi-Shane Liu, Von-Won Soo, Chia-Ning Chiang, Chen-Yu Lee, and Chun-Yu Lin. Gaz-Guide: Agentmediated information retrieval for official gazettes. Lecture Notes in Computer Science, 2132:154–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2132/21320154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2132/21320154.
                                        pdf.

Lee:2001:TIH

[763] Hao Lee and Hiroshi Deguchi. Technological innovation of high-tech industry and patent policy — agent based simulation with double loop learning. Lecture Notes in Computer Science, 2132: 168–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2132/21320168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2132/21320168.
                                        pdf.

Murakami:2001:CFD

[764] Eiji Murakami and Takao Terano. Collaborative filtering for a distributed smart IC card system. Lecture Notes in Computer Science, 2132:183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2132/21320183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2132/21320183.
pdf.

Funakoshi:2001:ICB

[765] Kaname Funakoshi, Koji Kamei, Sen Yoshida, and Kazuhiro Kuwabara. Incorporating content-based collaborative filtering in a community support system. Lecture Notes in Computer Science, 2132:198–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2132/21320198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2132/21320198.
pdf.

Kim:2001:PAB

[766] In-Cheol Kim. A personal agent for bookmark classification. Lecture Notes in Computer Science, 2132:210– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2132/21320210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2132/21320210.
pdf.

Yoo:2001:SWS

[767] Jeong-Joon Yoo, Doheon Lee, Young-Ho Suh, and Dong-Ik Lee. Scalable workflow system model based on mobile agents. Lecture Notes in Computer Science, 2132:222–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2132/21320222.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2132/21320222.
                                        pdf.

Anonymous:2001:AIfu

[768] Anonymous. Author index. Lecture Notes in Computer Science, 2132:237– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2132/2132ind.pdf.

Anonymous:2001:KAS

[769] Anonymous. Keynote address: Security protocols and the Swiss Army knife (transcript of discussion). Lecture Notes in Computer Science, 2133: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2133/21330001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330001.
                                        pdf.

Gollmann:2001:MP

[770] Dieter Gollmann. Mergers and principals. Lecture Notes in Computer Science, 2133:5–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2133/21330005.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2133/21330005.pdf.

Anonymous:2001:MPT

[771] Anonymous. Mergers and principals (transcript of discussion). Lecture Notes in Computer Science, 2133:14– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330014.
pdf.

Anonymous:2001:ANT

[772] Anonymous. Authentication and naming (transcript of discussion). Lecture Notes in Computer Science, 2133:20– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330020.
pdf.

Nikander:2001:UTC

[773] Pekka Nikander and Kristiina Karvonen. Users and trust in cyberspace. Lecture Notes in Computer Science, 2133:24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2133/21330024.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2133/21330024.
pdf.

Anonymous:2001:UTC

[774] Anonymous. Users and trust in cyberspace (transcript of discussion). Lecture Notes in Computer Science, 2133:36–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2133/21330036.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330036.
                                        pdf.

Anonymous:2001:IIP

[775] Anonymous. Interactive identification protocols (transcript of discussion). Lecture Notes in Computer Science, 2133:43–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2133/21330043.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330043.
                                        pdf.

Anonymous:2001:OQT

[776] Anonymous. Open questions (transcript of discussion). Lecture Notes in Computer Science, 2133:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2133/21330049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330049.
                                        pdf.

Anonymous:2001:LBS

[777] Anonymous. Looking on the bright side of black-box cryptography (transcript of discussion). Lecture Notes in Computer Science, 2133:54–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330054.
pdf.

Roe:2001:GAK

[778] Michael Roe, Ross Anderson, Bill Harbison, and Mark Lomas. Government access to keys — panel discussion. Lecture Notes in Computer Science, 2133: 62–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330062.
pdf.

Bella:2001:MSS

[779] Giampaolo Bella, Fabio Massacci, Lawrence C. Paulson, and Piero Tramontano. Making sense of specifications: The formalization of SET. Lecture Notes in Computer Science, 2133: 74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330074.
pdf.

Anonymous:2001:MSS

[780] Anonymous. Making sense of specifications: The formalization of SET (transcript of discussion). Lecture Notes in Computer Science, 2133:82– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2133/21330082.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330082.
                                        pdf.

Bella:2001:LES

[781] Giampaolo Bella. Lack of explicitness strikes back. Lecture Notes in Computer Science, 2133:87–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2133/21330087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330087.
                                        pdf.

Anonymous:2001:LES

[782] Anonymous. Lack of explicitness strikes back (transcript of discussion). Lecture Notes in Computer Science, 2133:94–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2133/21330094.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330094.
                                        pdf.

Khurana:2001:RRA

[783] Himanshu Khurana and Virgil D. Gligor. Review and revocation of access privileges distributed with PKI certificates. Lecture Notes in Computer Science, 2133:100–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330100.
pdf.

Anonymous:2001:RRA

[784] Anonymous. Review and revocation of access privileges distributed with PKI certificates (transcript of discussion). Lecture Notes in Computer Science, 2133:113–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2133/21330113.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2133/21330113.
pdf.

Anderson:2001:CCT

[785] Ross Anderson. The correctness of crypto transaction sets. Lecture Notes in Computer Science, 2133:125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330125.
pdf.

Anonymous:2001:CCT

[786] Anonymous. The correctness of crypto transaction sets (discussion). Lecture Notes in Computer Science, 2133:128– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2133/21330128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330128.
                                        pdf.

Anonymous:2001:MMR

[787] Anonymous. Micro-management of risk in a trust-based billing system (transcript of discussion). Lecture Notes in Computer Science, 2133:142– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2133/21330142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330142.
                                        pdf.

Price:2001:BSF

[788] Geraint Price. Broadening the scope of fault tolerance within secure services. Lecture Notes in Computer Science, 2133:155–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2133/21330155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330155.
                                        pdf.

Anonymous:2001:BSF

[789] Anonymous. Broadening the scope of fault tolerance within secure services (transcript of discussion). Lecture Notes in Computer Science, 2133: 165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2133/21330165.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2133/21330165.
pdf.

Aura:2001:RAC

[790] Tuomas Aura, Pekka Nikander, and Jussipekka Leiwo. DOS-resistant authentication with client puzzles. Lecture Notes in Computer Science, 2133: 170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330170.
pdf.

Anonymous:2001:RAC

[791] Anonymous. DOS-resistant authentication with client puzzles (transcript of discussion). Lecture Notes in Computer Science, 2133:178–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330178.
pdf.

Christianson:2001:PKC

[792] Bruce Christianson, Bruno Crispo, and James A. Malcolm. Publickey crypto-systems using symmetrickey crypto-algorithms. Lecture Notes in Computer Science, 2133:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330182.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330182.
                                        pdf.

Anonymous:2001:PKC

[793] Anonymous. Public-key cryptosystems using symmetric-key cryptoalgorithms (transcript of discussion). Lecture Notes in Computer Science, 2133:184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2133/21330184.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330184.
                                        pdf.

Gligor:2001:DSP

[794] Virgil Gligor, Matt Blaze, and John Ioannidis. Denial of service — panel discussion. Lecture Notes in Computer Science, 2133:194–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2133/21330194.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2133/21330194.
                                        pdf.

Stajano:2001:RDW

[795] Frank Stajano. The resurrecting duckling — what next? Lecture Notes in Computer Science, 2133:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2133/21330204.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2133/21330204.
pdf.

Anonymous:2001:RDW

[796] Anonymous. The resurrecting duckling — what next? (transcript of discussion). Lecture Notes in Computer Science, 2133:215–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2133/21330215.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2133/21330215.pdf.

Anonymous:2001:AAPb

[797] Anonymous. An anonymous auction protocol using “money escrow” (transcript of discussion). Lecture Notes in Computer Science, 2133:223– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330223.
pdf.

Anonymous:2001:SCS

[798] Anonymous. Short certification of secure RSA modulus (transcript of discussion). Lecture Notes in Computer Science, 2133:234–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2133/21330234.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2133/21330234.pdf.

Yoshiura:2001:AWB

[799] Hiroshi Yoshiura, Takaaki Shigematsu, Seiichi Susaki, Tsukasa Saitoh, Hisashi Toyoshima, Chikako Kurita, Satoru Tezuka, and Ryoichi Sasaki. Authenticating Web-based virtual shops using signature-embedded marks — A practical analysis. Lecture Notes in Computer Science, 2133:238–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330238.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330238.
pdf.

Anonymous:2001:AWB

[800] Anonymous. Authentication Webbased virtual shops using signatureembedded marks — A practical analysis — (transcript of discussion). Lecture Notes in Computer Science, 2133: 249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330249.
pdf.

Anonymous:2001:CTL

[801] Anonymous. I cannot tell a lie (transcript of discussion). Lecture Notes in Computer Science, 2133:253–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330253.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2133/21330253.
pdf.

Anonymous:2001:A

[802] Anonymous. Afterward. Lecture Notes in Computer Science, 2133:256– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2133/21330256.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2133/21330256.
pdf.

Anonymous:2001:AIfv

[803] Anonymous. Author index. Lecture Notes in Computer Science, 2133:257– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2133/2133auth.pdf.

Yuille:2001:DLA

[804] Alan Yuille. A double-loop algorithm to minimize the Bethe free energy. Lecture Notes in Computer Science, 2134: 3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340003.
pdf.

Hamza:2001:VAM

[805] A. Ben Hamza and Hamid Krim. A variational approach to maximum a posteriori estimation for image denoising. Lecture Notes in Computer Science, 2134:19–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340019.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340019.
pdf.

Aguiar:2001:MLE

[806] Pedro M. Q. Aguiar and José M. F. Moura. Maximum likelihood estimation of the template of a rigid moving object. Lecture Notes in Computer Science, 2134:34–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340034.
pdf.

Trouve:2001:MSL

[807] Alain Trouvé and Yong Yu. Metric similarities learning through examples: An application to shape retrieval. Lecture Notes in Computer Science, 2134:50– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340050.
pdf.

Sanches:2001:FMA

[808] João M. Sanches and Jorge S. Marques. A fast MAP algorithm for 3D ultrasound. Lecture Notes in Computer Science, 2134:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340063.
pdf.

Bicego:2001:DMS

[809] Manuele Bicego, Agostino Dovier, and Vittorio Murino. Designing the minimal structure of hidden Markov model by bisimulation. Lecture Notes in Computer Science, 2134:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340075.
pdf.

Fusiello:2001:RSM

[810] Andrea Fusiello, Umberto Castellani, and Vittorio Murino. Relaxing symmetric multiple windows stereo using Markov random fields. Lecture Notes in Computer Science, 2134:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340091.
pdf.

Morris:2001:MIM

[811] Robin D. Morris, Vadim N. Smelyanskiy, and Peter C. Cheeseman. Matching images to models — camera calibration for 3-D surface reconstruction. Lecture Notes in Computer Science, 2134:105–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340105.
                                        pdf.

Yu:2001:HMR

[812] Stella X. Yu, Tai Sing Lee, and Takeo Kanade. A hierarchical Markov random field model for figureground segregation. Lecture Notes in Computer Science, 2134:118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340118.
                                        pdf.

Rocha:2001:AOT

[813] Jairo Rocha and Arnau Mir. Articulated object tracking via a genetic algorithm. Lecture Notes in Computer Science, 2134:134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340134.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340134.
                                        pdf.

Rangarajan:2001:LMS

[814] Anand Rangarajan. Learning matrix space image representations. Lecture Notes in Computer Science, 2134:153– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340153.
pdf.

Gimelfarb:2001:STS

[815] Georgy Gimel’farb. Supervised texture segmentation by maximising conditional likelihood. Lecture Notes in Computer Science, 2134:169–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340169.
pdf.

Lebanon:2001:DMP

[816] Guy Lebanon and Alfred M. Bruckstein. Designing moiré patterns. Lecture Notes in Computer Science, 2134: 185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340185.
pdf.

Sziranyi:2001:OPR

[817] Tamás Szirányi and Zoltán Tóth. Optimization of paintbrush rendering of images by dynamic MCMC methods. Lecture Notes in Computer Science, 2134:

201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340201.
                                        pdf.

Al-Rawi:2001:IIR

[818] Mohammed Al-Rawi and Yang Jie. Illumination invariant recognition of color texture using correlation and covariance functions. Lecture Notes in Computer Science, 2134:216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340216.
                                        pdf.

Fischer:2001:PBP

[819] Bernd Fischer, Thomas Zöller, and Joachim M. Buhmann. Path based pairwise data clustering with application to texture segmentation. Lecture Notes in Computer Science, 2134:235– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340235.
                                        pdf.

Robles-Kelly:2001:MLF

[820] Antonio Robles-Kelly and Edwin R. Hancock. A maximum likelihood framework for grouping and segmentation. Lecture Notes in Computer Science, 2134:251–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2134/21340251.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2134/21340251.pdf.

Schellewald:2001:ILG

[821] Christian Schellewald, Jens Keuchel, and Christoph Schnörr. Image labeling and grouping by minimizing linear functionals over cones. Lecture Notes in Computer Science, 2134:267– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340267.
pdf.

Yu:2001:GDR

[822] Stella X. Yu and Jianbo Shi. Grouping with directed relationships. Lecture Notes in Computer Science, 2134: 283–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340283.
pdf.

Kamijo:2001:SST

[823] Shunsuke Kamijo, Katsushi Ikeuchi, and Masao Sakauchi. Segmentations of spatio-temporal images by spatio-temporal Markov random field model. Lecture Notes in Computer Science, 2134:298–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340298.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340298.
                                        pdf.

Fieguth:2001:HSI

[824] Paul Fieguth and Slawo Wesolkowski. Highlight and shading invariant color image segmentation using simulated annealing. Lecture Notes in Computer Science, 2134:314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340314.
                                        pdf.

Kubota:2001:EBP

[825] Toshiro Kubota, Terry Huntsberger, and Jeffrey T. Martin. Edge based probabilistic relaxation for sub-pixel contour extraction. Lecture Notes in Computer Science, 2134:328–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340328.
                                        pdf.

Samson:2001:TVM

[826] Christophe Samson, Laure Blanc-Féraud, Gilles Aubert, and Josiane Zerubia. Two variational models for multispectral image classification. Lecture Notes in Computer Science, 2134: 344–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340344.
pdf.

Boykov:2001:ECM

[827] Yuri Boykov and Vladimir Kolmogorov. An experimental comparison of min-cut/max-flow algorithms for energy minimization in vision. Lecture Notes in Computer Science, 2134:359– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340359.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340359.
pdf.

Dias:2001:DCM

[828] José M. B. Dias and José M. N. Leitão. A discrete/continuous minimization method in interferometric image processing. Lecture Notes in Computer Science, 2134:375–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340375.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340375.
pdf.

Toh:2001:GEM

[829] Kar-Ann Toh. Global energy minimization: A transformation approach. Lecture Notes in Computer Science, 2134: 391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340391.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340391.
                                        pdf.

Toh:2001:GFN

[830] Kar-Ann Toh, Juwei Lu, and Wei-Yun Yau. Global feedforward neural network learning for classification and regression. Lecture Notes in Computer Science, 2134:407–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340407.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340407.
                                        pdf.

Pelillo:2001:MFT

[831] Marcello Pelillo. Matching free trees, maximal cliques, and monotone game dynamics. Lecture Notes in Computer Science, 2134:423–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340423.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340423.
                                        pdf.

Torsello:2001:ECW

[832] Andrea Torsello and Edwin R. Hancock. Efficiently computing weighted tree edit distance using relaxation labeling. Lecture Notes in Computer Science, 2134:438–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340438.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340438.
pdf.

Bengoetxea:2001:EDA

[833] Endika Bengoetxea, Pedro Larrañaga, Isabelle Bloch, and Aymeric Perchant. Estimation of distribution algorithms: A new evolutionary computation approach for graph matching problems. Lecture Notes in Computer Science, 2134:454–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340454.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340454.
pdf.

Massaro:2001:CPA

[834] Alessio Massaro and Marcello Pelillo. A complementary pivoting approach to graph matching. Lecture Notes in Computer Science, 2134:469–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340469.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2134/21340469.
                                        pdf.

Woo:2001:AGA

[835] Sanghyuk Woo, Albert Dipanda, and Franck Marzani. Application of genetic algorithms to 3-D shape reconstruction in an active stereo vision system. Lecture Notes in Computer Science, 2134: 480–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340480.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340480.
                                        pdf.

August:2001:MPU

[836] Jonas August and Steven W. Zucker. A Markov process using curvature for filtering curve images. Lecture Notes in Computer Science, 2134:497– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340497.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340497.
                                        pdf.

Camion:2001:GIS

[837] Vincent Camion and Laurent Younes. Geodesic interpolating splines. Lecture Notes in Computer Science, 2134:513– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340513.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2134/21340513.
pdf.

Hirani:2001:ATM

[838] Anil N. Hirani, Jerrold E. Marsden, and James Arvo. Averaged template matching equations. Lecture Notes in Computer Science, 2134:528– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340528.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340528.
pdf.

Pao:2001:CSD

[839] Hsing-Kuo Pao and Davi Geiger. A continuous shape descriptor by orientation diffusion. Lecture Notes in Computer Science, 2134:544–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340544.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340544.
pdf.

Cohen:2001:MCF

[840] Laurent D. Cohen and Thomas Deschamps. Multiple contour finding and perceptual grouping as a set of energy minimizing paths. Lecture Notes in Computer Science, 2134:560– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340560.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2134/21340560.
                                        pdf.

Abrantes:2001:STU

[841] Arnaldo J. Abrantes and Jorge S. Marques. Shape tracking using centroid-based methods. Lecture Notes in Computer Science, 2134:576–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340576.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340576.
                                        pdf.

Lefebure:2001:OFI

[842] Martin Lefébure and Laurent D. Cohen. Optical flow and image registration: A new local rigidity approach for global minimization. Lecture Notes in Computer Science, 2134: 592–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340592.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2134/21340592.
                                        pdf.

Matula:2001:SOR

[843] Pavel Matula and David Svoboda. Spherical object reconstruction using star-shaped simplex meshes. Lecture Notes in Computer Science, 2134:608– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2134/21340608.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2134/21340608.
pdf.

Sagiv:2001:GFS

[844] Chen Sagiv, Nir A. Sochen, and Yehoshua Y. Zeevi. Gabor feature space diffusion via the minimal weighted area method. Lecture Notes in Computer Science, 2134:621– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340621.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340621.
pdf.

Siddiqi:2001:FMF

[845] Kaleem Siddiqi and Alexander Vasilevskiy. 3D flux maximizing flows. Lecture Notes in Computer Science, 2134:636– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2134/21340636.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2134/21340636.
pdf.

Anonymous:2001:AIfw

[846] Anonymous. Author index. Lecture Notes in Computer Science, 2134:651– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2134/2134auth.pdf.

Kirby:2001:SO

[847] Graham N. C. Kirby. Session

1:  Overview.  Lecture Notes

in Computer Science, 2135:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350001.
pdf.

Whitlock:2001:FPE

[848] David Whitlock and Antony L. Hosking. A framework for persistenceenabled optimization of Java object stores. Lecture Notes in Computer Science, 2135:4–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2135/21350004.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2135/21350004.pdf.

Lewis:2001:APH

[849] Brian Lewis, Bernd Mathiske, and Neal Gafter. Architecture of the PEVM: A high-performance orthogonally Persistent Java^{T M} Virtual Machine. Lecture Notes in Computer Science, 2135:18– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350018.
pdf.

Blackburn:2001:SO

[850] Stephen M. Blackburn. Session 2: Overview. Lecture Notes in Computer Science, 2135:34–35, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350034.
pdf.

Voges:2001:SMB

[851] Erik Voges and Sonia Berman. A spatiotemporal model as the basis for a persistent GIS. Lecture Notes in Computer Science, 2135:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350036.
pdf.

Shapiro:2001:EPL

[852] Marc Shapiro, Paulo Ferreira, and Nicolas Richer. Experience with the PerDiS large-scale data-sharing middleware. Lecture Notes in Computer Science, 2135:55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2135/21350055.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2135/21350055.pdf.

Kaplan:2001:TPP

[853] Alan Kaplan, John V. E. Ridgway, Bradley R. Schmerl, Krishnan Sridhar, and Jack C. Wileden. Toward pure polylingual persistence. Lecture Notes in Computer Science, 2135:70– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350070.
                                        pdf.

Jones:2001:SO

[854] Richard Jones. Session 3: Overview. Lecture Notes in Computer Science, 2135:84–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2135/21350084.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350084.
                                        pdf.

Bjornsson:2001:TRG

[855] Magnus E. Bjornsson and Liuba Shrira. Transactional remote group caching in distributed object systems. Lecture Notes in Computer Science, 2135:87– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350087.
                                        pdf.

He:2001:PDI

[856] Zhen He, Stephen M. Blackburn, Luke Kirby, and John Zigman. Platypus: Design and implementation of a flexible high performance object store. Lecture Notes in Computer Science, 2135: 100–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350100.
pdf.

Munro:2001:EPS

[857] David S. Munro and Alfred L. Brown. Evaluating partition selection policies using the PMOS garbage collector. Lecture Notes in Computer Science, 2135: 125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350125.
pdf.

Zigman:2001:TTG

[858] John Zigman, Stephen M. Blackburn, and J. Eliot B. Moss. TMOS: A transactional garbage collector. Lecture Notes in Computer Science, 2135:138– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350138.
pdf.

Hosking:2001:SO

[859] Antony L. Hosking. Session 4: Overview. Lecture Notes in Computer Science, 2135:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2135/21350157.

                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2135/21350157.pdf.

Richer:2001:MBW

[860] Nicolas Richer and Marc Shapiro. The memory behavior of the WWW, or the WWW considered as a persistent store. Lecture Notes in Computer Science, 2135:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350161.
                                        pdf.

Garratt:2001:CTP

[861] Andrea Garratt, Mike Jackson, Peter Burden, and Jon Wallis. A comparison of two persistent storage tools for implementing a search engine. Lecture Notes in Computer Science, 2135:177– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350177.
                                        pdf.

Shrira:2001:SO

[862] Liuba Shrira. Session 5: Overview. Lecture Notes in Computer Science, 2135: 187–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350187.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2135/21350187.
pdf.

Bem:2001:AIP

[863] Ewa Z. Bem and John Rosenberg. An approach to implementing persistent computations. Lecture Notes in Computer Science, 2135:189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350189.
pdf.

Skoglund:2001:TOC

[864] Espen Skoglund, Christian Ceelen, and Jochen Liedtke. Transparent orthogonal checkpointing through user-level pagers. Lecture Notes in Computer Science, 2135:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350201.
pdf.

Dini:2001:OUD

[865] Gianluca Dini, Giuseppe Lettieri, and Lanfranco Lopriore. An overview of ulisse, a distributed single address space system. Lecture Notes in Computer Science, 2135:215–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350215.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350215.
                                        pdf.

Dearle:2001:SO

[866] Alan Dearle. Session 6: Overview. Lecture Notes in Computer Science, 2135: 228–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350228.
                                        pdf.

Zirintsis:2001:HCR

[867] E. Zirintsis, Graham N. C. Kirby, and Ron Morrison. Hyper-code revisited: Unifying program source, executable, and data. Lecture Notes in Computer Science, 2135:232–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350232.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350232.
                                        pdf.

Marquez:2001:IOP

[868] Alonso Marquez, Stephen M. Blackburn, Gavin Mercer, and John Zigman. Implementing orthogonally persistent Java. Lecture Notes in Computer Science, 2135:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350247.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2135/21350247.
pdf.

Berman:2001:SO

[869] Sonia Berman. Session 7: Overview. Lecture Notes in Computer Science, 2135:262–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2135/21350262.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2135/21350262.
pdf.

Bacon:2001:ESF

[870] Jean Bacon, Alexis Hombrecher, Chaoying Ma, Ken Moody, and Walt Yao. Event storage and federation using ODMG. Lecture Notes in Computer Science, 2135:265–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350265.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350265.
pdf.

Gruber:2001:AME

[871] Olivier Gruber and Ravi Konuru. An application model and environment for personal information appliances. Lecture Notes in Computer Science, 2135: 282–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2135/21350282.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2135/21350282.
pdf.

Atkinson:2001:SRI

[872] M. P. Atkinson, M. Dmitriev, C. Hamilton, and T. Printezis. Scalable and recoverable implementation of object evolution for the PJama 1 platform. Lecture Notes in Computer Science, 2135:292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350292.
                                        pdf.

Morrison:2001:E

[873] Ron Morrison. Epilogue. Lecture Notes in Computer Science, 2135:315– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2135/21350315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2135/21350315.
                                        pdf.

Anonymous:2001:AIfx

[874] Anonymous. Author index. Lecture Notes in Computer Science, 2135:321– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2135/2135auth.pdf.

Scott:2001:NCS

[875] Dana S. Scott. A new category for semantics. Lecture Notes in Computer Science, 2136:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360001.
pdf.

Burgisser:2001:IBP

[876] Peter Bürgisser. On implications between P-NP-Hypotheses: Decision versus computation in algebraic complexity. Lecture Notes in Computer Science, 2136:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360003.
pdf.

Demaine:2001:PGA

[877] Erik D. Demaine. Playing games with algorithms: Algorithmic combinatorial game theory. Lecture Notes in Computer Science, 2136:18– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360018.
pdf.

Fiat:2001:SRR

[878] Amos Fiat. Some recent results on data mining and search. Lecture Notes in Computer Science, 2136:33– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360033.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360033.
                                        pdf.

Gottlob:2001:HDS

[879] Georg Gottlob, Nicola Leone, and Francesco Scarcello. Hypertree decompositions: A survey. Lecture Notes in Computer Science, 2136:37– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360037.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360037.
                                        pdf.

Hofmann:2001:SNS

[880] Martin Hofmann. The strength of non-size-increasing computation (introduction and summary). Lecture Notes in Computer Science, 2136:58– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360058.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360058.
                                        pdf.

Hoyer:2001:IRQ

[881] Peter Høyer. Introduction to recent quantum algorithms. Lecture Notes in Computer Science, 2136:62– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2136/21360062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360062.
pdf.

Randall:2001:DMS

[882] Dana Randall. Decomposition methods and sampling circuits in the Cartesian lattice. Lecture Notes in Computer Science, 2136:74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360074.
pdf.

Schoning:2001:NAK

[883] Uwe Schöning. New algorithms for k-SAT based on the local search principle. Lecture Notes in Computer Science, 2136:87–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360087.
pdf.

Wilke:2001:LTL

[884] Thomas Wilke. Linear temporal logic and finite semigroups. Lecture Notes in Computer Science, 2136:96– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360096.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2136/21360096.
                                        pdf.

Alber:2001:RST

[885] Jochen Alber, Hongbing Fan, Michael R. Fellows, Henning Fernau, Rolf Niedermeier, Fran Rosamond, and Ulrike Stege. Refined search tree technique for DOMINATING SET on planar graphs. Lecture Notes in Computer Science, 2136:111–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360111.
                                        pdf.

Amano:2001:CPF

[886] Kazuyuki Amano, Tsukuru Hirosawa, Yusuke Watanabe, and Akira Maruoka. The computational power of a family of decision forests. Lecture Notes in Computer Science, 2136:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360123.
                                        pdf.

Ambainis:2001:ERA

[887] Andris Ambainis and Arnolds Kikusts. Exact results for accepting probabilities of quantum automata. Lecture Notes in Computer Science, 2136:135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2136/21360135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360135.
pdf.

Atserias:2001:IBW

[888] Albert Atserias. Improved bounds on the weak pigeonhole principle and infinitely many primes from weaker axioms. Lecture Notes in Computer Science, 2136:148–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360148.
pdf.

Barrett:2001:APS

[889] Chris Barrett, Harry B. Hunt III, Madhav V. Marathe, S. S. Ravi, Daniel J. Rosenkrantz, and Richard E. Stearns. Analysis problems for sequential dynamical systems and communicating state machines. Lecture Notes in Computer Science, 2136:159– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360159.
pdf.

Beaudry:2001:CTC

[890] Martin Beaudry and Markus Holzer. The complexity of tensor circuit evaluation. Lecture Notes in Computer Science, 2136:173–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2136/21360173.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2136/21360173.pdf.

Blaser:2001:CRB

[891] Markus Bläser. Computing reciprocals of bivariate power series. Lecture Notes in Computer Science, 2136:186– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360186.
                                        pdf.

Bouajjani:2001:AVR

[892] Ahmed Bouajjani, Peter Habermehl, and Richard Mayr. Automatic verification of recursive procedures with one integer parameter. Lecture Notes in Computer Science, 2136:198–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360198.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360198.
                                        pdf.

Brosenne:2001:GDF

[893] Henrik Brosenne, Matthias Homeister, and Stephan Waack. Graphdriven free parity BDDs: Algorithms and lower bounds. Lecture Notes in Computer Science, 2136:212–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360212.
pdf.

Brattka:2001:CVB

[894] Vasco Brattka. Computable versions of Baire’s category theorem. Lecture Notes in Computer Science, 2136:224– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360224.
pdf.

Bruyere:2001:ALO

[895] Véronique Bruyère and Olivier Carton. Automata on linear orderings. Lecture Notes in Computer Science, 2136:236– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360236.
pdf.

Cervelle:2001:AIT

[896] Julien Cervelle, Bruno Durand, and Enrico Formenti. Algorithmic information theory and cellular automata dynamics. Lecture Notes in Computer Science, 2136:248–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360248.
                                        pdf.

Chrobak:2001:KMP

[897] Marek Chrobak, Lawrence L. Larmore, and Wojciech Rytter. The k-median problem for directed trees. Lecture Notes in Computer Science, 2136:260– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360260.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360260.
                                        pdf.

Cryan:2001:PGN

[898] Mary Cryan and Peter Bro Miltersen. On pseudorandom generators in NC. Lecture Notes in Computer Science, 2136:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360272.
                                        pdf.

Cucker:2001:TNS

[899] Felipe Cucker and Dima Grigoriev. There are no sparse NP _{W} -hard sets. Lecture Notes in Computer Science, 2136:285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2136/21360285.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2136/21360285.
pdf.

Crescenzo:2001:SOS

[900] Giovanni Di Crescenzo. Sharing one secret vs. sharing many secrets: Tight bounds for the max improvement ratio. Lecture Notes in Computer Science, 2136:292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360292.
pdf.

Diaz:2001:CFE

[901] Josep Díaz, Maria Serna, and Dimitrios M. Thilikos. (H, C, K)-coloring: Fast, easy, and hard cases. Lecture Notes in Computer Science, 2136:304– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360304.
pdf.

Downey:2001:RR

[902] Rod G. Downey, Denis R. Hirschfeldt, and Geoff LaForte. Randomness and reducibility. Lecture Notes in Computer Science, 2136:316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360316.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360316.
                                        pdf.

Duris:2001:CCI

[903] Pavol Ďuriš and Ján Maňuch. On the computational complexity of infinite words. Lecture Notes in Computer Science, 2136:328–337, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360328.
                                        pdf.

Epstein:2001:LBL

[904] Leah Epstein and Rob van Stee. Lower bounds for on-line singlemachine scheduling. Lecture Notes in Computer Science, 2136:338–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360338.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360338.
                                        pdf.

Erlebach:2001:AAC

[905] Thomas Erlebach. Approximation algorithms and complexity results for path problems in trees of rings. Lecture Notes in Computer Science, 2136: 351–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360351.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2136/21360351.
pdf.

Espelage:2001:AAM

[906] Wolfgang Espelage and Egon Wanke. A 3-approximation algorithm for movement minimization in conveyor flow shop processing. Lecture Notes in Computer Science, 2136:363–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360363.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360363.
pdf.

Fournier:2001:QRP

[907] Hervé Fournier. Quantifier rank for parity of embedded finite models. Lecture Notes in Computer Science, 2136: 375–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360375.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360375.
pdf.

Geffert:2001:SHT

[908] Viliam Geffert. Space hierarchy theorem revised. Lecture Notes in Computer Science, 2136:387–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360387.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360387.
pdf.

Geffert:2001:CTW

[909] Viliam Geffert, Carlo Mereghetti, and Giovanni Pighizzini. Converting twoway nondeterministic unary automata into simpler automata. Lecture Notes in Computer Science, 2136:398–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360398.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360398.
                                        pdf.

Hoang:2001:CMP

[910] Thanh Minh Hoang and Thomas Thierauf. The complexity of the minimal polynomial. Lecture Notes in Computer Science, 2136:408–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360408.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360408.
                                        pdf.

Jiraskova:2001:NMF

[911] Galina Jirásková. Note on minimal finite automata. Lecture Notes in Computer Science, 2136:421–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360421.
                                        pdf.

Kari:2001:SFA

[912] Jarkko Kari. Synchronizing finite automata on Eulerian digraphs. Lecture Notes in Computer Science, 2136:432– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360432.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360432.
pdf.

Klein:2001:THB

[913] Andreas Klein and Martin Kutrib. A time hierarchy for bounded oneway cellular automata. Lecture Notes in Computer Science, 2136:439–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360439.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360439.
pdf.

Klin:2001:CAC

[914] Bartek Klin, Piotr Hoffman, Andrzej Tarlecki, Lutz Schröder, and Till Mossakowski. Checking amalgamability conditions for C ASL architectural specifications. Lecture Notes in Computer Science, 2136:451–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360451.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360451.
pdf.

Koo:2001:LST

[915] Chiu-Yuen Koo, Tak-Wah Lam, Tsuen-Wan Ngan, and Kar-Keung To. On-line scheduling with tight deadlines. Lecture Notes in Computer Science, 2136: 464–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360464.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360464.
                                        pdf.

Kral:2001:CNM

[916] Daniel Král’, Jan Kratochvíl, and Heinz-Jürgen Voss. Complexity note on mixed hypergraphs. Lecture Notes in Computer Science, 2136:474–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360474.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360474.
                                        pdf.

Krumke:2001:NOT

[917] Sven O. Krumke, Willem E. de Paepe, Diana Poensgen, and Leen Stougie. News from the online traveling repairman. Lecture Notes in Computer Science, 2136:487–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2136/21360487.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2136/21360487.pdf.

Lohrey:2001:WPH

[918] Markus Lohrey. Word problems for 2-homogeneous monoids and symmetric logspace. Lecture Notes in Computer Science, 2136:500–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360500.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360500.
pdf.

Mignosi:2001:VTF

[919] Filippo Mignosi, Jeffrey Shallit, and Ming wei Wang. Variations on a theorem of fine & wilf. Lecture Notes in Computer Science, 2136:512– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360512.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360512.
pdf.

Monien:2001:UBB

[920] Burkhard Monien and Robert Preis. Upper bounds on the bisection width of 3- and 4-regular graphs. Lecture Notes in Computer Science, 2136:524– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360524.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360524.
pdf.

Moore:2001:SSE

[921] Cristopher Moore, Pascal Tesson, and Denis Thérien. Satisfiability of systems of equations over finite monoids. Lecture Notes in Computer Science, 2136: 537–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360537.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360537.
                                        pdf.

Morvan:2001:RGT

[922] Christophe Morvan and Colin Stirling. Rational graphs trace contextsensitive languages. Lecture Notes in Computer Science, 2136:548–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360548.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360548.
                                        pdf.

Neven:2001:TRL

[923] Frank Neven, Thomas Schwentick, and Victor Vianu. Towards regular languages over infinite alphabets. Lecture Notes in Computer Science, 2136:560– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360560.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360560.
                                        pdf.

Nickelsen:2001:PIS

[924] Arfst Nickelsen. Partial information and special case algorithms. Lecture Notes in Computer Science, 2136:573– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360573.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360573.
pdf.

Ogihara:2001:CCN

[925] Mitsunori Ogihara and Seinosuke Toda. The complexity of computing the number of self-avoiding walks in two-dimensional grid graphs and in hypercube graphs. Lecture Notes in Computer Science, 2136:585–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2136/21360585.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2136/21360585.
pdf.

Piterman:2001:BA

[926] Nir Piterman and Moshe Y. Vardi. From bidirectionality to alternation. Lecture Notes in Computer Science, 2136:598–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2136/21360598.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2136/21360598.
pdf.

Polak:2001:SSL

[927] Libor Polák. Syntactic semiring of a language. Lecture Notes in Computer Science, 2136:611–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360611.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360611.
                                        pdf.

Pudlak:2001:RSD

[928] Pavel Pudlák. On reducibility and symmetry of disjoint NP-pairs. Lecture Notes in Computer Science, 2136:621– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360621.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360621.
                                        pdf.

Rettinger:2001:HMC

[929] Robert Rettinger and Xizhong Zheng. Hierarchy of monotonically computable real numbers. Lecture Notes in Computer Science, 2136:633–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360633.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360633.
                                        pdf.

Santocanale:2001:EDL

[930] Luigi Santocanale. On the equational definition of the least prefixed point.

Lecture Notes in Computer Science, 2136:645–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2136/21360645.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2136/21360645.
pdf.

Shur:2001:PPW

[931] Arseny M. Shur and Yulia V. Konovalova. On the periods of partial words. Lecture Notes in Computer Science, 2136:657–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2136/21360657.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2136/21360657.
pdf.

Sutner:2001:SPA

[932] Klaus Sutner. The size of power automata. Lecture Notes in Computer Science, 2136:666–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2136/21360666.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2136/21360666.pdf.

Thimm:2001:AST

[933] Martin Thimm. On the approximability of the Steiner tree problem. Lecture Notes in Computer Science, 2136: 678–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2136/21360678.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360678.
                                        pdf.

Wang:2001:ABT

[934] Zhuozhi Wang and Kaizhong Zhang. Alignment between two RNA structures. Lecture Notes in Computer Science, 2136:690–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2136/21360690.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2136/21360690.pdf.

Wich:2001:CCF

[935] Klaus Wich. Characterization of context-free languages with polynomially bounded ambiguity. Lecture Notes in Computer Science, 2136:703– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2136/21360703.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2136/21360703.
                                        pdf.

Anonymous:2001:AIfy

[936] Anonymous. Author index. Lecture Notes in Computer Science, 2136:715– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2136/2136ind.pdf.

Bern:2001:TPD

[937] Marshall Bern, Jeff Breidenbach, and David Goldberg. Trustworthy paper documents. Lecture Notes in Computer Science, 2137:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370001.
pdf.

Sharp:2001:IKB

[938] Toby Sharp. An implementation of key-based digital signal steganography. Lecture Notes in Computer Science, 2137:13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2137/21370013.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2137/21370013.
pdf.

Goljan:2001:DFD

[939] Miroslav Goljan, Jessica J. Fridrich, and Rui Du. Distortion-free data embedding for images. Lecture Notes in Computer Science, 2137:27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370027.
pdf.

Korjik:2001:IHT

[940] Valeri Korjik and Guillermo Morales-Luna. Information hiding through noisy channels. Lecture Notes in Computer Science, 2137:42–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370042.
pdf.

Mihcak:2001:PAH

[941] M. Kıvanç Mıhçak and Ramarathnam Venkatesan. A perceptual audio hashing algorithm: A tool for robust audio identification and information hiding. Lecture Notes in Computer Science, 2137:51–65, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370051.
pdf.

Wong:2001:CFT

[942] Jennifer L. Wong, Darko Kirovski, and Miodrag Potkonjak. Computational forensic techniques for intellectual property protection. Lecture Notes in Computer Science, 2137:66– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370066.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2137/21370066.
pdf.

Koushanfar:2001:IPM

[943] Farinaz Koushanfar, Gang Qu, and Miodrag Potkonjak. Intellectual property metering. Lecture Notes in Computer Science, 2137:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370081.
pdf.

Qu:2001:KPW

[944] Gang Qu. Keyless public watermarking for intellectual property authentication. Lecture Notes in Computer Science, 2137:96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370096.
pdf.

Berthold:2001:EIP

[945] Oliver Berthold, Sebastian Clauß, Stefan Köpsell, and Andreas Pfitzmann. Efficiency improvements of the private message service. Lecture Notes in Computer Science, 2137:112–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370112.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2137/21370112.
                                        pdf.

Dingledine:2001:RSI

[946] Roger Dingledine, Michael J. Freedman, David Hopwood, and David Molnar. A reputation system to increase MIX-net reliability. Lecture Notes in Computer Science, 2137:126– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2137/21370126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370126.
                                        pdf.

Picard:2001:NNF

[947] Justin Picard and Arnaud Robert. Neural networks functions for public key watermarking. Lecture Notes in Computer Science, 2137:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2137/21370142.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370142.
                                        pdf.

Venkatesan:2001:GTA

[948] Ramarathnam Venkatesan, Vijay Vazirani, and Saurabh Sinha. A graph theoretic approach to software watermarking. Lecture Notes in Computer Science, 2137:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2137/21370157.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2137/21370157.
pdf.

Thanos:2001:CVM

[949] Dimitris Thanos. COiN-Video: A model for the dissemination of copyrighted video streams over open networks. Lecture Notes in Computer Science, 2137:169–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370169.
pdf.

Atallah:2001:NLW

[950] Mikhail J. Atallah, Victor Raskin, Michael Crogan, Christian Hempelmann, Florian Kerschbaum, Dina Mohamed, and Sanket Naik. Natural language watermarking: Design, analysis, and a proof-of-concept implementation. Lecture Notes in Computer Science, 2137:185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370185.
pdf.

Eggers:2001:DWC

[951] Joachim J. Eggers, Wolf-Dietrich Ihlenfeldt, and Bernd Girod. Digital watermarking of chemical structure sets. Lecture Notes in Computer Science, 2137:200–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2137/21370200.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370200.
                                        pdf.

Benes:2001:SES

[952] Tonda Beneš. The strong eternity service. Lecture Notes in Computer Science, 2137:215–229, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2137/21370215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370215.
                                        pdf.

Clayton:2001:RWP

[953] Richard Clayton, George Danezis, and Markus G. Kuhn. Real world patterns of failure in anonymity systems. Lecture Notes in Computer Science, 2137: 230–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2137/21370230.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370230.
                                        pdf.

Back:2001:TAA

[954] Adam Back, Ulf Möller, and Anton Stiglic. Traffic analysis attacks and trade-offs in anonymity providing systems. Lecture Notes in Computer Science, 2137:245–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370245.
pdf.

Stern:2001:ADW

[955] Julien P. Stern and Jean-Pierre Tillich. Automatic detection of a watermarked document using a private key. Lecture Notes in Computer Science, 2137:258– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370258.
pdf.

Adelsbach:2001:ZKW

[956] André Adelsbach and Ahmad-Reza Sadeghi. Zero-knowledge watermark detection and proof of ownership. Lecture Notes in Computer Science, 2137: 273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370273.
pdf.

Westfeld:2001:FSA

[957] Andreas Westfeld. F5 — A steganographic algorithm. Lecture Notes in Computer Science, 2137:289–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2137/21370289.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370289.
                                        pdf.

Muratani:2001:CSF

[958] Hirofumi Muratani. A collusionsecure fingerprinting code reduced by Chinese remaindering and its random-error resilience. Lecture Notes in Computer Science, 2137:303–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2137/21370303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370303.
                                        pdf.

Sugihara:2001:PCD

[959] Ryo Sugihara. Practical capacity of digital watermarks. Lecture Notes in Computer Science, 2137:316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2137/21370316.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370316.
                                        pdf.

Voloshynovskiy:2001:BDA

[960] Sviatoslav Voloshynovskiy, Alexander Herrigel, and Thierry Pun. Blur/deblur attack against document protection systems based on digital watermarking. Lecture Notes in Computer Science, 2137:330–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2137/21370330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370330.
pdf.

Pereira:2001:SGB

[961] Shelby Pereira, Sviatoslav Voloshynovskiy, Maribel Madueno, Stéphan Marchand-Maillet, and Thierry Pun. Second generation benchmarking and application oriented evaluation. Lecture Notes in Computer Science, 2137:340– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370340.
pdf.

Kirovski:2001:RCC

[962] Darko Kirovski and Henrique Malvar. Robust covert communication over a public audio channel using spread spectrum. Lecture Notes in Computer Science, 2137:354–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2137/21370354.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2137/21370354.
pdf.

Dasseni:2001:HAR

[963] Elena Dasseni, Vassilios S. Verykios, Ahmed K. Elmagarmid, and Elisa Bertino. Hiding association rules by using confidence and support. Lecture Notes in Computer Science, 2137:369– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2137/21370369.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370369.
                                        pdf.

Sadeghi:2001:HBS

[964] Ahmad-Reza Sadeghi. How to break a semi-anonymous fingerprinting scheme. Lecture Notes in Computer Science, 2137:384–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2137/21370384.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370384.
                                        pdf.

Boeuf:2001:AOS

[965] Julien Boeuf and Julien P. Stern. An analysis of one of the SDMI candidates. Lecture Notes in Computer Science, 2137:395–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2137/21370395.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2137/21370395.
                                        pdf.

Anonymous:2001:AIfz

[966] Anonymous. Author index. Lecture Notes in Computer Science, 2137:411– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/2137/2137ind.pdf.

Barzdins:2001:TAB

[967] Jānis Bārzdiņš, Rūsiņš Freivalds, and Carl H. Smith. Towards axiomatic basis of inductive inference. Lecture Notes in Computer Science, 2138:1– 13, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380001.
pdf.

Jansen:2001:AAF

[968] Klaus Jansen. Approximation algorithms for fractional covering and packing problems, and applications. Lecture Notes in Computer Science, 2138:14– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380014.
pdf.

Karhumaki:2001:CC

[969] Juhani Karhumäki. Challenges of commutation. Lecture Notes in Computer Science, 2138:15–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380015.
pdf.

Karpinski:2001:ABD

[970] Marek Karpinski. Approximating bounded degree instances of NP-hard problems. Lecture Notes in Computer Science, 2138:24–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380024.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380024.
                                        pdf.

Plotkin:2001:UAC

[971] Boris Plotkin and Tanya Plotkin. Universal algebra and computer science. Lecture Notes in Computer Science, 2138:35–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2138/21380035.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380035.
                                        pdf.

Vazirani:2001:QA

[972] Umesh Vazirani. Quantum algorithms. Lecture Notes in Computer Science, 2138:45–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2138/21380045.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380045.
                                        pdf.

Ablayev:2001:DAC

[973] Farid Ablayev and Svetlana Ablayeva. A discrete approximation and communication complexity approach to the superposition problem. Lecture Notes in Computer Science, 2138:47– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380047.
pdf.

Ablayev:2001:CPQ

[974] Farid Ablayev, Aida Gainutdinova, and Marek Karpinski. On computational power of quantum branching programs. Lecture Notes in Computer Science, 2138:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2138/21380059.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2138/21380059.
pdf.

Baier:2001:ECS

[975] Harald Baier. Efficient computation of singular moduli with application in cryptography. Lecture Notes in Computer Science, 2138:71– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380071.
pdf.

Berzina:2001:AFA

[976] Aija Bērziņa and Richard Bonner. Ambainis-Freivalds’ algorithm for measure-once automata. Lecture Notes in Computer Science, 2138:83– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380083.
                                        pdf.

Cirulis:2001:TEI

[977] Jānis Cīrulis. Are there essentially incomplete knowledge representation systems? Lecture Notes in Computer Science, 2138:94–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380094.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380094.
                                        pdf.

Ciura:2001:BIA

[978] Marcin Ciura. Best increments for the average case of shellsort. Lecture Notes in Computer Science, 2138:106– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380106.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380106.
                                        pdf.

Fomin:2001:AMC

[979] Fedor V. Fomin, Dieter Kratsch, and Jean-Christophe Novelli. Approximating minimum cocolourings. Lecture Notes in Computer Science, 2138:118– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380118.
pdf.

Freivalds:2001:CER

[980] Kārlis Freivalds. Curved edge routing. Lecture Notes in Computer Science, 2138:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2138/21380126.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2138/21380126.
pdf.

Gasieniec:2001:TSE

[981] Leszek Gasieniec and Igor Potapov. Time/space efficient compressed pattern matching. Lecture Notes in Computer Science, 2138:138–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380138.
pdf.

Heinemann:2001:MCA

[982] Bernhard Heinemann. Modelling change with the aid of knowledge and time. Lecture Notes in Computer Science, 2138:150–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2138/21380150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380150.
                                        pdf.

Hemaspaandra:2001:ITS

[983] Lane A. Hemaspaandra, Kari Pasanen, and Jörg Rothe. If P ≠ NP then some strongly noninvertible functions are invertible. Lecture Notes in Computer Science, 2138:162–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380162.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380162.
                                        pdf.

Hirata:2001:PPR

[984] Kouichi Hirata and Hiroshi Sakamoto. Prediction-preserving reducibility with membership queries on formal languages. Lecture Notes in Computer Science, 2138:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2138/21380172.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2138/21380172.pdf.

Jarvinen:2001:DFK

[985] Jouni Järvinen. Dense families and key functions of database relation instances. Lecture Notes in Computer Science, 2138:184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2138/21380184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380184.
pdf.

Karhumaki:2001:CDC

[986] Juhani Karhumäki, Wojciech Plandowski, and Wojciech Rytter. On the complexity of decidable cases of commutation problem for languages. Lecture Notes in Computer Science, 2138:193– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380193.
pdf.

Kuich:2001:CSA

[987] Werner Kuich. Cones, Semi-AFPs, and AFPs of algebraic power series. Lecture Notes in Computer Science, 2138: 204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380204.
pdf.

Kudlek:2001:NSU

[988] Manfred Kudlek and Yurii Rogozhin. New small universal circular post machines. Lecture Notes in Computer Science, 2138:217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380217.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380217.
                                        pdf.

Kuske:2001:DMP

[989] Dietrich Kuske. Divisibility monoids: Presentation, word problem, and rational languages. Lecture Notes in Computer Science, 2138:227–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380227.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380227.
                                        pdf.

Lanotte:2001:CTA

[990] Ruggero Lanotte, Andrea Maggiolo-Schettini, and Simone Tini. Concurrency in timed automata. Lecture Notes in Computer Science, 2138:240– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380240.
                                        pdf.

Lafitte:2001:HPI

[991] Grégory Lafitte. How powerful are infinite time machines? Lecture Notes in Computer Science, 2138:252– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380252.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2138/21380252.
pdf.

Linde:2001:EPC

[992] Ģirts Linde. Equivalence problem of composite class diagrams. Lecture Notes in Computer Science, 2138:264– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380264.
pdf.

Monnot:2001:DAR

[993] Jérôme Monnot, Vangelis Th. Paschos, and Sophie Toulouse. Differential approximation results for the traveling salesman problem with distances 1 and 2. Lecture Notes in Computer Science, 2138:275–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380275.
pdf.

Moskaljova:2001:CES

[994] Nataly S. Moskaljova and Irina B. Virbitskaite. On the category of event structures with dense time. Lecture Notes in Computer Science, 2138:287– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380287.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2138/21380287.
                                        pdf.

Nickelsen:2001:CPT

[995] Arfst Nickelsen and Till Tantau. Closure of polynomial time partial information classes under polynomial time reductions. Lecture Notes in Computer Science, 2138:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380299.
                                        pdf.

Rettinger:2001:MCP

[996] Robert Rettinger and Rutger Verbeek. Monte-Carlo polynomial versus linear time — the truth-table case. Lecture Notes in Computer Science, 2138:311– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380311.
                                        pdf.

Selivanov:2001:RAT

[997] V. L. Selivanov. Relating automatatheoretic hierarchies to complexitytheoretic hierarchies. Lecture Notes in Computer Science, 2138:323–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380323.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2138/21380323.
pdf.

Shoudai:2001:PTA

[998] Takayoshi Shoudai, Tomoyuki Uchida, and Tetsuhiro Miyahara. Polynomial time algorithms for finding unordered tree patterns with internal variables. Lecture Notes in Computer Science, 2138:335–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380335.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380335.
pdf.

Trahtman:2001:PLT

[999] A. N. Trahtman. Piecewise and local threshold testability of DFA. Lecture Notes in Computer Science, 2138:347– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380347.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380347.
pdf.

Walicki:2001:CHR

[1000] Michał Walicki, Adis Hodzic, and Sigurd Meldal. Compositional homomorphisms of relational structures. Lecture Notes in Computer Science, 2138: 359–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380359.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2138/21380359.
                                        pdf.

Buls:2001:RAA

[1001] Jānis Buls, Vaira Buža, and Roberts Glaudiņš. Representation of autonomous automata. Lecture Notes in Computer Science, 2138:372–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380372.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380372.
                                        pdf.

Ciamarra:2001:QRN

[1002] Massimo Pica Ciamarra. Quantum reversibility and a new model of quantum automaton. Lecture Notes in Computer Science, 2138:376–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380376.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380376.
                                        pdf.

Dubrovsky:2001:SEW

[1003] Andrej Dubrovsky. Space-efficient 1.5way quantum Turing machine. Lecture Notes in Computer Science, 2138:380– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380380.
                                        pdf.

Gambin:2001:CAA

[1004] Anna Gambin and Piotr Pokarowski.

A combinatorial aggregation algorithm for stationary distribution of a large Markov chain. Lecture Notes in Computer Science, 2138:384–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380384.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380384.
pdf.

Kiltz:2001:PPS

[1005] Eike Kiltz. A primitive for proving the security of every bit and about universal hash functions & hard core bits. Lecture Notes in Computer Science, 2138:388–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380388.
pdf.

Lipyanski:2001:PTU

[1006] Ruvim Lipyanski. Pythagorean triples in unification theory of nilpotent rings. Lecture Notes in Computer Science, 2138:392–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2138/21380392.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2138/21380392.
pdf.

Ollinger:2001:TSB

[1007] Nicolas Ollinger. Two-states bilinear intrinsically universal cellular automata. Lecture Notes in Computer Science, 2138:396–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380396.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380396.
                                        pdf.

Papazian:2001:LTR

[1008] Christophe Papazian and Eric Rémila. Linear time recognizer for subsets of Z^{2}. Lecture Notes in Computer Science, 2138:400–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2138/21380400.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2138/21380400.pdf.

Plotkin:2001:FSA

[1009] Tanya Plotkin. Fuzzy sets and algorithms of distributed task allocation for cooperative agents. Lecture Notes in Computer Science, 2138:404– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380404.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380404.
                                        pdf.

Rozenblat:2001:RES

[1010] Bella V. Rozenblat. On recursively enumerable subsets of N and Rees matrix semigroups over (Z_{3}; +). Lecture Notes in Computer Science, 2138:408– 411, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380408.
pdf.

Scegulnaja:2001:QRT

[1011] Oksana Scegulnaja. Quantum realtime Turing machine. Lecture Notes in Computer Science, 2138:412–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380412.
pdf.

Sokolov:2001:MMO

[1012] Andrew V. Sokolov. Mathematical models and optimal algorithms of dynamic data structure control. Lecture Notes in Computer Science, 2138:416– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380416.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380416.
pdf.

Sokratova:2001:LAR

[1013] Olga Sokratova. Linear automata and recognizable subsets in free semirings. Lecture Notes in Computer Science, 2138:420–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2138/21380420.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380420.
                                        pdf.

Tombak:2001:LMC

[1014] Mati Tombak, Ain Isotamm, and Tõnu Tamme. On logical method for counting Dedekind numbers. Lecture Notes in Computer Science, 2138:424– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380424.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380424.
                                        pdf.

Valiente:2001:GMG

[1015] Gabriel Valiente. A general method for graph isomorphism. Lecture Notes in Computer Science, 2138:428–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380428.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380428.
                                        pdf.

Afrati:2001:DPM

[1016] F. Afrati and I. Milis. Designing PTASs for MIN–SUM scheduling problems.

Lecture Notes in Computer Science, 2138:432–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2138/21380432.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2138/21380432.
pdf.

Brandstadt:2001:RAM

[1017] Andreas Brandstädt. On robust algorithms for the maximum weight stable set problem. Lecture Notes in Computer Science, 2138:445–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380445.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380445.
pdf.

Gargano:2001:MON

[1018] Luisa Gargano. Multicasting in optical networks. Lecture Notes in Computer Science, 2138:459–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380459.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380459.
pdf.

Doerr:2001:SRR

[1019] Benjamin Doerr. Structured randomized rounding and coloring. Lecture Notes in Computer Science, 2138:461– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380461.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380461.
                                        pdf.

Epstein:2001:OOF

[1020] Leah Epstein and Rob van Stee. Optimal online flow time with resource augmentation. Lecture Notes in Computer Science, 2138:472–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380472.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380472.
                                        pdf.

Erlebach:2001:NRP

[1021] Thomas Erlebach and Danica Vukadinović. New results for path problems in generalized stars, complete graphs, and brick wall graphs. Lecture Notes in Computer Science, 2138: 483–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380483.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380483.
                                        pdf.

Fishkin:2001:MAWb

[1022] Aleksei V. Fishkin, Klaus Jansen, and Lorant Porkolab. On minimizing average weighted completion time: A PTAS for scheduling general multiprocessor tasks. Lecture Notes in Computer Science, 2138:495– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380495.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380495.
pdf.

Fomin:2001:AAT

[1023] Fedor V. Fomin and Andrzej Lingas. Approximation algorithms for time-dependent orienteering. Lecture Notes in Computer Science, 2138:508– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380508.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380508.
pdf.

Kral:2001:CCM

[1024] Daniel Král’. On complexity of colouring mixed hypertrees. Lecture Notes in Computer Science, 2138:516–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2138/21380516.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2138/21380516.
pdf.

Mastrolilli:2001:CAG

[1025] Monaldo Mastrolilli. Combining arithmetic and geometric rounding techniques for knapsack problems. Lecture Notes in Computer Science, 2138:525– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380525.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380525.
                                        pdf.

Mielikainen:2001:CMM

[1026] Taneli Mielikäinen and Esko Ukkonen. The complexity of maximum matroid-greedoid intersection. Lecture Notes in Computer Science, 2138:535– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2138/21380535.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2138/21380535.
                                        pdf.

Anonymous:2001:AIga

[1027] Anonymous. Author index. Lecture Notes in Computer Science, 2138:541– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2138/2138auth.pdf.

Barak:2001:IPO

[1028] Boaz Barak, Oded Goldreich, Rusell Impagliazzo, Steven Rudich, Amit Sahai, Salil Vadhan, and Ke Yang. On the (im)possibility of obfuscating programs. Lecture Notes in Computer Science, 2139:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2139/21390001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2139/21390001.
pdf.

Canetti:2001:UCC

[1029] Ran Canetti and Marc Fischlin. Universally composable commitments. Lecture Notes in Computer Science, 2139:19–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2139/21390019.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2139/21390019.
pdf.

Naor:2001:RTS

[1030] Dalit Naor, Moni Naor, and Jeff Lotspiech. Revocation and tracing schemes for stateless receivers. Lecture Notes in Computer Science, 2139:41– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390041.
pdf.

Kiayias:2001:SPP

[1031] Aggelos Kiayias and Moti Yung.

Self protecting pirates and blackbox traitor tracing. Lecture Notes in Computer Science, 2139:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390063.
pdf.

Fitzi:2001:MCP

[1032] Matthias Fitzi, Juan A. Garay, Ueli Maurer, and Rafail Ostrovsky. Minimal complete primitives for secure multi-party computation. Lecture Notes in Computer Science, 2139:80– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2139/21390080.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2139/21390080.
                                        pdf.

Hirt:2001:RFU

[1033] Martin Hirt and Ueli Maurer. Robustness for free in unconditional multiparty computation. Lecture Notes in Computer Science, 2139:101–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2139/21390101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2139/21390101.
                                        pdf.

Cramer:2001:SDL

[1034] Ronald Cramer and Ivan Damgård. Secure distributed linear algebra in a constant number of rounds. Lecture Notes in Computer Science, 2139:119– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2139/21390119.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2139/21390119.
                                        pdf.

MacKenzie:2001:TPG

[1035] Philip MacKenzie and Michael K.

Reiter. Two-party generation of DSA signatures. Lecture Notes in Computer Science, 2139:137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390137.
pdf.

Ding:2001:OTB

[1036] Yan Zong Ding. Oblivious transfer in the bounded storage model. Lecture Notes in Computer Science, 2139:155– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390155.
pdf.

Lindell:2001:PCT

[1037] Yehuda Lindell. Parallel cointossing and constant-round secure twoparty computation. Lecture Notes in Computer Science, 2139:171–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390171.
pdf.

Gallant:2001:FPM

[1038] Robert P. Gallant, Robert J. Lambert, and Scott A. Vanstone. Faster point multiplication on elliptic curves with efficient endomorphisms. Lecture Notes in Computer Science, 2139:190– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390190.
pdf.

Boneh:2001:UBE

[1039] Dan Boneh and Igor E. Shparlinski. On the unpredictability of bits of the elliptic curve Diffie–Hellman scheme. Lecture Notes in Computer Science, 2139: 201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390201.
pdf.

Boneh:2001:IBE

[1040] Dan Boneh and Matt Franklin. Identity-based encryption from the Weil pairing. Lecture Notes in Computer Science, 2139:213–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2139/21390213.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2139/21390213.pdf.

Manger:2001:CCA

[1041] James Manger. A chosen ciphertext attack on RSA optimal asymmetric encryption padding (OAEP) as standardized in PKCS #1 v2.0. Lecture Notes in Computer Science, 2139:230– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390230.
pdf.

Shoup:2001:OR

[1042] Victor Shoup. OAEP reconsidered.

Lecture Notes in Computer Science, 2139:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2139/21390239.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2139/21390239.
pdf.

Fujisaki:2001:ROS

[1043] Eiichiro Fujisaki, Tatsuaki Okamoto, David Pointcheval, and Jacques Stern. RSA–OAEP is secure under the RSA assumption. Lecture Notes in Computer Science, 2139:260–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390260.
pdf.

Boneh:2001:SOR

[1044] Dan Boneh. Simplified OAEP for the RSA and Rabin functions. Lecture Notes in Computer Science, 2139:275– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390275.
pdf.

Bellare:2001:OCH

[1045] Mihir Bellare, Alexandra Boldyreva, Lars Knudsen, and Chanathip Namprempre. Online ciphers and the hash-CBC construction. Lecture Notes in Computer Science, 2139:292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390292.
pdf.

Krawczyk:2001:OEA

[1046] Hugo Krawczyk. The order of encryption and authentication for protecting communications (or: How secure is SSL?). Lecture Notes in Computer Science, 2139:310–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390310.
pdf.

Itkis:2001:FSS

[1047] Gene Itkis and Leonid Reyzin.

Forward-secure signatures with optimal signing and verifying. Lecture Notes in Computer Science, 2139:332– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390332.
pdf.

Shamir:2001:IOO

[1048] Adi Shamir and Yael Tauman.

Improved online/offline signature schemes. Lecture Notes in Computer Science, 2139:355–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2139/21390355.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2139/21390355.pdf.

Furukawa:2001:ESP

[1049] Jun Furukawa and Kazue Sako. An efficient scheme for proving a shuffle. Lecture Notes in Computer Science, 2139: 368–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390368.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390368.
pdf.

Camenisch:2001:IES

[1050] Jan Camenisch and Anna Lysyanskaya. An identity escrow scheme with appointed verifiers. Lecture Notes in Computer Science, 2139:388– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390388.
pdf.

Goldreich:2001:SKG

[1051] Oded Goldreich and Yehuda Lindell. Session-key generation using human passwords only. Lecture Notes in Computer Science, 2139:408–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390408.
pdf.

Brier:2001:CRS

[1052] Eric Brier, Christophe Clavier, Jean-Sébastien Coron, and David Naccache. Cryptanalysis of RSA signatures with fixed-pattern padding. Lecture Notes in Computer Science, 2139: 433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390433.
pdf.

Golic:2001:CAS

[1053] Jovan D. Golić. Correlation analysis of the shrinking generator. Lecture Notes in Computer Science, 2139:440– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390440.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390440.
pdf.

Cheon:2001:NVR

[1054] Jung Hee Cheon. Nonlinear vector resilient functions. Lecture Notes in Computer Science, 2139:458–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390458.
pdf.

Paeng:2001:NPK

[1055] Seong-Hun Paeng, Kil-Chan Ha, Jae Heon Kim, Seongtaek Chee, and Choonsik Park. New public key cryptosystem using finite non Abelian groups. Lecture Notes in Computer Science, 2139:470–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390470.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390470.
pdf.

Lee:2001:PBG

[1056] Eonkyung Lee, Sang Jin Lee, and Sang Geun Hahn. Pseudorandomness from braid groups. Lecture Notes in Computer Science, 2139:486– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2139/21390486.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2139/21390486.
                                        pdf.

Cramer:2001:CRS

[1057] Ronald Cramer, Ivan Damgård, and Serge Fehr. On the cost of reconstructing a secret, or VSS with optimal reconstruction phase. Lecture Notes in Computer Science, 2139:503– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2139/21390503.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2139/21390503.
                                        pdf.

Cachin:2001:SEA

[1058] Christian Cachin, Klaus Kursawe, Frank Petzold, and Victor Shoup. Secure and efficient asynchronous broadcast protocols. Lecture Notes in Computer Science, 2139:524–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2139/21390524.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2139/21390524.
                                        pdf.

Micali:2001:SPK

[1059] Silvio Micali and Leonid Reyzin. Soundness in the public-key model.

Lecture Notes in Computer Science, 2139:542–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2139/21390542.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2139/21390542.
pdf.

DeSantis:2001:RNI

[1060] Alfredo De Santis, Giovanni Di Crescenzo, Rafail Ostrovsky, Giuseppe Persiano, and Amit Sahai. Robust non-interactive zero knowledge. Lecture Notes in Computer Science, 2139: 566–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2139/21390566.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2139/21390566.
pdf.

Anonymous:2001:AIgb

[1061] Anonymous. Author index. Lecture Notes in Computer Science, 2139:599– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2139/2139auth.pdf.

Moore:2001:PCS

[1062] Simon Moore. Protecting consumer security devices. Lecture Notes in Computer Science, 2140:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2140/21400001.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400001.
                                        pdf.

Barthe:2001:JTR

[1063] G. Barthe, G. Dufay, M. Huisman, and S. Melo de Sousa. Jakarta: A toolset for reasoning about JavaCard. Lecture Notes in Computer Science, 2140: 2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400002.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400002.
                                        pdf.

Bella:2001:MPS

[1064] Giampaolo Bella. Mechanising a protocol for smart cards. Lecture Notes in Computer Science, 2140:19– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400019.
                                        pdf.

Campo:2001:JFC

[1065] Celeste Campo, Andrés Marm, Arturo García, Ignacio Díaz, Peter T. Breuer, Carlos Delgado, and Carlos García. JCCM: Flexible certificates for smartcards with Java card. Lecture Notes in Computer Science, 2140: 34–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400034.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2140/21400034.
pdf.

Caromel:2001:CIS

[1066] Denis Caromel, Ludovic Henrio, and Bernard Serpette. Context inference for static analysis of Java card object sharing. Lecture Notes in Computer Science, 2140:43–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2140/21400043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2140/21400043.
pdf.

Clarke:2001:ATO

[1067] Duncan Clarke, Thierry Jéron, Vlad Rusu, and Elena Zinovieva. Automated test and oracle generation for smart-card applications. Lecture Notes in Computer Science, 2140:58– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2140/21400058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2140/21400058.
pdf.

Deswarte:2001:IAS

[1068] Yves Deswarte, Noreddine Abghour, Vincent Nicomette, and David Powell. An Internet authorization scheme using smart-card-based security kernels. Lecture Notes in Computer Science, 2140: 71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400071.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400071.
                                        pdf.

Donsez:2001:TMA

[1069] Didier Donsez, Sébastien Jean, Sylvain Lecomte, and Olivier Thomas. Turning multi-applications smart cards services available from anywhere at anytime: A SOAP /MOM approach in the context of Java cards. Lecture Notes in Computer Science, 2140:83– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400083.
                                        pdf.

Eluard:2001:OSJ

[1070] Marc Éluard, Thomas Jensen, and Ewen Denne. An operational semantics of the Java card firewall. Lecture Notes in Computer Science, 2140:95– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400095.
                                        pdf.

Gore:2001:CMT

[1071] Rajeev Prabhakar Goré and Phuong Thê Nguyên. CardS4: Modal theorem proving on Java smartcards. Lecture Notes in Computer Science, 2140:111– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2140/21400111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2140/21400111.
pdf.

Henderson:2001:IEV

[1072] Neil J. Henderson, Neil M. White, and Pieter H. Hartel. iButton enrolment and verification requirements for the pressure sequence smartcard biometric. Lecture Notes in Computer Science, 2140:124–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2140/21400124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2140/21400124.
pdf.

Kehr:2001:STO

[1073] Roger Kehr and Hendrik Mieves. SIMspeak — towards an open and secure application platform for GSM SIMs. Lecture Notes in Computer Science, 2140: 135–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2140/21400135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2140/21400135.
pdf.

Leroy:2001:CBV

[1074] Xavier Leroy. On-card bytecode verification for Java card. Lecture Notes in Computer Science, 2140:150–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400150.
                                        pdf.

Meijer:2001:TFF

[1075] Hans Meijer and Erik Poll. Towards a full formal specification of the JavaCard API. Lecture Notes in Computer Science, 2140:165–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400165.
                                        pdf.

Meister:2001:PPG

[1076] Gisela Meister and Michael Vogel. Protection profiles and generic security targets for smart cards as secure signature creation devices — existing solutions for the payment sector. Lecture Notes in Computer Science, 2140:179– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400179.
                                        pdf.

Montgomery:2001:FIF

[1077] Michael Montgomery and Ksheerabdhi Krishna. A flexible invocation framework for Java card. Lecture Notes in Computer Science, 2140:188– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2140/21400188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2140/21400188.
pdf.

Quisquater:2001:EAE

[1078] Jean-Jacques Quisquater and David Samyde. ElectroMagnetic analysis (EMA): Measures and countermeasures for smart cards. Lecture Notes in Computer Science, 2140:200– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2140/21400200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2140/21400200.
pdf.

Romer:2001:ILA

[1079] Tanja Römer and Jean-Pierre Seifert.

Information leakage attacks against smart card implementations of the elliptic curve digital signature algorithm. Lecture Notes in Computer Science, 2140:211–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2140/21400211.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2140/21400211.
pdf.

Struif:2001:UBU

[1080] Bruno Struif. Use of biometrics for user verification in electronic signature smartcards. Lecture Notes in Computer Science, 2140:220–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400220.
                                        pdf.

Urien:2001:PIS

[1081] Pascal Urien. Programming Internet smartcard with XML scripts. Lecture Notes in Computer Science, 2140:228– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400228.
                                        pdf.

Yamamoto:2001:PKB

[1082] Hideo Yamamoto, Tetsutaro Kobayashi, Masahiro Morita, and Ryuji Yamada. Public-key-based high-speed payment (electronic money) system using contact-less smart cards. Lecture Notes in Computer Science, 2140:242– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2140/21400242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2140/21400242.
                                        pdf.

Anonymous:2001:AIgc

[1083] Anonymous. Author index. Lecture Notes in Computer Science, 2140:255– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2140/2140ind.pdf.

Navarro:2001:CDR

[1084] Gonzalo Navarro and Mathieu Raffinot. Compact DFA representation for fast regular expression search. Lecture Notes in Computer Science, 2141: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2141/21410001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2141/21410001.
pdf.

Iliopoulos:2001:MSA

[1085] Costas S. Iliopoulos, Laurent Mouchard, and Yoan J. Pinzon. The max-shift algorithm for approximate string matching. Lecture Notes in Computer Science, 2141:13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2141/21410013.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2141/21410013.pdf.

Bilardi:2001:FMM

[1086] Gianfranco Bilardi, Paolo D’Alberto, and Alex Nicolau. Fractal matrix multiplication: A case study on portability of cache performance. Lecture Notes in Computer Science, 2141:26– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2141/21410026.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2141/21410026.
                                        pdf.

Katajainen:2001:EDI

[1087] Jyrki Katajainen and Bjarke Buur Mortensen. Experiences with the design and implementation of spaceefficient deques. Lecture Notes in Computer Science, 2141:39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2141/21410039.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2141/21410039.
                                        pdf.

Bronnimann:2001:DIG

[1088] Hervé Brönnimann. Designing and implementing a general purpose halfedge data structure. Lecture Notes in Computer Science, 2141:51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2141/21410051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2141/21410051.
                                        pdf.

Rahman:2001:OPD

[1089] Naila Rahman, Richard Cole, and Rajeev Raman. Optimised predecessor data structures for internal memory. Lecture Notes in Computer Science, 2141:67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2141/21410067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2141/21410067.
pdf.

Hert:2001:AEG

[1090] Susan Hert, Michael Hoffmann, Lutz Kettner, Sylvain Pion, and Michael Seel. An adaptable and extensible geometry kernel. Lecture Notes in Computer Science, 2141:79–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2141/21410079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2141/21410079.
pdf.

Andersson:2001:ERA

[1091] Arne Andersson, Per Carlsson, and Fredrik Ygge. Efficient resource allocation with noisy functions. Lecture Notes in Computer Science, 2141:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2141/21410091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2141/21410091.
pdf.

Goldengorin:2001:IEB

[1092] Boris Goldengorin, Diptesh Ghosh, and Gerard Sierksma. Improving the efficiency of branch and bound algorithms for the simple plant location problem. Lecture Notes in Computer Science, 2141:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2141/21410106.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2141/21410106.
                                        pdf.

Iwama:2001:EPK

[1093] Kazuo Iwama and Suguru Tamaki. Exploiting partial knowledge of satisfying assignments. Lecture Notes in Computer Science, 2141:118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2141/21410118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2141/21410118.
                                        pdf.

Bader:2001:UPA

[1094] David A. Bader, Ajith K. Illendula, Bernard M. E. Moret, and Nina R. Weisse-Bernstein. Using PRAM algorithms on a uniform-memory-access shared-memory architecture. Lecture Notes in Computer Science, 2141:129– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2141/21410129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2141/21410129.
                                        pdf.

Anderson:2001:ESD

[1095] Eric Anderson, Joe Hall, Jason Hartline, Michael Hobbs, Anna R. Karlin, Jared Saia, Ram Swaminathan, and John Wilkes. An experimental study of data migration algorithms. Lecture Notes in Computer Science, 2141:145– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2141/21410145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2141/21410145.
pdf.

Chatzigiannakis:2001:ESB

[1096] Ioannis Chatzigiannakis, Sotiris Nikoletseas, Nearchos Paspallis, Paul Spirakis, and Christos Zaroliagis. An experimental study of basic communication protocols in ad-hoc mobile networks. Lecture Notes in Computer Science, 2141:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2141/21410159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2141/21410159.
pdf.

Barrett:2001:EAA

[1097] Chris Barrett, Doug Cook, Gregory Hicks, Vance Faber, Achla Marathe, Madhav Marathe, Aravind Srinivasan, Yoram J. Sussmann, and Heidi Thornquist. Experimental analysis of algorithms for bilateral-contract clearing mechanisms arising in deregulated power industry. Lecture Notes in Computer Science, 2141:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2141/21410172.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2141/21410172.
                                        pdf.

Muller-Hannemann:2001:PSP

[1098] Matthias Müller-Hannemann and Karsten Weihe. Pareto shortest paths is often feasible in practice. Lecture Notes in Computer Science, 2141:185– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2141/21410185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2141/21410185.
                                        pdf.

Anonymous:2001:AIgd

[1099] Anonymous. Author index. Lecture Notes in Computer Science, 2141:199– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2141/2141ind.pdf.

OHearn:2001:LRA

[1100] Peter O’Hearn, John Reynolds, and Hongseok Yang. Local reasoning about programs that alter data structures. Lecture Notes in Computer Science, 2142:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420001.
                                        pdf.

VandenBussche:2001:AAT

[1101] Jan Van den Bussche. Applications of Alfred Tarski’s ideas in database theory. Lecture Notes in Computer Science, 2142:20–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420020.
pdf.

Girard:2001:LSR

[1102] Jean-Yves Girard. Locus solum:

From the rules of logic to the logic of rules. Lecture Notes in Computer Science, 2142:38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420038.
pdf.

Kanovich:2001:EPH

[1103] Max Kanovich. The expressive power of horn monadic linear logic. Lecture Notes in Computer Science, 2142:39– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420039.
pdf.

Guglielmi:2001:NCM

[1104] Alessio Guglielmi and Lutz Straßburger. Non-commutativity and MELL in the calculus of structures. Lecture Notes in Computer Science, 2142:54– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420054.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420054.
                                        pdf.

Mogbil:2001:QCC

[1105] Virgile Mogbil. Quadratic correctness criterion for non-commutative logic. Lecture Notes in Computer Science, 2142:69–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2142/21420069.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420069.
                                        pdf.

Chen:2001:CCP

[1106] Yijia Chen and Enshao Shen. Capture complexity by partition. Lecture Notes in Computer Science, 2142:84– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420084.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420084.
                                        pdf.

Grohe:2001:ELT

[1107] Martin Grohe and Stefan Wöhrle. An existential locality theorem. Lecture Notes in Computer Science, 2142:99– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420099.
pdf.

Marion:2001:AAF

[1108] Jean-Yves Marion. Actual arithmetic and feasibility. Lecture Notes in Computer Science, 2142:115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420115.
pdf.

Schweikardt:2001:NOG

[1109] Nicole Schweikardt. The natural ordergeneric collapse for ω-representable databases over the rational and the real ordered group. Lecture Notes in Computer Science, 2142:130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420130.
pdf.

Power:2001:AFH

[1110] John Power and Konstantinos Tourlas.

An algebraic foundation for higraphs. Lecture Notes in Computer Science, 2142:145–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2142/21420145.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420145.
                                        pdf.

Korovina:2001:SCS

[1111] M. V. Korovina and O. V. Kudinov. Semantic characterisations of second-order computability over the real numbers. Lecture Notes in Computer Science, 2142:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420160.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420160.
                                        pdf.

Robinson:2001:ALR

[1112] Edmund Robinson and Giuseppe Rosolini. An abstract look at realizability. Lecture Notes in Computer Science, 2142:173–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2142/21420173.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420173.
                                        pdf.

Danos:2001:AI

[1113] Vincent Danos and Russell Harmer. The anatomy of innocence. Lecture Notes in Computer Science, 2142:188– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2142/21420188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420188.
pdf.

Bezem:2001:IEC

[1114] Marc Bezem. An improved extensionality criterion for higher-order logic programs. Lecture Notes in Computer Science, 2142:203–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420203.
pdf.

Stark:2001:LAS

[1115] Robert F. Stärk and Stanislas Nanchen.

A logic for abstract state machines. Lecture Notes in Computer Science, 2142:217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2142/21420217.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2142/21420217.
pdf.

vanEijck:2001:CHT

[1116] Jan van Eijck. Constrained hyper tableaux. Lecture Notes in Computer Science, 2142:232–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2142/21420232.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2142/21420232.pdf.

Lutz:2001:MLT

[1117] Carsten Lutz, Ulrike Sattler, and Frank Wolter. Modal logic and the two-variable fragment. Lecture Notes in Computer Science, 2142:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420247.
                                        pdf.

Koriche:2001:LAF

[1118] Frédéric Koriche. A logic for approximate first-order reasoning. Lecture Notes in Computer Science, 2142:262– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420262.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420262.
                                        pdf.

Dawar:2001:IFP

[1119] A. Dawar, E. Grädel, and S. Kreutzer. Inflationary fixed points in modal logic. Lecture Notes in Computer Science, 2142:277–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2142/21420277.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420277.
                                        pdf.

Alechina:2001:CKS

[1120] Natasha Alechina, Michael Mendler, Valeria de Paiva, and Eike Ritter. Categorical and Kripke semantics for constructive S4 modal logic. Lecture Notes in Computer Science, 2142:292– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420292.
pdf.

Rasmussen:2001:LND

[1121] Thomas Marthedal Rasmussen. Labelled natural deduction for interval logics. Lecture Notes in Computer Science, 2142:308–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2142/21420308.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2142/21420308.pdf.

deBoer:2001:DNL

[1122] Frank S. de Boer and Rogier M.

van Eijk. Decidable navigation logics for object structures. Lecture Notes in Computer Science, 2142:324– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420324.
pdf.

Charatonik:2001:DMC

[1123] Witold Charatonik and Jean-Marc Talbot. The decidability of model checking mobile ambients. Lecture Notes in Computer Science, 2142:339– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420339.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420339.
                                        pdf.

Galota:2001:GBE

[1124] Matthias Galota and Heribert Vollmer. A generalization of the Büchi-Elgot-Trakhtenbrot theorem. Lecture Notes in Computer Science, 2142:355–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420355.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420355.
                                        pdf.

Finkel:2001:EEW

[1125] Olivier Finkel. An effective extension of the wagner hierarchy to blind counter automata. Lecture Notes in Computer Science, 2142:369–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420369.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420369.
                                        pdf.

Klaedtke:2001:DPE

[1126] Felix Klaedtke. Decision procedure for an extension of WS1S. Lecture Notes in Computer Science, 2142:384– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420384.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420384.
pdf.

Akama:2001:LPC

[1127] Yohji Akama. Limiting partial combinatory algebras towards infinitary lambda-calculi and classical logic. Lecture Notes in Computer Science, 2142: 399–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420399.
pdf.

Rocca:2001:IL

[1128] Simona Ronchi Della Rocca and Luca Roversi. Intersection logic. Lecture Notes in Computer Science, 2142:414– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420414.
pdf.

Schroder:2001:LTT

[1129] Lutz Schröder. Life without the terminal type. Lecture Notes in Computer Science, 2142:429–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420429.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420429.
pdf.

Abramsky:2001:FCM

[1130] Samson Abramsky and Marina Lenisa. Fully complete minimal PER models for the simply typed λ-calculus. Lecture Notes in Computer Science, 2142: 443–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420443.
pdf.

Martin:2001:PI

[1131] Keye Martin. A principle of induction. Lecture Notes in Computer Science, 2142:458–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2142/21420458.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2142/21420458.
pdf.

Baaz:2001:GHT

[1132] Matthias Baaz and Georg Moser. On a generalisation of Herbrand’s theorem. Lecture Notes in Computer Science, 2142:469–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2142/21420469.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2142/21420469.
pdf.

Goubault-Larrecq:2001:WFR

[1133] Jean Goubault-Larrecq. Well-founded recursive relations. Lecture Notes in Computer Science, 2142:484–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420484.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420484.
pdf.

Schmidt-Schauss:2001:SCU

[1134] Manfred Schmidt-Schauß. Stratified context unification is in PSPACE. Lecture Notes in Computer Science, 2142: 498–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420498.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420498.
pdf.

Armando:2001:UDD

[1135] Alessandro Armando, Silvio Ranise, and Michaël Rusinowitch. Uniform derivation of decision procedures by superposition. Lecture Notes in Computer Science, 2142:513–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420513.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420513.
                                        pdf.

Rosu:2001:CCE

[1136] Grigore Roşu. Complete categorical equational deduction. Lecture Notes in Computer Science, 2142:528– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420528.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420528.
                                        pdf.

Ohsaki:2001:BRE

[1137] Hitoshi Ohsaki. Beyond regularity: Equational tree automata for associative and commutative theories. Lecture Notes in Computer Science, 2142: 539–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2142/21420539.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420539.
                                        pdf.

Courtieu:2001:NT

[1138] Pierre Courtieu. Normalized types. Lecture Notes in Computer Science, 2142:554–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2142/21420554.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2142/21420554.
                                        pdf.

Kopylov:2001:MPP

[1139] Alexei Kopylov and Aleksey Nogin. Markov’s principle for propositional type theory. Lecture Notes in Computer Science, 2142:570–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420570.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420570.
pdf.

Schurmann:2001:RHO

[1140] Carsten Schürmann. Recursion for higher-order encodings. Lecture Notes in Computer Science, 2142:585–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420585.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420585.
pdf.

Matthes:2001:MIC

[1141] Ralph Matthes. Monotone inductive and coinductive constructors of rank 2. Lecture Notes in Computer Science, 2142:600–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2142/21420600.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2142/21420600.
pdf.

Anonymous:2001:AIge

[1142] Anonymous. Author index. Lecture Notes in Computer Science, 2142:615– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2142/2142auth.pdf.

Jensen:2001:GML

[1143] Finn Verner Jensen. Graphical models as languages for computer assisted diagnosis and decision making. Lecture Notes in Computer Science, 2143: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430001.
                                        pdf.

Geffner:2001:PUI

[1144] Hector Geffner. Planning with uncertainty and incomplete information. Lecture Notes in Computer Science, 2143:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2143/21430016.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430016.
                                        pdf.

Schaub:2001:WYP

[1145] Torsten Schaub. What’s your preference? and how to express and implement it in logic programming! Lecture Notes in Computer Science, 2143:17– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430017.
pdf.

Grabisch:2001:PRO

[1146] Michel Grabisch. On preference representation on an ordinal scale. Lecture Notes in Computer Science, 2143:18– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430018.
pdf.

Greco:2001:RBD

[1147] Salvatore Greco, Benedetto Matarazzo, and Roman Slowinski. Rule-based decision support in multicriteria choice and ranking. Lecture Notes in Computer Science, 2143:29–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430029.
pdf.

Lafage:2001:PDP

[1148] Céline Lafage and Jérôme Lang.

Propositional distances and preference representation. Lecture Notes in Computer Science, 2143:48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430048.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430048.
                                        pdf.

Zhang:2001:VIB

[1149] Weihong Zhang. Value iteration over belief subspace. Lecture Notes in Computer Science, 2143:60–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430060.
                                        pdf.

Zhang:2001:SPV

[1150] Nevin L. Zhang and Weihong Zhang. Space-progressive value iteration: An anytime algorithm for a class of POMDPs. Lecture Notes in Computer Science, 2143:72–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2143/21430072.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2143/21430072.pdf.

Schut:2001:RAI

[1151] Martijn Schut, Michael Wooldridge, and Simon Parsons. Reasoning about intentions in uncertain domains. Lecture Notes in Computer Science, 2143: 84–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2143/21430084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430084.
pdf.

Vomlel:2001:TSM

[1152] Jiří Vomlel and Claus Skaanning. Troubleshooting with simultaneous models. Lecture Notes in Computer Science, 2143:96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2143/21430096.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2143/21430096.
pdf.

Bernardi:2001:RCU

[1153] Silvia Bernardi and Giulianella Coletti. A rational conditional utility model in a coherent framework. Lecture Notes in Computer Science, 2143: 108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430108.
pdf.

Coletti:2001:PRG

[1154] Giulianella Coletti, Romano Scozzafava, and Barbara Vantaggi. Probabilistic reasoning as a general unifying tool. Lecture Notes in Computer Science, 2143:120–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2143/21430120.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2143/21430120.pdf.

Capotorti:2001:OVC

[1155] Andrea Capotorti and Tania Paneni. An operational view of coherent conditional previsions. Lecture Notes in Computer Science, 2143:132–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430132.
                                        pdf.

Nielsen:2001:DID

[1156] Thomas D. Nielsen. Decomposition of influence diagrams. Lecture Notes in Computer Science, 2143:144–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430144.
                                        pdf.

Moral:2001:MTE

[1157] Serafín Moral, Rafael Rumí, and Antonio Salmerón. Mixtures of truncated exponentials in hybrid Bayesian networks. Lecture Notes in Computer Science, 2143:156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430156.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2143/21430156.
pdf.

Salmeron:2001:ISB

[1158] Antonio Salmerón and Serafín Moral.

Importance sampling in Bayesian networks using antithetic variables. Lecture Notes in Computer Science, 2143: 168–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430168.
pdf.

Darwiche:2001:URD

[1159] Adnan Darwiche and Mark Hopkins. Using recursive decomposition to construct elimination orders, jointrees, and dtrees. Lecture Notes in Computer Science, 2143:180–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430180.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430180.
pdf.

Dash:2001:CCR

[1160] Denver Dash and Marek Druzdzel.

Caveats for causal reasoning with equilibrium models. Lecture Notes in Computer Science, 2143:192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430192.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2143/21430192.
                                        pdf.

Lu:2001:SCS

[1161] Tsai-Ching Lu and Marek J. Druzdzel. Supporting changes in structure in causal model construction. Lecture Notes in Computer Science, 2143:204– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430204.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430204.
                                        pdf.

Acid:2001:SCO

[1162] Silvia Acid, Luis M. de Campos, and Juan F. Huete. The search of causal orderings: A short cut for learning belief networks. Lecture Notes in Computer Science, 2143:216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430216.
                                        pdf.

deCampos:2001:SLA

[1163] Luis M. de Campos and J. Miguel Puerta. Stochastic local algorithms for learning belief networks: Searching in the space of the orderings. Lecture Notes in Computer Science, 2143: 228–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430228.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2143/21430228.
pdf.

Borgelt:2001:EIK

[1164] Christian Borgelt and Rudolf Kruse.

An empirical investigation of the K2 metric. Lecture Notes in Computer Science, 2143:240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2143/21430240.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2143/21430240.pdf.

Demirer:2001:SVN

[1165] Riza Demirer and Prakash P. Shenoy.

Sequential valuation networks: A new graphical technique for asymmetric decision problems. Lecture Notes in Computer Science, 2143:252– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430252.
pdf.

Amor:2001:TSA

[1166] Nahla Ben Amor, Salem Benferhat, and Khaled Mellouli. A two-steps algorithm for min-based possibilistic causal networks. Lecture Notes in Computer Science, 2143:266–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430266.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430266.
                                        pdf.

Cano:2001:CIP

[1167] Andrés Cano and Serafín Moral. Computing intervals of probabilities with simulated annealing and probability trees. Lecture Notes in Computer Science, 2143:278–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430278.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430278.
                                        pdf.

Biazzo:2001:PLU

[1168] Veronica Biazzo, Angelo Gilio, Thomas Lukasiewicz, and Giuseppe Sanfilippo. Probabilistic logic under coherence, model-theoretic probabilistic logic, and default reasoning. Lecture Notes in Computer Science, 2143:290– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430290.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430290.
                                        pdf.

Kramosil:2001:BFP

[1169] Ivan Kramosil. Belief functions with partially ordered values. Lecture Notes in Computer Science, 2143:303– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2143/21430303.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430303.
pdf.

Monney:2001:DSM

[1170] Paul-André Monney. Dempster specialization matrices and the combination of belief functions. Lecture Notes in Computer Science, 2143:316– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430316.
pdf.

Baroni:2001:CSB

[1171] Pietro Baroni and Paolo Vicig.

On the conceptual status of belief functions with respect to coherent lower probabilities. Lecture Notes in Computer Science, 2143:328–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430328.
pdf.

Yaghlane:2001:ACB

[1172] Boutheina Ben Yaghlane, Philippe Smets, and Khaled Mellouli. About conditional belief function independence. Lecture Notes in Computer Science, 2143:340–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/

                                        series/0558/bibs/2143/21430340.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2143/21430340.pdf.

Elouedi:2001:ESR

[1173] Zied Elouedi, Khaled Mellouli, and Philippe Smets. The evaluation of sensors’ reliability and their tuning for multisensor data fusion within the transferable belief model. Lecture Notes in Computer Science, 2143:350– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430350.
                                        pdf.

Yaghlane:2001:CAB

[1174] Amel Ben Yaghlane, Thierry Denœux, and Khaled Mellouli. Coarsening approximations of belief functions. Lecture Notes in Computer Science, 2143: 362–373, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2143/21430362.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430362.
                                        pdf.

Lawry:2001:LSF

[1175] Jonathan Lawry. Label semantics: A formal framework for modeling with words. Lecture Notes in Computer Science, 2143:374–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2143/21430374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430374.
pdf.

Liu:2001:RAK

[1176] Weiru Liu. Reasoning about knowledge using rough sets. Lecture Notes in Computer Science, 2143:385–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430385.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430385.
pdf.

Sgarro:2001:CPC

[1177] Andrea Sgarro. The capacity of a possibilistic channel. Lecture Notes in Computer Science, 2143:398–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430398.
pdf.

Dubois:2001:NSQ

[1178] Didier Dubois, Henri Prade, and Philippe Smets. New semantics for quantitative possibility theory. Lecture Notes in Computer Science, 2143:410– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430410.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2143/21430410.
                                        pdf.

Benferhat:2001:BLC

[1179] Salem Benferhat, Didier Dubois, Souhila Kaci, and Henri Prade. Bridging logical, comparative, and graphical possibilistic representation frameworks. Lecture Notes in Computer Science, 2143:422–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430422.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430422.
                                        pdf.

Dufrenois:2001:EFU

[1180] Franck Dufrenois. Ellipse fitting with uncertainty and fuzzy decision stage for detection. application in videomicroscopy. Lecture Notes in Computer Science, 2143:432–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430432.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430432.
                                        pdf.

Fenton:2001:PMS

[1181] Norman Fenton, Paul Krause, and Martin Neil. Probabilistic modelling for software quality control. Lecture Notes in Computer Science, 2143:444– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2143/21430444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430444.
pdf.

Wurbel:2001:SIR

[1182] Éric Würbel, Robert Jeansoulin, and Odile Papini. Spatial information revision: A comparison between 3 approaches. Lecture Notes in Computer Science, 2143:454–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430454.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430454.
pdf.

Meyer:2001:SCM

[1183] Thomas Meyer, Aditya Ghose, and Samir Chopra. Social choice, merging, and elections. Lecture Notes in Computer Science, 2143:466–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430466.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430466.
pdf.

Cholvy:2001:DMT

[1184] Laurence Cholvy. Data merging: Theory of evidence vs. knowledge-bases merging operators. Lecture Notes in Computer Science, 2143:478–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2143/21430478.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430478.
                                        pdf.

DupindeSaint-Cyr:2001:PR

[1185] Florence Dupin de Saint-Cyr, Béatrice Duval, and Stéphane Loiseau. A priori revision. Lecture Notes in Computer Science, 2143:488–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430488.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430488.
                                        pdf.

Konieczny:2001:SOI

[1186] Sébastien Konieczny and Ramón Pino Pérez. Some operators for iterated revision. Lecture Notes in Computer Science, 2143:498–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430498.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430498.
                                        pdf.

Delgrande:2001:CSB

[1187] James P. Delgrande, Torsten Schaub, Hans Tompits, and Stefan Woltran. On computing solutions to belief change scenarios. Lecture Notes in Computer Science, 2143:510–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2143/21430510.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430510.
pdf.

Dubois:2001:SIV

[1188] Didier Dubois, Henri Prade, and Philippe Smets. “Not Impossible” vs. “Guaranteed Possible” in fusion and revision. Lecture Notes in Computer Science, 2143:522–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430522.
pdf.

Moinard:2001:GPE

[1189] Yves Moinard. General preferential entailments as circumscriptions. Lecture Notes in Computer Science, 2143:532– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430532.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430532.
pdf.

Hunter:2001:STV

[1190] Anthony Hunter. A semantic tableau version of first-order quasi-classical logic. Lecture Notes in Computer Science, 2143:544–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2143/21430544.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2143/21430544.pdf.

Koriche:2001:ACB

[1191] Frédéric Koriche. On anytime coherence-based reasoning. Lecture Notes in Computer Science, 2143:556– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430556.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430556.
                                        pdf.

Broersen:2001:RCB

[1192] Jan Broersen, Mehdi Dastani, and Leendert van der Torre. Resolving conflicts between beliefs, obligations, intentions, and desires. Lecture Notes in Computer Science, 2143:568– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430568.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430568.
                                        pdf.

Mercer:2001:CPW

[1193] Robert E. Mercer, Lionel Forget, and Vincent Risch. Comparing a pair-wise compatibility heuristic and relaxed stratification: Some preliminary results. Lecture Notes in Computer Science, 2143:580–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430580.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2143/21430580.
pdf.

Delgrande:2001:HRC

[1194] James P. Delgrande and Torsten Schaub. How to reason credulously and skeptically within a single extension. Lecture Notes in Computer Science, 2143:592–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430592.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430592.
pdf.

Kern-Isberner:2001:HCA

[1195] Gabriele Kern-Isberner. Handling conditionals adequately in uncertain reasoning. Lecture Notes in Computer Science, 2143:604–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430604.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430604.
pdf.

Weydert:2001:RWP

[1196] Emil Weydert. Rankings we prefer.

Lecture Notes in Computer Science, 2143:616–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2143/21430616.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2143/21430616.
pdf.

Bonnefon:2001:FHU

[1197] Jean-François Bonnefon and Denis J. Hilton. Formalizing human uncertain reasoning with default rules: A psychological conundrum and a pragmatic suggestion. Lecture Notes in Computer Science, 2143:628–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430628.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430628.
pdf.

Laming:2001:SIU

[1198] Donald Laming. Statistical information, uncertainty, and Bayes’ theorem: Some applications in experimental psychology. Lecture Notes in Computer Science, 2143:635–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430635.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430635.
pdf.

Neves:2001:PHJ

[1199] Rui Da Silva Neves and Eric Raufaste. Polymorphism of human judgment under uncertainty. Lecture Notes in Computer Science, 2143:647–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2143/21430647.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430647.
pdf.

Politzer:2001:HDA

[1200] Guy Politzer. How to doubt about a conditional. Lecture Notes in Computer Science, 2143:659–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430659.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430659.
pdf.

Cayrol:2001:DPT

[1201] Claudette Cayrol, Sylvie Doutre, and Jérôme Mengin. Dialectical proof theories for the credulous preferred semantics of argumentation frameworks. Lecture Notes in Computer Science, 2143: 668–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430668.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430668.
pdf.

Tamma:2001:AQP

[1202] Valentina Tamma and Simon Parsons. Argumentation and qualitative probabilistic reasoning using the kappa calculus. Lecture Notes in Computer Science, 2143:680–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2143/21430680.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430680.
                                        pdf.

Anrig:2001:IMR

[1203] Bernhard Anrig. Importance measures from reliability theory for probabilistic assumption-based reasoning. Lecture Notes in Computer Science, 2143: 692–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430692.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430692.
                                        pdf.

Khelfallah:2001:RNM

[1204] Mahat Khelfallah and Aïcha Mokhtari. Ramification in the normative method of causality. Lecture Notes in Computer Science, 2143:704–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2143/21430704.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2143/21430704.pdf.

Bell:2001:SEC

[1205] John Bell. Simultaneous events: Conflicts and preferences. Lecture Notes in Computer Science, 2143:714–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430714.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2143/21430714.
pdf.

Kshemkalyani:2001:ORR

[1206] Ajay Kshemkalyani and Roshan Kamath. Orthogonal relations for reasoning about abstract events. Lecture Notes in Computer Science, 2143:726– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430726.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430726.
pdf.

Bloch:2001:ERB

[1207] Isabelle Bloch, Ramón Pino-Pérez, and Carlos Uzcátegui. Explanatory relations based on mathematical morphology. Lecture Notes in Computer Science, 2143:736–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430736.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430736.
pdf.

Damasio:2001:MRL

[1208] Carlos Viegas Damásio and Luís Moniz Pereira. Monotonic and residuated logic programs. Lecture Notes in Computer Science, 2143:748–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430748.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2143/21430748.
                                        pdf.

Alsinet:2001:PPP

[1209] Teresa Alsinet and Lluís Godo. A proof procedure for possibilistic logic programming with fuzzy constants. Lecture Notes in Computer Science, 2143: 760–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430760.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430760.
                                        pdf.

Balbiani:2001:FOC

[1210] Philippe Balbiani and Dimiter Vakarelov. First-order characterization and modal analysis of indiscernibility and complementarity in information systems. Lecture Notes in Computer Science, 2143: 772–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430772.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2143/21430772.
                                        pdf.

Ferre:2001:CIK

[1211] Sébastien Ferré. Complete and incomplete knowledge in logical information systems. Lecture Notes in Computer Science, 2143:782–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2143/21430782.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2143/21430782.
pdf.

Altamirano:2001:EPC

[1212] E. Altamirano and G. Escalada-Imaz. Extending polynomiality to a class of non-clausal many-valued horn-like formulas. Lecture Notes in Computer Science, 2143:792–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430792.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430792.
pdf.

Ognjanovic:2001:GAS

[1213] Zoran Ognjanović, Jozef Kratica, and Miloš Milovanović. A genetic algorithm for satisfiability problem in a probabilistic logic: A first report. Lecture Notes in Computer Science, 2143:805– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2143/21430805.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2143/21430805.
pdf.

Anonymous:2001:AIgf

[1214] Anonymous. Author index. Lecture Notes in Computer Science, 2143:817– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2143/2143auth.pdf.

Johnson:2001:VFFa

[1215] Steven D. Johnson. View from the fringe of the fringe. Lecture Notes in Computer Science, 2144:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440001.
pdf.

Mycroft:2001:HSU

[1216] Alan Mycroft and Richard Sharp. Hardware synthesis using SAFL and application to processor design. Lecture Notes in Computer Science, 2144: 13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440013.
pdf.

Beers:2001:AHV

[1217] Robert Beers, Rajnish Ghughal, and Mark Aagaard. Applications of hierarchical verification in model checking. Lecture Notes in Computer Science, 2144:40–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440040.
pdf.

Shtrichman:2001:PTS

[1218] Ofer Shtrichman. Pruning techniques for the SAT-based bounded model checking problem. Lecture Notes in Computer Science, 2144:58– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440058.
pdf.

Moller:2001:HHP

[1219] M. Oliver Möller and Rajeev Alur.

Heuristics for hierarchical partitioning with application to model checking. Lecture Notes in Computer Science, 2144:71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2144/21440071.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2144/21440071.
pdf.

Beyer:2001:ERA

[1220] Dirk Beyer. Efficient reachability analysis and refinement checking of timed automata using BDDs. Lecture Notes in Computer Science, 2144:86– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440086.
pdf.

Siewe:2001:DRT

[1221] François Siewe and Dang Van Hung. Deriving real-time programs from duration calculus specifications. Lecture Notes in Computer Science, 2144:92– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440092.
                                        pdf.

Yorav:2001:RSB

[1222] Karen Yorav, Sagi Katz, and Ron Kiper. Reproducing synchronization bugs with model checking. Lecture Notes in Computer Science, 2144:98– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440098.
                                        pdf.

Turner:2001:FBD

[1223] Kenneth J. Turner and Ji He. Formally-based design evaluation. Lecture Notes in Computer Science, 2144: 104–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440104.
                                        pdf.

Berry:2001:ME

[1224] Gérard Berry and Ellen Sentovich.

Multiclock esterel. Lecture Notes in Computer Science, 2144:110–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440110.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440110.
pdf.

Albrecht:2001:RTM

[1225] Alvin R. Albrecht and Alan J. Hu.

Register transformations with multiple clock domains. Lecture Notes in Computer Science, 2144:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440126.
pdf.

Winstanley:2001:TPS

[1226] Anthony Winstanley and Mark Greenstreet. Temporal properties of self-timed rings. Lecture Notes in Computer Science, 2144:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440140.
pdf.

Ratzaby:2001:CAU

[1227] Gil Ratzaby, Shmuel Ur, and Yaron Wolfsthal. Coverability analysis using symbolic model checking. Lecture Notes in Computer Science, 2144:155– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440155.
                                        pdf.

He:2001:SHT

[1228] Ji He and Kenneth J. Turner. Specifying hardware timing with ET–Lotos. Lecture Notes in Computer Science, 2144:161–166, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440161.
                                        pdf.

Seceleanu:2001:FPD

[1229] Tiberiu Seceleanu and Juha Plosila. Formal pipeline design. Lecture Notes in Computer Science, 2144:167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440167.
                                        pdf.

Radhakrishnan:2001:VBB

[1230] Rajesh Radhakrishnan, Elena Teica, and Ranga Vemuri. Verification of basic block schedules using RTL transformations. Lecture Notes in Computer Science, 2144:173–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440173.
pdf.

McMillan:2001:PVF

[1231] K. L. McMillan. Parameterized verification of the FLASH cache coherence protocol by compositional model checking. Lecture Notes in Computer Science, 2144:179–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440179.
pdf.

Kaivola:2001:PEL

[1232] Roope Kaivola and Katherine Kohatsu. Proof engineering in the large: Formal verification of Pentium(R)4 floating-point divider. Lecture Notes in Computer Science, 2144:196–211, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440196.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2144/21440196.
                                        pdf.

McKeever:2001:TPC

[1233] Steve McKeever and Wayne Luk. Towards provably-correct hardware compilation tools based on pass separation techniques. Lecture Notes in Computer Science, 2144:212–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440212.
                                        pdf.

Sharp:2001:HLL

[1234] Richard Sharp and Alan Mycroft. A higher-level language for hardware synthesis. Lecture Notes in Computer Science, 2144:228–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440228.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440228.
                                        pdf.

Kort:2001:HVU

[1235] Iskander Kort, Sofiene Tahar, and Paul Curzon. Hierarchical verification using an MDG–HOL hybrid tool. Lecture Notes in Computer Science, 2144: 244–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440244.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2144/21440244.
pdf.

EnricoTronci:2001:ETL

[1236] Enrico Tronci, Giuseppe Della Penna, Benedetto Intrigila, and Marisa Venturini Zilli. Exploiting transition locality in automatic verification. Lecture Notes in Computer Science, 2144: 259–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440259.
pdf.

Copty:2001:EDF

[1237] Fady Copty, Amitai Irron, Osnat Weissberg, Nathan Kropp, and Gila Kamhi. Efficient debugging in a formal verification environment. Lecture Notes in Computer Science, 2144:275– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440275.
pdf.

Chauhan:2001:UCO

[1238] P. Chauhan, E. Clarke, S. Jha, J. Kukula, H. Veith, and D. Wang. Using combinatorial optimization methods for quantification scheduling. Lecture Notes in Computer Science, 2144: 293–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2144/21440293.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440293.
                                        pdf.

Esparza:2001:NRL

[1239] Javier Esparza and Claus Schröter. Net reductions for LTL model-checking. Lecture Notes in Computer Science, 2144:310–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2144/21440310.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440310.
                                        pdf.

Berg:2001:FVV

[1240] Christoph Berg and Christian Jacobi. Formal verification of the VAMP floating point unit. Lecture Notes in Computer Science, 2144:325–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440325.
                                        pdf.

Shimizu:2001:SMC

[1241] Kanna Shimizu, David L. Dill, and Ching-Tsun Chou. A specification methodology by a collection of compact properties as applied to the Intel(R) Itanium^{T M} processor bus protocol. Lecture Notes in Computer Science, 2144:340–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2144/21440340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440340.
pdf.

Claessen:2001:DVS

[1242] Koen Claessen, Mary Sheeran, and Satnam Singh. The design and verification of a sorter core. Lecture Notes in Computer Science, 2144:355– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440355.
pdf.

Kong:2001:RBF

[1243] Xiaohua Kong, Radu Negulescu, and Larry Weidong Ying. Refinementbased formal verification of asynchronous wrappers for independently clocked domains in systems on chip. Lecture Notes in Computer Science, 2144:370–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2144/21440370.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2144/21440370.
pdf.

Bhadra:2001:UAS

[1244] Jayanta Bhadra, Andrew Martin, Jacob Abraham, and Magdy Abadir. Using abstract specifications to verify PowerPC TM custom memories by symbolic trajectory evaluation. Lecture Notes in Computer Science, 2144:386– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440386.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440386.
                                        pdf.

Butler:2001:FVC

[1245] Ricky Butler, Víctor Carreño, Gilles Dowek, and César Muñoz. Formal verification of conflict detection algorithms. Lecture Notes in Computer Science, 2144:403–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2144/21440403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440403.
                                        pdf.

Gascard:2001:IOF

[1246] Eric Gascard and Laurence Pierre. Induction-oriented formal verification in symmetric interconnection networks. Lecture Notes in Computer Science, 2144:418–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2144/21440418.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2144/21440418.
                                        pdf.

Aagaard:2001:FMC

[1247] Mark D. Aagaard, Byron Cook, Nancy A. Day, and Robert B. Jones. A framework for microprocessor correctness statements. Lecture Notes in Computer Science, 2144:433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440433.
pdf.

Huibiao:2001:OSD

[1248] Zhu Huibiao, Jonathan P. Bowen, and He Jifeng. From operational semantics to denotational semantics for verilog. Lecture Notes in Computer Science, 2144:449–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440449.
pdf.

Xuandong:2001:EVC

[1249] Li Xuandong, Pei Yu, Zhao Jianhua, Li Yong, Zheng Tao, and Zheng Guoliang. Efficient verification of a class of linear hybrid automata using linear programming. Lecture Notes in Computer Science, 2144:465– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2144/21440465.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2144/21440465.
pdf.

Anonymous:2001:AIgg

[1250] Anonymous. Author index. Lecture Notes in Computer Science, 2144:481– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2144/2144auth.pdf.

Leyton:2001:CT

[1251] Michael Leyton. Chapter 1. transfer. Lecture Notes in Computer Science, 2145:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2145/21450001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2145/21450001.
                                        pdf.

Leyton:2001:CR

[1252] Michael Leyton. Chapter 2. recoverability. Lecture Notes in Computer Science, 2145:35–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2145/21450035.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2145/21450035.pdf.

Leyton:2001:CMTa

[1253] Michael Leyton. Chapter 3. mathematical theory of transfer, I. Lecture Notes in Computer Science, 2145:77– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2145/21450077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2145/21450077.
                                        pdf.

Leyton:2001:CMTb

[1254] Michael Leyton. Chapter 4. mathematical theory of transfer, II. Lecture Notes in Computer Science, 2145:115– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450115.
pdf.

Leyton:2001:CTG

[1255] Michael Leyton. Chapter 5. theory of grouping. Lecture Notes in Computer Science, 2145:135–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450135.
pdf.

Leyton:2001:CRMa

[1256] Michael Leyton. Chapter 6. robot manipulators. Lecture Notes in Computer Science, 2145:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2145/21450161.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2145/21450161.pdf.

Leyton:2001:CAT

[1257] Michael Leyton. Chapter 7. algebraic theory of inheritance. Lecture Notes in Computer Science, 2145:175– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450175.
pdf.

Leyton:2001:CRF

[1258] Michael Leyton. Chapter 8. reference frames. Lecture Notes in Computer Science, 2145:185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450185.
pdf.

Leyton:2001:CRMb

[1259] Michael Leyton. Chapter 9. relative motion. Lecture Notes in Computer Science, 2145:213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2145/21450213.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2145/21450213.pdf.

Leyton:2001:CSP

[1260] Michael Leyton. Chapter 10. surface primitives. Lecture Notes in Computer Science, 2145:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450229.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2145/21450229.
pdf.

Leyton:2001:CUGa

[1261] Michael Leyton. Chapter 11. unfolding groups, I. Lecture Notes in Computer Science, 2145:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450239.
pdf.

Leyton:2001:CUGb

[1262] Michael Leyton. Chapter 12. unfolding groups, II. Lecture Notes in Computer Science, 2145:257–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450257.
pdf.

Leyton:2001:CUGc

[1263] Michael Leyton. Chapter 13. unfolding groups, III. Lecture Notes in Computer Science, 2145:271–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450271.
pdf.

Leyton:2001:CMD

[1264] Michael Leyton. Chapter 14. mechanical design and manufacturing. Lecture Notes in Computer Science, 2145:299– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2145/21450299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2145/21450299.
                                        pdf.

Leyton:2001:CMTc

[1265] Michael Leyton. Chapter 15. A mathematical theory of architecture. Lecture Notes in Computer Science, 2145: 365–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2145/21450365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2145/21450365.
                                        pdf.

Leyton:2001:CSS

[1266] Michael Leyton. Chapter 16. solid structure. Lecture Notes in Computer Science, 2145:397–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2145/21450397.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2145/21450397.
                                        pdf.

Leyton:2001:CWFa

[1267] Michael Leyton. Chapter 17. wreath formulation of splines. Lecture Notes in Computer Science, 2145:423–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450423.
pdf.

Leyton:2001:CWFb

[1268] Michael Leyton. Chapter 18. wreath formulation of sweep representations. Lecture Notes in Computer Science, 2145:443–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2145/21450443.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2145/21450443.
pdf.

Leyton:2001:CPG

[1269] Michael Leyton. Chapter 19.

process grammar. Lecture Notes in Computer Science, 2145:455–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450455.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450455.
pdf.

Leyton:2001:CCL

[1270] Michael Leyton. Chapter 20. conservation laws of physics. Lecture Notes in Computer Science, 2145:467– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2145/21450467.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2145/21450467.
                                        pdf.

Leyton:2001:CM

[1271] Michael Leyton. Chapter 21. music. Lecture Notes in Computer Science, 2145:477–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2145/21450477.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2145/21450477.
                                        pdf.

Leyton:2001:CAE

[1272] Michael Leyton. Chapter 22. against the Erlanger program. Lecture Notes in Computer Science, 2145:495–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2145/21450495.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2145/21450495.
                                        pdf.

Leyton:2001:ASD

[1273] Michael Leyton. Appendix A. semidirect products. Lecture Notes in Computer Science, 2145:531–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2145/21450531.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2145/21450531.
                                        pdf.

Leyton:2001:ABS

[1274] Michael Leyton. Appendix B. symbols.

Lecture Notes in Computer Science, 2145:539–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2145/21450539.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2145/21450539.
pdf.

Leyton:2001:R

[1275] Michael Leyton. References. Lecture Notes in Computer Science, 2145:541– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2145/21450541.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2145/21450541.
pdf.

Anonymous:2001:In

[1276] Anonymous. Index. Lecture Notes in Computer Science, 2145:549–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2145/21450549.pdf.

Ajtai:2001:OSA

[1277] Miklós Ajtai, Ravi Kumar, and Dandapani Sivakumar. An overview of the sieve algorithm for the shortest lattice vector problem. Lecture Notes in Computer Science, 2146:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2146/21460001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460001.
pdf.

Blomer:2001:LSE

[1278] Johannes Blömer and Alexander May. Low secret exponent RSA revisited. Lecture Notes in Computer Science, 2146:4–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2146/21460004.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2146/21460004.
pdf.

Coppersmith:2001:FSS

[1279] Don Coppersmith. Finding small solutions to small degree polynomials. Lecture Notes in Computer Science, 2146: 20–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460020.
pdf.

Eisenbrand:2001:FRT

[1280] Friedrich Eisenbrand and Günter Rote. Fast reduction of ternary quadratic forms. Lecture Notes in Computer Science, 2146:32–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460032.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2146/21460032.
pdf.

vanHoeij:2001:FPV

[1281] Mark van Hoeij. Factoring polynomials and 0 − 1 vectors. Lecture Notes in Computer Science, 2146:45– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460045.
pdf.

Howgrave-Graham:2001:AIC

[1282] Nick Howgrave-Graham. Approximate integer common divisors. Lecture Notes in Computer Science, 2146:51– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460051.
pdf.

Koy:2001:SLRa

[1283] Henrik Koy and Claus Peter Schnorr.

Segment LLL-reduction of lattice bases. Lecture Notes in Computer Science, 2146:67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2146/21460067.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2146/21460067.pdf.

Koy:2001:SLRb

[1284] Henrik Koy and Claus Peter Schnorr. Segment LLL-reduction with floating point orthogonalization. Lecture Notes in Computer Science, 2146:81– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460081.
pdf.

Mahassni:2001:INR

[1285] Edwin El Mahassni, Phong Q. Nguyen, and Igor E. Shparlinski. The insecurity of Nyberg-Rueppel and other DSA-like signature schemes with partially known nonces. Lecture Notes in Computer Science, 2146:97–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460097.
pdf.

May:2001:DRM

[1286] Alexander May and Joseph H. Silverman. Dimension reduction methods for convolution modular lattices. Lecture Notes in Computer Science, 2146:110– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460110.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460110.
pdf.

Micciancio:2001:ILB

[1287] Daniele Micciancio. Improving lattice based cryptosystems using the Hermite Normal Form. Lecture Notes in Computer Science, 2146:126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460126.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460126.
pdf.

Nguyen:2001:TFLa

[1288] Phong Q. Nguyen and Jacques Stern. The two faces of lattices in cryptology. Lecture Notes in Computer Science, 2146:146–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460146.
pdf.

Semaev:2001:DLR

[1289] Igor Semaev. A 3-dimensional lattice reduction algorithm. Lecture Notes in Computer Science, 2146:181– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2146/21460181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2146/21460181.
pdf.

Trolin:2001:SVP

[1290] Mårten Trolin. The shortest vector problem in lattices with many cycles. Lecture Notes in Computer Science, 2146:194–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2146/21460194.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2146/21460194.
                                        pdf.

Wang:2001:MSI

[1291] Li ping Wang and Yue fei Zhu. Multisequence synthesis over an integral domain. Lecture Notes in Computer Science, 2146:206–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2146/21460206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2146/21460206.
                                        pdf.

Anonymous:2001:AIgh

[1292] Anonymous. Author index. Lecture Notes in Computer Science, 2146:219– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2146/21460219.pdf.

Flynn:2001:TTA

[1293] Michael J. Flynn and Albert A. Liddicoat. Technology trends and adaptive computing. Lecture Notes in Computer Science, 2147:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470001.
pdf.

Sawitzki:2001:PFR

[1294] Sergej Sawitzki, Steffen Köhler, and Rainer G. Spallek. Prototyping framework for reconfigurable processors. Lecture Notes in Computer Science, 2147: 6–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470006.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470006.
pdf.

Fisher:2001:EER

[1295] Chris Fisher, Kevin Rennie, Guanbin Xing, Stefan G. Berg, Kevin Bolding, John Naegle, Daniel Parshall, Dmitriy Portnov, Adnan Sulejmanpasic, and Carl Ebeling. An emulator for exploring RaPiD configurable computing architectures. Lecture Notes in Computer Science, 2147: 17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470017.
pdf.

Abke:2001:NPM

[1296] Joerg Abke and Erich Barke. A new placement method for direct mapping into LUT-based FPGAs. Lecture Notes in Computer Science, 2147:27– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470027.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470027.
                                        pdf.

Kannan:2001:FFG

[1297] Parivallal Kannan, Shankar Balachandran, and Dinesh Bhatia. fGREP — fast generic routing demand estimation for placed FPGA circuits. Lecture Notes in Computer Science, 2147: 37–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470037.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470037.
                                        pdf.

Lin:2001:MAP

[1298] Ernie Lin and Steven J. E. Wilton. Macrocell architectures for product term embedded memory arrays. Lecture Notes in Computer Science, 2147: 48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470048.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470048.
                                        pdf.

Goda:2001:GRC

[1299] Bryan S. Goda, Russell P. Kraft, Steven R. Carlough, Thomas W.

Krawczyk Jr., and John F. McDonald. Gigahertz reconfigurable computing using SiGe HBT BiCMOS FPGAs. Lecture Notes in Computer Science, 2147:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2147/21470059.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2147/21470059.
pdf.

Kasat:2001:MSF

[1300] Amit Kasat, Iyad Ouaiss, and Ranga Vemuri. Memory synthesis for FPGAbased reconfigurable computers. Lecture Notes in Computer Science, 2147: 70–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470070.
pdf.

Melnikoff:2001:IHM

[1301] Stephen J. Melnikoff, Steven F.

Quigley, and Martin J. Russell. Implementing a hidden Markov model speech recognition system in programmable logic. Lecture Notes in Computer Science, 2147:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470081.
pdf.

Albu:2001:INR

[1302] Felix Albu, Jiri Kadlec, Chris Softley, Rudolf Matousek, Antonin Hermanek, Nick Coleman, and Anthony Fagan. Implementation of (normalised) RLS lattice on virtex. Lecture Notes in Computer Science, 2147: 91–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470091.
                                        pdf.

Amira:2001:AMP

[1303] Abbes Amira, Ahmed Bouridane, and Peter Milligan. Accelerating matrix product on reconfigurable hardware for signal processing. Lecture Notes in Computer Science, 2147:101– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470101.
                                        pdf.

Cadambi:2001:SPD

[1304] Srihari Cadambi and Seth Copen Goldstein. Static profile-driven compilation for FPGAs. Lecture Notes in Computer Science, 2147:112–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470112.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2147/21470112.
pdf.

Wirthlin:2001:SRH

[1305] Michael J. Wirthlin, Brad L. Hutchings, and Carl Worth. Synthesizing RTL hardware from Java byte codes. Lecture Notes in Computer Science, 2147:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2147/21470123.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2147/21470123.
pdf.

Harbich:2001:PBS

[1306] Klaus Harbich and Erich Barke. PuMA ++: From behavioral specification to multi-FPGA-prototype. Lecture Notes in Computer Science, 2147:133– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470133.
pdf.

Kean:2001:SCF

[1307] Tom Kean. Secure configuration of field programmable gate arrays. Lecture Notes in Computer Science, 2147: 142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470142.
pdf.

McLoone:2001:SCF

[1308] Máire McLoone and John V. Mc-Canny. Single-chip FPGA implementation of the advanced encryption standard algorithm. Lecture Notes in Computer Science, 2147:152– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470152.
pdf.

McMillan:2001:JAE

[1309] Scott McMillan and Cameron Patterson. JBits^{T M} implementations of the Advanced Encryption Standard (Rijndael). Lecture Notes in Computer Science, 2147:162–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470162.
pdf.

Weinhardt:2001:TPP

[1310] Markus Weinhardt and Wayne Luk. Task-parallel programming of reconfigurable systems. Lecture Notes in Computer Science, 2147:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470172.
pdf.

Brebner:2001:CBR

[1311] Gordon Brebner and Oliver Diessel. Chip-based reconfigurable task management. Lecture Notes in Computer Science, 2147:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470182.
pdf.

Sudhir:2001:CCS

[1312] Suraj Sudhir, Suman Nath, and Seth Copen Goldstein. Configuration caching and swapping. Lecture Notes in Computer Science, 2147:192– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470192.
pdf.

Arias-Estrada:2001:MSM

[1313] Miguel Arias-Estrada and Juan M. Xicotencatl. Multiple stereo matching using an extended architecture. Lecture Notes in Computer Science, 2147:203– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470203.
pdf.

Mallon:2001:INB

[1314] Paula N. Mallón, Montserrat Bóo, and Javier D. Bruguera. Implementation of a NURBS to Bézier conversor with constant latency. Lecture Notes in Computer Science, 2147:213– 222, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470213.
pdf.

Cuenca:2001:RFG

[1315] Sergio A. Cuenca, Francisco Ibarra, and Rafael Alvarez. Reconfigurable frame-grabber for real-time automated visual inspection (RT-AVI) systems. Lecture Notes in Computer Science, 2147:223–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2147/21470223.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2147/21470223.
pdf.

Lawrence:2001:PMN

[1316] Jeff Lawrence. Processing models for the next generation network [abstract]. Lecture Notes in Computer Science, 2147:232–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470232.
pdf.

Kannan:2001:TIP

[1317] Parivallal Kannan and Dinesh Bhatia. Tightly integrated placement and routing for FPGAs. Lecture Notes in Computer Science, 2147:233– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470233.
pdf.

Karro:2001:GTS

[1318] John Karro and James Cohoon.

Gambit: A tool for the simultaneous placement and detailed routing of gate-arrays. Lecture Notes in Computer Science, 2147:243–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470243.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470243.
pdf.

Braun:2001:RRM

[1319] Florian Braun, John Lockwood, and Marcel Waldvogel. Reconfigurable router modules using network protocol wrappers. Lecture Notes in Computer Science, 2147:254–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470254.
pdf.

Ha:2001:DDF

[1320] Yajun Ha, Bingfeng Mei, Patrick Schaumont, Serge Vernalde, Rudy Lauwereins, and Hugo De Man. Development of a design framework for platform-independent networked reconfiguration of software and hardware. Lecture Notes in Computer Science, 2147:264–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470264.
                                        pdf.

Vassiliadis:2001:MCP

[1321] Stamatis Vassiliadis, Stephan Wong, and Sorin Cotofana. The MOLEN ρμ-coded processor. Lecture Notes in Computer Science, 2147:275–285, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470275.
                                        pdf.

Iliopoulos:2001:RTO

[1322] Marios Iliopoulos and Theodore Antonakopoulos. Run-time optimized reconfiguration using instruction forecasting. Lecture Notes in Computer Science, 2147:286–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470286.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2147/21470286.
pdf.

OpdeBeeck:2001:CTR

[1323] Pieter Op de Beeck, Francisco Barat, Murali Jayapala, and Rudy Lauwereins. CRISP: A template for reconfigurable instruction set processors. Lecture Notes in Computer Science, 2147: 296–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470296.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470296.
pdf.

Schafer:2001:EFI

[1324] Benjamin Carrion Schafer, Steven F.

Quigley, and Andrew H. C. Chan. Evaluation of an FPGA implementation of the discrete element method. Lecture Notes in Computer Science, 2147:306– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470306.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470306.
pdf.

Dandalis:2001:RTP

[1325] Andreas Dandalis, Viktor K. Prasanna, and Bharani Thiruvengadam. Runtime performance optimization of an FPGA-based deduction engine for SAT solvers. Lecture Notes in Computer Science, 2147:315–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470315.
                                        pdf.

Dollas:2001:REI

[1326] Apostolos Dollas, Kyprianos Papademetriou, Nikolaos Aslanides, and Tom Kean. A reconfigurable embedded input device for kinetically challenged persons. Lecture Notes in Computer Science, 2147:326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470326.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470326.
                                        pdf.

Wolz:2001:BPL

[1327] Frank Wolz and Reiner Kolla. Bubble partitioning for LUT-based sequential circuits. Lecture Notes in Computer Science, 2147:336–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470336.
                                        pdf.

Singh:2001:RCP

[1328] Satnam Singh and Philip James-Roxby. Rapid construction of partial configuration datastreams from high-level constructs using JBits. Lecture Notes in Computer Science, 2147:346–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470346.
pdf.

Lagadec:2001:PRE

[1329] Loïc Lagadec, Dominique Lavenier, Erwan Fabiani, and Bernard Pottier. Placing, routing, and editing virtual FPGAs. Lecture Notes in Computer Science, 2147:357–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470357.
pdf.

Ting:2001:VIP

[1330] Lok-Kee Ting, Roger Woods, and Colin Cowan. Virtex implementation of pipelined adaptive LMS predictor in electronic support measures receiver. Lecture Notes in Computer Science, 2147:367–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2147/21470367.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2147/21470367.
pdf.

Saito:2001:MSF

[1331] Takashi Saito, Tsutomu Maruyama, Tsutomu Hoshino, and Saburo Hirano. A music synthesizer on FPGA. Lecture Notes in Computer Science, 2147: 377–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470377.
                                        pdf.

Sheidaei:2001:MFR

[1332] Shervin Sheidaei, Hamid Noori, Ahmad Akbari, and Hosein Pedram. Motivation from a full-rate specific design to a DSP core approach for GSM vocoders. Lecture Notes in Computer Science, 2147:388–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470388.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470388.
                                        pdf.

Derrien:2001:LTR

[1333] Steven Derrien and Sanjay Rajopadhye. Loop tiling for reconfigurable accelerators. Lecture Notes in Computer Science, 2147:398–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470398.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470398.
                                        pdf.

Sassatelli:2001:SRD

[1334] Gilles Sassatelli, Lionel Torres, Jerome Galy, Gaston Cambon, and Camille Diou. The systolic ring: A dynamically reconfigurable architecture for embedded systems. Lecture Notes in Computer Science, 2147:409–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470409.
pdf.

Cadenas:2001:BRS

[1335] Oswaldo Cadenas and Graham Megson. A n-bit reconfigurable scalar quantiser. Lecture Notes in Computer Science, 2147:420–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2147/21470420.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2147/21470420.pdf.

Kasperek:2001:RTM

[1336] Jerzy Kasperek. Real time morphological image contrast enhancement in virtex FPGA. Lecture Notes in Computer Science, 2147:430–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470430.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470430.
pdf.

Simpson:2001:DRT

[1337] Albert Simpson, Jill Hunter, Moira Wylie, Yi Hu, and David Mann.

Demonstrating real-time JPEG image compression-decompression using standard component IP cores on a programmable logic based platform for DSP and image processing. Lecture Notes in Computer Science, 2147:441– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470441.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470441.
pdf.

Voss:2001:DIA

[1338] Nikolaus Voß and Bärbel Mertsching. Design and implementation of an accelerated Gabor filter bank using parallel hardware. Lecture Notes in Computer Science, 2147:451–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470451.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470451.
pdf.

Carter:2001:EPL

[1339] Bill Carter. The evolution of programmable logic: Past, present, and future predictions [abstract]. Lecture Notes in Computer Science, 2147:461– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470461.
pdf.

MacBeth:2001:DRC

[1340] John MacBeth and Patrick Lysaght.

Dynamically reconfigurable cores. Lecture Notes in Computer Science, 2147: 462–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470462.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470462.
pdf.

Price:2001:RBC

[1341] Tim Price and Cameron Patterson. Reconfigurable breakpoints for co-debug. Lecture Notes in Computer Science, 2147:473–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2147/21470473.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2147/21470473.
pdf.

Wheeler:2001:UDL

[1342] Timothy Wheeler, Paul Graham, Brent Nelson, and Brad Hutchings. Using design-level scan to improve FPGA design observability and controllability for functional verification. Lecture Notes in Computer Science, 2147:483– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470483.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470483.
pdf.

Civera:2001:FBF

[1343] Pierluigi Civera, Luca Macchiarulo, Maurizio Rebaudengo, Matteo Sonza Reorda, and Massimo Violante. FPGAbased fault injection techniques for fast evaluation of fault tolerance in VLSI circuits. Lecture Notes in Computer Science, 2147:493–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470493.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470493.
pdf.

Neumann:2001:GLA

[1344] Tilman Neumann and Andreas Koch. A generic library for adaptive computing environments. Lecture Notes in Computer Science, 2147:503–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470503.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470503.
pdf.

Ruhl:2001:GDS

[1345] Stephan Rühl, Peter Dillinger, Stefan Hezel, and Reinhard Männer. Generative development system for FPGA processors with active components. Lecture Notes in Computer Science, 2147:513–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2147/21470513.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2147/21470513.
pdf.

Cardoso:2001:CIS

[1346] João M. P. Cardoso and Horácio C.

Neto. Compilation increasing the scheduling scope for multi-memory-FPGA-based custom computing machines. Lecture Notes in Computer Science, 2147:523–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2147/21470523.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2147/21470523.pdf.

Hwang:2001:SLT

[1347] James Hwang, Brent Milne, Nabeel Shirazi, and Jeffrey D. Stroomer. System level tools for DSP in FPGAs. Lecture Notes in Computer Science, 2147: 534–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470534.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470534.
pdf.

Mencer:2001:PFE

[1348] Oskar Mencer, Nicolas Boullis, Wayne Luk, and Henry Styles. Parameterized function evaluation for FPGAs. Lecture Notes in Computer Science, 2147: 544–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470544.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2147/21470544.
                                        pdf.

Wirthlin:2001:ECC

[1349] Michael J. Wirthlin and Brian Mc-Murtrey. Efficient constant coefficient multiplication using advanced FPGA architectures. Lecture Notes in Computer Science, 2147:555–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470555.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470555.
                                        pdf.

Visavakul:2001:DSS

[1350] Chakkapas Visavakul, Peter Y. K. Cheung, and Wayne Luk. A digit-serial structure for reconfigurable multipliers. Lecture Notes in Computer Science, 2147:565–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470565.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470565.
                                        pdf.

Wires:2001:FRR

[1351] Kent E. Wires, Michael J. Schulte, and Don McCarley. FPGA resource reduction through truncated multiplication. Lecture Notes in Computer Science, 2147:574–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2147/21470574.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2147/21470574.
pdf.

Becker:2001:EMP

[1352] Jürgen Becker, Nicolas Liebau, Thilo Pionteck, and Manfred Glesner. Efficient mapping of pre-synthesized IPcores onto dynamically reconfigurable array architectures. Lecture Notes in Computer Science, 2147:584–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470584.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470584.
pdf.

Ciressan:2001:FBS

[1353] Cristian Ciressan, Eduardo Sanchez, Martin Rajman, and Jean-Cédric Chappelier. An FPGA-based syntactic parser for real-life almost unrestricted context-free grammars. Lecture Notes in Computer Science, 2147: 590–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470590.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470590.
pdf.

Harkin:2001:HSP

[1354] Jim Harkin, Thomas M. McGinnity, and Liam P. Maguire. Hardwaresoftware partitioning: A reconfigurable and evolutionary computing approach. Lecture Notes in Computer Science, 2147:595–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470595.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470595.
                                        pdf.

Maruyama:2001:ART

[1355] Tsutomu Maruyama, Yoshiki Yamaguchi, and Atsushi Kawase. An approach to real-time visualization of PIV method with FPGA. Lecture Notes in Computer Science, 2147:601– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470601.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470601.
                                        pdf.

Nibouche:2001:FBD

[1356] Mokhtar Nibouche, Ahmed Bouridane, Fionn Murtagh, and Omar Nibouche. FPGA-based discrete wavelet transforms system. Lecture Notes in Computer Science, 2147:607–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470607.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470607.
                                        pdf.

Nunez:2001:XMP

[1357] José Luis Núñez, Claudia Feregrino, Simon Jones, and Stephen Bateman. X-MatchPRO: A ProASICbased 200 Mbytes/s full-duplex lossless data compressor. Lecture Notes in Computer Science, 2147:613–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470613.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470613.
pdf.

Yamauchi:2001:AOO

[1358] Tsukasa Yamauchi, Shogo Nakaya, Takeshi Inuo, and Nobuki Kajihara. Arithmetic operation oriented reconfigurable chip: RHW. Lecture Notes in Computer Science, 2147:618–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470618.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470618.
pdf.

Dales:2001:IAP

[1359] Michael Dales. Initial analysis of the Proteus architecture. Lecture Notes in Computer Science, 2147:623– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470623.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470623.
pdf.

Keller:2001:BAC

[1360] Eric Keller. Building asynchronous circuits with JBits. Lecture Notes in Computer Science, 2147:628–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470628.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470628.
                                        pdf.

Lehmann:2001:CSI

[1361] Thomas Lehmann and Andreas Schreckenberg. Case study of integration of reconfigurable logic as a coprocessor into a SCI-cluster under RT-Linux. Lecture Notes in Computer Science, 2147:633– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470633.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470633.
                                        pdf.

Sinnappan:2001:RAP

[1362] Raymond Sinnappan and Scott Hazelhurst. A reconfigurable approach to packet filtering. Lecture Notes in Computer Science, 2147:638–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470638.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470638.
                                        pdf.

Stefo:2001:FBM

[1363] Riad Stefo, José Luis Núñez, Claudia Feregrino, Sudipta Mahapatra, and Simon Jones. FPGA-based modelling unit for high speed lossless arithmetic coding. Lecture Notes in Computer Science, 2147:643–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470643.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470643.
pdf.

Subramanian:2001:DRU

[1364] Ram Subramanian and Santosh Pande.

A data re-use based compiler optimization for FPGAs. Lecture Notes in Computer Science, 2147:648–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470648.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470648.
pdf.

Tommiska:2001:DSP

[1365] Matti Tommiska and Jorma Skyttä.

Dijkstra’s shortest path routing algorithm in reconfigurable hardware. Lecture Notes in Computer Science, 2147: 653–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2147/21470653.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2147/21470653.
pdf.

Urriza:2001:SCP

[1366] Isidoro Urriza, José I. García-Nicolás, Alfredo Sanz, and Antonio Valdovinos.

A system on chip for power line communications according to European home systems specifications. Lecture Notes in Computer Science, 2147:658– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2147/21470658.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2147/21470658.
                                        pdf.

Anonymous:2001:AIgi

[1367] Anonymous. Author index. Lecture Notes in Computer Science, 2147:663– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2147/2147indx.pdf.

Voss:2001:MHS

[1368] Stefan Voß. Meta-heuristics: The state of the art. Lecture Notes in Computer Science, 2148:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2148/21480001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2148/21480001.
                                        pdf.

Hamiez:2001:SSL

[1369] Jean-Philippe Hamiez and Jin-Kao Hao. Solving the sports league scheduling problem with tabu search. Lecture Notes in Computer Science, 2148:24– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2148/21480024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2148/21480024.
pdf.

Gerevini:2001:LML

[1370] Alfonso Gerevini and Ivan Serina.

Lagrange multipliers for local search on planning graphs. Lecture Notes in Computer Science, 2148:37–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2148/21480037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2148/21480037.
pdf.

Nareyek:2001:BPL

[1371] Alexander Nareyek. Beyond the plan-length criterion. Lecture Notes in Computer Science, 2148:55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2148/21480055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2148/21480055.
pdf.

Chien:2001:EEE

[1372] Steve Chien, Russell Knight, and Gregg Rabideau. An empirical evaluation of the effectiveness of local search for replanning. Lecture Notes in Computer Science, 2148:79– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2148/21480079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2148/21480079.
                                        pdf.

Knight:2001:BLT

[1373] Russell Knight, Gregg Rabideau, and Steve Chien. Board-laying techniques improve local search in mixed planning and scheduling. Lecture Notes in Computer Science, 2148:95– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2148/21480095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2148/21480095.
                                        pdf.

Engelhardt:2001:EEL

[1374] Barbara Engelhardt and Steve Chien. Empirical evaluation of local search methods for adapting planning policies in a stochastic environment. Lecture Notes in Computer Science, 2148:108– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2148/21480108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2148/21480108.
                                        pdf.

Refanidis:2001:GPN

[1375] Ioannis Refanidis and Ioannis Vlahavas. The GRT planner: New results. Lecture Notes in Computer Science, 2148:120–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2148/21480120.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2148/21480120.
pdf.

Onaindia:2001:ILS

[1376] Eva Onaindia, Laura Sebastia, and Eliseo Marzal. Incremental local search for planning problems. Lecture Notes in Computer Science, 2148:139– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2148/21480139.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2148/21480139.
pdf.

Hiraishi:2001:MDB

[1377] Hironori Hiraishi and Fumio Mizoguchi. Map drawing based on a resource-constrained search for a navigation system. Lecture Notes in Computer Science, 2148:158–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2148/21480158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2148/21480158.
pdf.

Anonymous:2001:AIgj

[1378] Anonymous. Author index. Lecture Notes in Computer Science, 2148:171– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2148/2148auth.pdf.

Miklos:2001:IMS

[1379] István Miklós and Zoltán Toroczkai. An improved model for statistical alignment. Lecture Notes in Computer Science, 2149:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490001.
                                        pdf.

vonOhsen:2001:IPP

[1380] Niklas von Öhsen and Ralf Zimmer. Improving profile-profile alignments via log average scoring. Lecture Notes in Computer Science, 2149:11– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490011.
                                        pdf.

Anantharaman:2001:FPG

[1381] Thomas Anantharaman and Bud Mishra. False positives in genomic map assembly and sequence validation. Lecture Notes in Computer Science, 2149: 27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490027.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490027.
                                        pdf.

Schiex:2001:BER

[1382] Thomas Schiex, Patrick Chabrier, Martin Bouchez, and Denis Milan. Boosting EM for radiation hybrid and genetic mapping. Lecture Notes in Computer Science, 2149:41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2149/21490041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2149/21490041.
pdf.

Casey:2001:PPA

[1383] Will Casey, Bud Mishra, and Mike Wigler. Placing probes along the genome using pairwise distance data. Lecture Notes in Computer Science, 2149:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2149/21490052.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2149/21490052.
pdf.

Jagota:2001:CHM

[1384] Arun Jagota, Rune B. Lyngsø, and Christian N. S. Pedersen. Comparing a hidden Markov model and a stochastic context-free grammar. Lecture Notes in Computer Science, 2149: 69–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2149/21490069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2149/21490069.
pdf.

Denise:2001:ASS

[1385] Alain Denise, Mireille Régnier, and Mathias Vandenbogaert. Assessing the statistical significance of overrepresented oligonucleotides. Lecture Notes in Computer Science, 2149:85– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490085.
                                        pdf.

Viksna:2001:PMP

[1386] Juris Vīksna and David Gilbert. Pattern matching and pattern discovery algorithms for protein topologies. Lecture Notes in Computer Science, 2149: 98–111, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490098.
                                        pdf.

Edelsbrunner:2001:CLN

[1387] Herbert Edelsbrunner and Afra Zomorodian. Computing linking numbers of a filtration. Lecture Notes in Computer Science, 2149:112–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490112.
                                        pdf.

Eriksson:2001:SCP

[1388] Olivia Eriksson, Yishao Zhou, and Arne Elofsson. Side chain-positioning as an integer programming problem. Lecture Notes in Computer Science, 2149:128–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2149/21490128.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2149/21490128.
pdf.

Pupko:2001:CDB

[1389] Tal Pupko, Roded Sharan, Masami Hasegawa, Ron Shamir, and Dan Graur. A chemical-distance-based test for positive darwinian selection. Lecture Notes in Computer Science, 2149: 142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2149/21490142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2149/21490142.
pdf.

Ganapathysaravanabavan:2001:FMC

[1390] Ganeshkumar Ganapathysaravanabavan and Tandy Warnow. Finding a maximum compatible tree for a bounded number of trees with bounded degree is solvable in polynomial time. Lecture Notes in Computer Science, 2149:156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2149/21490156.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2149/21490156.
                                        pdf.

Bergeron:2001:ECS

[1391] Anne Bergeron and François Strasbourg. Experiments in computing sequences of reversals. Lecture Notes in Computer Science, 2149:164–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490164.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490164.
                                        pdf.

Wang:2001:EIN

[1392] Li-San Wang. Exact-IEBP: A new technique for estimating evolutionary distances between whole genomes. Lecture Notes in Computer Science, 2149: 175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490175.
                                        pdf.

Siepel:2001:FOI

[1393] Adam C. Siepel and Bernard M. E. Moret. Finding an optimal inversion median: Experimental results. Lecture Notes in Computer Science, 2149:189– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490189.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2149/21490189.
pdf.

Chor:2001:AST

[1394] Benny Chor, Michael Hendy, and David Penny. Analytic solutions for three-taxon ML MC trees with variable rates across sites. Lecture Notes in Computer Science, 2149:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2149/21490204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2149/21490204.
pdf.

Nakhleh:2001:PPM

[1395] Luay Nakhleh, Usman Roshan, Katherine St. John, Jerry Sun, and Tandy Warnow. The performance of phylogenetic methods on trees of bounded diameter. Lecture Notes in Computer Science, 2149:214–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2149/21490214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2149/21490214.
pdf.

Eriksen:2001:ASR

[1396] Niklas Eriksen. (1 + ϵ)-approximation of sorting by reversals and transpositions. Lecture Notes in Computer Science, 2149:227–237, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2149/21490227.

                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2149/21490227.pdf.

Caprara:2001:PSR

[1397] Alberto Caprara. On the practical solution of the reversal median problem. Lecture Notes in Computer Science, 2149:238–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2149/21490238.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490238.
                                        pdf.

Heber:2001:AFG

[1398] Steffen Heber and Jens Stoye. Algorithms for finding gene clusters. Lecture Notes in Computer Science, 2149: 252–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490252.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2149/21490252.
                                        pdf.

vanderVeen:2001:DBA

[1399] Peter J. van der Veen, L. F. A. Wessels, J. W. Slootstra, R. H. Meloen, M. J. T. Reinders, and J. Hellendoorn. Determination of binding amino acids based on random peptide array screening data. Lecture Notes in Computer Science, 2149:264– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2149/21490264.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2149/21490264.
pdf.

Barash:2001:SHG

[1400] Yoseph Barash, Gill Bejerano, and Nir Friedman. A simple hyper-geometric approach for discovering putative transcription factor binding sites. Lecture Notes in Computer Science, 2149:278– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2149/21490278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2149/21490278.
pdf.

Huson:2001:CAU

[1401] Daniel H. Huson, Aaron L. Halpern, Zhongwu Lai, Eugene W. Myers, Knut Reinert, and Granger G. Sutton. Comparing assemblies using fragments and mate-pairs. Lecture Notes in Computer Science, 2149:294– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2149/21490294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2149/21490294.
pdf.

Anonymous:2001:AIgk

[1402] Anonymous. Author index. Lecture Notes in Computer Science, 2149:307– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2149/21490307.pdf.

Foster:2001:AGE

[1403] Ian Foster. The anatomy of the Grid: Enabling scalable virtual organizations. Lecture Notes in Computer Science, 2150:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500001.
                                        pdf.

Gannon:2001:SCT

[1404] Dennis Gannon. Software component technology for high performance parallel and Grid computing. Lecture Notes in Computer Science, 2150:5– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500005.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500005.
                                        pdf.

Kersten:2001:MMP

[1405] Martin Kersten, Stefan Manegold, Peter Boncz, and Niels Nes. Macro- and micro-parallelism in a DBMS. Lecture Notes in Computer Science, 2150:6– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500006.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500006.
                                        pdf.

Sterling:2001:IGP

[1406] Thomas Sterling. An introduction to the gilgamesh PIM architecture. Lecture Notes in Computer Science, 2150: 16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500016.
pdf.

Dongarra:2001:HPC

[1407] Jack Dongarra. High performance computing and trends: Connecting computational requirements with computing resources. Lecture Notes in Computer Science, 2150:33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500033.
pdf.

Gerndt:2001:TST

[1408] Michael Gerndt. Topic 01 support tools and environments. Lecture Notes in Computer Science, 2150:34– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500034.
pdf.

Morajko:2001:DPT

[1409] Anna Morajko, Eduardo César, Tomàs Margalef, Joan Sorribes, and Emilio Luque. Dynamic performance tuning environment. Lecture Notes in Computer Science, 2150:36–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500036.
pdf.

Ward:2001:SOH

[1410] Paul A. S. Ward and David J. Taylor. Self-organizing hierarchical cluster timestamps. Lecture Notes in Computer Science, 2150:46–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500046.
pdf.

Broberg:2001:TBT

[1411] Magnus Broberg, Lars Lundberg, and Håkan Grahn. A tool for binding threads to processors. Lecture Notes in Computer Science, 2150:57– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500057.
pdf.

Lowe:2001:VFV

[1412] Welf Löwe and Alex Liebrich. VizzScheduler — A framework for the visualization of scheduling algorithms. Lecture Notes in Computer Science, 2150:62–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2150/21500062.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2150/21500062.
pdf.

Muralidhar:2001:DOI

[1413] Rajeev Muralidhar and Manish Parashar.

A distributed object infrastructure for interaction and steering. Lecture Notes in Computer Science, 2150:67– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500067.
pdf.

Cardinale:2001:CFM

[1414] Yudith Cardinale and Emilio Hernández.

Checkpointing facility on a metasystem. Lecture Notes in Computer Science, 2150:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2150/21500075.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2150/21500075.pdf.

Booth:2001:OML

[1415] Stephen Booth. Optimising the MPI library for the T3E. Lecture Notes in Computer Science, 2150:80– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500080.
pdf.

Malony:2001:TPE

[1416] Allen D. Malony, Graham D. Riley, Bernd Mohr, Mark Bull, and Tomàs Margalef. Topic 02 performance evaluation and prediction. Lecture Notes in Computer Science, 2150: 84–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500084.
pdf.

Ponomarev:2001:OPL

[1417] Dmitry Ponomarev, Kanad Ghose, and Eugeny Saksonov. Optimal polling for latency-throughput tradeoffs in queue-based network interfaces for clusters. Lecture Notes in Computer Science, 2150:86–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500086.
pdf.

Gonzalez:2001:PPO

[1418] Jesús A. González, Coromoto León, Fabiana Piccoli, Marcela Printista, José L. Roda, Casiano Rodríguez, and Francisco de Sande. Performance prediction of oblivious BSP programs. Lecture Notes in Computer Science, 2150: 96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500096.
pdf.

Gautama:2001:PPD

[1419] Hasyim Gautama and Arjan J. C.

van Gemund. Performance prediction of data-dependent task parallel programs. Lecture Notes in Computer Science, 2150:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500106.
pdf.

Moreno:2001:TPP

[1420] Luz Marina Moreno, Francisco Almeida, Daniel González, and Casiano Rodríguez. The tuning problem on pipelines. Lecture Notes in Computer Science, 2150: 117–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500117.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2150/21500117.
                                        pdf.

DeRose:2001:HPM

[1421] Luiz A. DeRose. The hardware performance monitor toolkit. Lecture Notes in Computer Science, 2150:122– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500122.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500122.
                                        pdf.

Baker:2001:CPG

[1422] Mark Baker, Paul A. Farrell, Hong Ong, and Stephen L. Scott. VIA communication performance on a Gigabit Ethernet cluster. Lecture Notes in Computer Science, 2150:132–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500132.
                                        pdf.

Strey:2001:PAI

[1423] Alfred Strey and Martin Bange. Performance analysis of Intel’s MMX and SSE: A case study. Lecture Notes in Computer Science, 2150:142– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500142.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2150/21500142.
pdf.

Brunst:2001:GBP

[1424] Holger Brunst, Wolfgang E. Nagel, and Hans-Christian Hoppe. Group-based performance analysis for multithreaded SMP cluster applications. Lecture Notes in Computer Science, 2150:148– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500148.
pdf.

Ahmad:2001:TSL

[1425] Ishfaq Ahmad, Henri Casanova, Rupert Ford, and Yves Robert. Topic 03 scheduling and load balancing. Lecture Notes in Computer Science, 2150: 154–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500154.
pdf.

Boeres:2001:MPR

[1426] Cristina Boeres, Gerson N. da Cunha, and Vinod E. F. Rebello. On minimising the processor requirements of LogP schedules. Lecture Notes in Computer Science, 2150:156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500156.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500156.
                                        pdf.

Sinnen:2001:EUT

[1427] Oliver Sinnen and Leonel Sousa. Exploiting unused time slots in list scheduling considering communication contention. Lecture Notes in Computer Science, 2150:166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500166.
                                        pdf.

Chandra:2001:EPP

[1428] Sumir Chandra and Manish Parashar. An evaluation of partitioners for parallel SAMR applications. Lecture Notes in Computer Science, 2150:171– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500171.
                                        pdf.

Bahi:2001:LBN

[1429] Jacques M. Bahi and Jaafar Gaber. Load balancing on networks with dynamically changing topology. Lecture Notes in Computer Science, 2150:175– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500175.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2150/21500175.
pdf.

Cheung:2001:FLB

[1430] Lap-Sun Cheung and Yu-Kwong Kwok.

A fuzzy load balancing service for network computing based on jini. Lecture Notes in Computer Science, 2150:183– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500183.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500183.
pdf.

Blazewicz:2001:AAS

[1431] J. Błażewicz, M. Machowiak, G. Mounié, and D. Trystram. Approximation algorithms for scheduling independent malleable tasks. Lecture Notes in Computer Science, 2150:191–197, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500191.
pdf.

Costen:2001:WPQ

[1432] Fumie Costen and John Brooke.

The way to produce the quasiworkload in a cluster. Lecture Notes in Computer Science, 2150:198–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500198.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500198.
                                        pdf.

Knoop:2001:TCH

[1433] Jens Knoop, Manish Gupta, Keshav K. Pingali, and Michael F. P. O’Boyle. Topic 04 compilers for high performance. Lecture Notes in Computer Science, 2150:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500204.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500204.
                                        pdf.

Unger:2001:HIL

[1434] Sebastian Unger and Frank Mueller. Handling irreducible loops: Optimized node splitting vs. DJ-graphs. Lecture Notes in Computer Science, 2150:207– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500207.
                                        pdf.

Fernandez:2001:LRE

[1435] Manel Fernández, Roger Espasa, and Saumya Debray. Load redundancy elimination on executable code. Lecture Notes in Computer Science, 2150: 221–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500221.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2150/21500221.
pdf.

Faber:2001:LCC

[1436] Peter Faber, Martin Griebl, and Christian Lengauer. Loop-carried code placement. Lecture Notes in Computer Science, 2150:230–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500230.
pdf.

Qasem:2001:USI

[1437] Apan Qasem, David Whalley, Xin Yuan, and Robert van Engelen. Using a swap instruction to coalesce loads and stores. Lecture Notes in Computer Science, 2150:235–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500235.
pdf.

Chavarria-Miranda:2001:DPC

[1438] Daniel Chavarría-Miranda, John Mellor-Crummey, and Trushar Sarang. Data-parallel compiler support for multipartitioning. Lecture Notes in Computer Science, 2150:241–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2150/21500241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500241.
                                        pdf.

Knijnenburg:2001:CMI

[1439] Peter M. W. Knijnenburg, Toru Kisuki, and Kyle Gallivan. Cache models for iterative compilation. Lecture Notes in Computer Science, 2150:254– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500254.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500254.
                                        pdf.

Loechner:2001:DSL

[1440] Vincent Loechner, Benoît Meister, and Philippe Clauss. Data sequence locality: A generalization of temporal locality. Lecture Notes in Computer Science, 2150:262–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500262.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500262.
                                        pdf.

Sarkar:2001:EDA

[1441] Vivek Sarkar and Stephen Fink. Efficient dependence analysis for Java arrays. Lecture Notes in Computer Science, 2150:273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2150/21500273.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2150/21500273.pdf.

Kosch:2001:TPD

[1442] Harald Kosch, Pedro R. Falcone Sampaio, Abdelkader Hameurlain, and Lionel Brunie. Topic 05 parallel and distributed databases, data mining and knowledge discovery. Lecture Notes in Computer Science, 2150:278– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500278.
pdf.

deFMendesSampaio:2001:EPE

[1443] Sandra de F. Mendes Sampaio, Jim Smith, Norman W. Paton, and Paul Watson. An experimental performance evaluation of join algorithms for parallel object databases. Lecture Notes in Computer Science, 2150:280– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500280.
pdf.

Martens:2001:CSE

[1444] Holger Märtens. A classification of skew effects in parallel database systems. Lecture Notes in Computer Science, 2150:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500291.
                                        pdf.

Boukerche:2001:ICC

[1445] Azzedine Boukerche and Terry Tuck. Improving concurrency control in distributed databases with predeclared tables. Lecture Notes in Computer Science, 2150:301–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2150/21500301.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2150/21500301.pdf.

Guralnik:2001:PTP

[1446] Valerie Guralnik, Nivea Garg, and George Karypis. Parallel tree projection algorithm for sequence mining. Lecture Notes in Computer Science, 2150:310–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2150/21500310.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500310.
                                        pdf.

Gursoy:2001:PPK

[1447] Attila Gürsoy and Ilker Cengiz. Parallel pruning for K-means clustering on shared memory architectures. Lecture Notes in Computer Science, 2150:321– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2150/21500321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500321.
pdf.

Arlia:2001:EPC

[1448] Domenica Arlia and Massimo Coppola. Experiments in parallel clustering with DBSCAN. Lecture Notes in Computer Science, 2150:326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500326.
pdf.

Bilardi:2001:TCT

[1449] Gianfranco Bilardi, Rainer Feldmann, Kieran Herley, and Bruce Maggs. Topic 06 complexity theory and algorithms. Lecture Notes in Computer Science, 2150:332–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500332.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500332.
pdf.

Keller:2001:BEC

[1450] Jörg Keller and Jop F. Sibeyn. Beyond external computing: Analysis of the cycle structure of permutations. Lecture Notes in Computer Science, 2150:333– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500333.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500333.
                                        pdf.

Meyer:2001:HBT

[1451] Ulrich Meyer. Heaps are better than buckets: Parallel shortest paths on unbalanced graphs. Lecture Notes in Computer Science, 2150:343–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500343.
                                        pdf.

Lodha:2001:ESA

[1452] Sandeep Lodha, Punit Chandra, Ajay Kshemkalyani, and Mayank Rawat. Efficient synchronization of asynchronous processes. Lecture Notes in Computer Science, 2150:352–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500352.
                                        pdf.

Muraoka:2001:TAH

[1453] Yoichi Muraoka, Randall Bramley, David F. Snelling, and Harry Wijshoff. Topic 07 applications on high-performance computers. Lecture Notes in Computer Science, 2150:358– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500358.
pdf.

Schmidt:2001:SBD

[1454] Bertil Schmidt, Heiko Schröder, and Manfred Schimmler. Scanning biosequence databases on a hybrid parallel architecture. Lecture Notes in Computer Science, 2150:360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500360.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500360.
pdf.

Fung:2001:PCP

[1455] Y. F. Fung, M. F. Ercan, T. K. Ho, and W. L. Cheung. A parallel computation of power system equations. Lecture Notes in Computer Science, 2150: 371–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500371.
pdf.

Frankenfeld:2001:LTH

[1456] U. Frankenfeld, H. Helstrup, J. Lien, V. Lindenstruth, D. Röhrich, M. Schulz, B. Skaali, T. Steinbeck, K. Ullaland, A. Vestbø, and A. Wiebalck. Level-3 trigger for a heavy ion experiment at LHC. Lecture Notes in Computer Science, 2150:375–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500375.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500375.
pdf.

Allsopp:2001:EUM

[1457] Nicholas K. Allsopp, John F. Hague, and Jean-Pierre Prost. Experiences in using MPI–IO on top of GPFS for the IFS weather forecast code. Lecture Notes in Computer Science, 2150: 380–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500380.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500380.
pdf.

Ayguade:2001:TIL

[1458] Eduard Ayguadé, Fredrik Dahlgren, Christine Eisenbeis, Roger Espasa, Guang R. Gao, Henk Muller, Rizos Sakellariou, and André Seznec. Topic 08+13 instruction-level parallelism and computer architecture. Lecture Notes in Computer Science, 2150: 385–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500385.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500385.
pdf.

Ramirez:2001:BPU

[1459] Alex Ramirez, Josep L. Larriba-Pey, and Mateo Valero. Branch prediction using profile data. Lecture Notes in Computer Science, 2150:386–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500386.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500386.
pdf.

Chu:2001:EIB

[1460] Yul Chu and M. R. Ito. An efficient indirect branch predictor. Lecture Notes in Computer Science, 2150: 394–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500394.
pdf.

Ertl:2001:BEV

[1461] M. Anton Ertl and David Gregg. The behavior of efficient virtual machine interpreters on modern architectures. Lecture Notes in Computer Science, 2150:403–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2150/21500403.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2150/21500403.
pdf.

Iwama:2001:ICB

[1462] Chitaka Iwama, Niko Demus Barli, Shuichi Sakai, and Hidehiko Tanaka. Improving conditional branch prediction on speculative multithreading architectures. Lecture Notes in Computer Science, 2150:413–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500413.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500413.
pdf.

Onder:2001:IWW

[1463] Soner Önder and Rajiv Gupta. Instruction wake-up in wide issue superscalars. Lecture Notes in Computer Science, 2150:418–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500418.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500418.
pdf.

Sato:2001:ELR

[1464] Toshinori Sato and Itsujiro Arita. Execution latency reduction via variable latency pipeline and instruction reuse. Lecture Notes in Computer Science, 2150:428–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500428.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2150/21500428.
pdf.

Sebot:2001:MBT

[1465] Julien Sebot and Nathalie Drach-Temam. Memory bandwidth: The true bottleneck of SIMD multimedia performance on a superscalar processor. Lecture Notes in Computer Science, 2150: 439–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500439.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500439.
pdf.

Bulic:2001:MES

[1466] Patricio Bulić and Veselko Guštin.

Macro extension for SIMD processing. Lecture Notes in Computer Science, 2150:448–451, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2150/21500448.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2150/21500448.pdf.

Niar:2001:PDT

[1467] Smaïl Niar and Mahamed Adda. Performances of a dynamic threads scheduler. Lecture Notes in Computer Science, 2150:452–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2150/21500452.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2150/21500452.pdf.

Folliot:2001:TDS

[1468] Bertil Folliot, Giovanni Chiola, Peter Druschel, and Anne-Marie Kermarrec. Topic 09 distributed systems and algorithms. Lecture Notes in Computer Science, 2150:457–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500457.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500457.
                                        pdf.

Gradinariu:2001:SSN

[1469] Maria Gradinariu and Colette Johnen. Self-stabilizing neighborhood unique naming under unfair scheduler. Lecture Notes in Computer Science, 2150: 458–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500458.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500458.
                                        pdf.

Dahl:2001:ELM

[1470] Jörgen Dahl, Malolan Chetlur, and Philip A. Wilsey. Event list management in distributed simulation. Lecture Notes in Computer Science, 2150: 466–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500466.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500466.
                                        pdf.

Torres-Rojas:2001:PEP

[1471] Francisco J. Torres-Rojas. Performance evaluation of plausible clocks. Lecture Notes in Computer Science, 2150:476– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500476.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500476.
pdf.

Ezhilchelvan:2001:BTB

[1472] Paul Ezhilchelvan, Jean-Michel Hélary, and Michel Raynal. Building TMRbased reliable servers despite bounded input lifetimes. Lecture Notes in Computer Science, 2150:482–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500482.
pdf.

Klintskog:2001:FWR

[1473] Erik Klintskog, Anna Neiderud, Per Brand, and Seif Haridi. Fractional weighted reference counting. Lecture Notes in Computer Science, 2150:486– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500486.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500486.
pdf.

Baden:2001:TPP

[1474] Scott B. Baden, Paul H. J. Kelly, Sergei Gorlatch, and Calvin Lin. Topic 10 parallel programming: Models, methods and programming languages. Lecture Notes in Computer Science, 2150: 491–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500491.
pdf.

Christiaens:2001:ACL

[1475] Mark Christiaens and Koen De Bosschere. Accordion clocks: Logical clocks for data race detection. Lecture Notes in Computer Science, 2150: 494–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500494.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500494.
pdf.

Martel:2001:PEC

[1476] Matthieu Martel and Marc Gengler. Partial evaluation of concurrent programs. Lecture Notes in Computer Science, 2150:504–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500504.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500504.
pdf.

Venetis:2001:TOS

[1477] Ioannis E. Venetis, Dimitrios S.

Nikolopoulos, and Theodore S. Papatheodorou. A transparent operating system infrastructure for embedding adaptability to thread-based programming models. Lecture Notes in Computer Science, 2150:514–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500514.
pdf.

Chakravarty:2001:NND

[1478] Manuel M. T. Chakravarty, Gabriele Keller, Roman Lechtchinsky, and Wolf Pfannenstiel. Nepal — nested data parallelism in Haskell. Lecture Notes in Computer Science, 2150:524–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500524.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500524.
pdf.

Goodman:2001:ISL

[1479] Joy Goodman and John O’Donnell. Introduction of static load balancing in incremental parallel programming. Lecture Notes in Computer Science, 2150: 535–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500535.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2150/21500535.
                                        pdf.

Furmento:2001:CFH

[1480] Nathalie Furmento, Anthony Mayer, Stephen McGough, Steven Newhouse, and John Darlington. A component framework for HPC applications. Lecture Notes in Computer Science, 2150: 540–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500540.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500540.
                                        pdf.

Stewart:2001:TFR

[1481] Alan Stewart, Maurice Clint, Joquim Gabarró, and Maria J. Serna. Towards formally refining BSP barriers into explicit two-sided communications. Lecture Notes in Computer Science, 2150: 549–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500549.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500549.
                                        pdf.

Boulenouar:2001:SBK

[1482] Benamar Sidi Boulenouar. Solving bi-knapsack problem using tiling approach for dynamic programming. Lecture Notes in Computer Science, 2150: 560–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500560.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2150/21500560.
pdf.

vanderVorst:2001:TNA

[1483] Henk A. van der Vorst, Rob Bisseling, Iain S. Duff, and Bernard J. Philippe. Topic 11 numerical algorithms. Lecture Notes in Computer Science, 2150: 566–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500566.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500566.
pdf.

Cheng:2001:PIB

[1484] Sheung Hun Cheng and Nicholas J.

Higham. Parallel implementation of a block algorithm for matrix 1norm estimation. Lecture Notes in Computer Science, 2150:568–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500568.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500568.
pdf.

Thomas:2001:ESE

[1485] Ken S. Thomas, Simon J. Cox, Duan H.

Beckett, Ben P. Hiett, Jasek Generowicz, and Geoffrey J. Daniell. Eigenvalue spectrum estimation and photonic crystals. Lecture Notes in Computer Science, 2150:578–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500578.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500578.
                                        pdf.

Liang:2001:PPS

[1486] Y. Liang, J. Weston, and M. Szularz. Polynomial preconditioning for specially structured linear systems of equations. Lecture Notes in Computer Science, 2150:587–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500587.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500587.
                                        pdf.

Jimack:2001:PAN

[1487] Peter K. Jimack and Sarfraz A. Nadeem. Parallel application of a novel domain decomposition preconditioner for the stable finite element solution of three-dimensional convection-dominated PDEs. Lecture Notes in Computer Science, 2150:592– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500592.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500592.
                                        pdf.

Holmgren:2001:PHA

[1488] Sverker Holmgren and Dan Wallin. Performance of high-accuracy PDE solvers on a self-optimizing NUMA architecture. Lecture Notes in Computer Science, 2150:602–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2150/21500602.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2150/21500602.pdf.

Beivide:2001:TRC

[1489] Ramón Beivide, Chris Jesshope, Antonio Robles, and Cruz Izu. Topic 12 routing and communication in interconnection networks. Lecture Notes in Computer Science, 2150:611–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500611.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500611.
pdf.

Loucif:2001:AMD

[1490] Samia Loucif and Mohamed Ould-Khaoua. An analytical model of deterministic routing in the presence of hot-spot traffic. Lecture Notes in Computer Science, 2150:613–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500613.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500613.
pdf.

Alcover:2001:IAR

[1491] Rosa Alcover, Vicente Chirivella, and José Duato. Improving the accuracy of reliability models for direct interconnection networks. Lecture Notes in Computer Science, 2150:621– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500621.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500621.
                                        pdf.

Fernandez:2001:DFD

[1492] Lorenzo Fernández, José M. García, and Rafael Casado. On deadlock frequency during dynamic reconfiguration in NOWs. Lecture Notes in Computer Science, 2150:630–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500630.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500630.
                                        pdf.

Shahrabi:2001:ABC

[1493] A. Shahrabi, M. Ould-Khaoua, and L. M. Mackenzie. Analysis of broadcast communication in 2D tori. Lecture Notes in Computer Science, 2150: 639–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500639.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500639.
                                        pdf.

Pietracaprina:2001:OMO

[1494] Andrea Pietracaprina and Geppino Pucci. Optimal many-to-one routing on the mesh with constant queues. Lecture Notes in Computer Science, 2150: 645–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500645.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500645.
pdf.

Vassiliadis:2001:TME

[1495] Stamatis Vassiliadis, Francky Catthoor, Mateo Valero, and Sorin Cotofana. Topic 15+20 multimedia and embedded systems. Lecture Notes in Computer Science, 2150:651–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500651.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500651.
pdf.

Seinstra:2001:SAU

[1496] Frank Seinstra, Dennis Koelma, and Jan-Mark Geusebroek. A software architecture for user transparent parallel image processing on MIMD computers. Lecture Notes in Computer Science, 2150:653–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500653.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500653.
pdf.

Assarsson:2001:CSL

[1497] Ulf Assarsson and Per Stenström. A case study of load distribution in parallel view frustum culling and collision detection. Lecture Notes in Computer Science, 2150:663–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500663.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500663.
                                        pdf.

Kutil:2001:PZT

[1498] Rade Kutil. Parallelisable zero-tree image coding with significance maps. Lecture Notes in Computer Science, 2150: 674–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500674.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500674.
                                        pdf.

DmitriTcheressiz:2001:PCS

[1499] Dmitri Tcheressiz, Ben Juurlink, Stamatis Vassiliadis, and Harry Wijshoff. Performance of the complex streamed instruction set on image processing kernels. Lecture Notes in Computer Science, 2150:678–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500678.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500678.
                                        pdf.

El-Mahdy:2001:TDV

[1500] Ahmed El-Mahdy and Ian Watson.

A two dimensional vector architecture for multimedia. Lecture Notes in Computer Science, 2150:687–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500687.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500687.
pdf.

Kianzad:2001:MCE

[1501] Vida Kianzad and Shuvra S. Bhattacharyya. Multiprocessor clustering for embedded systems. Lecture Notes in Computer Science, 2150:697– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500697.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500697.
pdf.

Baker:2001:TCC

[1502] Mark Baker, John Brooke, Ken Hawick, and Rajkumar Buyya. Topic 16 cluster computing. Lecture Notes in Computer Science, 2150:702–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500702.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500702.
pdf.

Hansen:2001:PNE

[1503] Jørgen S. Hansen and Eric Jul. Prioritizing network event handling in clusters of workstations. Lecture Notes in Computer Science, 2150:704–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500704.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500704.
pdf.

Schreiner:2001:FTC

[1504] Wolfgang Schreiner, Gabor Kusper, and Karoly Bosa. Fault tolerance for cluster computing based on functional tasks. Lecture Notes in Computer Science, 2150:712–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500712.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500712.
pdf.

Renault:2001:PMP

[1505] Eric Renault and Pierre David. PAPI message passing library: Comparison of performance in user and kernel level messaging. Lecture Notes in Computer Science, 2150:717–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500717.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500717.
pdf.

Aridor:2001:IJC

[1506] Yariv Aridor, Michael Factor, and Avi Teperman. Implementing Java on clusters. Lecture Notes in Computer Science, 2150:722–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2150/21500722.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2150/21500722.pdf.

Solsona:2001:PCI

[1507] Francesc Solsona, Francesc Giné, Porfidio Hernández, and Emilio Luque. Predictive coscheduling implementation in a non-dedicated Linux cluster. Lecture Notes in Computer Science, 2150:732–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2150/21500732.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2150/21500732.
pdf.

Heymann:2001:SAS

[1508] Elisa Heymann, Miquel A. Senar, Emilio Luque, and Miron Livny. Selfadjusting scheduling of master-worker applications on distributed clusters. Lecture Notes in Computer Science, 2150:742–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2150/21500742.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2150/21500742.
pdf.

Kermarrec:2001:SEI

[1509] Anne-Marie Kermarrec and Christine Morin. Smooth and efficient integration of high-availability in a parallel single level store system. Lecture Notes in Computer Science, 2150:752– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500752.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500752.
pdf.

He:2001:OSA

[1510] Ligang He, Hai Jin, Ying Chen, and Zongfen Han. Optimal scheduling of aperiodic jobs on cluster. Lecture Notes in Computer Science, 2150:764– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500764.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500764.
pdf.

Munoz-Escoi:2001:HCM

[1511] Francesc D. Muñoz-Escoí, Óscar Gomis, Pablo Galdámez, and José M. Bernabéu-Aubán. HMM: A cluster membership service. Lecture Notes in Computer Science, 2150:773–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500773.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500773.
pdf.

Rose:2001:DPA

[1512] César A. F. De Rose and Hans-Ulrich Heiss. Dynamic processor allocation in large mesh-connected multicomputers. Lecture Notes in Computer Science, 2150:783–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500783.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500783.
pdf.

Ibanez:2001:NCM

[1513] Andres Ibañez, Valentin Puente, Jose Angel Gregorio, and Ramón Beivide. A new communication mechanism for cluster computing. Lecture Notes in Computer Science, 2150:793– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500793.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500793.
pdf.

Kalantar:2001:IDC

[1514] Michael Kalantar and Jun Fong. Isolated dynamic clusters for Web hosting. Lecture Notes in Computer Science, 2150:801–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2150/21500801.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2150/21500801.
pdf.

Reinefeld:2001:TMG

[1515] Alexander Reinefeld, Omer F. Rana, Jarek Nabrzyski, and David W. Walker. Topic 17 metacomputing and Grid computing. Lecture Notes in Computer Science, 2150:805–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500805.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500805.
pdf.

Ripeanu:2001:CAP

[1516] Matei Ripeanu, Adriana Iamnitchi, and Ian Foster. Cactus application: Performance predictions in Grid environments. Lecture Notes in Computer Science, 2150:807–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500807.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500807.
pdf.

Allen:2001:CGC

[1517] Gabrielle Allen, Werner Benger, Thomas Dramlitsch, Tom Goodale, Hans-Christian Hege, Gerd Lanfermann, André Merzky, Thomas Radke, and Edward Seidel. Cactus Grid computing: Review of current development. Lecture Notes in Computer Science, 2150:817–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2150/21500817.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500817.
pdf.

Erwin:2001:UGC

[1518] Dietmar W. Erwin and David F.

Snelling. UNICORE: A Grid computing environment. Lecture Notes in Computer Science, 2150:825–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500825.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500825.
pdf.

Denis:2001:PPC

[1519] Alexandre Denis, Christian Pérez, and Thierry Priol. Portable parallel CORBA objects: An approach to combine parallel and distributed programming for Grid computing. Lecture Notes in Computer Science, 2150: 835–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500835.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500835.
pdf.

Sevilla:2001:CLC

[1520] Diego Sevilla, José M. García, and Antonio Gómez. CORBA lightweight components: A model for distributed component-based heterogeneous computation. Lecture Notes in Computer Science, 2150:845–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500845.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500845.
                                        pdf.

Furmento:2001:BCC

[1521] Nathalie Furmento, Steven Newhouse, and John Darlington. Building computational communities from federated resources. Lecture Notes in Computer Science, 2150:855–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500855.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500855.
                                        pdf.

Bhatia:2001:SCM

[1522] Karan Bhatia, Keith Marzullo, and Lorenzo Alvisi. Scalable causal message logging for wide-area environments. Lecture Notes in Computer Science, 2150:864–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2150/21500864.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500864.
                                        pdf.

Balaton:2001:CMG

[1523] Zoltán Balaton, Péter Kacsuk, Norbert Podhorszki, and Ferenc Vajda. From cluster monitoring to Grid monitoring based on GRM. Lecture Notes in Computer Science, 2150:874–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500874.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500874.
pdf.

Cao:2001:UAB

[1524] Junwei Cao, Darren J. Kerbyson, and Graham R. Nudd. Use of agent-based service discovery for resource management in metacomputing environment. Lecture Notes in Computer Science, 2150:882–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500882.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500882.
pdf.

Brezany:2001:TPS

[1525] Peter Brezany, Marianne Winslett, Denis A. Nicole, and Toni Cortes. Topic 18 parallel I/O and storage technology. Lecture Notes in Computer Science, 2150:887–888, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500887.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500887.
pdf.

Ferhatosmanoglu:2001:OPE

[1526] Hakan Ferhatosmanoglu, Divyakant Agrawal, and Amr El Abbadi. Optimal partitioning for efficient I/O in spatial databases. Lecture Notes in Computer Science, 2150:889–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500889.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500889.
pdf.

Molero:2001:INP

[1527] Xavier Molero, Federico Silla, Vicente Santonja, and José Duato. Improving network performance by efficiently dealing with short control messages in fibre channel SANs. Lecture Notes in Computer Science, 2150:901– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500901.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500901.
pdf.

Ilroy:2001:IMP

[1528] Jonathan Ilroy, Cyrille Randriamaro, and Gil Utard. Improving MPI-I/O performance on PVFS. Lecture Notes in Computer Science, 2150:911– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500911.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500911.
pdf.

Walker:2001:TPS

[1529] David W. Walker, Ken Hawick, Domenico Laforenza, and Efstratios Gallopoulos. Topic 19 problem solving environments. Lecture Notes in Computer Science, 2150:916–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2150/21500916.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2150/21500916.
pdf.

Engquist:2001:RVD

[1530] Erik Engquist. Remote visualization of distributed electro-magnetic simulations. Lecture Notes in Computer Science, 2150:918–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2150/21500918.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2150/21500918.pdf.

Petcu:2001:SIV

[1531] Dana Petcu. Solving initial value problems with parallel Maple processes. Lecture Notes in Computer Science, 2150:926–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2150/21500926.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2150/21500926.
pdf.

Bunnin:2001:DPS

[1532] F. Oliver Bunnin, Yike Guo, and John Darlington. Design of problemsolving environment for contingent claim valuation. Lecture Notes in Computer Science, 2150:935–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500935.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2150/21500935.
                                        pdf.

Anonymous:2001:AIgl

[1533] Anonymous. Author index. Lecture Notes in Computer Science, 2150:939– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2150/2150ind.pdf.

Garzotto:2001:UWA

[1534] Franca Garzotto. Ubiquitous Web applications. Lecture Notes in Computer Science, 2151:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2151/21510001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2151/21510001.pdf.

Rusinkiewicz:2001:WSC

[1535] Marek Rusinkiewicz. From workflows to service composition in virtual enterprises. Lecture Notes in Computer Science, 2151:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510002.
pdf.

Schmidt:2001:SOW

[1536] Joachim W. Schmidt, Hans-Werner Sehring, Michael Skusa, and Axel Wienberg. Subject-oriented work: Lessons learned from an interdisciplinary content management project. Lecture Notes in Computer Science, 2151:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2151/21510003.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2151/21510003.
pdf.

Plodzien:2001:QOT

[1537] Jacek Płodzien and Kazimierz Subieta. Query optimization through removing dead subqueries. Lecture Notes in Computer Science, 2151:27– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510027.
pdf.

Corral:2001:IBC

[1538] Antonio Corral, Michael Vassilakopoulos, and Yannis Manolopoulos. The impact of buffering on closest pairs queries using R-trees. Lecture Notes in Computer Science, 2151:41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2151/21510041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2151/21510041.
                                        pdf.

Slivinskas:2001:EEQ

[1539] Giedrius Slivinskas and Christian S. Jensen. Enhancing an extensible query optimizer with support for multiple equivalence types. Lecture Notes in Computer Science, 2151:55– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2151/21510055.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2151/21510055.
                                        pdf.

Briukhov:2001:ISR

[1540] Dmitry O. Briukhov, Leonid A. Kalinichenko, and Nikolay A. Skvortsov. Information sources registration at a subject mediator as compositional development. Lecture Notes in Computer Science, 2151:70–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2151/21510070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2151/21510070.
                                        pdf.

Kang:2001:ETM

[1541] Yong-Kyoon Kang, Kyong-I Ku, and Yoo-Sung Kim. Extracting theme melodies by using a graphical clustering algorithm for content-based music information retrieval. Lecture Notes in Computer Science, 2151:84– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510084.
pdf.

Roussey:2001:MIS

[1542] Catherine Roussey, Sylvie Calabretto, and Jean-Marie Pinon. A multilingual information system based on knowledge representation. Lecture Notes in Computer Science, 2151:98– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510098.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510098.
pdf.

Pfoser:2001:CFU

[1543] Dieter Pfoser and Nectaria Tryfona.

Capturing fuzziness and uncertainty of spatiotemporal objects. Lecture Notes in Computer Science, 2151:112– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510112.
pdf.

Kang:2001:PBT

[1544] Yong-Kyoon Kang, Ki-Chang Kim, and Yoo-Sung Kim. Probability-based tile pre-fetching and cache replacement algorithms for Web geographical information systems. Lecture Notes in Computer Science, 2151:127– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2151/21510127.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2151/21510127.
                                        pdf.

Morzy:2001:OPQ

[1545] Tadeusz Morzy, Marek Wojciechowski, and Maciej Zakrzewicz. Optimizing pattern queries for Web access logs. Lecture Notes in Computer Science, 2151:141–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2151/21510141.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2151/21510141.
                                        pdf.

Puuronen:2001:EFSb

[1546] Seppo Puuronen, Iryna Skrypnyk, and Alexey Tsymbal. Ensemble feature selection based on contextual merit and correlation heuristics. Lecture Notes in Computer Science, 2151:155– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2151/21510155.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2151/21510155.
pdf.

Wojciechowski:2001:ICB

[1547] Marek Wojciechowski. Interactive constraint-based sequential pattern mining. Lecture Notes in Computer Science, 2151:169–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2151/21510169.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2151/21510169.pdf.

Karakaya:2001:EBS

[1548] Murat Karakaya and Özgür Ulusoy. Evaluation of a broadcast scheduling algorithm. Lecture Notes in Computer Science, 2151:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510182.
pdf.

Wietrzyk:2001:AWF

[1549] Vlad Ingar Wietrzyk, Makoto Takizawa, and Vijay Khandelwal. An architecture for work flows interoperability supporting electronic commerce. Lecture Notes in Computer Science, 2151:196–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2151/21510196.htm;   http:
//link.springer-ny.com/link/service/

series/0558/papers/2151/21510196.
pdf.

Norvaag:2001:OLM

[1550] Kjetil Nørvåg. Object and log management in temporal log-only object database systems. Lecture Notes in Computer Science, 2151:210–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510210.
pdf.

Christensen:2001:OCS

[1551] Helle L. Christensen, Mads L. Haslund, Henrik N. Nielsen, and Nectaria Tryfona. Operations for conceptual schema manipulation: Definitions and semantics. Lecture Notes in Computer Science, 2151:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510225.
pdf.

Lellahi:2001:OOD

[1552] Kazem Lellahi and Alexandre Zamulin. Object-oriented database as a dynamic system with implicit state. Lecture Notes in Computer Science, 2151:239– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510239.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2151/21510239.
pdf.

Pranevicius:2001:UAZ

[1553] Henrikas Pranevicius. The use of aggregate and Z formal methods for specification and analysis of distributed systems. Lecture Notes in Computer Science, 2151:253–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510253.
pdf.

Ray:2001:DTA

[1554] Indrakshi Ray and Indrajit Ray. Detecting termination of active database rules using symbolic model checking. Lecture Notes in Computer Science, 2151:266–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2151/21510266.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2151/21510266.
pdf.

Pokorny:2001:DMF

[1555] Jaroslav Pokorný and Peter Vojtáš. A data model for flexible querying. Lecture Notes in Computer Science, 2151: 280–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510280.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2151/21510280.
                                        pdf.

Kapopoulos:2001:ATN

[1556] Dimitris G. Kapopoulos and Michael Hatzopoulos. The arc-tree: A novel symmetric access method for multidimensional data. Lecture Notes in Computer Science, 2151:294–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2151/21510294.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2151/21510294.
                                        pdf.

Josifovski:2001:EJS

[1557] Vanja Josifovski, Timour Katchaounov, and Tore Risch. Evaluation of join strategies for distributed mediation. Lecture Notes in Computer Science, 2151:308–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2151/21510308.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2151/21510308.
                                        pdf.

Frasincar:2001:RBM

[1558] Flavius Frasincar, Geert Jan Houben, and Richard Vdovjak. An RMMbased methodology for hypermedia presentation design. Lecture Notes in Computer Science, 2151:323–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2151/21510323.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2151/21510323.
pdf.

Yang:2001:EMI

[1559] Xiaochun Yang, Ge Yu, and Guoren Wang. Efficiently mapping integrity constraints from relational database to XML document. Lecture Notes in Computer Science, 2151:338–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510338.
pdf.

Gergatsoulis:2001:WBS

[1560] Manolis Gergatsoulis, Yannis Stavrakas, Dimitris Karteris, Athina Mouzaki, and Dimitris Sterpis. A Web-based system for handling multidimensional information through MXML. Lecture Notes in Computer Science, 2151:352– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2151/21510352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2151/21510352.
pdf.

Biskup:2001:ADM

[1561] Joachim Biskup and Ralf Menzel.

An abstract database machine for cost driven design of object-oriented database schemas. Lecture Notes in Computer Science, 2151:366–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2151/21510366.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2151/21510366.
                                        pdf.

Anonymous:2001:AIgm

[1562] Anonymous. Author index. Lecture Notes in Computer Science, 2151:379– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2151/2151auth.pdf.

Jacobs:2001:JPV

[1563] Bart Jacobs. JavaCard program verification. Lecture Notes in Computer Science, 2152:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2152/21520001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2152/21520001.pdf.

Johnson:2001:VFFb

[1564] Steven D. Johnson. View from the fringe of the fringe (joint with CHARME 2001). Lecture Notes in Computer Science, 2152:4–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2152/21520004.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2152/21520004.
                                        pdf.

Shankar:2001:UDP

[1565] Natarajan Shankar. Using decision procedures with a higher-order logic. Lecture Notes in Computer Science, 2152: 5–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520005.
pdf.

Adams:2001:CAM

[1566] Andrew Adams, Martin Dunstan, Hanne Gottliebsen, Tom Kelsey, Ursula Martin, and Sam Owre. Computer algebra meets automated theorem proving: Integrating Maple and PVS. Lecture Notes in Computer Science, 2152: 27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520027.
pdf.

Arthan:2001:ICR

[1567] Rob D. Arthan. An irrational construction of R from Z. Lecture Notes in Computer Science, 2152:43– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520043.
pdf.

Asperti:2001:HSM

[1568] Andrea Asperti, Luca Padovani, Claudio Sacerdoti Coen, and Irene Schena. HELM and the semantic Math-Web. Lecture Notes in Computer Science, 2152:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520059.
pdf.

Bauer:2001:CRR

[1569] Gertrud Bauer and Markus Wenzel. Calculational reasoning revisited (an Isabelle/Isar experience). Lecture Notes in Computer Science, 2152:75– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520075.
pdf.

Bella:2001:MPA

[1570] Giampaolo Bella and Lawrence C. Paulson. Mechanical proofs about a non-repudiation protocol. Lecture Notes in Computer Science, 2152:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520091.
pdf.

Bickford:2001:PHP

[1571] Mark Bickford, Christoph Kreitz, Robbert van Renesse, and Xiaoming Liu. Proving hybrid protocols correct. Lecture Notes in Computer Science, 2152: 105–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520105.
pdf.

Bove:2001:NGR

[1572] Ana Bove and Venanzio Capretta.

Nested general recursion and partiality in type theory. Lecture Notes in Computer Science, 2152:121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520121.
pdf.

Caccamo:2001:HOC

[1573] Mario Cáccamo and Glynn Winskel.

A higher-order calculus for categories. Lecture Notes in Computer Science, 2152:136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2152/21520136.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2152/21520136.
pdf.

Capretta:2001:CFF

[1574] Venanzio Capretta. Certifying the Fast Fourier Transform with Coq. Lecture Notes in Computer Science, 2152:154– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520154.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520154.
pdf.

Daumas:2001:GLF

[1575] Marc Daumas, Laurence Rideau, and Laurent Théry. A generic library for floating-point numbers and its application to exact computing. Lecture Notes in Computer Science, 2152:169– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520169.
pdf.

Dennis:2001:OAC

[1576] Louise A. Dennis and Alan Smaill. Ordinal arithmetic: A case study for rippling in a higher order domain. Lecture Notes in Computer Science, 2152: 185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520185.
pdf.

Fairtlough:2001:ARH

[1577] Matt Fairtlough, Michael Mendler, and Xiaochun Cheng. Abstraction and refinement in higher order logic. Lecture Notes in Computer Science, 2152:201– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520201.
pdf.

Gay:2001:FFP

[1578] Simon J. Gay. A framework for the formalisation of pi calculus type systems in Isabelle/HOL. Lecture Notes in Computer Science, 2152:217– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520217.
pdf.

Helke:2001:RHA

[1579] Steffen Helke and Florian Kammüller.

Representing hierarchical automata in interactive theorem provers. Lecture Notes in Computer Science, 2152:233– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520233.
pdf.

Hemer:2001:RCL

[1580] David Hemer, Ian Hayes, and Paul Strooper. Refinement calculus for logic programming in Isabelle/HOL. Lecture Notes in Computer Science, 2152:249– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2152/21520249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2152/21520249.
                                        pdf.

Hurd:2001:PSP

[1581] Joe Hurd. Predicate subtyping with predicate sets. Lecture Notes in Computer Science, 2152:265–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2152/21520265.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2152/21520265.
                                        pdf.

Kellomaki:2001:SEO

[1582] Pertti Kellomäki. A structural embedding of ocsid in PVS. Lecture Notes in Computer Science, 2152:281– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2152/21520281.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2152/21520281.
                                        pdf.

Medina-Bulo:2001:CPB

[1583] Inmaculada Medina-Bulo, Francisco Palomo-Lozano, and José A. Alonso-Jiménez. A certified polynomial-based decision procedure for propositional logic. Lecture Notes in Computer Science, 2152:297–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2152/21520297.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2152/21520297.pdf.

Moore:2001:FST

[1584] J. Strother Moore. Finite set theory in ACL2. Lecture Notes in Computer Science, 2152:313–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520313.
pdf.

Naumov:2001:HNP

[1585] Pavel Naumov, Mark-Oliver Stehr, and José Meseguer. The HOL/NuPRL proof translator (A practical approach to formal interoperability). Lecture Notes in Computer Science, 2152:329– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2152/21520329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2152/21520329.
pdf.

Pichardie:2001:FCH

[1586] David Pichardie and Yves Bertot. Formalizing convex hull algorithms. Lecture Notes in Computer Science, 2152: 346–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2152/21520346.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2152/21520346.
                                        pdf.

Rival:2001:EFT

[1587] Xavier Rival and Jean Goubault-Larrecq. Experiments with finite tree automata in coq. Lecture Notes in Computer Science, 2152:362–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2152/21520362.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2152/21520362.
                                        pdf.

Wiedijk:2001:MLH

[1588] Freek Wiedijk. Mizar light for HOL light. Lecture Notes in Computer Science, 2152:378–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2152/21520378.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2152/21520378.pdf.

Anonymous:2001:AIgn

[1589] Anonymous. Author index. Lecture Notes in Computer Science, 2152:395– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/2152/21520395.pdf.

Fekete:2001:SHP

[1590] Sándor P. Fekete, Henk Meijer, André Rohe, and Walter Tietze. Solving a “hard” problem to approximate an “easy” one: Heuristics for maximum matchings and maximum traveling salesman problems. Lecture Notes in Computer Science, 2153:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2153/21530001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2153/21530001.
pdf.

Mehlhorn:2001:CPC

[1591] K. Mehlhorn and M. Ziegelmann.

CNOP — A package for constrained network optimization. Lecture Notes in Computer Science, 2153:17–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2153/21530017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2153/21530017.
pdf.

Cirasella:2001:ATS

[1592] Jill Cirasella, David S. Johnson, Lyle A. McGeoch, and Weixiong Zhang. The asymmetric traveling salesman problem: Algorithms, instance generators, and tests. Lecture Notes in Computer Science, 2153:32– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2153/21530032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2153/21530032.
                                        pdf.

Towsley:2001:NTT

[1593] Don Towsley. Network tomography through end-to-end measurements. Lecture Notes in Computer Science, 2153:60–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2153/21530060.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2153/21530060.
                                        pdf.

Leung:2001:ERS

[1594] Vitus Leung and Sandy Irani. Experimental results on statistical approaches to page replacement policies. Lecture Notes in Computer Science, 2153:61– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2153/21530061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2153/21530061.
                                        pdf.

Mitzenmacher:2001:ERM

[1595] Michael Mitzenmacher and Sean Owen. Estimating resemblance of MIDI documents. Lecture Notes in Computer Science, 2153:78–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2153/21530078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2153/21530078.
pdf.

Demaine:2001:EAS

[1596] Erik D. Demaine, Alejandro López-Ortiz, and J. Ian Munro. Experiments on adaptive set intersections for text retrieval systems. Lecture Notes in Computer Science, 2153:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2153/21530091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2153/21530091.
pdf.

Sethia:2001:PSI

[1597] Saurabh Sethia, Martin Held, and Joseph S. B. Mitchell. PVD: A stable implementation for computing Voronoi diagrams of polygonal pockets. Lecture Notes in Computer Science, 2153: 105–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2153/21530105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2153/21530105.
pdf.

Finocchi:2001:HCT

[1598] Irene Finocchi and Rossella Petreschi.

Hierarchical clustering of trees: Algorithms and experiments. Lecture Notes in Computer Science, 2153:117– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2153/21530117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2153/21530117.
                                        pdf.

Brandes:2001:TPS

[1599] Ulrik Brandes, Frank Schulz, Dorothea Wagner, and Thomas Willhalm. Travel planning with self-made maps. Lecture Notes in Computer Science, 2153:132– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2153/21530132.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2153/21530132.
                                        pdf.

Willinger:2001:NAC

[1600] Walter Willinger. New algorithmic challenges arising in measurementdriven networking research. Lecture Notes in Computer Science, 2153:145– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2153/21530145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2153/21530145.
                                        pdf.

Chavez:2001:PSC

[1601] Edgar Chávez and Gonzalo Navarro. A probabilistic spell for the curse of dimensionality. Lecture Notes in Computer Science, 2153:147–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2153/21530147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2153/21530147.
pdf.

Breimer:2001:EEH

[1602] Eric Breimer, Mark Goldberg, Brian Kolstad, and Malik Magdon-Ismail. Experimental evaluation of the height of a random set of points in a ddimensional cube. Lecture Notes in Computer Science, 2153:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2153/21530161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2153/21530161.
pdf.

Maneewongvatana:2001:ESN

[1603] Songrit Maneewongvatana and David M.

Mount. An empirical study of a new approach to nearest neighbor searching. Lecture Notes in Computer Science, 2153:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2153/21530172.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2153/21530172.
pdf.

Karlin:2001:SAD

[1604] Anna R. Karlin. Spectral analysis for data mining. Lecture Notes in Computer Science, 2153:188–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2153/21530188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2153/21530188.
                                        pdf.

Makinen:2001:TBC

[1605] Veli Mäkinen. Trade off between compression and search times in compact suffix array. Lecture Notes in Computer Science, 2153:189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2153/21530189.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2153/21530189.
                                        pdf.

Hepner:2001:IPS

[1606] Clint Hepner and Cliff Stein. Implementation of a PTAS for scheduling with release dates. Lecture Notes in Computer Science, 2153:202–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2153/21530202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2153/21530202.
                                        pdf.

Ergun:2001:BSL

[1607] Funda Ergun, S. Cenk Sahinalp, Jonathan Sharp, and Rakesh K. Sinha. Biased skip lists for highly skewed access patterns. Lecture Notes in Computer Science, 2153:216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2153/21530216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2153/21530216.
pdf.

Anonymous:2001:AIgo

[1608] Anonymous. Author index. Lecture Notes in Computer Science, 2153:231– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2153/2153auth.pdf.

Abdulla:2001:CRP

[1609] Parosh Aziz Abdulla and Bengt Jonsson. Channel representations in protocol verification. Lecture Notes in Computer Science, 2154:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540001.
pdf.

Milner:2001:BRS

[1610] Robin Milner. Bigraphical reactive systems. Lecture Notes in Computer Science, 2154:16–35, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540016.
pdf.

Sastry:2001:CNU

[1611] Shankar Sastry. Control of networks of unmanned vehicles. Lecture Notes in Computer Science, 2154:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540036.
pdf.

Schneider:2001:PAS

[1612] Steve Schneider. Process algebra and security. Lecture Notes in Computer Science, 2154:37–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540037.
pdf.

Hatcliff:2001:UBT

[1613] John Hatcliff and Matthew Dwyer. Using the bandera tool set to model-check properties of concurrent Java software. Lecture Notes in Computer Science, 2154:39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540039.
pdf.

Hermanns:2001:PEP

[1614] Holger Hermanns and Joost-Pieter Katoen. Performance evaluation : = (process algebra + model checking) × Markov chains. Lecture Notes in Computer Science, 2154:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540059.
pdf.

Castagna:2001:TMS

[1615] Giuseppe Castagna, Giorgio Ghelli, and Francesco Zappa Nardelli. Typing mobility in the Seal Calculus. Lecture Notes in Computer Science, 2154:82– 101, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540082.
pdf.

Bugliesi:2001:RAS

[1616] Michele Bugliesi, Giuseppe Castagna, and Silvia Crafa. Reasoning about security in mobile ambients. Lecture Notes in Computer Science, 2154:102– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540102.
pdf.

Hirsch:2001:SHR

[1617] Dan Hirsch and Ugo Montanari. Synchronized hyperedge replacement with name mobility. Lecture Notes in Computer Science, 2154:121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540121.
                                        pdf.

Attie:2001:DIO

[1618] Paul C. Attie and Nancy A. Lynch. Dynamic input/output automata: A formal model for dynamic systems. Lecture Notes in Computer Science, 2154: 137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540137.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540137.
                                        pdf.

Aldini:2001:PIF

[1619] Alessandro Aldini. Probabilistic information flow in a process algebra. Lecture Notes in Computer Science, 2154: 152–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540152.
                                        pdf.

Kwiatkowska:2001:SCM

[1620] Marta Kwiatkowska, Gethin Norman, and Jeremy Sproston. Symbolic computation of maximal probabilistic reachability. Lecture Notes in Computer Science, 2154:169–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540169.
pdf.

Volzer:2001:RNS

[1621] Hagen Völzer. Randomized nonsequential processes. Lecture Notes in Computer Science, 2154:184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540184.
pdf.

Puhakka:2001:LFP

[1622] Antti Puhakka and Antti Valmari.

Liveness and fairness in processalgebraic verification. Lecture Notes in Computer Science, 2154:202–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540202.
pdf.

Heljanko:2001:BRC

[1623] Keijo Heljanko. Bounded reachability checking with process semantics. Lecture Notes in Computer Science, 2154: 218–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540218.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540218.
                                        pdf.

Geldenhuys:2001:TSI

[1624] Jaco Geldenhuys and Antti Valmari. Techniques for smaller intermediary BDDs. Lecture Notes in Computer Science, 2154:233–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540233.
                                        pdf.

Bouyer:2001:ACD

[1625] Patricia Bouyer, Antoine Petit, and Denis Thérien. An algebraic characterization of data and timed languages. Lecture Notes in Computer Science, 2154:248–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540248.
                                        pdf.

Luttgen:2001:FTR

[1626] Gerald Lüttgen and Walter Vogler.

A faster-than relation for asynchronous processes. Lecture Notes in Computer Science, 2154:262–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540262.
pdf.

Srba:2001:PLT

[1627] Jiří Srba. On the power of labels in transition systems. Lecture Notes in Computer Science, 2154:277– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540277.
pdf.

Sangiorgi:2001:BEC

[1628] Davide Sangiorgi and David Walker.

On barbed equivalences in π-calculus. Lecture Notes in Computer Science, 2154:292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2154/21540292.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2154/21540292.
pdf.

Phillips:2001:CPG

[1629] Iain Phillips. CCS with priority guards.

Lecture Notes in Computer Science, 2154:305–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2154/21540305.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2154/21540305.
pdf.

Lopez:2001:TTG

[1630] Natalia López and Manuel Núñez. A testing theory for generally distributed stochastic processes. Lecture Notes in Computer Science, 2154:321– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540321.
pdf.

vanBreugel:2001:AQV

[1631] Franck van Breugel and James Worrell. An algorithm for quantitative verification of probabilistic transition systems. Lecture Notes in Computer Science, 2154:336–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540336.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540336.
pdf.

deAlfaro:2001:CMP

[1632] Luca de Alfaro, Thomas A. Henzinger, and Ranjit Jhala. Compositional methods for probabilistic systems. Lecture Notes in Computer Science, 2154:351– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540351.
pdf.

Khomenko:2001:TEA

[1633] Victor Khomenko and Maciej Koutny.

Towards an efficient algorithm for unfolding Petri nets. Lecture Notes in Computer Science, 2154:366–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540366.
pdf.

Baldan:2001:SAT

[1634] Paolo Baldan, Andrea Corradini, and Barbara König. A static analysis technique for graph transformation systems. Lecture Notes in Computer Science, 2154:381–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540381.
pdf.

Niebert:2001:LFS

[1635] Peter Niebert, Michaela Huhn, Sarah Zennou, and Denis Lugiez. Local first search — A new paradigm for partial order reductions. Lecture Notes in Computer Science, 2154:396–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540396.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540396.
                                        pdf.

Glusman:2001:EMC

[1636] Marcelo Glusman and Shmuel Katz. Extending memory consistency of finite prefixes to infinite computations. Lecture Notes in Computer Science, 2154: 411–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540411.
                                        pdf.

Godefroid:2001:ABM

[1637] Patrice Godefroid, Michael Huth, and Radha Jagadeesan. Abstractionbased model checking using modal transition systems. Lecture Notes in Computer Science, 2154:426–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540426.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540426.
                                        pdf.

Chechik:2001:EMV

[1638] Marsha Chechik, Benet Devereux, Steve Easterbrook, Albert Y. C. Lai, and Victor Petrovykh. Efficient multiple-valued model-checking using lattice representations. Lecture Notes in Computer Science, 2154:441– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540441.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540441.
pdf.

Wang:2001:DCS

[1639] Chao Wang, Roderick Bloem, Gary D.

Hachtel, Kavita Ravi, and Fabio Somenzi. Divide and compose: SCC refinement for language emptiness. Lecture Notes in Computer Science, 2154: 456–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540456.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540456.
pdf.

Duflot:2001:UCP

[1640] Marie Duflot, Laurent Fribourg, and Ulf Nilsson. Unavoidable configurations of parameterized rings of processes. Lecture Notes in Computer Science, 2154:472–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2154/21540472.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2154/21540472.pdf.

Chen:2001:LGS

[1641] Yifeng Chen and J. W. Sanders.

Logic of global synchrony. Lecture Notes in Computer Science, 2154:487– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540487.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540487.
                                        pdf.

Baldan:2001:CMR

[1642] P. Baldan, A. Corradini, H. Ehrig, and R. Heckel. Compositional modeling of reactive systems using open nets. Lecture Notes in Computer Science, 2154: 502–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540502.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540502.
                                        pdf.

Kupferman:2001:ETL

[1643] Orna Kupferman, Nir Piterman, and Moshe Y. Vardi. Extended temporal logic revisited. Lecture Notes in Computer Science, 2154:519–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2154/21540519.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2154/21540519.
                                        pdf.

deAlfaro:2001:SAI

[1644] Luca de Alfaro, Thomas A. Henzinger, and Rupak Majumdar. Symbolic algorithms for infinite-state games. Lecture Notes in Computer Science, 2154: 536–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540536.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540536.
pdf.

Kremer:2001:GBV

[1645] Steve Kremer and Jean-François Raskin. A game-based verification of non-repudiation and fair exchange protocols. Lecture Notes in Computer Science, 2154:551–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540551.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540551.
pdf.

deAlfaro:2001:CSS

[1646] Luca de Alfaro, Thomas A. Henzinger, and Freddy Y. C. Mang. The control of synchronous systems, Part II. Lecture Notes in Computer Science, 2154: 566–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2154/21540566.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2154/21540566.
pdf.

Anonymous:2001:AIgp

[1647] Anonymous. Author index. Lecture Notes in Computer Science, 2154:583– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2154/2154auth.pdf.

Beun:2001:MCC

[1648] Robbert-Jan Beun and Harry Bunt. Multimodal cooperative communication. Lecture Notes in Computer Science, 2155:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2155/21550001.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2155/21550001.pdf.

Scott:2001:GTD

[1649] Donia Scott and Richard Power. Generating textual diagrams and diagrammatic texts. Lecture Notes in Computer Science, 2155:13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2155/21550013.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2155/21550013.
                                        pdf.

vanMulken:2001:PAP

[1650] Susanne van Mulken. Pedro: Assessing presentation decodability on the basis of empirically validated models. Lecture Notes in Computer Science, 2155:30–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2155/21550030.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2155/21550030.
pdf.

Zhou:2001:MAG

[1651] Michelle X. Zhou and Steven K. Feiner.

Improvise: Automated generation of animated graphics for coordinated multimedia presentations. Lecture Notes in Computer Science, 2155:43– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2155/21550043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2155/21550043.
pdf.

Beun:2001:MRO

[1652] Robbert-Jan Beun and Anita Cremers. Multimodal reference to objects: An empirical approach. Lecture Notes in Computer Science, 2155:64– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2155/21550064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2155/21550064.
pdf.

Stock:2001:AES

[1653] Oliviero Stock, Carlo Strapparava, and Massimo Zancanaro. Augmenting and executing SharedPlans for multimodal communication. Lecture Notes in Computer Science, 2155:89– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2155/21550089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2155/21550089.
                                        pdf.

Allwood:2001:CFM

[1654] Jens Allwood. Cooperation and flexibility in multimodal communication. Lecture Notes in Computer Science, 2155:113–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2155/21550113.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2155/21550113.
                                        pdf.

Hurault-Plantet:2001:CMA

[1655] Martine Hurault-Plantet and Cecile Balkanski. Communication and manipulation acts in a collaborative dialogue model. Lecture Notes in Computer Science, 2155:125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2155/21550125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2155/21550125.
                                        pdf.

Piwek:2001:RIA

[1656] Paul Piwek. Relating imperatives to action. Lecture Notes in Computer Science, 2155:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2155/21550140.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2155/21550140.
pdf.

Pouteau:2001:IGS

[1657] Xavier Pouteau. Interpretation of gestures and speech: A practical approach to multimodal communication. Lecture Notes in Computer Science, 2155:159– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2155/21550159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2155/21550159.
pdf.

Streit:2001:WMS

[1658] Michael Streit. Why are multimodal systems so difficult to build? — about the difference between deictic gestures and direct manipulation. Lecture Notes in Computer Science, 2155: 176–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2155/21550176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2155/21550176.
pdf.

Kievit:2001:MCR

[1659] Leen Kievit, Paul Piwek, Robbert-Jan Beun, and Harry Bunt. Multimodal cooperative resolution of referential expressions in the DenK system. Lecture Notes in Computer Science, 2155:197– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2155/21550197.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2155/21550197.
                                        pdf.

Brondsted:2001:IWE

[1660] Tom Brøndsted, Paul Dalsgaard, Lars Bo Larsen, Michael Manthey, Paul Mc Kevitt, Thomas B. Moeslund, and Kristian G. Olesen. The Intelli-Media WorkBench — an environment for building multimodal systems. Lecture Notes in Computer Science, 2155: 217–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2155/21550217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2155/21550217.
                                        pdf.

Cheyer:2001:UFC

[1661] Adam Cheyer, Luc Julia, and Jean-Claude Martin. A unified framework for constructing multimodal experiments and applications. Lecture Notes in Computer Science, 2155:234– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2155/21550234.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2155/21550234.
                                        pdf.

Anonymous:2001:Io

[1662] Anonymous. Index. Lecture Notes in Computer Science, 2155:243–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
papers/2155/21550243.pdf.

Anonymous:2001:AIgq

[1663] Anonymous. Author index. Lecture Notes in Computer Science, 2155:251– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2155/21550243.pdf.

Laoutaris:2001:DOP

[1664] N. Laoutaris, G. Boukeas, and I. Stavrakakis. Design of optimal playout schedulers for packet video receivers. Lecture Notes in Computer Science, 2156:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560001.
pdf.

Sisalem:2001:CTF

[1665] D. Sisalem and A. Wolisz. Constrained TCP-friendly congestion control for multimedia communication. Lecture Notes in Computer Science, 2156:17– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560017.
pdf.

Kassler:2001:AWV

[1666] A. Kassler, C. Kücherer, and A. Schrader. Adaptive wavelet video filtering. Lecture Notes in Computer Science, 2156: 32–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560032.
pdf.

Altman:2001:UFM

[1667] E. Altman, C. Barakat, and V. M. Ramos R. On the utility of FEC mechanisms for audio applications. Lecture Notes in Computer Science, 2156:45– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560045.
pdf.

Szabo:2001:CAC

[1668] I. Szabó. A call admission control method for supporting telephony sessions in a best effort IP network. Lecture Notes in Computer Science, 2156: 57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560057.
pdf.

Benameur:2001:IAC

[1669] N. Benameur, S. Ben Fredj, F. Delcoigne, S. Oueslati-Boulahia, and J. W. Roberts. Integrated admission control for streaming and elastic traffic. Lecture Notes in Computer Science, 2156: 69–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560069.
pdf.

Marquetant:2001:NEL

[1670] A. Marquetant, O. Pop, R. Szabó, G. Dinnyés, and Z. Turányi. Novel enhancements to load control — A soft-state, lightweight admission control protocol. Lecture Notes in Computer Science, 2156:82–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560082.
pdf.

Ivars:2001:PPB

[1671] I. M. Ivars and G. Karlsson. PBAC:

Probe-Based Admission Control. Lecture Notes in Computer Science, 2156: 97–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560097.
pdf.

Kuipers:2001:QRA

[1672] F. A. Kuipers and P. Van Mieghem. QoS routing: Average complexity and hopcount in m dimensions. Lecture Notes in Computer Science, 2156:110– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2156/21560110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2156/21560110.
                                        pdf.

Levendovszky:2001:QRI

[1673] J. Levendovszky, A. Fancsali, C. Végsö, and G. Rétvári. QoS routing with incomplete information by analog computing algorithms. Lecture Notes in Computer Science, 2156:127–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2156/21560127.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2156/21560127.
                                        pdf.

Suri:2001:PBR

[1674] S. Suri, M. Waldvogel, and P. R. Warkhede. Profile-based routing: A new framework for MPLS traffic engineering. Lecture Notes in Computer Science, 2156:138–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2156/21560138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2156/21560138.
                                        pdf.

Bless:2001:TBS

[1675] R. Bless, D. Holzhausen, and K. Wehrle.

Towards better support of transaction oriented communication in differentiated services networks. Lecture Notes in Computer Science, 2156:158– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560158.
pdf.

Banchs:2001:OSM

[1676] A. Banchs, O. Leon, and S. Sallent. The Olympic service model: Issues and architecture. Lecture Notes in Computer Science, 2156:170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560170.
pdf.

Siris:2001:SDE

[1677] V. A. Siris, C. Courcoubetis, and G. Margetis. Service differentiation in ECN networks using weighted windowbased congestion control for various packet marking algorithms. Lecture Notes in Computer Science, 2156:190– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560190.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2156/21560190.
                                        pdf.

Lee:2001:TDM

[1678] S.-H. Lee, S.-J. Seok, S.-J. Lee, S.- K. Youm, and C.-H. Kang. Twodifferentiated marking strategies for TCP flows in a differentiated services network. Lecture Notes in Computer Science, 2156:207–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2156/21560207.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2156/21560207.
                                        pdf.

Cho:2001:AAB

[1679] K. Cho, R. Kaizaki, and A. Kato. Aguri: An aggregation-based traffic profiler. Lecture Notes in Computer Science, 2156:222–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2156/21560222.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2156/21560222.pdf.

Bak:2001:THA

[1680] A. Bak, W. Burakowski, F. Ricciato, S. Salsano, and H. Tarasiuk. Traffic handling in AQUILA QoS IP network. Lecture Notes in Computer Science, 2156:243–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2156/21560243.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2156/21560243.
pdf.

Savoric:2001:TCB

[1681] M. Savoric. The TCP control block interdependence in fixed networks — some performance results. Lecture Notes in Computer Science, 2156:261– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560261.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560261.
pdf.

Ryu:2001:ASD

[1682] S.-H. Ryu, J.-Y. Kim, and J. W.-K.

Hong. Approaches to support differentiated quality of Web service. Lecture Notes in Computer Science, 2156: 273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2156/21560273.
pdf.

Uhlig:2001:ULT

[1683] S. Uhlig and O. Bonaventure. Understanding the long-term self-similarity of Internet traffic. Lecture Notes in Computer Science, 2156:286–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2156/21560286.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2156/21560286.
                                        pdf.

Cardoso:2001:NDM

[1684] L. Cardoso. Network dimensioning with MPLS. Lecture Notes in Computer Science, 2156:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2156/21560299.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2156/21560299.
                                        pdf.

Urvoy-Keller:2001:DDS

[1685] G. Urvoy-Keller and E. W. Biersack. DSS: A deterministic and scalable QoS provisioning scheme. Lecture Notes in Computer Science, 2156:310– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2156/21560310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2156/21560310.
                                        pdf.

Akyildiz:2001:EIQ

[1686] I. F. Akyildiz. Experience with an IP QoS physical testbed: Problems and research issues. Lecture Notes in Computer Science, 2156:324–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2156/21560324.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2156/21560324.
                                        pdf.

Zander:2001:AQF

[1687] J. Zander. Affordable QoS in future wireless networks: Myth or reality? Lecture Notes in Computer Science, 2156:325–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2156/21560325.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2156/21560325.
pdf.

Anonymous:2001:AIgr

[1688] Anonymous. Author index. Lecture Notes in Computer Science, 2156: ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2156/2156auth.pdf.

Badea:2001:ROT

[1689] Liviu Badea. A refinement operator for theories. Lecture Notes in Computer Science, 2157:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570001.
pdf.

Basilio:2001:LLP

[1690] Rodrigo Basilio, Gerson Zaverucha, and Valmir C. Barbosa. Learning logic programs with neural networks. Lecture Notes in Computer Science, 2157: 15–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570015.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570015.
pdf.

Braud:2001:GAP

[1691] Agnès Braud and Christel Vrain. A genetic algorithm for propositionalization. Lecture Notes in Computer Science, 2157:27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2157/21570027.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2157/21570027.pdf.

Eineborg:2001:CUE

[1692] Martin Eineborg and Henrik Boström. Classifying uncovered examples by rule stretching. Lecture Notes in Computer Science, 2157:41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570041.
pdf.

Hoche:2001:RLU

[1693] Susanne Hoche and Stefan Wrobel. Relational learning using constrained confidence-rated boosting. Lecture Notes in Computer Science, 2157:51– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2157/21570051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570051.
pdf.

Inoue:2001:IAC

[1694] Katsumi Inoue. Induction, abduction, and consequence-finding. Lecture Notes in Computer Science, 2157:65– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570065.
pdf.

Jacobs:2001:SLS

[1695] Nico Jacobs and Hendrik Blockeel.

From shell logs to shell scripts. Lecture Notes in Computer Science, 2157: 80–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570080.
pdf.

Karwath:2001:AIS

[1696] Andreas Karwath and Ross D. King.

An automated ILP server in the field of bioinformatics. Lecture Notes in Computer Science, 2157:91–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570091.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2157/21570091.
                                        pdf.

Kersting:2001:ABL

[1697] Kristian Kersting and Luc De Raedt. Adaptive Bayesian logic programs. Lecture Notes in Computer Science, 2157: 104–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2157/21570104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2157/21570104.
                                        pdf.

Kersting:2001:TCI

[1698] Kristian Kersting and Luc De Raedt. Towards combining inductive logic programming with Bayesian networks. Lecture Notes in Computer Science, 2157:118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2157/21570118.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2157/21570118.
                                        pdf.

Kramer:2001:DDC

[1699] Stefan Kramer. Demand-driven construction of structural features in ILP. Lecture Notes in Computer Science, 2157:132–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2157/21570132.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2157/21570132.
                                        pdf.

Krogel:2001:TBL

[1700] Mark-A. Krogel and Stefan Wrobel.

Transformation-based learning using multirelational aggregation. Lecture Notes in Computer Science, 2157:142– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570142.
pdf.

Malerba:2001:DAB

[1701] Donato Malerba and Francesca A. Lisi.

Discovering associations between spatial objects: An ILP application. Lecture Notes in Computer Science, 2157: 156–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570156.
pdf.

Maloberti:2001:SCS

[1702] Jérôme Maloberti and Michèle Sebag.

θ-subsumption in a constraint satisfaction perspective. Lecture Notes in Computer Science, 2157:164–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570164.
pdf.

Nepil:2001:LPT

[1703] Miloslav Nepil. Learning to parse from a treebank: Combining TBL and ILP. Lecture Notes in Computer Science, 2157:179–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2157/21570179.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2157/21570179.
                                        pdf.

Otero:2001:ISM

[1704] Ramón P. Otero. Induction of stable models. Lecture Notes in Computer Science, 2157:193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2157/21570193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2157/21570193.
                                        pdf.

Ozaki:2001:APT

[1705] Tomonobu Ozaki and Koichi Furukawa. Application of pruning techniques for propositional learning to progol. Lecture Notes in Computer Science, 2157:206–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2157/21570206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2157/21570206.
                                        pdf.

Quiniou:2001:AIC

[1706] René Quiniou, Marie-Odile Cordier, Guy Carrault, and Feng Wang. Application of ILP to cardiac arrhythmia characterization for chronicle recognition. Lecture Notes in Computer Science, 2157:220–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570220.
pdf.

Struyf:2001:ECV

[1707] Jan Struyf and Hendrik Blockeel. Efficient cross-validation in ILP. Lecture Notes in Computer Science, 2157:228– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2157/21570228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2157/21570228.
pdf.

Yamamoto:2001:MSS

[1708] Akihiro Yamamoto, Kimihito Ito, Akira Ishino, and Hiroki Arimura. Modelling semi-structured documents with hedges for deduction and induction. Lecture Notes in Computer Science, 2157:240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2157/21570240.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2157/21570240.pdf.

Zelezny:2001:LFI

[1709] Filip Železný. Learning functions from imperfect positive data. Lecture Notes in Computer Science, 2157:248– 259, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2157/21570248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2157/21570248.
                                        pdf.

Anonymous:2001:AIgs

[1710] Anonymous. Author index. Lecture Notes in Computer Science, 2157:261– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2157/2157auth.pdf.

McAuley:2001:QMW

[1711] Derek McAuley. QoS for multimedia — what’s going to make it pay? Lecture Notes in Computer Science, 2158: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580001.
                                        pdf.

Roscoe:2001:NRC

[1712] Timothy Roscoe and Prashant Shenoy. New resource control issues in shared clusters. Lecture Notes in Computer Science, 2158:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2158/21580002.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2158/21580002.pdf.

Ott:2001:TLP

[1713] David E. Ott and Ketan Mayer-Patel.

Transport-level protocol coordination in cluster-to-cluster applications. Lecture Notes in Computer Science, 2158: 10–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580010.
pdf.

Cortes:2001:DPM

[1714] Mauricio Cortes and J. Robert Ensor. Data to the people — it’s a matter of control. Lecture Notes in Computer Science, 2158:23–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580023.
pdf.

Schmid:2001:ACA

[1715] Stefan Schmid, Joe Finney, Maomao Wu, Adrian Friday, Andrew Scott, and Doug Shepherd. An access control architecture for metropolitan area wireless networks. Lecture Notes in Computer Science, 2158:29– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580029.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580029.
                                        pdf.

On:2001:DIQ

[1716] Giwon On, Jens B. Schmitt, and Ralf Steinmetz. Design and implementation of a QoS-aware replication mechanism for a distributed multimedia system. Lecture Notes in Computer Science, 2158:38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580038.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580038.
                                        pdf.

Segarra:2001:DVD

[1717] Juan Segarra and Vicent Cholvi. Distribution of video-on-demand in residential networks. Lecture Notes in Computer Science, 2158:50–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580050.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580050.
                                        pdf.

Song:2001:QNS

[1718] Minseok Song, Heonshik Shin, and Naehyuck Chang. A QoS negotiation scheme for efficient failure recovery in multi-resolution video servers. Lecture Notes in Computer Science, 2158:62– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580062.
pdf.

Holmes:2001:THD

[1719] Peter Holmes, Lars Aarhus, and Eirik Maus. Tolerance of highly degraded network conditions for an H.323-based VoIP service. Lecture Notes in Computer Science, 2158:74– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580074.
pdf.

Garcia:2001:CIE

[1720] Fabien Garcia, Christophe Chassot, Andre Lozes, Michel Diaz, Pascal Anelli, and Emmanuel Lochin. Conception, implementation, and evaluation of a QoS-based architecture for an IP environment supporting differentiated services. Lecture Notes in Computer Science, 2158:86–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580086.
pdf.

Cotroneo:2001:SDS

[1721] Domenico Cotroneo, Massimo Ficco, and Giorgio Ventre. A service differentiation scheme for the end-system. Lecture Notes in Computer Science, 2158: 99–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580099.
                                        pdf.

Hofmann:2001:EIP

[1722] Markus Hofmann. Enabling the Internet to provide multimedia services. Lecture Notes in Computer Science, 2158:110–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2158/21580110.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580110.
                                        pdf.

Fitzpatrick:2001:DAT

[1723] Tom Fitzpatrick, Julian Gallop, Gordon Blair, Christopher Cooper, Geoff Coulson, David Duce, and Ian Johnson. Design and application of TOAST: An adaptive distributed multimedia middleware platform. Lecture Notes in Computer Science, 2158:111– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580111.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580111.
                                        pdf.

Ecklund:2001:QMM

[1724] Denise Ecklund, Vera Goebel, Thomas Plagemann, Earl F. Ecklund Jr., Carsten Griwodz, Jan Øyvind Aagedal, Ketil Lund, and Arne-Jørgen Berre. QoS management middleware: A separable, reusable solution. Lecture Notes in Computer Science, 2158:124– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580124.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580124.
pdf.

Orvalho:2001:STM

[1725] João Orvalho, Pedro Ferreira, and Fernando Boavida. State transmission mechanisms for a collaborative virtual environment middleware platform. Lecture Notes in Computer Science, 2158: 138–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580138.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580138.
pdf.

Khayat:2001:SFT

[1726] Ibtissam El Khayat and Guy Leduc.

A stable and flexible TCP-friendly congestion control protocol for layered multicast transmission. Lecture Notes in Computer Science, 2158:154– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580154.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580154.
                                        pdf.

Miras:2001:CAQ

[1727] Dimitrios Miras, Richard J. Jacobs, and Vicky Hardman. Content-aware quality adaptation for IP sessions with multiple streams. Lecture Notes in Computer Science, 2158:168–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580168.
                                        pdf.

Li:2001:MBR

[1728] Kang Li, Charles Krasic, Jonathan Walpole, Molly H. Shor, and Calton Pu. The minimal buffering requirements of congestion controlled interactive multimedia applications. Lecture Notes in Computer Science, 2158:181– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580181.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580181.
                                        pdf.

Takahashi:2001:VCM

[1729] Nozomu Takahashi, Yuki Wakita, Shigeki Ouchi, and Takayuki Kunieda. Video content management using logical content. Lecture Notes in Computer Science, 2158:193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580193.
pdf.

Sugar:2001:ACU

[1730] Robert Sugar and Sandor Imre. Adaptive clustering using mobile agents in wireless ad-hoc networks. Lecture Notes in Computer Science, 2158:199– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580199.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580199.
pdf.

Finney:2001:MNM

[1731] Joe Finney and Greg O’Shea. Mobile 4-in-6: A novel mechanism for IPv4/v6 transitioning. Lecture Notes in Computer Science, 2158:205–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580205.
pdf.

Krasic:2001:CSM

[1732] Charles Krasic, Kang Li, and Jonathan Walpole. The case for streaming multimedia with TCP. Lecture Notes in Computer Science, 2158:213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580213.
                                        pdf.

Pereira:2001:MSW

[1733] Fernando Pereira. The MPEG-21 standard: Why an open multimedia framework? Lecture Notes in Computer Science, 2158:219–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580219.
                                        pdf.

El-Khatib:2001:SQP

[1734] Khalil El-Khatib, Gregor v. Bochmann, and Yu Zhong. Selecting the QoS parameters for multicast applications based on user profile and device capability. Lecture Notes in Computer Science, 2158:221–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2158/21580221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2158/21580221.
                                        pdf.

Meissner:2001:NGI

[1735] Andreas Meissner, Lars Wolf, and Ralf Steinmetz. A novel group integrity concept for multimedia multicasting. Lecture Notes in Computer Science, 2158: 233–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580233.
pdf.

Krishnakumar:2001:CBC

[1736] Krish T. Krishnakumar and Morris Sloman. Constraint-based configuration of proxylets for programmable networks. Lecture Notes in Computer Science, 2158:245–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2158/21580245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2158/21580245.
pdf.

Anonymous:2001:AIgt

[1737] Anonymous. Author index. Lecture Notes in Computer Science, 2158:257– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2158/2158indx.pdf.

Adamatzky:2001:CNM

[1738] Andrew Adamatzky. Computing in nonlinear media: Make waves, study collisions. Lecture Notes in Computer Science, 2159:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590001.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2159/21590001.
                                        pdf.

Dorigo:2001:AAS

[1739] Marco Dorigo. Ant algorithms solve difficult optimization problems. Lecture Notes in Computer Science, 2159: 11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590011.
                                        pdf.

Husbands:2001:SNV

[1740] Phil Husbands, Andy Philippides, Tom Smith, and Michael O’Shea. The shifting network: Volume signalling in real and robot nervous systems. Lecture Notes in Computer Science, 2159:23– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590023.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590023.
                                        pdf.

Kvasnicka:2001:SRH

[1741] V. Kvasnička, J. Pospíchal, and T. Kaláb. A study of replicators and hypercycles by typogenetics. Lecture Notes in Computer Science, 2159:37– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590037.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2159/21590037.
pdf.

Wiedermann:2001:EST

[1742] Jiří Wiedermann and Jan van Leeuwen.

Emergence of a super-Turing computational potential in artificial living systems. Lecture Notes in Computer Science, 2159:55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590055.
pdf.

Anchorena:2001:EGM

[1743] Sergio O. Anchorena and Blanca Cases. Eco-grammars to model biological systems: Adding probabilities to agents. Lecture Notes in Computer Science, 2159:66–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590066.
pdf.

Bedau:2001:DEA

[1744] Mark A. Bedau. Dynamics of the environment for adaptation in static resource models. Lecture Notes in Computer Science, 2159:76–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590076.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590076.
                                        pdf.

Bellas:2001:ABT

[1745] F. Bellas, A. Lamas, and R. J. Duro. Adaptive behavior through a darwinist machine. Lecture Notes in Computer Science, 2159:86–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590086.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590086.
                                        pdf.

Bunzli:2001:FTS

[1746] Daniel C. Bünzli and Mathieu S. Capcarrere. Fault-tolerant structures: Towards robust self-replication in a probabilistic environment. Lecture Notes in Computer Science, 2159:90– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590090.
                                        pdf.

Dittrich:2001:SUS

[1747] Peter Dittrich and Wolfgang Banzhaf. Survival of the unfittest? — the seceder model and its fitness landscape. Lecture Notes in Computer Science, 2159: 100–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590100.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2159/21590100.
pdf.

Glotzmann:2001:EMA

[1748] T. Glotzmann, H. Lange, M. Hauhs, and A. Lamm. Evolving multi-agent networks in structured environments. Lecture Notes in Computer Science, 2159:110–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2159/21590110.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2159/21590110.
pdf.

Mascaro:2001:SES

[1749] Steven Mascaro, Kevin B. Korb, and Ann E. Nicholson. Suicide as an evolutionarily stable strategy. Lecture Notes in Computer Science, 2159:120– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590120.
pdf.

McCormack:2001:EES

[1750] Jon McCormack. Eden: An evolutionary sonic ecosystem. Lecture Notes in Computer Science, 2159:133– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590133.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2159/21590133.
                                        pdf.

Kadlecek:2001:NHA

[1751] David Kadleček and Pavel Nahodil. New hybrid architecture in artificial life simulation. Lecture Notes in Computer Science, 2159:143–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590143.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590143.
                                        pdf.

Kadukova:2001:SEH

[1752] Jana Kadukova, Marian Mihalo, Martin Palko, and Marek Polak. “in silico” experiments on heavy metal sorption by algal biomass. Lecture Notes in Computer Science, 2159:147– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590147.
                                        pdf.

Seth:2001:SEM

[1753] Anil K. Seth. Spatially explicit models of forager interference. Lecture Notes in Computer Science, 2159:151– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590151.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2159/21590151.
pdf.

Stow:2001:AAW

[1754] Daniel Stow and Chris Roadknight.

Antigens, antibodies, and the World Wide Web. Lecture Notes in Computer Science, 2159:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2159/21590161.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2159/21590161.pdf.

Todd:2001:LWK

[1755] Peter M. Todd and Simon Kirby. I like what I know: How recognitionbased decisions can structure the environment. Lecture Notes in Computer Science, 2159:166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2159/21590166.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2159/21590166.pdf.

Baranda:2001:BLC

[1756] Angel V. Baranda, Juan Castellanos, Fernando Arroyo, and Carmen Luengo. Bio-language for computing with membranes. Lecture Notes in Computer Science, 2159:176–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590176.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2159/21590176.
pdf.

Ono:2001:ACC

[1757] Naoaki Ono and Takashi Ikegami. Artificial chemistry: Computational studies on the emergence of selfreproducing units. Lecture Notes in Computer Science, 2159:186–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590186.
pdf.

SperonidiFenizio:2001:SMB

[1758] Pietro Speroni di Fenizio and Wolfgang Banzhaf. Stability of metabolic and balanced organisations. Lecture Notes in Computer Science, 2159:196– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590196.
pdf.

SperonidiFenizio:2001:SFP

[1759] Pietro Speroni di Fenizio, Peter Dittrich, and Wolfgang Banzhaf. Spontaneous formation of proto-cells in an universal artificial chemistry on a planar graph. Lecture Notes in Computer Science, 2159:206–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2159/21590206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590206.
pdf.

Aharonov:2001:UAB

[1760] Ranit Aharonov, Isaac Meilijson, and Eytan Ruppin. Understanding the agent’s brain: A quantitative approach. Lecture Notes in Computer Science, 2159:216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590216.
pdf.

Bilotta:2001:OCM

[1761] Eleonora Bilotta and Pietro Pantano. Observations on complex multi-state CAs. Lecture Notes in Computer Science, 2159:226–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590226.
pdf.

Kelemen:2001:ANN

[1762] Eduard Kelemen. Artificial neural networks and artificial evolution as tools to study organization of spatial behavior in rats. Lecture Notes in Computer Science, 2159:236–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590236.
                                        pdf.

Kim:2001:TGF

[1763] Jan T. Kim. transsys: A generic formalism for modelling regulatory networks in morphogenesis. Lecture Notes in Computer Science, 2159:242– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590242.
                                        pdf.

Niv:2001:ERL

[1764] Yael Niv, Daphna Joel, Isaac Meilijson, and Eytan Ruppin. Evolution of reinforcement learning in uncertain environments: Emergence of risk-aversion and matching. Lecture Notes in Computer Science, 2159:252– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590252.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590252.
                                        pdf.

Oliveira:2001:SOD

[1765] Gina M. B. Oliveira, Pedro P. B. de Oliveira, and N. Omar. Searching for one-dimensional cellular automata in the absence of a priori information. Lecture Notes in Computer Science, 2159:262–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590262.
pdf.

Smith:2001:NNE

[1766] Tom Smith, Phil Husbands, and Michael O’Shea. Neutral networks and evolvability with complex genotypephenotype mapping. Lecture Notes in Computer Science, 2159:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590272.
pdf.

Stauffer:2001:ECD

[1767] André Stauffer and Moshe Sipper.

Externally controllable and destructible self-replicating loops. Lecture Notes in Computer Science, 2159:282– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590282.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590282.
pdf.

Tokura:2001:ENA

[1768] Seiji Tokura, Akio Ishiguro, Hiroki Kawai, and Peter Eggenberger. The effect of neuromodulations on the adaptability of evolved neurocontrollers. Lecture Notes in Computer Science, 2159: 292–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590292.
                                        pdf.

Balanescu:2001:TCA

[1769] Tudor Bălănescu, Marian Gheorghe, Mike Holcombe, and Florentin Ipate. Testing collaborative agents defined as stream X-machines with distributed grammars. Lecture Notes in Computer Science, 2159:296–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590296.
                                        pdf.

Ebner:2001:TDE

[1770] Marc Ebner. A three-dimensional environment for self-reproducing programs. Lecture Notes in Computer Science, 2159:306–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2159/21590306.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590306.
                                        pdf.

Ficici:2001:POC

[1771] Sevan G. Ficici and Jordan B. Pollack. Pareto optimality in coevolutionary learning. Lecture Notes in Computer Science, 2159:316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590316.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590316.
pdf.

Kubik:2001:EEM

[1772] Aleš Kubík. On emergence in evolutionary multiagent systems. Lecture Notes in Computer Science, 2159:326– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590326.
pdf.

Larsen:2001:DLS

[1773] Jesper Bach Larsen. Division of labour in simulated ant colonies under spatial constraints. Lecture Notes in Computer Science, 2159:338–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590338.
pdf.

Marshall:2001:EOC

[1774] I. W. Marshall and C. M. Roadknight. Emergent organisation in colonies of simple automata. Lecture Notes in Computer Science, 2159:349– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590349.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590349.
                                        pdf.

Quinn:2001:ECD

[1775] Matt Quinn. Evolving communication without dedicated communication channels. Lecture Notes in Computer Science, 2159:357–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590357.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590357.
                                        pdf.

Quinn:2001:MAB

[1776] Matt Quinn and Jason Noble. Modelling animal behaviour in contests: Conventions for resource allocation. Lecture Notes in Computer Science, 2159:367–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2159/21590367.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590367.
                                        pdf.

Simao:2001:MHM

[1777] Jorge Simão and Peter M. Todd. A model of human mate choice with courtship that predicts population patterns. Lecture Notes in Computer Science, 2159:377–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590377.
pdf.

Smith:2001:ECS

[1778] Andrew D. M. Smith. Establishing communication systems without explicit meaning transmission. Lecture Notes in Computer Science, 2159:381– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590381.
pdf.

Yamauchi:2001:DBA

[1779] Hajime Yamauchi. The difficulty of the baldwinian account of linguistic innateness. Lecture Notes in Computer Science, 2159:391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590391.
pdf.

Bongard:2001:MEO

[1780] Josh C. Bongard and Chandana Paul.

Making evolution an offer it can’t refuse: Morphology and the extradimensional bypass. Lecture Notes in Computer Science, 2159:401–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590401.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590401.
                                        pdf.

Burtsev:2001:MEE

[1781] Mikhail S. Burtsev, Vladimir G. Red’ko, and Roman V. Gusarev. Model of evolutionary emergence of purposeful adaptive behavior. the role of motivation. Lecture Notes in Computer Science, 2159:413–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590413.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590413.
                                        pdf.

Channon:2001:PAT

[1782] Alastair Channon. Passing the ALife test: Activity statistics classify evolution in geb as unbounded. Lecture Notes in Computer Science, 2159:417– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590417.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590417.
                                        pdf.

Jones:2001:EAC

[1783] Stephen Jones. On the evolution of artificial consciousness. Lecture Notes in Computer Science, 2159:427– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590427.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590427.
pdf.

Jung:2001:SEI

[1784] Tobias Jung, Peter Dauscher, and Thomas Uthmann. Some effects of individual learning on the evolution of sensors. Lecture Notes in Computer Science, 2159:432–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590432.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590432.
pdf.

Lenaerts:2001:TSE

[1785] T. Lenaerts, A. Defaweux, P. Beyens, and B. Manderick. Transitions in a simple evolutionary model. Lecture Notes in Computer Science, 2159:436– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590436.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590436.
pdf.

McMullin:2001:TIE

[1786] Barry McMullin and Dominique Groß.

Towards the implementation of evolving autopoietic artificial agents. Lecture Notes in Computer Science, 2159: 440–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590440.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590440.
                                        pdf.

Sato:2001:VTT

[1787] Yuji Sato, Masafumi Kawamoto, and Motoi Chizaki. Verification of text transcription history by using evolutionary algorithms. Lecture Notes in Computer Science, 2159:444–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590444.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590444.
                                        pdf.

Schwardy:2001:GAR

[1788] Eva Schwardy. Genetic algorithm as a result of phenomenological reduction of natural evolution. Lecture Notes in Computer Science, 2159:454– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590454.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590454.
                                        pdf.

Suzuki:2001:SRG

[1789] Hideaki Suzuki. String rewriting grammar optimized using an evolvability measure. Lecture Notes in Computer Science, 2159:458–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590458.
pdf.

Vardy:2001:VBE

[1790] Andrew Vardy. A visually-based evolvable control architecture for agents in interactive entertainment applications. Lecture Notes in Computer Science, 2159:469–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2159/21590469.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2159/21590469.
pdf.

Watson:2001:SCE

[1791] Richard A. Watson and Jordan B.

Pollack. Symbiotic composition and evolvability. Lecture Notes in Computer Science, 2159:480–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590480.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590480.
pdf.

Balaam:2001:N

[1792] Andy Balaam. nBrains. Lecture Notes in Computer Science, 2159:491– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2159/21590491.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590491.
                                        pdf.

Bird:2001:CWR

[1793] Jon Bird. Can wheeled robots illuminate adaptive behaviour? Lecture Notes in Computer Science, 2159:495– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590495.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590495.
                                        pdf.

Dahl:2001:EAB

[1794] T. S. Dahl and C. Giraud-Carrier. Evolution, adaption, and behavioural holism in artificial intelligence. Lecture Notes in Computer Science, 2159: 499–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590499.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590499.
                                        pdf.

Fujii:2001:EBL

[1795] Akinobu Fujii, Akio Ishiguro, Takeshi Aoki, and Peter Eggenberger. Evolving bipedal locomotion with a dynamicallyrearranging neural network. Lecture Notes in Computer Science, 2159:509– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590509.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2159/21590509.
pdf.

Kelly:2001:SRP

[1796] Ian Kelly and Chris Melhuish. Slug-Bot: A robot predator. Lecture Notes in Computer Science, 2159:519– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590519.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590519.
pdf.

Kim:2001:MRC

[1797] DaeEun Kim and John C. T. Hallam.

Mobile robot control based on Boolean logic with internal memory. Lecture Notes in Computer Science, 2159:529– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590529.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590529.
pdf.

Kumada:2001:ECT

[1798] Yoichiro Kumada and Kazuhiro Ueda.

Emergence of cooperative tactics by soccer agents with ability of prediction and learning. Lecture Notes in Computer Science, 2159:539–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590539.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2159/21590539.
                                        pdf.

Melhuish:2001:PSM

[1799] Chris Melhuish, Matt Wilson, and Ana Sendova-Franks. Patch sorting: Multi-object clustering using minimalist robots. Lecture Notes in Computer Science, 2159:543–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590543.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590543.
                                        pdf.

Svatos:2001:BFM

[1800] Václav Svatoš. Behavioural formation management in robotic soccer. Lecture Notes in Computer Science, 2159:553– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590553.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590553.
                                        pdf.

Takahashi:2001:CSF

[1801] Nobuyuki Takahashi, Takashi Nagai, Hiroshi Yokoi, and Yukinori Kakazu. Control system of flexible structure multi-cell robot using amoeboid selforganization mode. Lecture Notes in Computer Science, 2159:563–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590563.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2159/21590563.
pdf.

Wessnitzer:2001:TSO

[1802] Jan Wessnitzer, Andrew Adamatzky, and Chris Melhuish. Towards selforganising structure formations: A decentralized approach. Lecture Notes in Computer Science, 2159:573–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590573.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590573.
pdf.

Breazeal:2001:AIB

[1803] Cynthia Breazeal. Affective interaction between humans and robots. Lecture Notes in Computer Science, 2159:582– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590582.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590582.
pdf.

Brighton:2001:SSS

[1804] Henry Brighton and Simon Kirby.

The survival of the smallest: Stability conditions for the cultural evolution of compositional language. Lecture Notes in Computer Science, 2159: 592–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590592.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590592.
                                        pdf.

Bullock:2001:SOU

[1805] Seth Bullock. Smooth operator? understanding and visualising mutation bias. Lecture Notes in Computer Science, 2159:602–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590602.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590602.
                                        pdf.

Edmonds:2001:ULS

[1806] Ian R. Edmonds. The use of latent semantic indexing to identify evolutionary trajectories in behaviour space. Lecture Notes in Computer Science, 2159:613–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590613.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590613.
                                        pdf.

Mikami:2001:DVM

[1807] Tsuyoshi Mikami and Mitsuo Wada. Data visualization method for growing self-organizing networks with ant clustering algorithm. Lecture Notes in Computer Science, 2159:623–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2159/21590623.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590623.
pdf.

Neumann:2001:IIV

[1808] Titus R. Neumann and Heinrich H.

Bülthoff. Insect inspired visual control of translatory flight. Lecture Notes in Computer Science, 2159:627– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590627.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590627.
pdf.

Smith:2001:IRC

[1809] Kenny Smith. The importance of rapid cultural convergence in the evolution of learned symbolic communication. Lecture Notes in Computer Science, 2159: 637–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590637.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590637.
pdf.

Zuidema:2001:ESU

[1810] Willem H. Zuidema. Emergent syntax: The unremitting value of computational modeling for understanding the origins of complex language. Lecture Notes in Computer Science, 2159: 641–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2159/21590641.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590641.
                                        pdf.

DHondt:2001:AG

[1811] Ellie D’Hondt and Theo D’Hondt. Amorphous geometry. Lecture Notes in Computer Science, 2159:645–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590645.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590645.
                                        pdf.

Paolo:2001:ALH

[1812] Ezequiel A. Di Paolo. Artificial life and historical processes. Lecture Notes in Computer Science, 2159:649– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590649.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2159/21590649.
                                        pdf.

Dorin:2001:AFA

[1813] Alan Dorin. Aesthetic fitness and artificial evolution for the selection of imagery from the mythical infinite library. Lecture Notes in Computer Science, 2159:659–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2159/21590659.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2159/21590659.
pdf.

Humphrys:2001:DMI

[1814] Mark Humphrys. Distributing a mind on the Internet: The World-Wide-Mind. Lecture Notes in Computer Science, 2159:669–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590669.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590669.
pdf.

Jog:2001:DCU

[1815] Ninad Jog. The dimensions of the cyber universe. Lecture Notes in Computer Science, 2159:681–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590681.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590681.
pdf.

Komosinski:2001:TAM

[1816] Maciej Komosinski and Marek Kubiak. Taxonomy in alife. measures of similarity for complex artificial organisms. Lecture Notes in Computer Science, 2159:685–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590685.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2159/21590685.
pdf.

Masumoto:2001:GSA

[1817] Gen Masumoto and Takashi Ikegami. The λ-game system: An approach to a meta-game. Lecture Notes in Computer Science, 2159:695–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590695.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590695.
pdf.

Nomura:2001:FDA

[1818] Tatsuya Nomura. Formal description of autopoiesis based on the theory of category. Lecture Notes in Computer Science, 2159:700–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590700.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590700.
pdf.

Polani:2001:ITA

[1819] Daniel Polani, Thomas Martinetz, and Jan Kim. An informationtheoretic approach for the quantification of relevance. Lecture Notes in Computer Science, 2159:704–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590704.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2159/21590704.
pdf.

Walshe:2001:FIW

[1820] Ray Walshe and Mark Humphrys. First implementation of the World-Wide-Mind. Lecture Notes in Computer Science, 2159:714–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2159/21590714.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2159/21590714.pdf.

Wood:2001:ELI

[1821] Rachel Wood. Evolving lives:

The individual historical dimension in evolution. Lecture Notes in Computer Science, 2159:719–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2159/21590719.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2159/21590719.
pdf.

Anonymous:2001:AIgu

[1822] Anonymous. Author index. Lecture Notes in Computer Science, 2159:723– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2159/2159auth.pdf.

Lu:2001:ORC

[1823] Ruqian Lu and Songmao Zhang. 1 overview of research on computer animation and related topics. Lecture Notes in Computer Science, 2160:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2160/21600001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2160/21600001.
pdf.

Lu:2001:SFL

[1824] Ruqian Lu and Songmao Zhang. 2 SWAN: Full life cycle automation of computer animation. Lecture Notes in Computer Science, 2160:29–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2160/21600029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2160/21600029.
pdf.

Lu:2001:ULC

[1825] Ruqian Lu and Songmao Zhang. 3 understanding the limited Chinese natural language. Lecture Notes in Computer Science, 2160:65–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2160/21600065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2160/21600065.
pdf.

Lu:2001:SUT

[1826] Ruqian Lu and Songmao Zhang. 4 story understanding: The theory. Lecture Notes in Computer Science, 2160: 93–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2160/21600093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2160/21600093.
pdf.

Lu:2001:SUP

[1827] Ruqian Lu and Songmao Zhang. 5 story understanding: The practice. Lecture Notes in Computer Science, 2160:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2160/21600123.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2160/21600123.
pdf.

Lu:2001:PPA

[1828] Ruqian Lu and Songmao Zhang. 6 plot planning and act planning. Lecture Notes in Computer Science, 2160: 173–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2160/21600173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2160/21600173.
pdf.

Lu:2001:DP

[1829] Ruqian Lu and Songmao Zhang. 7 director’s planning. Lecture Notes in Computer Science, 2160:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2160/21600229.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2160/21600229.
                                        pdf.

Lu:2001:CP

[1830] Ruqian Lu and Songmao Zhang. 8 camera planning. Lecture Notes in Computer Science, 2160:255–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2160/21600255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2160/21600255.
                                        pdf.

Lu:2001:LCR

[1831] Ruqian Lu and Songmao Zhang. 9 light, color, and role planning. Lecture Notes in Computer Science, 2160: 287–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2160/21600287.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2160/21600287.
                                        pdf.

Lu:2001:KBL

[1832] Ruqian Lu and Songmao Zhang. 10 knowledge base and libraries. Lecture Notes in Computer Science, 2160:315– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2160/21600315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2160/21600315.
                                        pdf.

Lu:2001:A

[1833] Ruqian Lu and Songmao Zhang. Appendix. Lecture Notes in Computer Science, 2160:351–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2160/21600351.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2160/21600351.pdf.

Lu:2001:R

[1834] Ruqian Lu and Songmao Zhang. References. Lecture Notes in Computer Science, 2160:363–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2160/21600363.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2160/21600363.pdf.

Anonymous:2001:Ip

[1835] Anonymous. Index. Lecture Notes in Computer Science, 2160:375–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2160/2160indx.pdf.

Arge:2001:EMD

[1836] Lars Arge. External memory data structures. Lecture Notes in Computer Science, 2161:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610001.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2161/21610001.
pdf.

Albers:2001:SAP

[1837] Susanne Albers. Some algorithmic problems in large networks. Lecture Notes in Computer Science, 2161:30– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610030.
pdf.

Zwick:2001:EAD

[1838] Uri Zwick. Exact and approximate distances in graphs — A survey. Lecture Notes in Computer Science, 2161:33– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610033.
pdf.

Gaysinsky:2001:SCA

[1839] Alexander Gaysinsky, Alon Itai, and Hadas Shachnai. Strongly competitive algorithms for caching with pipelined prefetching. Lecture Notes in Computer Science, 2161:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610049.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2161/21610049.
pdf.

Hutchinson:2001:DBP

[1840] David A. Hutchinson, Peter Sanders, and Jeffrey Scott Vitter. Duality between prefetching and queued writing with parallel disks. Lecture Notes in Computer Science, 2161:62– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610062.
pdf.

Krumke:2001:OC

[1841] Sven O. Krumke, Willem E. de Paepe, Jörg Rambau, and Leen Stougie. Online bin coloring. Lecture Notes in Computer Science, 2161:74–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610074.
pdf.

Seiden:2001:GDT

[1842] Steven S. Seiden. A general decomposition theorem for the kserver problem. Lecture Notes in Computer Science, 2161:86–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610086.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2161/21610086.
                                        pdf.

Csirik:2001:BCC

[1843] János Csirik, Csanád Imreh, John Noga, Steve S. Seiden, and Gerhard J. Woeginger. Buying a constant competitive ratio for paging. Lecture Notes in Computer Science, 2161:98– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610098.
                                        pdf.

Dietzfelbinger:2001:SMP

[1844] Martin Dietzfelbinger and Torben Hagerup. Simple minimal perfect hashing in less space. Lecture Notes in Computer Science, 2161:109–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610109.
                                        pdf.

Pagh:2001:CH

[1845] Rasmus Pagh and Flemming Friche Rodler. Cuckoo hashing. Lecture Notes in Computer Science, 2161:121– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610121.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2161/21610121.
pdf.

Sellmann:2001:CVF

[1846] Meinolf Sellmann and Torsten Fahle.

Coupling variable fixing algorithms for the automatic recording problem. Lecture Notes in Computer Science, 2161: 134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610134.
pdf.

Lepere:2001:AAS

[1847] Renaud Lepère, Denis Trystram, and Gerhard J. Woeginger. Approximation algorithms for scheduling malleable tasks under precedence constraints. Lecture Notes in Computer Science, 2161:146–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610146.
pdf.

Halldorsson:2001:AMT

[1848] Bjarni V. Halldórsson, Magnús M.

Halldórsson, and R. Ravi. On the approximability of the minimum test collection problem. Lecture Notes in Computer Science, 2161:158–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2161/21610158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610158.
                                        pdf.

Kolpakov:2001:FAR

[1849] Roman Kolpakov and Gregory Kucherov. Finding approximate repetitions under Hamming distance. Lecture Notes in Computer Science, 2161:170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610170.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610170.
                                        pdf.

Lancia:2001:SPC

[1850] Giuseppe Lancia, Vineet Bafna, Sorin Istrail, Ross Lippert, and Russell Schwartz. SNPs problems, complexity, and algorithms. Lecture Notes in Computer Science, 2161:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610182.
                                        pdf.

Angel:2001:FAU

[1851] Eric Angel, Evripidis Bampis, and Alexander Kononov. A FPTAS for approximating the unrelated parallel machines scheduling problem with costs. Lecture Notes in Computer Science, 2161:194–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610194.
pdf.

Fishkin:2001:GTS

[1852] Aleksei V. Fishkin, Klaus Jansen, and Monaldo Mastrolilli. Grouping techniques for scheduling problems: Simpler and faster. Lecture Notes in Computer Science, 2161:206– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610206.
pdf.

Karuno:2001:AAM

[1853] Yoshiyuki Karuno and Hiroshi Nagamochi. A 2-approximation algorithm for the multi-vehicle scheduling problem on a path with release and handling times. Lecture Notes in Computer Science, 2161:218–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610218.
pdf.

Goldberg:2001:SSP

[1854] Andrew V. Goldberg. A simple shortest path algorithm with linear average time. Lecture Notes in Computer Science, 2161:230–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610230.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610230.
                                        pdf.

Mehlhorn:2001:HDA

[1855] Kurt Mehlhorn and Guido Schäfer. A heuristic for Dijkstra’s algorithm with many targets and its use in weighted matching algorithms. Lecture Notes in Computer Science, 2161:242– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610242.
                                        pdf.

Burnikel:2001:SBR

[1856] Christoph Burnikel, Stefan Funke, Kurt Mehlhorn, Stefan Schirra, and Susanne Schmitt. A separation bound for real algebraic expressions. Lecture Notes in Computer Science, 2161:254– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610254.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610254.
                                        pdf.

Czumaj:2001:PTG

[1857] Artur Czumaj and Christian Sohler. Property testing with geometric queries. Lecture Notes in Computer Science, 2161:266–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2161/21610266.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2161/21610266.pdf.

Abellanas:2001:SCS

[1858] Manuel Abellanas, Ferran Hurtado, Christian Icking, Rolf Klein, Elmar Langetepe, Lihong Ma, Belén Palop, and Vera Sacristán. Smallest color-spanning objects. Lecture Notes in Computer Science, 2161:278– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610278.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610278.
pdf.

Radhakrishnan:2001:EDC

[1859] Jaikumar Radhakrishnan, Venkatesh Raman, and S. Srinivasa Rao. Explicit deterministic constructions for membership in the bitprobe model. Lecture Notes in Computer Science, 2161: 290–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610290.
pdf.

Pagh:2001:LD

[1860] Rasmus Pagh and Flemming Friche Rodler. Lossy dictionaries. Lecture Notes in Computer Science, 2161:300– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610300.
                                        pdf.

Chazelle:2001:SDT

[1861] Bernard Chazelle, Olivier Devillers, Ferran Hurtado, Mercè Mora, Vera Sacristán, and Monique Teillaud. Splitting a Delaunay triangulation in linear time. Lecture Notes in Computer Science, 2161:312–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610312.
                                        pdf.

Ebbers-Baumann:2001:FAA

[1862] Annette Ebbers-Baumann, Rolf Klein, Elmar Langetepe, and Andrzej Lingas. A fast algorithm for approximating the detour of a polygonal chain. Lecture Notes in Computer Science, 2161:321– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610321.
                                        pdf.

Eidenbenz:2001:AAM

[1863] Stephan Eidenbenz and Peter Widmayer. An approximation algorithm for MINIMUM CONVEX COVER with logarithmic performance guarantee. Lecture Notes in Computer Science, 2161:333–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610333.
pdf.

Czygrinow:2001:DEC

[1864] A. Czygrinow, M. Hanćkowiak, and M. Karoński. Distributed O(Δ log n)- edge-coloring algorithm. Lecture Notes in Computer Science, 2161:345–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610345.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610345.
pdf.

Douceur:2001:MRP

[1865] John R. Douceur and Roger P. Wattenhofer. Modeling replica placement in a distributed file system: Narrowing the gap between analysis and simulation. Lecture Notes in Computer Science, 2161:356–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610356.
pdf.

Blaser:2001:CCC

[1866] Markus Bläser and Bodo Siebert. Computing cycle covers without short cycles. Lecture Notes in Computer Science, 2161:368–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2161/21610368.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2161/21610368.pdf.

Thilikos:2001:PTA

[1867] Dimitrios M. Thilikos, Maria J. Serna, and Hans L. Bodlaender. A polynomial time algorithm for the cutwidth of bounded degree graphs with small treewidth. Lecture Notes in Computer Science, 2161:380–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610380.
                                        pdf.

Sensen:2001:LBE

[1868] Norbert Sensen. Lower bounds and exact algorithms for the graph partitioning problem using multicommodity flows. Lecture Notes in Computer Science, 2161:391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610391.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610391.
                                        pdf.

Akcoglu:2001:FPE

[1869] Karhan Akcoglu, Ming-Yang Kao, and Shuba V. Raghavan. Fast pricing of European Asian options with provable accuracy: Single-stock and basket options. Lecture Notes in Computer Science, 2161:404–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610404.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610404.
pdf.

Goldberg:2001:CAM

[1870] Andrew V. Goldberg and Jason D.

Hartline. Competitive auctions for multiple digital goods. Lecture Notes in Computer Science, 2161:416–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610416.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610416.
pdf.

Langerman:2001:AEF

[1871] Stefan Langerman, Sachin Lodha, and Rahul Shah. Algorithms for efficient filtering in content-based multicast. Lecture Notes in Computer Science, 2161: 428–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610428.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610428.
pdf.

Fraigniaud:2001:AAM

[1872] Pierre Fraigniaud. Approximation algorithms for minimum-time broadcast under the vertex-disjoint paths mode. Lecture Notes in Computer Science, 2161:440–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2161/21610440.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610440.
                                        pdf.

Clementi:2001:RRO

[1873] Andrea E. F. Clementi, Angelo Monti, and Riccardo Silvestri. Round robin is optimal for fault-tolerant broadcasting on wireless networks. Lecture Notes in Computer Science, 2161:452– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610452.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610452.
                                        pdf.

Fiala:2001:OOD

[1874] Jiří Fiala, Aleksei V. Fishkin, and Fedor V. Fomin. Online and offline distance constrained labeling of disk graphs. Lecture Notes in Computer Science, 2161:464–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610464.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610464.
                                        pdf.

Gavoille:2001:ADL

[1875] Cyril Gavoille, Michal Katz, Nir A.

Katz, Christophe Paul, and David Peleg. Approximate distance labeling schemes. Lecture Notes in Computer Science, 2161:476–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610476.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610476.
pdf.

Dujmovic:2001:PCL

[1876] V. Dujmović, M. Fellows, M. Hallett, M. Kitching, G. Liotta, C. McCartin, N. Nishimura, P. Ragde, F. Rosamond, M. Suderman, S. Whitesides, and D. R. Wood. On the parameterized complexity of layered graph drawing. Lecture Notes in Computer Science, 2161:488– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2161/21610488.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2161/21610488.
pdf.

Cooper:2001:GMU

[1877] Colin Cooper and Alan M. Frieze.

A general model of undirected Web graphs. Lecture Notes in Computer Science, 2161:500–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2161/21610500.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2161/21610500.pdf.

Caprara:2001:PCC

[1878] Alberto Caprara, Alessandro Panconesi, and Romeo Rizzi. Packing cycles and cuts in undirected graphs. Lecture Notes in Computer Science, 2161: 512–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610512.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610512.
                                        pdf.

Zito:2001:GAM

[1879] Michele Zito. Greedy algorithms for minimisation problems in random regular graphs. Lecture Notes in Computer Science, 2161:524–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2161/21610524.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2161/21610524.
                                        pdf.

Anonymous:2001:AIgv

[1880] Anonymous. Author index. Lecture Notes in Computer Science, 2161:537– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2161/2161auth.pdf.

Anderson:2001:PES

[1881] R. Anderson. Protecting embedded systems — the next ten years. Lecture Notes in Computer Science, 2162: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620001.
pdf.

Goubin:2001:SMS

[1882] L. Goubin. A sound method for switching between Boolean and arithmetic masking. Lecture Notes in Computer Science, 2162:3–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620003.
pdf.

Brier:2001:FPI

[1883] E. Brier, H. Handschuh, and C. Tymen. Fast primitives for internal data scrambling in tamper resistant hardware. Lecture Notes in Computer Science, 2162:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620016.
pdf.

May:2001:RRR

[1884] D. May, H. L. Muller, and N. P.

Smart. Random register renaming to foil DPA. Lecture Notes in Computer Science, 2162:28–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620028.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620028.
                                        pdf.

Oswald:2001:RAS

[1885] E. Oswald and M. Aigner. Randomized addition-subtraction chains as a countermeasure against power attacks. Lecture Notes in Computer Science, 2162: 39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620039.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620039.
                                        pdf.

Kuo:2001:AOS

[1886] H. Kuo and I. Verbauwhede. Architectural optimization for a 1.82gbits/ sec VLSI implementation of the AES Rijndael algorithm. Lecture Notes in Computer Science, 2162:51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620051.
                                        pdf.

McLoone:2001:HPS

[1887] M. McLoone and J. V. McCanny. High performance single-chip FPGA Rijndael algorithm implementations. Lecture Notes in Computer Science, 2162: 65–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620065.
pdf.

Fischer:2001:TMR

[1888] V. Fischer and M. Drutarovský. Two methods of Rijndael implementation in reconfigurable hardware. Lecture Notes in Computer Science, 2162:77– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620077.
pdf.

Howgrave-Graham:2001:PRN

[1889] N. Howgrave-Graham, J. Dyer, and R. Gennaro. Pseudo-random number generation on the IBM 4758 Secure Crypto Coprocessor. Lecture Notes in Computer Science, 2162:93– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620093.
pdf.

Schindler:2001:EOT

[1890] W. Schindler. Efficient online tests for true random number generators. Lecture Notes in Computer Science, 2162:

103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620103.
                                        pdf.

Smart:2001:HFE

[1891] N. P. Smart. The Hessian form of an elliptic curve. Lecture Notes in Computer Science, 2162:118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620118.
                                        pdf.

Okeya:2001:EEC

[1892] K. Okeya and K. Sakurai. Efficient elliptic curve cryptosystems from a scalar multiplication algorithm with recovery of the y-coordinate on a Montgomery-form elliptic curve. Lecture Notes in Computer Science, 2162: 126–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620126.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620126.
                                        pdf.

Savas:2001:GEC

[1893] E. Savas, T. A. Schmidt, and Ç.K. Koç. Generating elliptic curves of prime order. Lecture Notes in Computer Science, 2162:142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620142.
pdf.

Shamir:2001:NDC

[1894] A. Shamir. New directions in cryptography. Lecture Notes in Computer Science, 2162:159–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2162/21620159.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2162/21620159.pdf.

Leone:2001:NLC

[1895] M. Leone. A new low complexity parallel multiplier for a class of finite fields. Lecture Notes in Computer Science, 2162:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620160.
pdf.

Rudra:2001:ERE

[1896] A. Rudra, P. K. Dubey, C. S. Jutla, V. Kumar, J. R. Rao, and P. Rohatgi. Efficient Rijndael encryption implementation with composite field arithmetic. Lecture Notes in Computer Science, 2162:171–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620171.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620171.
                                        pdf.

Tenca:2001:HRD

[1897] A. F. Tenca, G. Todorov, and Ç.K. Koç. High-radix design of a scalable modular multiplier. Lecture Notes in Computer Science, 2162:185– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620185.
                                        pdf.

Grossschadl:2001:BSU

[1898] J. Großschädl. A bit-serial unified multiplier architecture for finite fields GF(p) and GF(2^{m}). Lecture Notes in Computer Science, 2162:202–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620202.
                                        pdf.

Bond:2001:ACT

[1899] M. Bond. Attacks on cryptoprocessor transaction sets. Lecture Notes in Computer Science, 2162:220–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2162/21620220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620220.
pdf.

Young:2001:BOK

[1900] A. Young and M. Yung. Bandwidthoptimal kleptographic attacks. Lecture Notes in Computer Science, 2162:235– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620235.
pdf.

Gandolfi:2001:EAC

[1901] K. Gandolfi, C. Mourtel, and F. Olivier.

Electromagnetic analysis: Concrete results. Lecture Notes in Computer Science, 2162:251–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2162/21620251.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2162/21620251.pdf.

Bailey:2001:NCD

[1902] D. V. Bailey, D. Coffin, A. Elbirt, J. H. Silverman, and A. D. Woodbury. NTRU in constrained devices. Lecture Notes in Computer Science, 2162:262– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620262.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2162/21620262.
                                        pdf.

Pornin:2001:THE

[1903] T. Pornin. Transparent harddisk encryption. Lecture Notes in Computer Science, 2162:273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620273.
                                        pdf.

Walter:2001:SWS

[1904] C. D. Walter. Sliding windows succumbs to Big Mac attack. Lecture Notes in Computer Science, 2162:286– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620286.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620286.
                                        pdf.

Clavier:2001:UEA

[1905] C. Clavier and M. Joye. Universal exponentiation algorithm. Lecture Notes in Computer Science, 2162:300– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620300.
                                        pdf.

Akkar:2001:IAS

[1906] M.-L. Akkar and C. Giraud. An implementation of DES and AES, secure against some attacks. Lecture Notes in Computer Science, 2162:309– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620309.
pdf.

Sarkar:2001:EIS

[1907] P. Sarkar and S. Maitra. Efficient implementation of “large” stream cipher systems. Lecture Notes in Computer Science, 2162:319–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620319.
pdf.

Cheung:2001:TPS

[1908] O. Y. H. Cheung, K. H. Tsoi, P. H. W.

Leong, and M. P. Leong. Tradeoffs in parallel and serial implementations of the international data encryption algorithm IDEA. Lecture Notes in Computer Science, 2162:333– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620333.
pdf.

Orlando:2001:SGP

[1909] G. Orlando and C. Paar. A scalable GF ( p ) elliptic curve processor architecture for programmable hardware. Lecture Notes in Computer Science, 2162: 348–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620348.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620348.
                                        pdf.

Nozaki:2001:IRA

[1910] H. Nozaki, M. Motoyama, A. Shimbo, and S. Kawamura. Implementation of RSA algorithm based on RNS Montgomery multiplication. Lecture Notes in Computer Science, 2162:364– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620364.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620364.
                                        pdf.

Joye:2001:PAD

[1911] M. Joye and C. Tymen. Protections against differential analysis for elliptic curve cryptography. Lecture Notes in Computer Science, 2162:377– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2162/21620377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2162/21620377.
                                        pdf.

Liardet:2001:PSD

[1912] P.-Y. Liardet and N. P. Smart. Preventing SPA/DPA in ECC systems using the Jacobi form. Lecture Notes in Computer Science, 2162:391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2162/21620391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2162/21620391.
pdf.

Joye:2001:HEC

[1913] M. Joye and J.-J. Quisquater. Hessian elliptic curves and side-channel attacks. Lecture Notes in Computer Science, 2162:402–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2162/21620402.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2162/21620402.
pdf.

Anonymous:2001:AIgw

[1914] Anonymous. Author index. Lecture Notes in Computer Science, 2162: ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2162/2162auth.pdf.

Wilson:2001:EET

[1915] Ruth Wilson and Monica Landoni. Evaluating electronic textbooks: A methodology. Lecture Notes in Computer Science, 2163:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630001.
pdf.

Mahoui:2001:SBR

[1916] Malika Mahoui and Sally Jo Cunningham. Search behavior in a research-oriented digital library. Lecture Notes in Computer Science, 2163: 13–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630013.
pdf.

Paynter:2001:CPT

[1917] Gordon W. Paynter and Ian H. Witten. A combined phrase and thesaurus browser for large document collections. Lecture Notes in Computer Science, 2163:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630025.
pdf.

Brocks:2001:CRF

[1918] Holger Brocks, Ulrich Thiel, Adelheit Stein, and Andrea Dirsch-Weigand. Customizable retrieval functions based on user tasks in the cultural heritage domain. Lecture Notes in Computer Science, 2163:37– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630037.
pdf.

Mehrabi:2001:DW

[1919] Hamid Reza Mehrabi. Digital watermark. Lecture Notes in Computer Science, 2163:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2163/21630049.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2163/21630049.pdf.

Semeraro:2001:DCI

[1920] Giovanni Semeraro, Stefano Ferilli, Nicola Fanizzi, and Floriana Esposito. Document classification and interpretation through the inference of logic-based models. Lecture Notes in Computer Science, 2163:59–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630059.
pdf.

Furuta:2001:CPS

[1921] Richard Furuta, Siddarth S. Kalasapur, Rajiv Kochumman, Eduardo Urbina, and Ricardo Vivancos-Pérez. The cervantes project: Steps to a customizable and interlinked online electronic variorum edition supporting scholarship. Lecture Notes in Computer Science, 2163:71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630071.
pdf.

Mandl:2001:FAM

[1922] Thomas Mandl and Christa Womser-Hacker. Fusion approaches for mappings between heterogeneous ontologies. Lecture Notes in Computer Science, 2163:83–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2163/21630083.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2163/21630083.pdf.

Bauer:2001:EDL

[1923] Michael G. Bauer and Günther Specht. Enhancing digital library documents by A posteriori cross linking using XSLT. Lecture Notes in Computer Science, 2163:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630095.
pdf.

Zaslavsky:2001:UCD

[1924] Arkady Zaslavsky, Alejandro Bia, and Krisztian Monostori. Using copydetection and text comparison algorithms for cross-referencing multiple editions of literary works. Lecture Notes in Computer Science, 2163:103– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630103.
pdf.

Bergmark:2001:AAR

[1925] Donna Bergmark and Carl Lagoze.

An architecture for automatic reference linking. Lecture Notes in Computer Science, 2163:115–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630115.
pdf.

Smith:2001:DGN

[1926] David A. Smith and Gregory Crane.

Disambiguating geographic names in a historical digital library. Lecture Notes in Computer Science, 2163:127– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630127.
pdf.

Bainbridge:2001:GPD

[1927] David Bainbridge, George Buchanan, John McPherson, Steve Jones, Abdelaziz Mahoui, and Ian H. Witten. Greenstone: A platform for distributed digital library applications. Lecture Notes in Computer Science, 2163:137– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630137.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630137.
                                        pdf.

Aalberg:2001:LID

[1928] Trond Aalberg. Linking information with distributed objects. Lecture Notes in Computer Science, 2163:149– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630149.
                                        pdf.

Day:2001:MDP

[1929] Michael Day. Metadata for digital preservation: A review of recent developments. Lecture Notes in Computer Science, 2163:161–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630161.
                                        pdf.

Goncalves:2001:MFI

[1930] Marcos André Gonçalves, Robert K.

France, and Edward A. Fox. MAR-IAN: Flexible interoperability for federated digital libraries. Lecture Notes in Computer Science, 2163:173–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630173.
pdf.

Fuhr:2001:DLG

[1931] Norbert Fuhr, Preben Hansen, Michael Mabe, Andras Micsik, and Ingeborg Sølvberg. Digital libraries: A generic classification and evaluation scheme. Lecture Notes in Computer Science, 2163:187–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630187.
pdf.

Noronha:2001:DDC

[1932] Norman Noronha, João P. Campos, Daniel Gomes, Mário J. Silva, and José Borbinha. A deposit for digital collections. Lecture Notes in Computer Science, 2163:200–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630200.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2163/21630200.
                                        pdf.

Adams:2001:DLC

[1933] Anne Adams and Ann Blandford. Digital libraries in a clinical setting: Friend or foe? Lecture Notes in Computer Science, 2163:213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630213.
                                        pdf.

Radev:2001:IDI

[1934] Dragomir R. Radev, Sasha Blair-Goldensohn, Zhu Zhang, and Revathi Sundara Raghavan. Interactive, domain-independent identification and summarization of topically related news articles. Lecture Notes in Computer Science, 2163:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630225.
                                        pdf.

Meyyappan:2001:DWE

[1935] Narayanan Meyyappan, Suliman Al-Hawamdeh, and Schubert Foo. Digital work environment (DWE): Using tasks to organize digital resources. Lecture Notes in Computer Science, 2163: 239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2163/21630239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630239.
pdf.

Coleman:2001:LSD

[1936] Anita S. Coleman, Terence R. Smith, Olha A. Buchel, and Richard E. Mayer. Learning spaces in digital libraries. Lecture Notes in Computer Science, 2163: 251–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630251.
pdf.

Khoo:2001:EED

[1937] Michael Khoo. Ethnography, evaluation, and design as integrated strategies: A case study from WES. Lecture Notes in Computer Science, 2163: 263–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630263.
pdf.

Kim:2001:DME

[1938] DaeEun Kim and Sea Woo Kim. Dynamic models of expert groups to recommend Web documents. Lecture Notes in Computer Science, 2163:275– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2163/21630275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630275.
                                        pdf.

Mutschke:2001:EIR

[1939] Peter Mutschke. Enhancing information retrieval in federated bibliographic data sources using author network based stratagems. Lecture Notes in Computer Science, 2163:287– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630287.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630287.
                                        pdf.

Kholief:2001:AEB

[1940] Mohamed Kholief, Stewart Shen, and Kurt Maly. Architecture for eventbased retrieval from data streams in digital libraries. Lecture Notes in Computer Science, 2163:300–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630300.
                                        pdf.

Kanazawa:2001:ERB

[1941] Teruhito Kanazawa, Akiko Aizawa, Atsuhiro Takasu, and Jun Adachi. The effects of the relevance-based superimposition model in cross-language information retrieval. Lecture Notes in Computer Science, 2163:312–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630312.
pdf.

Ishikawa:2001:LDC

[1942] Yoshiharu Ishikawa, Yibing Chen, and Hiroyuki Kitagawa. An on-line document clustering method based on forgetting factors. Lecture Notes in Computer Science, 2163:325–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630325.
pdf.

Cheney:2001:TTI

[1943] James Cheney, Carl Lagoze, and Peter Botticelli. Towards a theory of information preservation. Lecture Notes in Computer Science, 2163:340– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630340.
pdf.

Matthes:2001:CMT

[1944] Florian Matthes, Claudia Niederée, and Ulrike Steffens. C-Merge: A tool for policy-based merging of resource classifications. Lecture Notes in Computer Science, 2163:352–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630352.
                                        pdf.

Benel:2001:TDL

[1945] Aurélien Bénel, Elöd Egyed-Zsigmond, Yannick Prié, Sylvie Calabretto, Alain Mille, Andréa Iacovella, and Jean-Marie Pinon. Truth in the digital library: From ontological to hermeneutical systems. Lecture Notes in Computer Science, 2163:366–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630366.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630366.
                                        pdf.

Hong:2001:XBC

[1946] Jen-Shin Hong, Bai-Hsuen Chen, and Jieh Hsiang. XSL-based content management for multi-presentation digital museum exhibitions. Lecture Notes in Computer Science, 2163:378– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630378.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630378.
                                        pdf.

Borgman:2001:IDE

[1947] Christine L. Borgman, Gregory H.

Leazer, Anne J. Gilliland-Swetland, and Rich Gazan. Iterative design and evaluation of a geographic digital library for university students: A case study of the Alexandria Digital Earth Prototype (ADEPT). Lecture Notes in Computer Science, 2163:390– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630390.
pdf.

Rauber:2001:AAO

[1948] Andreas Rauber and Markus Frühwirth.

Automatically analyzing and organizing music archives. Lecture Notes in Computer Science, 2163:402–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630402.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630402.
pdf.

Hunter:2001:BID

[1949] Jane Hunter and Suzanne Little. Building and indexing a distributed multimedia presentation archive using SMIL. Lecture Notes in Computer Science, 2163:415–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630415.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2163/21630415.
                                        pdf.

Spencer:2001:DCC

[1950] George Andrew Spencer. Digitization, coded character sets, and optical character recognition for multi-script information resources: The case of the Letopis’ Zhurnal’nykh Statei. Lecture Notes in Computer Science, 2163:429– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630429.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630429.
                                        pdf.

Muresan:2001:DCL

[1951] Gheorghe Muresan and David J. Harper. Document clustering and language models for system-mediated information access. Lecture Notes in Computer Science, 2163:438–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630438.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630438.
                                        pdf.

Li:2001:RDD

[1952] Guohui Li and Michael Bailou Huang. Research and development of digital libraries in china: Major issues and trends. Lecture Notes in Computer Science, 2163:450–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2163/21630450.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630450.
pdf.

Freeston:2001:WHD

[1953] Michael Freeston and Linda L. Hill.

What’s holding up the development of georeferenced DLs? Lecture Notes in Computer Science, 2163:458–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630458.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630458.
pdf.

Niggemann:2001:OAI

[1954] Elisabeth Niggemann and Matthias Hemmje. Open archive initiative, publishers and scientific societies: Future of publishing — next generation publishing models. Lecture Notes in Computer Science, 2163:459–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2163/21630459.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2163/21630459.
pdf.

Neuhold:2001:DLP

[1955] Erich J. Neuhold. Digital library programs: Current status and future plans. Lecture Notes in Computer Science, 2163:460–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2163/21630460.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2163/21630460.
                                        pdf.

Anonymous:2001:AIgx

[1956] Anonymous. Author index. Lecture Notes in Computer Science, 2163:461– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2163/2163indx.pdf.

Rouhana:2001:DCA

[1957] Nicolas Rouhana and Eric Horlait. Dynamic congestion avoidance using multi-agents systems. Lecture Notes in Computer Science, 2164:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2164/21640001.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2164/21640001.pdf.

Khoumsi:2001:UAR

[1958] Ahmed Khoumsi and Soumaya Cherkaoui. Use of agents for resolving feature interactions. Lecture Notes in Computer Science, 2164:11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2164/21640011.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2164/21640011.pdf.

Li:2001:IPB

[1959] Xining Li. IMAGO: A Prologbased system for intelligent mobile agents. Lecture Notes in Computer Science, 2164:21–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640021.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640021.pdf.

Puliafito:2001:MAL

[1960] Antonio Puliafito, Orazio Tomarchio, and Lorenzo Vita. Mobile agents and legacy systems: How to integrate alternative communication paradigms. Lecture Notes in Computer Science, 2164:31–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640031.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640031.pdf.

Amin:2001:ABD

[1961] Kaizar A. Amin, John T. Mayes, and Armin R. Mikler. Agent-based distance vector routing. Lecture Notes in Computer Science, 2164: 41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2164/21640041.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2164/21640041.
pdf.

Yewen:2001:HAC

[1962] LI Yewen and MENG Luoming. High availability connection management via software bus and mobile agent in network management environment. Lecture Notes in Computer Science, 2164:51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640051.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640051.pdf.

Gonzalez-Valenzuela:2001:MPR

[1963] Sergio Gonzalez-Valenzuela, Victor C. M. Leung, and Son T. Vuong. Multipoint-to-point routing with QoS guarantees using mobile agents. Lecture Notes in Computer Science, 2164:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640063.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640063.pdf.

Campadello:2001:DCE

[1964] Stefano Campadello. Dynamic composition of execution environment for adaptive nomadic applications. Lecture Notes in Computer Science, 2164:73–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640073.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640073.pdf.

Satoh:2001:NPM

[1965] Ichiro Satoh. Network processing of mobile agents, by mobile agents, for mobile agents. Lecture Notes in Computer Science, 2164:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640081.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640081.pdf.

Suh:2001:DMA

[1966] Young-Ho Suh, Han Namgoong, Jeong-Joon Yoo, and Dong-Ik Lee. Design of a mobile agent-based workflow management system. Lecture Notes in Computer Science, 2164: 93–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2164/21640093.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2164/21640093.
pdf.

Khan:2001:TAB

[1967] Bilal Khan, Dardo D. Kleiner, and David Talmage. Towards an agentbased distributed hierarchical network management system for all-optical networks. Lecture Notes in Computer Science, 2164:103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640103.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640103.pdf.

Lakhdissi:2001:PMS

[1968] Mouhsine Lakhdissi, Hamid Harroud, Ahmed Karmouch, and Cliff Grossner. A policy management system for mobile agent-based services. Lecture Notes in Computer Science, 2164:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2164/21640114.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2164/21640114.pdf.

Goutet:2001:MAA

[1969] Sylvain Goutet, Samuel Pierre, and Roch Glitho. A multi-agent architecture for intelligent mobile agents. Lecture Notes in Computer Science, 2164:124–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2164/21640124.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2164/21640124.pdf.

Liotta:2001:SAA

[1970] Antonio Liotta, George Pavlou, and Graham Knight. A self-adaptable agent system for efficient information gathering. Lecture Notes in Computer Science, 2164:139–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2164/21640139.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2164/21640139.pdf.

Fornaciari:2001:ABA

[1971] William Fornaciari, Vincenzo Piuri, Andrea Prestileo, and Vittorio Zaccaria. An agent-based approach to full interoperability and allocation transparency in distributed file systems. Lecture Notes in Computer Science, 2164:153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640153.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640153.pdf.

Cardoso:2001:PPC

[1972] André Ribeiro Cardoso and Joaquim Celestino, Jr. A proposal of PVCs configuration methodology employing mobile agents. Lecture Notes in Computer Science, 2164:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640163.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640163.pdf.

Perry:2001:SPI

[1973] Mark Perry and Qin Zhang. SITA: Protecting Internet trade agents from malicious hosts. Lecture Notes in Computer Science, 2164:173–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640173.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640173.pdf.

Adler:2001:AMA

[1974] Mark Adler, Jeffrey M. Bradshaw, Michael Mahan, and Niranjan Suri. Applying mobile agents to enable dynamic, context-aware interactions for mobile phone users. Lecture Notes in Computer Science, 2164:184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640184.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640184.pdf.

Quintero:2001:UMA

[1975] Alejandro Quintero, Dougoukolo Konaré, and Samuel Pierre. Using mobile agents for managing control and alarms in urban infrastructures. Lecture Notes in Computer Science, 2164:195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640195.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640195.pdf.

Huang:2001:SMA

[1976] Freeman Y. Huang and David B. Skillicorn. The spider model of agents. Lecture Notes in Computer Science, 2164:209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640209.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640209.pdf.

Muscutariu:2001:MMA

[1977] Florin Muscutariu and Marie-Pierre Gervais. On the modeling of mobile agent-based systems. Lecture Notes in Computer Science, 2164: 219–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/bibs/2164/21640219.htm;   http:
//link.springer.de/link/service/
series/0558/papers/2164/21640219.
pdf.

Agoulmine:2001:MDP

[1978] Nazim Agoulmine, Mauro Fonseca, and Alan Marshall. Multi-domain policy based management using mobile agents. Lecture Notes in Computer Science, 2164:235–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640235.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640235.pdf.

Velazquez:2001:MAP

[1979] Elio Velazquez, Nicola Santoro, and Amiya Nayak. A mobile agent prototype for distributed search. Lecture Notes in Computer Science, 2164:245–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2164/21640245.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2164/21640245.pdf.

Helvik:2001:UCE

[1980] Bjarne E. Helvik and Otto Wittner. Using the cross-entropy method to guide/ govern mobile agent’s path finding in networks. Lecture Notes in Computer Science, 2164:255–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2164/21640255.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2164/21640255.pdf.

Jamadagni:2001:CGM

[1981] N. S. Satish Jamadagni. Context gathering mobile agents to assist in resource reservation for inter-technology hand-off. Lecture Notes in Computer Science, 2164:269–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2164/21640269.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2164/21640269.pdf.

Zhang:2001:TSA

[1982] Min Zhang, Ahmed Karmouch, and Roger Impey. Towards a secure agent platform based on FIPA. Lecture Notes in Computer Science, 2164:277–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2164/21640277.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2164/21640277.pdf.

Anonymous:2001:AI

[1983] Anonymous. Author index. Lecture Notes in Computer Science, 2164: 291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.de/link/service/series/
0558/papers/2164/2164ind.pdf.

Siegle:2001:AMR

[1984] Markus Siegle. Advances in model representations. Lecture Notes in Computer Science, 2165:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2165/21650001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2165/21650001.
pdf.

Katoen:2001:FSC

[1985] Joost-Pieter Katoen, Marta Kwiatkowska, Gethin Norman, and David Parker. Faster and symbolic CTMC model checking. Lecture Notes in Computer Science, 2165:23–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2165/21650023.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2165/21650023.pdf.

DArgenio:2001:RAP

[1986] Pedro R. D’Argenio, Bertrand Jeannet, Henrik E. Jensen, and Kim G. Larsen. Reachability analysis of probabilistic systems by successive refinements. Lecture Notes in Computer Science, 2165:

39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2165/21650039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2165/21650039.
pdf.

Lopez:2001:BMD

[1987] Gabriel G. Infante López, Holger Hermanns, and Joost-Pieter Katoen. Beyond memoryless distributions: Model checking semi-Markov chains. Lecture Notes in Computer Science, 2165:57– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2165/21650057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2165/21650057.
pdf.

Folegati:2001:CLR

[1988] Katia Folegati and Roberto Segala. Coin lemmas with random variables. Lecture Notes in Computer Science, 2165:71–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2165/21650071.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2165/21650071.
pdf.

DArgenio:2001:MDL

[1989] Pedro R. D’Argenio, Holger Hermanns, Joost-Pieter Katoen, and Ric Klaren. MoDeST — A Modelling and Description Language for Stochastic Timed Systems. Lecture Notes in Computer Science, 2165:87–104, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2165/21650087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2165/21650087.
pdf.

Brim:2001:RHL

[1990] Luboš Brim, Ivana Černá, and Martin Nečesal. Randomization helps in LTL model checking. Lecture Notes in Computer Science, 2165:105– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2165/21650105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2165/21650105.
pdf.

Hillston:2001:EKR

[1991] Jane Hillston and Leïla Kloul. An efficient Kronecker representation for PEPA models. Lecture Notes in Computer Science, 2165:120–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2165/21650120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2165/21650120.
pdf.

Bernardo:2001:RBC

[1992] Marco Bernardo and Mario Bravetti.

Reward based congruences: Can we aggregate more? Lecture Notes in Computer Science, 2165:136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2165/21650136.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2165/21650136.
                                        pdf.

Cloth:2001:UMP

[1993] Lucia Cloth, Henrik Bohnenkamp, and Boudewijn Haverkort. Using max-plus algebra for the evaluation of stochastic process algebra prefixes. Lecture Notes in Computer Science, 2165:152– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2165/21650152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2165/21650152.
                                        pdf.

Bravetti:2001:EPD

[1994] Mario Bravetti and Alessandro Aldini. Expressing processes with different action durations through probabilities. Lecture Notes in Computer Science, 2165:168–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2165/21650168.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2165/21650168.
                                        pdf.

Buchholz:2001:QDB

[1995] Peter Buchholz and Peter Kemper. Quantifying the dynamic behavior of process algebras. Lecture Notes in Computer Science, 2165:184–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2165/21650184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2165/21650184.
pdf.

Clark:2001:ISP

[1996] Graham Clark and William H. Sanders.

Implementing a stochastic process algebra within the Möbius modeling framework. Lecture Notes in Computer Science, 2165:200–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2165/21650200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2165/21650200.
pdf.

Anonymous:2001:AIgy

[1997] Anonymous. Author index. Lecture Notes in Computer Science, 2165:217– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2165/2165auth.pdf.

Agirre:2001:KSW

[1998] E. Agirre and D. Martinez. Knowledge sources for word sense disambiguation. Lecture Notes in Computer Science, 2166:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2166/21660001.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2166/21660001.
                                        pdf.

Hajicova:2001:CSP

[1999] Eva Hajicová, Jan Hajic, Barbora Hladká, Martin Holub, P. Pajas, V. Reznícková, and P. Sgall. The current status of the Prague Dependency Treebank. Lecture Notes in Computer Science, 2166:11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660011.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660011.
                                        pdf.

Cermak:2001:LCC

[2000] F. Cermák. Language corpora: The Czech case. Lecture Notes in Computer Science, 2166:21–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660021.
                                        pdf.

Kruijff:2001:HLF

[2001] G.-J. M. Kruijff and I. Kruijff-Korbayová. A hybrid logic formalization of information structure sensitive discourse interpretation. Lecture Notes in Computer Science, 2166:31– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660031.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2166/21660031.
pdf.

Oliva:2001:PAD

[2002] K. Oliva. The possibilities of automatic detection/correction of errors in tagged corpora: A pilot study on a German corpus. Lecture Notes in Computer Science, 2166:39– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660039.
pdf.

Petkevic:2001:GAA

[2003] Vladimír Petkevič. Grammatical agreement and automatic morphological disambiguation of inflectional languages. Lecture Notes in Computer Science, 2166:47–53, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660047.
pdf.

Vaillant:2001:MSA

[2004] P. Vaillant. Modelling semantic association and conceptual inheritance for semantic analysis. Lecture Notes in Computer Science, 2166:54–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2166/21660054.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660054.
                                        pdf.

Kermanidis:2001:ICI

[2005] K. Kermanidis, M. Maragoudakis, N. Fakotakis, and G. Kokkinakis. Influence of conditional independence assumption on verb subcategorization detection. Lecture Notes in Computer Science, 2166:62–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660062.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660062.
                                        pdf.

Hlavacova:2001:MGC

[2006] J. Hlavácová. Morphological guesser of Czech words. Lecture Notes in Computer Science, 2166:70–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660070.
                                        pdf.

Munoz:2001:CTB

[2007] R. Muñoz and M. Palomar. Clustering technique based on semantics for definite description resolution. Lecture Notes in Computer Science, 2166:76– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2166/21660076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660076.
pdf.

Saiz-Noeda:2001:PRS

[2008] M. Saiz-Noeda, M. Palomar, and L. Moreno. Pronoun resolution in Spanish from full parsing. Lecture Notes in Computer Science, 2166:84– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660084.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660084.
pdf.

Platek:2001:MAR

[2009] V. Kubonand M. Plátek. A method of accurate robust parsing of Czech. Lecture Notes in Computer Science, 2166: 92–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660092.
pdf.

Sedlacek:2001:NCM

[2010] R. Sedláček and P. Smrž. A new Czech morphological analyser ajka. Lecture Notes in Computer Science, 2166:100– 107, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660100.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2166/21660100.
pdf.

Smrz:2001:FSR

[2011] Pavel Smrž and Pavel Rychlý. Finding semantically related words in large corpora. Lecture Notes in Computer Science, 2166:108–115, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660108.
pdf.

Gamallo:2001:SBM

[2012] P. Gamallo, C. Gasperin, A. Agustini, and G. P. Lopes. Syntactic-based methods for measuring word similarity. Lecture Notes in Computer Science, 2166: 116–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660116.
pdf.

Mima:2001:AWA

[2013] H. Mima, S. Ananiadou, and G. Nenadić. The ATRACT workbench: Automatic term recognition and clustering for terms. Lecture Notes in Computer Science, 2166:126–133, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660126.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2166/21660126.
pdf.

Xydas:2001:AAR

[2014] G. Xydas and G. Kouroupetroglou.

Augmented auditory representation of e-texts for text-to-speech systems. Lecture Notes in Computer Science, 2166: 134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660134.
pdf.

Skoumalova:2001:EVD

[2015] Hana Skoumalová, Markéta Stran̆áková-Lopatková, and Zdeněk Žabokrtský. Enhancing the valency dictionary of Czech verbs: Tectogrammatical annotation. Lecture Notes in Computer Science, 2166:142–149, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660142.
pdf.

Quaresma:2001:IGB

[2016] P. Quaresma and I. P. Rodrigues.

An interactive graph based legal information retrieval system. Lecture Notes in Computer Science, 2166:150– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660150.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660150.
                                        pdf.

Bolshakov:2001:TSP

[2017] I. A. Bolshakov and A. Gelbukh. Text segmentation into paragraphs based on local text cohesion. Lecture Notes in Computer Science, 2166:158– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660158.
                                        pdf.

Larouk:2001:BDC

[2018] O. Larouk. Building a digital collection of Web-pages: Access and filtering information with textual expansion. Lecture Notes in Computer Science, 2166: 167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660167.
                                        pdf.

Kral:2001:TAW

[2019] R. Král. Three approaches to word sense disambiguation for Czech. Lecture Notes in Computer Science, 2166: 174–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660174.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2166/21660174.
pdf.

Montoyo:2001:MWE

[2020] A. Montoyo, M. Palomar, and G. Rigau. Method for WordNet enrichment using WSD. Lecture Notes in Computer Science, 2166:180–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660180.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660180.
pdf.

Hermansky:2001:HSP

[2021] Hynek Heřmanský. Human speech perception: Some lessons from automatic speech recognition. Lecture Notes in Computer Science, 2166:187– 196, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660187.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660187.
pdf.

Mangold:2001:STR

[2022] H. Mangold. Speech technology in reality — applications, their challenges and solutions. Lecture Notes in Computer Science, 2166:197–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660197.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2166/21660197.
                                        pdf.

Vintsiuk:2001:GPT

[2023] T. K. Vintsiuk. Generative phonemethreephone model for ASR. Lecture Notes in Computer Science, 2166:201– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660201.
                                        pdf.

Kopecek:2001:AMS

[2024] I. Kopeček. Algebraic models of speech segment databases. Lecture Notes in Computer Science, 2166:208– 213, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660208.
                                        pdf.

Potamitis:2001:BNC

[2025] I. Potamitis, N. Fakotakis, and G. Kokkinakis. Bayesian noise compensation of time trajectories of spectral coefficients for robust speech recognition. Lecture Notes in Computer Science, 2166:214–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660214.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2166/21660214.
pdf.

Psutka:2001:IFS

[2026] J. Psutka, L. Müller, and J. V. Psutka.

The influence of a filter shape in telephone-based recognition module using PLP parameterization. Lecture Notes in Computer Science, 2166:222– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660222.
pdf.

Machovikov:2001:SSC

[2027] A. Machovikov and K. Stolyarov.

System for speech communication through Internet. Lecture Notes in Computer Science, 2166:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660229.
pdf.

Marcinek:2001:PBA

[2028] M. Marcinek, J. Juhar, and A. Cizmar. Phoneme based ASR system for Slovak SpeechDat database. Lecture Notes in Computer Science, 2166:237– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660237.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2166/21660237.
                                        pdf.

Zgank:2001:LVC

[2029] Andrej Žgank, Zdravko Kačič, and Bogomir Horvat. Large vocabulary continuous speech recognizer for Slovenian language. Lecture Notes in Computer Science, 2166:242–248, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660242.
                                        pdf.

Kocsor:2001:NDA

[2030] A. Kocsor, L. Tóth, and D. Paczolay. A nonlinearized discriminant analysis and its application to speech impediment therapy. Lecture Notes in Computer Science, 2166:249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660249.
                                        pdf.

Ordelman:2001:SRI

[2031] R. Ordelman, A. van Hessen, and F. de Jong. Speech recognition issues for Dutch spoken document retrieval. Lecture Notes in Computer Science, 2166:258–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2166/21660258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660258.
pdf.

Istrate:2001:MSA

[2032] D. Istrate and E. Castelli. Multichannel sound acquisition with stress situations determination for medical supervision in a smart house. Lecture Notes in Computer Science, 2166:266– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660266.
pdf.

Ircing:2001:TPR

[2033] P. Ircing and J. Psutka. Two-pass recognition of Czech speech using adaptive vocabulary. Lecture Notes in Computer Science, 2166:273–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660273.
pdf.

Kruschke:2001:SSS

[2034] H. Kruschke. Simulation of speaking styles with adapted prosody. Lecture Notes in Computer Science, 2166:278– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660278.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660278.
                                        pdf.

Volskaya:2001:UTS

[2035] N. Volskaya and P. Skrelin. The utterance type and the size of the intonation unit. Lecture Notes in Computer Science, 2166:285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660285.
                                        pdf.

Cmejla:2001:EBB

[2036] Roman Čmejla and Pavel Sovka. Estimation of boundaries between speech units using Bayesian changepoint detectors. Lecture Notes in Computer Science, 2166:291–298, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660291.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660291.
                                        pdf.

Burget:2001:DDD

[2037] L. Burget and H. Heřmanský. Data driven design of filter bank for speech recognition. Lecture Notes in Computer Science, 2166:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660299.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2166/21660299.
pdf.

Motlicek:2001:MTN

[2038] Petr Motlícek, Geneviève Baudoin, Jan Černocký, and Gérard Chollet. Minimization of transition noise and HNM synthesis in very low bit rate speech coding. Lecture Notes in Computer Science, 2166:305–312, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660305.
pdf.

Wong:2001:CRT

[2039] S. H. S. Wong and K. Pala.

Chinese radicals and top ontology in EuroWordNet. Lecture Notes in Computer Science, 2166:313–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660313.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660313.
pdf.

Stemmer:2001:TDA

[2040] G. Stemmer, V. Zeissler, E. Nöth, and H. Niemann. Towards a dynamic adjustment of the language weight. Lecture Notes in Computer Science, 2166: 323–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2166/21660323.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660323.
                                        pdf.

Potamitis:2001:SRB

[2041] I. Potamitis, N. Fakotakis, and G. Kokkinakis. Speech recognition based on feature extraction with variable rate frequency sampling. Lecture Notes in Computer Science, 2166:329– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660329.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660329.
                                        pdf.

Gros:2001:PSR

[2042] Jerneja Gros, France Mihelič, Nikola Pavešić, Mario Žganec, Aleš Mihelič, Marko Knez, Aleš Merčun, and Darko Škerl. The phonectic SMS reader. Lecture Notes in Computer Science, 2166:334–340, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660334.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660334.
                                        pdf.

Noth:2001:RIN

[2043] E. Nöth, M. Boros, J. Fischer, F. Gallwitz, J. Haas, R. Huber, H. Niemann, G. Stemmer, and V. Warnke. Research issues for the next generation spoken dialogue systems revisited. Lecture Notes in Computer Science, 2166:

341–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660341.
pdf.

Ton:2001:DUI

[2044] L. Ton and L. J. M. Rothkrantz. Determining user interface semantics using communicating agents. Lecture Notes in Computer Science, 2166:349– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660349.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660349.
pdf.

Turunen:2001:ABA

[2045] M. Turunen and J. Hakulinen. Agentbased adaptive interaction and dialogue management architecture for speech applications. Lecture Notes in Computer Science, 2166:357–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660357.
pdf.

Kruschwitz:2001:DWS

[2046] U. Kruschwitz. Dialogue for Web search utilizing automatically acquired domain knowledge. Lecture Notes in Computer Science, 2166:365–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660365.
                                        pdf.

Kogure:2001:DTS

[2047] S. Kogure and S. Nakagawa. A development tool for spoken dialogue systems and its evaluation. Lecture Notes in Computer Science, 2166:373– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660373.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660373.
                                        pdf.

Boella:2001:UXR

[2048] G. Boella, R. Damiano, M. Danieli, and L. Lesmo. Using XML for representing domain dependent knowledge in dialogos. Lecture Notes in Computer Science, 2166:381–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660381.
                                        pdf.

Nygryn:2001:DGP

[2049] Pavel Nygrýn and Robert Batůšek. Dialogue generation of program source codes. Lecture Notes in Computer Science, 2166:388–395, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660388.
pdf.

Navarretta:2001:SAI

[2050] C. Navarretta. Semi-automatic identification of Danish discourse deictics. Lecture Notes in Computer Science, 2166: 396–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660396.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660396.
pdf.

Pla:2001:LUU

[2051] F. Pla, A. Molina, E. Sanchis, E. Segarra, and F. García. Language understanding using two-level stochastic models with POS and semantic units. Lecture Notes in Computer Science, 2166:403–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660403.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660403.
pdf.

Lager:2001:SPC

[2052] T. Lager. Shallow processing and cautious incrementality in a dialogue system front end: Two steps towards robustness and reactivity. Lecture Notes in Computer Science, 2166:410– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660410.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660410.
                                        pdf.

Schwarz:2001:CCT

[2053] Jana Schwarz and Václav Matoušek. Creation of a corpus of training sentences based on automated dialogue analysis. Lecture Notes in Computer Science, 2166:418–426, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660418.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660418.
                                        pdf.

Tauser:2001:DMD

[2054] Karel Taušer, Václav Matoušek, and Roman Mouček. Dialogue manager in dialogue system for visually impaired programmers. Lecture Notes in Computer Science, 2166:427–434, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2166/21660427.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2166/21660427.
                                        pdf.

Moucek:2001:HSD

[2055] Roman Mouček, Karel Tauser, and Pavel Nygrýn. Help system in dialogue grammar based programming environment for visually impaired programmers. Lecture Notes in Computer Science, 2166:435–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660435.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660435.
pdf.

Bartek:2001:AGD

[2056] L. Bártek. Automatic generation of dialogue interfaces for Webbased applications. Lecture Notes in Computer Science, 2166:443–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2166/21660443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2166/21660443.
pdf.

Anonymous:2001:AIgz

[2057] Anonymous. Author index. Lecture Notes in Computer Science, 2166: ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2166/2166auth.pdf.

Ait-Kaci:2001:AAF

[2058] Hassan Aït-Kaci and Yutaka Sasaki.

An axiomatic approach to feature term generalization. Lecture Notes in Computer Science, 2167:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670001.
                                        pdf.

Armengol:2001:LID

[2059] Eva Armengol and Enric Plaza. Lazy induction of descriptions for relational case-based learning. Lecture Notes in Computer Science, 2167:13– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670013.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670013.
                                        pdf.

Bensusan:2001:EPA

[2060] Hilan Bensusan and Alexandros Kalousis. Estimating the predictive accuracy of a classifier. Lecture Notes in Computer Science, 2167:25– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670025.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670025.
                                        pdf.

Camacho:2001:IRE

[2061] Rui Camacho and Pavel Brazdil. Improving the robustness and encoding complexity of behavioural clones. Lecture Notes in Computer Science, 2167: 37–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2167/21670037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670037.
pdf.

Chevaleyre:2001:FLR

[2062] Yann Chevaleyre and Jean-Daniel Zucker. A framework for learning rules from multiple instance data. Lecture Notes in Computer Science, 2167:49– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670049.
pdf.

Chidlovskii:2001:WWI

[2063] Boris Chidlovskii. Wrapping Web information providers by transducer induction. Lecture Notes in Computer Science, 2167:61–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2167/21670061.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2167/21670061.pdf.

Dahl:2001:LWE

[2064] Fredrik A. Dahl and Ole Martin Halck.

Learning while exploring: Bridging the gaps in the eligibility traces. Lecture Notes in Computer Science, 2167:73– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670073.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670073.
                                        pdf.

Dahl:2001:RLA

[2065] Fredrik A. Dahl. A reinforcement learning algorithm applied to simplified twoplayer Texas hold’em poker. Lecture Notes in Computer Science, 2167:85– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670085.
                                        pdf.

Driessens:2001:SRR

[2066] Kurt Driessens, Jan Ramon, and Hendrik Blockeel. Speeding up relational reinforcement learning through the use of an incremental first order decision tree learner. Lecture Notes in Computer Science, 2167:97– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670097.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670097.
                                        pdf.

Eibl:2001:APA

[2067] Günther Eibl and Karl Peter Pfeiffer. Analysis of the performance of AdaBoost.M2 for the simulated digitrecognition-example. Lecture Notes in Computer Science, 2167:109–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2167/21670109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670109.
pdf.

El-Yaniv:2001:IDC

[2068] Ran El-Yaniv and Oren Souroujon. Iterative double clustering for unsupervised and semisupervised learning. Lecture Notes in Computer Science, 2167:121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670121.
pdf.

Elomaa:2001:PBP

[2069] Tapio Elomaa and Matti Kääriäinen.

On the practice of branching program boosting. Lecture Notes in Computer Science, 2167:133–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670133.
pdf.

Frank:2001:SAO

[2070] Eibe Frank and Mark Hall. A simple approach to ordinal classification. Lecture Notes in Computer Science, 2167: 145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2167/21670145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670145.
                                        pdf.

Gallagher:2001:FDC

[2071] Marcus Gallagher. Fitness distance correlation of neural network error surfaces: A scalable, continuous optimization problem. Lecture Notes in Computer Science, 2167:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670157.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670157.
                                        pdf.

Ganascia:2001:ERP

[2072] Jean-Gabriel Ganascia. Extraction of recurrent patterns from stratified ordered trees. Lecture Notes in Computer Science, 2167:167–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670167.
                                        pdf.

Garg:2001:UPC

[2073] Ashutosh Garg and Dan Roth. Understanding probabilistic classifiers. Lecture Notes in Computer Science, 2167: 179–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670179.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2167/21670179.
pdf.

Gu:2001:EDS

[2074] Baohua Gu, Bing Liu, Feifang Hu, and Huan Liu. Efficiently determining the starting sample size for progressive sampling. Lecture Notes in Computer Science, 2167:192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670192.
pdf.

Hoffmann:2001:USI

[2075] Achim Hoffmann, Rex Kwok, and Paul Compton. Using subclasses to improve classification learning. Lecture Notes in Computer Science, 2167:203– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670203.
pdf.

Hofmann:2001:LWP

[2076] Thomas Hofmann. Learning what people (don’t) want. Lecture Notes in Computer Science, 2167:214–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670214.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2167/21670214.
                                        pdf.

Hutter:2001:TUT

[2077] Marcus Hutter. Towards a universal theory of artificial intelligence based on algorithmic probability and sequential decisions. Lecture Notes in Computer Science, 2167:226–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670226.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670226.
                                        pdf.

Hutter:2001:CEB

[2078] Marcus Hutter. Convergence and error bounds for universal prediction of nonbinary sequences. Lecture Notes in Computer Science, 2167:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670239.
                                        pdf.

Kavsek:2001:CDT

[2079] Branko Kavšek, Nada Lavrač, and Anuška Ferligoj. Consensus decision trees: Using consensus hierarchical clustering for data relabelling and reduction. Lecture Notes in Computer Science, 2167:251–262, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2167/21670251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670251.
pdf.

Keysers:2001:LVI

[2080] Daniel Keysers, Wolfgang Macherey, Jörg Dahmen, and Hermann Ney. Learning of variability for invariant statistical pattern recognition. Lecture Notes in Computer Science, 2167:263– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670263.
pdf.

Korb:2001:EPL

[2081] Kevin B. Korb, Lucas R. Hope, and Michelle J. Hughes. The evaluation of predictive learners: Some theoretical and empirical results. Lecture Notes in Computer Science, 2167:276– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670276.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670276.
pdf.

Kwedlo:2001:EAC

[2082] Wojciech Kwedlo and Marek Krȩtowski.

An evolutionary algorithm for costsensitive decision rule learning. Lecture Notes in Computer Science, 2167: 288–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670288.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670288.
                                        pdf.

Lauer:2001:MAN

[2083] Martin Lauer. A mixture approach to novelty detection using training data with outliers. Lecture Notes in Computer Science, 2167:300–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670300.
                                        pdf.

Law:2001:ABE

[2084] Martin H. Law and James T. Kwok. Applying the Bayesian evidence framework to ν-support vector regression. Lecture Notes in Computer Science, 2167:312–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670312.
                                        pdf.

Mariano:2001:DNU

[2085] Carlos E. Mariano and Eduardo F. Morales. DQL: A new updating strategy for reinforcement learning based on Q-learning. Lecture Notes in Computer Science, 2167:324–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670324.
pdf.

Martin:2001:LBS

[2086] Lionel Martin and Frédéric Moal.

A language-based similarity measure. Lecture Notes in Computer Science, 2167:336–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2167/21670336.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2167/21670336.
pdf.

Medina-Chico:2001:BDT

[2087] Victor Medina-Chico, Alberto Suárez, and James F. Lutsko. Backpropagation in decision trees for regression. Lecture Notes in Computer Science, 2167: 348–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670348.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670348.
pdf.

Melluish:2001:CBT

[2088] Thomas Melluish, Craig Saunders, Ilia Nouretdinov, and Volodya Vovk. Comparing the Bayes and typicalness frameworks. Lecture Notes in Computer Science, 2167:360–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670360.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670360.
                                        pdf.

Moore:2001:SDA

[2089] Jason H. Moore, Joel S. Parker, and Lance W. Hahn. Symbolic discriminant analysis for mining gene expression patterns. Lecture Notes in Computer Science, 2167:372–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670372.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670372.
                                        pdf.

Nowe:2001:SAP

[2090] Ann Nowé, Johan Parent, and Katja Verbeeck. Social agents playing a periodical policy. Lecture Notes in Computer Science, 2167:382–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670382.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670382.
                                        pdf.

Ontanon:2001:LWC

[2091] Santiago Ontañón and Enric Plaza. Learning when to collaborate among learning agents. Lecture Notes in Computer Science, 2167:394–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2167/21670394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670394.
pdf.

Ragg:2001:BCC

[2092] Thomas Ragg. Building committees by clustering models based on pairwise similarity values. Lecture Notes in Computer Science, 2167:406– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670406.
pdf.

Raskutti:2001:SOF

[2093] Bhavani Raskutti, Herman Ferrá, and Adam Kowalczyk. Second order features for maximising text classification performance. Lecture Notes in Computer Science, 2167:419–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670419.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670419.
pdf.

Sanz-Gonzalez:2001:IST

[2094] José L. Sanz-González and Diego Andina. Importance sampling techniques in neural detector training. Lecture Notes in Computer Science, 2167:431– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670431.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670431.
                                        pdf.

Suc:2001:IQT

[2095] Dorian Šuc and Ivan Bratko. Induction of qualitative trees. Lecture Notes in Computer Science, 2167:442– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670442.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670442.
                                        pdf.

Taira:2001:TCU

[2096] Hirotoshi Taira and Masahiko Haruno. Text categorization using transductive boosting. Lecture Notes in Computer Science, 2167:454–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670454.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670454.
                                        pdf.

Tang:2001:UMC

[2097] Lappoon R. Tang and Raymond J. Mooney. Using multiple clause constructors in inductive logic programming for semantic parsing. Lecture Notes in Computer Science, 2167:466– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2167/21670466.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670466.
pdf.

Todorovski:2001:UDK

[2098] Ljupčo Todorovski and Sašo Džeroski.

Using domain knowledge on population dynamics modeling for equation discovery. Lecture Notes in Computer Science, 2167:478–490, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670478.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670478.
pdf.

Turney:2001:MWS

[2099] Peter D. Turney. Mining the Web for synonyms: PMI–IR versus LSA on TOEFL. Lecture Notes in Computer Science, 2167:491–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670491.
pdf.

Vilalta:2001:UFE

[2100] Ricardo Vilalta, Mark Brodie, Daniel Oblinger, and Irina Rish. A unified framework for evaluation metrics in classification using decision trees. Lecture Notes in Computer Science, 2167: 503–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670503.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670503.
                                        pdf.

Vivaldi:2001:ITE

[2101] Jordi Vivaldi, Lluís Màrquez, and Horacio Rodríguez. Improving term extraction by system combination using boosting. Lecture Notes in Computer Science, 2167:515–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670515.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670515.
                                        pdf.

Vucetic:2001:CDB

[2102] Slobodan Vucetic and Zoran Obradovic. Classification on data with biased class distribution. Lecture Notes in Computer Science, 2167:527–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670527.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670527.
                                        pdf.

Washio:2001:DAS

[2103] Takashi Washio, Hiroshi Motoda, and Yuji Niwa. Discovering admissible simultaneous equation models from observed data. Lecture Notes in Computer Science, 2167:539–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670539.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670539.
pdf.

Widmer:2001:DSP

[2104] Gerhard Widmer. Discovering strong principles of expressive music performance with the PLCG rule learning strategy. Lecture Notes in Computer Science, 2167:552–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670552.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670552.
pdf.

Yang:2001:PID

[2105] Ying Yang and Geoffrey I. Webb.

Proportional k-interval discretization for naive-Bayes classifiers. Lecture Notes in Computer Science, 2167:564– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670564.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670564.
pdf.

Zenobi:2001:UDP

[2106] Gabriele Zenobi and Pádraig Cunningham. Using diversity in preparing ensembles of classifiers based on different feature subsets to minimize generalization error. Lecture Notes in Computer Science, 2167:576–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670576.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670576.
                                        pdf.

Zhang:2001:GPN

[2107] Huajie Zhang and Charles X. Ling. Geometric properties of naive Bayes in nominal domains. Lecture Notes in Computer Science, 2167:588–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670588.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670588.
                                        pdf.

Dietterich:2001:SVRa

[2108] Thomas G. Dietterich and Xin Wang. Support vectors for reinforcement learning. Lecture Notes in Computer Science, 2167:600–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670600.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670600.
                                        pdf.

Mannila:2001:CDAa

[2109] Heikki Mannila. Combining discrete algorithmic and probabilistic approaches in data mining. Lecture Notes in Computer Science, 2167:601– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670601.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670601.
pdf.

Unwin:2001:SMNa

[2110] Antony Unwin. Statistification or mystification? the need for statistical thought in visual data mining. Lecture Notes in Computer Science, 2167: 602–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670602.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670602.
pdf.

Widmer:2001:MEPa

[2111] Gerhard Widmer. The musical expression project: A challenge for machine learning and knowledge discovery. Lecture Notes in Computer Science, 2167: 603–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2167/21670603.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2167/21670603.
pdf.

Wrobel:2001:SSSa

[2112] Stefan Wrobel. Scalability, search, and sampling: From smart algorithms to active discovery. Lecture Notes in Computer Science, 2167:615– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2167/21670615.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2167/21670615.
                                        pdf.

Anonymous:2001:AIha

[2113] Anonymous. Author index. Lecture Notes in Computer Science, 2167:617– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2167/2167auth.pdf.

Adibi:2001:SSL

[2114] Jafar Adibi and Wei-Min Shen. Selfsimilar layered hidden Markov models. Lecture Notes in Computer Science, 2168:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2168/21680001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680001.
                                        pdf.

Amini:2001:ATS

[2115] Massih-Reza Amini and Patrick Gallinari. Automatic text summarization using unsupervised and semisupervised learning. Lecture Notes in Computer Science, 2168:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680016.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680016.
                                        pdf.

Blockeel:2001:DTC

[2116] H. Blockeel, J. Fürnkranz, A. Prskawetz, and F. C. Billari. Detecting temporal change in event sequences: An application to demographic data. Lecture Notes in Computer Science, 2168:29– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680029.
pdf.

Clare:2001:KDM

[2117] Amanda Clare and Ross D. King.

Knowledge discovery in multi-label phenotype data. Lecture Notes in Computer Science, 2168:42–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680042.
pdf.

Coenen:2001:CAR

[2118] Frans Coenen, Graham Goulbourne, and Paul Leng. Computing association rules using partial totals. Lecture Notes in Computer Science, 2168:54– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680054.
pdf.

Congdon:2001:GGA

[2119] Clare Bates Congdon. Gaphyl: A genetic algorithms approach to cladistics. Lecture Notes in Computer Science, 2168:67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2168/21680067.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680067.
                                        pdf.

Egecioglu:2001:PAA

[2120] Ömer Eğecioğlu. Parametric approximation algorithms for high-dimensional Euclidean similarity. Lecture Notes in Computer Science, 2168:79–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680079.
                                        pdf.

Estivill-Castro:2001:DSM

[2121] Vladimir Estivill-Castro and Michael E. Houle. Data structures for minimization of total within-group distance for spatio-temporal clustering. Lecture Notes in Computer Science, 2168:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680091.
                                        pdf.

Estivill-Castro:2001:NCC

[2122] Vladimir Estivill-Castro and Jianhua Yang. Non-crisp clustering by fast, convergent, and robust algorithms. Lecture Notes in Computer Science, 2168: 103–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680103.
pdf.

Geurts:2001:PET

[2123] Pierre Geurts. Pattern extraction for time series classification. Lecture Notes in Computer Science, 2168:115– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680115.
pdf.

Giannotti:2001:SMA

[2124] Fosca Giannotti, Giuseppe Manco, and Franco Turini. Specifying mining algorithms with iterative user-defined aggregates: A case study. Lecture Notes in Computer Science, 2168:128– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680128.
pdf.

deGraaf:2001:IFA

[2125] Jeannette M. de Graaf, Walter A. Kosters, and Jeroen J. W. Witteman. Interesting fuzzy association rules in quantitative databases. Lecture Notes in Computer Science, 2168: 140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680140.
                                        pdf.

Gyenesei:2001:IMF

[2126] Attila Gyenesei. Interestingness measures for fuzzy association rules. Lecture Notes in Computer Science, 2168: 152–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680152.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680152.
                                        pdf.

Halkidi:2001:DSO

[2127] Maria Halkidi and Michalis Vazirgiannis. A data set oriented approach for clustering algorithm selection. Lecture Notes in Computer Science, 2168:165– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680165.
                                        pdf.

Hilario:2001:FMK

[2128] Melanie Hilario and Alexandros Kalousis. Fusion of meta-knowledge and meta-data for case-based model selection. Lecture Notes in Computer Science, 2168:180–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2168/21680180.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2168/21680180.pdf.

Hoppner:2001:DTP

[2129] Frank Höppner. Discovery of temporal patterns. learning rules about the qualitative behaviour of time series. Lecture Notes in Computer Science, 2168:192– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680192.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680192.
pdf.

Honda:2001:TRD

[2130] Rie Honda and Osamu Konishi.

Temporal rule discovery for timeseries satellite images and integration with RDB. Lecture Notes in Computer Science, 2168:204–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680204.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680204.
pdf.

Hong:2001:UGI

[2131] Theodore W. Hong and Keith L. Clark. Using grammatical inference to automate information extraction from the Web. Lecture Notes in Computer Science, 2168:216–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680216.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680216.
                                        pdf.

Hu:2001:BSD

[2132] Yuh-Jyh Hu. Biological sequence data mining. Lecture Notes in Computer Science, 2168:228–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2168/21680228.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2168/21680228.pdf.

Hullermeier:2001:IBF

[2133] Eyke Hüllermeier. Implication-based fuzzy association rules. Lecture Notes in Computer Science, 2168:241–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680241.
                                        pdf.

Jaroszewicz:2001:GMR

[2134] Szymon Jaroszewicz and Dan A. Simovici. A general measure of rule interestingness. Lecture Notes in Computer Science, 2168:253–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680253.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680253.
pdf.

Kindermann:2001:ECC

[2135] Jörg Kindermann, Gerhard Paass, and Edda Leopold. Error correcting codes with optimized Kullback–Leibler distances for text categorization. Lecture Notes in Computer Science, 2168:266– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680266.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680266.
pdf.

Knobbe:2001:PA

[2136] Arno J. Knobbe, Marc de Haas, and Arno Siebes. Propositionalisation and aggregates. Lecture Notes in Computer Science, 2168:277–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680277.
pdf.

Kuznetsov:2001:ACC

[2137] Sergei O. Kuznetsov and Sergei A.

Obiedkov. Algorithms for the construction of concept lattices and their diagram graphs. Lecture Notes in Computer Science, 2168:289–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680289.
pdf.

Lazarevic:2001:DRU

[2138] Aleksandar Lazarevic and Zoran Obradovic. Data reduction using multiple models integration. Lecture Notes in Computer Science, 2168:301– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680301.
pdf.

Mendes:2001:DFC

[2139] Roberto R. F. Mendes, Fabricio de B. Voznika, Alex A. Freitas, and Julio C. Nievola. Discovering fuzzy classification rules with genetic programming and co-evolution. Lecture Notes in Computer Science, 2168:314– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680314.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680314.
pdf.

Nedellec:2001:SFI

[2140] Claire Nédellec, Mohamed Ould Abdel Vetah, and Philippe Bessières. Sentence filtering for information extraction in genomics, a classification problem. Lecture Notes in Computer Science, 2168:326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680326.
pdf.

Ontrup:2001:TCS

[2141] Jörg Ontrup and Helge Ritter. Text categorization and semantic browsing with self-organizing maps on non-Euclidean spaces. Lecture Notes in Computer Science, 2168:338–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680338.
pdf.

Oyang:2001:SHD

[2142] Yen-Jen Oyang, Chien-Yu Chen, and Tsui-Wei Yang. A study on the hierarchical data clustering algorithm based on gravity theory. Lecture Notes in Computer Science, 2168:350– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680350.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2168/21680350.
                                        pdf.

Park:2001:IDF

[2143] Laurence A. F. Park, Marimuthu Palaniswami, and Ramamohanarao Kotagiri. Internet document filtering using Fourier domain scoring. Lecture Notes in Computer Science, 2168:362– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680362.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680362.
                                        pdf.

Pereira:2001:DNL

[2144] Francisco Pereira, Marcel Just, and Tom Mitchell. Distinguishing natural language processes on the basis of fMRI-measured brain activation. Lecture Notes in Computer Science, 2168: 374–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680374.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680374.
                                        pdf.

Pernelle:2001:ACR

[2145] Nathalie Pernelle, Marie-Christine Rousset, and Véronique Ventos. Automatic construction and refinement of a class hierarchy over multi-valued data. Lecture Notes in Computer Science, 2168:386–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2168/21680386.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680386.
pdf.

Robardet:2001:CTO

[2146] Céline Robardet and Fabien Feschet.

Comparison of three objective functions for conceptual clustering. Lecture Notes in Computer Science, 2168: 399–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680399.
pdf.

Roberts:2001:IEA

[2147] Felice M. Roberts, Richard J. Povinelli, and Kristina M. Ropella. Identification of ECG arrhythmias using phase space reconstruction. Lecture Notes in Computer Science, 2168:411– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680411.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680411.
pdf.

Scheffer:2001:FAR

[2148] Tobias Scheffer. Finding association rules that trade support optimally against confidence. Lecture Notes in Computer Science, 2168:424– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680424.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680424.
                                        pdf.

Suzuki:2001:BDT

[2149] Einoshin Suzuki, Masafumi Gotoh, and Yuta Choki. Bloomy decision tree for multi-objective classification. Lecture Notes in Computer Science, 2168:436– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680436.
                                        pdf.

Tsumoto:2001:DTK

[2150] Shusaku Tsumoto. Discovery of temporal knowledge in medical timeseries databases using moving average, multiscale matching, and rule induction. Lecture Notes in Computer Science, 2168:448–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680448.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680448.
                                        pdf.

Tsumoto:2001:MPN

[2151] Shusaku Tsumoto. Mining positive and negative knowledge in clinical databases based on rough set model. Lecture Notes in Computer Science, 2168:460–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2168/21680460.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2168/21680460.
pdf.

Unwin:2001:TPM

[2152] Antony Unwin, Heike Hofmann, and Klaus Bernt. The TwoKey plot for multiple association rules control. Lecture Notes in Computer Science, 2168: 472–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680472.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680472.
pdf.

Weiss:2001:LCF

[2153] Sholom M. Weiss and Nitin Indurkhya.

Lightweight collaborative filtering method for binary-encoded data. Lecture Notes in Computer Science, 2168: 484–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680484.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680484.
pdf.

Dietterich:2001:SVRb

[2154] Thomas G. Dietterich and Xin Wang. Support vectors for reinforcement learning. Lecture Notes in Computer Science, 2168:492–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680492.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680492.
                                        pdf.

Mannila:2001:CDAb

[2155] Heikki Mannila. Combining discrete algorithmic and probabilistic approaches in data mining. Lecture Notes in Computer Science, 2168:493– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680493.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680493.
                                        pdf.

Unwin:2001:SMNb

[2156] Antony Unwin. Statistification or mystification? the need for statistical thought in visual data mining. Lecture Notes in Computer Science, 2168: 494–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2168/21680494.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2168/21680494.
                                        pdf.

Widmer:2001:MEPb

[2157] Gerhard Widmer. The musical expression project: A challenge for machine learning and knowledge discovery. Lecture Notes in Computer Science, 2168: 495–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2168/21680495.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680495.
pdf.

Wrobel:2001:SSSb

[2158] Stefan Wrobel. Scalability, search, and sampling: From smart algorithms to active discovery. Lecture Notes in Computer Science, 2168:507– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2168/21680507.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2168/21680507.
pdf.

Anonymous:2001:AIhb

[2159] Anonymous. Author index. Lecture Notes in Computer Science, 2168:509– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2168/2168auth.pdf.

Jaedicke:2001:CI

[2160] Michael Jaedicke. Chapter 1: Introduction. Lecture Notes in Computer Science, 2169:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2169/21690001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2169/21690001.pdf.

Jaedicke:2001:CBU

[2161] Michael Jaedicke. Chapter 2: Background on user-defined routines. Lecture Notes in Computer Science, 2169: 5–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2169/21690005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2169/21690005.
pdf.

Jaedicke:2001:CPP

[2162] Michael Jaedicke. Chapter 3: Parallel processing of user-defined functions. Lecture Notes in Computer Science, 2169:14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2169/21690014.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2169/21690014.
pdf.

Jaedicke:2001:CIF

[2163] Michael Jaedicke. Chapter 4: Intrafunction parallelism. Lecture Notes in Computer Science, 2169:33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2169/21690033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2169/21690033.
pdf.

Jaedicke:2001:CMO

[2164] Michael Jaedicke. Chapter 5: The multi-operator method. Lecture Notes in Computer Science, 2169:45–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2169/21690045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2169/21690045.
pdf.

Jaedicke:2001:CUD

[2165] Michael Jaedicke. Chapter 6: Userdefined table operators. Lecture Notes in Computer Science, 2169:67– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2169/21690067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2169/21690067.
pdf.

Jaedicke:2001:CIU

[2166] Michael Jaedicke. Chapter 7: Implementation of UDTO. Lecture Notes in Computer Science, 2169:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2169/21690106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2169/21690106.
pdf.

Jaedicke:2001:CSC

[2167] Michael Jaedicke. Chapter 8: Summary, conclusion, and future work. Lecture Notes in Computer Science, 2169: 145–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2169/21690145.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2169/21690145.
                                        pdf.

Jaedicke:2001:CA

[2168] Michael Jaedicke. Chapter: Appendix A. Lecture Notes in Computer Science, 2169:157–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2169/21690157.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2169/21690157.
                                        pdf.

Jaedicke:2001:R

[2169] Michael Jaedicke. References. Lecture Notes in Computer Science, 2169:157– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2169/21690151.pdf.

Vandenhoute:2001:PIO

[2170] M. Vandenhoute, F. Masetti, A. Jourdan, and D. Chiaroni. Progressive introduction of optical packet switching techniques in WDM networks. Lecture Notes in Computer Science, 2170: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700001.
                                        pdf.

Bregni:2001:OPS

[2171] S. Bregni, G. Guerra, and A. Pattavina. Optical packet switching for IPover-WDM transport networks. Lecture Notes in Computer Science, 2170: 10–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700010.
pdf.

Callegati:2001:MOP

[2172] F. Callegati, W. Cerroni, G. Corazza, and C. Raffaelli. MPLS over optical packet switching. Lecture Notes in Computer Science, 2170:26–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700026.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700026.
pdf.

Bianco:2001:ACP

[2173] A. Bianco, G. Galante, E. Leonardi, F. Neri, and M. Rundo. Access control protocols for interconnected WDM rings in the DAVID metro network. Lecture Notes in Computer Science, 2170:38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700038.
pdf.

Battiti:2001:RST

[2174] R. Battiti and M. Brunato. Reactive search for traffic grooming in WDM networks. Lecture Notes in Computer Science, 2170:56–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700056.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700056.
                                        pdf.

Soma-Reddy:2001:MSI

[2175] J. Soma-Reddy and A. Acampora. Micromobility strategies for IP based cellular networks. Lecture Notes in Computer Science, 2170:67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700067.
                                        pdf.

Loreti:2001:SSP

[2176] P. Loreti, M. Luglio, R. Kapoor, J. Stepanek, M. Gerla, F. Vatalaro, and M. A. Vázquez-Castro. Satellite systems performance with TCP–IP applications. Lecture Notes in Computer Science, 2170:76–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700076.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700076.
                                        pdf.

Zorzi:2001:PTW

[2177] M. Zorzi, M. Rossi, and G. Mazzini. Performance of TCP on a wideband CDMA air interface. Lecture Notes in Computer Science, 2170:91– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700091.
pdf.

Soursos:2001:DSG

[2178] S. Soursos, C. Courcoubetis, and G. C. Polyzos. Differentiated services in the GPRS wireless access environment. Lecture Notes in Computer Science, 2170:108–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700108.
pdf.

Bruno:2001:WAI

[2179] R. Bruno, M. Conti, and E. Gregori. Wireless access to Internet via IEEE 802.11: An optimal control strategy for minimizing the energy consumption. Lecture Notes in Computer Science, 2170:120–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700120.
pdf.

Lopez:2001:SQP

[2180] A. López, J. Manner, A. Mihailovic, H. Velayos, E. Hepworth, and Y. Khouaja. A study on QoS provision for IP-based radio access networks. Lecture Notes in Computer Science, 2170:139–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700139.
                                        pdf.

Salamatian:2001:MBM

[2181] K. Salamatian and S. Fdida. Measurement based modeling of quality of service in the Internet: A methodological approach. Lecture Notes in Computer Science, 2170:158– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700158.
                                        pdf.

Bianchi:2001:PEM

[2182] G. Bianchi, N. Blefari-Melazzi, M. Femminella, and F. Pugini. Performance evaluation of a measurement-based algorithm for distributed admission control in a DiffServ framework. Lecture Notes in Computer Science, 2170:175– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700175.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2170/21700175.
pdf.

Sargento:2001:RSE

[2183] S. Sargento, R. Valadas, and E. Knightly.

Resource stealing in endpoint controlled multi-class networks. Lecture Notes in Computer Science, 2170:195– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700195.
pdf.

Lin:2001:PBC

[2184] X. Lin and N. B. Shroff. Pricingbased control of large networks. Lecture Notes in Computer Science, 2170: 212–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700212.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700212.
pdf.

Bianchi:2001:EAC

[2185] G. Bianchi, N. Blefari-Melazzi, and V. Mancuso. Endpoint admission control over assured forwarding PHBs and its performance over RED implementations. Lecture Notes in Computer Science, 2170:232–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700232.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700232.
                                        pdf.

Davoli:2001:BBA

[2186] F. Davoli and P. Maryni. A bandwidth broker assignment scheme in DiffServ networks. Lecture Notes in Computer Science, 2170:251–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700251.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700251.
                                        pdf.

Thiran:2001:PBA

[2187] P. Thiran, N. Taft, C. Diot, H. Zang, and R. Mac Donald. A protectionbased approach to QoS in packet over fiber networks. Lecture Notes in Computer Science, 2170:266–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700266.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700266.
                                        pdf.

Hristea:2001:IRM

[2188] C. Hristea and F. Tobagi. IP routing and mobility. Lecture Notes in Computer Science, 2170:279–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700279.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2170/21700279.
pdf.

Gerla:2001:AMS

[2189] M. Gerla, A. Fei, J.-H. Cui, and M. Faloutsos. Aggregated multicast for scalable QoS multicast provisioning. Lecture Notes in Computer Science, 2170:295–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2170/21700295.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2170/21700295.
pdf.

Iannaccone:2001:FSR

[2190] G. Iannaccone and L. Rizzo. Fairness of a single-rate multicast congestion control scheme. Lecture Notes in Computer Science, 2170:309–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700309.
pdf.

Graves:2001:DCW

[2191] E. E. Graves, R. Srikant, and D. Towsley. Decentralized computation of weighted max-min fair bandwidth allocation in networks with multicast flows. Lecture Notes in Computer Science, 2170:326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700326.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700326.
                                        pdf.

Almeroth:2001:CIA

[2192] K. C. Almeroth, S. Bhattacharyya, and C. Diot. Challenges of integrating ASM and SSM IP multicast protocol architectures. Lecture Notes in Computer Science, 2170:343– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700343.
                                        pdf.

Francini:2001:IPQ

[2193] A. Francini and F. M. Chiussi. Integrated provision of QoS guarantees to unicast and multicast traffic in packet switches. Lecture Notes in Computer Science, 2170:361–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700361.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700361.
                                        pdf.

Tobagi:2001:SDI

[2194] F. Tobagi, W. Noureddine, B. Chen, A. Markopoulou, Chuck Fraleigh, Mansour Karam, Jose-Miguel Pulido, and Jun ichi Kimura. Service differentiation in the Internet to support multimedia traffic. Lecture Notes in Computer Science, 2170:381–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700381.
pdf.

Liu:2001:TIS

[2195] H. Liu, X. Wei, and M. El Zarki.

A transport infrastructure supporting real time interactive MPEG-4 clientserver applications over IP networks. Lecture Notes in Computer Science, 2170:401–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2170/21700401.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2170/21700401.
pdf.

Cernuto:2001:AFE

[2196] A. Cernuto, A. Lombardo, and G. Schembra. An analytical framework to evaluate the effects of a TCP-friendly congestion control over adaptive-rate MPEG video sources. Lecture Notes in Computer Science, 2170:413–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2170/21700413.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2170/21700413.
pdf.

Rejaie:2001:DIL

[2197] R. Rejaie and A. Reibman. Design issues for layered quality-adaptive Internet video playback. Lecture Notes in Computer Science, 2170:433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700433.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700433.
                                        pdf.

Balaouras:2001:MSC

[2198] P. Balaouras and I. Stavrakakis. A multi-state congestion control scheme for rate adaptable unicast continuous media flows. Lecture Notes in Computer Science, 2170:452–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700452.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700452.
                                        pdf.

Bianchi:2001:PDE

[2199] G. Bianchi, F. Borgonovo, A. Capone, L. Fratta, and C. Petrioli. PCP-DV: An end-to-end admission control mechanism for IP telephony. Lecture Notes in Computer Science, 2170:470– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700470.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700470.
                                        pdf.

Garroppo:2001:DAS

[2200] R. G. Garroppo, S. Giordano, S. Niccolini, and F. Russo. DiffServ aggregation strategies of real time services in a WFQ+ schedulers network. Lecture Notes in Computer Science, 2170: 481–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700481.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700481.
pdf.

Senac:2001:TNG

[2201] P. Sénac, E. Exposito, and M. Diaz.

Towards a new generation of generic transport protocols. Lecture Notes in Computer Science, 2170:492–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700492.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700492.
pdf.

Ji:2001:ADR

[2202] P. Ji, J. Kurose, and D. Towsley. Activating and deactivating repair servers in active multicast trees. Lecture Notes in Computer Science, 2170:507– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700507.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700507.
pdf.

Grossglauser:2001:LSA

[2203] M. Grossglauser and B. Krishnamurthy. Looking for science in the art of network measurement. Lecture Notes in Computer Science, 2170:524– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700524.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700524.
                                        pdf.

Garetto:2001:UQN

[2204] M. Garetto, R. Lo Cigno, M. Meo, and M. Ajmone Marsan. On the use of queueing network models to predict the performance of TCP connections. Lecture Notes in Computer Science, 2170: 536–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700536.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700536.
                                        pdf.

Fraleigh:2001:DDP

[2205] C. Fraleigh, C. Diot, B. Lyles, S. Moon, P. Owezarski, D. Papagiannaki, and F. Tobagi. Design and deployment of a passive monitoring infrastructure. Lecture Notes in Computer Science, 2170: 556–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700556.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700556.
                                        pdf.

Duffield:2001:NDT

[2206] N. G. Duffield, J. Horowitz, F. Lo Presti, and D. Towsley. Network delay tomography from end-to-end unicast measurements. Lecture Notes in Computer Science, 2170:576–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700576.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700576.
pdf.

Sikdar:2001:CTS

[2207] B. Sikdar and K. S. Vastola. On the contribution of TCP to the selfsimilarity of network traffic. Lecture Notes in Computer Science, 2170:596– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700596.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700596.
pdf.

Capone:2001:BET

[2208] A. Capone and F. Martignon. Bandwidth estimates in the TCP congestion control scheme. Lecture Notes in Computer Science, 2170:614–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700614.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700614.
pdf.

Fantacci:2001:IPE

[2209] R. Fantacci, G. Vannuccini, A. Mannelli, and L. Pancani. Implementation and performance evaluation of a differentiated services test-bed for voice over IP support. Lecture Notes in Computer Science, 2170:627– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700627.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700627.
pdf.

DArienzo:2001:DSB

[2210] M. D’Arienzo, M. Esposito, S. P. Romano, and G. Ventre. Dynamic SLA-based management of virtual private networks. Lecture Notes in Computer Science, 2170:642–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700642.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700642.
pdf.

Lehtonen:2001:DSP

[2211] R. Lehtonen and J. Harju. Distributed service platform for managing stream services. Lecture Notes in Computer Science, 2170:660–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700660.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700660.
pdf.

Koch:2001:IQW

[2212] B. Koch and S. Salsano. IP QoS at work: Definition and implementation of the AQUILA architecture. Lecture Notes in Computer Science, 2170:674– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700674.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700674.
pdf.

Choi:2001:APF

[2213] T.-U. Choi, M.-K. Ji, S.-H. Park, and K. d. Chung. An adaptive periodic FEC scheme for Internet video applications. Lecture Notes in Computer Science, 2170:691–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700691.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700691.
pdf.

Bernardini:2001:EIV

[2214] R. Bernardini and G. M. Cortelazzo. Efficient interaction with virtual worlds over IP. Lecture Notes in Computer Science, 2170:703–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2170/21700703.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2170/21700703.
pdf.

Bartolini:2001:MVD

[2215] F. Bartolini, V. Cappellini, R. Caldelli, A. De Rosa, A. Piva, and M. Barni. MPEG-4 video data protection for Internet distribution. Lecture Notes in Computer Science, 2170:713–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2170/21700713.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2170/21700713.
                                        pdf.

Anonymous:2001:AIhc

[2216] Anonymous. Author index. Lecture Notes in Computer Science, 2170:721– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2170/2170auth.pdf.

Ryan:2001:MMC

[2217] Peter Y. A. Ryan. Mathematical models of computer security. Lecture Notes in Computer Science, 2171:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2171/21710001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2171/21710001.
                                        pdf.

Syverson:2001:LAP

[2218] Paul Syverson and Iliano Cervesato. The logic of authentication protocols. Lecture Notes in Computer Science, 2171:63–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2171/21710063.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2171/21710063.
pdf.

Samarati:2001:ACP

[2219] Pierangela Samarati and Sabrina de Capitani di Vimercati. Access control: Policies, models, and mechanisms. Lecture Notes in Computer Science, 2171:137–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2171/21710137.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2171/21710137.
pdf.

Guttman:2001:SGP

[2220] Joshua D. Guttman. Security goals:

Packet trajectories and strand spaces. Lecture Notes in Computer Science, 2171:197–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2171/21710197.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2171/21710197.
pdf.

Gordon:2001:NNC

[2221] Andrew D. Gordon. Notes on nominal calculi for security and mobility. Lecture Notes in Computer Science, 2171: 262–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2171/21710262.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2171/21710262.
                                        pdf.

Focardi:2001:CSP

[2222] Riccardo Focardi and Roberto Gorrieri. Classification of security properties (Part I: Information flow). Lecture Notes in Computer Science, 2171: 331–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2171/21710331.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2171/21710331.
                                        pdf.

Anonymous:2001:AIhd

[2223] Anonymous. Author index. Lecture Notes in Computer Science, 2171:397– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2171/2171ind.pdf.

Bernstein:2001:GMM

[2224] Philip A. Bernstein. Generic model management: A database infrastructure for schema manipulation. Lecture Notes in Computer Science, 2172:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720001.
                                        pdf.

Cobb:2001:EDC

[2225] Edward E. Cobb. The evolution of distributed component architectures. Lecture Notes in Computer Science, 2172: 7–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720007.
pdf.

Lenzerini:2001:DIH

[2226] Maurizio Lenzerini. Data integration is harder than you thought (extended abstract). Lecture Notes in Computer Science, 2172:22–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720022.
pdf.

Blanzieri:2001:ICM

[2227] Enrico Blanzieri, Paolo Giorgini, Paolo Massa, and Sabrina Recla. Implicit culture for multi-agent interaction support. Lecture Notes in Computer Science, 2172:27–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720027.
pdf.

Busetta:2001:EMA

[2228] Paolo Busetta, Luciano Serafini, and Dhirendra Singh Floriano Zini. Extending multi-agent cooperation by overhearing. Lecture Notes in Computer Science, 2172:40–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720040.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720040.
                                        pdf.

Guan:2001:MAB

[2229] Jihong Guan, Shuigeng Zhou, and Aoying Zhou. Mobile-agent based distributed Web GIS. Lecture Notes in Computer Science, 2172:53–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720053.
                                        pdf.

Ogston:2001:LDA

[2230] Elth Ogston and Stamatis Vassiliadis. Local distributed agent matchmaking. Lecture Notes in Computer Science, 2172:67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2172/21720067.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720067.
                                        pdf.

Schimkat:2001:DDS

[2231] Ralf-Dieter Schimkat, Michael Friedrich, and Wolfgang Küchlin. Deploying distributed state information in mobile agent systems. Lecture Notes in Computer Science, 2172:80–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720080.
pdf.

Shintani:2001:CMS

[2232] Toramatsu Shintani and Takayuki Ito. Cooperative meeting scheduling among agents based on multiple negotiations. Lecture Notes in Computer Science, 2172:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720095.
pdf.

Berlin:2001:AAD

[2233] Jacob Berlin and Amihai Motro. Autoplex: Automated discovery of content for virtual databases. Lecture Notes in Computer Science, 2172:108– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720108.
pdf.

Panti:2001:CSI

[2234] Maurizio Panti, Luca Spalazzi, and Loris Penserini. Cooperation strategies for information integration. Lecture Notes in Computer Science, 2172: 123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720123.
                                        pdf.

Patel:2001:POS

[2235] Shuchi Patel and Amit Sheth. Planning and optimizing semantic information requests using domain modeling and resource characteristics. Lecture Notes in Computer Science, 2172:135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720135.
                                        pdf.

Roantree:2001:IVS

[2236] Mark Roantree, Jessie B. Kennedy, and Peter J. Barclay. Integrating view schemata using an extended object definition language. Lecture Notes in Computer Science, 2172:150– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720150.
                                        pdf.

Rosaci:2001:DSS

[2237] Domenico Rosaci, Giorgio Terracina, and Domenico Ursino. Deriving “sub-source” similarities from heterogeneous, semi-structured information sources. Lecture Notes in Computer Science, 2172:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720163.
pdf.

Aberer:2001:PGS

[2238] Karl Aberer. P-Grid: A selforganizing access structure for P2P information systems. Lecture Notes in Computer Science, 2172:179–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720179.
pdf.

Cilia:2001:MAF

[2239] Mariano Cilia, Christof Bornhövd, and Alejandro P. Buchmann. Moving active functionality from centralized to open distributed heterogeneous environments. Lecture Notes in Computer Science, 2172:195–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720195.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2172/21720195.
                                        pdf.

Muhl:2001:GCC

[2240] Gero Mühl. Generic constraints for content-based publish/subscribe. Lecture Notes in Computer Science, 2172: 211–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720211.
                                        pdf.

Pinto:2001:SHU

[2241] Monica Pinto, Mercedes Amor, Lidia Fuentes, and Jose M. Troya. Supporting heterogeneous users in collaborative virtual environments using AOP. Lecture Notes in Computer Science, 2172: 226–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720226.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720226.
                                        pdf.

Baina:2001:PSM

[2242] Karim Baïna, Khalid Benali, and Claude Godart. A process service model for dynamic enterprise process interconnection. Lecture Notes in Computer Science, 2172:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720239.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2172/21720239.
pdf.

Haber:2001:EMI

[2243] Cornelia Haber. Employing multiuser interactions in the development of synchronous applications. Lecture Notes in Computer Science, 2172:255– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720255.
pdf.

vandenHeuvel:2001:SRD

[2244] Willem-Jan van den Heuvel, Jian Yang, and Mike P. Papazoglou. Service representation, discovery, and composition for E-marketplaces. Lecture Notes in Computer Science, 2172:270– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720270.
pdf.

Berthold:2001:SDQ

[2245] Henrike Berthold and Klaus Meyer-Wegener. Schema design and query processing in a federated multimedia database system. Lecture Notes in Computer Science, 2172:285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720285.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720285.
                                        pdf.

Brayner:2001:GSS

[2246] Angelo Brayner and Theo Härder. Global semantic serializability: An approach to increase concurrency in multidatabase systems. Lecture Notes in Computer Science, 2172:301–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720301.
                                        pdf.

Doucet:2001:CIC

[2247] Anne Doucet, Stephane Gançarski, Claudia León, and Marta Rukoz. Checking integrity constraints in multidatabase systems with nested transactions. Lecture Notes in Computer Science, 2172:316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2172/21720316.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2172/21720316.pdf.

Cameron:2001:IMT

[2248] Mark A. Cameron, Kerry L. Taylor, and David J. Abel. The Internet marketplace template: An architecture template for inter-enterprise information systems. Lecture Notes in Computer Science, 2172:329–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720329.
pdf.

Debenham:2001:IEE

[2249] John Debenham and Simeon Simoff.

Investigating the evolution of electronic markets. Lecture Notes in Computer Science, 2172:344–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720344.
pdf.

Tolksdorf:2001:CWB

[2250] Robert Tolksdorf and Dirk Glaubitz.

Coordinating Web-based systems with documents in XMLSpaces. Lecture Notes in Computer Science, 2172:356– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720356.
pdf.

Zadorozhny:2001:VAC

[2251] Vladimir Zadorozhny, Louiqa Raschid, Tao Zhan, and Laura Bright. Validating an access cost model for wide area applications. Lecture Notes in Computer Science, 2172:371–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720371.
                                        pdf.

Christophides:2001:QSX

[2252] Vassilis Christophides, Richard Hull, and Akhil Kumar. Querying and splicing of XML workflows. Lecture Notes in Computer Science, 2172:386– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720386.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720386.
                                        pdf.

Zeng:2001:DBB

[2253] Liangzhao Zeng, Boualem Benatallah, and Anne H. H. Ngu. On demand business-to-business integration. Lecture Notes in Computer Science, 2172: 403–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2172/21720403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2172/21720403.
                                        pdf.

Shahabi:2001:YAS

[2254] Cyrus Shahabi, Farnoush Banaei-Kashani, Yi-Shin Chen, and Dennis McLeod. Yoda: An accurate and scalable Web-based recommendation system. Lecture Notes in Computer Science, 2172:418–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720418.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720418.
pdf.

Modica:2001:UMG

[2255] Giovanni Modica, Avigdor Gal, and Hasan M. Jamil. The use of machinegenerated ontologies in dynamic information seeking. Lecture Notes in Computer Science, 2172:433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2172/21720433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2172/21720433.
pdf.

Anonymous:2001:AIhe

[2256] Anonymous. Author index. Lecture Notes in Computer Science, 2172:449– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2172/2172ind.pdf.

Dix:2001:CLA

[2257] Jürgen Dix. A computational logic approach to heterogeneous agent systems. Lecture Notes in Computer Science, 2173:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730001.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2173/21730001.
                                        pdf.

Baumgartner:2001:DIE

[2258] Robert Baumgartner, Sergio Flesca, and Georg Gottlob. Declarative information extraction, Web crawling, and recursive wrapping with lixto. Lecture Notes in Computer Science, 2173:21– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730021.
                                        pdf.

Kirousis:2001:CMC

[2259] Lefteris M. Kirousis and Phokion G. Kolaitis. On the complexity of model checking and inference in minimal models. Lecture Notes in Computer Science, 2173:42–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730042.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730042.
                                        pdf.

Lenzerini:2001:DIN

[2260] Maurizio Lenzerini. Data integration needs reasoning. Lecture Notes in Computer Science, 2173:54–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730054.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2173/21730054.
pdf.

Sakama:2001:NIL

[2261] Chiaki Sakama. Nonmonotonic inductive logic programming. Lecture Notes in Computer Science, 2173:62– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730062.
pdf.

Turner:2001:SEL

[2262] Hudson Turner. Strong equivalence for logic programs and default theories (made easy). Lecture Notes in Computer Science, 2173:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730081.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730081.
pdf.

Janhunen:2001:EDN

[2263] Tomi Janhunen. On the effect of default negation on the expressiveness of disjunctive rules. Lecture Notes in Computer Science, 2173:93– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730093.
pdf.

Marek:2001:ESL

[2264] V. W. Marek and J. B. Remmel. On the expressibility of stable logic programming. Lecture Notes in Computer Science, 2173:107–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730107.
                                        pdf.

Brewka:2001:RBD

[2265] Gerhard Brewka. On the relationship between defeasible logic and wellfounded semantics. Lecture Notes in Computer Science, 2173:121–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730121.
                                        pdf.

Wang:2001:CSW

[2266] Kewen Wang. A comparative study of well-founded semantics for disjunctive logic programs. Lecture Notes in Computer Science, 2173:133–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730133.
                                        pdf.

Bonatti:2001:ROL

[2267] Piero A. Bonatti. Reasoning with open logic programs. Lecture Notes in Computer Science, 2173:147–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730147.
pdf.

Nicolas:2001:RIK

[2268] Pascal Nicolas and Béatrice Duval. Representation of incomplete knowledge by induction of default theories. Lecture Notes in Computer Science, 2173:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730160.
pdf.

Ventos:2001:EUD

[2269] Véronique Ventos, Pierre Brézellec, and Henry Soldano. Explicitly using default knowledge in concept learning: An extended description logics plus strict and default rules. Lecture Notes in Computer Science, 2173:173– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730173.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730173.
pdf.

Baral:2001:DSS

[2270] Chitta Baral and Cenk Uyan. Declarative specification and solution of combinatorial auctions using logic programming. Lecture Notes in Computer Science, 2173:186–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2173/21730186.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2173/21730186.pdf.

Heljanko:2001:BLM

[2271] Keijo Heljanko and Ilkka Niemelä. Bounded LTL model checking with stable models. Lecture Notes in Computer Science, 2173:200–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730200.
                                        pdf.

Gelfond:2001:DPS

[2272] Michael Gelfond, Marcello Balduccini, and Joel Galloway. Diagnosing physical systems in A-prolog. Lecture Notes in Computer Science, 2173:213– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730213.
                                        pdf.

Son:2001:PDF

[2273] Tran Cao Son, Chitta Baral, and Sheila McIlraith. Planning with different forms of domain-dependent control knowledge — an answer set programming approach. Lecture Notes in Computer Science, 2173:226–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730226.
pdf.

Foo:2001:ESF

[2274] Norman Foo, Dongmo Zhang, Yan Zhang, Samir Chopra, and Bao Quoc Vo. Encoding solutions of the frame problem in dynamic logic. Lecture Notes in Computer Science, 2173:240– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730240.
pdf.

Kakas:2001:RRA

[2275] Antonis Kakas, Rob Miller, and Francesca Toni. E-RES: Reasoning about actions, events and observations. Lecture Notes in Computer Science, 2173:254–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2173/21730254.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2173/21730254.
                                        pdf.

Syrjanen:2001:ORL

[2276] Tommi Syrjänen. Omega-restricted logic programs. Lecture Notes in Computer Science, 2173:267–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730267.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730267.
                                        pdf.

Leone:2001:IAI

[2277] Nicola Leone, Simona Perri, and Francesco Scarcello. Improving ASP instantiators by join-ordering methods. Lecture Notes in Computer Science, 2173:280–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2173/21730280.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730280.
                                        pdf.

Faber:2001:OCH

[2278] Wolfgang Faber, Nicola Leone, and Gerald Pfeifer. Optimizing the computation of heuristics for answer set programming systems. Lecture Notes in Computer Science, 2173:295–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730295.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2173/21730295.
pdf.

Nicolas:2001:NGS

[2279] Pascal Nicolas, Frédéric Saubion, and Igor Stéphan. New generation systems for non-monotonic reasoning. Lecture Notes in Computer Science, 2173:309– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730309.
pdf.

Avin:2001:ACX

[2280] Chen Avin and Rachel Ben-Eliyahu-Zohary. Algorithms for computing X-minimal models. Lecture Notes in Computer Science, 2173:322–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730322.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730322.
pdf.

Lukasiewicz:2001:FCM

[2281] Thomas Lukasiewicz. Fixpoint characterizations for many-valued disjunctive logic programs with probabilistic semantics. Lecture Notes in Computer Science, 2173:336–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730336.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2173/21730336.
                                        pdf.

Medina:2001:MAL

[2282] Jesús Medina, Manuel Ojeda-Aciego, and Peter Vojtáš. Multi-adjoint logic programming with continuous semantics. Lecture Notes in Computer Science, 2173:351–364, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730351.
                                        pdf.

Leite:2001:MDD

[2283] João Alexandre Leite, José Júlio Alferes, and Luís Moniz Pereira. Multi-dimensional dynamic knowledge representation. Lecture Notes in Computer Science, 2173:365–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730365.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730365.
                                        pdf.

Damasio:2001:ALP

[2284] Carlos Viegas Damásio and Luís Moniz Pereira. Antitonic logic programs. Lecture Notes in Computer Science, 2173: 379–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730379.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2173/21730379.
pdf.

VanNuffelen:2001:SDP

[2285] Bert Van Nuffelen and Antonis Kakas.

A-system: Declarative programming with abduction. Lecture Notes in Computer Science, 2173:393–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730393.
pdf.

Eiter:2001:UFE

[2286] Thomas Eiter, Michael Fink, Giuliana Sabbatini, and Hans Tompits. An update front-end for extended logic programs. Lecture Notes in Computer Science, 2173:397–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730397.
pdf.

East:2001:AIA

[2287] Deborah East and Mirosław Truszczyński.

aspps — an implementation of answer-set programming with propositional schemata. Lecture Notes in Computer Science, 2173:402–405, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730402.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730402.
                                        pdf.

Anger:2001:NSNb

[2288] Christian Anger, Kathrin Konczak, and Thomas Linke. NoMoRe : A system for non-monotonic reasoning under answer set semantics. Lecture Notes in Computer Science, 2173:406– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730406.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730406.
                                        pdf.

Delgrande:2001:PGC

[2289] James P. Delgrande, Torsten Schaub, and Hans Tompits. plp: A generic compiler for ordered logic programs. Lecture Notes in Computer Science, 2173: 411–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730411.
                                        pdf.

Bonatti:2001:PRI

[2290] Piero A. Bonatti. Prototypes for reasoning with infinite stable models and function symbols. Lecture Notes in Computer Science, 2173:416– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730416.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2173/21730416.
pdf.

Otero:2001:PSM

[2291] Ramón P. Otero. psmodels: Stable models based on pertinence. Lecture Notes in Computer Science, 2173:420– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730420.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2173/21730420.
pdf.

DellArmi:2001:SDD

[2292] Tina Dell’Armi, Wolfgang Faber, Giuseppe Ielpa, Christoph Koch, Nicola Leone, Simona Perri, and Gerald Pfeifer. System description: DLV. Lecture Notes in Computer Science, 2173:424–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2173/21730424.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2173/21730424.
pdf.

Eiter:2001:SDD

[2293] Thomas Eiter, Wolfgang Faber, Nicola Leone, Gerald Pfeifer, and Axel Polleres. System description: The DLV^{K} planning system. Lecture Notes in Computer Science, 2173:429–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2173/21730429.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730429.
                                        pdf.

Syrjanen:2001:SS

[2294] Tommi Syrjänen and Ilkka Niemelä. The smodels system. Lecture Notes in Computer Science, 2173:434–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730434.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730434.
                                        pdf.

Balduccini:2001:UAC

[2295] Marcello Balduccini, Michael Gelfond, R. Watson, and M. Nogueira. The USA-Advisor: A case study in answer set planning. Lecture Notes in Computer Science, 2173:439–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2173/21730439.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2173/21730439.
                                        pdf.

Anonymous:2001:AIhf

[2296] Anonymous. Author index. Lecture Notes in Computer Science, 2173:443– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2173/2173ind.pdf.

Kearns:2001:CGT

[2297] Michael Kearns. Computational game theory and AI. Lecture Notes in Computer Science, 2174:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740001.
pdf.

Ozcan:2001:OAS

[2298] Fatma Özcan, V. S. Subrahmanian, and Leana Golubchik. Optimal agent selection. Lecture Notes in Computer Science, 2174:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740002.
pdf.

Degtyarev:2001:TFO

[2299] Anatoli Degtyarev and Michael Fisher.

Towards first-order temporal resolution. Lecture Notes in Computer Science, 2174:18–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2174/21740018.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2174/21740018.pdf.

Kusters:2001:AMS

[2300] Ralf Küsters and Ralf Molitor. Approximating most specific concepts in description logics with existential restrictions. Lecture Notes in Computer Science, 2174:33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740033.
pdf.

Ragg:2001:BLE

[2301] Thomas Ragg. Bayesian learning and evolutionary parameter optimization. Lecture Notes in Computer Science, 2174:48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2174/21740048.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2174/21740048.
pdf.

Badea:2001:APO

[2302] Liviu Badea and Doina Tilivea. Abductive partial order planning with dependent fluents. Lecture Notes in Computer Science, 2174:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740063.
pdf.

Bennewitz:2001:CBO

[2303] Maren Bennewitz, Wolfram Burgard, and Sebastian Thrun. Constraintbased optimization of priority schemes for decoupled path planning techniques. Lecture Notes in Computer Science, 2174:78–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2174/21740078.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2174/21740078.pdf.

Bochman:2001:PWS

[2304] Alexander Bochman. Possible worlds semantics for credulous and contraction inference. Lecture Notes in Computer Science, 2174:94–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740094.
pdf.

Broxvall:2001:PAB

[2305] Mathias Broxvall. The point algebra for branching time revisited. Lecture Notes in Computer Science, 2174:106– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740106.
pdf.

Bruning:2001:ECE

[2306] Stefan Brüning. Exploiting conditional equivalences in connection calculi. Lecture Notes in Computer Science, 2174: 122–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740122.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740122.
                                        pdf.

East:2001:PSA

[2307] Deborah East and Mirosław Truszczyński. Propositional satisfiability in answerset programming. Lecture Notes in Computer Science, 2174:138–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740138.
                                        pdf.

Edelkamp:2001:PRS

[2308] Stefan Edelkamp. Prediction of regular search tree growth by spectral analysis. Lecture Notes in Computer Science, 2174:154–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2174/21740154.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740154.
                                        pdf.

Edelkamp:2001:TPT

[2309] Stefan Edelkamp and Ulrich Meyer. Theory and practice of time-space trade-offs in memory limited search. Lecture Notes in Computer Science, 2174:169–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2174/21740169.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2174/21740169.
pdf.

Felfernig:2001:HDL

[2310] Alexander Felfernig, Gerhard E.

Friedrich, Dietmar Jannach, Markus Stumptner, and Markus Zanker. Hierarchical diagnosis of large configurator knowledge bases. Lecture Notes in Computer Science, 2174:185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740185.
pdf.

Felfernig:2001:TDC

[2311] Alexander Felfernig, Gerhard E.

Friedrich, Dietmar Jannach, and Markus Zanker. Towards distributed configuration. Lecture Notes in Computer Science, 2174:198–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740198.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740198.
pdf.

Grosskreutz:2001:BUP

[2312] Henrik Grosskreutz and Gerhard Lakemeyer. Belief update in the pGOLOG framework. Lecture Notes in Computer Science, 2174:213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740213.
                                        pdf.

Huffner:2001:FOS

[2313] Falk Hüffner, Stefan Edelkamp, Henning Fernau, and Rolf Niedermeier. Finding optimal solutions to atomix. Lecture Notes in Computer Science, 2174:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740229.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740229.
                                        pdf.

Iwan:2001:HBD

[2314] Gero Iwan. History-based diagnosis templates in the framework of the situation calculus. Lecture Notes in Computer Science, 2174:244–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740244.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740244.
                                        pdf.

Jamroga:2001:DMG

[2315] Wojciech Jamroga. A defense model for games with incomplete information. Lecture Notes in Computer Science, 2174:260–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2174/21740260.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2174/21740260.
pdf.

Kiziltan:2001:TIL

[2316] Zeynep Kiziltan, Pierre Flener, and Brahim Hnich. Towards inferring labelling heuristics for CSP application domains. Lecture Notes in Computer Science, 2174:275–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740275.
pdf.

Martin:2001:AQP

[2317] Yves Martin and Michael Thielscher.

Addressing the qualification problem in FLUX. Lecture Notes in Computer Science, 2174:290–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740290.
pdf.

Schonherr:2001:ESF

[2318] Frank Schönherr, Mihaela Cistelecan, Joachim Hertzberg, and Thomas Christaller. Extracting situation facts from activation value histories in behavior-based robots. Lecture Notes in Computer Science, 2174:305–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740305.
                                        pdf.

Schulz:2001:LSC

[2319] Stephan Schulz. Learning search control knowledge for equational theorem proving. Lecture Notes in Computer Science, 2174:320–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740320.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740320.
                                        pdf.

Stumme:2001:ISR

[2320] Gerd Stumme, Rafik Taouil, Yves Bastide, Nicolas Pasquier, and Lotfi Lakhal. Intelligent structuring and reducing of association rules with formal concept analysis. Lecture Notes in Computer Science, 2174:335–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740335.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740335.
                                        pdf.

Stumptner:2001:CTM

[2321] Markus Stumptner, Dominik Wieland, and Franz Wotawa. Comparing two models for software debugging. Lecture Notes in Computer Science, 2174: 351–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740351.
pdf.

Thielscher:2001:IIS

[2322] Michael Thielscher. Inferring implicit state knowledge and plans with sensing actions. Lecture Notes in Computer Science, 2174:366–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740366.
pdf.

Anastassakis:2001:MAS

[2323] George Anastassakis, Tim Ritchings, and Themis Panayiotopoulos. Multi-agent systems as intelligent virtual environments. Lecture Notes in Computer Science, 2174:381–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740381.
pdf.

Bechhofer:2001:ORA

[2324] Sean Bechhofer, Ian Horrocks, Carole Goble, and Robert Stevens. OilEd: A reason-able ontology editor for the semantic Web. Lecture Notes in Computer Science, 2174:396–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740396.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740396.
                                        pdf.

Benzmuller:2001:EAO

[2325] Christoph Benzmüller, Mateja Jamnik, Manfred Kerber, and Volker Sorge. Experiments with an agentoriented reasoning system. Lecture Notes in Computer Science, 2174:409– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740409.
                                        pdf.

Belker:2001:LEN

[2326] Thorsten Belker and Michael Beetz. Learning to execute navigation plans. Lecture Notes in Computer Science, 2174:425–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2174/21740425.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740425.
                                        pdf.

Fleischanderl:2001:DMB

[2327] Gerhard Fleischanderl, Thomas Havelka, Herwig Schreiner, Markus Stumptner, and Franz Wotawa. DiKe — A model-based diagnosis kernel and its application. Lecture Notes in Computer Science, 2174:440–454, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740440.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740440.
pdf.

Fleischanderl:2001:CAC

[2328] Gerhard Fleischanderl. Constraints applied to configurations. Lecture Notes in Computer Science, 2174:455– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740455.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740455.
pdf.

Koehler:2001:TPA

[2329] Jana Koehler. From theory to practice: AI planning for high performance elevator control. Lecture Notes in Computer Science, 2174:459–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2174/21740459.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2174/21740459.
pdf.

Lebeth:2001:SNK

[2330] Kai Lebeth. Semantic networks in a knowledge management portal. Lecture Notes in Computer Science, 2174:

463–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740463.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740463.
                                        pdf.

Sundermeyer:2001:CSN

[2331] Kurt Sundermeyer. Collaborative supply net management. Lecture Notes in Computer Science, 2174:467–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2174/21740467.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2174/21740467.
                                        pdf.

Anonymous:2001:AIhg

[2332] Anonymous. Author index. Lecture Notes in Computer Science, 2174:471– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2174/2174auth.pdf.

Serra:2001:MCA

[2333] A. Serra and A. Giordana. A Monte Carlo approach to hard relational learning problems. Lecture Notes in Computer Science, 2175:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2175/21750001.
pdf.

Esposito:2001:BMC

[2334] Roberto Esposito and Lorenza Saitta.

Boosting as a Monte Carlo algorithm. Lecture Notes in Computer Science, 2175:11–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2175/21750011.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2175/21750011.
pdf.

Malerba:2001:SIM

[2335] Donato Malerba, Annalisa Appice, Antonia Bellino, Michelangelo Ceci, and Domenico Pallotta. Stepwise induction of model trees. Lecture Notes in Computer Science, 2175:20– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750020.
pdf.

Passerini:2001:EMF

[2336] Andrea Passerini, Paolo Frasconi, and Giovanni Soda. Evaluation methods for focused crawling. Lecture Notes in Computer Science, 2175:33– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750033.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2175/21750033.
pdf.

Castellano:2001:KBN

[2337] Giovanna Castellano and Anna Maria Fanelli. A knowledge-based neurocomputing approach to extract refined linguistic rules from data. Lecture Notes in Computer Science, 2175:40– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750040.
pdf.

Blonda:2001:RNE

[2338] Palma Blonda, Andrea Baraldi, Annarita D’Addabbo, Cristina Tarantino, and Roberto De Blasi. RBF networks exploiting supervised data in the adaptation of hidden neuron parameters. Lecture Notes in Computer Science, 2175:51–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750051.
pdf.

Yao:2001:NML

[2339] Yuan Yao, Gian Luca Marcialis, Massimiliano Pontil, Paolo Frasconi, and Fabio Roli. A new machine learning approach to fingerprint classification. Lecture Notes in Computer Science, 2175: 57–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750057.
pdf.

Orio:2001:AAB

[2340] Nicola Orio. An automatic accompanist based on hidden Markov models. Lecture Notes in Computer Science, 2175:64–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2175/21750064.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2175/21750064.
pdf.

Botta:2001:RVR

[2341] Marco Botta. Resampling vs reweighting in boosting a relational weak learner. Lecture Notes in Computer Science, 2175:70–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750070.
pdf.

Ferilli:2001:LLM

[2342] Stefano Ferilli, Nicola Fanizzi, and Giovanni Semeraro. Learning logic models for automated text categorization. Lecture Notes in Computer Science, 2175: 81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2175/21750081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750081.
                                        pdf.

Abbattista:2001:UPA

[2343] Fabio Abbattista, Nicola Fanizzi, Stefano Ferilli, Pasquale Lops, and Giovanni Semeraro. User profiling in an application of electronic commerce. Lecture Notes in Computer Science, 2175:87–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750087.
                                        pdf.

Aiello:2001:CSS

[2344] Marco Aiello. Computing spatial similarity by games. Lecture Notes in Computer Science, 2175:99–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750099.
                                        pdf.

Giunchiglia:2001:ABT

[2345] Enrico Giunchiglia, Massimo Narizzano, and Armando Tacchella. An analysis of backjumping and trivial truth in quantified Boolean formulas satisfiability. Lecture Notes in Computer Science, 2175:111–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750111.
pdf.

Leone:2001:APL

[2346] Nicola Leone, Simona Perri, and Francesco Scarcello. Abduction with penalization in logic programming. Lecture Notes in Computer Science, 2175:123–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2175/21750123.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2175/21750123.
pdf.

Panati:2001:CSD

[2347] Andrea Panati and Daniele Theseider Dupré. Causal simulation and diagnosis of dynamic systems. Lecture Notes in Computer Science, 2175:135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750135.
pdf.

Roli:2001:CPL

[2348] Andrea Roli and Christian Blum.

Critical parallelization of local search for MAX–SAT. Lecture Notes in Computer Science, 2175:147–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750147.
                                        pdf.

Bandini:2001:PDP

[2349] Stefania Bandini and Sara Manzoni. Product design as product revise: The case of chemical compounds. Lecture Notes in Computer Science, 2175:159– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750159.
                                        pdf.

Giordano:2001:BRR

[2350] Laura Giordano, Valentina Gliozzi, and Nicola Olivetti. Belief revision and the Ramsey test: A solution. Lecture Notes in Computer Science, 2175:165– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750165.
                                        pdf.

Magro:2001:SPC

[2351] Diego Magro and Pietro Torasso. Supporting product configuration in a virtual store. Lecture Notes in Computer Science, 2175:176–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2175/21750176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750176.
pdf.

Tamma:2001:CCP

[2352] Valentina A. M. Tamma and Trevor J. M. Bench-Capon. Characterising concept’s properties in ontologies. Lecture Notes in Computer Science, 2175: 189–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750189.
pdf.

Remagnino:2001:RAD

[2353] Paolo Remagnino, Graeme A. Jones, and Ndedi Monekosso. Reasoning about dynamic scenes using autonomous agents. Lecture Notes in Computer Science, 2175:201–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750201.
pdf.

Falcone:2001:TCL

[2354] Rino Falcone and Cristiano Castelfranchi. Tuning the collaboration level with autonomous agents: A principled theory. Lecture Notes in Computer Science, 2175:212–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750212.
                                        pdf.

Boella:2001:DLR

[2355] Guido Boella, Leonardo Lesmo, and Lyda Favali. The definition of legal relations in a BDI multiagent framework. Lecture Notes in Computer Science, 2175:225–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750225.
                                        pdf.

Giordano:2001:RAA

[2356] Laura Giordano, Alberto Martelli, and Camilla Schwind. Reasoning about actions in a multiagent domain. Lecture Notes in Computer Science, 2175:237– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750237.
                                        pdf.

Bandini:2001:LLS

[2357] Stefania Bandini, Sara Manzoni, Giulio Pavesi, and Carla Simone. L*MASS: A language for situated multi-agent systems. Lecture Notes in Computer Science, 2175:249–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750249.
pdf.

Carofiglio:2001:ISG

[2358] Valeria Carofiglio, Cristiano Castelfranchi, Fiorella de Rosis, and Roberto Grassano. An interactive system for generating arguments in deceptive communication. Lecture Notes in Computer Science, 2175:255–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750255.
pdf.

Neri:2001:ABA

[2359] Filippo Neri. An agent based approach to virtual market place simulation. Lecture Notes in Computer Science, 2175: 267–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750267.
pdf.

Magnini:2001:ODQ

[2360] Bernardo Magnini, Matteo Negri, and Roberto Prevete. Open domain question/answering on the Web. Lecture Notes in Computer Science, 2175:273– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750273.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750273.
                                        pdf.

Cassotta:2001:UAI

[2361] Maria Teresa Cassotta, Berardina De Carolis, Fiorella de Rosis, Chiara Andreoli, and M. Luisa De Cicco. Useradapted image descriptions from annotated knowledge sources. Lecture Notes in Computer Science, 2175:285– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750285.
                                        pdf.

Costa:2001:WCI

[2362] Fabrizio Costa, Vincenzo Lombardo, Paolo Frasconi, and Giovanni Soda. Wide coverage incremental parsing by learning attachment preferences. Lecture Notes in Computer Science, 2175: 297–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750297.
                                        pdf.

Basili:2001:FPA

[2363] Roberto Basili, Maria Teresa Pazienza, and Fabio Massimo Zanzotto. Flexible parsing architectures for NLP applications. Lecture Notes in Computer Science, 2175:308–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2175/21750308.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2175/21750308.pdf.

Carolis:2001:IPA

[2364] Berardina De Carolis, Pietro Di Maggio, and Sebastiano Pizzutilo. Information presentation adapted to the “user in context”. Lecture Notes in Computer Science, 2175:314–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750314.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750314.
pdf.

Basili:2001:HAO

[2365] Roberto Basili, Alessandro Moschitti, and Maria Teresa Pazienza. A hybrid approach to optimize feature selection process in text classification. Lecture Notes in Computer Science, 2175:320– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750320.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750320.
pdf.

Bonarini:2001:CAR

[2366] Andrea Bonarini, Matteo Matteucci, and Marcello Restelli. Concepts for anchoring in robotics. Lecture Notes in Computer Science, 2175:327–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750327.
pdf.

Chella:2001:SCR

[2367] Antonio Chella, Marcello Frixione, and Salvatore Gaglio. Symbolic and conceptual representation of dynamic scenes: Interpreting situation calculus on conceptual spaces. Lecture Notes in Computer Science, 2175:333– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750333.
pdf.

Adorni:2001:NTO

[2368] Giovanni Adorni, Luca Bolognini, Stefano Cagnoni, and Monica Mordonini. A non-traditional omnidirectional vision system with stereo capabilities for autonomous robots. Lecture Notes in Computer Science, 2175:344–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750344.
pdf.

Infantino:2001:ASR

[2369] Ignazio Infantino and Antonio Chella.

Architectural scenes reconstruction from uncalibrated photos and map based model knowledge. Lecture Notes in Computer Science, 2175:356–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750356.
pdf.

Chella:2001:SAH

[2370] Antonio Chella, Maria Donatella Guarino, and Roberto Pirrone. A SOM/ ARSOM hierarchy for the description of dynamic scenes. Lecture Notes in Computer Science, 2175:362– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750362.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2175/21750362.
pdf.

Cesta:2001:CBA

[2371] Amedeo Cesta, Gabriella Cortellessa, Angelo Oddi, Nicola Policella, and Angelo Susi. A constraint-based architecture for flexible support to activity scheduling. Lecture Notes in Computer Science, 2175:369–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2175/21750369.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2175/21750369.
                                        pdf.

Barruffi:2001:PED

[2372] Rosy Barruffi and Michela Milano. Planning and execution in dynamic environments. Lecture Notes in Computer Science, 2175:382–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750382.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750382.
                                        pdf.

Armano:2001:AAP

[2373] Giuliano Armano, Giancarlo Cherchi, and Eloisa Vargiu. An agent architecture for planning in a dynamic environment. Lecture Notes in Computer Science, 2175:388–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2175/21750388.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2175/21750388.
                                        pdf.

Anonymous:2001:AIhh

[2374] Anonymous. Author index. Lecture Notes in Computer Science, 2175:395– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2175/2175indx.pdf.

Feldmann:2001:SLS

[2375] Raimund L. Feldmann and Klaus-Dieter Althoff. On the status of learning software organizations in the year 2001. Lecture Notes in Computer Science, 2176:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760002.
pdf.

Henninger:2001:OLD

[2376] Scott Henninger. Organizational learning in dynamic domains. Lecture Notes in Computer Science, 2176:8– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760008.
pdf.

Lehner:2001:KAH

[2377] Franz Lehner. Keynote address: How do companies learn? selected applications from the IT sector. Lecture Notes in Computer Science, 2176:17– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760017.
pdf.

Janetzko:2001:PKC

[2378] Dietmar Janetzko. Panel: Knowledge creating communities. Lecture Notes in Computer Science, 2176:18– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760018.
pdf.

Trittmann:2001:OMF

[2379] Ralph Trittmann. The organic and the mechanistic form of managing knowledge in software development. Lecture Notes in Computer Science, 2176:22– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760022.
pdf.

Pfahl:2001:CTM

[2380] Dietmar Pfahl, Niniek Angkasaputra, Christiane Differding, and Günther Ruhe. CORONET-Train: A methodology for Web-based collaborative learning in software organisations. Lecture Notes in Computer Science, 2176:37– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760037.
pdf.

GressevonWangenheim:2001:SKM

[2381] Christiane Gresse von Wangenheim, Daniel Lichtnow, Aldo von Wangenheim, and Eros Comunello. Supporting knowledge management in university software R&D groups. Lecture Notes in Computer Science, 2176:52– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760052.
pdf.

Segal:2001:OLS

[2382] Judith Segal. Organisational learning and software process improvement: A case study. Lecture Notes in Computer Science, 2176:68–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760068.
pdf.

Voss:2001:CSV

[2383] Angi Voß, Gernot Richter, Steve Moyle, and Alípio Jorge. Collaboration support for virtual data mining enterprises. Lecture Notes in Computer Science, 2176:83–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760083.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2176/21760083.
                                        pdf.

Dingsoyr:2001:SMK

[2384] Torgeir Dingsøyr and Emil Røyrvik. Skills management as knowledge technology in a software consultancy company. Lecture Notes in Computer Science, 2176:96–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2176/21760096.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2176/21760096.
                                        pdf.

Lindvall:2001:LLA

[2385] Mikael Lindvall, Michael Frey, Patricia Costa, and Roseanne Tesoriero. Lessons learned about structuring and describing experience for three experience bases. Lecture Notes in Computer Science, 2176:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2176/21760106.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2176/21760106.
                                        pdf.

Krishnaswamy:2001:KET

[2386] Shonali Krishnaswamy, Seng Wai Loke, and Arkady Zaslavsky. Knowledge elicitation through Web-based data mining services. Lecture Notes in Computer Science, 2176:120–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2176/21760120.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760120.
pdf.

Lewerentz:2001:MAM

[2387] Claus Lewerentz, Heinrich Rust, and Frank Simon. A model for analyzing measurement based feedback loops in software development projects. Lecture Notes in Computer Science, 2176:135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760135.
pdf.

Starkloff:2001:PIL

[2388] Philipp Starkloff and Katja Pook.

Process-integrated learning: The AD-VISOR approach for corporate development. Lecture Notes in Computer Science, 2176:152–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2176/21760152.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2176/21760152.pdf.

Holz:2001:TSK

[2389] Harald Holz, Arne Könnecker, and Frank Maurer. Task-specific knowledge management in a process-centred SEE. Lecture Notes in Computer Science, 2176:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2176/21760163.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2176/21760163.
                                        pdf.

Brandt:2001:CSR

[2390] Matthias Brandt and Markus Nick. Computer-supported reuse of project management experience with an experience base. Lecture Notes in Computer Science, 2176:178–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2176/21760178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2176/21760178.
                                        pdf.

Decker:2001:ICI

[2391] Björn Decker and Andreas Jedlitschka. The integrated corporate information network iCoIN: A comprehensive, Webbased experience factory. Lecture Notes in Computer Science, 2176:192– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2176/21760192.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2176/21760192.
                                        pdf.

Differding:2001:RMP

[2392] Christiane M. Differding. Reuse of measurement plans based on process and quality models. Lecture Notes in Computer Science, 2176:207–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760207.
pdf.

Nick:2001:EEB

[2393] Markus Nick and Klaus-Dieter Althoff. Engineering experience base maintenance knowledge. Lecture Notes in Computer Science, 2176:222–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760222.
pdf.

LWAC:2001:ALD

[2394] LSO Workshop Advisory Committee. Announcement of LSO discussion forum. Lecture Notes in Computer Science, 2176:238–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2176/21760238.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2176/21760238.
pdf.

Henninger:2001:AL

[2395] Scott Henninger. Announcement for LSO 2002. Lecture Notes in Computer Science, 2176:239–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2176/21760239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2176/21760239.
                                        pdf.

Anonymous:2001:AIhi

[2396] Anonymous. Author index. Lecture Notes in Computer Science, 2176:241– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2176/2176auth.pdf.

Bassett:2001:TPA

[2397] Paul G. Bassett. The theory and practice of adaptive components. Lecture Notes in Computer Science, 2177:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2177/21770001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2177/21770001.
                                        pdf.

Tilman:2001:DCD

[2398] Michel Tilman. Designing for change, a dynamic perspective. Lecture Notes in Computer Science, 2177:15–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2177/21770015.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2177/21770015.
                                        pdf.

Rashid:2001:AP

[2399] Awais Rashid. On to aspect persistence. Lecture Notes in Computer Science, 2177:26–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2177/21770026.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2177/21770026.pdf.

Coplien:2001:SBS

[2400] James O. Coplien and Liping Zhao.

Symmetry breaking in software patterns. Lecture Notes in Computer Science, 2177:37–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2177/21770037.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2177/21770037.pdf.

Klaeren:2001:ACA

[2401] Herbert Klaeren, Elke Pulvermüller, Awais Rashid, and Andreas Speck. Aspect composition applying the design by contract principle. Lecture Notes in Computer Science, 2177:57– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2177/21770057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2177/21770057.
pdf.

Teschke:2001:TFC

[2402] Thorsten Teschke and Jörg Ritter.

Towards a foundation of componentoriented software reference models. Lecture Notes in Computer Science, 2177:70–??, 2001. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2177/21770070.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2177/21770070.
                                        pdf.

deJonge:2001:GC

[2403] Merijn de Jonge and Joost Visser. Grammars as contracts. Lecture Notes in Computer Science, 2177:85– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2177/21770085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2177/21770085.
                                        pdf.

Becker:2001:GCS

[2404] Martin Becker. Generic components: A symbiosis of paradigms. Lecture Notes in Computer Science, 2177:100– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2177/21770100.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2177/21770100.
                                        pdf.

Goedicke:2001:DIC

[2405] Michael Goedicke, Gustaf Neumann, and Uwe Zdun. Design and implementation constructs for the development of flexible, component-oriented software architectures. Lecture Notes in Computer Science, 2177:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2177/21770114.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2177/21770114.
pdf.

deBruin:2001:SBA

[2406] Hans de Bruin. Scenario-based analysis of component compositions. Lecture Notes in Computer Science, 2177: 129–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2177/21770129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2177/21770129.
pdf.

Bosch:2001:PIS

[2407] Jan Bosch and Mattias Högström.

Product instantiation in software product lines: A case study. Lecture Notes in Computer Science, 2177:147– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2177/21770147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2177/21770147.
pdf.

Smaragdakis:2001:MBP

[2408] Yannis Smaragdakis and Don Batory.

Mixin-based programming in C++. Lecture Notes in Computer Science, 2177:163–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2177/21770163.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2177/21770163.
                                        pdf.

Ludwig:2001:ML

[2409] Andreas Ludwig and Dirk Heuzeroth. Metaprogramming in the large. Lecture Notes in Computer Science, 2177: 178–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2177/21770178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2177/21770178.
                                        pdf.

Veldhuizen:2001:JWY

[2410] Todd L. Veldhuizen. Just when you thought your little language was safe: “expression templates” in Java. Lecture Notes in Computer Science, 2177: 188–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2177/21770188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2177/21770188.
                                        pdf.

Anonymous:2001:AIhj

[2411] Anonymous. Author index. Lecture Notes in Computer Science, 2177:203– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2177/21770203.pdf.

Buchberger:2001:GBS

[2412] Bruno Buchberger. Gröbner bases: A short introduction for systems theorists. Lecture Notes in Computer Science, 2178:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2178/21780001.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2178/21780001.pdf.

Borger:2001:DRS

[2413] Egon Börger. Design for reuse via structuring techniques for ASMs. Lecture Notes in Computer Science, 2178: 20–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780020.
pdf.

Affenzeller:2001:CFC

[2414] Michael Affenzeller, Franz Pichler, and Rudolf Mittelmann. On CAST.FSM computation of hierarchical multilayer networks of automata. Lecture Notes in Computer Science, 2178:36– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780036.
pdf.

Meine:2001:HOD

[2415] Sebastian Meine. A holarchical organized design assistant for hierarchical decompositions. Lecture Notes in Computer Science, 2178:45–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780045.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780045.
                                        pdf.

Jacak:2001:CMI

[2416] Witold Jacak, Karin Pröll, and Stephan Dreiseitl. Conflict management in an intelligent multiagent robotic system based on FSM approach. Lecture Notes in Computer Science, 2178:52–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780052.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780052.
                                        pdf.

Nardo:2001:CAS

[2417] Elvira Di Nardo, Amelia G. Nobile, Enrica Pirozzi, and Luigi M. Ricciardi. Computer-aided simulations of Gaussian processes and related asymptotic properties. Lecture Notes in Computer Science, 2178:67– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780067.
                                        pdf.

Stauner:2001:HSP

[2418] Thomas Stauner. Hybrid systems’ properties — classification and relation to computer science. Lecture Notes in Computer Science, 2178:79– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780079.
pdf.

Abraham-Mumm:2001:ABA

[2419] Erika Ábrahám-Mumm, Ulrich Hannemann, and Martin Steffen. Assertionbased analysis of hybrid systems with PVS. Lecture Notes in Computer Science, 2178:94–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2178/21780094.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2178/21780094.pdf.

Issendorff:2001:ADP

[2420] Hermann von Issendorff. Algebraic description of physical systems. Lecture Notes in Computer Science, 2178:110– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780110.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780110.
pdf.

Shimogawa:2001:ISL

[2421] Takuhei Shimogawa. Information structure and logical design of social rules. Lecture Notes in Computer Science, 2178:125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780125.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780125.
                                        pdf.

Nemeth:2001:TGM

[2422] Gábor Németh. Towards a generic model of information processing systems. Lecture Notes in Computer Science, 2178:139–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2178/21780139.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2178/21780139.pdf.

Hurtado:2001:EID

[2423] Maria V. Hurtado and Jose Parets. Evolutionary information and decision support systems: An integration based on ontologies. Lecture Notes in Computer Science, 2178:146– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780146.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780146.
                                        pdf.

Aguila:2001:PPM

[2424] Isabel María del Aguila, Samuel Túnez, Joaquín Cañadas, Alfonso Bosch, and Roque Marín. A proposal for project management using CommonKADS. Lecture Notes in Computer Science, 2178:160–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780160.
pdf.

Mauerkirchner:2001:GPM

[2425] Manfred Mauerkirchner. A general planning method for allocation of human resource groups. Lecture Notes in Computer Science, 2178:172–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780172.
pdf.

Schwaninger:2001:OOS

[2426] Markus Schwaninger. Optimizing organizational structure: A Hausdorff benchmark for complex social systems. Lecture Notes in Computer Science, 2178:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780182.
pdf.

Paniagua-Aris:2001:PUU

[2427] Enrique Paniagua-Arís, José T.

Palma-Méndez, and Fernando Martín-Rubio. A planning under uncertainty model. Lecture Notes in Computer Science, 2178:196–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780196.
pdf.

Bienvenido:2001:KBM

[2428] Jose F. Bienvenido, Isabel M. Flores-Parra, Rafael Guirado, and Roque L. Marín. Knowledge based modeling of the design process as a base of design tools. application to the development of agricultural structures. Lecture Notes in Computer Science, 2178: 209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780209.
pdf.

Dierneder:2001:CAS

[2429] Stefan Dierneder and Rudolf Scheidl. Complexity analysis of systems from a functional and technical viewpoint. Lecture Notes in Computer Science, 2178:223–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2178/21780223.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2178/21780223.
pdf.

Cobas:2001:CTA

[2430] Juan David González Cobas and José Antonio López Brugos. A complexity-theoretic approach to the design of good measures of cryptographic strength. Lecture Notes in Computer Science, 2178:233–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780233.
pdf.

Campani:2001:CSD

[2431] Carlos A. P. Campani and Paulo Blauth Menezes. Characterizing the software development process: A new approach based on Kolmogorov complexity. Lecture Notes in Computer Science, 2178: 242–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780242.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780242.
pdf.

Scharinger:2001:ASK

[2432] Josef Scharinger. Application of signed Kolmogorov hashes to provide integrity and authenticity in Webbased software distribution. Lecture Notes in Computer Science, 2178:257– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780257.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2178/21780257.
                                        pdf.

Garcia-Cabrera:2001:TFE

[2433] Lina García-Cabrera, María José Rodríguez-Fórtiz, and José Parets-Llorca. Toward a formalisation of evolutionary hypermedia systems based on system theory. Lecture Notes in Computer Science, 2178:270–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780270.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780270.
                                        pdf.

Leal:2001:OPC

[2434] Liara Aparecida dos Santos Leal, Paulo Blauth Menezes, Dalcidio Moraes Claudio, and Laira Vieira Toscani. Optimization problems categories. Lecture Notes in Computer Science, 2178: 285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780285.
                                        pdf.

Moreno-Diaz:2001:SMC

[2435] Roberto Moreno-Díaz and Gabriel de Blasio. Systems methodology for classical neural nets at various levels. Lecture Notes in Computer Science, 2178:300–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2178/21780300.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2178/21780300.
pdf.

Quesada-Arencibia:2001:BBC

[2436] Alexis Quesada-Arencibia, Roberto Moreno-Díaz Jr., and Miguel Aleman-Flores. Biologically based CASTmechanism for visual motion analysis. Lecture Notes in Computer Science, 2178:316–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2178/21780316.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2178/21780316.
pdf.

Araujo:2001:TMV

[2437] Carmen Paz Suarez Araujo, Pablo Fernandez Lopez, and Patricio Garcia Baez. Towards a model of volume transmission in biological and artificial neural networks: A CAST approach. Lecture Notes in Computer Science, 2178:328–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780328.
pdf.

Zorrilla:2001:OIR

[2438] Marta E. Zorrilla, José Luis Crespo, and Eduardo Mora. An online information retrieval systems by means of artificial neural networks. Lecture Notes in Computer Science, 2178:343– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780343.
                                        pdf.

Resconi:2001:KRM

[2439] Germano Resconi, Alberto Borboni, Rodolfo Faglia, and Monica Tiboni. Kinematics in robotics by the morphogenetic neuron. Lecture Notes in Computer Science, 2178:352–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780352.
                                        pdf.

Ceska:2001:AVQ

[2440] Milan Ceska, Vladimír Janousek, and Tomás Vojnar. Analysis and verification queries over object-oriented petri nets. Lecture Notes in Computer Science, 2178:369–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2178/21780369.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2178/21780369.pdf.

Albertos:2001:IDI

[2441] Pedro Albertos and Alfons Crespo. Integrated design and implementation of digital controllers. Lecture Notes in Computer Science, 2178:385– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780385.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780385.
pdf.

Rodriguez-Millan:2001:SCT

[2442] Jesús Rodríguez-Millán. Symbolic computation tools for dynamical nonlinear control systems. Lecture Notes in Computer Science, 2178:393–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780393.
pdf.

Rodriguez-Millan:2001:ISG

[2443] Jesús Rodríguez-Millán. Integrated symbolic-graphic-numeric analysis and design in nonlinear control through notebooks in mathematica. Lecture Notes in Computer Science, 2178:405– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780405.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780405.
pdf.

Frappier:2001:TCA

[2444] Marc Frappier and Richard St-Denis.

Towards a computer-aided design of reactive systems. Lecture Notes in Computer Science, 2178:421–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780421.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780421.
                                        pdf.

Raya:2001:CCS

[2445] Agustin Raya and Dario Maravall. Contributions to the control and stabilization of the pole-cart system. Lecture Notes in Computer Science, 2178: 437–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780437.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780437.
                                        pdf.

Rosa:2001:QSF

[2446] Ricardo García Rosa and Teresa de Pedro. A qualitative system as a frame to control unmanned vehicles. Lecture Notes in Computer Science, 2178:450– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780450.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780450.
                                        pdf.

Machado:2001:HAS

[2447] Júlio P. Machado, Márcia R. Notare, Simone A. Costa, Tiarajú A. Diverio, and P. Blauth Menezes. Hyperautomaton system applied to geometry demonstration environment. Lecture Notes in Computer Science, 2178: 457–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780457.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780457.
pdf.

Buchberger:2001:PPT

[2448] Bruno Buchberger. The PCS prover in THEOREMA. Lecture Notes in Computer Science, 2178:469–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780469.
pdf.

Ida:2001:HOL

[2449] Tetsuo Ida, Mircea Marin, and Taro Suzuki. Higher-order lazy narrowing calculus: A solver for higherorder equations. Lecture Notes in Computer Science, 2178:479–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780479.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780479.
pdf.

Meier:2001:CIR

[2450] Andreas Meier, Martin Pollet, and Volker Sorge. Classifying isomorphic residue classes. Lecture Notes in Computer Science, 2178:494–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780494.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780494.
                                        pdf.

Sato:2001:CCA

[2451] Masahiko Sato, Yukiyoshi Kameyama, and Takeuti Izumi. CAL: A computer assisted learning system for computation and logic. Lecture Notes in Computer Science, 2178:509–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780509.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780509.
                                        pdf.

Windsteiger:2001:STP

[2452] Wolfgang Windsteiger. A set theory prover within theorema. Lecture Notes in Computer Science, 2178:525– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780525.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780525.
                                        pdf.

Barreiro:2001:TLF

[2453] Miguel Barreiro, Víctor M. Gulías, Juan J. Sánchez, and Santiago Jorge. The tertiary level in a functional cluster-based hierarchical VoD server. Lecture Notes in Computer Science, 2178:540–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780540.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780540.
pdf.

Dominguez:2001:HSF

[2454] César Domínguez, Laureano Lambán, Vico Pascual, and Julio Rubio. Hidden specification of a functional system. Lecture Notes in Computer Science, 2178:555–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2178/21780555.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2178/21780555.
pdf.

Bois:2001:DEF

[2455] André Rauber Du Bois and Antônio Carlos da Rocha Costa. Distributed execution of functional programs using the JVM. Lecture Notes in Computer Science, 2178:570–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780570.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780570.
pdf.

Nistal:2001:FC

[2456] José L. Freire Nistal, José E. Freire Brañas, Antonio Blanco Ferro, and Juan J. Sánchez Penas. Fusion in coq.

Lecture Notes in Computer Science, 2178:583–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2178/21780583.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780583.
                                        pdf.

Gayo:2001:LPT

[2457] Jose E. Labra Gayo, Juan M. Cueva Lovelle, María C. Luengo Díez, and Bernardo M. González Rodríguez. A language prototyping tool based on semantic building blocks. Lecture Notes in Computer Science, 2178:597– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780597.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780597.
                                        pdf.

Martin-Mateos:2001:VAA

[2458] Francisco J. Martín-Mateos, Jose A. Alonso, María J. Hidalgo, and José L. Ruiz-Reina. Verifying an applicative ATP using multiset relations. Lecture Notes in Computer Science, 2178:612– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2178/21780612.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2178/21780612.
                                        pdf.

Pareja:2001:ATL

[2459] Cristóbal Pareja, Ricardo Peña, Fernando Rubio, and Clara Segura. Adding traces to a lazy monadic evaluator. Lecture Notes in Computer Science, 2178:627–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780627.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780627.
pdf.

Pointon:2001:FVO

[2460] Robert F. Pointon, Steffen Priebe, Hans-Wolfgang Loidl, Rita Loogen, and Phil W. Trinder. Functional vs object-oriented distributed languages. Lecture Notes in Computer Science, 2178:642–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2178/21780642.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2178/21780642.
pdf.

Quintela:2001:PH

[2461] Juan Quintela and Juan J. Sánchez.

Persistent haskell. Lecture Notes in Computer Science, 2178:657–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2178/21780657.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2178/21780657.
pdf.

Anonymous:2001:AIa

[2462] Anonymous. Author index. Lecture Notes in Computer Science, 2178:669– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2178/2178indx.pdf.

Axelsson:2001:OTL

[2463] O. Axelsson and I. Kaporin. Optimizing two-level preconditionings for the conjugate gradient method. Lecture Notes in Computer Science, 2179:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790003.
                                        pdf.

Garcke:2001:PSG

[2464] J. Garcke and M. Griebel. On the parallelization of the sparse grid approach for data mining. Lecture Notes in Computer Science, 2179:22– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790022.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790022.
                                        pdf.

Getov:2001:JCL

[2465] V. Getov and M. Philippsen. Java communications for large-scale parallel computing. Lecture Notes in Computer Science, 2179:33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2179/21790033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790033.
pdf.

Given:2001:CPB

[2466] J. A. Given, M. Mascagni, and C.-O. Hwang. Continuous path Brownian trajectories for diffusion Monte Carlo via first- and lastpassage distributions. Lecture Notes in Computer Science, 2179:46–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790046.
pdf.

Heinrich:2001:MMC

[2467] S. Heinrich. Multilevel Monte Carlo methods. Lecture Notes in Computer Science, 2179:58–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790058.
pdf.

Marek:2001:IAD

[2468] I. Marek and P. Mayer. Iterative aggregation/disaggregation methods for computing some characteristics of Markov chains. Lecture Notes in Computer Science, 2179:68– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790068.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790068.
                                        pdf.

Zlatev:2001:TIA

[2469] Z. Zlatev. Time-integration algorithms for the computer treatment of the horizontal advection in air pollution models. Lecture Notes in Computer Science, 2179:81–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790081.
                                        pdf.

Georgiev:2001:MPR

[2470] I. Georgiev and S. Margenov. MIC(0) preconditioning of rotated trilinear FEM elliptic systems. Lecture Notes in Computer Science, 2179:95–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790095.
                                        pdf.

Karatson:2001:SSP

[2471] J. Karátson and I. Faragó. Sobolev space preconditioning for mixed nonlinear elliptic boundary value problems. Lecture Notes in Computer Science, 2179:104–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2179/21790104.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2179/21790104.
pdf.

Neytcheva:2001:SCA

[2472] M. Neytcheva and O. Axelsson. On a Schur complement approach for solving two-level finite element systems. Lecture Notes in Computer Science, 2179: 113–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790113.
pdf.

Angelova:2001:LSE

[2473] D. Angelova, E. Semerdjiev, Tz.

Semerdjiev, and P. Konstantinova. On-line state estimation of maneuvering objects by sequential Monte Carlo algorithm. Lecture Notes in Computer Science, 2179:125–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790125.
pdf.

Atanassov:2001:DDS

[2474] E. I. Atanassov. On the dyadic diaphony of the Sobol’ sequences. Lecture Notes in Computer Science, 2179: 133–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790133.
                                        pdf.

Dimov:2001:IMC

[2475] I. T. Dimov, E. I. Atanassov, and M. K. Durchova. An improved Monte Carlo algorithm for elastic electron backscattering from surfaces. Lecture Notes in Computer Science, 2179:141– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790141.
                                        pdf.

Gurov:2001:SAS

[2476] T. V. Gurov and P. A. Whitlock. Statistical algorithms for simulation of electron quantum kinetics in semiconductors — Part I. Lecture Notes in Computer Science, 2179:149–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790149.
                                        pdf.

Karaivanova:2001:QMC

[2477] A. Karaivanova, I. Dimov, and S. Ivanovska. A quasi-Monte Carlo method for integration with improved convergence. Lecture Notes in Computer Science, 2179:158–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790158.
pdf.

Karaivanova:2001:SSL

[2478] A. Karaivanova and R. Georgieva.

Solving systems of linear algebraic equations using quasirandom numbers. Lecture Notes in Computer Science, 2179:166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2179/21790166.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2179/21790166.
pdf.

Kosina:2001:MCA

[2479] H. Kosina, M. Nedjalkov, and S. Selberherr. Monte Carlo analysis of the small-signal response of charge carriers. Lecture Notes in Computer Science, 2179:175–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790175.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790175.
pdf.

Nedjalkov:2001:SAS

[2480] M. Nedjalkov, T. V. Gurov, H. Kosina, and P. A. Whitlock. Statistical algorithms for simulation of electron quantum kinetics in semiconductors — Part II. Lecture Notes in Computer Science, 2179:183–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790183.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790183.
                                        pdf.

Baude:2001:IIC

[2481] F. Baude, A. Bergel, D. Caromel, F. Huet, O. Nano, and J. Vayssière. IC2D: Interactive control and debugging of distribution. Lecture Notes in Computer Science, 2179:193–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790193.
                                        pdf.

Budimlic:2001:JJC

[2482] Zoran Budimlić and Ken Kennedy. JaMake: A Java compiler environment. Lecture Notes in Computer Science, 2179:201–209, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2179/21790201.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2179/21790201.pdf.

Chapman:2001:PDE

[2483] B. Chapman, O. Hernandez, A. Patil, and A. Prabhakar. Program development environment for OpenMP programs on ccNUMA architectures. Lecture Notes in Computer Science, 2179: 210–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790210.
pdf.

Germain:2001:GCS

[2484] C. Germain, G. Fedak, V. Néri, and F. Cappello. Global computing systems. Lecture Notes in Computer Science, 2179:218–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790218.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790218.
pdf.

Krall:2001:JLS

[2485] A. Krall and P. Tomsich. Java for largescale scientific computations? Lecture Notes in Computer Science, 2179:228– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790228.
pdf.

Sips:2001:JSC

[2486] H. J. Sips and K. van Reeuwijk. Java for scientific computation: Prospects and problems. Lecture Notes in Computer Science, 2179:236–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790236.
                                        pdf.

Antonov:2001:OOF

[2487] A. Antonov. Object-oriented framework for large scale air pollution modeling. Lecture Notes in Computer Science, 2179:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790247.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790247.
                                        pdf.

Ebel:2001:ERM

[2488] A. Ebel. Evaluation and reliability of meso-scale air pollution simulations. Lecture Notes in Computer Science, 2179:255–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790255.
                                        pdf.

Farago:2001:MBO

[2489] I. Faragó and Á. Havasi. The mathematical background of operator splitting and the effect of non-commutativity. Lecture Notes in Computer Science, 2179:264–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790264.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790264.
pdf.

Georgiev:2001:FGR

[2490] K. Georgiev and Z. Zlatev. Fine-grid resolution in Danish Eulerian model and an implementation on SGI Origin 2000 computer. Lecture Notes in Computer Science, 2179:272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790272.
pdf.

Kallos:2001:MFA

[2491] G. Kallos, A. Voudouri, I. Pytharoulis, and O. Kakaliagou. Modelling framework for atmospheric mercury over the Mediterranean region: Model development and applications. Lecture Notes in Computer Science, 2179:281– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790281.
pdf.

Karagiorgos:2001:ILB

[2492] G. Karagiorgos and N. M. Missirlis.

Iterative load balancing schemes for air pollution models. Lecture Notes in Computer Science, 2179:291–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790291.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790291.
pdf.

Mensink:2001:CAA

[2493] C. Mensink, K. De Ridder, N. Lewyckyj, L. Delobbe, L. Janssen, and Ph. Van Haver. Computational aspects of air quality modelling in urban regions using an optimal resolution approach (AURORA). Lecture Notes in Computer Science, 2179:299–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790299.
pdf.

Ostromsky:2001:PIL

[2494] T. Ostromsky and Z. Zlatev. Parallel implementation of a large-scale 3-D air pollution model. Lecture Notes in Computer Science, 2179:309– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790309.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790309.
pdf.

Syrakov:2001:LTE

[2495] D. Syrakov, M. Prodanova, Ch. Zerefos, K. Ganev, and N. Miloshev. Longterm estimates of sulfur deposition in the region of Southeastern Europe. Lecture Notes in Computer Science, 2179: 317–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790317.
pdf.

Wang:2001:CEA

[2496] S. W. Wang, P. G. Georgopoulos, G. Li, and H. Rabitz. Computationally efficient atmospheric chemical kinetic modeling by means of high dimensional model representation (HDMR). Lecture Notes in Computer Science, 2179: 326–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790326.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790326.
pdf.

Denev:2001:CSA

[2497] J. A. Denev, D. G. Markov, and P. Stankov. Computer simulation of the air flow and the distribution of combustion generated pollutants around buildings. Lecture Notes in Computer Science, 2179:337– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790337.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790337.
                                        pdf.

Drikakis:2001:MMC

[2498] D. Drikakis, O. Iliev, and D. Vassileva. On multigrid methods for the compressible Navier–Stokes equations. Lecture Notes in Computer Science, 2179: 344–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790344.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790344.
                                        pdf.

Hoppe:2001:SOB

[2499] R. H. W. Hoppe and S. I. Petrova. Structural optimization of biomorphic microcellular ceramics by homogenization approach. Lecture Notes in Computer Science, 2179:353–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790353.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790353.
                                        pdf.

Iliev:2001:MAL

[2500] Oleg Iliev and Dimitar Stoyanov. Multigrid — adaptive local refinement solver for incompressible flows. Lecture Notes in Computer Science, 2179: 361–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790361.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2179/21790361.
pdf.

Langemann:2001:HSN

[2501] D. Langemann. Handling systems from non-linear theory of elasticity. Lecture Notes in Computer Science, 2179:369– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790369.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790369.
pdf.

Mitkova:2001:NMF

[2502] T. Mitkova and L. Tobiska. Numerical modelling of the flow in magnetic liquid seals. Lecture Notes in Computer Science, 2179:378–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790378.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790378.
pdf.

Sieber:2001:DMS

[2503] R. Sieber and M. Schäfer. Dynamic mesh schemes for fluid– structure interaction. Lecture Notes in Computer Science, 2179:387–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790387.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2179/21790387.
pdf.

Bazhlekov:2001:BIM

[2504] I. B. Bazhlekov, F. N. van de Vosse, and H. E. H. Meijer. Boundary integral method for 3D simulation of foam dynamics. Lecture Notes in Computer Science, 2179:401–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790401.
pdf.

Diaz-Goano:2001:LMF

[2505] C. Diaz-Goano, P. Minev, and K. Nandakumar. A Lagrange multipliers/fictitious domain approach for particulate flow. Lecture Notes in Computer Science, 2179:409–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790409.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790409.
pdf.

Maryska:2001:PVD

[2506] Jiří Maryška, Miroslav Rozložník, and Miroslav Tůma. Primal vs. dual variable approach for mixedhybrid finite element approximation of the potential fluid flow problem in porous media. Lecture Notes in Computer Science, 2179:417–424, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790417.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790417.
pdf.

Miladinova:2001:WEH

[2507] S. Miladinova. Wave evolution of heated falling films. numerical analysis using finite-difference method. Lecture Notes in Computer Science, 2179: 425–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790425.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790425.
pdf.

Vijfvinkel:2001:PUS

[2508] L. W. Vijfvinkel. Picard-Uzawa schemes: Errors, convergence and stopping criterion. Lecture Notes in Computer Science, 2179:433–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790433.
pdf.

Andreev:2001:FEM

[2509] A. B. Andreev, J. T. Maximov, and M. R. Racheva. Finite element method for plates with dynamic loads. Lecture Notes in Computer Science, 2179: 445–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790445.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790445.
                                        pdf.

Bencheva:2001:MPI

[2510] G. Bencheva. MPI parallel implementation of a fast separable solver. Lecture Notes in Computer Science, 2179: 454–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790454.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790454.
                                        pdf.

Brandts:2001:CSM

[2511] J. Brandts. A comparison of subspace methods for Sylvester equations. Lecture Notes in Computer Science, 2179: 462–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2179/21790462.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2179/21790462.
                                        pdf.

Dent:2001:SPN

[2512] Deborah Dent, Marcin Paprzycki, Anna Kucaba-Pietal, and Ludomir Laudański. Studying the performance nonlinear systems solvers applied to the random vibration test. Lecture Notes in Computer Science, 2179:471– 478, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2179/21790471.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790471.
pdf.

Dimitriu:2001:PIT

[2513] G. Dimitriu. Parameter identification in a two-dimensional parabolic equation using an ADI based solver. Lecture Notes in Computer Science, 2179: 479–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790479.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790479.
pdf.

Geiser:2001:NSM

[2514] J. Geiser. Numerical simulation of a model for transport and reaction of radionuclides. Lecture Notes in Computer Science, 2179:487–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2179/21790487.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2179/21790487.
pdf.

Anonymous:2001:AIhk

[2515] Anonymous. Author index. Lecture Notes in Computer Science, 2179:497– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2179/2179auth.pdf.

Kim:2001:TCB

[2516] Yong-Jik Kim and James H. Anderson. A time complexity bound for adaptive mutual exclusion (extended abstract). Lecture Notes in Computer Science, 2180:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800001.
pdf.

Joung:2001:QBA

[2517] Yuh-Jzer Joung. Quorum-based algorithms for group mutual exclusion. Lecture Notes in Computer Science, 2180: 16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800016.
pdf.

Boldi:2001:ECC

[2518] Paolo Boldi and Sebastiano Vigna. An effective characterization of computability in anonymous networks. Lecture Notes in Computer Science, 2180: 33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800033.
pdf.

Douceur:2001:CHC

[2519] John R. Douceur and Roger P. Wattenhofer. Competitive hill-climbing strategies for replica placement in a distributed file system. Lecture Notes in Computer Science, 2180:48– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800048.
pdf.

Malkhi:2001:OUI

[2520] Dahlia Malkhi, Elan Pavlov, and Yaron Sella. Optimal unconditional information diffusion. Lecture Notes in Computer Science, 2180:63–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800063.
pdf.

Mittal:2001:CST

[2521] Neeraj Mittal and Vijay K. Garg.

Computation slicing: Techniques and theory. Lecture Notes in Computer Science, 2180:78–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2180/21800078.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2180/21800078.pdf.

Jimenez-Peris:2001:LLN

[2522] R. Jiménez-Peris, M. Patiño-Martínez, G. Alonso, and S. Arévalo. A lowlatency non-blocking commit service. Lecture Notes in Computer Science, 2180:93–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2180/21800093.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2180/21800093.
pdf.

Aguilera:2001:SLE

[2523] Marcos K. Aguilera, Carole Delporte-Gallet, Hugues Fauconnier, and Sam Toueg. Stable leader election. Lecture Notes in Computer Science, 2180:108– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800108.
pdf.

Inoue:2001:ALL

[2524] Michiko Inoue, Shinya Umetani, Toshimitsu Masuzawa, and Hideo Fujiwara. Adaptive long-lived O(k^{2})- renaming with O(k^{2}) steps. Lecture Notes in Computer Science, 2180:123– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800123.
pdf.

Herlihy:2001:NSL

[2525] Maurice Herlihy, Sergio Rajsbaum, and Mark Tuttle. A new synchronous lower bound for set agreement. Lecture Notes in Computer Science, 2180: 136–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800136.
pdf.

Georgiou:2001:CSI

[2526] Chryssis Georgiou, Alexander Russell, and Alex A. Shvartsman. The complexity of synchronous iterative do-all with crashes. Lecture Notes in Computer Science, 2180:151–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800151.
pdf.

Dobrev:2001:MSB

[2527] Stefan Dobrev, Paola Flocchini, Giuseppe Prencipe, and Nicola Santoro. Mobile search for a black hole in an anonymous ring. Lecture Notes in Computer Science, 2180:166– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800166.
pdf.

Hoepman:2001:RMS

[2528] Jaap-Henk Hoepman. Randomised mutual search for k > 2 agents. Lecture Notes in Computer Science, 2180:180– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2180/21800180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2180/21800180.
                                        pdf.

Higham:2001:SSM

[2529] Lisa Higham and Zhiying Liang. Selfstabilizing minimum spanning tree construction on message-passing networks. Lecture Notes in Computer Science, 2180:194–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2180/21800194.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2180/21800194.
                                        pdf.

Herlihy:2001:SSD

[2530] Maurice Herlihy and Srikanta Tirthapura. Self stabilizing distributed queuing. Lecture Notes in Computer Science, 2180:209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2180/21800209.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2180/21800209.pdf.

Fich:2001:SOD

[2531] Faith E. Fich and Colette Johnen. A space optimal, deterministic, selfstabilizing, leader election algorithm for unidirectional rings. Lecture Notes in Computer Science, 2180:224–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800224.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800224.
pdf.

Duflot:2001:RFS

[2532] Marie Duflot, Laurent Fribourg, and Claudine Picaronny. Randomized finite-state distributed algorithms as Markov chains. Lecture Notes in Computer Science, 2180:240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800240.
pdf.

Peleg:2001:AHC

[2533] David Peleg and Uri Pincas. The average hop count measure for virtual path layouts. Lecture Notes in Computer Science, 2180:255–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800255.
pdf.

Barriere:2001:ERN

[2534] Lali Barrière, Pierre Fraigniaud, Evangelos Kranakis, and Danny Krizanc. Efficient routing in networks with long range contacts. Lecture Notes in Computer Science, 2180:270–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2180/21800270.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2180/21800270.
                                        pdf.

Chatzigiannakis:2001:ECS

[2535] I. Chatzigiannakis, S. Nikoletseas, and P. Spirakis. An efficient communication strategy for ad-hoc mobile networks. Lecture Notes in Computer Science, 2180:285–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2180/21800285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2180/21800285.
                                        pdf.

Harris:2001:PIN

[2536] Timothy L. Harris. A pragmatic implementation of non-blocking linked-lists. Lecture Notes in Computer Science, 2180:300–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2180/21800300.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2180/21800300.
                                        pdf.

Herman:2001:SRS

[2537] Ted Herman and Toshimitsu Masuzawa. Stabilizing replicated search trees. Lecture Notes in Computer Science, 2180:315–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2180/21800315.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2180/21800315.pdf.

Fatourou:2001:AN

[2538] Panagiota Fatourou and Maurice Herlihy. Adding networks. Lecture Notes in Computer Science, 2180:330– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2180/21800330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2180/21800330.
pdf.

Anonymous:2001:AIhl

[2539] Anonymous. Author index. Lecture Notes in Computer Science, 2180:343– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2180/2180auth.pdf.

Brunner:2001:ZAE

[2540] Albert Brunner. Zürich Airport Extension Project: Digital support for earthwork construction. Lecture Notes in Computer Science, 2181:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810001.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2181/21810001.
                                        pdf.

Ervin:2001:DL

[2541] Stephen M. Ervin. Designed landforms. Lecture Notes in Computer Science, 2181:2–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2181/21810002.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2181/21810002.
                                        pdf.

Huber:2001:TAC

[2542] Gary Huber. Terraffectors and the art of consensus building. Lecture Notes in Computer Science, 2181:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2181/21810003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2181/21810003.
                                        pdf.

Koehrsen:2001:GBE

[2543] Craig L. Koehrsen, William C. Sahm, and Claude W. Keefer. GPS-based earthmoving for construction. Lecture Notes in Computer Science, 2181:4– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2181/21810004.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2181/21810004.
                                        pdf.

Raviola:2001:FLR

[2544] Alessandro Raviola, Michela Spagnuolo, and Giuseppe Patané. Feature lines reconstruction for reverse engineering. Lecture Notes in Computer Science, 2181:18–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2181/21810018.
pdf.

Mezosi:2001:GMR

[2545] Gábor Mezősi and Richárd Kiss. Geomorphometrical mapping of reliefdissection using GIS. Lecture Notes in Computer Science, 2181:31–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2181/21810031.
pdf.

Schoenstein:2001:ALD

[2546] Michael Schoenstein. AutoCAD Land Development Desktop Release 2i. Lecture Notes in Computer Science, 2181: 39–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2181/21810039.
pdf.

Widmark:2001:PGR

[2547] Mattias Widmark. Procedural geometry for real-time terrain visualisation in Blueberry3D. Lecture Notes in Computer Science, 2181:43– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2181/21810043.
pdf.

Hansen:2001:QFD

[2548] Henning Sten Hansen. A quasifour dimensional database for the built environment. Lecture Notes in Computer Science, 2181:48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2181/21810048.
pdf.

Weber:2001:IGD

[2549] Angelika Weber and Joachim Benner. Interactive generation of digital terrain models using multiple data sources. Lecture Notes in Computer Science, 2181:60–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810060.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2181/21810060.
pdf.

Hennig:2001:SRT

[2550] Thomas A. Hennig, Jeffrey L. Kretsch, Charles J. Pessagno, Paul H. Salamonowicz, and William L. Stein. The shuttle radar topography mission. Lecture Notes in Computer Science, 2181: 65–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2181/21810065.
pdf.

Westort:2001:CEO

[2551] Caroline Y. Westort. Corner, end, and overlap “extrusion junctures”: Parameters for geometric control. Lecture Notes in Computer Science, 2181:78– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2181/21810078.
pdf.

Rognant:2001:CLG

[2552] L. Rognant, J. G. Planès, M. Memier, and J. M. Chassery. Contour lines and DEM: Generation and extraction. Lecture Notes in Computer Science, 2181: 87–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2181/21810087.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2181/21810087.
pdf.

Hammes:2001:MED

[2553] Johan Hammes. Modeling of ecosystems as a data source for real-time terrain rendering. Lecture Notes in Computer Science, 2181:98–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2181/21810098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2181/21810098.
                                        pdf.

Korpershoek:2001:SIN

[2554] Ewout Korpershoek. Seamless integration of new control technologies. Lecture Notes in Computer Science, 2181: 112–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2181/21810112.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2181/21810112.
                                        pdf.

Anonymous:2001:AIhm

[2555] Anonymous. Author index. Lecture Notes in Computer Science, 2181:117– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2181/2181auth.pdf.

Kitamura:2001:III

[2556] Yasuhiko Kitamura, Teruhiro Yamada, Takashi Kokubo, Yasuhiro Mawarimichi, Taizo Yamamoto, and Toru Ishida. Interactive integration of information agents on the Web. Lecture Notes in Computer Science, 2182:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820001.
pdf.

Imbert:2001:ICV

[2557] Ricardo Imbert, Angélica de Antonio, and Javier Segovia. Improving communication in 3D virtual environments by means of task delegation in agents. Lecture Notes in Computer Science, 2182: 14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820014.
pdf.

Okamoto:2001:WOM

[2558] Masayuki Okamoto, Yeonsoo Yang, and Toru Ishida. Wizard of Oz method for learning dialog agents. Lecture Notes in Computer Science, 2182:20– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820020.
pdf.

Palopoli:2001:SUP

[2559] Luigi Palopoli, Domenico Rosaci, Giorgio Terracina, and Domenico Ursino. Supporting user-profiled semantic Web-oriented search. Lecture Notes in Computer Science, 2182:26– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820026.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820026.
                                        pdf.

Soo:2001:RTP

[2560] Von-Wun Soo and Shu-Hau Liang. Recommending a trip plan by negotiation with a software travel agent. Lecture Notes in Computer Science, 2182:32– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820032.
                                        pdf.

Kanawati:2001:CCB

[2561] Rushed Kanawati and Maria Malek. CoWing: A collaborative bookmark management system. Lecture Notes in Computer Science, 2182:38–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820038.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820038.
                                        pdf.

Plaza:2001:CAA

[2562] Enric Plaza and Josep-Lluís Arcos. Context aware agents for personal information services. Lecture Notes in Computer Science, 2182:44–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820044.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820044.
pdf.

Jarke:2001:DWQ

[2563] Matthias Jarke. Data warehouse quality and agent technology. Lecture Notes in Computer Science, 2182:56– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820056.
pdf.

Lee:2001:DTM

[2564] Doheon Lee, Mina Jeong, and Yong-Kwan Won. Decision trees for multiple abstraction levels of data. Lecture Notes in Computer Science, 2182:76– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820076.
pdf.

Bergamaschi:2001:SII

[2565] S. Bergamaschi, G. Cabri, F. Guerra, L. Leonardi, M. Vincini, and F. Zambonelli. Supporting information integration with autonomous agents. Lecture Notes in Computer Science, 2182:

88–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820088.
                                        pdf.

Nitto:2001:UAP

[2566] Elisabetta Di Nitto, Carlo Ghezzi, Maurizio Sabba, and Paolo Selvini. Using agents in performing multi-site queries. Lecture Notes in Computer Science, 2182:100–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2182/21820100.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2182/21820100.pdf.

Decker:2001:EMA

[2567] Keith Decker, Salim Khan, Carl Schmidt, and Dennis Michaud. Extending a multi-agent system for genomic annotation. Lecture Notes in Computer Science, 2182:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820106.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820106.
                                        pdf.

Gomez:2001:DIO

[2568] Mario Gomez, Chema Abasolo, and Enric Plaza. Domain-independent ontologies for cooperative information agents. Lecture Notes in Computer Science, 2182:118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820118.
pdf.

Fornara:2001:ABA

[2569] Nicoletta Fornara and Luca Maria Gambardella. An autonomous bidding agent for simultaneous auctions. Lecture Notes in Computer Science, 2182: 130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820130.
pdf.

Shehory:2001:ORP

[2570] Onn Shehory. Optimality and risk in purchase from multiple auctions. Lecture Notes in Computer Science, 2182: 142–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820142.
pdf.

Brandt:2001:CPS

[2571] Felix Brandt. Cryptographic protocols for secure second-price auctions. Lecture Notes in Computer Science, 2182: 154–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820154.
                                        pdf.

Goldman:2001:ESS

[2572] Claudia V. Goldman, Sarit Kraus, and Onn Shehory. Equilibria strategies for selecting sellers and satisfying buyers. Lecture Notes in Computer Science, 2182:166–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820166.
                                        pdf.

Brito:2001:LAA

[2573] Luís Brito, Paulo Novais, and José Neves. On the logical aspects of argument-based negotiation among agents. Lecture Notes in Computer Science, 2182:178–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820178.
                                        pdf.

Manouselis:2001:IMA

[2574] Nikos Manouselis and Nikos F. Matsatsinis. Introducing a multi-agent, multi-criteria methodology for modeling electronic consumer’s behavior: The case of Internet radio. Lecture Notes in Computer Science, 2182:190– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820190.
pdf.

Molina:2001:MCK

[2575] Martin Molina. Modeling commercial knowledge to develop advanced agentbased marketplaces for E-commerce. Lecture Notes in Computer Science, 2182:196–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2182/21820196.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2182/21820196.
pdf.

Carlsson:2001:ARW

[2576] Bengt Carlsson and Rune Gustavsson. Arms race within information ecosystems. Lecture Notes in Computer Science, 2182:202–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820202.
pdf.

Castelfranchi:2001:IAS

[2577] Cristiano Castelfranchi. Information agents: The social nature of information and the role of trust. Lecture Notes in Computer Science, 2182:208– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820208.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820208.
                                        pdf.

Iwao:2001:FEI

[2578] Tadashige Iwao, Yuji Wada, Makoto Okada, and Makoto Amamiya. A framework for the exchange and installation of protocols in a multi-agent system. Lecture Notes in Computer Science, 2182:211–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2182/21820211.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2182/21820211.pdf.

Bailin:2001:ONB

[2579] Sidney C. Bailin and Walt Truszkowski. Ontology negotiation as a basis for opportunistic cooperation between intelligent information agents. Lecture Notes in Computer Science, 2182:223– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820223.
                                        pdf.

Hadad:2001:MTR

[2580] Meirav Hadad and Sarit Kraus. A mechanism for temporal reasoning by collaborative agents. Lecture Notes in Computer Science, 2182:229–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820229.
pdf.

Ricci:2001:ACI

[2581] Alessandro Ricci, Enrico Denti, and Andrea Omicini. Agent coordination infrastructures for virtual enterprises and workflow management. Lecture Notes in Computer Science, 2182:235– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2182/21820235.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2182/21820235.
pdf.

Picco:2001:MAS

[2582] Gian Pietro Picco. Mobile agents:

State of the art and research opportunities. Lecture Notes in Computer Science, 2182:247–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2182/21820247.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2182/21820247.pdf.

Adorni:2001:EFA

[2583] Giovanni Adorni, Federico Bergenti, Agostino Poggi, and Giovanni Rimassa. Enabling FIPA agents on small devices. Lecture Notes in Computer Science, 2182:248–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820248.
                                        pdf.

Helin:2001:TER

[2584] Heikki Helin and Mikko Laukkanen. Towards efficient and reliable agent communication in wireless environments. Lecture Notes in Computer Science, 2182:258–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820258.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820258.
                                        pdf.

Finin:2001:IAM

[2585] Tim Finin, Anupam Joshi, Lalana Kagal, Olga Ratsimore, Vlad Korolev, and Harry Chen. Information agents for mobile and embedded devices. Lecture Notes in Computer Science, 2182:264– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2182/21820264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2182/21820264.
                                        pdf.

Anonymous:2001:AIhn

[2586] Anonymous. Author index. Lecture Notes in Computer Science, 2182:287– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2182/2182auth.pdf.

Anonymous:2001:AIho

[2587] Anonymous. Author index. Lecture Notes in Computer Science, 2183:000– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2183/2183auth.pdf.

Aehlig:2001:LRH

[2588] Klaus Aehlig, Jan Johannsen, Helmut Schwichtenberg, and Sebastiaan A. Terwijn. Linear ramified higher type recursion and parallel complexity. Lecture Notes in Computer Science, 2183: 1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830001.
pdf.

Alt:2001:RC

[2589] Jesse Alt and Sergei Artemov. Reflective λ-calculus. Lecture Notes in Computer Science, 2183:22–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830022.
pdf.

Baaz:2001:NPT

[2590] Matthias Baaz and Christian G. Fermüller. A note on the prooftheoretic strength of a single application of the schema of identity. Lecture Notes in Computer Science, 2183:38– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830038.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830038.
pdf.

Baaz:2001:CCC

[2591] Matthias Baaz and Alexander Leitsch. Comparing the complexity of cutelimination methods. Lecture Notes in Computer Science, 2183:49–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830049.
pdf.

Berger:2001:PEG

[2592] Ulrich Berger. Program extraction from Gentzen’s proof of transfinite induction up to ϵ_{0}. Lecture Notes in Computer Science, 2183:68– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830068.
pdf.

Dosen:2001:CBS

[2593] Kosta Došen and Zoran Petrić. Coherent bicartesian and sesquicartesian categories. Lecture Notes in Computer Science, 2183:78–92, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830078.
pdf.

Dybjer:2001:IIR

[2594] Peter Dybjer and Anton Setzer. Indexed induction-recursion. Lecture Notes in Computer Science, 2183:93– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830093.
pdf.

Elbl:2001:MML

[2595] Birgit Elbl. Modeling meta-logical features in a calculus with frozen variables. Lecture Notes in Computer Science, 2183:114–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2183/21830114.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2183/21830114.pdf.

Gordeew:2001:PTP

[2596] Lew Gordeew. Proof theory and Post-Turing analysis. Lecture Notes in Computer Science, 2183:130–152, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830130.
pdf.

Matthes:2001:IND

[2597] Ralph Matthes. Interpolation for natural deduction with generalized eliminations. Lecture Notes in Computer Science, 2183:153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830153.
pdf.

Oitavem:2001:ICP

[2598] Isabel Oitavem. Implicit characterizations of pspace. Lecture Notes in Computer Science, 2183:170–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830170.
pdf.

Schmitt:2001:IL

[2599] Peter H. Schmitt. Iterate logic. Lecture Notes in Computer Science, 2183: 191–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2183/21830191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830191.
pdf.

Studer:2001:CFF

[2600] Thomas Studer. Constructive foundations for featherweight Java. Lecture Notes in Computer Science, 2183:202– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2183/21830202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2183/21830202.
pdf.

Chang:2001:MOG

[2601] Shi-Kuo Chang. MAWC operations for the growing book. Lecture Notes in Computer Science, 2184:3– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840003.
pdf.

Sclaroff:2001:MM

[2602] Stan Sclaroff, George Kollios, Margrit Betke, and Romer Rosales. Motion mining. Lecture Notes in Computer Science, 2184:16–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2184/21840016.
htm;   http://link.springer-ny.

com/link/service/series/0558/papers/
2184/21840016.pdf.

Abate:2001:RMR

[2603] Andrea F. Abate, Giuseppe Sasso, Anthony C. Donadio, and Francesco S. Sasso. The riddles of murecine: The role of anthropological research by images and visual computing. Lecture Notes in Computer Science, 2184:33– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840033.
pdf.

Balestrieri:2001:IRM

[2604] Angelo Balestrieri, Antonella Balestrieri, Aristide Barone, Andrea Casanova, and Matteo Fraschini. Information retrieval from medical database. Lecture Notes in Computer Science, 2184:42– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840042.
pdf.

Costagliola:2001:MCT

[2605] Gennaro Costagliola, Rita Francese, Annamaria Musto, and Giuseppe Polese. M-CODD: A technique for the conceptual design of multimedia databases. Lecture Notes in Computer Science, 2184:55–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840055.
pdf.

Jun:2001:CCA

[2606] Woochun Jun. Controlling concurrent accesses in multimedia database systems. Lecture Notes in Computer Science, 2184:67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2184/21840067.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2184/21840067.pdf.

Maresca:2001:TDM

[2607] Paolo Maresca, Angela Guercio Tim Arndt, and Pasquale Donadio. Transformation dataflow in multimedia software engineering using TAO XML: A component-based approach. Lecture Notes in Computer Science, 2184:77– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840077.
pdf.

Paolino:2001:MEP

[2608] Luca Paolino, Monica Sebillo, Genoveffa Tortora, and Giuliana Vitiello. Monitoring electromagnetic pollution: A GIS-based visual approach. Lecture Notes in Computer Science, 2184:90– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2184/21840090.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2184/21840090.
                                        pdf.

Celentano:2001:QBM

[2609] Augusto Celentano and Ombretta Gaggi. Querying and browsing multimedia presentations. Lecture Notes in Computer Science, 2184:105–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2184/21840105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2184/21840105.
                                        pdf.

Cibelli:2001:CIR

[2610] Maurizio Cibelli, Michele Nappi, and Maurizio Tucci. Content image retrieval: A new query paradigm toward analogy based indexing. Lecture Notes in Computer Science, 2184:117– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2184/21840117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2184/21840117.
                                        pdf.

Distasi:2001:RIR

[2611] Riccardo Distasi and Sergio Vitulano. Robust image retrieval based on texture information. Lecture Notes in Computer Science, 2184:130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840130.
pdf.

Jungert:2001:STC

[2612] Erland Jungert and Fredrik Lantz.

Symbolic tile categories for spatial reasoning on objects represented in multiple-resolution. Lecture Notes in Computer Science, 2184:140–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840140.
pdf.

Kosch:2001:CBI

[2613] Harald Kosch, Mario Döller, and László Böszörményi. Content-based indexing and retrieval supported by mobile agent technology. Lecture Notes in Computer Science, 2184:152–166, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840152.
pdf.

Tian:2001:DOI

[2614] Qi Tian, Baback Moghaddam, and Thomas S. Huang. Display optimization for image browsing. Lecture Notes in Computer Science, 2184:167– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2184/21840167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2184/21840167.
                                        pdf.

Assfalg:2001:SCV

[2615] Jürgen Assfalg, Marco Bertini, Carlo Colombo, and Alberto Del Bimbo. Semantic characterization of visual content for sports videos annotation. Lecture Notes in Computer Science, 2184: 179–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2184/21840179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2184/21840179.
                                        pdf.

Santo:2001:DSD

[2616] Massimo De Santo, Gennaro Percannella, Carlo Sansone, and Mario Vento. Dialogue scenes detection in MPEG movies: A multi-expert approach. Lecture Notes in Computer Science, 2184: 192–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2184/21840192.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2184/21840192.
                                        pdf.

Kato:2001:PQC

[2617] Yuka Kato and Katsuya Hakozaki. A proposal of a QoS control method for digital video libraries. Lecture Notes in Computer Science, 2184:202– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840202.
pdf.

Lee:2001:EVC

[2618] Insuh Lee and Byeungwoo Jeon. Enhanced video communication using gaze correction with simplified 3D warping and single camera. Lecture Notes in Computer Science, 2184:213– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2184/21840213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2184/21840213.
pdf.

Anonymous:2001:AIhp

[2619] Anonymous. Author index. Lecture Notes in Computer Science, 2184:225– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2184/21840225.pdf.

Rumbaugh:2001:PA

[2620] Jim Rumbaugh. The preacher at arrakeen. Lecture Notes in Computer Science, 2185:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2185/21850001.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2185/21850001.pdf.

Alvarez:2001:ASM

[2621] José M. Álvarez, Tony Clark, Andy Evans, and Paul Sammut. An action semantics for MML. Lecture Notes in Computer Science, 2185:2– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2185/21850002.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2185/21850002.
                                        pdf.

Atkinson:2001:EMM

[2622] Colin Atkinson and Thomas Kühne. The essence of multilevel metamodeling. Lecture Notes in Computer Science, 2185:19–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2185/21850019.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2185/21850019.pdf.

Alvarez:2001:MBL

[2623] José Álvarez, Andy Evans, and Paul Sammut. Mapping between levels in the metamodel architecture. Lecture Notes in Computer Science, 2185:34– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2185/21850034.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2185/21850034.
                                        pdf.

Eshuis:2001:EAU

[2624] Rik Eshuis and Roel Wieringa.

An execution algorithm for UML activity graphs. Lecture Notes in Computer Science, 2185:47–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850047.
pdf.

Xuandong:2001:TAU

[2625] Li Xuandong, Cui Meng, Pei Yu, Zhao Jianhua, and Zheng Guoliang. Timing analysis of UML activity diagrams. Lecture Notes in Computer Science, 2185:62–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850062.
pdf.

Dumas:2001:UAD

[2626] Marlon Dumas and Arthur H. M. ter Hofstede. UML activity diagrams as a workflow specification language. Lecture Notes in Computer Science, 2185: 76–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850076.
pdf.

Akehurst:2001:QUD

[2627] D. H. Akehurst and B. Bordbar. On querying UML data models with OCL. Lecture Notes in Computer Science, 2185:91–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2185/21850091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2185/21850091.
                                        pdf.

Demuth:2001:OSL

[2628] Birgit Demuth, Heinrich Hussmann, and Sten Loecher. OCL as a specification language for business rules in database applications. Lecture Notes in Computer Science, 2185:104– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2185/21850104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2185/21850104.
                                        pdf.

Cengarle:2001:FSO

[2629] María Victoria Cengarle and Alexander Knapp. A formal semantics for OCL 1.4. Lecture Notes in Computer Science, 2185:118–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2185/21850118.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2185/21850118.
                                        pdf.

Sunye:2001:RUM

[2630] Gerson Sunyé, Damien Pollet, Yves Le Traon, and Jean-Marc Jézéquel. Refactoring UML models. Lecture Notes in Computer Science, 2185:134–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850134.
pdf.

Yacoub:2001:USD

[2631] Sherif M. Yacoub and Hany H. Ammar. UML support for designing software systems as a composition of design patterns. Lecture Notes in Computer Science, 2185:149–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850149.
pdf.

Crettaz:2001:ICA

[2632] Valentin Crettaz, Mohamed Mancona Kandé, Shane Sendall, and Alfred Strohmeier. Integrating the Concern-BASE approach with SADL. Lecture Notes in Computer Science, 2185:166– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850166.
pdf.

Devos:2001:MPO

[2633] Frank Devos and Eric Steegmans. The message paradigm in objectoriented analysis. Lecture Notes in Computer Science, 2185:182–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2185/21850182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2185/21850182.
                                        pdf.

Briand:2001:UBA

[2634] Lionel Briand and Yvan Labiche. A UML-based approach to system testing. Lecture Notes in Computer Science, 2185:194–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2185/21850194.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2185/21850194.pdf.

Grassi:2001:UMP

[2635] Vincenzo Grassi and Raffaela Mirandola. UML modelling and performance analysis of mobile software architectures. Lecture Notes in Computer Science, 2185:209–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2185/21850209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2185/21850209.
                                        pdf.

Marcos:2001:EUO

[2636] E. Marcos, B. Vela, and J. M.

Cavero. Extending UML for objectrelational database design. Lecture Notes in Computer Science, 2185:225– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850225.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850225.
pdf.

Damm:2001:UUP

[2637] Werner Damm. Understanding UML — pains and rewards. Lecture Notes in Computer Science, 2185:240–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850240.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850240.
pdf.

Kuske:2001:FSU

[2638] Sabine Kuske. A formal semantics of UML state machines based on structured graph transformation. Lecture Notes in Computer Science, 2185:241– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850241.
pdf.

Bottoni:2001:VOU

[2639] Paolo Bottoni, Manuel Koch, Francesco Parisi-Presicce, and Gabriele Taentzer. A visualization of OCL using collaborations. Lecture Notes in Computer Science, 2185:257–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2185/21850257.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2185/21850257.pdf.

Engels:2001:RBS

[2640] Gregor Engels, Reiko Heckel, and Jochen Malte Küster. Rule-based specification of behavioral consistency based on the UML meta-model. Lecture Notes in Computer Science, 2185: 272–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850272.
pdf.

Apvrille:2001:NUP

[2641] L. Apvrille, P. de Saqui-Sannes, C. Lohr, P. Sénac, and J.-P. Courtiat. A new UML profile for real-time system formal design and validation. Lecture Notes in Computer Science, 2185: 287–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850287.
pdf.

Latronico:2001:RES

[2642] Elizabeth Latronico and Philip Koopman. Representing embedded system sequence diagrams as a formal language. Lecture Notes in Computer Science, 2185:302–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2185/21850302.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2185/21850302.pdf.

Lettrari:2001:SBM

[2643] Marc Lettrari and Jochen Klose.

Scenario-based monitoring and testing of real-time UML models. Lecture Notes in Computer Science, 2185:317– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850317.
pdf.

Genova:2001:SMM

[2644] Gonzalo Génova, Juan Llorens, and Paloma Martínez. Semantics of the minimum multiplicity in ternary associations in UML. Lecture Notes in Computer Science, 2185:329–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850329.
pdf.

Baclawski:2001:EUS

[2645] Kenneth Baclawski, Mieczyslaw K. Kokar, Paul A. Kogut, Lewis Hart, Jeffrey Smith, William S. Holmes III, Jerzy Letkowski, and Michael L. Aronson. Extending UML to support ontology engineering for the semantic Web. Lecture Notes in Computer Science, 2185:342–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850342.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850342.
pdf.

Stevens:2001:AUM

[2646] Perdita Stevens. On associations in the unified modelling language. Lecture Notes in Computer Science, 2185: 361–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850361.
pdf.

Sekerinski:2001:IST

[2647] Emil Sekerinski and Rafik Zurob. iState: A statechart translator. Lecture Notes in Computer Science, 2185:376– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850376.
pdf.

Sendall:2001:SCS

[2648] Shane Sendall and Alfred Strohmeier.

Specifying concurrent system behavior and timing constraints using OCL and UML. Lecture Notes in Computer Science, 2185:391–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850391.
pdf.

vonderBeeck:2001:FUS

[2649] Michael von der Beeck. Formalization of UML-statecharts. Lecture Notes in Computer Science, 2185:406– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850406.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850406.
pdf.

Mylopoulos:2001:UAO

[2650] John Mylopoulos, Manuel Kolp, and Jaelson Castro. UML for agentoriented software development: The tropos proposal. Lecture Notes in Computer Science, 2185:422–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850422.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850422.
pdf.

Weis:2001:UMM

[2651] Torben Weis, Christian Becker, Kurt Geihs, and Noël Plouzeau. A UML meta-model for contract aware components. Lecture Notes in Computer Science, 2185:442–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850442.
pdf.

Roubtsova:2001:SMI

[2652] E. E. Roubtsova, L. C. M. van Gool, R. Kuiper, and H. B. M. Jonkers. A specification model for interface suites. Lecture Notes in Computer Science, 2185:457–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850457.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850457.
pdf.

Metz:2001:AUC

[2653] Pierre Metz, John O’Brien, and Wolfgang Weber. Against use case interleaving. Lecture Notes in Computer Science, 2185:472–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850472.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850472.
pdf.

Anda:2001:ESD

[2654] Bente Anda, Hege Dreiem, Dag I. K.

Sjøberg, and Magne Jørgensen. Estimating software development effort based on use cases-experiences from industry. Lecture Notes in Computer Science, 2185:487–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2185/21850487.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2185/21850487.
pdf.

Hussmann:2001:WTU

[2655] Heinrich Hussmann. Workshops and tutorials at the UML 2001 conference. Lecture Notes in Computer Science, 2185:503–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2185/21850503.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2185/21850503.
pdf.

Anonymous:2001:AIhq

[2656] Anonymous. Author index. Lecture Notes in Computer Science, 2185:509– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2185/2185auth.pdf.

Biggerstaff:2001:CGC

[2657] Ted J. Biggerstaff. A characterization of generator and component reuse technologies. Lecture Notes in Computer Science, 2186:1– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2186/21860001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2186/21860001.
pdf.

Lopez-Herrejon:2001:SPE

[2658] Roberto E. Lopez-Herrejon and Don Batory. A standard problem for evaluating product-line methodologies. Lecture Notes in Computer Science, 2186: 10–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2186/21860010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2186/21860010.
pdf.

Wijnstra:2001:CII

[2659] Jan Gerben Wijnstra. Components, interfaces and information models within a platform architecture. Lecture Notes in Computer Science, 2186:25– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2186/21860025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2186/21860025.
pdf.

Zhang:2001:XAS

[2660] Hongyu Zhang, Stan Jarzabek, and Soe Myat Swe. XVCL approach to separating concerns in product family assets. Lecture Notes in Computer Science, 2186:36–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2186/21860036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2186/21860036.
pdf.

Vranic:2001:APM

[2661] Valentino Vranić. AspectJ paradigm model: A basis for multi-paradigm design for AspectJ. Lecture Notes in Computer Science, 2186:48–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2186/21860048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2186/21860048.
pdf.

Heuzeroth:2001:AOC

[2662] Dirk Heuzeroth, Welf Löwe, Andreas Ludwig, and Uwe Aßmann. Aspectoriented configuration and adaptation of component communication. Lecture Notes in Computer Science, 2186:58– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2186/21860058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2186/21860058.
pdf.

Pulvermuller:2001:VMA

[2663] Elke Pulvermüller, Andreas Speck, and James O. Coplien. A version model for aspect dependency management. Lecture Notes in Computer Science, 2186: 70–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2186/21860070.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2186/21860070.
                                        pdf.

Hanenberg:2001:OMG

[2664] Stefan Hanenberg, Boris Bachmendo, and Rainer Unland. An object model for general-purpose aspect languages. Lecture Notes in Computer Science, 2186:80–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2186/21860080.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2186/21860080.
                                        pdf.

Nicolet:2001:GVF

[2665] Jean-Daniel Nicolet. Generic visitor framework computing statistical estimators. Lecture Notes in Computer Science, 2186:92–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2186/21860092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2186/21860092.
                                        pdf.

Gregor:2001:BCI

[2666] Douglas Gregor, Sibylle Schupp, and David Musser. Base class injection. Lecture Notes in Computer Science, 2186:106–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2186/21860106.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2186/21860106.
pdf.

Attardi:2001:RSM

[2667] Giuseppe Attardi and Antonio Cisternino. Reflection support by means of template metaprogramming. Lecture Notes in Computer Science, 2186:118– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2186/21860118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2186/21860118.
pdf.

deBruijn:2001:SBG

[2668] Hans de Bruijn and Hans van Vliet.

Scenario-based generation and evaluation of software architectures. Lecture Notes in Computer Science, 2186:128– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2186/21860128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2186/21860128.
pdf.

Kim:2001:RDC

[2669] Jaehyoun Kim and C. Robert Carlson. The role of design components in test plan generation. Lecture Notes in Computer Science, 2186:140– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2186/21860140.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2186/21860140.
                                        pdf.

Washizaki:2001:RSC

[2670] Hironori Washizaki and Yoshiaki Fukazawa. Retrieving software components using directed replaceability distance. Lecture Notes in Computer Science, 2186:153–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2186/21860153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2186/21860153.
                                        pdf.

Hakala:2001:GAD

[2671] Markku Hakala, Juha Hautamäki, Kai Koskimies, Jukka Paakki, Antti Viljamaa, and Jukka Viljamaa. Generating application development environments for Java frameworks. Lecture Notes in Computer Science, 2186:163– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2186/21860163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2186/21860163.
                                        pdf.

Anonymous:2001:AIhr

[2672] Anonymous. Author index. Lecture Notes in Computer Science, 2186:177– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2186/2186auth.pdf.

Fox:2001:DSM

[2673] John Fox. Designing safety into medical decisions and clinical processes. Lecture Notes in Computer Science, 2187:1–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2187/21870001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2187/21870001.
pdf.

Winther:2001:SAS

[2674] Rune Winther, Ole-Arnt Johnsen, and Bjørn Axel Gran. Security assessments of safety critical systems using HAZOPs. Lecture Notes in Computer Science, 2187:14–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870014.
pdf.

Naedele:2001:NSS

[2675] Martin Naedele, Dacfey Dzung, and Michael Stanimirov. Network security for substation automation systems. Lecture Notes in Computer Science, 2187:25–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2187/21870025.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2187/21870025.
                                        pdf.

Gran:2001:BBN

[2676] Bjørn Axel Gran and Atte Helminen. A Bayesian belief network for reliability assessment. Lecture Notes in Computer Science, 2187:35– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2187/21870035.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2187/21870035.
                                        pdf.

Pap:2001:CGS

[2677] Zsigmond Pap, István Majzik, and András Pataricza. Checking general safety criteria on UML statecharts. Lecture Notes in Computer Science, 2187:46–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2187/21870046.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2187/21870046.
                                        pdf.

Nordland:2001:PSC

[2678] Odd Nordland. Presenting a safety case — A case study. Lecture Notes in Computer Science, 2187:56– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2187/21870056.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2187/21870056.
pdf.

Fankhauser:2001:SFV

[2679] Hans R. Fankhauser. Safety functions versus control functions. Lecture Notes in Computer Science, 2187:66– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870066.
pdf.

Laible:2001:FSD

[2680] Ulrich Laible, Thomas Bürger, and Günter Pritschow. A fail-safe dual channel robot control for surgery applications. Lecture Notes in Computer Science, 2187:75–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2187/21870075.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2187/21870075.pdf.

Rushby:2001:MHH

[2681] John Rushby. Modeling the human in human factors. Lecture Notes in Computer Science, 2187:86– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870086.
pdf.

Mackel:2001:AHM

[2682] Oliver Mäckel and Georg Hoever. Analyzing human-machine interactions in safety-critical systems: A basic applicable approach. Lecture Notes in Computer Science, 2187:92– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870092.
pdf.

Pasquini:2001:AII

[2683] Alberto Pasquini, Antonio Rizzo, and Luca Save. Analysis of incidents involving interactive systems. Lecture Notes in Computer Science, 2187:100– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870100.
pdf.

Gawkowski:2001:EEF

[2684] Piotr Gawkowski and Janusz Sosnowski. Experimental evaluation of fault handling mechanisms. Lecture Notes in Computer Science, 2187:109– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870109.
pdf.

Redmill:2001:CDP

[2685] Felix Redmill. The COTS debate in perspective. Lecture Notes in Computer Science, 2187:119–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870119.
pdf.

Kaaniche:2001:IMS

[2686] Mohamed Kaâniche, Yannick Le Guédart, Jean Arlat, and Thierry Boyer. An investigation on mutation strategies for fault injection into RDD-100 models. Lecture Notes in Computer Science, 2187:130–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870130.
pdf.

Vakalis:2001:CSB

[2687] Ioannis Vakalis. A comparison study of the behavior of equivalent algorithms in fault injection experiments in parallel superscalar architectures. Lecture Notes in Computer Science, 2187:145– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870145.
pdf.

Kuball:2001:EST

[2688] Silke Kuball, Gordon Hughes, John H. R. May, Julio Gallardo, Andrew D. John, and Roy B. Carter. The effectiveness of statistical testing when applied to logic systems. Lecture Notes in Computer Science, 2187:156– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2187/21870156.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2187/21870156.
                                        pdf.

Zoffmann:2001:CSS

[2689] Günther Zoffmann, Manfred Gingerl, Christian Reumann, and Gerald Sonneck. A classification scheme for software verification tools with regard to RTCA/DO-178B. Lecture Notes in Computer Science, 2187:166– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2187/21870166.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2187/21870166.
                                        pdf.

Bitsch:2001:SPK

[2690] Friedemann Bitsch. Safety patterns — the key to formal specification of safety requirements. Lecture Notes in Computer Science, 2187:176–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2187/21870176.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2187/21870176.
pdf.

Cichocki:2001:FSF

[2691] Tadeusz Cichocki and Janusz Górski.

Formal support for fault modelling and analysis. Lecture Notes in Computer Science, 2187:190–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870190.
pdf.

Faller:2001:PEI

[2692] Rainer Faller. Project experience with IEC 61508 and its consequences. Lecture Notes in Computer Science, 2187: 200–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2187/21870200.
pdf.

Caspi:2001:ADD

[2693] Paul Caspi, Christine Mazuet, and Natacha Reynaud Paligot. About the design of distributed control systems: The quasi-synchronous approach. Lecture Notes in Computer Science, 2187: 215–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2187/21870215.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2187/21870215.
                                        pdf.

Betous-Almeida:2001:DE

[2694] Cláudia Betous-Almeida and Karama Kanoun. Dependability evaluation. Lecture Notes in Computer Science, 2187:227–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2187/21870227.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2187/21870227.
                                        pdf.

Porcarelli:2001:TDA

[2695] Stefano Porcarelli, Felicita Di Giandomenico, Amine Chohra, and Andrea Bondavalli. Tuning of database audits to improve scheduled maintenance in communication systems. Lecture Notes in Computer Science, 2187:238– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2187/21870238.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2187/21870238.
                                        pdf.

Anonymous:2001:AIhs

[2696] Anonymous. Author index. Lecture Notes in Computer Science, 2187:249– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2187/2187auth.pdf.